GNU Scientific Library

Reference Manual

Edition 1.0, for GSL Version 1.0 1 November 2001

Mark Galassi

Los Alamos National Laboratory

Jim Davies

Department of Computer Science, Georgia Institute of Technology

James Theiler

Astrophysics and Radiation Measurements Group, Los Alamos National Laboratory

Brian Gough

Network Theory Limited

Gerard Jungman

Theoretical Fluid Dynamics Group, Los Alamos National Laboratory

Michael Booth

Department of Physics and Astronomy, The Johns Hopkins University

Fabrice Rossi

University of Paris-Dauphine Copyright \begin{matrix}c \\ ◯ \end{matrix} 1996, 1997, 1998, 1999, 2000, 2001 The GSL Team. Permission is granted to copy, distribute and/or modify this document under the terms of the GNU Free Documentation License, Version 1.1 or any later version published by the Free Software Foundation; with no Invariant Sections, no Front-Cover Texts, and no Back-Cover Texts. A copy of the license is included in the section entitled “GNU Free Documentation License”.

1 Introduction

The GNU Scientific Library (GSL) is a collection of routines for numerical computing. The routines have been written from scratch in C, and are meant to present a modern Applications Programming Interface (API) for C programmers, while allowing wrappers to be written for very high level languages. The source code is distributed under the GNU General Public License.

1.1 Routines available in GSL

The library covers a wide range of topics in numerical computing. Routines are available for the following areas,

Complex Numbers  Roots of Polynomials
Special Functions  Vectors and Matrices
Permutations  Sorting
BLAS Support  Linear Algebra
Eigensystems  Fast Fourier Transforms
Quadrature  Random Numbers
Quasi-Random Sequences  Random Distributions
Statistics  Histograms
N-Tuples  Monte Carlo Integration
Simulated Annealing  Differential Equations
Interpolation  Numerical Differentiation
Chebyshev Approximations  Series Acceleration
Discrete Hankel Transforms  Root-Finding
Minimization  Least-Squares Fitting
Physical Constants  IEEE Floating-Point

The use of these routines is described in this manual. Each chapter provides detailed definitions of the functions, followed by example programs and references to the articles on which the algorithms are based.

1.2 GSL is Free Software

The subroutines in the GNU Scientific Library are “free software”; this means that everyone is free to use them, and to redistribute them in other free programs. The library is not in the public domain; it is copyrighted and there are conditions on its distribution. These conditions are designed to permit everything that a good cooperating citizen would want to do. What is not allowed is to try to prevent others from further sharing any version of the software that they might get from you.

Specifically, we want to make sure that you have the right to give away copies of any programs related to the GNU Scientific Library, that you receive their source code or else can get it if you want it, that you can change these programs or use pieces of them in new free programs, and that you know you can do these things. The library should not be redistributed in proprietary programs.

To make sure that everyone has such rights, we have to forbid you to deprive anyone else of these rights. For example, if you distribute copies of any related code, you must give the recipients all the rights that you have. You must make sure that they, too, receive or can get the source code. And you must tell them their rights.

Also, for our own protection, we must make certain that everyone finds out that there is no warranty for the GNU Scientific Library. If these programs are modified by someone else and passed on, we want their recipients to know that what they have is not what we distributed, so that any problems introduced by others will not reflect on our reputation.

The precise conditions for the distribution of software related to the GNU Scientific Library are found in the GNU General Public License (see [GNU General Public License], page 392). Further information about this license is available from the GNU Project webpage Frequently Asked Questions about the GNU GPL,

    http://www.gnu.org/copyleft/gpl-faq.html

1.3 Obtaining GSL

The source code for the library can be obtained in different ways, by copying it from a friend, purchasing it on cdrom or downloading it from the internet. A list of public ftp servers which carry the source code can be found on the development website,

    http://sources.redhat.com/gsl/

The preferred platform for the library is a GNU system, which allows it to take advantage of additional features. The library is portable and compiles on most Unix platforms. It is also available for Microsoft Windows. Precompiled versions of the library can be purchased from commercial redistributors listed on the website.

Announcements of new releases, updates and other relevant events are made on the gslannounce mailing list. To subscribe to this low-volume list, send an email of the following form,

     To: gsl-announce-request@sources.redhat.com
     Subject: subscribe

You will receive a response asking to you to reply in order to confirm your subscription.

1.4 An Example Program

The following short program demonstrates the use of the library by computing the value of the Bessel function J_{0}(x) for x = 5,

     #include <stdio.h>
     #include <gsl/gsl_sf_bessel.h>
     int
     main (void)
     {
       double x = 5.0;
       double y = gsl_sf_bessel_J0 (x);
       printf("J0(%g) = %.18e\n", x, y);
       return 0;
     }

The output is shown below, and should be correct to double-precision accuracy,

     J0(5) = -1.775967713143382920e-01

The steps needed to compile programs which use the library are described in the next chapter.

1.5 No Warranty

The software described in this manual has no warranty, it is provided "as is". It is your responsibility to validate the behavior of the routines and their accuracy using the source code provided. Consult the GNU General Public license for further details (see [GNU General Public License], page 392).

1.6 Further Information

Additional information, including online copies of this manual, links to related projects, and mailing list archives are available from the development website mentioned above. The developers of the library can be reached via the project’s public mailing list,

    gsl-discuss@sources.redhat.com

This mailing list can be used to report bugs or to ask questions not covered by this manual.

2 Using the library

This chapter describes how to compile programs that use GSL, and introduces its conventions.

2.1 ANSI C Compliance

The library is written in ANSI C and is intended to conform to the ANSI C standard. It should be portable to any system with a working ANSI C compiler.

The library does not rely on any non-ANSI extensions in the interface it exports to the user. Programs you write using GSL can be ANSI compliant. Extensions which can be used in a way compatible with pure ANSI C are supported, however, via conditional compilation. This allows the library to take advantage of compiler extensions on those platforms which support them.

When an ANSI C feature is known to be broken on a particular system the library will exclude any related functions at compile-time. This should make it impossible to link a program that would use these functions and give incorrect results.

To avoid namespace conflicts all exported function names and variables have the prefix gsl_, while exported macros have the prefix GSL_.

2.2 Compiling and Linking

The library header files are installed in their own ‘gsl’ directory. You should write any preprocessor include statements with a ‘gsl/’ directory prefix thus,

     #include <gsl/gsl_math.h>

If the directory is not installed on the standard search path of your compiler you will also need to provide its location to the preprocessor as a command line flag. The default location

of the ‘gsl’ directory is ‘/usr/local/include/gsl’.

The library is installed as a single file, ‘libgsl.a’. A shared version of the library is also installed on systems that support shared libraries. The default location of these files is ‘/usr/local/lib’. To link against the library you need to specify both the main library and a supporting cblas library, which provides standard basic linear algebra subroutines. A suitable cblas implementation is provided in the library ‘libgslcblas.a’ if your system does not provide one. The following example shows how to link an application with the library,

     gcc app.o -lgsl -lgslcblas -lm

The following command line shows how you would link the same application with an alternative blas library called ‘libcblas’,

     gcc app.o -lgsl -lcblas -lm

For the best performance an optimized platform-specific cblas library should be used for -lcblas. The library must conform to the cblas standard. The atlas package provides a portable high-performance blas library with a cblas interface. It is free software and should be installed for any work requiring fast vector and matrix operations. The following command line will link with the atlas library and its cblas interface,

     gcc app.o -lgsl -lcblas -latlas -lm

For more information see Chapter 11 [BLAS Support], page 105.

The program gsl-config provides information on the local version of the library. For example, the following command shows that the library has been installed under the direc-

tory ‘/usr/local’,
     bash$ gsl-config --prefix
     /usr/local

Further information is available using the command gsl-config --help.

2.3 Shared Libraries

To run a program linked with the shared version of the library it may be necessary to define the shell variable LD_LIBRARY_PATH to include the directory where the library is installed. For example,

     LD_LIBRARY_PATH=/usr/local/lib:$LD_LIBRARY_PATH ./app

To compile a statically linked version of the program instead, use the -static flag in gcc,

     gcc -static app.o -lgsl -lgslcblas -lm

2.4 Autoconf macros

For applications using autoconf the standard macro AC_CHECK_LIB can be used to link with the library automatically from a configure script. The library itself depends on the presence of a cblas and math library as well, so these must also be located before linking with the main libgsl file. The following commands should be placed in the ‘configure.in’ file to perform these tests,

     AC_CHECK_LIB(m,main)
     AC_CHECK_LIB(gslcblas,main)
     AC_CHECK_LIB(gsl,main)

Assuming the libraries are found the output during the configure stage looks like this,

     checking for main in -lm... yes
     checking for main in -lgslcblas... yes
     checking for main in -lgsl... yes

If the library is found then the tests will define the macros HAVE_LIBGSL, HAVE_

LIBGSLCBLAS, HAVE_LIBM and add the options -lgsl -lgslcblas -lm to the variable
LIBS.

The tests above will find any version of the library. They are suitable for general use, where the versions of the functions are not important. An alternative macro is available in the file ‘gsl.m4’ to test for a specific version of the library. To use this macro simply add the following line to your ‘configure.in’ file instead of the tests above:

     AM_PATH_GSL(GSL_VERSION,
                [action-if-found],
                [action-if-not-found])

The argument GSL_VERSION should be the two or three digit major.minor or major.minor.micro version number of the release you require. A suitable choice for

action-if-not-found is,

     AC_MSG_ERROR(could not find required version of GSL)

Then you can add the variables GSL_LIBS and GSL_CFLAGS to your Makefile.am files to obtain the correct compiler flags. GSL_LIBS is equal to the output of the gsl-config -- libs command and GSL_CFLAGS is equal to gsl-config --cflags command. For example,

     libgsdv_la_LDFLAGS =    \
             $(GTK_LIBDIR) \
             $(GTK_LIBS) -lgsdvgsl $(GSL_LIBS) -lgslcblas

Note that the macro AM_PATH_GSL needs to use the C compiler so it should appear in the ‘configure.in’ file before the macro AC_LANG_CPLUSPLUS for programs that use C++.

2.5 Inline functions

The inline keyword is not part of ANSI C and the library does not export any inline function definitions by default. However, the library provides optional inline versions of performance-critical functions by conditional compilation. The inline versions of these functions can be included by defining the macro HAVE_INLINE when compiling an application.

     gcc -c -DHAVE_INLINE app.c

If you use autoconf this macro can be defined automatically. The following test should be placed in your ‘configure.in’ file,

     AC_C_INLINE
     if test "$ac_cv_c_inline" != no ; then
       AC_DEFINE(HAVE_INLINE,1)
       AC_SUBST(HAVE_INLINE)
     fi

and the macro will then be defined in the compilation flags or by including the file ‘config.h’ before any library headers. If you do not define the macro HAVE_INLINE then the slower non-inlined versions of the functions will be used instead.

Note that the actual usage of the inline keyword is extern inline, which eliminates unnecessary function definitions in gcc. If the form extern inline causes problems with other compilers a stricter autoconf test can be used, see Appendix C [Autoconf Macros], page 377.

2.6 Long double

The extended numerical type long double is part of the ANSI C standard and should be available in every modern compiler. However, the precision of long double is platform dependent, and this should be considered when using it. The IEEE standard only specifies the minimum precision of extended precision numbers, while the precision of double is the same on all platforms.

In some system libraries the stdio.h formatted input/output functions printf and scanf are not implemented correctly for long double. Undefined or incorrect results are avoided by testing these functions during the configure stage of library compilation and eliminating certain GSL functions which depend on them if necessary. The corresponding line in the configure output looks like this,

     checking whether printf works with long double... no

Consequently when long double formatted input/output does not work on a given system it should be impossible to link a program which uses GSL functions dependent on this.

If it is necessary to work on a system which does not support formatted long double input/output then the options are to use binary formats or to convert long double results into double for reading and writing.

2.7 Portability functions

To help in writing portable applications GSL provides some implementations of functions that are found in other libraries, such as the BSD math library. You can write your application to use the native versions of these functions, and substitute the GSL versions via a preprocessor macro if they are unavailable on another platform. The substitution can be made automatically if you use autoconf. For example, to test whether the BSD function hypot is available you can include the following line in the configure file ‘configure.in’ for your application,

     AC_CHECK_FUNCS(hypot)

and place the following macro definitions in the file ‘config.h.in’,

     /* Substitute gsl_hypot for missing system hypot */
     #ifndef HAVE_HYPOT
     #define hypot gsl_hypot
     #endif

The application source files can then use the include command #include <config.h> to substitute gsl_hypot for each occurrence of hypot when hypot is not available.

In most circumstances the best strategy is to use the native versions of these functions when available, and fall back to GSL versions otherwise, since this allows your application to take advantage of any platform-specific optimizations in the system library. This is the strategy used within GSL itself.

2.8 Alternative optimized functions

The main implementation of some functions in the library will not be optimal on all architectures. For example, there are several ways to compute a Gaussian random variate and their relative speeds are platform-dependent. In cases like this the library provides alternate implementations of these functions with the same interface. If you write your application using calls to the standard implementation you can select an alternative version later via a preprocessor definition. It is also possible to introduce your own optimized functions this way while retaining portability. The following lines demonstrate the use of a platform-dependent choice of methods for sampling from the Gaussian distribution,

     #ifdef SPARC
     #define gsl_ran_gaussian gsl_ran_gaussian_ratio_method
     #endif
     #ifdef INTEL
     #define gsl_ran_gaussian my_gaussian
     #endif

These lines would be placed in the configuration header file ‘config.h’ of the application, which should then be included by all the source files. Note that the alternative implementations will not produce bit-for-bit identical results, and in the case of random number distributions will produce an entirely different stream of random variates.

2.9 Support for different numeric types

Many functions in the library are defined for different numeric types. This feature is implemented by varying the name of the function with a type-related modifier — a primitive form of C++ templates. The modifier is inserted into the function name after the initial module prefix. The following table shows the function names defined for all the numeric types of an imaginary module gsl_foo with function fn,

gsl_foo_fn               double
gsl_foo_long_double_fn   long double
gsl_foo_float_fn         float
gsl_foo_long_fn          long
gsl_foo_ulong_fn         unsigned long
gsl_foo_int_fn           int
gsl_foo_uint_fn          unsigned int
gsl_foo_short_fn         short
gsl_foo_ushort_fn        unsigned short
gsl_foo_char_fn          char
gsl_foo_uchar_fn         unsigned char

The normal numeric precision double is considered the default and does not require a suffix. For example, the function gsl_stats_mean computes the mean of double precision numbers, while the function gsl_stats_int_mean computes the mean of integers.

A corresponding scheme is used for library defined types, such as gsl_vector and gsl_ matrix. In this case the modifier is appended to the type name. For example, if a module defines a new type-dependent struct or typedef gsl_foo it is modified for other types in the following way,

gsl_foo                  double
gsl_foo_long_double      long double
gsl_foo_float            float
gsl_foo_long             long
gsl_foo_ulong            unsigned long
gsl_foo_int              int
gsl_foo_uint             unsigned int
gsl_foo_short            short
gsl_foo_ushort           unsigned short
gsl_foo_char             char
gsl_foo_uchar            unsigned char

When a module contains type-dependent definitions the library provides individual header files for each type. The filenames are modified as shown in the below. For convenience the default header includes the definitions for all the types. To include only the double precision header, or any other specific type, file use its individual filename.

     #include <gsl/gsl_foo.h>               All types
     #include <gsl/gsl_foo_double.h>        double
     #include <gsl/gsl_foo_long_double.h>   long double
     #include <gsl/gsl_foo_float.h>         float
     #include <gsl/gsl_foo_long.h>          long
     #include <gsl/gsl_foo_ulong.h>         unsigned long
     #include <gsl/gsl_foo_int.h>           int
     #include <gsl/gsl_foo_uint.h>          unsigned int
     #include <gsl/gsl_foo_short.h>         short
     #include <gsl/gsl_foo_ushort.h>        unsigned short
     #include <gsl/gsl_foo_char.h>          char
     #include <gsl/gsl_foo_uchar.h>         unsigned char

2.10 Compatibility with C++

The library header files automatically define functions to have extern "C" linkage when included in C++ programs.

2.11 Aliasing of arrays

The library assumes that arrays, vectors and matrices passed as modifiable arguments are not aliased and do not overlap with each other. This removes the need for the library to handle overlapping memory regions as a special case, and allows additional optimizations to be used. If overlapping memory regions are passed as modifiable arguments then the results of such functions will be undefined. If the arguments will not be modified (for example, if a function prototype declares them as const arguments) then overlapping or aliased memory regions can be safely used.

2.12 Code Reuse

Where possible the routines in the library have been written to avoid dependencies between modules and files. This should make it possible to extract individual functions for use in your own applications, without needing to have the whole library installed. You may need to define certain macros such as GSL_ERROR and remove some #include statements in order to compile the files as standalone units. Reuse of the library code in this way is encouraged, subject to the terms of the GNU General Public License.

3 Error Handling

This chapter describes the way that GSL functions report and handle errors. By examining the status information returned by every function you can determine whether it succeeded or failed, and if it failed you can find out what the precise cause of failure was. You can also define your own error handling functions to modify the default behavior of the library.

The functions described in this section are declared in the header file ‘gsl_errno.h’.

3.1 Error Reporting

The library follows the thread-safe error reporting conventions of the posix Threads library. Functions return a non-zero error code to indicate an error and 0 to indicate success.

     int status = gsl_function(...)
     if (status) { /* an error occurred */
       .....
       /* status value specifies the type of error */
     }

The routines report an error whenever they cannot perform the task requested of them. For example, a root-finding function would return a non-zero error code if could not converge to the requested accuracy, or exceeded a limit on the number of iterations. Situations like this are a normal occurrence when using any mathematical library and you should check the return status of the functions that you call.

Whenever a routine reports an error the return value specifies the type of error. The return value is analogous to the value of the variable errno in the C library. However, the C library’s errno is a global variable, which is not thread-safe (There can be only one instance of a global variable per program. Different threads of execution may overwrite errno simultaneously). Returning the error number directly avoids this problem. The caller can examine the return code and decide what action to take, including ignoring the error if it is not considered serious.

The error code numbers are defined in the file ‘gsl_errno.h’. They all have the prefix GSL_and expand to non-zero constant integer values. Many of the error codes use the same base name as a corresponding error code in C library. Here are some of the most common error codes,

Macro int GSL EDOM

Domain error; used by mathematical functions when an argument value does not fall into the domain over which the function is defined (like EDOM in the C library)

Macro int GSL ERANGE

Range error; used by mathematical functions when the result value is not representable because of overflow or underflow (like ERANGE in the C library)

Macro int GSL ENOMEM

No memory available. The system cannot allocate more virtual memory because its capacity is full (like ENOMEM in the C library). This error is reported when a GSL routine encounters problems when trying to allocate memory with malloc.

Macro int GSL EINVAL

Invalid argument. This is used to indicate various kinds of problems with passing the wrong argument to a library function (like EINVAL in the C library).

Here is an example of some code which checks the return value of a function where an error might be reported,

     int status = gsl_fft_complex_radix2_forward (data, n);
     if (status) {
       if (status == GSL_EINVAL) {
          fprintf (stderr, "invalid argument, n=%d\n", n);
       } else {
          fprintf (stderr, "failed, gsl_errno=%d\n",
                           status);
       }
       exit (-1);
     }

The function gsl_fft_complex_radix2 only accepts integer lengths which are a power of two. If the variable n is not a power of two then the call to the library function will return GSL_EINVAL, indicating that the length argument is invalid. The else clause catches any other possible errors.

The error codes can be converted into an error message using the function gsl_strerror.

Function const char * gsl strerror (const int gsl errno)

This function returns a pointer to a string describing the error code gsl errno. For example,

          printf("error: %s\n", gsl_strerror (status));

would print an error message like error: output range error for a status value of

     GSL_ERANGE.

3.2 Error Handlers

In addition to reporting errors the library also provides an optional error handler. The error handler is called by library functions when they report an error, just before they return to the caller. The purpose of the handler is to provide a function where a breakpoint can be set that will catch library errors when running under the debugger. It is not intended for use in production programs, which should handle any errors using the error return codes described above.

The default behavior of the error handler is to print a short message and call abort() whenever an error is reported by the library. If this default is not turned off then any program using the library will stop with a core-dump whenever a library routine reports an error. This is intended as a fail-safe default for programs which do not check the return status of library routines (we don’t encourage you to write programs this way). If you turn off the default error handler it is your responsibility to check the return values of the GSL routines. You can customize the error behavior by providing a new error handler. For example, an alternative error handler could log all errors to a file, ignore certain error conditions (such as underflows), or start the debugger and attach it to the current process when an error occurs.

All GSL error handlers have the type gsl_error_handler_t, which is defined in

‘gsl_errno.h’,

Data Type gsl error handler t

This is the type of GSL error handler functions. An error handler will be passed four arguments which specify the reason for the error (a string), the name of the source file in which it occurred (also a string), the line number in that file (an integer) and the error number (an integer). The source file and line number are set at compile time using the __FILE__and __LINE__directives in the preprocessor. An error handler function returns type void. Error handler functions should be defined like this,

     void handler (const char * reason,
                   const char * file,
                   int line,
                   int gsl_errno)

To request the use of your own error handler you need to call the function gsl_set_error_ handler which is also declared in ‘gsl_errno.h’,

Function gsl_error_handler_t gsl set error handler

    (gsl_error_handler_t new handler)

This functions sets a new error handler, new handler, for the GSL library routines. The previous handler is returned (so that you can restore it later). Note that the pointer to a user defined error handler function is stored in a static variable, so there can only be one error handler per program. This function should be not be used in multi-threaded programs except to set up a program-wide error handler from a master thread. The following example shows how to set and restore a new error handler,

     /* save original handler, install new handler */
     old_handler = gsl_set_error_handler (&my_handler);
     /* code uses new handler */
     .....
     /* restore original handler */
     gsl_set_error_handler (old_handler);

To use the default behavior (abort on error) set the error handler to NULL,

     old_handler = gsl_set_error_handler (NULL);

Function gsl_error_handler_t gsl set error handler off ()

This function turns off the error handler by defining an error handler which does nothing. This will cause the program to continue after any error, so the return values from any library routines must be checked. This is the recommended behavior for production programs. The previous handler is returned (so that you can restore it later).

The error behavior can be changed for specific applications by recompiling the library with a customized definition of the GSL_ERROR macro in the file ‘gsl_errno.h’.

3.3 Using GSL error reporting in your own functions

If you are writing numerical functions in a program which also uses GSL code you may find it convenient to adopt the same error reporting conventions as in the library.

To report an error you need to call the function gsl_error with a string describing the error and then return an appropriate error code from gsl_errno.h, or a special value, such as NaN. For convenience the file ‘gsl_errno.h’ defines two macros which carry out these steps:

Macro GSL ERROR (reason, gsl errno)

This macro reports an error using the GSL conventions and returns a status value of gsl_errno. It expands to the following code fragment,

          gsl_error (reason, __FILE__, __LINE__, gsl_errno);
          return gsl_errno;

The macro definition in ‘gsl_errno.h’ actually wraps the code in a do { ... } while (0) block to prevent possible parsing problems.

Here is an example of how the macro could be used to report that a routine did not achieve a requested tolerance. To report the error the routine needs to return the error

code GSL_ETOL.
     if (residual > tolerance)
       {
         GSL_ERROR("residual exceeds tolerance", GSL_ETOL);
       }

Macro GSL ERROR VAL (reason, gsl errno, value)

This macro is the same as GSL_ERROR but returns a user-defined status value of value instead of an error code. It can be used for mathematical functions that return a floating point value.

Here is an example where a function needs to return a NaN because of a mathematical singularity,

     if (x == 0)
       {
         GSL_ERROR_VAL("argument lies on singularity",
                       GSL_ERANGE, GSL_NAN);
       }

Chapter 4: Mathematical Functions 14

4 Mathematical Functions

This chapter describes basic mathematical functions. Some of these functions are present in system libraries, but the alternative versions given here can be used as a substitute when the system functions are not available.

The functions and macros described in this chapter are defined in the header file

‘gsl_math.h’.

4.1 Mathematical Constants

The library ensures that the standard bsd mathematical constants are defined. For reference here is a list of the constants. M_E The base of exponentials, e M_LOG2E The base-2 logarithm of e, log_{2}(e) M_LOG10E The base-10 logarithm of e, log_{10}(e) M_SQRT2 The square root of two, \sqrt{2}

M_SQRT1_2

The square root of one-half, \sqrt{1/2}

M_SQRT3 The square root of three, \sqrt{3} M_PI The constant pi, π M_PI_2 Pi divided by two, π/2 M_PI_4 Pi divided by four, π/4 M_SQRTPI The square root of pi, \sqrt{π}

M_2_SQRTPI

Two divided by the square root of pi, 2/\sqrt{π}

M_1_PI The reciprocal of pi, 1/π M_2_PI Twice the reciprocal of pi, 2/π M_LN10 The natural logarithm of ten, ln(10) M_LN2 The natural logarithm of two, ln(2) M_LNPI The natural logarithm of pi, ln(π) M_EULER Euler’s constant, γ

4.2 Infinities and Not-a-number

Macro GSL POSINF

This macro contains the IEEE representation of positive infinity, +∞. It is computed from the expression +1.0/0.0.

Macro GSL NEGINF

This macro contains the IEEE representation of negative infinity, −∞. It is computed from the expression -1.0/0.0.

Macro GSL NAN

This macro contains the IEEE representation of the Not-a-Number symbol, NaN. It is computed from the ratio 0.0/0.0.

Function int gsl isnan (const double x)

This function returns 1 if x is not-a-number.

Function int gsl isinf (const double x)

This function returns +1 if x is positive infinity, −1 if x is negative infinity and 0 otherwise.

Function int gsl finite (const double x)

This function returns 1 if x is a real number, and 0 if it is infinite or not-a-number.

4.3 Elementary Functions

The following routines provide portable implementations of functions found in the BSD math library. When native versions are not available the functions described here can be used instead. The substitution can be made automatically if you use autoconf to compile your application (see Section 2.7 [Portability functions], page 7).

Function double gsl log1p (const double x)

This function computes the value of log(1 + x) in a way that is accurate for small x. It provides an alternative to the BSD math function log1p(x).

Function double gsl expm1 (const double x)

This function computes the value of exp(x) − 1 in a way that is accurate for small x. It provides an alternative to the BSD math function expm1(x).

Function double gsl hypot (const double x, const double y)

This function computes the value of \sqrt{x^{2} + y^{2}} in a way that avoids overflow. It provides an alternative to the BSD math function hypot(x,y).

Function double gsl acosh (const double x)

This function computes the value of arccosh(x). It provides an alternative to the standard math function acosh(x).

Function double gsl asinh (const double x)

This function computes the value of arcsinh(x). It provides an alternative to the standard math function asinh(x).

Function double gsl atanh (const double x)

This function computes the value of arctanh(x). It provides an alternative to the standard math function atanh(x).

4.4 Small integer powers

A common complaint about the standard C library is its lack of a function for calculating (small) integer powers. GSL provides a simple functions to fill this gap. For reasons of efficiency, these functions do not check for overflow or underflow conditions.

Function double gsl pow int (double x, int n)

This routine computes the power x^{n} for integer n. The power is computed using the minimum number of multiplications. For example, x^{8} is computed as ((x^{2})^{2})^{2}, requiring only 3 multiplications. A version of this function which also computes the numerical error in the result is available as gsl_sf_pow_int_e.

Function double gsl pow 2 (const double x) Function double gsl pow 3 (const double x) Function double gsl pow 4 (const double x) Function double gsl pow 5 (const double x) Function double gsl pow 6 (const double x) Function double gsl pow 7 (const double x) Function double gsl pow 8 (const double x) Function double gsl pow 9 (const double x)

These functions can be used to compute small integer powers x^{2}, x^{3}, etc. efficiently. The functions will be inlined when possible so that use of these functions should be as efficient as explicitly writing the corresponding product expression.

#include <gsl/gsl_math.h>
double y = gsl_pow_4 (3.141)  /* compute 3.141**4 */

4.5 Testing the Sign of Numbers

Macro GSL SIGN (x)

This macro returns the sign of x. It is defined as ((x) >= 0 ? 1 : -1). Note that with this definition the sign of zero is positive (regardless of its ieee sign bit).

4.6 Testing for Odd and Even Numbers

Macro GSL IS ODD (n)

This macro evaluates to 1 if n is odd and 0 if n is even. The argument n must be of integer type.

Macro GSL IS EVEN (n)

This macro is the opposite of GSL_IS_ODD(n). It evaluates to 1 if n is even and 0 if n is odd. The argument n must be of integer type.

4.7 Maximum and Minimum functions

Macro GSL MAX (a, b)

This macro returns the maximum of a and b. It is defined as ((a) > (b) ? (a):(b)).

Macro GSL MIN (a, b)

This macro returns the minimum of a and b. It is defined as ((a) < (b) ? (a):(b)).

Function extern inline double GSL MAX DBL (double a, double b)

This function returns the maximum of the double precision numbers a and b using an inline function. The use of a function allows for type checking of the arguments as an extra safety feature. On platforms where inline functions are not available the macro GSL_MAX will be automatically substituted.

Function extern inline double GSL MIN DBL (double a, double b)

This function returns the minimum of the double precision numbers a and b using an inline function. The use of a function allows for type checking of the arguments as an extra safety feature. On platforms where inline functions are not available the macro GSL_MIN will be automatically substituted.

Function extern inline int GSL MAX INT (int a, int b)
Function extern inline int GSL MIN INT (int a, int b)

These functions return the maximum or minimum of the integers a and b using an inline function. On platforms where inline functions are not available the macros GSL_MAX or GSL_MIN will be automatically substituted.

Function extern inline long double GSL MAX LDBL (long double a,
    long double b)
Function extern inline long double GSL MIN LDBL (long double a,
    long double b)

These functions return the maximum or minimum of the long doubles a and b using an inline function. On platforms where inline functions are not available the macros GSL_MAX or GSL_MIN will be automatically substituted.

5 Complex Numbers

The functions described in this chapter provide support for complex numbers. The algorithms take care to avoid unnecessary intermediate underflows and overflows, allowing the functions to evaluated over the as much of the complex plane as possible.

For multiple-valued functions the branch cuts have been chosen to follow the conventions of Abramowitz and Stegun in the Handbook of Mathematical Functions. The functions return principal values which are the same as those in GNU Calc, which in turn are the same as those in Common Lisp, The Language (Second Edition) (n.b. The second edition uses different definitions from the first edition) and the HP-28/48 series of calculators.

The complex types are defined in the header file ‘gsl_complex.h’, while the corresponding complex functions and arithmetic operations are defined in ‘gsl_complex_math.h’.

5.1 Complex numbers

Complex numbers are represented using the type gsl_complex. The internal representation of this type may vary across platforms and should not be accessed directly. The functions and macros described below allow complex numbers to be manipulated in a portable way.

For reference, the default form of the gsl_complex type is given by the following struct,

     typedef struct
     {
       double dat[2];
     } gsl_complex;

The real and imaginary part are stored in contiguous elements of a two element array. This eliminates any padding between the real and imaginary parts, dat[0] and dat[1], allowing the struct to be mapped correctly onto packed complex arrays.

Function gsl_complex gsl complex rect (double x, double y)

This function uses the rectangular cartesian components (x,y) to return the complex number z = x + iy.

Function gsl_complex gsl complex polar (double r, double theta)

This function returns the complex number z = r exp(iθ) = r(cos(θ) + i sin(θ)) from the polar representation (r,theta).

Macro GSL REAL (z) Macro GSL IMAG (z)

These macros return the real and imaginary parts of the complex number z.

Macro GSL SET COMPLEX (zp, x, y)

This macro uses the cartesian components (x,y) to set the real and imaginary parts of the complex number pointed to by zp. For example,

          GSL_SET_COMPLEX(&z, 3, 4)

sets z to be 3 + 4i.

Macro GSL SET REAL (zp,x) Macro GSL SET IMAG (zp,y)

These macros allow the real and imaginary parts of the complex number pointed to by zp to be set independently.

5.2 Properties of complex numbers

Function double gsl complex arg (gsl_complex z)

This function returns the argument of the complex number z, arg(z), where −π < arg(z) ≤ π.

Function double gsl complex abs (gsl_complex z)

This function returns the magnitude of the complex number z, |z|.

Function double gsl complex abs2 (gsl_complex z)

This function returns the squared magnitude of the complex number z, |z|^{2}.

Function double gsl complex logabs (gsl_complex z)

This function returns the natural logarithm of the magnitude of the complex number z, log |z|. It allows an accurate evaluation of log |z| when |z| is close to one. The direct evaluation of log(gsl_complex_abs(z)) would lead to a loss of precision in this case.

5.3 Complex arithmetic operators

Function gsl_complex gsl complex add (gsl_complex a, gsl_complex b)

This function returns the sum of the complex numbers a and b, z = a + b.

Function gsl_complex gsl complex sub (gsl_complex a, gsl_complex b)

This function returns the difference of the complex numbers a and b, z = a − b.

Function gsl_complex gsl complex mul (gsl_complex a, gsl_complex b)

This function returns the product of the complex numbers a and b, z = ab.

Function gsl_complex gsl complex div (gsl_complex a, gsl_complex b)

This function returns the quotient of the complex numbers a and b, z = a/b.

Function gsl_complex gsl complex add real (gsl_complex a, double x)

This function returns the sum of the complex number a and the real number x, z = a + x.

Function gsl_complex gsl complex sub real (gsl_complex a, double x)

This function returns the difference of the complex number a and the real number x, z = a − x.

Function gsl_complex gsl complex mul real (gsl_complex a, double x)

This function returns the product of the complex number a and the real number x, z = ax.

Function gsl_complex gsl complex div real (gsl_complex a, double x)

This function returns the quotient of the complex number a and the real number x, z = a/x.

Function gsl_complex gsl complex add imag (gsl_complex a, double y)

This function returns the sum of the complex number a and the imaginary number iy, z = a + iy.

Function gsl_complex gsl complex sub imag (gsl_complex a, double y)

This function returns the difference of the complex number a and the imaginary number iy, z = a − iy.

Function gsl_complex gsl complex mul imag (gsl_complex a, double y)

This function returns the product of the complex number a and the imaginary number iy, z = a ∗ (iy).

Function gsl_complex gsl complex div imag (gsl_complex a, double y)

This function returns the quotient of the complex number a and the imaginary number iy, z = a/(iy).

Function gsl_complex gsl complex conjugate (gsl_complex z)

This function returns the complex conjugate of the complex number z, z^{∗} = x − iy.

Function gsl_complex gsl complex inverse (gsl_complex z)

This function returns the inverse, or reciprocal, of the complex number z, 1/z = (x − iy)/(x^{2} + y^{2}).

Function gsl_complex gsl complex negative (gsl_complex z)

This function returns the negative of the complex number z, −z = (−x) + i(−y).

5.4 Elementary Complex Functions

Function gsl_complex gsl complex sqrt (gsl_complex z)

This function returns the square root of the complex number z, \sqrt{z}. The branch cut is the negative real axis. The result always lies in the right half of the complex plane.

Function gsl_complex gsl complex sqrt real (double x)

This function returns the complex square root of the real number x, where x may be negative.

Function gsl_complex gsl complex pow (gsl_complex z, gsl_complex a)

The function returns the complex number z raised to the complex power a, z^{a}. This is computed as exp(log(z) ∗ a) using complex logarithms and complex exponentials.

Function gsl_complex gsl complex pow real (gsl_complex z, double x)

This function returns the complex number z raised to the real power x, z^{x}.

Function gsl_complex gsl complex exp (gsl_complex z)

This function returns the complex exponential of the complex number z, exp(z).

Function gsl_complex gsl complex log (gsl_complex z)

This function returns the complex natural logarithm (base e) of the complex number z, log(z). The branch cut is the negative real axis.

Function gsl_complex gsl complex log10 (gsl_complex z)

This function returns the complex base-10 logarithm of the complex number z, log_{10}(z).

Function gsl_complex gsl complex log b (gsl_complex z, gsl_complex b)

This function returns the complex base-b logarithm of the complex number z, log_{b}(z). This quantity is computed as the ratio log(z)/ log(b).

5.5 Complex Trigonometric Functions

Function gsl_complex gsl complex sin (gsl_complex z)

This function returns the complex sine of the complex number z, sin(z) = (exp(iz) − exp(−iz))/(2i).

Function gsl_complex gsl complex cos (gsl_complex z)

This function returns the complex cosine of the complex number z, cos(z) = (exp(iz)+ exp(−iz))/2.

Function gsl_complex gsl complex tan (gsl_complex z)

This function returns the complex tangent of the complex number z, tan(z) = sin(z)/ cos(z).

Function gsl_complex gsl complex sec (gsl_complex z)

This function returns the complex secant of the complex number z, sec(z) = 1/ cos(z).

Function gsl_complex gsl complex csc (gsl_complex z)

This function returns the complex cosecant of the complex number z, csc(z) = 1/ sin(z).

Function gsl_complex gsl complex cot (gsl_complex z)

This function returns the complex cotangent of the complex number z, cot(z) = 1/ tan(z).

5.6 Inverse Complex Trigonometric Functions

Function gsl_complex gsl complex arcsin (gsl_complex z)

This function returns the complex arcsine of the complex number z, arcsin(z). The branch cuts are on the real axis, less than −1 and greater than 1.

Function gsl_complex gsl complex arcsin real (double z)

This function returns the complex arcsine of the real number z, arcsin(z). For z between −1 and 1, the function returns a real value in the range (−π, π]. For z less than −1 the result has a real part of −π/2 and a positive imaginary part. For z greater than 1 the result has a real part of π/2 and a negative imaginary part.

Function gsl_complex gsl complex arccos (gsl_complex z)

This function returns the complex arccosine of the complex number z, arccos(z). The branch cuts are on the real axis, less than −1 and greater than 1.

Function gsl_complex gsl complex arccos real (double z)

This function returns the complex arccosine of the real number z, arccos(z). For z between −1 and 1, the function returns a real value in the range [0, π]. For z less than −1 the result has a real part of π/2 and a negative imaginary part. For z greater than 1 the result is purely imaginary and positive.

Function gsl_complex gsl complex arctan (gsl_complex z)

This function returns the complex arctangent of the complex number z, arctan(z). The branch cuts are on the imaginary axis, below −i and above i.

Function gsl_complex gsl complex arcsec (gsl_complex z)

This function returns the complex arcsecant of the complex number z, arcsec(z) = arccos(1/z).

Function gsl_complex gsl complex arcsec real (double z)

This function returns the complex arcsecant of the real number z, arcsec(z) = arccos(1/z).

Function gsl_complex gsl complex arccsc (gsl_complex z)

This function returns the complex arccosecant of the complex number z, arccsc(z) = arcsin(1/z).

Function gsl_complex gsl complex arccsc real (double z)

This function returns the complex arccosecant of the real number z, arccsc(z) = arcsin(1/z).

Function gsl_complex gsl complex arccot (gsl_complex z)

This function returns the complex arccotangent of the complex number z, arccot(z) = arctan(1/z).

5.7 Complex Hyperbolic Functions

Function gsl_complex gsl complex sinh (gsl_complex z)

This function returns the complex hyperbolic sine of the complex number z, sinh(z) = (exp(z) − exp(−z))/2.

Function gsl_complex gsl complex cosh (gsl_complex z)

This function returns the complex hyperbolic cosine of the complex number z, cosh(z) = (exp(z) + exp(−z))/2.

Function gsl_complex gsl complex tanh (gsl_complex z)

This function returns the complex hyperbolic tangent of the complex number z, tanh(z) = sinh(z)/ cosh(z).

Function gsl_complex gsl complex sech (gsl_complex z)

This function returns the complex hyperbolic secant of the complex number z, sech(z) = 1/ cosh(z).

Function gsl_complex gsl complex csch (gsl_complex z)

This function returns the complex hyperbolic cosecant of the complex number z, csch(z) = 1/ sinh(z).

Function gsl_complex gsl complex coth (gsl_complex z)

This function returns the complex hyperbolic cotangent of the complex number z, coth(z) = 1/ tanh(z).

5.8 Inverse Complex Hyperbolic Functions

Function gsl_complex gsl complex arcsinh (gsl_complex z)

This function returns the complex hyperbolic arcsine of the complex number z, arcsinh(z). The branch cuts are on the imaginary axis, below −i and above i.

Function gsl_complex gsl complex arccosh (gsl_complex z)

This function returns the complex hyperbolic arccosine of the complex number z, arccosh(z). The branch cut is on the real axis, less than 1.

Function gsl_complex gsl complex arccosh real (double z)

This function returns the complex hyperbolic arccosine of the real number z, arccosh(z).

Function gsl_complex gsl complex arctanh (gsl_complex z)

This function returns the complex hyperbolic arctangent of the complex number z, arctanh(z). The branch cuts are on the real axis, less than −1 and greater than 1.

Function gsl_complex gsl complex arctanh real (double z)

This function returns the complex hyperbolic arctangent of the real number z, arctanh(z).

Function gsl_complex gsl complex arcsech (gsl_complex z)

This function returns the complex hyperbolic arcsecant of the complex number z, arcsech(z) = arccosh(1/z).

Function gsl_complex gsl complex arccsch (gsl_complex z)

This function returns the complex hyperbolic arccosecant of the complex number z, arccsch(z) = arcsin(1/z).

Function gsl_complex gsl complex arccoth (gsl_complex z)

This function returns the complex hyperbolic arccotangent of the complex number z, arccoth(z) = arctanh(1/z).

5.9 References and Further Reading

The implementations of the elementary and trigonometric functions are based on the following papers,

T. E. Hull, Thomas F. Fairgrieve, Ping Tak Peter Tang, "Implementing Complex Elementary Functions Using Exception Handling", ACM Transactions on Mathematical Software, Volume 20 (1994), pp 215-244, Corrigenda, p553 T. E. Hull, Thomas F. Fairgrieve, Ping Tak Peter Tang, "Implementing the complex arcsin and arccosine functions using exception handling", ACM Transactions on Mathematical Software, Volume 23 (1997) pp 299-335

The general formulas and details of branch cuts can be found in the following books, Abramowitz and Stegun, Handbook of Mathematical Functions, "Circular Functions in Terms of Real and Imaginary Parts", Formulas 4.3.55–58, "Inverse Circular Functions in Terms of Real and Imaginary Parts", Formulas 4.4.37–39, "Hyperbolic Functions in Terms of Real and Imaginary Parts", Formulas 4.5.49–52, "Inverse Hyperbolic Functions – relation to Inverse Circular Functions", Formulas 4.6.14–19. Dave Gillespie, Calc Manual, Free Software Foundation, ISBN 1-882114-18-3

6 Roots of Polynomials

This chapter describes functions for evaluating and solving polynomials. There are routines for finding real and complex roots of quadratic and cubic equations using analytic methods. An iterative polynomial solver is also available for finding the roots of general polynomials with real coefficients (of any order). The functions are declared in the header

file gsl_poly.h.

6.1 Polynomial evaluation

Function double gsl poly eval (const double c[], const int len, const
    double x)

This function evaluates the polynomial c[0] + c[1]x + c[2]x^{2} + … + c[len − 1]x^{len−1} using Horner’s method for stability. The function is inlined when possible.

6.2 Quadratic equations

Function int gsl poly solve quadratic (double a, double b, double c,

    double *x0, double *x1)

This function finds the real roots of the quadratic equation,

ax^{2} + bx + c = 0

The number of real roots (either zero or two) is returned, and their locations are stored in x0 and x1. If no real roots are found then x0 and x1 are not modified. When two real roots are found they are stored in x0 and x1 in ascending order. The case of coincident roots is not considered special. For example (x − 1)^{2} = 0 will have two roots, which happen to have exactly equal values. The number of roots found depends on the sign of the discriminant b^{2} −4ac. This will be subject to rounding and cancellation errors when computed in double precision, and will also be subject to errors if the coefficients of the polynomial are inexact. These errors may cause a discrete change in the number of roots. However, for polynomials with small integer coefficients the discriminant can always be computed exactly.

Function int gsl poly complex solve quadratic (double a, double b,

    double c, gsl_complex *z0, gsl_complex *z1)

This function finds the complex roots of the quadratic equation,

az^{2} + bz + c = 0

The number of complex roots is returned (always two) and the locations of the roots are stored in z0 and z1. The roots are returned in ascending order, sorted first by their real components and then by their imaginary components.

6.3 Cubic equations

Function int gsl poly solve cubic (double a, double b, double c, double
    *x0, double *x1, double *x2)

This function finds the real roots of the cubic equation,

x^{3} + ax^{2} + bx + c = 0

with a leading coefficient of unity. The number of real roots (either one or three) is returned, and their locations are stored in x0, x1 and x2. If one real root is found then only x0 is modified. When three real roots are found they are stored in x0, x1 and x2 in ascending order. The case of coincident roots is not considered special. For example, the equation (x − 1)^{3} = 0 will have three roots with exactly equal values.

Function int gsl poly complex solve cubic (double a, double b, double c,

    gsl_complex *z0, gsl_complex *z1, gsl_complex *z2)

This function finds the complex roots of the cubic equation,

z^{3} + az^{2} + bz + c = 0

The number of complex roots is returned (always three) and the locations of the roots are stored in z0, z1 and z2. The roots are returned in ascending order, sorted first by their real components and then by their imaginary components.

6.4 General polynomial equations

The roots of polynomial equations cannot be found analytically beyond the special cases of the quadratic, cubic and quartic equation. The algorithm described in this section uses an iterative method to find the approximate locations of roots of higher order polynomials.

Function gsl_poly_complex_workspace *

gsl poly complex workspace alloc (size_t n)

This function allocates space for a gsl_poly_complex_workspace struct and a workspace suitable for solving a polynomial with n coefficients using the routine

gsl_poly_complex_solve.

The function returns a pointer to the newly allocated gsl_poly_complex_workspace if no errors were detected, and a null pointer in the case of error.

Function void gsl poly complex workspace free

    (gsl_poly_complex_workspace * w)

This function frees all the memory associated with the workspace w.

Function int gsl poly complex solve (const double * a, size_t n,

    gsl_poly_complex_workspace * w, gsl_complex_packed_ptr z)

This function computes the roots of the general polynomial P (x) = a_{0} + a_{1}x + a_{2}x^{2} + ... + a_{n−1}x^{n−1} using balanced-QR reduction of the companion matrix. The parameter n specifies the length of the coefficient array. The coefficient of the highest order term must be non-zero. The function requires a workspace w of the appropriate size. The n − 1 roots are returned in the packed complex array z of length 2(n − 1), alternating real and imaginary parts. The function returns GSL_SUCCESS if all the roots are found and GSL_EFAILED if the QR reduction does not converge.

6.5 Examples

To demonstrate the use of the general polynomial solver we will take the polynomial P (x) = x^{5} − 1 which has the following roots,

1, e^{2πi/5}, e^{4πi/5}, e^{6πi/5}, e^{8πi/5}

The following program will find these roots.

#include <stdio.h>
#include <gsl/gsl_poly.h>
int
main (void)
{
  int i;
  /* coefficient of P(x) =  -1 + x^5  */
  double a[6] = { -1, 0, 0, 0, 0, 1 };
  double z[10];
  gsl_poly_complex_workspace * w
      = gsl_poly_complex_workspace_alloc (6);
  gsl_poly_complex_solve (a, 6, w, z);
  gsl_poly_complex_workspace_free (w);
  for (i = 0; i < 5; i++)
    {
      printf("z%d = %+.18f %+.18f\n",
             i, z[2*i], z[2*i+1]);
    }
  return 0;
}

The output of the program is,

bash$ ./a.out
z0 = -0.809016994374947451 +0.587785252292473137
z1 = -0.809016994374947451 -0.587785252292473137
z2 = +0.309016994374947451 +0.951056516295153642
z3 = +0.309016994374947451 -0.951056516295153642
z4 = +1.000000000000000000 +0.000000000000000000

which agrees with the analytic result, z_{n} = exp(2πni/5).

Chapter 6: Roots of Polynomials 29

6.6 References and Further Reading

The balanced-QR method and its error analysis is described in the following papers. R.S. Martin, G. Peters and J.H. Wilkinson, “The QR Algorithm for Real Hessenberg Matrices”, Numerische Mathematik, 14 (1970), 219–231. B.N. Parlett and C. Reinsch, “Balancing a Matrix for Calculation of Eigenvalues and Eigenvectors”, Numerische Mathematik, 13 (1969), 293–304. A. Edelman and H. Murakami, “Polynomial roots from companion matrix eigenvalues”, Mathematics of Computation, Vol. 64 No. 210 (1995), 763–776.

7 Special Functions

This chapter describes the GSL special function library. The library includes routines for calculating the values of Airy functions, Bessel functions, Clausen functions, Coulomb wave functions, Coupling coefficients, the Dawson function, Debye functions, Dilogarithms, Elliptic integrals, Jacobi elliptic functions, Error functions, Exponential integrals, Fermi-Dirac functions, Gamma functions, Gegenbauer functions, Hypergeometric functions, Laguerre functions, Legendre functions and Spherical Harmonics, the Psi (Digamma) Function, Synchrotron functions, Transport functions, Trigonometric functions and Zeta functions. Each routine also computes an estimate of the numerical error in the calculated value of the function.

The functions are declared in individual header files, such as ‘gsl_sf_airy.h’, ‘gsl_sf_bessel.h’, etc. The complete set of header files can be included using the file

‘gsl_sf.h’.

7.1 Usage

The special functions are available in two calling conventions, a natural form which returns the numerical value of the function and an error-handling form which returns an error code. The two types of function provide alternative ways of accessing the same underlying code.

The natural form returns only the value of the function and can be used directly in mathematical expressions.. For example, the following function call will compute the value of the Bessel function J_{0}(x),

     double y = gsl_sf_bessel_J0 (x);

There is no way to access an error code or to estimate the error using this method. To allow access to this information the alternative error-handling form stores the value and error in a modifiable argument,

     gsl_sf_result result;
     int status = gsl_sf_bessel_J0_e (x, &result);

The error-handling functions have the suffix _e. The returned status value indicates error conditions such as overflow, underflow or loss of precision. If there are no errors the errorhandling functions return GSL_SUCCESS.

7.2 The gsl sf result struct

The error handling form of the special functions always calculate an error estimate along with the value of the result. Therefore, structures are provided for amalgamating a value and error estimate. These structures are declared in the header file ‘gsl_sf_result.h’.

The gsl_sf_result struct contains value and error fields.

     typedef struct
     {
       double val;
       double err;
     } gsl_sf_result;

The field val contains the value and the field err contains an estimate of the absolute error in the value.

In some cases, an overflow or underflow can be detected and handled by a function. In this case, it may be possible to return a scaling exponent as well as an error/value pair in order to save the result from exceeding the dynamic range of the built-in types. The gsl_sf_result_e10 struct contains value and error fields as well as an exponent field such that the actual result is obtained as result * 10^(e10).

     typedef struct
     {
       double val;
       double err;
       int    e10;
     } gsl_sf_result_e10;

7.3 Modes

The goal of the library is to achieve double precision accuracy wherever possible. However the cost of evaluating some special functions to double precision can be significant, particularly where very high order terms are required. In these cases a mode argument allows the accuracy of the function to be reduced in order to improve performance. The following precision levels are available for the mode argument,

GSL_PREC_DOUBLE

Double-precision, a relative accuracy of approximately 2 × 10^{−16}.

GSL_PREC_SINGLE

Single-precision, a relative accuracy of approximately 1 × 10^{−7}.

GSL_PREC_APPROX

Approximate values, a relative accuracy of approximately 5 × 10^{−4}.

The approximate mode provides the fastest evaluation at the lowest accuracy.

7.4 Airy Functions and Derivatives

The Airy functions Ai(x) and Bi(x) are defined by the integral representations,

\begin{aligned}Ai(x) & = \frac{1}{π} ∫_{0}^{∞} cos(\frac{1}{3}t^{3} + xt)dt, \\ Bi(x) & = \frac{1}{π} ∫_{0}^{∞}(e^{−t^{3}/3} + sin(\frac{1}{3}t^{3} + xt))dt. \end{aligned}

For further information see Abramowitz & Stegun, Section 10.4. The Airy functions are defined in the header file ‘gsl_sf_airy.h’.

7.4.1 Airy Functions

Function double gsl sf airy Ai (double x, gsl_mode_t mode)

Function int gsl sf airy Ai e (double x, gsl_mode_t mode, gsl_sf_result *

result) These routines compute the Airy function Ai(x) with an accuracy specified by mode.

Function double gsl sf airy Bi (double x, gsl_mode_t mode)

Function int gsl sf airy Bi e (double x, gsl_mode_t mode, gsl_sf_result *

result) These routines compute the Airy function Bi(x) with an accuracy specified by mode.

Function double gsl sf airy Ai scaled (double x, gsl_mode_t mode) Function int gsl sf airy Ai scaled e (double x, gsl_mode_t mode,

         gsl_sf_result * result)

These routines compute a scaled version of the Airy function S_{A}(x)Ai(x). For x > 0 the scaling factor S_{A}(x) is exp(+(2/3)x^{3/2}), and is 1 for x < 0.

Function double gsl sf airy Bi scaled (double x, gsl_mode_t mode) Function int gsl sf airy Bi scaled e (double x, gsl_mode_t mode,

         gsl_sf_result * result)

These routines compute a scaled version of the Airy function S_{B}(x)Bi(x). For x > 0 the scaling factor S_{B}(x) is exp(−(2/3)x^{3/2}), and is 1 for x < 0.

7.4.2 Derivatives of Airy Functions

Function double gsl sf airy Ai deriv (double x, gsl_mode_t mode) Function int gsl sf airy Ai deriv e (double x, gsl_mode_t mode,

         gsl_sf_result * result)

These routines compute the Airy function derivative Ai′(x) with an accuracy specified by mode.

Function double gsl sf airy Bi deriv (double x, gsl_mode_t mode) Function int gsl sf airy Bi deriv e (double x, gsl_mode_t mode,

         gsl_sf_result * result)

These routines compute the Airy function derivative Bi′(x) with an accuracy specified by mode.

Function double gsl sf airy Ai deriv scaled (double x, gsl_mode_t mode) Function int gsl sf airy Ai deriv scaled e (double x, gsl_mode_t mode,

         gsl_sf_result * result)

These routines compute the derivative of the scaled Airy function S_{A}(x)Ai(x).

Function double gsl sf airy Bi deriv scaled (double x, gsl_mode_t mode) Function int gsl sf airy Bi deriv scaled e (double x, gsl_mode_t mode,

         gsl_sf_result * result)

These routines compute the derivative of the scaled Airy function S_{B}(x)Bi(x).

7.4.3 Zeros of Airy Functions

Function double gsl sf airy zero Ai (unsigned int s)

Function int gsl sf airy zero Ai e (unsigned int s, gsl_sf_result * result)

These routines compute the location of the s-th zero of the Airy function Ai(x).

Function double gsl sf airy zero Bi (unsigned int s)

Function int gsl sf airy zero Bi e (unsigned int s, gsl_sf_result * result)

These routines compute the location of the s-th zero of the Airy function Bi(x).

7.4.4 Zeros of Derivatives of Airy Functions

Function double gsl sf airy zero Ai deriv (unsigned int s)

Function int gsl sf airy zero Ai deriv e (unsigned int s, gsl_sf_result *

result) These routines compute the location of the s-th zero of the Airy function derivative Ai′(x).

Function double gsl sf airy zero Bi deriv (unsigned int s)

Function int gsl sf airy zero Bi deriv e (unsigned int s, gsl_sf_result *

result) These routines compute the location of the s-th zero of the Airy function derivative Bi′(x).

7.5 Bessel Functions

The routines described in this section compute the Cylindrical Bessel functions J_{n}(x), Y_{n}(x), Modified cylindrical Bessel functions I_{n}(x), K_{n}(x), Spherical Bessel functions j_{l}(x), y_{l}(x), and Modified Spherical Bessel functions i_{l}(x), k_{l}(x). For more information see Abramowitz & Stegun, Chapters 9 and 10. The Bessel functions are defined in the header

file ‘gsl_sf_bessel.h’.

7.5.1 Regular Cylindrical Bessel Functions

Function double gsl sf bessel J0 (double x)

Function int gsl sf bessel J0 e (double x, gsl_sf_result * result)

These routines compute the regular cylindrical Bessel function of zeroth order, J_{0}(x).

Function double gsl sf bessel J1 (double x)

Function int gsl sf bessel J1 e (double x, gsl_sf_result * result)

These routines compute the regular cylindrical Bessel function of first order, J_{1}(x).

Function double gsl sf bessel Jn (int n, double x)

Function int gsl sf bessel Jn e (int n, double x, gsl_sf_result * result)

These routines compute the regular cylindrical Bessel function of order n, J_{n}(x).

Function int gsl sf bessel Jn array (int nmin, int nmax, double x, double

result array[]) This routine computes the values of the regular cylindrical Bessel functions J_{n}(x) for n from nmin to nmax inclusive, storing the results in the array result array. The values are computed using recurrence relations, for efficiency, and therefore may differ slightly from the exact values.

7.5.2 Irregular Cylindrical Bessel Functions

Function double gsl sf bessel Y0 (double x)

Function int gsl sf bessel Y0 e (double x, gsl_sf_result * result)

These routines compute the irregular cylindrical Bessel function of zeroth order, Y_{0}(x), for x > 0.

Function double gsl sf bessel Y1 (double x)

Function int gsl sf bessel Y1 e (double x, gsl_sf_result * result)

These routines compute the irregular cylindrical Bessel function of first order, Y_{1}(x), for x > 0.

Function double gsl sf bessel Yn (int n,double x)

Function int gsl sf bessel Yn e (int n,double x, gsl_sf_result * result)

These routines compute the irregular cylindrical Bessel function of order n, Y_{n}(x), for x > 0.

Function int gsl sf bessel Yn array (int nmin, int nmax, double x,

double result array[]) This routine computes the values of the irregular cylindrical Bessel functions Y_{n}(x) for n from nmin to nmax inclusive, storing the results in the array result array. The domain of the function is x > 0. The values are computed using recurrence relations, for efficiency, and therefore may differ slightly from the exact values.

7.5.3 Regular Modified Cylindrical Bessel Functions

Function double gsl sf bessel I0 (double x)

Function int gsl sf bessel I0 e (double x, gsl_sf_result * result)

These routines compute the regular modified cylindrical Bessel function of zeroth order, I_{0}(x).

Function double gsl sf bessel I1 (double x)

Function int gsl sf bessel I1 e (double x, gsl_sf_result * result)

These routines compute the regular modified cylindrical Bessel function of first order, I_{1}(x).

Function double gsl sf bessel In (int n, double x)

Function int gsl sf bessel In e (int n, double x, gsl_sf_result * result)

These routines compute the regular modified cylindrical Bessel function of order n, I_{n}(x).

Function int gsl sf bessel In array (int nmin, int nmax, double x, double

result array[]) This routine computes the values of the regular modified cylindrical Bessel functions I_{n}(x) for n from nmin to nmax inclusive, storing the results in the array result array. The start of the range nmin must be positive or zero. The values are computed using recurrence relations, for efficiency, and therefore may differ slightly from the exact values.

Function double gsl sf bessel I0 scaled (double x)

Function int gsl sf bessel I0 scaled e (double x, gsl_sf_result * result)

These routines compute the scaled regular modified cylindrical Bessel function of zeroth order exp(−|x|)I_{0}(x).

Function double gsl sf bessel I1 scaled (double x)

Function int gsl sf bessel I1 scaled e (double x, gsl_sf_result * result)

These routines compute the scaled regular modified cylindrical Bessel function of first order exp(−|x|)I_{1}(x).

Function double gsl sf bessel In scaled (int n, double x)

Function int gsl sf bessel In scaled e (int n, double x, gsl_sf_result *

result) These routines compute the scaled regular modified cylindrical Bessel function of order n, exp(−|x|)I_{n}(x)

Function int gsl sf bessel In scaled array (int nmin, int nmax, double x,

double result array[]) This routine computes the values of the scaled regular cylindrical Bessel functions exp(−|x|)I_{n}(x) for n from nmin to nmax inclusive, storing the results in the array result array. The start of the range nmin must be positive or zero. The values are computed using recurrence relations, for efficiency, and therefore may differ slightly from the exact values.

7.5.4 Irregular Modified Cylindrical Bessel Functions

Function double gsl sf bessel K0 (double x)

Function int gsl sf bessel K0 e (double x, gsl_sf_result * result)

These routines compute the irregular modified cylindrical Bessel function of zeroth order, K_{0}(x), for x > 0.

Function double gsl sf bessel K1 (double x)

Function int gsl sf bessel K1 e (double x, gsl_sf_result * result)

These routines compute the irregular modified cylindrical Bessel function of first order, K_{1}(x), for x > 0.

Function double gsl sf bessel Kn (int n, double x)

Function int gsl sf bessel Kn e (int n, double x, gsl_sf_result * result)

These routines compute the irregular modified cylindrical Bessel function of order n, K_{n}(x), for x > 0.

Function int gsl sf bessel Kn array (int nmin, int nmax, double x,

double result array[]) This routine computes the values of the irregular modified cylindrical Bessel functions K_{n}(x) for n from nmin to nmax inclusive, storing the results in the array result array. The start of the range nmin must be positive or zero. The domain of the function is x > 0. The values are computed using recurrence relations, for efficiency, and therefore may differ slightly from the exact values.

Function double gsl sf bessel K0 scaled (double x)

Function int gsl sf bessel K0 scaled e (double x, gsl_sf_result * result)

These routines compute the scaled irregular modified cylindrical Bessel function of zeroth order exp(x)K_{0}(x) for x > 0.

Function double gsl sf bessel K1 scaled (double x)

Function int gsl sf bessel K1 scaled e (double x, gsl_sf_result * result)

These routines compute the scaled irregular modified cylindrical Bessel function of first order exp(x)K_{1}(x) for x > 0.

Function double gsl sf bessel Kn scaled (int n, double x)

Function int gsl sf bessel Kn scaled e (int n, double x, gsl_sf_result *

result) These routines compute the scaled irregular modified cylindrical Bessel function of order n, exp(x)K_{n}(x), for x > 0.

Function int gsl sf bessel Kn scaled array (int nmin, int nmax, double

x, double result array[]) This routine computes the values of the scaled irregular cylindrical Bessel functions exp(x)K_{n}(x) for n from nmin to nmax inclusive, storing the results in the array result array. The start of the range nmin must be positive or zero. The domain of the function is x > 0. The values are computed using recurrence relations, for efficiency, and therefore may differ slightly from the exact values.

7.5.5 Regular Spherical Bessel Functions

Function double gsl sf bessel j0 (double x)

Function int gsl sf bessel j0 e (double x, gsl_sf_result * result)

These routines compute the regular spherical Bessel function of zeroth order, j_{0}(x) = sin(x)/x.

Function double gsl sf bessel j1 (double x)

Function int gsl sf bessel j1 e (double x, gsl_sf_result * result)

These routines compute the regular spherical Bessel function of first order, j_{1}(x) = (sin(x)/x − cos(x))/x.

Function double gsl sf bessel j2 (double x)

Function int gsl sf bessel j2 e (double x, gsl_sf_result * result)

These routines compute the regular spherical Bessel function of second order, j_{2}(x) = ((3/x^{2} − 1) sin(x) − 3 cos(x)/x)/x.

Function double gsl sf bessel jl (int l, double x)

Function int gsl sf bessel jl e (int l, double x, gsl_sf_result * result)

These routines compute the regular spherical Bessel function of order l, j_{l}(x), for l ≥ 0 and x ≥ 0.

Function int gsl sf bessel jl array (int lmax, double x, double

result array[]) This routine computes the values of the regular spherical Bessel functions j_{l}(x) for l from 0 to lmax inclusive for lmax ≥ 0 and x ≥ 0, storing the results in the array result array. The values are computed using recurrence relations, for efficiency, and therefore may differ slightly from the exact values.

Function int gsl sf bessel jl steed array (int lmax, double x, double *

jl x array) This routine uses Steed’s method to compute the values of the regular spherical Bessel functions j_{l}(x) for l from 0 to lmax inclusive for lmax ≥ 0 and x ≥ 0, storing the results in the array result array. The Steed/Barnett algorithm is described in Comp. Phys. Comm. 21, 297 (1981). Steed’s method is more stable than the recurrence used in the other functions but is also slower.

7.5.6 Irregular Spherical Bessel Functions

Function double gsl sf bessel y0 (double x)

Function int gsl sf bessel y0 e (double x, gsl_sf_result * result)

These routines compute the irregular spherical Bessel function of zeroth order, y_{0}(x) = − cos(x)/x.

Function double gsl sf bessel y1 (double x)

Function int gsl sf bessel y1 e (double x, gsl_sf_result * result)

These routines compute the irregular spherical Bessel function of first order, y_{1}(x) = −(cos(x)/x + sin(x))/x.

Function double gsl sf bessel y2 (double x)

Function int gsl sf bessel y2 e (double x, gsl_sf_result * result)

These routines compute the irregular spherical Bessel function of second order, y_{2}(x) = (−3/x^{2} + 1/x) cos(x) − (3/x^{2}) sin(x).

Function double gsl sf bessel yl (int l, double x)

Function int gsl sf bessel yl e (int l, double x, gsl_sf_result * result)

These routines compute the irregular spherical Bessel function of order l, y_{l}(x), for l ≥ 0.

Function int gsl sf bessel yl array (int lmax, double x, double

result array[]) This routine computes the values of the irregular spherical Bessel functions y_{l}(x) for l from 0 to lmax inclusive for lmax ≥ 0, storing the results in the array result array. The values are computed using recurrence relations, for efficiency, and therefore may differ slightly from the exact values.

7.5.7 Regular Modified Spherical Bessel Functions

The regular modified spherical Bessel functions i_{l}(x) are related to the modified Bessel functions of fractional order, i_{l}(x) = \sqrt{π/(2x)}I_{l+1/2}(x)

Function double gsl sf bessel i0 scaled (double x)

Function int gsl sf bessel i0 scaled e (double x, gsl_sf_result * result)

These routines compute the scaled regular modified spherical Bessel function of zeroth order, exp(−|x|)i_{0}(x).

Function double gsl sf bessel i1 scaled (double x)

Function int gsl sf bessel i1 scaled e (double x, gsl_sf_result * result)

These routines compute the scaled regular modified spherical Bessel function of first order, exp(−|x|)i_{1}(x).

Function double gsl sf bessel i2 scaled (double x)

Function int gsl sf bessel i2 scaled e (double x, gsl_sf_result * result)

These routines compute the scaled regular modified spherical Bessel function of second order, exp(−|x|)i_{2}(x)

Function double gsl sf bessel il scaled (int l, double x)

Function int gsl sf bessel il scaled e (int l, double x, gsl_sf_result *

result) These routines compute the scaled regular modified spherical Bessel function of order l, exp(−|x|)i_{l}(x)

Function int gsl sf bessel il scaled array (int lmax, double x, double

result array[]) This routine computes the values of the scaled regular modified cylindrical Bessel functions exp(−|x|)i_{l}(x) for l from 0 to lmax inclusive for lmax ≥ 0, storing the results in the array result array. The values are computed using recurrence relations, for efficiency, and therefore may differ slightly from the exact values.

7.5.8 Irregular Modified Spherical Bessel Functions

The irregular modified spherical Bessel functions k_{l}(x) are related to the irregular modified Bessel functions of fractional order, k_{l}(x) = \sqrt{π/(2x)}K_{l+1/2}(x).

Function double gsl sf bessel k0 scaled (double x)

Function int gsl sf bessel k0 scaled e (double x, gsl_sf_result * result)

These routines compute the scaled irregular modified spherical Bessel function of zeroth order, exp(x)k_{0}(x), for x > 0.

Function double gsl sf bessel k1 scaled (double x)

Function int gsl sf bessel k1 scaled e (double x, gsl_sf_result * result)

These routines compute the scaled irregular modified spherical Bessel function of first order, exp(x)k_{1}(x), for x > 0.

Function double gsl sf bessel k2 scaled (double x)

Function int gsl sf bessel k2 scaled e (double x, gsl_sf_result * result)

These routines compute the scaled irregular modified spherical Bessel function of second order, exp(x)k_{2}(x), for x > 0.

Function double gsl sf bessel kl scaled (int l, double x)

Function int gsl sf bessel kl scaled e (int l, double x, gsl_sf_result *

result) These routines compute the scaled irregular modified spherical Bessel function of order l, exp(x)k_{l}(x), for x > 0.

Function int gsl sf bessel kl scaled array (int lmax, double x, double

result array[]) This routine computes the values of the scaled irregular modified spherical Bessel functions exp(x)k_{l}(x) for l from 0 to lmax inclusive for lmax ≥ 0 and x > 0, storing the results in the array result array. The values are computed using recurrence relations, for efficiency, and therefore may differ slightly from the exact values.

7.5.9 Regular Bessel Function - Fractional Order

Function double gsl sf bessel Jnu (double nu, double x)

Function int gsl sf bessel Jnu e (double nu, double x, gsl_sf_result *

result) These routines compute the regular cylindrical Bessel function of fractional order nu, J_{ν}(x).

Function int gsl sf bessel sequence Jnu e (double nu, gsl_mode_t mode,

    size_t size, double v[])

This function computes the regular cylindrical Bessel function of fractional order ν, J_{ν}(x), evaluated at a series of x values. The array v of length size contains the x values. They are assumed to be strictly ordered and positive. The array is overwritten with the values of J_{ν}(x_{i}).

7.5.10 Irregular Bessel Functions - Fractional Order

Function double gsl sf bessel Ynu (double nu, double x)

Function int gsl sf bessel Ynu e (double nu, double x, gsl_sf_result *

result) These routines compute the irregular cylindrical Bessel function of fractional order nu, Y_{ν}(x).

7.5.11 Regular Modified Bessel Functions - Fractional Order

Function double gsl sf bessel Inu (double nu, double x)

Function int gsl sf bessel Inu e (double nu, double x, gsl_sf_result *

result) These routines compute the regular modified Bessel function of fractional order nu, I_{ν}(x) for x > 0, ν > 0.

Function double gsl sf bessel Inu scaled (double nu, double x) Function int gsl sf bessel Inu scaled e (double nu, double x,

    gsl_sf_result * result)

These routines compute the scaled regular modified Bessel function of fractional order nu, exp(−|x|)I_{ν}(x) for x > 0, ν > 0.

7.5.12 Irregular Modified Bessel Functions - Fractional Order

Function double gsl sf bessel Knu (double nu, double x)

Function int gsl sf bessel Knu e (double nu, double x, gsl_sf_result *

result) These routines compute the irregular modified Bessel function of fractional order nu, K_{ν}(x) for x > 0, ν > 0.

Function double gsl sf bessel lnKnu (double nu, double x)

Function int gsl sf bessel lnKnu e (double nu, double x, gsl_sf_result *

result) These routines compute the logarithm of the irregular modified Bessel function of fractional order nu, ln(K_{ν}(x)) for x > 0, ν > 0.

Function double gsl sf bessel Knu scaled (double nu, double x) Function int gsl sf bessel Knu scaled e (double nu, double x,

    gsl_sf_result * result)

These routines compute the scaled irregular modified Bessel function of fractional order nu, exp(+|x|)K_{ν}(x) for x > 0, ν > 0.

7.5.13 Zeros of Regular Bessel Functions

Function double gsl sf bessel zero J0 (unsigned int s)

Function int gsl sf bessel zero J0 e (unsigned int s, gsl_sf_result *

result) These routines compute the location of the s-th positive zero of the Bessel function J_{0}(x).

Function double gsl sf bessel zero J1 (unsigned int s)

Function int gsl sf bessel zero J1 e (unsigned int s, gsl_sf_result *

result) These routines compute the location of the s-th positive zero of the Bessel function J_{1}(x).

Function double gsl sf bessel zero Jnu (double nu, unsigned int s)

Function int gsl sf bessel zero Jnu e (double nu, unsigned int s,

         gsl_sf_result * result)

These routines compute the location of the s-th positive zero of the Bessel function J_{ν}(x).

7.6 Clausen Functions

The Clausen function is defined by the following integral,

Cl_{2}(x) = − ∫_{0}^{x} dt log(2 sin(t/2))

It is related to the dilogarithm by Cl_{2}(θ) = ImLi_{2}(exp(iθ)). The Clausen functions are declared in the header file ‘gsl_sf_clausen.h’.

Function double gsl sf clausen (double x)

Function int gsl sf clausen e (double x, gsl_sf_result * result)

These routines compute the Clausen integral Cl_{2}(x).

7.7 Coulomb Functions

The Coulomb functions are declared in the header file ‘gsl_sf_coulomb.h’. Both bound state and scattering solutions are available.

7.7.1 Normalized Hydrogenic Bound States

Function double gsl sf hydrogenicR 1 (double Z, double r)

Function int gsl sf hydrogenicR 1 e (double Z, double r, gsl_sf_result *

result) These routines compute the lowest-order normalized hydrogenic bound state radial wavefunction R_{1} := 2Z\sqrt{Z} exp(−Zr).

Function double gsl sf hydrogenicR (int n, int l, double Z, double r) Function int gsl sf hydrogenicR e (int n, int l, double Z, double r,

         gsl_sf_result * result)

These routines compute the n-th normalized hydrogenic bound state radial wavefunction,

R_{n} := \frac{2Z^{3/2}}{n^{2}} (\frac{2Z}{n} )^{l} \sqrt{\frac{(n − l − 1)!}{(n + l)!}} exp(−Zr/n)L_{n−l−1}^{2l+1}(2Z/nr).

The normalization is chosen such that the wavefunction ψ is given by ψ(n, l, r) = R_{n}Y_{lm}.

7.7.2 Coulomb Wave Functions

The Coulomb wave functions F_{L}(η, x), G_{L}(η, x) are described in Abramowitz & Stegun, Chapter 14. Because there can be a large dynamic range of values for these functions, overflows are handled gracefully. If an overflow occurs, GSL_EOVRFLW is signalled and exponent(s) are returned through the modifiable parameters exp F, exp G. The full solution can be reconstructed from the following relations,

F_{L}(η, x) = fc[k_{L}] ∗ exp(exp_{F})
G_{L}(η, x) = gc[k_{L}] ∗ exp(exp_{G})
F′_{L}(η, x) = fcp[k_{L}] ∗ exp(exp_{F})
G′_{L}(η, x) = gcp[k_{L}] ∗ exp(exp_{G})

Function int gsl sf coulomb wave FG e (double eta, double x, double

    L F, int k, gsl_sf_result * F, gsl_sf_result * Fp, gsl_sf_result * G,
    gsl_sf_result * Gp, double * exp F, double * exp G)

This function computes the coulomb wave functions F_{L}(η, x), G_{L−k}(η, x) and their derivatives with respect to x, F′_{L}(η, x) G′_{L−k}(η, x). The parameters are restricted to L, L − k > −1/2, x > 0 and integer k. Note that L itself is not restricted to being an integer. The results are stored in the parameters F, G for the function values and Fp, Gp for the derivative values. If an overflow occurs, GSL_EOVRFLW is returned and scaling exponents are stored in the modifiable parameters exp F, exp G.

Function int gsl sf coulomb wave F array (double L min, int kmax,

    double eta, double x, double fc array[], double * F exponent)

This function computes the function F_{L}(eta, x) for L = Lmin … Lmin+kmax storing the results in fc array. In the case of overflow the exponent is stored in F exponent.

Function int gsl sf coulomb wave FG array (double L min, int kmax,

    double eta, double x, double fc array[], double gc array[], double *

F exponent, double * G exponent) This function computes the functions F_{L}(η, x), G_{L}(η, x) for L = Lmin … Lmin + kmax storing the results in fc array and gc array. In the case of overflow the exponents are stored in F exponent and G exponent.

Function int gsl sf coulomb wave FGp array (double L min, int kmax,

    double eta, double x, double fc array[], double fcp array[], double

gc array[], double gcp array[], double * F exponent, double * G exponent) This function computes the functions F_{L}(η, x), G_{L}(η, x) and their derivatives F′_{L}(η, x), G′_{L}(η, x) for L = Lmin … Lmin + kmax storing the results in fc array, gc array, fcp array and gcp array. In the case of overflow the exponents are stored in F exponent and G exponent.

Function int gsl sf coulomb wave sphF array (double L min, int kmax,

    double eta, double x, double fc array[], double F exponent[])

This function computes the Coulomb wave function divided by the argument F_{L}(η, x)/x for L = Lmin … Lmin + kmax, storing the results in fc array. In the case of overflow the exponent is stored in F exponent. This function reduces to spherical Bessel functions in the limit η → 0.

7.7.3 Coulomb Wave Function Normalization Constant

The Coulomb wave function normalization constant is defined in Abramowitz 14.1.7.

Function int gsl sf coulomb CL e (double L, double eta, gsl_sf_result *

result) This function computes the Coulomb wave function normalization constant C_{L}(η) for L > −1.

Function int gsl sf coulomb CL array (double Lmin, int kmax, double

eta, double cl[]) This function computes the coulomb wave function normalization constant C_{L}(η) for L = Lmin … Lmin + kmax, Lmin > −1.

7.8 Coupling Coefficients

The Wigner 3-j, 6-j and 9-j symbols give the coupling coefficients for combined angular momentum vectors. Since the arguments of the standard coupling coefficient functions are integer or half-integer, the arguments of the following functions are, by convention, integers equal to twice the actual spin value. For information on the 3-j coefficients see Abramowitz & Stegun, Section 27.9. The functions described in this section are declared in the header

file ‘gsl_sf_coupling.h’.

7.8.1 3-j Symbols

Function double gsl sf coupling 3j (int two ja, int two jb, int two jc, int

two ma, int two mb, int two mc)

Function int gsl sf coupling 3j e (int two ja, int two jb, int two jc, int

         two ma, int two mb, int two mc, gsl_sf_result * result)

These routines compute the Wigner 3-j coefficient,

( \begin{matrix}ja & jb & jc \\ ma & mb & mc \end{matrix} )

where the arguments are given in half-integer units, ja = two ja/2, ma = two ma/2, etc.

7.8.2 6-j Symbols

Function double gsl sf coupling 6j (int two ja, int two jb, int two jc, int

two jd, int two je, int two jf )

Function int gsl sf coupling 6j e (int two ja, int two jb, int two jc, int

         two jd, int two je, int two jf, gsl_sf_result * result)

These routines compute the Wigner 6-j coefficient,

( \begin{matrix}ja & jb & jc \\ jd & je & jf \end{matrix} )

where the arguments are given in half-integer units, ja = two ja/2, ma = two ma/2, etc.

7.8.3 9-j Symbols

Function double gsl sf coupling 9j (int two ja, int two jb, int two jc, int

two jd, int two je, int two jf, int two jg, int two jh, int two ji)

Function int gsl sf coupling 9j e (int two ja, int two jb, int two jc, int two jd, int two je, int two jf, int two jg, int two jh, int two ji,

         gsl_sf_result * result)

These routines compute the Wigner 9-j coefficient,

( \begin{matrix}ja & jb & jc \\ jd & je & jf \\ jg & jh & ji \end{matrix} )

where the arguments are given in half-integer units, ja = two ja/2, ma = two ma/2, etc.

7.9 Dawson Function

The Dawson integral is defined by exp(−x^{2}) ∫_{0}^{x} dt exp(t^{2}). A table of Dawson’s integral can be found in Abramowitz & Stegun, Table 7.5. The Dawson functions are declared in

the header file ‘gsl_sf_dawson.h’.

Function double gsl sf dawson (double x)

Function int gsl sf dawson e (double x, gsl_sf_result * result)

These routines compute the value of Dawson’s integral for x.

7.10 Debye Functions

The Debye functions are defined by the integral D_{n}(x) = n/x^{n} ∫_{0}^{x} dt(t^{n}/(e^{t} − 1)). For further information see Abramowitz & Stegun, Section 27.1. The Debye functions are declared in the header file ‘gsl_sf_debye.h’.

Function double gsl sf debye 1 (double x)

Function int gsl sf debye 1 e (double x, gsl_sf_result * result)

These routines compute the first-order Debye function D_{1}(x) = (1/x) ∫_{0}^{x} dt(t/(e^{t}−1)).

Function double gsl sf debye 2 (double x)

Function int gsl sf debye 2 e (double x, gsl_sf_result * result)

These routines compute the second-order Debye function D_{2}(x) = (2/x^{2}) ∫_{0}^{x} dt(t^{2}/(e^{t} − 1)).

Function double gsl sf debye 3 (double x)

Function int gsl sf debye 3 e (double x, gsl_sf_result * result)

These routines compute the third-order Debye function D_{3}(x) = (3/x^{3}) ∫_{0}^{x} dt(t^{3}/(e^{t} − 1)).

Function double gsl sf debye 4 (double x)

Function int gsl sf debye 4 e (double x, gsl_sf_result * result)

These routines compute the fourth-order Debye function D_{4}(x) = (4/x^{4}) ∫_{0}^{x} dt(t^{4}/(e^{t}− 1)).

7.11 Dilogarithm

The functions described in this section are declared in the header file ‘gsl_sf_dilog.h’.

7.11.1 Real Argument

Function double gsl sf dilog (double x)

Function int gsl sf dilog e (double x, gsl_sf_result * result)

These routines compute the dilogarithm for a real argument. In Lewin’s notation this is Li_{2}(x), the real part of the dilogarithm of a real x. It is defined by the integral representation Li_{2}(x) = −Re ∫_{0}^{x} ds log(1 − s)/s. Note that Im(Li_{2}(x)) = 0 for x ≤ 1, and −π log(x) for x > 1.

7.11.2 Complex Argument

Function int gsl sf complex dilog e (double r, double theta,

    gsl_sf_result * result re, gsl_sf_result * result im)

This function computes the full complex-valued dilogarithm for the complex argument z = r exp(iθ). The real and imaginary parts of the result are returned in result re, result im.

7.12 Elementary Operations

The following functions allow for the propagation of errors when combining quantities by multiplication. The functions are declared in the header file ‘gsl_sf_elementary.h’.

Function int gsl sf multiply e (double x, double y, gsl_sf_result * result)

This function multiplies x and y storing the product and its associated error in result.

Function int gsl sf multiply err e (double x, double dx, double y, double
    dy, gsl_sf_result * result)

This function multiplies x and y with associated absolute errors dx and dy. The product xy ± xy\sqrt{(dx/x)^{2} + (dy/y)^{2}} is stored in result.

7.13 Elliptic Integrals

The functions described in this section are declared in the header file ‘gsl_sf_ellint.h’.

7.13.1 Definition of Legendre Forms

The Legendre forms of elliptic integrals F (ϕ, k), E(ϕ, k) and P (ϕ, k, n) are defined by,

F (ϕ, k) = ∫_{0}^{ϕ} dt\frac{1}{\sqrt{(1 − k^{2} sin^{2}(t))}}

E(ϕ, k) = ∫_{0}^{ϕ} dt\sqrt{(1 − k^{2} sin^{2}(t))}
P (ϕ, k, n) = ∫_{0}^{ϕ} dt\frac{1}{(1 + n sin^{2}(t))\sqrt{1 − k^{2} sin^{2}(t)}}

The complete Legendre forms are denoted by K(k) = F (π/2, k) and E(k) = E(π/2, k). Further information on the Legendre forms of elliptic integrals can be found in Abramowitz & Stegun, Chapter 17. The notation used here is based on Carlson, Numerische Mathematik 33 (1979) 1 and differs slightly from that used by Abramowitz & Stegun.

7.13.2 Definition of Carlson Forms

The Carlson symmetric forms of elliptical integrals RC(x, y), RD(x, y, z), RF (x, y, z) and RJ(x, y, z, p) are defined by,

\begin{aligned}RC(x, y) & = 1/2 ∫_{0}^{∞} dt(t + x)^{−1/2}(t + y)^{−1} \\ RD(x, y, z) & = 3/2 ∫_{0}^{∞} dt(t + x)^{−1/2}(t + y)^{−1/2}(t + z)^{−3/2} \\ RF (x, y, z) & = 1/2 ∫_{0}^{∞} dt(t + x)^{−1/2}(t + y)^{−1/2}(t + z)^{−1/2} \\ RJ(x, y, z, p) & = 3/2 ∫_{0}^{∞} dt(t + x)^{−1/2}(t + y)^{−1/2}(t + z)^{−1/2}(t + p)^{−1} \end{aligned}

7.13.3 Legendre Form of Complete Elliptic Integrals

Function double gsl sf ellint Kcomp (double k, gsl_mode_t mode) Function int gsl sf ellint Kcomp e (double k, gsl_mode_t mode,

         gsl_sf_result * result)

These routines compute the complete elliptic integral K(k) to the accuracy specified by the mode variable mode.

Function double gsl sf ellint Ecomp (double k, gsl_mode_t mode) Function int gsl sf ellint Ecomp e (double k, gsl_mode_t mode,

         gsl_sf_result * result)

These routines compute the complete elliptic integral E(k) to the accuracy specified by the mode variable mode.

7.13.4 Legendre Form of Incomplete Elliptic Integrals

Function double gsl sf ellint F (double phi, double k, gsl_mode_t mode)

Function int gsl sf ellint F e (double phi, double k, gsl_mode_t mode,

         gsl_sf_result * result)

These routines compute the incomplete elliptic integral F (ϕ, k) to the accuracy specified by the mode variable mode.

Function double gsl sf ellint E (double phi, double k, gsl_mode_t mode)

Function int gsl sf ellint E e (double phi, double k, gsl_mode_t mode,

         gsl_sf_result * result)

These routines compute the incomplete elliptic integral E(ϕ, k) to the accuracy specified by the mode variable mode.

Function double gsl sf ellint P (double phi, double k, double n, gsl_mode_t

mode)

Function int gsl sf ellint P e (double phi, double k, double n, gsl_mode_t
         mode, gsl_sf_result * result)

These routines compute the incomplete elliptic integral P (ϕ, k, n) to the accuracy specified by the mode variable mode.

Function double gsl sf ellint D (double phi, double k, double n, gsl_mode_t

mode)

Function int gsl sf ellint D e (double phi, double k, double n, gsl_mode_t
         mode, gsl_sf_result * result)

These functions compute the incomplete elliptic integral D(ϕ, k, n) which is defined through the Carlson form RD(x, y, z) by the following relation,

D(ϕ, k, n) = RD(1 − sin^{2}(ϕ), 1 − k^{2} sin^{2}(ϕ), 1).

7.13.5 Carlson Forms

Function double gsl sf ellint RC (double x, double y, gsl_mode_t mode)

Function int gsl sf ellint RC e (double x, double y, gsl_mode_t mode,

         gsl_sf_result * result)

These routines compute the incomplete elliptic integral RC(x, y) to the accuracy specified by the mode variable mode.

Function double gsl sf ellint RD (double x, double y, double z, gsl_mode_t

mode)

Function int gsl sf ellint RD e (double x, double y, double z, gsl_mode_t
         mode, gsl_sf_result * result)

These routines compute the incomplete elliptic integral RD(x, y, z) to the accuracy specified by the mode variable mode.

Function double gsl sf ellint RF (double x, double y, double z, gsl_mode_t

mode)

Function int gsl sf ellint RF e (double x, double y, double z, gsl_mode_t
         mode, gsl_sf_result * result)

These routines compute the incomplete elliptic integral RF (x, y, z) to the accuracy specified by the mode variable mode.

Function double gsl sf ellint RJ (double x, double y, double z, double p,
         gsl_mode_t mode)
Function int gsl sf ellint RJ e (double x, double y, double z, double p,
         gsl_mode_t mode, gsl_sf_result * result)

These routines compute the incomplete elliptic integral RJ(x, y, z, p) to the accuracy specified by the mode variable mode.

7.14 Elliptic Functions (Jacobi)

The Jacobian Elliptic functions are defined in Abramowitz & Stegun, Chapter 16. The functions are declared in the header file ‘gsl_sf_elljac.h’.

Function int gsl sf elljac e (double u, double m, double * sn, double * cn,
         double * dn)

This function computes the Jacobian elliptic functions sn(u|m), cn(u|m), dn(u|m) by descending Landen transformations.

7.15 Error Functions

The error function is described in Abramowitz & Stegun, Chapter 7. The functions in this section are declared in the header file ‘gsl_sf_erf.h’.

7.15.1 Error Function

Function double gsl sf erf (double x)

Function int gsl sf erf e (double x, gsl_sf_result * result)

These routines compute the error function erf(x) = (2/\sqrt{π}) ∫_{0}^{x} dt exp(−t^{2}).

7.15.2 Complementary Error Function

Function double gsl sf erfc (double x)

Function int gsl sf erfc e (double x, gsl_sf_result * result)

These routines compute the complementary error function erfc(x) = 1 − erf(x) = (2/\sqrt{π}) ∫_{x}^{∞} exp(−t^{2}).

7.15.3 Log Complementary Error Function

Function double gsl sf log erfc (double x)

Function int gsl sf log erfc e (double x, gsl_sf_result * result)

These routines compute the logarithm of the complementary error function log(erfc(x)).

7.15.4 Probability functions

The probability functions for the Normal or Gaussian distribution are described in Abramowitz & Stegun, Section 26.2.

Function double gsl sf erf Z (double x)

Function int gsl sf erf Z e (double x, gsl_sf_result * result)

These routines compute the Gaussian probability function Z(x) = (1/(2π)) exp(−x^{2}/2).

Function double gsl sf erf Q (double x)

Function int gsl sf erf Q e (double x, gsl_sf_result * result)

These routines compute the upper tail of the Gaussian probability function Q(x) = (1/(2π)) ∫_{x}^{∞} dt exp(−t^{2}/2).

7.16 Exponential Functions

The functions described in this section are declared in the header file ‘gsl_sf_exp.h’.

7.16.1 Exponential Function

Function double gsl sf exp (double x)

Function int gsl sf exp e (double x, gsl_sf_result * result)

These routines provide an exponential function exp(x) using GSL semantics and error checking.

Function int gsl sf exp e10 e (double x, gsl_sf_result_e10 * result)

This function computes the exponential exp(x) using the gsl_sf_result_e10 type to return a result with extended range. This function may be useful if the value of exp(x) would overflow the numeric range of double.

Function double gsl sf exp mult (double x, double y)

Function int gsl sf exp mult e (double x, double y, gsl_sf_result * result)

These routines exponentiate x and multiply by the factor y to return the product y exp(x).

Function int gsl sf exp mult e10 e (const double x, const double y,
    gsl_sf_result_e10 * result)

This function computes the product y exp(x) using the gsl_sf_result_e10 type to return a result with extended numeric range.

7.16.2 Relative Exponential Functions

Function double gsl sf expm1 (double x)

Function int gsl sf expm1 e (double x, gsl_sf_result * result)

These routines compute the quantity exp(x) − 1 using an algorithm that is accurate for small x.

Function double gsl sf exprel (double x)

Function int gsl sf exprel e (double x, gsl_sf_result * result)

These routines compute the quantity (exp(x) − 1)/x using an algorithm that is accurate for small x. For small x the algorithm is based on the expansion (exp(x)−1)/x = 1 + x/2 + x^{2}/(2 ∗ 3) + x^{3}/(2 ∗ 3 ∗ 4) + . . ..

Function double gsl sf exprel 2 (double x)

Function int gsl sf exprel 2 e (double x, gsl_sf_result * result)

These routines compute the quantity 2(exp(x) − 1 − x)/x^{2} using an algorithm that is accurate for small x. For small x the algorithm is based on the expansion 2(exp(x) − 1 − x)/x^{2} = 1 + x/3 + x^{2}/(3 ∗ 4) + x^{3}/(3 ∗ 4 ∗ 5) + . . ..

Function double gsl sf exprel n (int n, double x)

Function int gsl sf exprel n e (int n, double x, gsl_sf_result * result)

These routines compute the N-relative exponential, which is the n-th generalization of the functions gsl_sf_exprel and gsl_sf_exprel2. The N-relative exponential is given by,

exprel_{N}(x) = N!/x^{N} (exp(x) − ∑_{k=0}^{N−1} x^{k}/k!)
= 1 + x/(N + 1) + x^{2}/((N + 1)(N + 2)) + …
= _{1}F_{1}(1, 1 + N, x)

7.16.3 Exponentiation With Error Estimate

Function int gsl sf exp err e (double x, double dx, gsl_sf_result * result)

This function exponentiates x with an associated absolute error dx.

Function int gsl sf exp err e10 e (double x, double dx, gsl_sf_result_e10

* result) This functions exponentiate a quantity x with an associated absolute error dx using the gsl_sf_result_e10 type to return a result with extended range.

Function int gsl sf exp mult err e (double x, double dx, double y, double
    dy, gsl_sf_result * result)

This routine computes the product y exp(x) for the quantities x, y with associated absolute errors dx, dy.

Function int gsl sf exp mult err e10 e (double x, double dx, double y,

    double dy, gsl_sf_result_e10 * result)

This routine computes the product y exp(x) for the quantities x, y with associated absolute errors dx, dy using the gsl_sf_result_e10 type to return a result with extended range.

7.17 Exponential Integrals

Information on the exponential integrals can be found in Abramowitz & Stegun, Chapter 5. These functions are declared in the header file ‘gsl_sf_expint.h’.

7.17.1 Exponential Integral

Function double gsl sf expint E1 (double x)

Function int gsl sf expint E1 e (double x, gsl_sf_result * result)

These routines compute the exponential integral E_{1}(x),

E_{1}(x) := Re ∫_{1}^{∞} dt exp(−xt)/t.

Function double gsl sf expint E2 (double x)

Function int gsl sf expint E2 e (double x, gsl_sf_result * result)

These routines compute the second-order exponential integral E_{2}(x),

E_{2}(x) := Re ∫_{1}^{∞} dt exp(−xt)/t^{2}.

7.17.2 Ei(x)

Function double gsl sf expint Ei (double x)

Function int gsl sf expint Ei e (double x, gsl_sf_result * result)

These routines compute the exponential integral E_{i}(x),

Ei(x) := P V (∫_{−x}^{∞} dt exp(−t)/t)

where P V denotes the principal value of the integral.

7.17.3 Hyperbolic Integrals

Function double gsl sf Shi (double x)

Function int gsl sf Shi e (double x, gsl_sf_result * result)

These routines compute the integral Shi(x) = ∫_{0}^{x} dt sinh(t)/t.

Function double gsl sf Chi (double x)

Function int gsl sf Chi e (double x, gsl_sf_result * result)

These routines compute the integral Chi(x) := Re[γ_{E} + log(x) + ∫_{0}^{x} dt(cosh[t] − 1)/t], where γ_{E} is the Euler constant (available as the macro M_EULER).

7.17.4 Ei 3(x)

Function double gsl sf expint 3 (double x)

Function int gsl sf expint 3 e (double x, gsl_sf_result * result)

These routines compute the exponential integral Ei_{3}(x) = ∫_{0}^{x} dt exp(−t^{3}) for x ≥ 0.

7.17.5 Trigonometric Integrals

Function double gsl sf Si (const double x) Function int gsl sf Si e (double x, gsl_sf_result * result)

These routines compute the Sine integral Si(x) = ∫_{0}^{x} dt sin(t)/t.

Function double gsl sf Ci (const double x) Function int gsl sf Ci e (double x, gsl_sf_result * result)

These routines compute the Cosine integral Ci(x) = − ∫_{x}^{∞} dt cos(t)/t for x > 0.

7.17.6 Arctangent Integral

Function double gsl sf atanint (double x)

Function int gsl sf atanint e (double x, gsl_sf_result * result)

These routines compute the Arctangent integral AtanInt(x) = ∫_{0}^{x} dt arctan(t)/t.

7.18 Fermi-Dirac Function

The functions described in this section are declared in the header file

‘gsl_sf_fermi_dirac.h’.

7.18.1 Complete Fermi-Dirac Integrals

The complete Fermi-Dirac integral F_{j}(x) is given by,

F_{j}(x) := \frac{1}{Γ(j + 1)} ∫_{0}^{∞} \frac{t^{j}}{(exp(t − x) + 1)}

Function double gsl sf fermi dirac m1 (double x)

Function int gsl sf fermi dirac m1 e (double x, gsl_sf_result * result)

These routines compute the complete Fermi-Dirac integral with an index of −1. This integral is given by F_{−1}(x) = e^{x}/(1 + e^{x}).

Function double gsl sf fermi dirac 0 (double x)

Function int gsl sf fermi dirac 0 e (double x, gsl_sf_result * result)

These routines compute the complete Fermi-Dirac integral with an index of 0. This integral is given by F_{0}(x) = ln(1 + e^{x}).

Function double gsl sf fermi dirac 1 (double x)

Function int gsl sf fermi dirac 1 e (double x, gsl_sf_result * result)

These routines compute the complete Fermi-Dirac integral with an index of 1, F_{1}(x) = ∫_{0}^{∞} dt(t/(exp(t − x) + 1)).

Function double gsl sf fermi dirac 2 (double x)

Function int gsl sf fermi dirac 2 e (double x, gsl_sf_result * result)

These routines compute the complete Fermi-Dirac integral with an index of 2, F_{2}(x) = (1/2) ∫_{0}^{∞} dt(t^{2}/(exp(t − x) + 1)).

Function double gsl sf fermi dirac int (int j, double x)

Function int gsl sf fermi dirac int e (int j, double x, gsl_sf_result *

result) These routines compute the complete Fermi-Dirac integral with an integer index of j, F_{j}(x) = (1/Γ(j + 1)) ∫_{0}^{∞} dt(t^{j}/(exp(t − x) + 1)).

Function double gsl sf fermi dirac mhalf (double x)

Function int gsl sf fermi dirac mhalf e (double x, gsl_sf_result * result)

These routines compute the complete Fermi-Dirac integral F_{−1/2}(x).

Function double gsl sf fermi dirac half (double x)

Function int gsl sf fermi dirac half e (double x, gsl_sf_result * result)

These routines compute the complete Fermi-Dirac integral F_{1/2}(x).

Function double gsl sf fermi dirac 3half (double x)

Function int gsl sf fermi dirac 3half e (double x, gsl_sf_result * result)

These routines compute the complete Fermi-Dirac integral F_{3/2}(x).

7.18.2 Incomplete Fermi-Dirac Integrals

The incomplete Fermi-Dirac integral F_{j}(x, b) is given by,

F_{j}(x, b) := \frac{1}{Γ(j + 1)} ∫_{b}^{∞} \frac{t^{j}}{(exp(t − x) + 1)}

Function double gsl sf fermi dirac inc 0 (double x, double b) Function int gsl sf fermi dirac inc 0 e (double x, double b,

    gsl_sf_result * result)

These routines compute the incomplete Fermi-Dirac integral with an index of zero, F_{0}(x, b) = ln(1 + e^{b−x}) − (b − x).

7.19 Gamma Function

The Gamma function is defined by the following integral,

Γ(x) = ∫_{0}^{t} dt t^{x−1} exp(−t)

Further information on the Gamma function can be found in Abramowitz & Stegun, Chapter 6. The functions described in this section are declared in the header file ‘gsl_sf_gamma.h’.

Function double gsl sf gamma (double x)

Function int gsl sf gamma e (double x, gsl_sf_result * result)

These routines compute the Gamma function Γ(x), subject to x not being a negative integer. The function is computed using the real Lanczos method. The maximum value of x such that Γ(x) is not considered an overflow is given by the macro GSL_

SF_GAMMA_XMAX and is 171.0.

Function double gsl sf lngamma (double x)

Function int gsl sf lngamma e (double x, gsl_sf_result * result)

These routines compute the logarithm of the Gamma function, log(Γ(x)), subject to x not a being negative integer. For x < 0 the real part of log(Γ(x)) is returned, which is equivalent to log(|Γ(x)|). The function is computed using the real Lanczos method.

Function int gsl sf lngamma sgn e (double x, gsl_sf_result * result lg,

    double * sgn)

This routine computes the sign of the gamma function and the logarithm its magnitude, subject to x not being a negative integer. The function is computed using the real Lanczos method. The value of the gamma function can be reconstructed using the relation Γ(x) = sgn ∗ exp(resultlg).

Function double gsl sf gammastar (double x)

Function int gsl sf gammastar e (double x, gsl_sf_result * result)

These routines compute the regulated Gamma Function Γ^{∗}(x) for x > 0. The regulated gamma function is given by,

Γ^{∗}(x) = Γ(x)/(\sqrt{2π}x^{(x−1/2)} exp(−x))

= (1 + \frac{1}{12x} + ...) for x → ∞

and is a useful suggestion of Temme.

Function double gsl sf gammainv (double x)

Function int gsl sf gammainv e (double x, gsl_sf_result * result)

These routines compute the reciprocal of the gamma function, 1/Γ(x) using the real Lanczos method.

Function int gsl sf lngamma complex e (double zr, double zi,

    gsl_sf_result * lnr, gsl_sf_result * arg)

This routine computes log(Γ(z)) for complex z = z_{r} +iz_{i} and z not a negative integer, using the complex Lanczos method. The returned parameters are lnr = log |Γ(z)| and arg = arg(Γ(z)) in (−π, π]. Note that the phase part (arg) is not well-determined when |z| is very large, due to inevitable roundoff in restricting to (−π, π]. This will result in a GSL_ELOSS error when it occurs. The absolute value part (lnr), however, never suffers from loss of precision.

Function double gsl sf taylorcoeff (int n, double x)

Function int gsl sf taylorcoeff e (int n, double x, gsl_sf_result * result)

These routines compute the Taylor coefficient x^{n}/n! for x ≥ 0, n ≥ 0.

Function double gsl sf fact (unsigned int n)

Function int gsl sf fact e (unsigned int n, gsl_sf_result * result)

These routines compute the factorial n!. The factorial is related to the Gamma function by n! = Γ(n + 1).

Function double gsl sf doublefact (unsigned int n)

Function int gsl sf doublefact e (unsigned int n, gsl_sf_result * result)

These routines compute the double factorial n!! = n(n − 2)(n − 4) . . ..

Function double gsl sf lnfact (unsigned int n)

Function int gsl sf lnfact e (unsigned int n, gsl_sf_result * result)

These routines compute the logarithm of the factorial of n, log(n!). The algorithm is faster than computing ln(Γ(n + 1)) via gsl_sf_lngamma for n < 170, but defers for larger n.

Function double gsl sf lndoublefact (unsigned int n)

Function int gsl sf lndoublefact e (unsigned int n, gsl_sf_result * result)

These routines compute the logarithm of the double factorial of n, log(n!!).

Function double gsl sf choose (unsigned int n, unsigned int m)
Function int gsl sf choose e (unsigned int n, unsigned int m,
    gsl_sf_result * result)

These routines compute the combinatorial factor n choose m = n!/(m!(n − m)!)

Function double gsl sf lnchoose (unsigned int n, unsigned int m)
Function int gsl sf lnchoose e (unsigned int n, unsigned int m,
    gsl_sf_result * result)

These routines compute the logarithm of n choose m. This is equivalent to the sum log(n!) − log(m!) − log((n − m)!).

Function double gsl sf poch (double a, double x)

Function int gsl sf poch e (double a, double x, gsl_sf_result * result)

These routines compute the Pochhammer symbol (a)_{x} := Γ(a + x)/Γ(x), subject to a and a + x not being negative integers. The Pochhammer symbol is also known as the Apell symbol.

Function double gsl sf lnpoch (double a, double x)

Function int gsl sf lnpoch e (double a, double x, gsl_sf_result * result)

These routines compute the logarithm of the Pochhammer symbol, log((a)_{x}) = log(Γ(a + x)/Γ(a)) for a > 0, a + x > 0.

Function int gsl sf lnpoch sgn e (double a, double x, gsl_sf_result *

result, double * sgn) These routines compute the sign of the Pochhammer symbol and the logarithm of its magnitude. The computed parameters are result = log(|(a)_{x}|) and sgn = sgn((a)_{x}) where (a)_{x} := Γ(a + x)/Γ(a), subject to a, a + x not being negative integers.

Function double gsl sf pochrel (double a, double x)

Function int gsl sf pochrel e (double a, double x, gsl_sf_result * result)

These routines compute the relative Pochhammer symbol ((a, x)−1)/x where (a, x) = (a)_{x} := Γ(a + x)/Γ(a).

Function double gsl sf gamma inc Q (double a, double x)

Function int gsl sf gamma inc Q e (double a, double x, gsl_sf_result *

result) These routines compute the normalized incomplete Gamma Function P (a, x) = 1/Γ(a) ∫_{x}^{∞} dt t^{(a−1)} exp(−t) for a > 0, x ≥ 0.

Function double gsl sf gamma inc P (double a, double x)

Function int gsl sf gamma inc P e (double a, double x, gsl_sf_result *

result) These routines compute the complementary normalized incomplete Gamma Function P (a, x) = 1/Γ(a) ∫_{0}^{x} dt t^{(a−1)} exp(−t) for a > 0, x ≥ 0.

Function double gsl sf beta (double a, double b)

Function int gsl sf beta e (double a, double b, gsl_sf_result * result)

These routines compute the Beta Function, B(a, b) = Γ(a)Γ(b)/Γ(a + b) for a > 0, b > 0.

Function double gsl sf lnbeta (double a, double b)

Function int gsl sf lnbeta e (double a, double b, gsl_sf_result * result)

These routines compute the logarithm of the Beta Function, log(B(a, b)) for a > 0, b > 0.

Function double gsl sf beta inc (double a, double b, double x)

Function int gsl sf beta inc e (double a, double b, double x,

    gsl_sf_result * result)

These routines compute the normalize incomplete Beta function B_{x}(a, b)/B(a, b) for a > 0, b > 0, and 0 ≤ x ≤ 1.

7.20 Gegenbauer Functions

The Gegenbauer polynomials are defined in Abramowitz & Stegun, Chapter 22, where they are known as Ultraspherical polynomials. The functions described in this section are declared in the header file ‘gsl_sf_gegenbauer.h’.

Function double gsl sf gegenpoly 1 (double lambda, double x) Function double gsl sf gegenpoly 2 (double lambda, double x) Function double gsl sf gegenpoly 3 (double lambda, double x) Function int gsl sf gegenpoly 1 e (double lambda, double x,

         gsl_sf_result * result)

Function int gsl sf gegenpoly 2 e (double lambda, double x,

         gsl_sf_result * result)

Function int gsl sf gegenpoly 3 e (double lambda, double x,

         gsl_sf_result * result)

These functions evaluate the Gegenbauer polynomials C_{n}^{(λ)}(x) using explicit representations for n = 1, 2, 3.

Function double gsl sf gegenpoly n (int n, double lambda, double x) Function int gsl sf gegenpoly n e (int n, double lambda, double x,

         gsl_sf_result * result)

These functions evaluate the Gegenbauer polynomial C_{n}^{(λ)}(x) for a specific value of n, lambda, x subject to λ > −1/2, n ≥ 0.

Function int gsl sf gegenpoly array (int nmax, double lambda, double x,

double result array[]) This function computes an array of Gegenbauer polynomials C_{n}^{(λ)}(x) for n = 0, 1, 2, …, nmax, subject to λ > −1/2, nmax ≥ 0.

7.21 Hypergeometric Functions

Hypergeometric functions are described in Abramowitz & Stegun, Chapters 13 and 15. These functions are declared in the header file ‘gsl_sf_hyperg.h’.

Function double gsl sf hyperg 0F1 (double c, double x)

Function int gsl sf hyperg 0F1 e (double c, double x, gsl_sf_result *

result) These routines compute the hypergeometric function _{0}F_{1}(c, x).

Function double gsl sf hyperg 1F1 int (int m, int n, double x) Function int gsl sf hyperg 1F1 int e (int m, int n, double x,

         gsl_sf_result * result)

These routines compute the confluent hypergeometric function _{1}F_{1}(m, n, x) = M(m, n, x) for integer parameters m, n.

Function double gsl sf hyperg 1F1 (double a, double b, double x) Function int gsl sf hyperg 1F1 e (double a, double b, double x,

    gsl_sf_result * result)

These routines compute the confluent hypergeometric function _{1}F_{1}(a, b, x) = M(a, b, x) for general parameters a, b.

Function double gsl sf hyperg U int (int m, int n, double x) Function int gsl sf hyperg U int e (int m, int n, double x,

    gsl_sf_result * result)

These routines compute the confluent hypergeometric function U(m, n, x) for integer parameters m, n.

Function int gsl sf hyperg U int e10 e (int m, int n, double x,

    gsl_sf_result_e10 * result)

This routine computes the confluent hypergeometric function U(m, n, x) for integer parameters m, n using the gsl_sf_result_e10 type to return a result with extended range.

Function double gsl sf hyperg U (double a, double b, double x)

Function int gsl sf hyperg U e (double a, double b, double x)

These routines compute the confluent hypergeometric function U(a, b, x).

Function int gsl sf hyperg U e10 e (double a, double b, double x,

    gsl_sf_result_e10 * result)

This routine computes the confluent hypergeometric function U(a, b, x) using the gsl_sf_result_e10 type to return a result with extended range.

Function double gsl sf hyperg 2F1 (double a, double b, double c, double

x)

Function int gsl sf hyperg 2F1 e (double a, double b, double c, double x,
    gsl_sf_result * result)

These routines compute the Gauss hypergeometric function _{2}F_{1}(a, b, c, x) for |x| < 1.

Function double gsl sf hyperg 2F1 conj (double aR, double aI, double c,

    double x)

Function int gsl sf hyperg 2F1 conj e (double aR, double aI, double c,

    double x, gsl_sf_result * result)

These routines compute the Gauss hypergeometric function_{2}F_{1}(a_{R}+ia_{I}, aR−iaI, c, x) with complex parameters for |x| < 1. exceptions:

Function double gsl sf hyperg 2F1 renorm (double a, double b, double c,

    double x)

Function int gsl sf hyperg 2F1 renorm e (double a, double b, double c,

    double x, gsl_sf_result * result)

These routines compute the renormalized Gauss hypergeometric function 2F_{1}(a, b, c, x)/Γ(c) for |x| < 1.

Function double gsl sf hyperg 2F1 conj renorm (double aR, double aI,

         double c, double x)

Function int gsl sf hyperg 2F1 conj renorm e (double aR, double aI,

         double c, double x, gsl_sf_result * result)

These routines compute the renormalized Gauss hypergeometric function _{2}F_{1}(a_{R} + ia_{I}, a_{R} − ia_{I}, c, x)/Γ(c) for |x| < 1.

Function double gsl sf hyperg 2F0 (double a, double b, double x) Function int gsl sf hyperg 2F0 e (double a, double b, double x,

         gsl_sf_result * result)

These routines compute the hypergeometric function _{2}F_{0}(a, b, x). The series representation is a divergent hypergeometric series. However, for x < 0 we have _{2}F_{0}(a, b, x) = (−1/x)^{a}U(a, 1 + a − b, −1/x)

7.22 Laguerre Functions

The Laguerre polynomials are defined in terms of confluent hypergeometric functions as L_{n}^{a}(x) = ((a + 1)_{n}/n!)_{1}F_{1}(−n, a + 1, x). These functions are declared in the header file

‘gsl_sf_laguerre.h’.

Function double gsl sf laguerre 1 (double a, double x) Function double gsl sf laguerre 2 (double a, double x) Function double gsl sf laguerre 3 (double a, double x)

Function int gsl sf laguerre 1 e (double a, double x, gsl_sf_result *

result)

Function int gsl sf laguerre 2 e (double a, double x, gsl_sf_result *

result)

Function int gsl sf laguerre 3 e (double a, double x, gsl_sf_result *

result) These routines evaluate the generalized Laguerre polynomials L_{1}^{a}(x), L_{2}^{a}(x), L_{3}^{a}(x) using explicit representations.

Function double gsl sf laguerre n (const int n, const double a, const
         double x)
Function int gsl sf laguerre n e (int n, double a, double x, gsl_sf_result

* result) Thse routines evaluate the generalized Laguerre polynomials L_{n}^{a}(x) for a > −1, n ≥ 0.

7.23 Lambert W Functions

Lambert’s W functions, W (x), are defined to be solutions of the equation W (x) exp(W (x)) = x. This function has multiple branches for x < 0; however, it has only two real-valued branches. We define W_{0}(x) to be the principal branch, where W > −1 for x < 0, and W_{−1}(x) to be the other real branch, where W < −1 for x < 0. The Lambert functions are declared in the header file ‘gsl_sf_lambert.h’.

Function double gsl sf lambert W0 (double x)

Function int gsl sf lambert W0 e (double x, gsl_sf_result * result)

These compute the principal branch of the Lambert W function, W_{0}(x).

Function double gsl sf lambert Wm1 (double x)

Function int gsl sf lambert Wm1 e (double x, gsl_sf_result * result)

These compute the secondary real-valued branch of the Lambert W function, W_{−1}(x).

7.24 Legendre Functions and Spherical Harmonics

The Legendre Functions and Legendre Polynomials are described in Abramowitz & Stegun, Chapter 8. These functions are declared in the header file ‘gsl_sf_legendre.h’.

7.24.1 Legendre Polynomials

Function double gsl sf legendre P1 (double x) Function double gsl sf legendre P2 (double x) Function double gsl sf legendre P3 (double x)

Function int gsl sf legendre P1 e (double x, gsl_sf_result * result) Function int gsl sf legendre P2 e (double x, gsl_sf_result * result) Function int gsl sf legendre P3 e (double x, gsl_sf_result * result)

These functions evaluate the Legendre polynomials P_{l}(x) using explicit representations for l = 1, 2, 3.

Function double gsl sf legendre Pl (int l, double x)

Function int gsl sf legendre Pl e (int l, double x, gsl_sf_result * result)

These functions evaluate the Legendre polynomial P_{l}(x) for a specific value of l, x subject to l ≥ 0, |x| ≤ 1

Function int gsl sf legendre Pl array (int lmax, double x, double

result array[]) This function computes an array of Legendre polynomials P_{l}(x) for l = 0, …, lmax, |x| ≤ 1

Function double gsl sf legendre Q0 (double x)

Function int gsl sf legendre Q0 e (double x, gsl_sf_result * result)

These routines compute the Legendre function Q_{0}(x) for x > −1, x ≠ 1.

Function double gsl sf legendre Q1 (double x)

Function int gsl sf legendre Q1 e (double x, gsl_sf_result * result)

These routines compute the Legendre function Q_{1}(x) for x > −1, x ≠ 1.

Function double gsl sf legendre Ql (int l, double x)

Function int gsl sf legendre Ql e (int l, double x, gsl_sf_result * result)

These routines compute the Legendre function Q_{l}(x) for x > −1, x ≠ 1 and l ≥ 0.

7.24.2 Associated Legendre Polynomials and Spherical Harmonics

The following functions compute the associated Legendre Polynomials P_{l}^{m}(x). Note that this function grows combinatorially with l and can overflow for l larger than about 150. There is no trouble for small m, but overflow occurs when m and l are both large. Rather than allow overflows, these functions refuse to calculate P_{l}^{m}(x) and return GSL_EOVRFLW when they can sense that l and m are too big.

If you want to calculate a spherical harmonic, then do not use these functions. Instead use gsl_sf_legendre_sphPlm() below, which uses a similar recursion, but with the normalized functions.

Function double gsl sf legendre Plm (int l, int m, double x)

Function int gsl sf legendre Plm e (int l, int m, double x, gsl_sf_result

* result) These routines compute the associated Legendre polynomial P_{l}^{m}(x) for m ≥ 0, l ≥ m, |x| ≤ 1.

Function int gsl sf legendre Plm array (int lmax, int m, double x,

double result array[]) This function computes an array of Legendre polynomials P_{l}^{m}(x) for m ≥ 0, l = |m|, …, lmax, |x| ≤ 1.

Function double gsl sf legendre sphPlm (int l, int m, double x) Function int gsl sf legendre sphPlm e (int l, int m, double x,

    gsl_sf_result * result)

These routines compute the normalized associated Legendre polynomial \sqrt{(2l + 1)/(4π)}\sqrt{(l − m)!/(l + m)!}P_{l}^{m}(x) suitable for use in spherical harmonics. The parameters must satisfy m ≥ 0, l ≥ m, |x| ≤ 1. Theses routines avoid the overflows that occur for the standard normalization of P_{l}^{m}(x).

Function int gsl sf legendre sphPlm array (int lmax, int m, double x,

double result array[]) This function computes an array of normalized associated Legendre functions \sqrt{(2l + 1)/(4π)}\sqrt{(l − m)!/(l + m)!}P_{l}^{m}(x) for m ≥ 0, l = |m|, …, lmax, |x| ≤ 1

Function int gsl sf legendre array size (const int lmax, const int m)

This functions returns the size of result array[] needed for the array versions of P_{l}^{m}(x), lmax - m + 1.

7.24.3 Conical Functions

The Conical Functions P_{−(1/2)+iλ}^{μ}(x), Q_{−(1/2)+iλ}^{μ} are described in Abramowitz & Stegun, Section 8.12.

Function double gsl sf conicalP half (double lambda, double x)

Function int gsl sf conicalP half e (double lambda, double x,

         gsl_sf_result * result)

These routines compute the irregular Spherical Conical Function P_{−1/2+iλ}^{1/2}(x) for x > −1.

Function double gsl sf conicalP mhalf (double lambda, double x) Function int gsl sf conicalP mhalf e (double lambda, double x,

         gsl_sf_result * result)

These routines compute the regular Spherical Conical Function P_{−1/2+iλ}^{−1/2}(x) for x > −1.

Function double gsl sf conicalP 0 (double lambda, double x) Function int gsl sf conicalP 0 e (double lambda, double x, gsl_sf_result

* result) These routines compute the conical function P_{−1/2+iλ}^{0}(x) for x > −1.

Function double gsl sf conicalP 1 (double lambda, double x) Function int gsl sf conicalP 1 e (double lambda, double x, gsl_sf_result

* result) These routines compute the conical function P_{−1/2+iλ}^{1}(x) for x > −1.

Function double gsl sf conicalP sph reg (int l, double lambda, double x) Function int gsl sf conicalP sph reg e (int l, double lambda, double x,

         gsl_sf_result * result)

These routines compute the Regular Spherical Conical Function P_{−1/2+iλ}^{−1/2−l}(x) for x > −1, l ≥ −1.

Function double gsl sf conicalP cyl reg (int m, double lambda, double x) Function int gsl sf conicalP cyl reg e (int m, double lambda, double x,

         gsl_sf_result * result)

These routines compute the Regular Cylindrical Conical Function P_{−1/2+iλ}^{−m}(x) for x > −1, m ≥ −1.

7.24.4 Radial Functions for Hyperbolic Space

The following spherical functions are specializations of Legendre functions which give the regular eigenfunctions of the Laplacian on a 3-dimensional hyperbolic space H3d. Of particular interest is the flat limit, λ → ∞, η → 0, λη fixed.

Function double gsl sf legendre H3d 0 (double lambda, double eta) Function int gsl sf legendre H3d 0 e (double lambda, double eta,

         gsl_sf_result * result)

These routines compute the zeroth radial eigenfunction of the Laplacian on the 3dimensional hyperbolic space, L_{0}^{H3d}(λ, η) := sin(λη)/(λ sinh(η)) for η ≥ 0. In the flat limit this takes the form L_{0}^{H3d}(λ, η) = j_{0}(λη)

Function double gsl sf legendre H3d 1 (double lambda, double eta) Function int gsl sf legendre H3d 1 e (double lambda, double eta,

         gsl_sf_result * result)

These routines compute the first radial eigenfunction of the Laplacian on the 3dimensional hyperbolic space, L_{1}^{H3d}(λ, η) := 1/\sqrt{λ^{2} + 1} sin(λη)/(λ sinh(η))(coth(η) − λ cot(λη)) for η ≥ 0. In the flat limit this takes the form L_{1}^{H3d}(λ, η) = j_{1}(λη).

Function double gsl sf legendre H3d (int l, double lambda, double eta) Function int gsl sf legendre H3d e (int l, double lambda, double eta,

         gsl_sf_result * result)

These routines compute the l’th radial eigenfunction of the Laplacian on the 3dimensional hyperbolic space η ≥ 0, l ≥ 0. In the flat limit this takes the form L_{l}^{H3d}(λ, η) = j_{l}(λη).

Function int gsl sf legendre H3d array (int lmax, double lambda, double

eta, double result array[]) This function computes an array of radial eigenfunctions L_{l}^{H3d}(λ, η) for 0 ≤ l ≤ lmax.

7.25 Logarithm and Related Functions

Information on the properties of the Logarithm function can be found in Abramowitz & Stegun, Chapter 4. The functions described in this section are declared in the header file

‘gsl_sf_log.h’.

Function double gsl sf log (double x)

Function int gsl sf log e (double x, gsl_sf_result * result)

These routines compute the logarithm of x, log(x), for x > 0.

Function double gsl sf log abs (double x)

Function int gsl sf log abs e (double x, gsl_sf_result * result)

These routines compute the logarithm of the magnitude of x, log(|x|), for x ≠ 0.

Function int gsl sf complex log e (double zr, double zi, gsl_sf_result *
         lnr, gsl_sf_result * theta)

This routine computes the complex logarithm of z = z_{r} +iz_{i}. The results are returned as lnr, theta such that exp(lnr + iθ) = z_{r} + iz_{i}, where θ lies in the range [−π, π].

Function double gsl sf log 1plusx (double x)

Function int gsl sf log 1plusx e (double x, gsl_sf_result * result)

These routines compute log(1 + x) for x > −1 using an algorithm that is accurate for small x.

Function double gsl sf log 1plusx mx (double x)

Function int gsl sf log 1plusx mx e (double x, gsl_sf_result * result)

These routines compute log(1 + x) − x for x > −1 using an algorithm that is accurate for small x.

7.26 Power Function

The following functions are equivalent to the function gsl_pow_int (see Section 4.4 [Small integer powers], page 16) with an error estimate. These functions are declared in the

header file ‘gsl_sf_pow_int.h’.

Function double gsl sf pow int (double x, int n)

Function int gsl sf pow int e (double x, int n, gsl_sf_result * result)

These routines compute the power x^{n} for integer n. The power is computed using the minimum number of multiplications. For example, x^{8} is computed as ((x^{2})^{2})^{2}, requiring only 3 multiplications. For reasons of efficiency, these functions do not check for overflow or underflow conditions.

#include <gsl/gsl_sf_pow_int.h>
/* compute 3.0**12 */
double y = gsl_sf_pow_int(3.0, 12);

7.27 Psi (Digamma) Function

The polygamma functions of order m defined by ψ^{(m)}(x) = (d/dx)^{m}ψ(x) = (d/dx)^{m+1} log(Γ(x)), where ψ(x) = Γ′(x)/Γ(x) is known as the digamma function. These functions are declared in the header file ‘gsl_sf_psi.h’.

7.27.1 Digamma Function

Function double gsl sf psi int (int n)

Function int gsl sf psi int e (int n, gsl_sf_result * result)

These routines compute the digamma function ψ(n) for positive integer n. The digamma function is also called the Psi function.

Function double gsl sf psi (double x)

Function int gsl sf psi e (double x, gsl_sf_result * result)

These routines compute the digamma function ψ(x) for general x, x ≠ 0.

Function double gsl sf psi 1piy (double y)

Function int gsl sf psi 1piy e (double y, gsl_sf_result * result)

These routines compute the real part of the digamma function on the line 1 + iy, Re[ψ(1 + iy)].

7.27.2 Trigamma Function

Function double gsl sf psi 1 int (int n)

Function int gsl sf psi 1 int e (int n, gsl_sf_result * result)

These routines compute the Trigamma function ψ′(n) for positive integer n.

7.27.3 Polygamma Function

Function double gsl sf psi n (int m, double x)

Function int gsl sf psi n e (int m, double x, gsl_sf_result * result)

These routines compute the polygamma function ψ^{(m)}(x) for c ≥ 0, x > 0.

7.28 Synchrotron Functions

The functions described in this section are declared in the header file

‘gsl_sf_synchroton.h’.

Function double gsl sf synchrotron 1 (double x)

Function int gsl sf synchrotron 1 e (double x, gsl_sf_result * result)

These routines compute the first synchrotron function x ∫_{x}^{∞} dtK_{5/3}(t) for x ≥ 0.

Function double gsl sf synchrotron 2 (double x)

Function int gsl sf synchrotron 2 e (double x, gsl_sf_result * result)

These routines compute the second synchrotron function xK_{2/3}(x) for x ≥ 0.

7.29 Transport Functions

The transport functions J(n, x) are defined by the integral representations J(n, x) := ∫_{0}^{x} dt t^{n}e^{t}/(e^{t} − 1)^{2}. They are declared in the header file ‘gsl_sf_transport.h’.

Function double gsl sf transport 2 (double x)

Function int gsl sf transport 2 e (double x, gsl_sf_result * result)

These routines compute the transport function J(2, x).

Function double gsl sf transport 3 (double x)

Function int gsl sf transport 3 e (double x, gsl_sf_result * result)

These routines compute the transport function J(3, x).

Function double gsl sf transport 4 (double x)

Function int gsl sf transport 4 e (double x, gsl_sf_result * result)

These routines compute the transport function J(4, x).

Function double gsl sf transport 5 (double x)

Function int gsl sf transport 5 e (double x, gsl_sf_result * result)

These routines compute the transport function J(5, x).

7.30 Trigonometric Functions

The library includes its own trigonometric functions in order to provide consistency across platforms and reliable error estimates. These functions are declared in the header

file ‘gsl_sf_trig.h’.

7.30.1 Circular Trigonometric Functions

Function double gsl sf sin (double x)

Function int gsl sf sin e (double x, gsl_sf_result * result)

These routines compute the sine function sin(x).

Function double gsl sf cos (double x)

Function int gsl sf cos e (double x, gsl_sf_result * result)

These routines compute the cosine function cos(x).

Function double gsl sf hypot (double x, double y)

Function int gsl sf hypot e (double x, double y, gsl_sf_result * result)

These routines compute the hypotenuse function \sqrt{x^{2} + y^{2}} avoiding overflow and underflow.

Function double gsl sf sinc (double x)

Function int gsl sf sinc e (double x, gsl_sf_result * result)

These routines compute sinc(x) = sin(πx)/(πx) for any value of x.

7.30.2 Trigonometric Functions for Complex Arguments

Function int gsl sf complex sin e (double zr, double zi, gsl_sf_result *
         szr, gsl_sf_result * szi)

This function computes the complex sine, sin(z_{r} + iz_{i}) storing the real and imaginary parts in szr, szi.

Function int gsl sf complex cos e (double zr, double zi, gsl_sf_result *
         czr, gsl_sf_result * czi)

This function computes the complex cosine, cos(z_{r}+iz_{i}) storing the real and imaginary parts in szr, szi.

Function int gsl sf complex logsin e (double zr, double zi, gsl_sf_result

         * lszr, gsl_sf_result * lszi)

This function computes the logarithm of the complex sine, log(sin(z_{r} + iz_{i})) storing the real and imaginary parts in szr, szi.

7.30.3 Hyperbolic Trigonometric Functions

Function double gsl sf lnsinh (double x)

Function int gsl sf lnsinh e (double x, gsl_sf_result * result)

These routines compute log(sinh(x)) for x > 0.

Function double gsl sf lncosh (double x)

Function int gsl sf lncosh e (double x, gsl_sf_result * result)

These routines compute log(cosh(x)) for any x.

7.30.4 Conversion Functions

Function int gsl sf polar to rect (double r, double theta, gsl_sf_result *
    x, gsl_sf_result * y);

This function converts the polar coordinates (r,theta) to rectilinear coordinates (x,y), x = r cos(θ), y = r sin(θ).

Function int gsl sf rect to polar (double x, double y, gsl_sf_result * r,
    gsl_sf_result * theta)

This function converts the rectilinear coordinates (x,y) to polar coordinates (r,theta), such that x = r cos(θ), y = r sin(θ). The argument theta lies in the range [−π, π].

7.30.5 Restriction Functions

Function double gsl sf angle restrict symm (double theta) Function int gsl sf angle restrict symm e (double * theta)

These routines force the angle theta to lie in the range (−π, π].

Function double gsl sf angle restrict pos (double theta) Function int gsl sf angle restrict pos e (double * theta)

These routines force the angle theta to lie in the range [0, 2π).

7.30.6 Trigonometric Functions With Error Estimates

Function double gsl sf sin err (double x, double dx)

Function int gsl sf sin err e (double x, double dx, gsl_sf_result * result)

These routines compute the sine of an angle x with an associated absolute error dx, sin(x ± dx).

Function double gsl sf cos err (double x, double dx)

Function int gsl sf cos err e (double x, double dx, gsl_sf_result * result)

These routines compute the cosine of an angle x with an associated absolute error dx, cos(x ± dx).

7.31 Zeta Functions

The Riemann zeta function is defined in Abramowitz & Stegun, Section 23.2. The functions described in this section are declared in the header file ‘gsl_sf_zeta.h’.

7.31.1 Riemann Zeta Function

The Riemann zeta function is defined by the infinite sum ζ(s) = ∑_{k=1}^{∞} k^{−s}.

Function double gsl sf zeta int (int n)

Function int gsl sf zeta int e (int n, gsl_sf_result * result)

These routines compute the Riemann zeta function ζ(n) for integer n, n ≠ 1.

Function double gsl sf zeta (double s)

Function int gsl sf zeta e (double s, gsl_sf_result * result)

These routines compute the Riemann zeta function ζ(s) for arbitrary s, s ≠ 1.

7.31.2 Hurwitz Zeta Function

The Hurwitz zeta function is defined by ζ(s, q) = ∑_{0}^{∞}(k + q)^{−s}.

Function double gsl sf hzeta (double s, double q)

Function int gsl sf hzeta e (double s, double q, gsl_sf_result * result)

These routines compute the Hurwitz zeta function ζ(s, q) for s > 1, q > 0.

7.31.3 Eta Function

The eta function is defined by η(s) = (1 − 2^{1−s})ζ(s).

Function double gsl sf eta int (int n)

Function int gsl sf eta int e (int n, gsl_sf_result * result)

These routines compute the eta function η(n) for integer n.

Function double gsl sf eta (double s)

Function int gsl sf eta e (double s, gsl_sf_result * result)

These routines compute the eta function η(s) for arbitrary s.

7.32 Examples

The following example demonstrates the use of the error handling form of the special functions, in this case to compute the Bessel function J_{0}(5.0),

#include <stdio.h>
#include <gsl/gsl_sf_bessel.h>
int
main (void)
{
  double x = 5.0;
  gsl_sf_result result;
  double expected = -0.17759677131433830434739701;
  int status = gsl_sf_bessel_J0_e (x, &result);
  printf("status  = %s\n", gsl_strerror(status));
  printf("J0(5.0) = %.18f\n"
         "      +/- % .18f\n",
         result.val, result.err);
  printf("exact   = %.18f\n", expected);
  return status;
}

Here are the results of running the program,

$ ./a.out
status  = success
J0(5.0) = -0.177596771314338292
      +/-  0.000000000000000193
exact   = -0.177596771314338292

The next program computes the same quantity using the natural form of the function. In this case the error term result.err and return status are not accessible.

#include <stdio.h>
#include <gsl/gsl_sf_bessel.h>
int
main (void)
{
  double x = 5.0;
  double expected = -0.17759677131433830434739701;
  double y = gsl_sf_bessel_J0 (x);
  printf("J0(5.0) = %.18f\n", y);
  printf("exact   = %.18f\n", expected);
  return 0;
}

The results of the function are the same, Chapter 7: Special Functions 70

 $ ./a.out
 J0(5.0) = -0.177596771314338292
 exact   = -0.177596771314338292

7.33 References and Further Reading

The library follows the conventions of Abramowitz & Stegun where possible,

Abramowitz & Stegun (eds.), Handbook of Mathematical Functions The following papers contain information on the algorithms used to compute the special functions,

MISCFUN: A software package to compute uncommon special functions. ACM Trans. Math. Soft., vol. 22, 1996, 288-301 G.N. Watson, A Treatise on the Theory of Bessel Functions, 2nd Edition (Cambridge University Press, 1944). G. Nemeth, Mathematical Approximations of Special Functions, Nova Science Publishers, ISBN 1-56072-052-2 B.C. Carlson, Special Functions of Applied Mathematics (1977) W.J. Thompson, Atlas for Computing Mathematical Functions, John Wiley & Sons, New York (1997). Y.Y. Luke, Algorithms for the Computation of Mathematical Functions, Academic Press, New York (1977).

8 Vectors and Matrices

The functions described in this chapter provide a simple vector and matrix interface to ordinary C arrays. The memory management of these arrays is implemented using a single underlying type, known as a block. By writing your functions in terms of vectors and matrices you can pass a single structure containing both data and dimensions as an argument without needing additional function parameters. The structures are compatible with the vector and matrix formats used by blas routines.

8.1 Data types

All the functions are available for each of the standard data-types. The versions for double have the prefix gsl_block, gsl_vector and gsl_matrix. Similarly the versions for single-precision float arrays have the prefix gsl_block_float, gsl_vector_float and gsl_matrix_float. The full list of available types is given below,

gsl_block                       double
gsl_block_float                 float
gsl_block_long_double           long double
gsl_block_int                   int
gsl_block_uint                  unsigned int
gsl_block_long                  long
gsl_block_ulong                 unsigned long
gsl_block_short                 short
gsl_block_ushort                unsigned short
gsl_block_char                  char
gsl_block_uchar                 unsigned char
gsl_block_complex               complex double
gsl_block_complex_float         complex float
gsl_block_complex_long_double   complex long double

Corresponding types exist for the gsl_vector and gsl_matrix functions.

8.2 Blocks

For consistency all memory is allocated through a gsl_block structure. The structure contains two components, the size of an area of memory and a pointer to the memory. The gsl_block structure looks like this,

typedef struct
{
  size_t size;
  double * data;
} gsl_block;

Vectors and matrices are made by slicing an underlying block. A slice is a set of elements formed from an initial offset and a combination of indices and step-sizes. In the case of a matrix the step-size for the column index represents the row-length. The step-size for a vector is known as the stride.

The functions for allocating and deallocating blocks are defined in ‘gsl_block.h’

8.2.1 Block allocation

The functions for allocating memory to a block follow the style of malloc and free. In addition they also perform their own error checking. If there is insufficient memory available to allocate a block then the functions call the GSL error handler (with an error number of GSL_ENOMEM) in addition to returning a null pointer. Thus if you use the library error handler to abort your program then it isn’t necessary to check every alloc.

Function gsl_block * gsl block alloc (size_t n)

This function allocates memory for a block of n double-precision elements, returning a pointer to the block struct. The block is not initialized and so the values of its elements are undefined. Use the function gsl_block_calloc if you want to ensure that all the elements are initialized to zero. A null pointer is returned if insufficient memory is available to create the block.

Function gsl_block * gsl block calloc (size_t n)

This function allocates memory for a block and initializes all the elements of the block to zero.

Function void gsl block free (gsl_block * b)

This function frees the memory used by a block b previously allocated with gsl_

block_alloc or gsl_block_calloc.

8.2.2 Reading and writing blocks

The library provides functions for reading and writing blocks to a file as binary data or formatted text.

Function int gsl block fwrite (FILE * stream, const gsl_block * b)

This function writes the elements of the block b to the stream stream in binary format. The return value is 0 for success and GSL_EFAILED if there was a problem writing to the file. Since the data is written in the native binary format it may not be portable between different architectures.

Function int gsl block fread (FILE * stream, gsl_block * b)

This function reads into the block b from the open stream stream in binary format. The block b must be preallocated with the correct length since the function uses the size of b to determine how many bytes to read. The return value is 0 for success and GSL_EFAILED if there was a problem reading from the file. The data is assumed to have been written in the native binary format on the same architecture.

Function int gsl block fprintf (FILE * stream, const gsl_block * b, const

char * format) This function writes the elements of the block b line-by-line to the stream stream using the format specifier format, which should be one of the %g, %e or %f formats for floating point numbers and %d for integers. The function returns 0 for success and GSL_EFAILED if there was a problem writing to the file.

Function int gsl block fscanf (FILE * stream, gsl_block * b)

This function reads formatted data from the stream stream into the block b. The block b must be preallocated with the correct length since the function uses the size of b to determine how many numbers to read. The function returns 0 for success and GSL_EFAILED if there was a problem reading from the file.

8.2.3 Example programs for blocks

The following program shows how to allocate a block,

#include <stdio.h>
#include <gsl/gsl_block.h>
int
main (void)
{
  gsl_block * b = gsl_block_alloc (100);
  printf("length of block = %u\n", b->size);
  printf("block data address = %#x\n", b->data);
  gsl_block_free (b);
  return 0;
}

Here is the output from the program,

length of block = 100
block data address = 0x804b0d8

8.3 Vectors

Vectors are defined by a gsl_vector structure which describes a slice of a block. Different vectors can be created which point to the same block. A vector slice is a set of equally-spaced elements of an area of memory.

The gsl_vector structure contains five components, the size, the stride, a pointer to the memory where the elements are stored, data, a pointer to the block owned by the vector, block, if any, and an ownership flag, owner. The structure is very simple and looks like this,

typedef struct
{
  size_t size;
  size_t stride;
  double * data;
  gsl_block * block;
  int owner;
} gsl_vector;

The size is simply the number of vector elements. The range of valid indices runs from 0 to size-1. The stride is the step-size from one element to the next in physical memory, measured in units of the appropriate datatype. The pointer data gives the location of the first element of the vector in memory. The pointer block stores the location of the memory block in which the vector elements are located (if any). If the vector owns this block then the owner field is set to one and the block will be deallocated when the vector is freed. If the vector points to a block owned by another object then the owner field is zero and any underlying block will not be deallocated.

The functions for allocating and accessing vectors are defined in ‘gsl_vector.h’

8.3.1 Vector allocation

The functions for allocating memory to a vector follow the style of malloc and free. In addition they also perform their own error checking. If there is insufficient memory available to allocate a vector then the functions call the GSL error handler (with an error number of GSL_ENOMEM) in addition to returning a null pointer. Thus if you use the library error handler to abort your program then it isn’t necessary to check every alloc.

Function gsl_vector * gsl vector alloc (size_t n)

This function creates a vector of length n, returning a pointer to a newly initialized vector struct. A new block is allocated for the elements of the vector, and stored in the block component of the vector struct. The block is “owned” by the vector, and will be deallocated when the vector is deallocated.

Function gsl_vector * gsl vector calloc (size_t n)

This function allocates memory for a vector of length n and initializes all the elements of the vector to zero.

Function void gsl vector free (gsl_vector * v)

This function frees a previously allocated vector v. If the vector was created using gsl_vector_alloc then the block underlying the vector will also be deallocated. If the vector has been created from another object then the memory is still owned by that object and will not be deallocated.

8.3.2 Accessing vector elements

Unlike fortran compilers, C compilers do not usually provide support for range checking of vectors and matrices. Range checking is available in the GNU C Compiler extension checkergcc but it is not available on every platform. The functions gsl_vector_get and gsl_vector_set can perform portable range checking for you and report an error if you attempt to access elements outside the allowed range.

The functions for accessing the elements of a vector or matrix are defined in ‘gsl_vector.h’ and declared extern inline to eliminate function-call overhead. If necessary you can turn off range checking completely without modifying any source files by recompiling your program with the preprocessor definition GSL_RANGE_CHECK_OFF. Provided your compiler supports inline functions the effect of turning off range checking

is to replace calls to gsl_vector_get(v,i) by v->data[i*v->stride] and and calls
to gsl_vector_set(v,i,x) by v->data[i*v->stride]=x. Thus there should be no

performance penalty for using the range checking functions when range checking is turned off.

Function double gsl vector get (const gsl_vector * v, size_t i)

This function returns the i-th element of a vector v. If i lies outside the allowed range of 0 to n-1 then the error handler is invoked and 0 is returned.

Function void gsl vector set (gsl_vector * v, size_t i, double x)

This function sets the value of the i-th element of a vector v to x. If i lies outside the allowed range of 0 to n-1 then the error handler is invoked.

Function double * gsl vector ptr (gsl_vector * v, size_t i)
Function const double * gsl vector ptr (const gsl_vector * v, size_t i)

These functions return a pointer to the i-th element of a vector v. If i lies outside the allowed range of 0 to n-1 then the error handler is invoked and a null pointer is returned.

8.3.3 Initializing vector elements

Function void gsl vector set all (gsl_vector * v, double x)

This function sets all the elements of the vector v to the value x.

Function void gsl vector set zero (gsl_vector * v)

This function sets all the elements of the vector v to zero.

Function int gsl vector set basis (gsl_vector * v, size_t i)

This function makes a basis vector by setting all the elements of the vector v to zero except for the i-th element which is set to one.

8.3.4 Reading and writing vectors

The library provides functions for reading and writing vectors to a file as binary data or formatted text.

Function int gsl vector fwrite (FILE * stream, const gsl_vector * v)

This function writes the elements of the vector v to the stream stream in binary format. The return value is 0 for success and GSL_EFAILED if there was a problem writing to the file. Since the data is written in the native binary format it may not be portable between different architectures.

Function int gsl vector fread (FILE * stream, gsl_vector * v)

This function reads into the vector v from the open stream stream in binary format. The vector v must be preallocated with the correct length since the function uses the size of v to determine how many bytes to read. The return value is 0 for success and GSL_EFAILED if there was a problem reading from the file. The data is assumed to have been written in the native binary format on the same architecture.

Function int gsl vector fprintf (FILE * stream, const gsl_vector * v, const

char * format) This function writes the elements of the vector v line-by-line to the stream stream using the format specifier format, which should be one of the %g, %e or %f formats for floating point numbers and %d for integers. The function returns 0 for success and GSL_EFAILED if there was a problem writing to the file.

Function int gsl vector fscanf (FILE * stream, gsl_vector * v)

This function reads formatted data from the stream stream into the vector v. The vector v must be preallocated with the correct length since the function uses the size of v to determine how many numbers to read. The function returns 0 for success and GSL_EFAILED if there was a problem reading from the file.

8.3.5 Vector views

In addition to creating vectors from slices of blocks it is also possible to slice vectors and create vector views. For example, a subvector of another vector can be described with a view, or two views can be made which provide access to the even and odd elements of a vector.

A vector view is a temporary object, stored on the stack, which can be used to operate on a subset of vector elements. Vector views can be defined for both constant and nonconstant vectors, using separate types that preserve constness. A vector view has the type

gsl_vector_view and a constant vector view has the type gsl_vector_const_view. In

both cases the elements of the view can be accessed as a gsl_vector using the vector component of the view object. A pointer to a vector of type gsl_vector * or const gsl_ vector * can be obtained by taking the address of this component with the & operator.

Function gsl_vector_view gsl vector subvector (gsl_vector *v, size_t
    offset, size_t n)

Function gsl_vector_const_view gsl vector const subvector (const

    gsl_vector * v, size_t offset, size_t n)

These functions return a vector view of a subvector of another vector v. The start of the new vector is offset by offset elements from the start of the original vector. The new vector has n elements. Mathematically, the i-th element of the new vector v’ is given by,

     v'(i) = v->data[(offset + i)*v->stride]

where the index i runs from 0 to n-1. The data pointer of the returned vector struct is set to null if the combined parameters (offset,n) overrun the end of the original vector. The new vector is only a view of the block underlying the original vector, v. The block containing the elements of v is not owned by the new vector. When the view goes out of scope the original vector v and its block will continue to exist. The original memory can only be deallocated by freeing the original vector. Of course, the original vector should not be deallocated while the view is still in use.

The function gsl_vector_const_subvector is equivalent to gsl_vector_subvector

but can be used for vectors which are declared const.

Function gsl_vector gsl vector subvector with stride (gsl_vector *v,

    size_t offset, size_t stride, size_t n)
Function gsl_vector_const_view

gsl vector const subvector with stride (const gsl_vector * v,

    size_t offset, size_t stride, size_t n)

These functions return a vector view of a subvector of another vector v with an additional stride argument. The subvector is formed in the same way as for gsl_ vector_subvector but the new vector has n elements with a step-size of stride from one element to the next in the original vector. Mathematically, the i-th element of the new vector v’ is given by,

     v'(i) = v->data[(offset + i*stride)*v->stride]

where the index i runs from 0 to n-1. Note that subvector views give direct access to the underlying elements of the original vector. For example, the following code will zero the even elements of the vector v of length n, while leaving the odd elements untouched,

     gsl_vector_view v_even
       = gsl_vector_subvector_with_stride (v, 0, 2, n/2);
     gsl_vector_set_zero (&v_even.vector);

A vector view can be passed to any subroutine which takes a vector argument just as a directly allocated vector would be, using &view.vector. For example, the following code computes the norm of odd elements of v using the blas routine dnrm2,

     gsl_vector_view v_odd
       = gsl_vector_subvector_with_stride (v, 1, 2, n/2);
     double r = gsl_blas_dnrm2 (&v_odd.vector);
The function gsl_vector_const_subvector_with_stride is equivalent to
gsl_vector_subvector_with_stride but can be used for vectors which are

declared const.

Function gsl_vector_view gsl vector complex real (gsl_vector_complex

*v)

Function gsl_vector_const_view gsl vector complex const real (const

    gsl_vector_complex *v)

These functions return a vector view of the real parts of the complex vector v.

The function gsl_vector_complex_const_real is equivalent to gsl_vector_

complex_real but can be used for vectors which are declared const.

Function gsl_vector_view gsl vector complex imag (gsl_vector_complex

*v)

Function gsl_vector_const_view gsl vector complex const imag (const

    gsl_vector_complex *v)

These functions return a vector view of the imaginary parts of the complex vector v.

The function gsl_vector_complex_const_imag is equivalent to gsl_vector_

complex_imag but can be used for vectors which are declared const.

Function gsl_vector_view gsl vector view array (double *base, size_t n)

Function gsl_vector_const_view gsl vector const view array (const

    double *base, size_t n)

These functions return a vector view of an array. The start of the new vector is given by base and has n elements. Mathematically, the i-th element of the new vector v’ is given by,

     v'(i) = base[i]

where the index i runs from 0 to n-1. The array containing the elements of v is not owned by the new vector view. When the view goes out of scope the original array will continue to exist. The original memory can only be deallocated by freeing the original pointer base. Of course, the original array should not be deallocated while the view is still in use.

The function gsl_vector_const_view_array is equivalent to gsl_vector_view_

array but can be used for arrays which are declared const.

Function gsl_vector_view gsl vector view array with stride (double *

    base, size_t stride, size_t n)
Function gsl_vector_const_view

gsl vector const view array with stride (const double * base,

    size_t stride, size_t n)

These functions return a vector view of an array base with an additional stride argument. The subvector is formed in the same way as for gsl_vector_view_array but the new vector has n elements with a step-size of stride from one element to the next in the original array. Mathematically, the i-th element of the new vector v’ is given by,

     v'(i) = base[i*stride]

where the index i runs from 0 to n-1. Note that the view gives direct access to the underlying elements of the original array. A vector view can be passed to any subroutine which takes a vector argument just as a directly allocated vector would be, using &view.vector.

The function gsl_vector_const_view_array_with_stride is equivalent to gsl_

vector_view_array_with_stride but can be used for arrays which are declared

const.

8.3.6 Copying vectors

Common operations on vectors such as addition and multiplication are available in the blas part of the library (see Chapter 11 [BLAS Support], page 105). However, it is useful to have a small number of utility functions which do not require the full blas code. The following functions fall into this category.

Function int gsl vector memcpy (gsl_vector * dest, const gsl_vector *

src) This function copies the elements of the vector src into the vector dest. The two vectors must have the same length.

Function int gsl vector swap (gsl_vector * v, gsl_vector * w)

This function exchanges the elements of the vectors v and w by copying. The two vectors must have the same length.

8.3.7 Exchanging elements

The following function can be used to exchange, or permute, the elements of a vector.

Function int gsl vector swap elements (gsl_vector * v, size_t i, size_t

    j)

This function exchanges the i-th and j-th elements of the vector v in-place.

Function int gsl vector reverse (gsl_vector * v)

This function reverses the order of the elements of the vector v.

8.3.8 Vector operations

The following operations are only defined for real vectors.

Function int gsl vector add (gsl_vector * a, const gsl_vector * b)

This function adds the elements of vector b to the elements of vector a, a′_{i} = a_{i} + b_{i}. The two vectors must have the same length.

Function int gsl vector sub (gsl_vector * a, const gsl_vector * b)

This function subtracts the elements of vector b from the elements of vector a, a′_{i} = a_{i} − b_{i}. The two vectors must have the same length.

Function int gsl vector mul (gsl_vector * a, const gsl_vector * b)

This function multiplies the elements of vector a by the elements of vector b, a′_{i} = a_{i} ∗ b_{i}. The two vectors must have the same length.

Function int gsl vector div (gsl_vector * a, const gsl_vector * b)

This function divides the elements of vector a by the elements of vector b, a′_{i} = a_{i}/b_{i}. The two vectors must have the same length.

Function int gsl vector scale (gsl_vector * a, const double x)

This function multiplies the elements of vector a by the constant factor x, a′_{i} = xa_{i}.

Function int gsl vector add constant (gsl_vector * a, const double x)

This function adds the constant value x to the elements of the vector a, a′_{i} = a_{i} + x.

8.3.9 Finding maximum and minimum elements of vectors

Function double gsl vector max (const gsl_vector * v)

This function returns the maximum value in the vector v.

Function double gsl vector min (const gsl_vector * v)

This function returns the minimum value in the vector v.

Function void gsl vector minmax (const gsl_vector * v, double * min out,

    double * max out)

This function returns the minimum and maximum values in the vector v, storing them in min out and max out.

Function size_t gsl vector max index (const gsl_vector * v)

This function returns the index of the maximum value in the vector v. When there are several equal maximum elements then the lowest index is returned.

Function size_t gsl vector min index (const gsl_vector * v)

This function returns the index of the minimum value in the vector v. When there are several equal minimum elements then the lowest index is returned.

Function void gsl vector minmax index (const gsl_vector * v, size_t *

imin, size_t * imax) This function returns the indices of the minimum and maximum values in the vector v, storing them in imin and imax. When there are several equal minimum or maximum elements then the lowest indices are returned.

8.3.10 Vector properties

Function int gsl vector isnull (const gsl_vector * v)

This function returns 1 if all the elements of the vector v are zero, and 0 otherwise.

8.3.11 Example programs for vectors

This program shows how to allocate, initialize and read from a vector using the functions

gsl_vector_alloc, gsl_vector_set and gsl_vector_get.
#include <stdio.h>
#include <gsl/gsl_vector.h>
int
main (void)
{
  int i;
  gsl_vector * v = gsl_vector_alloc (3);
  for (i = 0; i < 3; i++)
    {
      gsl_vector_set (v, i, 1.23 + i);

    }
  for (i = 0; i < 100; i++)
    {
      printf("v_%d = %g\n", i, gsl_vector_get (v, i));
    }
  return 0;
}

Here is the output from the program. The final loop attempts to read outside the range of the vector v, and the error is trapped by the range-checking code in gsl_vector_get.

v_0 = 1.23
v_1 = 2.23
v_2 = 3.23
gsl: vector_source.c:12: ERROR: index out of range
IOT trap/Abort (core dumped)

The next program shows how to write a vector to a file.

#include <stdio.h>
#include <gsl/gsl_vector.h>
int
main (void)
{
  int i;
  gsl_vector * v = gsl_vector_alloc (100);
  for (i = 0; i < 100; i++)
    {
      gsl_vector_set (v, i, 1.23 + i);
    }
  {
     FILE * f = fopen("test.dat", "w");
     gsl_vector_fprintf (f, v, "%.5g");
     fclose (f);
  }
  return 0;
}

After running this program the file ‘test.dat’ should contain the elements of v, written using the format specifier %.5g. The vector could then be read back in using the function

gsl_vector_fscanf (f, v).

8.4 Matrices

Matrices are defined by a gsl_matrix structure which describes a generalized slice of a block. Like a vector it represents a set of elements in an area of memory, but uses two indices instead of one.

The gsl_matrix structure contains six components, the two dimensions of the matrix, a physical dimension, a pointer to the memory where the elements of the matrix are stored, data, a pointer to the block owned by the matrix block, if any, and an ownership flag, owner. The physical dimension determines the memory layout and can differ from the matrix dimension to allow the use of submatrices. The gsl_matrix structure is very simple and looks like this,

     typedef struct
     {
       size_t size1;
       size_t size2;
       size_t tda;
       double * data;
       gsl_block * block;
       int owner;
     } gsl_matrix;

Matrices are stored in row-major order, meaning that each row of elements forms a contiguous block in memory. This is the standard "C-language ordering" of two-dimensional arrays. Note that fortran stores arrays in column-major order. The number of rows is size1. The range of valid row indices runs from 0 to size1-1. Similarly size2 is the number of columns. The range of valid column indices runs from 0 to size2-1. The physical row dimension tda, or trailing dimension, specifies the size of a row of the matrix as laid out in memory.

For example, in the following matrix size1 is 3, size2 is 4, and tda is 8. The physical memory layout of the matrix begins in the top left hand-corner and proceeds from left to right along each row in turn.

     00 01 02 03 XX XX XX XX
     10 11 12 13 XX XX XX XX
     20 21 22 23 XX XX XX XX

Each unused memory location is represented by “XX”. The pointer data gives the location of the first element of the matrix in memory. The pointer block stores the location of the memory block in which the elements of the matrix are located (if any). If the matrix owns this block then the owner field is set to one and the block will be deallocated when the matrix is freed. If the matrix is only a slice of a block owned by another object then the owner field is zero and any underlying block will not be freed.

The functions for allocating and accessing matrices are defined in ‘gsl_matrix.h’

8.4.1 Matrix allocation

The functions for allocating memory to a matrix follow the style of malloc and free. They also perform their own error checking. If there is insufficient memory available to allocate a vector then the functions call the GSL error handler (with an error number of GSL_ENOMEM) in addition to returning a null pointer. Thus if you use the library error handler to abort your program then it isn’t necessary to check every alloc.

Function gsl_matrix * gsl matrix alloc (size_t n1, size_t n2)

This function creates a matrix of size n1 rows by n2 columns, returning a pointer to a newly initialized matrix struct. A new block is allocated for the elements of the matrix, and stored in the block component of the matrix struct. The block is “owned” by the matrix, and will be deallocated when the matrix is deallocated.

Function gsl_matrix * gsl matrix calloc (size_t n1, size_t n2)

This function allocates memory for a matrix of size n1 rows by n2 columns and initializes all the elements of the matrix to zero.

Function void gsl matrix free (gsl_matrix * m)

This function frees a previously allocated matrix m. If the matrix was created using gsl_matrix_alloc then the block underlying the matrix will also be deallocated. If the matrix has been created from another object then the memory is still owned by that object and will not be deallocated.

8.4.2 Accessing matrix elements

The functions for accessing the elements of a matrix use the same range checking system as vectors. You turn off range checking by recompiling your program with the preprocessor

definition GSL_RANGE_CHECK_OFF.

The elements of the matrix are stored in "C-order", where the second index moves continuously through memory. More precisely, the element accessed by the function gsl_

matrix_get(m,i,j) and gsl_matrix_set(m,i,j,x) is
m->data[i * m->tda + j]

where tda is the physical row-length of the matrix.

Function double gsl matrix get (const gsl_matrix * m, size_t i, size_t j)

This function returns the (i,j)th element of a matrix m. If i or j lie outside the allowed range of 0 to n1-1 and 0 to n2-1 then the error handler is invoked and 0 is returned.

Function void gsl matrix set (gsl_matrix * m, size_t i, size_t j, double

x) This function sets the value of the (i,j)th element of a matrix m to x. If i or j lies outside the allowed range of 0 to n1-1 and 0 to n2-1 then the error handler is invoked.

Function double * gsl matrix ptr (gsl_matrix * m, size_t i, size_t j)
Function const double * gsl matrix ptr (const gsl_matrix * m, size_t i,
    size_t j)

These functions return a pointer to the (i,j)th element of a matrix m. If i or j lie outside the allowed range of 0 to n1-1 and 0 to n2-1 then the error handler is invoked and a null pointer is returned.

8.4.3 Initializing matrix elements

Function void gsl matrix set all (gsl_matrix * m, double x)

This function sets all the elements of the matrix m to the value x.

Function void gsl matrix set zero (gsl_matrix * m)

This function sets all the elements of the matrix m to zero.

Function void gsl matrix set identity (gsl_matrix * m)

This function sets the elements of the matrix m to the corresponding elements of the identity matrix, m(i, j) = δ(i, j), i.e. a unit diagonal with all off-diagonal elements zero. This applies to both square and rectangular matrices.

8.4.4 Reading and writing matrices

The library provides functions for reading and writing matrices to a file as binary data or formatted text.

Function int gsl matrix fwrite (FILE * stream, const gsl_matrix * m)

This function writes the elements of the matrix m to the stream stream in binary format. The return value is 0 for success and GSL_EFAILED if there was a problem writing to the file. Since the data is written in the native binary format it may not be portable between different architectures.

Function int gsl matrix fread (FILE * stream, gsl_matrix * m)

This function reads into the matrix m from the open stream stream in binary format. The matrix m must be preallocated with the correct dimensions since the function uses the size of m to determine how many bytes to read. The return value is 0 for success and GSL_EFAILED if there was a problem reading from the file. The data is assumed to have been written in the native binary format on the same architecture.

Function int gsl matrix fprintf (FILE * stream, const gsl_matrix * m,

    const char * format)

This function writes the elements of the matrix m line-by-line to the stream stream using the format specifier format, which should be one of the %g, %e or %f formats for floating point numbers and %d for integers. The function returns 0 for success and GSL_EFAILED if there was a problem writing to the file.

Function int gsl matrix fscanf (FILE * stream, gsl_matrix * m)

This function reads formatted data from the stream stream into the matrix m. The matrix m must be preallocated with the correct dimensions since the function uses the size of m to determine how many numbers to read. The function returns 0 for success and GSL_EFAILED if there was a problem reading from the file.

8.4.5 Matrix views

A matrix view is a temporary object, stored on the stack, which can be used to operate on a subset of matrix elements. Matrix views can be defined for both constant and nonconstant matrices using separate types that preserve constness. A matrix view has the type

gsl_matrix_view and a constant matrix view has the type gsl_matrix_const_view. In

both cases the elements of the view can by accessed using the matrix component of the view object. A pointer gsl_matrix * or const gsl_matrix * can be obtained by taking the address of the matrix component with the & operator. In addition to matrix views it is also possible to create vector views of a matrix, such as row or column views.

Function gsl_matrix_view gsl matrix submatrix (gsl_matrix * m, size_t
    i, size_t j, size_t n1, size_t n2)

Function gsl_matrix_const_view gsl matrix const submatrix (const

    gsl_matrix * m, size_t i, size_t j, size_t n1, size_t n2)

These functions return a matrix view of a submatrix of the matrix m. The upper-left element of the submatrix is the element (k1,k2) of the original matrix. The submatrix has n1 rows and n2 columns. The physical number of columns in memory given by tda is unchanged. Mathematically, the (i,j)-th element of the new matrix is given by,

     m'(i,j) = m->data[(k1*m->tda + k2) + i*m->tda + j]

where the index i runs from 0 to n1-1 and the index j runs from 0 to n2-1. The data pointer of the returned matrix struct is set to null if the combined parameters (k1,k2,n1,n2,tda) overrun the ends of the original matrix. The new matrix view is only a view of the block underlying the existing matrix, m. The block containing the elements of m is not owned by the new matrix view. When the view goes out of scope the original matrix m and its block will continue to exist. The original memory can only be deallocated by freeing the original matrix. Of course, the original matrix should not be deallocated while the view is still in use.

The function gsl_matrix_const_submatrix is equivalent to gsl_matrix_submatrix

but can be used for matrices which are declared const.

Function gsl_matrix_view gsl matrix view array (double * base, size_t

    n1, size_t n2)

Function gsl_matrix_const_view gsl matrix const view array (const

    double * base, size_t n1, size_t n2)

These functions return a matrix view of the array base. The matrix has n1 rows and n2 columns. The physical number of columns in memory is also given by n2. Mathematically, the (i,j)-th element of the new matrix is given by,

     m'(i,j) = base[i*n2 + j]

where the index i runs from 0 to n1-1 and the index j runs from 0 to n2-1. The new matrix is only a view of the array base. When the view goes out of scope the original array base will continue to exist. The original memory can only be deallocated by freeing the original array. Of course, the original array should not be deallocated while the view is still in use.

The function gsl_matrix_const_view_array is equivalent to gsl_matrix_view_

array but can be used for matrices which are declared const.

Function gsl_matrix_view gsl matrix view array with tda (double *

    base, size_t n1, size_t n2, size_t tda)

Function gsl_matrix_const_view gsl matrix const view array with tda

    (const double * base, size_t n1, size_t n2, size_t tda)

These functions return a matrix view of the array base with a physical number of columns tda which may differ from corresponding the dimension of the matrix. The matrix has n1 rows and n2 columns, and the physical number of columns in memory is given by tda. Mathematically, the (i,j)-th element of the new matrix is given by,

     m'(i,j) = base[i*tda + j]

where the index i runs from 0 to n1-1 and the index j runs from 0 to n2-1. The new matrix is only a view of the array base. When the view goes out of scope the original array base will continue to exist. The original memory can only be deallocated by freeing the original array. Of course, the original array should not be deallocated while the view is still in use.

The function gsl_matrix_const_view_array_with_tda is equivalent to gsl_

matrix_view_array_with_tda but can be used for matrices which are declared

const.

Function gsl_matrix_view gsl matrix view vector (gsl_vector * v,

    size_t n1, size_t n2)

Function gsl_matrix_const_view gsl matrix const view vector (const

    gsl_vector * v, size_t n1, size_t n2)

These functions return a matrix view of the vector v. The matrix has n1 rows and n2 columns. The vector must have unit stride. The physical number of columns in memory is also given by n2. Mathematically, the (i,j)-th element of the new matrix is given by,

     m'(i,j) = v->data[i*n2 + j]

where the index i runs from 0 to n1-1 and the index j runs from 0 to n2-1. The new matrix is only a view of the vector v. When the view goes out of scope the original vector v will continue to exist. The original memory can only be deallocated by freeing the original vector. Of course, the original vector should not be deallocated while the view is still in use.

The function gsl_matrix_const_view_vector is equivalent to gsl_matrix_view_

vector but can be used for matrices which are declared const.

Function gsl_matrix_view gsl matrix view vector with tda (gsl_vector

    * v, size_t n1, size_t n2, size_t tda)
Function gsl_matrix_const_view

gsl matrix const view vector with tda (const gsl_vector * v,

    size_t n1, size_t n2, size_t tda)

These functions return a matrix view of the vector v with a physical number of columns tda which may differ from the corresponding matrix dimension. The vector must have unit stride. The matrix has n1 rows and n2 columns, and the physical number of columns in memory is given by tda. Mathematically, the (i,j)-th element of the new matrix is given by,

     m'(i,j) = v->data[i*tda + j]

where the index i runs from 0 to n1-1 and the index j runs from 0 to n2-1. The new matrix is only a view of the vector v. When the view goes out of scope the original vector v will continue to exist. The original memory can only be deallocated by freeing the original vector. Of course, the original vector should not be deallocated while the view is still in use.

The function gsl_matrix_const_view_vector_with_tda is equivalent to
gsl_matrix_view_vector_with_tda but can be used for matrices which are

declared const.

8.4.6 Creating row and column views

In general there are two ways to access an object, by reference or by copying. The functions described in this section create vector views which allow access to a row or column of a matrix by reference. Modifying elements of the view is equivalent to modifying the matrix, since both the vector view and the matrix point to the same memory block.

Function gsl_vector_view gsl matrix row (gsl_matrix * m, size_t i)
Function gsl_vector_const_view gsl matrix const row (const gsl_matrix
    * m, size_t i)

These functions return a vector view of the i-th row of the matrix m. The data pointer of the new vector is set to null if i is out of range.

The function gsl_vector_const_row is equivalent to gsl_matrix_row but can be

used for matrices which are declared const.

Function gsl_vector_view gsl matrix column (gsl_matrix * m, size_t j)

Function gsl_vector_const_view gsl matrix const column (const

    gsl_matrix * m, size_t j)

These functions return a vector view of the j-th column of the matrix m. The data pointer of the new vector is set to null if j is out of range.

The function gsl_vector_const_column equivalent to gsl_matrix_column but can

be used for matrices which are declared const.

Function gsl_vector_view gsl matrix diagonal (gsl_matrix * m) Function gsl_vector_const_view gsl matrix const diagonal (const

    gsl_matrix * m)

These functions returns a vector view of the diagonal of the matrix m. The matrix m is not required to be square. For a rectangular matrix the length of the diagonal is the same as the smaller dimension of the matrix.

The function gsl_matrix_const_diagonal is equivalent to gsl_matrix_diagonal

but can be used for matrices which are declared const.

Function gsl_vector_view gsl matrix subdiagonal (gsl_matrix * m,

    size_t k)

Function gsl_vector_const_view gsl matrix const subdiagonal (const

    gsl_matrix * m, size_t k)

These functions return a vector view of the k-th subdiagonal of the matrix m. The matrix m is not required to be square. The diagonal of the matrix corresponds to k = 0.

The function gsl_matrix_const_subdiagonal is equivalent to gsl_matrix_

subdiagonal but can be used for matrices which are declared const.

Function gsl_vector_view gsl matrix superdiagonal (gsl_matrix * m,

    size_t k)

Function gsl_vector_const_view gsl matrix const superdiagonal (const

    gsl_matrix * m, size_t k)

These functions return a vector view of the k-th superdiagonal of the matrix m. The matrix m is not required to be square. The diagonal of the matrix corresponds to k = 0.

The function gsl_matrix_const_superdiagonal is equivalent to gsl_matrix_

superdiagonal but can be used for matrices which are declared const.

8.4.7 Copying matrices

Function int gsl matrix memcpy (gsl_matrix * dest, const gsl_matrix *

src) This function copies the elements of the matrix src into the matrix dest. The two matrices must have the same size.

Function int gsl matrix swap (gsl_matrix * m1, const gsl_matrix * m2)

This function exchanges the elements of the matrices m1 and m2 by copying. The two matrices must have the same size.

8.4.8 Copying rows and columns

The functions described in this section copy a row or column of a matrix into a vector. This allows the elements of the vector and the matrix to be modified independently. Note that if the matrix and the vector point to overlapping regions of memory then the result will be undefined. The same effect can be achieved with more generality using gsl_vector_ memcpy with vector views of rows and columns.

Function int gsl matrix get row (gsl_vector * v, const gsl_matrix * m,
    size_t i)

This function copies the elements of the i-th row of the matrix m into the vector v. The length of the vector must be the same as the length of the row.

Function int gsl matrix get col (gsl_vector * v, const gsl_matrix * m,
    size_t j)

This function copies the elements of the i-th column of the matrix m into the vector v. The length of the vector must be the same as the length of the column.

Function int gsl matrix set row (gsl_matrix * m, size_t i, const
    gsl_vector * v)

This function copies the elements of the vector v into the i-th row of the matrix m. The length of the vector must be the same as the length of the row.

Function int gsl matrix set col (gsl_matrix * m, size_t j, const
    gsl_vector * v)

This function copies the elements of the vector v into the i-th column of the matrix m. The length of the vector must be the same as the length of the column.

8.4.9 Exchanging rows and columns

The following functions can be used to exchange the rows and columns of a matrix.

Function int gsl matrix swap rows (gsl_matrix * m, size_t i, size_t j)

This function exchanges the i-th and j-th rows of the matrix m in-place.

Function int gsl matrix swap columns (gsl_matrix * m, size_t i, size_t

j) This function exchanges the i-th and j-th columns of the matrix m in-place.

Function int gsl matrix swap rowcol (gsl_matrix * m, size_t i, size_t j)

This function exchanges the i-th row and j-th column of the matrix m in-place. The matrix must be square for this operation to be possible.

Function int gsl matrix transpose memcpy (gsl_matrix * dest,

    gsl_matrix * src)

This function makes the matrix dest the transpose of the matrix src by copying the elements of src into dest. This function works for all matrices provided that the dimensions of the matrix dest match the transposed dimensions of the matrix src.

Function int gsl matrix transpose (gsl_matrix * m)

This function replaces the matrix m by its transpose by copying the elements of the matrix in-place. The matrix must be square for this operation to be possible.

8.4.10 Matrix operations

The following operations are only defined for real matrices.

Function int gsl matrix add (gsl_matrix * a, const gsl_matrix * b)

This function adds the elements of matrix b to the elements of matrix a, a′(i, j) = a(i, j) + b(i, j). The two matrices must have the same dimensions.

Function int gsl matrix sub (gsl_matrix * a, const gsl_matrix * b)

This function subtracts the elements of matrix b from the elements of matrix a, a′(i, j) = a(i, j) − b(i, j). The two matrices must have the same dimensions.

Function int gsl matrix mul elements (gsl_matrix * a, const gsl_matrix

* b) This function multiplies the elements of matrix a by the elements of matrix b, a′(i, j) = a(i, j) ∗ b(i, j). The two matrices must have the same dimensions.

Function int gsl matrix div elements (gsl_matrix * a, const gsl_matrix *

b) This function divides the elements of matrix a by the elements of matrix b, a′(i, j) = a(i, j)/b(i, j). The two matrices must have the same dimensions.

Function int gsl matrix scale (gsl_matrix * a, const double x)

This function multiplies the elements of matrix a by the constant factor x, a′(i, j) = xa(i, j).

Function int gsl matrix add constant (gsl_matrix * a, const double x)

This function adds the constant value x to the elements of the matrix a, a′(i, j) = a(i, j) + x.

8.4.11 Finding maximum and minimum elements of matrices

Function double gsl matrix max (const gsl_matrix * m)

This function returns the maximum value in the matrix m.

Function double gsl matrix min (const gsl_matrix * m)

This function returns the minimum value in the matrix m.

Function void gsl matrix minmax (const gsl_matrix * m, double *

min out, double * max out) This function returns the minimum and maximum values in the matrix m, storing them in min out and max out.

Function void gsl matrix max index (const gsl_matrix * m, size_t *

imax, size_t * jmax) This function returns the indices of the maximum value in the matrix m, storing them in imax and jmax. When there are several equal maximum elements then the first element found is returned.

Function void gsl matrix min index (const gsl_matrix * m, size_t *

imax, size_t * jmax) This function returns the indices of the minimum value in the matrix m, storing them in imax and jmax. When there are several equal minimum elements then the first element found is returned.

Function void gsl matrix minmax index (const gsl_matrix * m, size_t *

imin, size_t * imax) This function returns the indices of the minimum and maximum values in the matrix m, storing them in (imin,jmin) and (imax,jmax). When there are several equal minimum or maximum elements then the first elements found are returned.

8.4.12 Matrix properties

Function int gsl matrix isnull (const gsl_matrix * m)

This function returns 1 if all the elements of the matrix m are zero, and 0 otherwise.

8.4.13 Example programs for matrices

The program below shows how to allocate, initialize and read from a matrix using the

functions gsl_matrix_alloc, gsl_matrix_set and gsl_matrix_get.
#include <stdio.h>
#include <gsl/gsl_matrix.h>
int
main (void)
{
  int i, j;
  gsl_matrix * m = gsl_matrix_alloc (10, 3);
  for (i = 0; i < 10; i++)
    for (j = 0; j < 3; j++)
      gsl_matrix_set (m, i, j, 0.23 + 100*i + j);
  for (i = 0; i < 100; i++)
    for (j = 0; j < 3; j++)
      printf("m(%d,%d) = %g\n", i, j,
             gsl_matrix_get (m, i, j));
  return 0;
}

Here is the output from the program. The final loop attempts to read outside the range of the matrix m, and the error is trapped by the range-checking code in gsl_matrix_get.

m(0,0) = 0.23
m(0,1) = 1.23
m(0,2) = 2.23
m(1,0) = 100.23
m(1,1) = 101.23
m(1,2) = 102.23
...
m(9,2) = 902.23
gsl: matrix_source.c:13: ERROR: first index out of range
IOT trap/Abort (core dumped)

The next program shows how to write a matrix to a file.

#include <stdio.h>
#include <gsl/gsl_matrix.h>
int
main (void)
{
  int i, j, k = 0;

gsl_matrix * m = gsl_matrix_alloc (100, 100);
gsl_matrix * a = gsl_matrix_alloc (100, 100);
for (i = 0; i < 100; i++)
  for (j = 0; j < 100; j++)
    gsl_matrix_set (m, i, j, 0.23 + i + j);
{
   FILE * f = fopen("test.dat", "w");
   gsl_matrix_fwrite (f, m);
   fclose (f);
}
{
   FILE * f = fopen("test.dat", "r");
   gsl_matrix_fread (f, a);
   fclose (f);
}
for (i = 0; i < 100; i++)
  for (j = 0; j < 100; j++)
    {
      double mij = gsl_matrix_get(m, i, j);
      double aij = gsl_matrix_get(a, i, j);
      if (mij != aij) k++;
    }
printf("differences = %d (should be zero)\n", k);
return (k > 0);
}

After running this program the file ‘test.dat’ should contain the elements of m, written in binary format. The matrix which is read back in using the function gsl_matrix_fread should be exactly equal to the original matrix.

The following program demonstrates the use of vector views. The program computes the column-norms of a matrix.

#include <math.h>
#include <stdio.h>
#include <gsl/gsl_matrix.h>
#include <gsl/gsl_blas.h>
int
main (void)
{
size_t i,j;
gsl_matrix *m = gsl_matrix_alloc (10, 10);
for (i = 0; i < 10; i++)
  for (j = 0; j < 10; j++)
    gsl_matrix_set (m, i, j, sin (i) + cos (j));

  for (j = 0; j < 10; j++)
    {
      gsl_vector_view column = gsl_matrix_column (m, j);
      double d;
      d = gsl_blas_dnrm2 (&column.vector);
      printf ("matrix column %d, norm = %g\n", j, d);
    }
  gsl_matrix_free (m);
}

Here is the output of the program, which can be confirmed using gnu Octave,

$ ./a.out
matrix column 0, norm = 4.31461
matrix column 1, norm = 3.1205
matrix column 2, norm = 2.19316
matrix column 3, norm = 3.26114
matrix column 4, norm = 2.53416
matrix column 5, norm = 2.57281
matrix column 6, norm = 4.20469
matrix column 7, norm = 3.65202
matrix column 8, norm = 2.08524
matrix column 9, norm = 3.07313
octave> m = sin(0:9)' * ones(1,10)
               + ones(10,1) * cos(0:9);
octave> sqrt(sum(m.^2))
ans =
  4.3146  3.1205  2.1932  3.2611  2.5342  2.5728
  4.2047  3.6520  2.0852  3.0731

8.5 References and Further Reading

The block, vector and matrix objects in GSL follow the valarray model of C++. A description of this model can be found in the following reference,

B. Stroustrup, The C++ Programming Language (3rd Ed), Section 22.4 Vector Arithmetic. Addison-Wesley 1997, ISBN 0-201-88954-4.

9 Permutations

This chapter describes functions for creating and manipulating permutations. A permutation p is represented by an array of n integers in the range 0 .. n − 1, where each value p_{i} occurs once and only once. The application of a permutation p to a vector v yields a new vector v′ where v′_{i} = v_{p_{i}}. For example, the array (0, 1, 3, 2) represents a permutation which exchanges the last two elements of a four element vector. The corresponding identity permutation is (0, 1, 2, 3).

Note that the permutations produced by the linear algebra routines correspond to the exchange of matrix columns, and so should be considered as applying to row-vectors in the form v′ = vP rather than column-vectors, when permuting the elements of a vector.

The functions described in this chapter are defined in the header file

‘gsl_permutation.h’.

9.1 The Permutation struct

A permutation is stored by a structure containing two components, the size of the permutation and a pointer to the permutation array. The elements of the permutation array are all of type size_t. The gsl_permutation structure looks like this,

typedef struct
{
  size_t size;
  size_t * data;
} gsl_permutation;

9.2 Permutation allocation

Function gsl_permutation * gsl permutation alloc (size_t n)

This function allocates memory for a new permutation of size n. The permutation is not initialized and its elements are undefined. Use the function gsl_permutation_ calloc if you want to create a permutation which is initialized to the identity. A null pointer is returned if insufficient memory is available to create the permutation.

Function gsl_permutation * gsl permutation calloc (size_t n)

This function allocates memory for a new permutation of size n and initializes it to the identity. A null pointer is returned if insufficient memory is available to create the permutation.

Function void gsl permutation init (gsl_permutation * p)

This function initializes the permutation p to the identity, i.e. (0, 1, 2, ..., n − 1).

Function void gsl permutation free (gsl_permutation * p)

This function frees all the memory used by the permutation p.

9.3 Accessing permutation elements

The following functions can be used to access and manipulate permutations.

Function size_t gsl permutation get (const gsl_permutation * p, const
    size_t i)

This function returns the value of the i-th element of the permutation p. If i lies outside the allowed range of 0 to n-1 then the error handler is invoked and 0 is returned.

Function int gsl permutation swap (gsl_permutation * p, const size_t i,
    const size_t j)

This function exchanges the i-th and j-th elements of the permutation p.

9.4 Permutation properties

Function size_t gsl permutation size (const gsl_permutation * p)

This function returns the size of the permutation p.

Function size_t * gsl permutation data (const gsl_permutation * p)

This function returns a pointer to the array of elements in the permutation p.

Function int gsl permutation valid (gsl_permutation * p)

This function checks that the permutation p is valid. The n elements should contain each of the numbers 0 .. n-1 once and only once.

9.5 Permutation functions

Function void gsl permutation reverse (gsl_permutation * p)

This function reverses the elements of the permutation p.

Function int gsl permutation inverse (gsl_permutation * inv, const

    gsl_permutation * p)

This function computes the inverse of the permutation p, storing the result in inv.

Function int gsl permutation next (gsl_permutation * p)

This function advances the permutation p to the next permutation in lexicographic order and returns GSL_SUCCESS. If no further permutations are available it returns GSL_FAILURE and leaves p unmodified. Starting with the identity permutation and repeatedly applying this function will iterate through all possible permutations of a given order.

Function int gsl permutation prev (gsl_permutation * p)

This function steps backwards from the permutation p to the previous permutation in lexicographic order, returning GSL_SUCCESS. If no previous permutation is available it returns GSL_FAILURE and leaves p unmodified.

9.6 Applying Permutations

Function int gsl permute (const size_t * p, double * data, size_t stride,
    size_t n)

This function applies the permutation p to the array data of size n with stride stride.

Function int gsl permute inverse (const size_t * p, double * data, size_t

stride, size_t n) This function applies the inverse of the permutation p to the array data of size n with stride stride.

Function int gsl permute vector (const gsl_permutation * p, gsl_vector *

v) This function applies the permutation p to the elements of the vector v, considered as a row-vector acted on by a permutation matrix from the right, v′ = vP. The j-th column of the permutation matrix P is given by the p_{j}-th column of the identity matrix. The permutation p and the vector v must have the same length.

Function int gsl permute vector inverse (const gsl_permutation * p,

    gsl_vector * v)

This function applies the inverse of the permutation p to the elements of the vector v, considered as a row-vector acted on by an inverse permutation matrix from the right, v′ = vP^{T}. Note that for permutation matrices the inverse is the same as the transpose. The j-th column of the permutation matrix P is given by the p_{j}-th column of the identity matrix. The permutation p and the vector v must have the same length.

9.7 Reading and writing permutations

The library provides functions for reading and writing permutations to a file as binary data or formatted text.

Function int gsl permutation fwrite (FILE * stream, const

    gsl_permutation * p)

This function writes the elements of the permutation p to the stream stream in binary format. The function returns GSL_EFAILED if there was a problem writing to the file. Since the data is written in the native binary format it may not be portable between different architectures.

Function int gsl permutation fread (FILE * stream, gsl_permutation * p)

This function reads into the permutation p from the open stream stream in binary format. The permutation p must be preallocated with the correct length since the function uses the size of p to determine how many bytes to read. The function returns GSL_EFAILED if there was a problem reading from the file. The data is assumed to have been written in the native binary format on the same architecture.

Function int gsl permutation fprintf (FILE * stream, const

  gsl_permutation * p, const char *format)

This function writes the elements of the permutation p line-by-line to the stream stream using the format specifier format, which should be suitable for a type of size t. On a GNU system the type modifier Z represents size_t, so "%Zu\n" is a suitable format. The function returns GSL_EFAILED if there was a problem writing to the file.

Function int gsl permutation fscanf (FILE * stream, gsl_permutation * p)

This function reads formatted data from the stream stream into the permutation p. The permutation p must be preallocated with the correct length since the function uses the size of p to determine how many numbers to read. The function returns GSL_EFAILED if there was a problem reading from the file.

9.8 Examples

The example program below creates a random permutation by shuffling and finds its inverse.

#include <stdio.h>
#include <gsl/gsl_rng.h>
#include <gsl/gsl_randist.h>
#include <gsl/gsl_permutation.h>
int
main (void)
{
const size_t N = 10;
const gsl_rng_type * T;
gsl_rng * r;
gsl_permutation * p = gsl_permutation_alloc (N);
gsl_permutation * q = gsl_permutation_alloc (N);
gsl_rng_env_setup();
T = gsl_rng_default;
r = gsl_rng_alloc (T);
printf("initial permutation:");
gsl_permutation_init (p);
gsl_permutation_fprintf (stdout, p, " %u");
printf("\n");
printf(" random permutation:");
gsl_ran_shuffle (r, p->data, N, sizeof(size_t));
gsl_permutation_fprintf (stdout, p, " %u");
printf("\n");
printf("inverse permutation:");
gsl_permutation_invert (q, p);
gsl_permutation_fprintf (stdout, q, " %u");
printf("\n");

  return 0;
}

Here is the output from the program,

bash$ ./a.out
initial permutation: 0 1 2 3 4 5 6 7 8 9
 random permutation: 1 3 5 2 7 6 0 4 9 8
inverse permutation: 6 0 3 1 7 2 5 4 9 8

The random permutation p[i] and its inverse q[i] are related through the identity p[q[i]] = i, which can be verified from the output.

The next example program steps forwards through all possible 3-rd order permutations, starting from the identity,

#include <stdio.h>
#include <gsl/gsl_permutation.h>
int
main (void)
{
  gsl_permutation * p = gsl_permutation_alloc (3);
  gsl_permutation_init (p);
  do
   {
      gsl_permutation_fprintf (stdout, p, " %u");
      printf("\n");
   }
  while (gsl_permutation_next(p) == GSL_SUCCESS);
  return 0;
}

Here is the output from the program,

bash$ ./a.out
 0 1 2
 0 2 1
 1 0 2
 1 2 0
 2 0 1
 2 1 0

All 6 permutations are generated in lexicographic order. To reverse the sequence, begin with the final permutation (which is the reverse of the identity) and replace gsl_permutation_

next with gsl_permutation_prev.

9.9 References and Further Reading

The subject of permutations is covered extensively in Knuth’s Sorting and Searching, Donald E. Knuth, The Art of Computer Programming: Sorting and Searching (Vol 3, 3rd Ed, 1997), Addison-Wesley, ISBN 0201896850.

10 Sorting

This chapter describes functions for sorting data, both directly and indirectly (using an index). All the functions use the heapsort algorithm. Heapsort is an O(N log N) algorithm which operates in-place. It does not require any additional storage and provides consistent performance. The running time for its worst-case (ordered data) is not significantly longer than the average and best cases. Note that the heapsort algorithm does not preserve the relative ordering of equal elements — it is an unstable sort. However the resulting order of equal elements will be consistent across different platforms when using these functions.

10.1 Sorting objects

The following function provides a simple alternative to the standard library function qsort. It is intended for systems lacking qsort, not as a replacement for it. The function qsort should be used whenever possible, as it will be faster and can provide stable ordering of equal elements. Documentation for qsort is available in the GNU C Library Reference Manual.

The functions described in this section are defined in the header file ‘gsl_heapsort.h’.

Function void gsl heapsort (void * array, size_t count, size_t size,

    gsl_comparison_fn_t compare)

This function sorts the count elements of the array array, each of size size, into ascending order using the comparison function compare. The type of the comparison function is defined by,

     int (*gsl_comparison_fn_t) (const void * a,
                                 const void * b)

A comparison function should return a negative integer if the first argument is less than the second argument, 0 if the two arguments are equal and a positive integer if the first argument is greater than the second argument. For example, the following function can be used to sort doubles into ascending numerical order.

     int
     compare_doubles (const double * a,
                      const double * b)
     {
         return (int) (*a - *b);
     }

The appropriate function call to perform the sort is,

     gsl_heapsort (array, count, sizeof(double),
                   compare_doubles);

Note that unlike qsort the heapsort algorithm cannot be made into a stable sort by pointer arithmetic. The trick of comparing pointers for equal elements in the comparison function does not work for the heapsort algorithm. The heapsort algorithm performs an internal rearrangement of the data which destroys its initial ordering.

Function int gsl heapsort index (size_t * p, const void * array, size_t

    count, size_t size, gsl_comparison_fn_t compare)

This function indirectly sorts the count elements of the array array, each of size size, into ascending order using the comparison function compare. The resulting permutation is stored in p, an array of length n. The elements of p give the index of the array element which would have been stored in that position if the array had been sorted in place. The first element of p gives the index of the least element in array, and the last element of p gives the index of the greatest element in array. The array itself is not changed.

10.2 Sorting vectors

The following functions will sort the elements of an array or vector, either directly or indirectly. They are defined for all real and integer types using the normal suffix rules. For example, the float versions of the array functions are gsl_sort_float and gsl_sort_float_ index. The corresponding vector functions are gsl_sort_vector_float and gsl_sort_ vector_float_index. The prototypes are available in the header files ‘gsl_sort_float.h’ ‘gsl_sort_vector_float.h’. The complete set of prototypes can be included using the

header files ‘gsl_sort.h’ and ‘gsl_sort_vector.h’.

There are no functions for sorting complex arrays or vectors, since the ordering of complex numbers is not uniquely defined. To sort a complex vector by magnitude compute a real vector containing the the magnitudes of the complex elements, and sort this vector indirectly. The resulting index gives the appropriate ordering of the original complex vector.

Function void gsl sort (double * data, size_t stride, size_t n)

This function sorts the n elements of the array data with stride stride into ascending numerical order.

Function void gsl sort vector (gsl_vector * v)

This function sorts the elements of the vector v into ascending numerical order.

Function int gsl sort index (size_t * p, const double * data, size_t stride,

    size_t n)

This function indirectly sorts the n elements of the array data with stride stride into ascending order, storing the resulting permutation in p. The array p must be allocated to a sufficient length to store the n elements of the permutation. The elements of p give the index of the array element which would have been stored in that position if the array had been sorted in place. The array data is not changed.

Function int gsl sort vector index (gsl_permutation * p, const

    gsl_vector * v)

This function indirectly sorts the elements of the vector v into ascending order, storing the resulting permutation in p. The elements of p give the index of the vector element which would have been stored in that position if the vector had been sorted in place. The first element of p gives the index of the least element in v, and the last element of p gives the index of the greatest element in v. The vector v is not changed.

10.3 Selecting the k-th smallest or largest elements

The functions described in this section select the k-th smallest or largest elements of a data set of size N. The routines use an O(kN) direct insertion algorithm which is suited to subsets that are small compared with the total size of the dataset. For example, the routines are useful for selecting the 10 largest values from one million data points, but not for selecting the largest 100,000 values. If the subset is a significant part of the total dataset it may be faster to sort all the elements of the dataset directly with an O(N log N) algorithm and obtain the smallest or largest values that way.

Function void gsl sort smallest (double * dest, size_t k, const double *
    src, size_t stride, size_t n)

This function copies the k-th smallest elements of the array src, of size n and stride stride, in ascending numerical order in dest. The size of the subset k must be less than or equal to n. The data src is not modified by this operation.

Function void gsl sort largest (double * dest, size_t k, const double * src,
    size_t stride, size_t n)

This function copies the k-th largest elements of the array src, of size n and stride stride, in descending numerical order in dest. The size of the subset k must be less than or equal to n. The data src is not modified by this operation.

Function void gsl sort vector smallest (double * dest, size_t k, const

    gsl_vector * v)

Function void gsl sort vector largest (double * dest, size_t k, const

    gsl_vector * v)

These functions copy the k-th smallest or largest elements of the vector v into the array dest. The size of the subset k must be less than or equal to the length of the vector v.

The following functions find the indices of the k-th smallest or largest elements of a dataset,

Function void gsl sort smallest index (size_t * p, size_t k, const double

    * src, size_t stride, size_t n)

This function stores the indices of the k-th smallest elements of the array src, of size n and stride stride, in the array p. The indices are chosen so that the corresponding data is in ascending numerical order. The size of the subset k must be less than or equal to n. The data src is not modified by this operation.

Function void gsl sort largest index (size_t * p, size_t k, const double *
    src, size_t stride, size_t n)

This function stores the indices of the k-th largest elements of the array src, of size n and stride stride, in the array p. The indices are chosen so that the corresponding data is in descending numerical order. The size of the subset k must be less than or equal to n. The data src is not modified by this operation.

Function void gsl sort vector smallest index (size_t * p, size_t k,

    const gsl_vector * v)

Function void gsl sort vector largest index (size_t * p, size_t k, const

    gsl_vector * v)

These functions store the indices of k-th smallest or largest elements of the vector v in the array p. The size of the subset k must be less than or equal to the length of the vector v.

10.4 Computing the rank

The rank of an element is its order in the sorted data. The rank is the inverse of the index permutation, p. It can be computed using the following algorithm,

for (i = 0; i < p->size; i++)
{
    size_t pi = p->data[i];
    rank->data[pi] = i;
}

This can be computed directly from the function gsl_permutation_invert(rank,p).

The following function will print the rank of each element of the vector v,

void
print_rank (gsl_vector * v)
{
  size_t i;
  size_t n = v->size;
  gsl_permutation * perm = gsl_permutation_alloc(n);
  gsl_permutation * rank = gsl_permutation_alloc(n);
  gsl_sort_vector_index (perm, v);
  gsl_permutation_invert (rank, perm);
  for (i = 0; i < n; i++)
   {
    double vi = gsl_vector_get(v, i);
    printf("element = %d, value = %g, rank = %d\n",
            i, vi, rank->data[i]);
   }
  gsl_permutation_free (perm);
  gsl_permutation_free (rank);
}

10.5 Examples

The following example shows how to use the permutation p to print the elements of the vector v in ascending order,

gsl_sort_vector_index (p, v);
for (i = 0; i < v->size; i++)
{
  double vpi = gsl_vector_get(v, p->data[i]);
  printf("order = %d, value = %g\n", i, vpi);
}

The next example uses the function gsl_sort_smallest to select the 5 smallest numbers from 100000 uniform random variates stored in an array,

#include <gsl/gsl_rng.h>
#include <gsl/gsl_sort_double.h>
int
main (void)
{
const gsl_rng_type * T;
gsl_rng * r;
int i, k = 5, N = 100000;
double * x = malloc (N * sizeof(double));
double * small = malloc (k * sizeof(double));
gsl_rng_env_setup();
T = gsl_rng_default;
r = gsl_rng_alloc (T);
for (i = 0; i < N; i++)
  {
    x[i] = gsl_rng_uniform(r);
  }
gsl_sort_smallest (small, k, x, 1, N);
printf("%d smallest values from %d\n", k, N);
for (i = 0; i < k; i++)
  {
    printf ("%d: %.18f\n", i, small[i]);
  }
return 0;
}

The output lists the 5 smallest values, in ascending order,

$ ./a.out
5 smallest values from 100000
0: 0.000005466630682349
1: 0.000012384494766593
2: 0.000017581274732947
3: 0.000025131041184068
4: 0.000031369971111417

10.6 References and Further Reading

The subject of sorting is covered extensively in Knuth’s Sorting and Searching, Donald E. Knuth, The Art of Computer Programming: Sorting and Searching (Vol 3, 3rd Ed, 1997), Addison-Wesley, ISBN 0201896850.

The Heapsort algorithm is described in the following book, Robert Sedgewick, Algorithms in C, Addison-Wesley, ISBN 0201514257.

11 BLAS Support

The Basic Linear Algebra Subprograms (blas) define a set of fundamental operations on vectors and matrices which can be used to create optimized higher-level linear algebra functionality.

The library provides a low-level layer which corresponds directly to the C-language blas standard, referred to here as “cblas”, and a higher-level interface for operations on GSL vectors and matrices. Users who are interested in simple operations on GSL vector and matrix objects should use the high-level layer, which is declared in the file gsl_blas.h. This should satisfy the needs of most users. Note that GSL matrices are implemented using dense-storage so the interface only includes the corresponding dense-storage blas functions. The full blas functionality for band-format and packed-format matrices is available through the low-level cblas interface.

The interface for the gsl_cblas layer is specified in the file gsl_cblas.h. This interface corresponds the blas Technical Forum’s draft standard for the C interface to legacy blas implementations. Users who have access to other conforming cblas implementations can use these in place of the version provided by the library. Note that users who have only a Fortran blas library can use a cblas conformant wrapper to convert it into a cblas library. A reference cblas wrapper for legacy Fortran implementations exists as part of the draft cblas standard and can be obtained from Netlib. The complete set of cblas functions is listed in an appendix (see Appendix D [GSL CBLAS Library], page 378).

There are three levels of blas operations,

Level 1 Vector operations, e.g. y = αx + y

Level 2 Matrix-vector operations, e.g. y = αAx + βy

Level 3 Matrix-matrix operations, e.g. C = αAB + C

Each routine has a name which specifies the operation, the type of matrices involved and their precisions. Some of the most common operations and their names are given below,

DOT scalar product, x^{T} y

AXPY vector sum, αx + y

MV matrix-vector product, Ax

SV matrix-vector solve, inv(A)x

MM matrix-matrix product, AB

SM matrix-matrix solve, inv(A)B

The type of matrices are,

GE general

GB general band

SY symmetric

SB symmetric band

SP symmetric packed

HE hermitian HB hermitian band

HP hermitian packed

TR triangular

TB triangular band

TP triangular packed

Each operation is defined for four precisions,

S single real

D double real

C single complex

Z double complex

Thus, for example, the name sgemm stands for “single-precision general matrix-matrix multiply” and zgemm stands for “double-precision complex matrix-matrix multiply”.

11.1 GSL BLAS Interface

GSL provides dense vector and matrix objects, based on the relevant built-in types. The library provides an interface to the blas operations which apply to these objects. The interface to this functionality is given in the file gsl_blas.h.

11.1.1 Level 1

Function int gsl blas sdsdot (float alpha, const gsl_vector_float * x,
         const gsl_vector_float * y, float * result)
Function int gsl blas dsdot (const gsl_vector_float * x, const
         gsl_vector_float * y, double * result)

These functions compute the sum α+x^{T} y for the vectors x and y, returning the result in result.

Function int gsl blas sdot (const gsl_vector_float * x, const
         gsl_vector_float * y, float * result)
Function int gsl blas ddot (const gsl_vector * x, const gsl_vector * y,
         double * result)

These functions compute the scalar product x^{T} y for the vectors x and y, returning the result in result.

Function int gsl blas cdotu (const gsl_vector_complex_float * x, const
         gsl_vector_complex_float * y, gsl_complex_float * dotu)
Function int gsl blas zdotu (const gsl_vector_complex * x, const
         gsl_vector_complex * y, gsl_complex * dotu)

These functions compute the complex scalar product x^{T} y for the vectors x and y, returning the result in result

Function int gsl blas cdotc (const gsl_vector_complex_float * x, const
         gsl_vector_complex_float * y, gsl_complex_float * dotc)
Function int gsl blas zdotc (const gsl_vector_complex * x, const
         gsl_vector_complex * y, gsl_complex * dotc)

These functions compute the complex conjugate scalar product x^{H}y for the vectors x and y, returning the result in result

Function float gsl blas snrm2 (const gsl_vector_float * x)

Function double gsl blas dnrm2 (const gsl_vector * x)

These functions compute the Euclidean norm ||x||_{2} = \sqrt{∑ x_{i}^{2}} of the vector x.

Function float gsl blas scnrm2 (const gsl_vector_complex_float * x)
Function double gsl blas dznrm2 (const gsl_vector_complex * x)

These functions compute the Euclidean norm of the complex vector x,

||x||_{2} = \sqrt{∑(Re(x_{i})^{2} + Im(x_{i})^{2})}.

Function float gsl blas sasum (const gsl_vector_float * x)

Function double gsl blas dasum (const gsl_vector * x)

These functions compute the absolute sum ∑ |x_{i}| of the elements of the vector x.

Function float gsl blas scasum (const gsl_vector_complex_float * x)
Function double gsl blas dzasum (const gsl_vector_complex * x)

These functions compute the absolute sum ∑ |Re(x_{i})| + |Im(x_{i})| of the elements of the vector x.

Function CBLAS_INDEX_t gsl blas isamax (const gsl_vector_float * x)
Function CBLAS_INDEX_t gsl blas idamax (const gsl_vector * x)
Function CBLAS_INDEX_t gsl blas icamax (const gsl_vector_complex_float

* x)

Function CBLAS_INDEX_t gsl blas izamax (const gsl_vector_complex * x)

These functions return the index of the largest element of the vector x. The largest element is determined by its absolute magnitude for real vector and by the sum of the magnitudes of the real and imaginary parts |Re(x_{i})| + |Im(x_{i})| for complex vectors. If the largest value occurs several times then the index of the first occurrence is returned.

Function int gsl blas sswap (gsl_vector_float * x, gsl_vector_float * y)
Function int gsl blas dswap (gsl_vector * x, gsl_vector * y)
Function int gsl blas cswap (gsl_vector_complex_float * x,
         gsl_vector_complex_float * y)
Function int gsl blas zswap (gsl_vector_complex * x, gsl_vector_complex

* y) These functions exchange the elements of the vectors x and y.

Function int gsl blas scopy (const gsl_vector_float * x, gsl_vector_float

* y)

Function int gsl blas dcopy (const gsl_vector * x, gsl_vector * y)
Function int gsl blas ccopy (const gsl_vector_complex_float * x,
         gsl_vector_complex_float * y)
Function int gsl blas zcopy (const gsl_vector_complex * x,
         gsl_vector_complex * y)

These functions copy the elements of the vector x into the vector y.

Function int gsl blas saxpy (float alpha, const gsl_vector_float * x,
         gsl_vector_float * y)
Function int gsl blas daxpy (double alpha, const gsl_vector * x,
         gsl_vector * y)
Function int gsl blas caxpy (const gsl_complex_float alpha, const
         gsl_vector_complex_float * x, gsl_vector_complex_float * y)

Function int gsl blas zaxpy (const gsl_complex alpha, const

         gsl_vector_complex * x, gsl_vector_complex * y)

These functions compute the sum y = αx + y for the vectors x and y.

Function void gsl blas sscal (float alpha, gsl_vector_float * x) Function void gsl blas dscal (double alpha, gsl_vector * x)

Function void gsl blas cscal (const gsl_complex_float alpha,
         gsl_vector_complex_float * x)
Function void gsl blas zscal (const gsl_complex alpha, gsl_vector_complex

* x)

Function void gsl blas csscal (float alpha, gsl_vector_complex_float * x)
Function void gsl blas zdscal (double alpha, gsl_vector_complex * x)

These functions rescale the vector x by the multiplicative factor alpha.

Function int gsl blas srotg (float a[], float b[], float c[], float s[])
Function int gsl blas drotg (double a[], double b[], double c[], double s[])

These functions compute a Givens rotation (c, s) which zeroes the vector (a, b),

( \begin{matrix}c & s \\ −s & c \end{matrix} ) ( \begin{matrix}a \\ b \end{matrix} ) = ( \begin{matrix}r′ \\ 0 \end{matrix} )

The variables a and b are overwritten by the routine.

Function int gsl blas srot (gsl_vector_float * x, gsl_vector_float * y,
         float c, float s)
Function int gsl blas drot (gsl_vector * x, gsl_vector * y, const double c,
         const double s)

These functions apply a Givens rotation (x′, y′) = (cx + sy, −sx + cy) to the vectors x, y.

Function int gsl blas srotmg (float d1[], float d2[], float b1[], float b2,
         float P[])
Function int gsl blas drotmg (double d1[], double d2[], double b1[], double
         b2, double P[])

These functions compute a modified Given’s transformation.

Function int gsl blas srotm (gsl_vector_float * x, gsl_vector_float * y,
const float P[])
Function int gsl blas drotm (gsl_vector * x, gsl_vector * y, const double

P[]) These functions apply a modified Given’s transformation.

11.1.2 Level 2

Function int gsl blas sgemv (CBLAS_TRANSPOSE_t TransA, float alpha,

const gsl_matrix_float * A, const gsl_vector_float * x, float beta,
gsl_vector_float * y)

Function int gsl blas dgemv (CBLAS_TRANSPOSE_t TransA, double alpha,

const gsl_matrix * A, const gsl_vector * x, double beta, gsl_vector *

y)

Function int gsl blas cgemv (CBLAS_TRANSPOSE_t TransA, const

gsl_complex_float alpha, const gsl_matrix_complex_float * A, const
gsl_vector_complex_float * x, const gsl_complex_float beta,
gsl_vector_complex_float * y)

Function int gsl blas zgemv (CBLAS_TRANSPOSE_t TransA, const

gsl_complex alpha, const gsl_matrix_complex * A, const
gsl_vector_complex * x, const gsl_complex beta, gsl_vector_complex *

y) These functions compute the matrix-vector product and sum y = αop(A)x+βy, where

op(A) = A, A^{T}, A^{H} for TransA = CblasNoTrans, CblasTrans, CblasConjTrans.
Function int gsl blas strmv (CBLAS_UPLO_t Uplo, CBLAS_TRANSPOSE_t
TransA, CBLAS_DIAG_t Diag, const gsl_matrix_float * A,
gsl_vector_float * x)
Function int gsl blas dtrmv (CBLAS_UPLO_t Uplo, CBLAS_TRANSPOSE_t
TransA, CBLAS_DIAG_t Diag, const gsl_matrix * A, gsl_vector * x)
Function int gsl blas ctrmv (CBLAS_UPLO_t Uplo, CBLAS_TRANSPOSE_t
TransA, CBLAS_DIAG_t Diag, const gsl_matrix_complex_float * A,
gsl_vector_complex_float * x)
Function int gsl blas ztrmv (CBLAS_UPLO_t Uplo, CBLAS_TRANSPOSE_t
TransA, CBLAS_DIAG_t Diag, const gsl_matrix_complex * A,
gsl_vector_complex * x)

These functions compute the matrix-vector product and sum y = αop(A)x + βy for the triangular matrix A, where op(A) = A, A^{T}, A^{H} for TransA = CblasNoTrans,

CblasTrans, CblasConjTrans. When Uplo is CblasUpper then the upper triangle of

A is used, and when Uplo is CblasLower then the lower triangle of A is used. If Diag is CblasNonUnit then the diagonal of the matrix is used, but if Diag is CblasUnit then the diagonal elements of the matrix A are taken as unity and are not referenced.

Function int gsl blas strsv (CBLAS_UPLO_t Uplo, CBLAS_TRANSPOSE_t TransA,
    CBLAS_DIAG_t Diag, const gsl_matrix_float * A, gsl_vector_float * x)
Function int gsl blas dtrsv (CBLAS_UPLO_t Uplo, CBLAS_TRANSPOSE_t
    TransA, CBLAS_DIAG_t Diag, const gsl_matrix * A, gsl_vector * x)
Function int gsl blas ctrsv (CBLAS_UPLO_t Uplo, CBLAS_TRANSPOSE_t TransA,
    CBLAS_DIAG_t Diag, const gsl_matrix_complex_float * A,
    gsl_vector_complex_float * x)
Function int gsl blas ztrsv (CBLAS_UPLO_t Uplo, CBLAS_TRANSPOSE_t TransA,
    CBLAS_DIAG_t Diag, const gsl_matrix_complex * A, gsl_vector_complex

*x) These functions compute inv(op(A))x for x, where op(A) = A, A^{T}, A^{H} for TransA =

CblasNoTrans, CblasTrans, CblasConjTrans. When Uplo is CblasUpper then the

upper triangle of A is used, and when Uplo is CblasLower then the lower triangle of A is used. If Diag is CblasNonUnit then the diagonal of the matrix is used, but if Diag is CblasUnit then the diagonal elements of the matrix A are taken as unity and are not referenced.

Function int gsl blas ssymv (CBLAS_UPLO_t Uplo, float alpha, const

    gsl_matrix_float * A, const gsl_vector_float * x, float beta,
    gsl_vector_float * y)

Function int gsl blas dsymv (CBLAS_UPLO_t Uplo, double alpha, const

    gsl_matrix * A, const gsl_vector * x, double beta, gsl_vector * y)

These functions compute the matrix-vector product and sum y = αAx + βy for the symmetric matrix A. Since the matrix A is symmetric only its upper half or lower half need to be stored. When Uplo is CblasUpper then the upper triangle and diagonal of A are used, and when Uplo is CblasLower then the lower triangle and diagonal of A are used.

Function int gsl blas chemv (CBLAS_UPLO_t Uplo, const gsl_complex_float
    alpha, const gsl_matrix_complex_float * A, const
    gsl_vector_complex_float * x, const gsl_complex_float beta,
    gsl_vector_complex_float * y)
Function int gsl blas zhemv (CBLAS_UPLO_t Uplo, const gsl_complex alpha,
    const gsl_matrix_complex * A, const gsl_vector_complex * x, const
    gsl_complex beta, gsl_vector_complex * y)

These functions compute the matrix-vector product and sum y = αAx + βy for the hermitian matrix A. Since the matrix A is hermitian only its upper half or lower half need to be stored. When Uplo is CblasUpper then the upper triangle and diagonal of A are used, and when Uplo is CblasLower then the lower triangle and diagonal of A are used. The imaginary elements of the diagonal are automatically assumed to be zero and are not referenced.

Function int gsl blas sger (float alpha, const gsl_vector_float * x, const
gsl_vector_float * y, gsl_matrix_float * A)
Function int gsl blas dger (double alpha, const gsl_vector * x, const
gsl_vector * y, gsl_matrix * A)
Function int gsl blas cgeru (const gsl_complex_float alpha, const
gsl_vector_complex_float * x, const gsl_vector_complex_float * y,
gsl_matrix_complex_float * A)

Function int gsl blas zgeru (const gsl_complex alpha, const

gsl_vector_complex * x, const gsl_vector_complex * y,
gsl_matrix_complex * A)

These functions compute the rank-1 update A = αxy^{T} + A of the matrix A.

Function int gsl blas cgerc (const gsl_complex_float alpha, const
gsl_vector_complex_float * x, const gsl_vector_complex_float * y,
gsl_matrix_complex_float * A)

Function int gsl blas zgerc (const gsl_complex alpha, const

gsl_vector_complex * x, const gsl_vector_complex * y,
gsl_matrix_complex * A)

These functions compute the conjugate rank-1 update A = αxy^{H} + A of the matrix A.

Function int gsl blas ssyr (CBLAS_UPLO_t Uplo, float alpha, const

gsl_vector_float * x, gsl_matrix_float * A)

Function int gsl blas dsyr (CBLAS_UPLO_t Uplo, double alpha, const

gsl_vector * x, gsl_matrix * A)

These functions compute the symmetric rank-1 update A = αxx^{T} +A of the symmetric matrix A. Since the matrix A is symmetric only its upper half or lower half need to be stored. When Uplo is CblasUpper then the upper triangle and diagonal of A are used, and when Uplo is CblasLower then the lower triangle and diagonal of A are used.

Function int gsl blas cher (CBLAS_UPLO_t Uplo, float alpha, const

gsl_vector_complex_float * x, gsl_matrix_complex_float * A)

Function int gsl blas zher (CBLAS_UPLO_t Uplo, double alpha, const

gsl_vector_complex * x, gsl_matrix_complex * A)

These functions compute the hermitian rank-1 update A = αxx^{H} +A of the hermitian matrix A. Since the matrix A is hermitian only its upper half or lower half need to be stored. When Uplo is CblasUpper then the upper triangle and diagonal of A are used, and when Uplo is CblasLower then the lower triangle and diagonal of A are used. The imaginary elements of the diagonal are automatically set to zero.

Function int gsl blas ssyr2 (CBLAS_UPLO_t Uplo, float alpha, const

gsl_vector_float * x, const gsl_vector_float * y, gsl_matrix_float *

A)

Function int gsl blas dsyr2 (CBLAS_UPLO_t Uplo, double alpha, const

gsl_vector * x, const gsl_vector * y, gsl_matrix * A)

These functions compute the symmetric rank-2 update A = αxy^{T} + αyx^{T} + A of the symmetric matrix A. Since the matrix A is symmetric only its upper half or lower half need to be stored. When Uplo is CblasUpper then the upper triangle and diagonal of A are used, and when Uplo is CblasLower then the lower triangle and diagonal of A are used.

Function int gsl blas cher2 (CBLAS_UPLO_t Uplo, const gsl_complex_float
alpha, const gsl_vector_complex_float * x, const
gsl_vector_complex_float * y, gsl_matrix_complex_float * A)
Function int gsl blas zher2 (CBLAS_UPLO_t Uplo, const gsl_complex alpha,
const gsl_vector_complex * x, const gsl_vector_complex * y,
gsl_matrix_complex * A)

These functions compute the hermitian rank-2 update A = αxy^{H} + α^{∗}yx^{H}A of the hermitian matrix A. Since the matrix A is hermitian only its upper half or lower half need to be stored. When Uplo is CblasUpper then the upper triangle and diagonal of A are used, and when Uplo is CblasLower then the lower triangle and diagonal of A are used. The imaginary elements of the diagonal are automatically set to zero.

11.1.3 Level 3

Function int gsl blas sgemm (CBLAS_TRANSPOSE_t TransA,

CBLAS_TRANSPOSE_t TransB, float alpha, const gsl_matrix_float * A,
const gsl_matrix_float * B, float beta, gsl_matrix_float * C)

Function int gsl blas dgemm (CBLAS_TRANSPOSE_t TransA,

CBLAS_TRANSPOSE_t TransB, double alpha, const gsl_matrix * A, const
gsl_matrix * B, double beta, gsl_matrix * C)

Function int gsl blas cgemm (CBLAS_TRANSPOSE_t TransA,

CBLAS_TRANSPOSE_t TransB, const gsl_complex_float alpha, const
gsl_matrix_complex_float * A, const gsl_matrix_complex_float * B,
const gsl_complex_float beta, gsl_matrix_complex_float * C)

Function int gsl blas zgemm (CBLAS_TRANSPOSE_t TransA,

CBLAS_TRANSPOSE_t TransB, const gsl_complex alpha, const
gsl_matrix_complex * A, const gsl_matrix_complex * B, const
gsl_complex beta, gsl_matrix_complex * C)

These functions compute the matrix-matrix product and sum C = αop(A)op(B)+βC where op(A) = A, A^{T}, A^{H} for TransA = CblasNoTrans, CblasTrans, CblasConjTrans and similarly for the parameter TransB.

Function int gsl blas ssymm (CBLAS_SIDE_t Side, CBLAS_UPLO_t Uplo, float
alpha, const gsl_matrix_float * A, const gsl_matrix_float * B, float
beta, gsl_matrix_float * C)
Function int gsl blas dsymm (CBLAS_SIDE_t Side, CBLAS_UPLO_t Uplo,
double alpha, const gsl_matrix * A, const gsl_matrix * B, double beta,
gsl_matrix * C)
Function int gsl blas csymm (CBLAS_SIDE_t Side, CBLAS_UPLO_t Uplo, const
gsl_complex_float alpha, const gsl_matrix_complex_float * A, const
gsl_matrix_complex_float * B, const gsl_complex_float beta,
gsl_matrix_complex_float * C)
Function int gsl blas zsymm (CBLAS_SIDE_t Side, CBLAS_UPLO_t Uplo, const
gsl_complex alpha, const gsl_matrix_complex * A, const
gsl_matrix_complex * B, const gsl_complex beta, gsl_matrix_complex *

C) These functions compute the matrix-matrix product and sum C = αAB + βC for Side is CblasLeft and C = αBA + βC for Side is CblasRight, where the matrix A is symmetric. When Uplo is CblasUpper then the upper triangle and diagonal of A are used, and when Uplo is CblasLower then the lower triangle and diagonal of A are used.

Function int gsl blas chemm (CBLAS_SIDE_t Side, CBLAS_UPLO_t Uplo, const
gsl_complex_float alpha, const gsl_matrix_complex_float * A, const
gsl_matrix_complex_float * B, const gsl_complex_float beta,
gsl_matrix_complex_float * C)
Function int gsl blas zhemm (CBLAS_SIDE_t Side, CBLAS_UPLO_t Uplo,
const gsl_complex alpha, const gsl_matrix_complex * A, const
gsl_matrix_complex * B, const gsl_complex beta, gsl_matrix_complex *

C) These functions compute the matrix-matrix product and sum C = αAB + βC for Side is CblasLeft and C = αBA + βC for Side is CblasRight, where the matrix A is hermitian. When Uplo is CblasUpper then the upper triangle and diagonal of A are used, and when Uplo is CblasLower then the lower triangle and diagonal of A are used. The imaginary elements of the diagonal are automatically set to zero.

Function int gsl blas strmm (CBLAS_SIDE_t Side, CBLAS_UPLO_t Uplo,
CBLAS_TRANSPOSE_t TransA, CBLAS_DIAG_t Diag, float alpha, const
gsl_matrix_float * A, gsl_matrix_float * B)
Function int gsl blas dtrmm (CBLAS_SIDE_t Side, CBLAS_UPLO_t Uplo,
CBLAS_TRANSPOSE_t TransA, CBLAS_DIAG_t Diag, double alpha, const
gsl_matrix * A, gsl_matrix * B)
Function int gsl blas ctrmm (CBLAS_SIDE_t Side, CBLAS_UPLO_t Uplo,
CBLAS_TRANSPOSE_t TransA, CBLAS_DIAG_t Diag, const gsl_complex_float
alpha, const gsl_matrix_complex_float * A, gsl_matrix_complex_float

* B)

Function int gsl blas ztrmm (CBLAS_SIDE_t Side, CBLAS_UPLO_t Uplo,
CBLAS_TRANSPOSE_t TransA, CBLAS_DIAG_t Diag, const gsl_complex alpha,
const gsl_matrix_complex * A, gsl_matrix_complex * B)

These functions compute the matrix-matrix product B = αop(A)B for Side is CblasLeft and B = αBop(A) for Side is CblasRight. The matrix A is triangular

and op(A) = A, A^{T}, A^{H} for TransA = CblasNoTrans, CblasTrans, CblasConjTrans

When Uplo is CblasUpper then the upper triangle of A is used, and when Uplo is CblasLower then the lower triangle of A is used. If Diag is CblasNonUnit then the diagonal of A is used, but if Diag is CblasUnit then the diagonal elements of the matrix A are taken as unity and are not referenced.

Function int gsl blas strsm (CBLAS_SIDE_t Side, CBLAS_UPLO_t Uplo,
    CBLAS_TRANSPOSE_t TransA, CBLAS_DIAG_t Diag, float alpha, const
    gsl_matrix_float * A, gsl_matrix_float * B)
Function int gsl blas dtrsm (CBLAS_SIDE_t Side, CBLAS_UPLO_t Uplo,
    CBLAS_TRANSPOSE_t TransA, CBLAS_DIAG_t Diag, double alpha, const
    gsl_matrix * A, gsl_matrix * B)
Function int gsl blas ctrsm (CBLAS_SIDE_t Side, CBLAS_UPLO_t Uplo,
    CBLAS_TRANSPOSE_t TransA, CBLAS_DIAG_t Diag, const gsl_complex_float
    alpha, const gsl_matrix_complex_float * A, gsl_matrix_complex_float

* B)

Function int gsl blas ztrsm (CBLAS_SIDE_t Side, CBLAS_UPLO_t Uplo,
    CBLAS_TRANSPOSE_t TransA, CBLAS_DIAG_t Diag, const gsl_complex alpha,
    const gsl_matrix_complex * A, gsl_matrix_complex * B)

These functions compute the matrix-matrix product B = αop(inv(A))B for Side is CblasLeft and B = αBop(inv(A)) for Side is CblasRight. The matrix A is triangular and op(A) = A, A^{T}, A^{H} for TransA = CblasNoTrans, CblasTrans, CblasConjTrans When Uplo is CblasUpper then the upper triangle of A is used, and when Uplo is CblasLower then the lower triangle of A is used. If Diag is CblasNonUnit then the diagonal of A is used, but if Diag is CblasUnit then the diagonal elements of the matrix A are taken as unity and are not referenced.

Function int gsl blas ssyrk (CBLAS_UPLO_t Uplo, CBLAS_TRANSPOSE_t Trans,
    float alpha, const gsl_matrix_float * A, float beta, gsl_matrix_float

* C)

Function int gsl blas dsyrk (CBLAS_UPLO_t Uplo, CBLAS_TRANSPOSE_t Trans,
    double alpha, const gsl_matrix * A, double beta, gsl_matrix * C)
Function int gsl blas csyrk (CBLAS_UPLO_t Uplo, CBLAS_TRANSPOSE_t Trans,
    const gsl_complex_float alpha, const gsl_matrix_complex_float * A,
    const gsl_complex_float beta, gsl_matrix_complex_float * C)
Function int gsl blas zsyrk (CBLAS_UPLO_t Uplo, CBLAS_TRANSPOSE_t Trans,
    const gsl_complex alpha, const gsl_matrix_complex * A, const
    gsl_complex beta, gsl_matrix_complex * C)

These functions compute a rank-k update of the symmetric matrix C, C = αAA^{T} +βC when Trans is CblasNoTrans and C = αA^{T} A + βC when Trans is CblasTrans. Since the matrix C is symmetric only its upper half or lower half need to be stored. When Uplo is CblasUpper then the upper triangle and diagonal of C are used, and when Uplo is CblasLower then the lower triangle and diagonal of C are used.

Function int gsl blas cherk (CBLAS_UPLO_t Uplo, CBLAS_TRANSPOSE_t Trans,
float alpha, const gsl_matrix_complex_float * A, float beta,
gsl_matrix_complex_float * C)
Function int gsl blas zherk (CBLAS_UPLO_t Uplo, CBLAS_TRANSPOSE_t Trans,
double alpha, const gsl_matrix_complex * A, double beta,
gsl_matrix_complex * C)

These functions compute a rank-k update of the hermitian matrix C, C = αAA^{H} +βC when Trans is CblasNoTrans and C = αA^{H}A+βC when Trans is CblasTrans. Since the matrix C is hermitian only its upper half or lower half need to be stored. When Uplo is CblasUpper then the upper triangle and diagonal of C are used, and when Uplo is CblasLower then the lower triangle and diagonal of C are used. The imaginary elements of the diagonal are automatically set to zero.

Function int gsl blas ssyr2k (CBLAS_UPLO_t Uplo, CBLAS_TRANSPOSE_t Trans,
float alpha, const gsl_matrix_float * A, const gsl_matrix_float * B,
float beta, gsl_matrix_float * C)
Function int gsl blas dsyr2k (CBLAS_UPLO_t Uplo, CBLAS_TRANSPOSE_t
Trans, double alpha, const gsl_matrix * A, const gsl_matrix * B,
double beta, gsl_matrix * C)
Function int gsl blas csyr2k (CBLAS_UPLO_t Uplo, CBLAS_TRANSPOSE_t Trans,
const gsl_complex_float alpha, const gsl_matrix_complex_float * A,
const gsl_matrix_complex_float * B, const gsl_complex_float beta,
gsl_matrix_complex_float * C)
Function int gsl blas zsyr2k (CBLAS_UPLO_t Uplo, CBLAS_TRANSPOSE_t Trans,
const gsl_complex alpha, const gsl_matrix_complex * A, const
gsl_matrix_complex * B, const gsl_complex beta, gsl_matrix_complex

*C) These functions compute a rank-2k update of the symmetric matrix C, C = αAB^{T} + αBA^{T} + βC when Trans is CblasNoTrans and C = αA^{T} B + αB^{T} A + βC when Trans is CblasTrans. Since the matrix C is symmetric only its upper half or lower half need to be stored. When Uplo is CblasUpper then the upper triangle and diagonal of C are used, and when Uplo is CblasLower then the lower triangle and diagonal of C are used.

Function int gsl blas cher2k (CBLAS_UPLO_t Uplo, CBLAS_TRANSPOSE_t Trans,
const gsl_complex_float alpha, const gsl_matrix_complex_float * A,
const gsl_matrix_complex_float * B, float beta,
gsl_matrix_complex_float * C)
Function int gsl blas zher2k (CBLAS_UPLO_t Uplo, CBLAS_TRANSPOSE_t
Trans, const gsl_complex alpha, const gsl_matrix_complex * A, const
gsl_matrix_complex * B, double beta, gsl_matrix_complex * C)

These functions compute a rank-2k update of the hermitian matrix C, C = αAB^{H} + α^{∗}BA^{H} + βC when Trans is CblasNoTrans and C = αA^{H}B + α^{∗}B^{H}A + βC when Trans is CblasTrans. Since the matrix C is hermitian only its upper half or lower half need to be stored. When Uplo is CblasUpper then the upper triangle and diagonal of C are used, and when Uplo is CblasLower then the lower triangle and diagonal of C are used. The imaginary elements of the diagonal are automatically set to zero.

11.2 Examples

The following program computes the product of two matrices using the Level-3 blas function dgemm,

( \begin{matrix}0.11 & 0.12 & 0.13 \\ 0.21 & 0.22 & 0.23 \end{matrix} ) ( \begin{matrix}1011 & 1012 \\ 1021 & 1022 \\ 1031 & 1031 \end{matrix} ) = ( \begin{matrix}367.76 & 368.12 \\ 674.06 & 674.72 \end{matrix} )

The matrices are stored in row major order, according to the C convention for arrays.

#include <stdio.h>
#include <gsl/gsl_blas.h>
int
main (void)
{
double a[] = { 0.11, 0.12, 0.13,
               0.21, 0.22, 0.23 };
double b[] = { 1011, 1012,
               1021, 1022,
               1031, 1032 };
double c[] = { 0.00, 0.00,
               0.00, 0.00 };
gsl_matrix_view A = gsl_matrix_view_array(a, 2, 3);
gsl_matrix_view B = gsl_matrix_view_array(b, 3, 2);
gsl_matrix_view C = gsl_matrix_view_array(c, 2, 2);
/* Compute C = A B */
gsl_blas_dgemm (CblasNoTrans, CblasNoTrans,
                1.0, &A.matrix, &B.matrix,
                0.0, &C.matrix);
printf("[ %g, %g\n", c[0], c[1]);
printf("  %g, %g ]\n", c[2], c[3]);
return 0;
}

Here is the output from the program,

$ ./a.out
[ 367.76, 368.12
674.06, 674.72 ]

11.3 References and Further Reading

Information on the blas standards, including both the legacy and draft interface standards, is available online from the blas Homepage and blas Technical Forum web-site.

BLAS Homepage http://www.netlib.org/blas/
BLAS Technical Forum http://www.netlib.org/cgi-bin/checkout/blast/blast.pl

The following papers contain the specifications for Level 1, Level 2 and Level 3 blas. C. Lawson, R. Hanson, D. Kincaid, F. Krogh, "Basic Linear Algebra Subprograms for Fortran Usage", ACM Transactions on Mathematical Software, Vol. 5 (1979), Pages 308-325. J.J. Dongarra, J. DuCroz, S. Hammarling, R. Hanson, "An Extended Set of Fortran Basic Linear Algebra Subprograms", ACM Transactions on Mathematical Software, Vol. 14, No. 1 (1988), Pages 1-32. J.J. Dongarra, I. Duff, J. DuCroz, S. Hammarling, "A Set of Level 3 Basic Linear Algebra Subprograms", ACM Transactions on Mathematical Software, Vol. 16 (1990), Pages 1-28.

Postscript versions of the latter two papers are available from http://www.netlib.org/blas/. A cblas wrapper for Fortran blas libraries is available from the same location.

12 Linear Algebra

This chapter describes functions for solving linear systems. The library provides simple linear algebra operations which operate directly on the gsl_vector and gsl_matrix objects. These are intended for use with "small" systems where simple algorithms are acceptable.

Anyone interested in large systems will want to use the sophisticated routines found in lapack. The Fortran version of lapack is recommended as the standard package for linear algebra. It supports blocked algorithms, specialized data representations and other optimizations.

The functions described in this chapter are declared in the header file ‘gsl_linalg.h’.

12.1 LU Decomposition

A general square matrix A has an LU decomposition into upper and lower triangular matrices,

P A = LU

where P is a permutation matrix, L is unit lower triangular matrix and U is upper triangular matrix. For square matrices this decomposition can be used to convert the linear system Ax = b into a pair of triangular systems (Ly = P b, Ux = y), which can be solved by forward and back-substitution.

Function int gsl linalg LU decomp (gsl_matrix * A, gsl_permutation * p,

int *signum)

Function int gsl linalg complex LU decomp (gsl_matrix_complex * A,

         gsl_permutation * p, int *signum)

These functions factorize the square matrix A into the LU decomposition P A = LU. On output the diagonal and upper triangular part of the input matrix A contain the matrix U. The lower triangular part of the input matrix (excluding the diagonal) contains L. The diagonal elements of L are unity, and are not stored. The permutation matrix P is encoded in the permutation p. The j-th column of the matrix P is given by the k-th column of the identity matrix, where k = p_{j} the j-th element of the permutation vector. The sign of the permutation is given by signum. It has the value (−1)^{n}, where n is the number of interchanges in the permutation. The algorithm used in the decomposition is Gaussian Elimination with partial pivoting (Golub & Van Loan, Matrix Computations, Algorithm 3.4.1).

Function int gsl linalg LU solve (const gsl_matrix * LU, const

         gsl_permutation * p, const gsl_vector * b, gsl_vector * x)

Function int gsl linalg complex LU solve (const gsl_matrix_complex *

         LU, const gsl_permutation * p, const gsl_vector_complex * b,
         gsl_vector_complex * x)

These functions solve the system Ax = b using the LU decomposition of A into (LU,

     p) given by gsl_linalg_LU_decomp or gsl_linalg_complex_LU_decomp.

Function int gsl linalg LU svx (const gsl_matrix * LU, const
    gsl_permutation * p, gsl_vector * x)

Function int gsl linalg complex LU svx (const gsl_matrix_complex * LU,

    const gsl_permutation * p, gsl_vector_complex * x)

These functions solve the system Ax = b in-place using the LU decomposition of A into (LU,p). On input x should contain the right-hand side b, which is replaced by the solution on output.

Function int gsl linalg LU refine (const gsl_matrix * A, const gsl_matrix
    * LU, const gsl_permutation * p, const gsl_vector * b, gsl_vector * x,
    gsl_vector * residual)

Function int gsl linalg complex LU refine (const gsl_matrix_complex *

    A, const gsl_matrix_complex * LU, const gsl_permutation * p, const
    gsl_vector_complex * b, gsl_vector_complex * x, gsl_vector_complex *

residual) These functions apply an iterative improvement to x, the solution of Ax = b, using the LU decomposition of A into (LU,p). The initial residual r = Ax − b is also computed and stored in residual.

Function int gsl linalg LU invert (const gsl_matrix * LU, const

    gsl_permutation * p, gsl_matrix * inverse)

Function int gsl complex linalg LU invert (const gsl_matrix_complex *

    LU, const gsl_permutation * p, gsl_matrix_complex * inverse)

These functions compute the inverse of a matrix A from its LU decomposition (LU,p), storing the result in the matrix inverse. The inverse is computed by solving the system Ax = b for each column of the identity matrix.

Function double gsl linalg LU det (gsl_matrix * LU, int signum)

Function gsl_complex gsl linalg complex LU det (gsl_matrix_complex *

LU, int signum) These functions compute the determinant of a matrix A from its LU decomposition, LU. The determinant is computed as the product of the diagonal elements of U and the sign of the row permutation signum.

Function double gsl linalg LU lndet (gsl_matrix * LU) Function double gsl linalg complex LU lndet (gsl_matrix_complex * LU)

These functions compute the logarithm of the absolute value of the determinant of a matrix A, ln |det(A)|, from its LU decomposition, LU. This function may be useful if the direct computation of the determinant would overflow or underflow.

Function int gsl linalg LU sgndet (gsl_matrix * LU, int signum)

Function gsl_complex gsl linalg complex LU sgndet

    (gsl_matrix_complex * LU, int signum)

These functions compute the sign or phase factor of the determinant of a matrix A, det(A)/|det(A)|, from its LU decomposition, LU.

12.2 QR Decomposition

A general rectangular M-by-N matrix A has a QR decomposition into the product of an orthogonal M-by-M square matrix Q (where Q^{T} Q = I) and an M-by-N right-triangular matrix R,

A = QR

This decomposition can be used to convert the linear system Ax = b into the triangular system Rx = Q^{T} b, which can be solved by back-substitution. Another use of the QR decomposition is to compute an orthonormal basis for a set of vectors. The first N columns of Q form an orthonormal basis for the range of A, ran(A), when A has full column rank.

Function int gsl linalg QR decomp (gsl_matrix * A, gsl_vector * tau)

This function factorizes the M-by-N matrix A into the QR decomposition A = QR. On output the diagonal and upper triangular part of the input matrix contain the matrix R. The vector tau and the columns of the lower triangular part of the matrix A contain the Householder coefficients and Householder vectors which encode the orthogonal matrix Q. The vector tau must be of length k = min(M, N). The matrix Q is related to these components by, Q = Q_{k}...Q_{2}Q_{1} where Q_{i} = I − τ_{i}v_{i}v_{i}^{T} and v_{i} is the Householder vector v_{i} = (0, ..., 1, A(i + 1, i), A(i + 2, i), ..., A(m, i)). This is the same storage scheme as used by lapack. The algorithm used to perform the decomposition is Householder QR (Golub & Van Loan, Matrix Computations, Algorithm 5.2.1).

Function int gsl linalg QR solve (const gsl_matrix * QR, const

    gsl_vector * tau, const gsl_vector * b, gsl_vector * x)

This function solves the system Ax = b using the QR decomposition of A into (QR,

tau) given by gsl_linalg_QR_decomp.
Function int gsl linalg QR svx (const gsl_matrix * QR, const gsl_vector
    * tau, gsl_vector * x)

This function solves the system Ax = b in-place using the QR decomposition of A into (QR,tau) given by gsl_linalg_QR_decomp. On input x should contain the right-hand side b, which is replaced by the solution on output.

Function int gsl linalg QR lssolve (const gsl_matrix * QR, const

    gsl_vector * tau, const gsl_vector * b, gsl_vector * x, gsl_vector *

residual) This function finds the least squares solution to the overdetermined system Ax = b where the matrix A has more rows than columns. The least squares solution minimizes the Euclidean norm of the residual, ||Ax−b||.The routine uses the QR decomposition of A into (QR, tau) given by gsl_linalg_QR_decomp. The solution is returned in x. The residual is computed as a by-product and stored in residual.

Function int gsl linalg QR QTvec (const gsl_matrix * QR, const

    gsl_vector * tau, gsl_vector * v)

This function applies the matrix Q^{T} encoded in the decomposition (QR,tau) to the vector v, storing the result Q^{T} v in v. The matrix multiplication is carried out directly using the encoding of the Householder vectors without needing to form the full matrix Q^{T}.

Function int gsl linalg QR Qvec (const gsl_matrix * QR, const

    gsl_vector * tau, gsl_vector * v)

This function applies the matrix Q encoded in the decomposition (QR,tau) to the vector v, storing the result Qv in v. The matrix multiplication is carried out directly using the encoding of the Householder vectors without needing to form the full matrix Q.

Function int gsl linalg QR Rsolve (const gsl_matrix * QR, const

    gsl_vector * b, gsl_vector * x)

This function solves the triangular system Rx = b for x. It may be useful if the product b′ = Q^{T} b has already been computed using gsl_linalg_QR_QTvec.

Function int gsl linalg QR Rsvx (const gsl_matrix * QR, gsl_vector * x)

This function solves the triangular system Rx = b for x in-place. On input x should contain the right-hand side b and is replaced by the solution on output. This function may be useful if the product b′ = Q^{T} b has already been computed using gsl_linalg_

QR_QTvec.

Function int gsl linalg QR unpack (const gsl_matrix * QR, const

    gsl_vector * tau, gsl_matrix * Q, gsl_matrix * R)

This function unpacks the encoded QR decomposition (QR,tau) into the matrices Q and R, where Q is M-by-M and R is M-by-N.

Function int gsl linalg QR QRsolve (gsl_matrix * Q, gsl_matrix * R,

    const gsl_vector * b, gsl_vector * x)

This function solves the system Rx = Q^{T} b for x. It can be used when the QR decomposition of a matrix is available in unpacked form as (Q,R).

Function int gsl linalg QR update (gsl_matrix * Q, gsl_matrix * R,

    gsl_vector * w, const gsl_vector * v)

This function performs a rank-1 update wv^{T} of the QR decomposition (Q, R). The update is given by Q′R′ = QR + wv^{T} where the output matrices Q′ and R′ are also orthogonal and right triangular. Note that w is destroyed by the update.

Function int gsl linalg R solve (const gsl_matrix * R, const gsl_vector *
    b, gsl_vector * x)

This function solves the triangular system Rx = b for the N-by-N matrix R.

Function int gsl linalg R svx (const gsl_matrix * R, gsl_vector * x)

This function solves the triangular system Rx = b in-place. On input x should contain the right-hand side b, which is replaced by the solution on output.

12.3 QR Decomposition with Column Pivoting

The QR decomposition can be extended to the rank deficient case by introducing a column permutation P,

AP = QR

The first r columns of this Q form an orthonormal basis for the range of A for a matrix with column rank r. This decomposition can also be used to convert the linear system Ax = b into the triangular system Ry = Q^{T} b, x = P y, which can be solved by back-substitution and permutation. We denote the QR decomposition with column pivoting by QRP^{T} since A = QRP^{T}.

Function int gsl linalg QRPT decomp (gsl_matrix * A, gsl_vector *

    tau, gsl_permutation * p, int *signum, gsl_vector * norm)

This function factorizes the M-by-N matrix A into the QRP^{T} decomposition A = QRP^{T}. On output the diagonal and upper triangular part of the input matrix contain the matrix R. The permutation matrix P is stored in the permutation p. The sign of the permutation is given by signum. It has the value (−1)^{n}, where n is the number of interchanges in the permutation. The vector tau and the columns of the lower triangular part of the matrix A contain the Householder coefficients and vectors which encode the orthogonal matrix Q. The vector tau must be of length k = min(M, N). The matrix Q is related to these components by, Q = Q_{k}...Q_{2}Q_{1} where Q_{i} = I−τ_{i}v_{i}v_{i}^{T} and v_{i} is the Householder vector v_{i} = (0, ..., 1, A(i + 1, i), A(i + 2, i), ..., A(m, i)). This is the same storage scheme as used by lapack. On output the norms of each column of R are stored in the vector norm. The algorithm used to perform the decomposition is Householder QR with column pivoting (Golub & Van Loan, Matrix Computations, Algorithm 5.4.1).

Function int gsl linalg QRPT decomp2 (const gsl_matrix * A,

    gsl_matrix * q, gsl_matrix * r, gsl_vector * tau, gsl_permutation * p,
    int *signum, gsl_vector * norm)

This function factorizes the matrix A into the decomposition A = QRP^{T} without modifying A itself and storing the output in the separate matrices q and r.

Function int gsl linalg QRPT solve (const gsl_matrix * QR, const

    gsl_vector * tau, const gsl_permutation * p, const gsl_vector * b,
    gsl_vector * x)

This function solves the system Ax = b using the QRP^{T} decomposition of A into

(QR, tau, p) given by gsl_linalg_QRPT_decomp.

Function int gsl linalg QRPT svx (const gsl_matrix * QR, const

    gsl_vector * tau, const gsl_permutation * p, gsl_vector * x)

This function solves the system Ax = b in-place using the QRP^{T} decomposition of A into (QR,tau,p). On input x should contain the right-hand side b, which is replaced by the solution on output.

Function int gsl linalg QRPT QRsolve (const gsl_matrix * Q, const

    gsl_matrix * R, const gsl_permutation * p, const gsl_vector * b,
    gsl_vector * x)

This function solves the system RP^{T} x = Q^{T} b for x. It can be used when the QR decomposition of a matrix is available in unpacked form as (Q,R).

Function int gsl linalg QRPT update (gsl_matrix * Q, gsl_matrix * R,

    const gsl_permutation * p, gsl_vector * u, const gsl_vector * v)

This function performs a rank-1 update wv^{T} of the QRP^{T} decomposition (Q, R,p). The update is given by Q′R′ = QR + wv^{T} where the output matrices Q′ and R′ are also orthogonal and right triangular. Note that w is destroyed by the update. The permutation p is not changed.

Function int gsl linalg QRPT Rsolve (const gsl_matrix * QR, const

    gsl_permutation * p, const gsl_vector * b, gsl_vector * x)

This function solves the triangular system RP^{T} x = b for the N-by-N matrix R contained in QR.

Function int gsl linalg QRPT Rsvx (const gsl_matrix * QR, const

    gsl_permutation * p, gsl_vector * x)

This function solves the triangular system RP^{T} x = b in-place for the N-by-N matrix R contained in QR. On input x should contain the right-hand side b, which is replaced by the solution on output.

12.4 Singular Value Decomposition

A general rectangular M-by-N matrix A has a singular value decomposition (svd) into the product of an M-by-N orthogonal matrix U, an N-by-N diagonal matrix of singular values S and the transpose of an M-by-M orthogonal square matrix V,

A = USV^{T}

The singular values σ_{i} = S_{ii} are all non-negative and are generally chosen to form a nonincreasing sequence σ_{1} ≥ σ_{2} ≥ ... ≥ σ_{N} ≥ 0.

The singular value decomposition of a matrix has many practical uses. The condition number of the matrix is given by the ratio of the largest singular value to the smallest singular value. The presence of a zero singular value indicates that the matrix is singular. The number of non-zero singular values indicates the rank of the matrix. In practice singular value decomposition of a rank-deficient matrix will not produce exact zeroes for singular values, due to finite numerical precision. Small singular values should be edited by choosing a suitable tolerance.

Function int gsl linalg SV decomp (gsl_matrix * A, gsl_matrix * V,

    gsl_vector * S, gsl_vector * work)

This function factorizes the M-by-N matrix A into the singular value decomposition A = USV^{T}. On output the matrix A is replaced by U. The diagonal elements of the singular value matrix S are stored in the vector S. The singular values are nonnegative and form a non-increasing sequence from S_{1} to S_{N}. The matrix V contains the elements of V in untransposed form. To form the product USV^{T} it is necessary to take the transpose of V. A workspace of length N is required in work. This routine uses the Golub-Reinsch SVD algorithm.

Function int gsl linalg SV decomp mod (gsl_matrix * A, gsl_matrix *

    X, gsl_matrix * V, gsl_vector * S, gsl_vector * work)

This function computes the SVD using the modified Golub-Reinsch algorithm, which is faster for M >> N. It requires the vector work and the N-by-N matrix X as additional working space.

Function int gsl linalg SV decomp jacobi (gsl_matrix * A, gsl_matrix *

    V, gsl_vector * S)

This function computes the SVD using one-sided Jacobi orthogonalization (see references for details). The Jacobi method can compute singular values to higher relative accuracy than Golub-Reinsch algorithms.

Function int gsl linalg SV solve (gsl_matrix * U, gsl_matrix * V,
    gsl_vector * S, const gsl_vector * b, gsl_vector * x)

This function solves the system Ax = b using the singular value decomposition (U, S,

V) of A given by gsl_linalg_SV_decomp.

Only non-zero singular values are used in computing the solution. The parts of the solution corresponding to singular values of zero are ignored. Other singular values can be edited out by setting them to zero before calling this function. In the over-determined case where A has more rows than columns the system is solved in the least squares sense, returning the solution x which minimizes ||Ax − b||_{2}.

12.5 Cholesky Decomposition

A symmetric, positive definite square matrix A has a Cholesky decomposition into a product of a lower triangular matrix L and its transpose L^{T},

A = LL^{T}

This is sometimes referred to as taking the square-root of a matrix. The Cholesky decomposition can only be carried out when all the eigenvalues of the matrix are positive. This decomposition can be used to convert the linear system Ax = b into a pair of triangular systems (Ly = b, L^{T} x = y), which can be solved by forward and back-substitution.

Function int gsl linalg cholesky decomp (gsl_matrix * A)

This function factorizes the positive-definite square matrix A into the Cholesky decomposition A = LL^{T}. On output the diagonal and lower triangular part of the input matrix A contain the matrix L. The upper triangular part of the input matrix contains L^{T}, the diagonal terms being identical for both L and L^{T}. If the matrix is not positive-definite then the decomposition will fail, returning the error code GSL_EDOM.

Function int gsl linalg cholesky solve (const gsl_matrix * cholesky, const

    gsl_vector * b, gsl_vector * x)

This function solves the system Ax = b using the Cholesky decomposition of A into

the matrix cholesky given by gsl_linalg_cholesky_decomp.

Function int gsl linalg cholesky svx (const gsl_matrix * cholesky,

    gsl_vector * x)

This function solves the system Ax = b in-place using the Cholesky decomposition of A into the matrix cholesky given by gsl_linalg_cholesky_decomp. On input x should contain the right-hand side b, which is replaced by the solution on output.

12.6 Tridiagonal Decomposition of Real Symmetric Matrices

A symmetric matrix A can be factorized by similarity transformations into the form,

A = QT Q^{T}

where Q is an orthogonal matrix and T is a symmetric tridiagonal matrix.

Function int gsl linalg symmtd decomp (gsl_matrix * A, gsl_vector *

tau) This function factorizes the symmetric square matrix A into the symmetric tridiagonal decomposition QT Q^{T}. On output the diagonal and subdiagonal part of the input matrix A contain the tridiagonal matrix T. The remaining lower triangular part of the input matrix contains the Householder vectors which, together with the Householder coefficients tau, encode the orthogonal matrix Q. This storage scheme is the same as used by lapack. The upper triangular part of A is not referenced.

Function int gsl linalg symmtd unpack (const gsl_matrix * A, const

    gsl_vector * tau, gsl_matrix * Q, gsl_vector * d, gsl_vector * sd)

This function unpacks the encoded symmetric tridiagonal decomposition (A, tau) obtained from gsl_linalg_symmtd_decomp into the orthogonal matrix Q, the vector of diagonal elements d and the vector of subdiagonal elements sd.

Function int gsl linalg symmtd unpack dsd (const gsl_matrix * A,

    gsl_vector * d, gsl_vector * sd)

This function unpacks the diagonal and subdiagonal of the encoded symmetric tridiagonal decomposition (A, tau) obtained from gsl_linalg_symmtd_decomp into the vectors d and sd.

12.7 Tridiagonal Decomposition of Hermitian Matrices

A hermitian matrix A can be factorized by similarity transformations into the form,

A = UT U^{T}

where U is an unitary matrix and T is a real symmetric tridiagonal matrix.

Function int gsl linalg hermtd decomp (gsl_matrix_complex * A,

    gsl_vector_complex * tau)

This function factorizes the hermitian matrix A into the symmetric tridiagonal decomposition UT U^{T}. On output the real parts of the diagonal and subdiagonal part of the input matrix A contain the tridiagonal matrix T. The remaining lower triangular part of the input matrix contains the Householder vectors which, together with the Householder coefficients tau, encode the orthogonal matrix Q. This storage scheme is the same as used by lapack. The upper triangular part of A and imaginary parts of the diagonal are not referenced.

Function int gsl linalg hermtd unpack (const gsl_matrix_complex * A,

    const gsl_vector_complex * tau, gsl_matrix_complex * Q, gsl_vector *
    d, gsl_vector * sd)

This function unpacks the encoded tridiagonal decomposition (A, tau) obtained from gsl_linalg_hermtd_decomp into the unitary matrix U, the real vector of diagonal elements d and the real vector of subdiagonal elements sd.

Function int gsl linalg hermtd unpack dsd (const gsl_matrix_complex *

    A, gsl_vector * d, gsl_vector * sd)

This function unpacks the diagonal and subdiagonal of the encoded tridiagonal decomposition (A, tau) obtained from gsl_linalg_hermtd_decomp into the real vectors d and sd.

12.8 Bidiagonalization

A general matrix A can be factorized by similarity transformations into the form,

A = UBV^{T}

where U and V are orthogonal matrices and B is a N-by-N bidiagonal matrix with non-zero entries only on the diagonal and superdiagonal. The size of U is M-by-N and the size of V is N-by-N.

Function int gsl linalg bidiag decomp (gsl_matrix * A, gsl_vector *

    tau U, gsl_vector * tau V)

This function factorizes the M-by-N matrix A into bidiagonal form UBV^{T}. The diagonal and superdiagonal of the matrix B are stored in the diagonal and superdiagonal of A. The orthogonal matrices U and V are stored as compressed Householder vectors in the remaining elements of A. The Householder coefficients are stored in the vectors tau U and tau V. The length of tau U must equal the number of elements in the diagonal of A and the length of tau V should be one element shorter.

Function int gsl linalg bidiag unpack (const gsl_matrix * A, const

    gsl_vector * tau U, gsl_matrix * U, const gsl_vector * tau V,
    gsl_matrix * V, gsl_vector * diag, gsl_vector * superdiag)

This function unpacks the bidiagonal decomposition of A given by gsl_linalg_ bidiag_decomp, (A, tau U, tau V) into the separate orthogonal matrices U, V and the diagonal vector diag and superdiagonal superdiag.

Function int gsl linalg bidiag unpack2 (gsl_matrix * A, gsl_vector *

    tau U, gsl_vector * tau V, gsl_matrix * V)

This function unpacks the bidiagonal decomposition of A given by gsl_linalg_ bidiag_decomp, (A, tau U, tau V) into the separate orthogonal matrices U, V and the diagonal vector diag and superdiagonal superdiag. The matrix U is stored in-place in A.

Function int gsl linalg bidiag unpack B (const gsl_matrix * A,

    gsl_vector * diag, gsl_vector * superdiag)

This function unpacks the diagonal and superdiagonal of the bidiagonal decomposition of A given by gsl_linalg_bidiag_decomp, into the diagonal vector diag and superdiagonal vector superdiag.

12.9 Householder solver for linear systems

Function int gsl linalg HH solve (gsl_matrix * A, const gsl_vector * b,
    gsl_vector * x)

This function solves the system Ax = b directly using Householder transformations. On output the solution is stored in x and b is not modified. The matrix A is destroyed by the Householder transformations.

Function int gsl linalg HH svx (gsl_matrix * A, gsl_vector * x)

This function solves the system Ax = b in-place using Householder transformations. On input x should contain the right-hand side b, which is replaced by the solution on output. The matrix A is destroyed by the Householder transformations.

12.10 Tridiagonal Systems

Function int gsl linalg solve symm tridiag (const gsl_vector * diag,

    const gsl_vector * e, const gsl_vector * b, gsl_vector * x)

This function solves the general N-by-N system Ax = b where A is symmetric tridiagonal. The form of A for the 4-by-4 case is shown below,

A = ( \begin{matrix}d_{0} & e_{0} \\ e_{0} & d_{1} & e_{1} \\ & e_{1} & d_{2} & e_{2} \\ & & e_{2} & d_{3} \end{matrix} )

Function int gsl linalg solve symm cyc tridiag (const gsl_vector * diag,

    const gsl_vector * e, const gsl_vector * b, gsl_vector * x)

This function solves the general N-by-N system Ax = b where A is symmetric cyclic tridiagonal. The form of A for the 4-by-4 case is shown below,

A = ( \begin{matrix}d_{0} & e_{0} & & e_{3} \\ e_{0} & d_{1} & e_{1} \\ & e_{1} & d_{2} & e_{2} \\ e_{3} & & e_{2} & d_{3} \end{matrix} )

12.11 Examples

The following program solves the linear system Ax = b. The system to be solved is,

( \begin{matrix}0.18 & 0.60 & 0.57 & 0.96 \\ 0.41 & 0.24 & 0.99 & 0.58 \\ 0.14 & 0.30 & 0.97 & 0.66 \\ 0.51 & 0.13 & 0.19 & 0.85 \end{matrix} ) ( \begin{matrix}x_{0} \\ x_{1} \\ x_{2} \\ x_{3} \end{matrix} ) = ( \begin{matrix}1.0 \\ 2.0 \\ 3.0 \\ 4.0 \end{matrix} )

and the solution is found using LU decomposition of the matrix A.

#include <stdio.h>
#include <gsl/gsl_linalg.h>
int
main (void)
{
double a_data[] = { 0.18, 0.60, 0.57, 0.96,
                    0.41, 0.24, 0.99, 0.58,
                    0.14, 0.30, 0.97, 0.66,
                    0.51, 0.13, 0.19, 0.85 };
double b_data[] = { 1.0, 2.0, 3.0, 4.0 };
gsl_matrix_view m
  = gsl_matrix_view_array(a_data, 4, 4);
gsl_vector_view b
  = gsl_vector_view_array(b_data, 4);
gsl_vector *x = gsl_vector_alloc (4);
int s;
gsl_permutation * p = gsl_permutation_alloc (4);
gsl_linalg_LU_decomp (&m.matrix, p, &s);
gsl_linalg_LU_solve (&m.matrix, p, &b.vector, x);
printf ("x = \n");
gsl_vector_fprintf(stdout, x, "%g");
gsl_permutation_free (p);
return 0;
}

Here is the output from the program,

x = -4.05205
-12.6056
1.66091
8.69377

This can be verified by multiplying the solution x by the original matrix A using gnu octave,

 octave> A = [ 0.18, 0.60, 0.57, 0.96;
               0.41, 0.24, 0.99, 0.58;
               0.14, 0.30, 0.97, 0.66;
               0.51, 0.13, 0.19, 0.85 ];
 octave> x = [ -4.05205; -12.6056; 1.66091; 8.69377];
 octave> A * x
 ans =
   1.0000
   2.0000
   3.0000
   4.0000

This reproduces the original right-hand side vector, b, in accordance with the equation Ax = b.

12.12 References and Further Reading

Further information on the algorithms described in this section can be found in the following book,

G. H. Golub, C. F. Van Loan, Matrix Computations (3rd Ed, 1996), Johns Hopkins University Press, ISBN 0-8018-5414-8.

The lapack library is described in, LAPACK Users’ Guide (Third Edition, 1999), Published by SIAM, ISBN 0-89871-447-8.

http://www.netlib.org/lapack

The lapack source code can be found at the website above, along with an online copy of the users guide. The Modified Golub-Reinsch algorithm is described in the following paper,

T.F. Chan, "An Improved Algorithm for Computing the Singular Value Decomposition", ACM Transactions on Mathematical Software, 8 (1982), pp 72–83.

The Jacobi algorithm for singular value decomposition is described in the following papers, J.C.Nash, "A one-sided transformation method for the singular value decomposition and algebraic eigenproblem", Computer Journal, Volume 18, Number 1 (1973), p 74— 76 James Demmel, Kresimir Veselic, "Jacobi’s Method is more accurate than QR", Lapack Working Note 15 (LAWN-15), October 1989. Available from netlib,

http://www.netlib.org/lapack/ in the lawns or lawnspdf directories.

13 Eigensystems

This chapter describes functions for computing eigenvalues and eigenvectors of matrices. There are routines for real symmetric and complex hermitian matrices, and eigenvalues can be computed with or without eigenvectors. The algorithms used are symmetric bidiagonalization followed by QR reduction.

These routines are intended for "small" systems where simple algorithms are acceptable. Anyone interested finding eigenvalues and eigenvectors of large matrices will want to use the sophisticated routines found in lapack. The Fortran version of lapack is recommended as the standard package for linear algebra.

The functions described in this chapter are declared in the header file ‘gsl_eigen.h’.

13.1 Real Symmetric Matrices

Function gsl_eigen_symm_workspace * gsl eigen symm alloc (const
    size_t n)

This function allocates a workspace for computing eigenvalues of n-by-n real symmetric matrices. The size of the workspace is O(2n).

Function void gsl eigen symm free (gsl_eigen_symm_workspace * w)

This function frees the memory associated with the workspace w.

Function int gsl eigen symm (gsl_matrix * A, gsl_vector * eval,
    gsl_eigen_symm_workspace * w)

This function computes the eigenvalues of the real symmetric matrix A. Additional workspace of the appropriate size must be provided in w. The diagonal and lower triangular part of A are destroyed during the computation, but the strict upper triangular part is not referenced. The eigenvalues are stored in the vector eval and are unordered.

Function gsl_eigen_symmv_workspace * gsl eigen symmv alloc (const
    size_t n)

This function allocates a workspace for computing eigenvalues and eigenvectors of n-by-n real symmetric matrices. The size of the workspace is O(4n).

Function void gsl eigen symmv free (gsl_eigen_symmv_workspace * w)

This function frees the memory associated with the workspace w.

Function int gsl eigen symmv (gsl_matrix * A, gsl_vector * eval,
    gsl_matrix * evec, gsl_eigen_symmv_workspace * w)

This function computes the eigenvalues and eigenvectors of the real symmetric matrix A. Additional workspace of the appropriate size must be provided in w. The diagonal and lower triangular part of A are destroyed during the computation, but the strict upper triangular part is not referenced. The eigenvalues are stored in the vector eval and are unordered. The corresponding eigenvectors are stored in the columns of the matrix evec. For example, the eigenvector in the first column corresponds to the first eigenvalue. The eigenvectors are guaranteed to be mutually orthogonal and normalised to unit magnitude.

13.2 Complex Hermitian Matrices

Function gsl_eigen_herm_workspace * gsl eigen herm alloc (const
    size_t n)

This function allocates a workspace for computing eigenvalues of n-by-n complex hermitian matrices. The size of the workspace is O(3n).

Function void gsl eigen herm free (gsl_eigen_herm_workspace * w)

This function frees the memory associated with the workspace w.

Function int gsl eigen herm (gsl_matrix_complex * A, gsl_vector * eval,
    gsl_eigen_herm_workspace * w)

This function computes the eigenvalues of the complex hermitian matrix A. Additional workspace of the appropriate size must be provided in w. The diagonal and lower triangular part of A are destroyed during the computation, but the strict upper triangular part is not referenced. The imaginary parts of the diagonal are assumed to be zero and are not referenced. The eigenvalues are stored in the vector eval and are unordered.

Function gsl_eigen_hermv_workspace * gsl eigen hermv alloc (const
    size_t n)

This function allocates a workspace for computing eigenvalues and eigenvectors of n-by-n complex hermitian matrices. The size of the workspace is O(5n).

Function void gsl eigen hermv free (gsl_eigen_hermv_workspace * w)

This function frees the memory associated with the workspace w.

Function int gsl eigen hermv (gsl_matrix_complex * A, gsl_vector * eval,
    gsl_matrix_complex * evec, gsl_eigen_hermv_workspace * w)

This function computes the eigenvalues and eigenvectors of the complex hermitian matrix A. Additional workspace of the appropriate size must be provided in w. The diagonal and lower triangular part of A are destroyed during the computation, but the strict upper triangular part is not referenced. The imaginary parts of the diagonal are assumed to be zero and are not referenced. The eigenvalues are stored in the vector eval and are unordered. The corresponding complex eigenvectors are stored in the columns of the matrix evec. For example, the eigenvector in the first column corresponds to the first eigenvalue. The eigenvectors are guaranteed to be mutually orthogonal and normalised to unit magnitude.

13.3 Sorting Eigenvalues and Eigenvectors

Function int gsl eigen symmv sort (gsl_vector * eval, gsl_matrix * evec,

    gsl_eigen_sort_t sort type)

This function simultaneously sorts the eigenvalues stored in the vector eval and the corresponding real eigenvectors stored in the columns of the matrix evec into ascending or descending order according to the value of the parameter sort type,

GSL_EIGEN_SORT_VAL_ASC

ascending order in numerical value

GSL_EIGEN_SORT_VAL_DESC

descending order in numerical value

GSL_EIGEN_SORT_ABS_ASC

ascending order in magnitude

GSL_EIGEN_SORT_ABS_DESC

descending order in magnitude

Function int gsl eigen hermv sort (gsl_vector * eval, gsl_matrix_complex
    * evec, gsl_eigen_sort_t sort type)

This function simultaneously sorts the eigenvalues stored in the vector eval and the corresponding complex eigenvectors stored in the columns of the matrix evec into ascending or descending order according to the value of the parameter sort type as shown above.

13.4 Examples

The following program computes the eigenvalues and eigenvectors of the 4-th order Hilbert matrix, H(i, j) = 1/(i + j + 1).

#include <stdio.h>
#include <gsl/gsl_math.h>
#include <gsl/gsl_eigen.h>
int
main (void)
{
  double data[] = { 1.0  , 1/2.0, 1/3.0, 1/4.0,
                    1/2.0, 1/3.0, 1/4.0, 1/5.0,
                    1/3.0, 1/4.0, 1/5.0, 1/6.0,
                    1/4.0, 1/5.0, 1/6.0, 1/7.0 };
  gsl_matrix_view m
    = gsl_matrix_view_array(data, 4, 4);
  gsl_vector *eval = gsl_vector_alloc (4);
  gsl_matrix *evec = gsl_matrix_alloc (4, 4);
  gsl_eigen_symmv_workspace * w =
    gsl_eigen_symmv_alloc (4);

  gsl_eigen_symmv (&m.matrix, eval, evec, w);
  gsl_eigen_symmv_free(w);
  gsl_eigen_symmv_sort (eval, evec,
                        GSL_EIGEN_SORT_ABS_ASC);
  {
    int i;
    for (i = 0; i < 4; i++)
      {
        double eval_i
           = gsl_vector_get(eval, i);
        gsl_vector_view evec_i
           = gsl_matrix_column(evec, i);
        printf("eigenvalue = %g\n", eval_i);
        printf("eigenvector = \n");
        gsl_vector_fprintf(stdout,
                           &evec_i.vector, "%g");
      }
  }
  return 0;
}

Here is the beginning of the output from the program,

$ ./a.out
eigenvalue = 9.67023e-05
eigenvector =
-0.0291933
0.328712
-0.791411
0.514553
...

This can be compared with the corresponding output from gnu octave,

octave> [v,d] = eig(hilb(4));
octave> diag(d)
ans =
   9.6702e-05
   6.7383e-03
   1.6914e-01
   1.5002e+00
octave> v
v =
   0.029193   0.179186  -0.582076   0.792608

       -0.328712  -0.741918   0.370502   0.451923
        0.791411   0.100228   0.509579   0.322416
       -0.514553   0.638283   0.514048   0.252161

Note that the eigenvectors can differ by a change of sign, since the sign of an eigenvector is arbitrary.

13.5 References and Further Reading

Further information on the algorithms described in this section can be found in the following book,

G. H. Golub, C. F. Van Loan, Matrix Computations (3rd Ed, 1996), Johns Hopkins University Press, ISBN 0-8018-5414-8.

The lapack library is described in, LAPACK Users’ Guide (Third Edition, 1999), Published by SIAM, ISBN 0-89871-447-8.

    http://www.netlib.org/lapack

The lapack source code can be found at the website above along with an online copy of the users guide.

14 Fast Fourier Transforms (FFTs)

This chapter describes functions for performing Fast Fourier Transforms (FFTs). The library includes radix-2 routines (for lengths which are a power of two) and mixed-radix routines (which work for any length). For efficiency there are separate versions of the routines for real data and for complex data. The mixed-radix routines are a reimplementation of the fftpack library by Paul Swarztrauber. Fortran code for fftpack is available on Netlib (fftpack also includes some routines for sine and cosine transforms but these are currently not available in GSL). For details and derivations of the underlying algorithms consult the document GSL FFT Algorithms (see Section 14.8 [FFT References and Further Reading], page 150)

14.1 Mathematical Definitions

Fast Fourier Transforms are efficient algorithms for calculating the discrete fourier transform (DFT),

x_{j} = ∑_{k=0}^{N−1} z_{k} exp(−2πijk/N)

The DFT usually arises as an approximation to the continuous fourier transform when functions are sampled at discrete intervals in space or time. The naive evaluation of the discrete fourier transform is a matrix-vector multiplication W z⃗. A general matrix-vector multiplication takes O(N^{2}) operations for N data-points. Fast fourier transform algorithms use a divide-and-conquer strategy to factorize the matrix W into smaller sub-matrices, corresponding to the integer factors of the length N. If N can be factorized into a product of integers f_{1}f_{2} … f_{n} then the DFT can be computed in O(N ∑ f_{i}) operations. For a radix-2 FFT this gives an operation count of O(N log_{2} N).

All the FFT functions offer three types of transform: forwards, inverse and backwards, based on the same mathematical definitions. The definition of the forward fourier transform, x = FFT(z), is,

x_{j} = ∑_{k=0}^{N−1} z_{k} exp(−2πijk/N)

and the definition of the inverse fourier transform, x = IFFT(z), is,

z_{j} = \frac{1}{N} ∑_{k=0}^{N−1} x_{k} exp(2πijk/N).

The factor of 1/N makes this a true inverse. For example, a call to gsl_fft_complex_ forward followed by a call to gsl_fft_complex_inverse should return the original data (within numerical errors).

In general there are two possible choices for the sign of the exponential in the transform/ inverse-transform pair. GSL follows the same convention as fftpack, using a negative exponential for the forward transform. The advantage of this convention is that the inverse transform recreates the original function with simple fourier synthesis. Numerical Recipes uses the opposite convention, a positive exponential in the forward transform.

The backwards FFT is simply our terminology for an unscaled version of the inverse FFT,

z_{j}^{backwards} = ∑_{k=0}^{N−1} x_{k} exp(2πijk/N).

When the overall scale of the result is unimportant it is often convenient to use the backwards FFT instead of the inverse to save unnecessary divisions.

14.2 Overview of complex data FFTs

The inputs and outputs for the complex FFT routines are packed arrays of floating point numbers. In a packed array the real and imaginary parts of each complex number are placed in alternate neighboring elements. For example, the following definition of a packed array of length 6,

     gsl_complex_packed_array data[6];

can be used to hold an array of three complex numbers, z[3], in the following way,

     data[0] = Re(z[0])
     data[1] = Im(z[0])
     data[2] = Re(z[1])
     data[3] = Im(z[1])
     data[4] = Re(z[2])
     data[5] = Im(z[2])

A stride parameter allows the user to perform transforms on the elements z[stride*i] instead of z[i]. A stride greater than 1 can be used to take an in-place FFT of the column of a matrix. A stride of 1 accesses the array without any additional spacing between elements.

The array indices have the same ordering as those in the definition of the DFT — i.e. there are no index transformations or permutations of the data.

For physical applications it is important to remember that the index appearing in the DFT does not correspond directly to a physical frequency. If the time-step of the DFT is Δ then the frequency-domain includes both positive and negative frequencies, ranging from −1/(2Δ) through 0 to +1/(2Δ). The positive frequencies are stored from the beginning of the array up to the middle, and the negative frequencies are stored backwards from the end of the array.

Here is a table which shows the layout of the array data, and the correspondence between the time-domain data z, and the frequency-domain data x.

     index    z               x = FFT(z)
     0        z(t = 0)        x(f = 0)
     1        z(t = 1)        x(f = 1/(N Delta))
     2        z(t = 2)        x(f = 2/(N Delta))
     .        ........        ..................
     N/2      z(t = N/2)      x(f = +1/(2 Delta),
                                    -1/(2 Delta))
     .        ........        ..................
     N-3      z(t = N-3)      x(f = -3/(N Delta))
     N-2      z(t = N-2)      x(f = -2/(N Delta))
     N-1      z(t = N-1)      x(f = -1/(N Delta))

When N is even the location N/2 contains the most positive and negative frequencies +1/(2Δ), −1/(2Δ)) which are equivalent. If N is odd then general structure of the table above still applies, but N/2 does not appear.

14.3 Radix-2 FFT routines for complex data

The radix-2 algorithms described in this section are simple and compact, although not necessarily the most efficient. They use the Cooley-Tukey algorithm to compute in-place complex FFTs for lengths which are a power of 2 — no additional storage is required. The corresponding self-sorting mixed-radix routines offer better performance at the expense of requiring additional working space.

All these functions are declared in the header file ‘gsl_fft_complex.h’.

Function int gsl fft complex radix2 forward (gsl_complex_packed_array

    data[], size_t stride, size_t n)

Function int gsl fft complex radix2 transform

    (gsl_complex_packed_array data[], size_t stride, size_t n)

Function int gsl fft complex radix2 backward

    (gsl_complex_packed_array data[], size_t stride, size_t n)

Function int gsl fft complex radix2 inverse (gsl_complex_packed_array

    data[], size_t stride, size_t n)

These functions compute forward, backward and inverse FFTs of length n with stride stride, on the packed complex array data using an in-place radix-2 decimation-in-time algorithm. The length of the transform n is restricted to powers of two. The functions return a value of GSL_SUCCESS if no errors were detected, or GSL_EDOM if the length of the data n is not a power of two.

Function int gsl fft complex radix2 dif forward

    (gsl_complex_packed_array data[], size_t stride, size_t n)

Function int gsl fft complex radix2 dif transform

    (gsl_complex_packed_array data[], size_t stride, size_t n)

Function int gsl fft complex radix2 dif backward

    (gsl_complex_packed_array data[], size_t stride, size_t n)

Function int gsl fft complex radix2 dif inverse

    (gsl_complex_packed_array data[], size_t stride, size_t n)

These are decimation-in-frequency versions of the radix-2 FFT functions.

Here is an example program which computes the FFT of a short pulse in a sample of length 128. To make the resulting fourier transform real the pulse is defined for equal positive and negative times (−10 … 10), where the negative times wrap around the end of the array.

#include <stdio.h>
#include <math.h>
#include <gsl/gsl_errno.h>
#include <gsl/gsl_fft_complex.h>
#define REAL(z,i) ((z)[2*(i)])
#define IMAG(z,i) ((z)[2*(i)+1])

int
main (void)
{
int i;
double data[2*128];
for (i = 0; i < 128; i++)
  {
     REAL(data,i) = 0.0;
     IMAG(data,i) = 0.0;
  }
REAL(data,0) = 1.0;
for (i = 1; i <= 10; i++)
  {
     REAL(data,i) = REAL(data,128-i) = 1.0;
  }
for (i = 0; i < 128; i++)
  {
    printf ("%d %e %e\n", i,
            REAL(data,i), IMAG(data,i));
  }
printf ("\n");
gsl_fft_complex_radix2_forward (data, 1, 128);
for (i = 0; i < 128; i++)
  {
    printf ("%d %e %e\n", i,
            REAL(data,i)/sqrt(128),
            IMAG(data,i)/sqrt(128));
  }
return 0;
}

Note that we have assumed that the program is using the default error handler (which calls abort for any errors). If you are not using a safe error handler you would need to check

the return status of gsl_fft_complex_radix2_forward.

The transformed data is rescaled by 1/\sqrt{N} so that it fits on the same plot as the input. Only the real part is shown, by the choice of the input data the imaginary part is zero. Allowing for the wrap-around of negative times at t = 128, and working in units of k/N, the DFT approximates the continuum fourier transform, giving a modulated sin function.

∫_{−a}^{+a} e^{−2πikx}dx = \frac{sin(2πka)}{πk}

A pulse and its discrete fourier transform, output from the example program.

14.4 Mixed-radix FFT routines for complex data

This section describes mixed-radix FFT algorithms for complex data. The mixed-radix functions work for FFTs of any length. They are a reimplementation of the Fortran fftpack library by Paul Swarztrauber. The theory is explained in the review article Self-sorting Mixed-radix FFTs by Clive Temperton. The routines here use the same indexing scheme and basic algorithms as fftpack.

The mixed-radix algorithm is based on sub-transform modules – highly optimized small length FFTs which are combined to create larger FFTs. There are efficient modules for factors of 2, 3, 4, 5, 6 and 7. The modules for the composite factors of 4 and 6 are faster than combining the modules for 2 ∗ 2 and 2 ∗ 3.

For factors which are not implemented as modules there is a fall-back to a general lengthn module which uses Singleton’s method for efficiently computing a DFT. This module is O(n^{2}), and slower than a dedicated module would be but works for any length n. Of course, lengths which use the general length-n module will still be factorized as much as possible. For example, a length of 143 will be factorized into 11∗13. Large prime factors are the worst case scenario, e.g. as found in n = 2 ∗ 3 ∗ 99991, and should be avoided because their O(n^{2}) scaling will dominate the run-time (consult the document GSL FFT Algorithms included in the GSL distribution if you encounter this problem).

The mixed-radix initialization function gsl_fft_complex_wavetable_alloc returns the list of factors chosen by the library for a given length N. It can be used to check how well the length has been factorized, and estimate the run-time. To a first approximation the run-time scales as N ∑ f_{i}, where the f_{i} are the factors of N. For programs under user control you may wish to issue a warning that the transform will be slow when the length is poorly factorized. If you frequently encounter data lengths which cannot be factorized using the existing small-prime modules consult GSL FFT Algorithms for details on adding support for other factors.

All these functions are declared in the header file ‘gsl_fft_complex.h’.

Function gsl_fft_complex_wavetable * gsl fft complex wavetable alloc

    (size_t n)

This function prepares a trigonometric lookup table for a complex FFT of length n. The function returns a pointer to the newly allocated gsl_fft_complex_wavetable if no errors were detected, and a null pointer in the case of error. The length n is factorized into a product of subtransforms, and the factors and their trigonometric coefficients are stored in the wavetable. The trigonometric coefficients are computed using direct calls to sin and cos, for accuracy. Recursion relations could be used to compute the lookup table faster, but if an application performs many FFTs of the same length then this computation is a one-off overhead which does not affect the final throughput. The wavetable structure can be used repeatedly for any transform of the same length. The table is not modified by calls to any of the other FFT functions. The same wavetable can be used for both forward and backward (or inverse) transforms of a given length.

Function void gsl fft complex wavetable free

    (gsl_fft_complex_wavetable * wavetable)

This function frees the memory associated with the wavetable wavetable. The wavetable can be freed if no further FFTs of the same length will be needed.

These functions operate on a gsl_fft_complex_wavetable structure which contains internal parameters for the FFT. It is not necessary to set any of the components directly but it can sometimes be useful to examine them. For example, the chosen factorization of the FFT length is given and can be used to provide an estimate of the run-time or numerical error.

The wavetable structure is declared in the header file ‘gsl_fft_complex.h’.

Data Type gsl fft complex wavetable

This is a structure that holds the factorization and trigonometric lookup tables for the mixed radix fft algorithm. It has the following components:

size_t n This is the number of complex data points

size_t nf This is the number of factors that the length n was decomposed into.

size_t factor[64]

This is the array of factors. Only the first nf elements are used.

gsl_complex * trig

This is a pointer to a preallocated trigonometric lookup table of n complex elements.

gsl_complex * twiddle[64]

This is an array of pointers into trig, giving the twiddle factors for each pass.

The mixed radix algorithms require an additional working space to hold the intermediate steps of the transform.

Function gsl_fft_complex_workspace * gsl fft complex workspace alloc

    (size_t n)

This function allocates a workspace for a complex transform of length n.

Function void gsl fft complex workspace free

    (gsl_fft_complex_workspace * workspace)

This function frees the memory associated with the workspace workspace. The workspace can be freed if no further FFTs of the same length will be needed.

The following functions compute the transform,

Function int gsl fft complex forward (gsl_complex_packed_array data[],

    size_t stride, size_t n, const gsl_fft_complex_wavetable * wavetable,
    gsl_fft_complex_workspace * work)

Function int gsl fft complex transform (gsl_complex_packed_array data[],

    size_t stride, size_t n, const gsl_fft_complex_wavetable * wavetable,
    gsl_fft_complex_workspace * work)

Function int gsl fft complex backward (gsl_complex_packed_array data[],

    size_t stride, size_t n, const gsl_fft_complex_wavetable * wavetable,
    gsl_fft_complex_workspace * work)

Function int gsl fft complex inverse (gsl_complex_packed_array data[],

    size_t stride, size_t n, const gsl_fft_complex_wavetable * wavetable,
    gsl_fft_complex_workspace * work)

These functions compute forward, backward and inverse FFTs of length n with stride stride, on the packed complex array data, using a mixed radix decimation-in-frequency algorithm. There is no restriction on the length n. Efficient modules are provided for subtransforms of length 2, 3, 4, 5, 6 and 7. Any remaining factors are computed with a slow, O(n^{2}), general-n module. The caller must supply a wavetable containing the trigonometric lookup tables and a workspace work. The functions return a value of 0 if no errors were detected. The following gsl_errno conditions are defined for these functions:

GSL_EDOM The length of the data n is not a positive integer (i.e. n is zero).

GSL_EINVAL

The length of the data n and the length used to compute the given wavetable do not match.

Here is an example program which computes the FFT of a short pulse in a sample of length 630 (= 2 ∗ 3 ∗ 3 ∗ 5 ∗ 7) using the mixed-radix algorithm.

#include <stdio.h>
#include <math.h>
#include <gsl/gsl_errno.h>
#include <gsl/gsl_fft_complex.h>
#define REAL(z,i) ((z)[2*(i)])
#define IMAG(z,i) ((z)[2*(i)+1])
int
main (void)
{
int i;
const int n = 630;
double data[2*n];
gsl_fft_complex_wavetable * wavetable;
gsl_fft_complex_workspace * workspace;
for (i = 0; i < n; i++)
  {
    REAL(data,i) = 0.0;
    IMAG(data,i) = 0.0;
  }
data[0].real = 1.0;
for (i = 1; i <= 10; i++)
  {
    REAL(data,i) = REAL(data,n-i) = 1.0;
  }
for (i = 0; i < n; i++)
  {
    printf ("%d: %e %e\n", i, REAL(data,i),
                              IMAG(data,i));
  }
printf ("\n");
wavetable = gsl_fft_complex_wavetable_alloc (n);
workspace = gsl_fft_complex_workspace_alloc (n);
for (i = 0; i < wavetable->nf; i++)
  {
     printf("# factor %d: %d\n", i,
            wavetable->factor[i]);
  }
gsl_fft_complex_forward (data, 1, n,
                         wavetable, workspace);

       for (i = 0; i < n; i++)
         {
           printf ("%d: %e %e\n", i, REAL(data,i),
                                     IMAG(data,i));
         }
       gsl_fft_complex_wavetable_free (wavetable);
       gsl_fft_complex_workspace_free (workspace);
       return 0;
     }

Note that we have assumed that the program is using the default gsl error handler (which calls abort for any errors). If you are not using a safe error handler you would need to check the return status of all the gsl routines.

14.5 Overview of real data FFTs

The functions for real data are similar to those for complex data. However, there is an important difference between forward and inverse transforms. The fourier transform of a real sequence is not real. It is a complex sequence with a special symmetry:

z_{k} = z_{N−k}^{∗}

A sequence with this symmetry is called conjugate-complex or half-complex. This different structure requires different storage layouts for the forward transform (from real to halfcomplex) and inverse transform (from half-complex back to real). As a consequence the routines are divided into two sets: functions in gsl_fft_real which operate on real sequences and functions in gsl_fft_halfcomplex which operate on half-complex sequences.

Functions in gsl_fft_real compute the frequency coefficients of a real sequence. The half-complex coefficients c of a real sequence x are given by fourier analysis,

c_{k} = ∑_{j=0}^{N−1} x_{k} exp(−2πijk/N)

Functions in gsl_fft_halfcomplex compute inverse or backwards transforms. They reconstruct real sequences by fourier synthesis from their half-complex frequency coefficients, c,

x_{j} = \frac{1}{N} ∑_{k=0}^{N−1} c_{k} exp(2πijk/N)

The symmetry of the half-complex sequence implies that only half of the complex numbers in the output need to be stored. The remaining half can be reconstructed using the halfcomplex symmetry condition. (This works for all lengths, even and odd. When the length is even the middle value, where k = N/2, is also real). Thus only N real numbers are required to store the half-complex sequence, and the transform of a real sequence can be stored in the same size array as the original data.

The precise storage arrangements depend on the algorithm, and are different for radix-2 and mixed-radix routines. The radix-2 function operates in-place, which constrain the locations where each element can be stored. The restriction forces real and imaginary parts to be stored far apart. The mixed-radix algorithm does not have this restriction, and it stores the real and imaginary parts of a given term in neighboring locations. This is desirable for better locality of memory accesses.

14.6 Radix-2 FFT routines for real data

This section describes radix-2 FFT algorithms for real data. They use the Cooley-Tukey algorithm to compute in-place FFTs for lengths which are a power of 2.

The radix-2 FFT functions for real data are declared in the header files ‘gsl_fft_real.h’

Function int gsl fft real radix2 transform (double data[], size_t stride,

size_t n)

This function computes an in-place radix-2 FFT of length n and stride stride on the real array data. The output is a half-complex sequence, which is stored in-place. The arrangement of the half-complex terms uses the following scheme: for k < N/2 the real part of the k-th term is stored in location k, and the corresponding imaginary part is stored in location N − k. Terms with k > N/2 can be reconstructed using the symmetry z_{k} = z_{N−k}^{∗}. The terms for k = 0 and k = N/2 are both purely real, and count as a special case. Their real parts are stored in locations 0 and N/2 respectively, while their imaginary parts which are zero are not stored. The following table shows the correspondence between the output data and the equivalent results obtained by considering the input data as a complex sequence with zero imaginary part,

complex[0].real    =    data[0]
complex[0].imag    =    0
complex[1].real    =    data[1]
complex[1].imag    =    data[N-1]
...............         ................
complex[k].real    =    data[k]
complex[k].imag    =    data[N-k]
...............         ................
complex[N/2].real  =    data[N/2]
complex[N/2].real  =    0
...............         ................
complex[k'].real   =    data[k]        k' = N - k
complex[k'].imag   =   -data[N-k]
...............         ................
complex[N-1].real  =    data[1]
complex[N-1].imag  =   -data[N-1]

The radix-2 FFT functions for halfcomplex data are declared in the header file

‘gsl_fft_halfcomplex.h’.

Function int gsl fft halfcomplex radix2 inverse (double data[], size_t

stride, size_t n)

Function int gsl fft halfcomplex radix2 backward (double data[], size_t

stride, size_t n) These functions compute the inverse or backwards in-place radix-2 FFT of length n and stride stride on the half-complex sequence data stored according the output scheme used by gsl_fft_real_radix2. The result is a real array stored in natural order.

14.7 Mixed-radix FFT routines for real data

This section describes mixed-radix FFT algorithms for real data. The mixed-radix functions work for FFTs of any length. They are a reimplementation of the real-FFT routines in the Fortran fftpack library by Paul Swarztrauber. The theory behind the algorithm is explained in the article Fast Mixed-Radix Real Fourier Transforms by Clive Temperton. The routines here use the same indexing scheme and basic algorithms as fftpack.

The functions use the fftpack storage convention for half-complex sequences. In this convention the half-complex transform of a real sequence is stored with frequencies in increasing order, starting at zero, with the real and imaginary parts of each frequency in neighboring locations. When a value is known to be real the imaginary part is not stored. The imaginary part of the zero-frequency component is never stored. It is known to be zero (since the zero frequency component is simply the sum of the input data (all real)). For a sequence of even length the imaginary part of the frequency n/2 is not stored either, since the symmetry z_{k} = z_{N−k}^{∗} implies that this is purely real too.

The storage scheme is best shown by some examples. The table below shows the output for an odd-length sequence, n = 5. The two columns give the correspondence between the 5 values in the half-complex sequence returned by gsl_fft_real_transform, halfcomplex[] and the values complex[] that would be returned if the same real input sequence were passed to gsl_fft_complex_backward as a complex sequence (with imaginary parts set to 0),

complex[0].real  =  halfcomplex[0]
complex[0].imag  =  0
complex[1].real  =  halfcomplex[1]
complex[1].imag  =  halfcomplex[2]
complex[2].real  =  halfcomplex[3]
complex[2].imag  =  halfcomplex[4]
complex[3].real  =  halfcomplex[3]
complex[3].imag  = -halfcomplex[4]
complex[4].real  =  halfcomplex[1]
complex[4].imag  = -halfcomplex[2]

The upper elements of the complex array, complex[3] and complex[4] are filled in using the symmetry condition. The imaginary part of the zero-frequency term complex[0].imag is known to be zero by the symmetry.

The next table shows the output for an even-length sequence, n = 5 In the even case both the there are two values which are purely real,

complex[0].real  =  halfcomplex[0]
complex[0].imag  =  0
complex[1].real  =  halfcomplex[1]
complex[1].imag  =  halfcomplex[2]
complex[2].real  =  halfcomplex[3]
complex[2].imag  =  halfcomplex[4]
complex[3].real  =  halfcomplex[5]
complex[3].imag  =  0
complex[4].real  =  halfcomplex[3]
complex[4].imag  = -halfcomplex[4]
complex[5].real  =  halfcomplex[1]
complex[5].imag  = -halfcomplex[2]

The upper elements of the complex array, complex[4] and complex[5] are filled in using the symmetry condition. Both complex[0].imag and complex[3].imag are known to be zero.

All these functions are declared in the header files ‘gsl_fft_real.h’ and

‘gsl_fft_halfcomplex.h’.

Function gsl_fft_real_wavetable * gsl fft real wavetable alloc (size_t

n)

Function gsl_fft_halfcomplex_wavetable *

gsl fft halfcomplex wavetable alloc (size_t n)

These functions prepare trigonometric lookup tables for an FFT of size n real elements. The functions return a pointer to the newly allocated struct if no errors were detected, and a null pointer in the case of error. The length n is factorized into a product of subtransforms, and the factors and their trigonometric coefficients are stored in the wavetable. The trigonometric coefficients are computed using direct calls to sin and cos, for accuracy. Recursion relations could be used to compute the lookup table faster, but if an application performs many FFTs of the same length then computing the wavetable is a one-off overhead which does not affect the final throughput. The wavetable structure can be used repeatedly for any transform of the same length. The table is not modified by calls to any of the other FFT functions. The appropriate type of wavetable must be used for forward real or inverse half-complex transforms.

Function void gsl fft real wavetable free (gsl_fft_real_wavetable *

wavetable)

Function void gsl fft halfcomplex wavetable free

    (gsl_fft_halfcomplex_wavetable * wavetable)

These functions free the memory associated with the wavetable wavetable. The wavetable can be freed if no further FFTs of the same length will be needed.

The mixed radix algorithms require an additional working space to hold the intermediate steps of the transform,

Function gsl_fft_real_workspace * gsl fft real workspace alloc (size_t

n) This function allocates a workspace for a real transform of length n. The same workspace is used for both forward real and inverse halfcomplex transforms.

Function void gsl fft real workspace free (gsl_fft_real_workspace *

workspace) This function frees the memory associated with the workspace workspace. The workspace can be freed if no further FFTs of the same length will be needed.

The following functions compute the transforms of real and half-complex data, Function int gsl fft real transform (double data[], size_t stride, size_t n,

    const gsl_fft_real_wavetable * wavetable, gsl_fft_real_workspace *

work)

Function int gsl fft halfcomplex transform (double data[], size_t stride,

    size_t n, const gsl_fft_halfcomplex_wavetable * wavetable,
    gsl_fft_real_workspace * work)

These functions compute the FFT of data, a real or half-complex array of length n, using a mixed radix decimation-in-frequency algorithm. For gsl_fft_real_transform data is an array of time-ordered real data. For gsl_fft_halfcomplex_transform data contains fourier coefficients in the half-complex ordering described above. There is no restriction on the length n. Efficient modules are provided for subtransforms of length 2, 3, 4 and 5. Any remaining factors are computed with a slow, O(n^{2}), generaln module. The caller must supply a wavetable containing trigonometric lookup tables and a workspace work.

Function int gsl fft real unpack (const double real coefficient[],

    gsl_complex_packed_array complex coefficient[], size_t stride, size_t n)

This function converts a single real array, real coefficient into an equivalent complex array, complex coefficient, (with imaginary part set to zero), suitable for gsl_fft_ complex routines. The algorithm for the conversion is simply,

     for (i = 0; i < n; i++)
       {
         complex_coefficient[i].real
           = real_coefficient[i];
         complex_coefficient[i].imag
           = 0.0;
       }

Function int gsl fft halfcomplex unpack (const double

halfcomplex coefficient[], gsl_complex_packed_array complex coefficient[],

    size_t stride, size_t n)

This function converts halfcomplex coefficient, an array of half-complex coefficients as returned by gsl_fft_real_transform, into an ordinary complex array, complex coefficient. It fills in the complex array using the symmetry z_{k} = z_{N−k}^{∗} to reconstruct the redundant elements. The algorithm for the conversion is,

     complex_coefficient[0].real
       = halfcomplex_coefficient[0];
     complex_coefficient[0].imag
       = 0.0;
     for (i = 1; i < n - i; i++)
       {
         double hc_real
           = halfcomplex_coefficient[2 * i - 1];
         double hc_imag
           = halfcomplex_coefficient[2 * i];
         complex_coefficient[i].real = hc_real;
         complex_coefficient[i].imag = hc_imag;
         complex_coefficient[n - i].real = hc_real;

       complex_coefficient[n - i].imag = -hc_imag;
     }
   if (i == n - i)
     {
       complex_coefficient[i].real
         = halfcomplex_coefficient[n - 1];
       complex_coefficient[i].imag
         = 0.0;
     }
Here is an example program using gsl_fft_real_transform and gsl_fft_

halfcomplex_inverse. It generates a real signal in the shape of a square pulse. The pulse is fourier transformed to frequency space, and all but the lowest ten frequency components are removed from the array of fourier coefficients returned by gsl_fft_real_transform.

The remaining fourier coefficients are transformed back to the time-domain, to give a filtered version of the square pulse. Since fourier coefficients are stored using the halfcomplex symmetry both positive and negative frequencies are removed and the final filtered signal is also real.

#include <stdio.h>
#include <math.h>
#include <gsl/gsl_errno.h>
#include <gsl/gsl_fft_real.h>
#include <gsl/gsl_fft_halfcomplex.h>
int
main (void)
{
int i, n = 100;
double data[n];
gsl_fft_real_wavetable * real;
gsl_fft_halfcomplex_wavetable * hc;
gsl_fft_real_workspace * work;
for (i = 0; i < n; i++)
  {
    data[i] = 0.0;
  }
for (i = n / 3; i < 2 * n / 3; i++)
  {
    data[i] = 1.0;
  }
for (i = 0; i < n; i++)
  {
    printf ("%d: %e\n", i, data[i]);
  }
printf ("\n");

work = gsl_fft_real_workspace_alloc (n);
real = gsl_fft_real_wavetable_alloc (n);
gsl_fft_real_transform (data, 1, n,
                        real, work);
gsl_fft_real_wavetable_free (real);
for (i = 11; i < n; i++)
  {
    data[i] = 0;
  }
hc = gsl_fft_halfcomplex_wavetable_alloc (n);
gsl_fft_halfcomplex_inverse (data, 1, n,
                             hc, work);
gsl_fft_halfcomplex_wavetable_free (hc);
for (i = 0; i < n; i++)
  {
    printf ("%d: %e\n", i, data[i]);
  }
gsl_fft_real_workspace_free (work);
return 0;
}

Low-pass filtered version of a real pulse, output from the example program.

14.8 References and Further Reading

A good starting point for learning more about the FFT is the review article Fast Fourier Transforms: A Tutorial Review and A State of the Art by Duhamel and Vetterli,

P. Duhamel and M. Vetterli. Fast fourier transforms: A tutorial review and a state of the art. Signal Processing, 19:259–299, 1990.

To find out about the algorithms used in the GSL routines you may want to consult the latex document GSL FFT Algorithms (it is included in GSL, as ‘doc/fftalgorithms.tex’). This has general information on FFTs and explicit derivations of the implementation for each routine. There are also references to the relevant literature. For convenience some of the more important references are reproduced below. There are several introductory books on the FFT with example programs, such as The Fast Fourier Transform by Brigham and DFT/FFT and Convolution Algorithms by Burrus and Parks,

E. Oran Brigham. The Fast Fourier Transform. Prentice Hall, 1974.

C. S. Burrus and T. W. Parks. DFT/FFT and Convolution Algorithms. Wiley, 1984.

Both these introductory books cover the radix-2 FFT in some detail. The mixed-radix algorithm at the heart of the fftpack routines is reviewed in Clive Temperton’s paper,

Clive Temperton. Self-sorting mixed-radix fast fourier transforms. Journal of Computational Physics, 52(1):1–23, 1983.

The derivation of FFTs for real-valued data is explained in the following two articles, Henrik V. Sorenson, Douglas L. Jones, Michael T. Heideman, and C. Sidney Burrus. Real-valued fast fourier transform algorithms. IEEE Transactions on Acoustics, Speech, and Signal Processing, ASSP-35(6):849–863, 1987. Clive Temperton. Fast mixed-radix real fourier transforms. Journal of Computational Physics, 52:340–350, 1983.

In 1979 the IEEE published a compendium of carefully-reviewed Fortran FFT programs in Programs for Digital Signal Processing. It is a useful reference for implementations of many different FFT algorithms,

Digital Signal Processing Committee and IEEE Acoustics, Speech, and Signal Processing Committee, editors. Programs for Digital Signal Processing. IEEE Press, 1979.

For serious FFT work we recommend the use of the dedicated FFTW library by Frigo and Johnson. The FFTW library is self-optimizing — it automatically tunes itself for each hardware platform in order to achieve maximum performance. It is available under the GNU GPL.

    FFTW Website, http://www.fftw.org/

Chapter 15: Numerical Integration 151

15 Numerical Integration

この章では 1 次元関数の数値積分 (求積) を実行するルーチンについて述べる. 一般の関数の適応的および非適応的積分ルーチンが実装されている. 無限領域および半無限領域での積分, 対数特異点を含む特異点の積分, コーシー主値の計算, そして振動積分も実装されている. このライブラリは

quadpack で使われている Piessens, Doncker-Kapenga, Uberhuber, Kahaner により実装され

た数値積分パッケージを移植したものである. quadpack の Fortran コードは Netlib で利用できる. この章で説明される関数はヘッダファイル ‘gsl_integration.h’で宣言されている.

15.1 Introduction

各アルゴリズムでは次の形の有界積分の近似値を計算する:

I = ∫_{a}^{b} f(x)w(x) dx

ここで w(x) は重み関数 (一般の被積分関数では w(x) = 1 ) である. ユーザーは次の精度要求を特定する相対誤差範囲 (epsabs, epsrel ) を指定できる:

|RESULT − I| ≤ max(epsabs, epsrel |I|)

ここでŖESULT はアルゴリズムにより得られた数値的近似値である. アルゴリズムは絶対誤差 A̧BSERR = |RESULT − I| を次の不等式をみたす形で評価する:

|RESULT − I| ≤ ABSERR ≤ max(epsabs, epsrel |I|)

ルーチンは誤差範囲が厳密すぎると収束しなくなる. その場合途中で得られる最良の近似値が返される.

quadpack にあるアルゴリズムは次のような名前付け規則に従っている:

Q - 求積ルーチン

N - 非適応的積分 A - 適応的積分

G - 一般積分 (ユーザー定義)

W - 被積分関数に重みづけ関数をかける

S - 発散を含む積分 P - 特異点を含む積分 I - 無限範囲での積分 O - cos または sin による振動関数での重みづけ F - フーリエ積分 C - コーシー主値

アルゴリズムには, 高次および低次の規則による 1 対の積分則が使われている. 高次規則は狭い範囲の積分の近似値を計算するのに使われる. 高次および低次の結果の差は近似の誤差評価に用いられる.

一般の (重みのない) 関数の積分アルゴリズムは Gauss-Kronrod 則に基づいている. Gauss-Kronrod 則は m 次の古典的ガウス積分から始まる. これに 2m + 1 次の高次 Kronrod 則を与えるように横軸に点を付加する. Kornrod 則はガウス則で評価した関数値を再利用するので効果的である. 高次 Kronrod 則は積分の最良近似値として用いられ, これら 2 つの規則の差は近似誤差として評価される.

重み関数のある被積分関数の場合は Clenshaw-Curtis 求積法が用いられる. Clenshaw-Curtis 則

は n 次 Chebyschev 多項式近似を被積分関数に適用する. この多項式は厳密に積分できるので, 元の被積分関数の積分近似値が求められる. Chebyschev 展開は近似を改良するために高次に拡張することができる. 被積分関数の特異点 (や他の特徴) は Chebyschev 近似では収束を遅くする. quadpack で使われている改良 Clenshaw-Curtis 則は収束を遅くするいくつかの汎用重み関数を分離している. これらの重み関数は Chebyschev 多項式に対して変形 Chebyschev モーメントとしてあらかじめ解析的に積分されている. このモーメントと関数の Chebyschev 近似を組みあわせることにより好きな積分を実行できる. 関数の特異部分に解析的積分を使うことで厳密に相殺することができ, 積分全体の収束性をかなり改善できる.

15.2 QNG non-adaptive Gauss-Kronrod integration

QNG アルゴリズムは, 最大で 87 点で被積分関数のサンプリングを行う固定 Gauss-Kronrod 則による非適応プロシージャである. これにより滑らかな関数を高速に積分できる.

Function int gsl integration qng (const gsl_function *f, double a, double
         b, double epsabs, double epsrel, double * result, double * abserr, size_t *

neval)

この関数は, (a, b) 上の f の積分の近似値が要求される絶対および相対誤差 epsabs および epsrel の範囲内にある限り 10 点, 21 点, 43 点および 87 点の Gauss-Kronrod 積分を実行する. この関数は最終的な近似値 result, 絶対誤差の見積り abserr, 用いられた関数評価数 neval を返す. Gauss-Kronrod 則は, 関数評価の総数を減らすため, 各段で前段の結果を利用するよう設計されている.

15.3 QAG adaptive integration

QAG アルゴリズムは簡単な適応的積分プロシージャである. 積分区間を分割し, 各段で最大見積り誤差を与える区間を分割する. これにより全体の誤差が急減し, 区間の分割は被積分関数の局所的難

点に集中することになる. この分割区間は gsl_integration_workspace構造体で管理され, 区間,

結果, そして評価誤差が格納される.

Function gsl_integration_workspace * gsl integration workspace alloc

         (size_t n)

この関数は n 個の倍精度の区間, 積分結果, 評価誤差を格納するのに十分な作業空間を確保する.

Function void gsl integration workspace free

         (gsl_integration_workspace * w)

この関数は作業空間 w に割りあてられていたメモリを解放する.

Function int gsl integration qag (const gsl_function *f, double a, double
         b, double epsabs, double epsrel, size_t limit, int key,
         gsl_integration_workspace * workspace, double * result, double * abserr)

この関数は (a, b) 上の f の積分の近似値が要求される絶対および相対誤差 epsabs および epsrel の範囲内にある限り適応的に積分を実行する. この関数は最終的な近似値 result, 絶対誤差の評価値 abserr を返す. 積分則は key の値により決定される. key は次のシンボル名から選ぶ:

     GSL_INTEG_GAUSS15  (key = 1)
     GSL_INTEG_GAUSS21  (key = 2)
     GSL_INTEG_GAUSS31  (key = 3)
     GSL_INTEG_GAUSS41  (key = 4)
     GSL_INTEG_GAUSS51  (key = 5)
     GSL_INTEG_GAUSS61  (key = 6)

これは 15, 21, 31, 41, 51, 61 点 Gauss-Kronrod 則に相当する. 高次則は滑らかな関数であ

れば精度がよくなる. 低次則は不連続のような局所的難点を含む関数で時間を節約できる. 積分の各段で, 適応的積分ストラテジに従い評価誤差が最大の区間を分割する. 区間分割とその結果は workspace で割りあてられるメモリに格納される. 区間分割数の最大値は limit で与えられる. これは割りあてた作業領域のサイズを越えてはならない.

15.4 QAGS adaptive integration with singularities

積分区間に可積分の特異点が存在すると, 適応的ルーチンの区間分割が特異点のまわりに集中してしまう. 分割された区間の幅が減少するとそれによる積分の近似値は限られた形でしか収束しない. この収束を外挿により加速させる. QAGS アルゴリズムは適応的区間分割に Wynn εアルゴリズムを融合させ, 様々な可積分特異点の積分をスピードアップさせる.

Function int gsl integration qags (const gsl_function * f, double a,
    double b, double epsabs, double epsrel, size_t limit,
    gsl_integration_workspace * workspace, double *result, double *abserr)

この関数は (a, b) 上の f の積分の近似値が要求される絶対および相対誤差 epsabs および epsrel の範囲内にある限り 21 点 Gauss-Kronrod 積分則を適応的に実行する. 結果は ϵ アルゴリズムにより外挿され, 不連続や可積分特異点の存在する積分の収束を加速させる. この関数は外挿による最終近似値 result, 絶対誤差の評価値 abserr を返す. 区間分割とその結果は workspace で割りあてられるメモリに格納される. 分割区間の最大数は limit で指定する. 作業空間の割りあてサイズを越えてはならない.

15.5 QAGP adaptive integration with known singular points

Function int gsl integration qagp (const gsl_function * f, double *pts,

    size_t npts, double epsabs, double epsrel, size_t limit,
    gsl_integration_workspace * workspace, double *result, double *abserr)

この関数はユーザーが提供する特異点の場所を考慮しながら適応的に積分を実行するアルゴリズム QAGS の実装である. 長さ npts の配列 pts には積分区間の端点と特異点の位置を格納する. 例えば, 区間 (a, b) 上で, x_{1}, x_{2}, x_{3} に特異点をもつ積分を実行する場合 (ただし

a < x_{1} < x_{2} < x_{3} < b ), 次のような pts 配列を与える:

     pts[0] = a
     pts[1] = x_1
     pts[2] = x_2
     pts[3] = x_3
     pts[4] = b

ここで npts = 5.

積分区間での特異点の位置を知っている場合は, このルーチンのほうが QAGSより早い.

15.6 QAGI adaptive integration on infinite intervals

Function int gsl integration qagi (gsl_function * f, double epsabs, double

epsrel, size_t limit, gsl_integration_workspace * workspace, double

*result, double *abserr)

この関数は非有界区間 (−∞, +∞) で関数 f を積分する. 積分は変数変換 x = (1 − t)/t により区間 (0, 1] に写像される:

∫_{−∞}^{+∞} dx f(x) = ∫_{0}^{1} dt (f((1 − t)/t) + f(−(1 − t)/t))/t^{2}.

そして QAGS アルゴリズムを用いて積分される. 変数変換により原点に可積分の特異点ができてしまうので, 通常の QAGS の 21 点 Gauss-Kronrod 則を 15 点則におきかえる. この場合低次則のほうがより効果的だからである.

Function int gsl integration qagiu (gsl_function * f, double a, double
epsabs, double epsrel, size_t limit, gsl_integration_workspace *

workspace, double *result, double *abserr)

この関数は半非有界区間 (a, +∞) 上で関数 f を積分する. 積分は変数変換 x = a + (1 − t)/t により区間 (0, 1] に写像される.

∫_{a}^{+∞} dx f(x) = ∫_{0}^{1} dt f(a + (1 − t)/t)/t^{2}

そして QAGS アルゴリズムを用いて積分される.

Function int gsl integration qagil (gsl_function * f, double b, double
epsabs, double epsrel, size_t limit, gsl_integration_workspace *

workspace, double *result, double *abserr)

この関数は半非有界区間 (−∞, b) 上で関数 f を積分する. 積分は変数変換 x = b − (1 − t)/t により区間 (0, 1] に写像される.

∫_{−∞}^{b} dx f(x) = ∫_{0}^{1} dt f(b − (1 − t)/t)/t^{2}

そして QAGS アルゴリズムを用いて積分される.

15.7 QAWC adaptive integration for Cauchy principal values

Function int gsl integration qawc (gsl_function *f, double a, double b,
double c, double epsabs, double epsrel, size_t limit,
gsl_integration_workspace * workspace, double * result, double * abserr)

この関数は (a, b) 上の f の積分の c での特異点のコーシー主値を求める.

I = ∫_{a}^{b} dx \frac{f(x)}{x − c} = lim_{ϵ→0} \{∫_{a}^{c−ϵ} dx \frac{f(x)}{x − c} + ∫_{c+ϵ}^{b} dx \frac{f(x)}{x − c}\}

QAG の適応的分割アルゴリズムが使われるが, 特異点 x = c で分割されないように工夫されている. 分割区間が点 x = c を含んでいたり, その点に近い場合は特別な 25 点変形 Clenshaw-Curtis 則が特異点を避けるために使われる. 特異点から離れた場所では通常の 15

点 Gauss-Kronrod 積分則が使われる.

15.8 QAWS adaptive integration for singular functions

QAWS アルゴリズムは, 被積分関数が積分領域の端点で対数的な発散をするときに用いられる. 効果的に計算するため, Chebyschev モーメントをあらかじめ計算しておく.

Function gsl_integration_qaws_table *

gsl integration qaws table alloc (double alpha, double beta, int mu,

int nu)

この関数は gsl_integration_gaws_tableの領域を確保し, 特異点の重み関数 W (x) をパ

ラメータ (α, β, μ, ν) で表現するための作業領域を割りあてる:

W (x) = (x − a)^{α}(b − x)^{β} log^{μ}(x − a) log^{ν}(b − x)

ここで α < −1, β < −1, μ = 0, 1, ν = 0, 1 である. 重み関数は μ, ν の値により以下の 4 つ

の形をとる:

W (x) = (x − a)^{α}(b − x)^{β} (μ = 0, ν = 0)

W (x) = (x − a)^{α}(b − x)^{β} log(x − a) (μ = 1, ν = 0)
W (x) = (x − a)^{α}(b − x)^{β} log(b − x) (μ = 0, ν = 1)
W (x) = (x − a)^{α}(b − x)^{β} log(x − a) log(b − x) (μ = 1, ν = 1)

特異点 (a, b) は積分が計算されるまで特定されなくてもよい. これらは積分領域の端点である. この関数は, エラーが検出されなければ, 新しく割りあてられた gsl_integration_qaws_ tableへのポインタを返す. エラーの場合には 0 を返す.

Function int gsl integration qaws table set (gsl_integration_qaws_table

* t, double alpha, double beta, int mu, int nu)

この関数は gsl_integration_qaws_table構造体 t のパラメータ (α, β, μ, ν) を変更する.

Function void gsl integration qaws table free

(gsl_integration_qaws_table * t)
この関数は gsl_integration_qaws_table構造体 t に割りあてられたメモリを解放する.
Function int gsl integration qaws (gsl_function * f, const double a,
const double b, gsl_integration_qaws_table * t, const double epsabs,
const double epsrel, const size_t limit, gsl_integration_workspace *

workspace, double *result, double *abserr) この関数は特異点重み関数 (x − a)^{α}(b − x)^{β} log^{μ}(x − a) log^{ν}(b − x) を用いて区間 (a, b) 上

で関数 f(x) を積分する. パラメータ (α, β, μ, ν) はテーブル t から取得する.

I = ∫_{a}^{b} dx f(x)(x − a)^{α}(b − x)^{β} log^{μ}(x − a) log^{ν}(b − x).

QAG の適応的分割アルゴリズムが用いられる. 分割区間に端点が含まれるときは特殊 25 点変形 Clenshaw-Curtis 則が特異点を避けるために用いられる. 端点を含まない区間では通常の

15 点 Gauss-Kronrod 則が用いられる.

15.9 QAWO adaptive integration for oscillatory functions

QAWO アルゴリズムは振動因子 sin(ωx), cos(ωx) をもつ積分を計算する. 効果的に計算するため, この因子をあらかじめ計算した Chebyschev モーメントのテーブルが必要となる.

Function gsl_integration_qawo_table *

gsl integration qawo table alloc (double omega, double L, enum

    gsl_integration_qawo_enum sine, size_t n)
この関数は gsl_integration_qawo_table構造体のために空間を割りあて, パラメータ

(ω, L) をもつサインまたはコサインの重み関数 W (x) のための作業領域となる.

W (x) = \{ \begin{matrix}sin(ωx) \\ cos(ωx) \end{matrix} \}

パラメータ L は関数の積分領域の長さ L = b − a を与える. サインかコサインの選択はパラメータ sine で行われる. 値には以下のシンボル値を用いる:

     GSL_INTEG_COSINE
     GSL_INTEG_SINE

gsl_integration_qawo_tableは積分過程で必要となる三角関数表である. パラメータ n に

より計算される係数のレベルが指定される. 各レベルは間隔 L の分割に相当するので, n レベ

ルは長さ L/2^{n} までの分割区間に対応できる. 積分ルーチン gsl_integration_qawoは, 要

求された精度に対してレベル数が不足している場合には GSL_ETABLEを返す.

Function int gsl integration qawo table set (gsl_integration_qawo_table

    * t, double omega, double L, enum gsl_integration_qawo_enum sine)

この関数は存在する作業領域 t のパラメータ omega, L, sine を変更する.

Function int gsl integration qawo table set length

    (gsl_integration_qawo_table * t, double L)

この関数は作業領域 t の長さのパラメータ L を変更する.

Function void gsl integration qawo table free

    (gsl_integration_qawo_table * t)

この関数は作業領域 t に割りあてられたメモリを解放する.

Function int gsl integration qawo (gsl_function * f, const double a,
    const double epsabs, const double epsrel, const size_t limit,
    gsl_integration_workspace * workspace, gsl_integration_qawo_table *
    wf, double *result, double *abserr)

この関数は, テーブル wf で定義される重み関数 sin(ωx), cos(ωx) を用い, 区間 (a, b) 上で関数 f の積分を計算する適応的アルゴリズムである.

I = ∫_{a}^{b} dx f(x) \{ \begin{matrix}sin(ωx) \\ cos(ωx) \end{matrix} \}

積分の収束加速するため, 結果は ϵ アルゴリズムを用いて外挿される. 関数は最終的な推定値 result, 絶対誤差の評価値 abserr を返す. 分割区間とその結果は workspace のメモリに格納される. 分割区間の最大数は limit で与えられる. これは作業領域の割りあてサイズを越えてはならない. 「大きな」幅 d, dω > 4 をもつ分割区間は 25 点 Clenshaw-Curtis 積分則を用いて計算し, 振動を処理する. 「小さな」幅 dω < 4 をもつ分割区間では 15 点 Gauss-Kronrod 積分を用いる.

15.10 QAWF adaptive integration for Fourier integrals

Function int gsl integration qawf (gsl_function * f, const double a,
    const double epsabs, const size_t limit, gsl_integration_workspace *
    workspace, gsl_integration_workspace * cycle workspace,
    gsl_integration_qawo_table * wf, double *result, double *abserr)

この関数は半非有界区間 [a, +∞) での関数 f の Fourier 積分を計算する.

I = ∫_{a}^{+∞} dx f(x) \{ \begin{matrix}sin(ωx) \\ cos(ωx) \end{matrix} \}

パラメータ ω はテーブル wf (長さ L は好きな値をとることができる. Fourier 積分に適切な値となるようにこの関数によりオーバーライドされるからである) からとられる. 積分は QAWO アルゴリズムを使って各分割区間で計算される.

\begin{aligned}C_{1} & = [a, a + c] \\ C_{2} & = [a + c, a + 2c] \\ … & = … \\ C_{k} & = [a + (k − 1)c, a + kc] \end{aligned}

ここでç = (2 floor(|ω|) + 1)π/|ω| は周期の奇数倍をカバーするように選ばれる. 各区間からの寄与は符号が交代し, f が正で単調減少する場合には単調減少する. この数列の和は ϵ アルゴリズムで加速する. この関数は全体の絶対誤差 abserr で押さえられる. 以下のストラテジが用いられる: 各区間 C_{k} でアルゴリズムは許容誤差を達成しようとする:

T OL_{k} = u_{k}abserr

ここでu̧_{k} = (1 − p)p^{k−1} および p = 9/10 である. 各区間の三角関数列の寄与の和は全体の許容誤差 abserr を与える. 分割区間の積分に困難が発生した場合は分割区間に対する精度要求を緩和する.

T OL_{k} = u_{k} max(abserr, max_{i<k}{E_{i}})

ここで E_{k} は区間 C_{k} での評価誤差である. 分割区間とその結果は workspace のメモリに格納される. 分割区間の最大数は limit で与えられる. 作業領域の割りあてサイズより大きくなくてはならない. 各分割区間での積分には cycle workspace のメモリが QAWO アルゴリズムのための作業領域として用いられる.

15.11 Error codes

不正引数に関する標準エラーコードの他, これらの関数は次の値を返す:

GSL_EMAXITER

区間分割の最大数を越えた.

GSL_EROUND

打ちきり誤差により許容範囲に届かなかった, もしくは外挿テーブルに打ちきり誤差が検出された.

GSL_ESING

可積分でない特異点もしくはその他の被積分関数の悪い振舞いが積分区間内に見つかった.

GSL_EDIVERGE

積分が発散した, もしくは積分の収束が遅すぎる.

15.12 Examples

積分ルーチン QAGSは有界積分の大くのクラスで計算できる. 例えば, 次のような積分を考える. これは原点に対数的な特異点をもつ:

∫_{0}^{1} x^{−1/2} log(x) dx = −4

以下のプログラムはこの積分を相対誤差 1e-7で積分するものである.

#include <stdio.h>
#include <math.h>
#include <gsl/gsl_integration.h>
double f (double x, void * params) {
double alpha = *(double *) params;
double f = log(alpha*x) / sqrt(x);
return f;
}
int
main (void)
{
gsl_integration_workspace * w
  = gsl_integration_workspace_alloc(1000);
double result, error;
double expected = -4.0;
double alpha = 1.0;
gsl_function F;
F.function = &f;
F.params = &alpha;
gsl_integration_qags (&F, 0, 1, 0, 1e-7, 1000,
                      w, &result, &error);
printf("result          = % .18f\n", result);
printf("exact result    = % .18f\n", expected);
printf("estimated error = % .18f\n", error);
printf("actual error    = % .18f\n", result - expected);
printf("intervals =  %d\n", w->size);
return 0;
}

以下に示す結果は, 8 回分割することにより望んだ精度が達成できたことを示している.

bash$ ./a.out
result          = -3.999999999999973799
exact result    = -4.000000000000000000
estimated error =  0.000000000000246025
actual error    =  0.000000000000026201
intervals =  8

実際, QAGSで用いられる外挿プロシージャは要求精度の倍近い精度をもつ. 外挿プロシージャの返す評価誤差は実際の誤差よりも大きい. 安全のため 1 桁のマージンをとっているからである.

15.13 References and Further Reading

以下の本は quadpack の参考書の決定版であり, 作者により書かれたものである. アルゴリズム, プログラム一覧, テストプログラム, そして例題が載せられている. 数値積分に関する有用なアドバイスや, quadpack の開発に用いられた数値積分に関する参考書も載せれている.

R. Piessens, E. de Doncker-Kapenga, C.W. Uberhuber, D.K. Kahaner. quadpack A subroutine package for automatic integration Springer Verlag, 1983.

16 Random Number Generation

The library provides a large collection of random number generators which can be accessed through a uniform interface. Environment variables allow you to select different generators and seeds at runtime, so that you can easily switch between generators without needing to recompile your program. Each instance of a generator keeps track of its own state, allowing the generators to be used in multi-threaded programs. Additional functions are available for transforming uniform random numbers into samples from continuous or discrete probability distributions such as the Gaussian, log-normal or Poisson distributions.

These functions are declared in the header file ‘gsl_rng.h’.

16.1 General comments on random numbers

In 1988, Park and Miller wrote a paper entitled “Random number generators: good ones are hard to find.” [Commun. ACM, 31, 1192–1201]. Fortunately, some excellent random number generators are available, though poor ones are still in common use. You may be happy with the system-supplied random number generator on your computer, but you should be aware that as computers get faster, requirements on random number generators increase. Nowadays, a simulation that calls a random number generator millions of times can often finish before you can make it down the hall to the coffee machine and back.

A very nice review of random number generators was written by Pierre L’Ecuyer, as Chapter 4 of the book: Handbook on Simulation, Jerry Banks, ed. (Wiley, 1997). The chapter is available in postscript from from L’Ecuyer’s ftp site (see references). Knuth’s volume on Seminumerical Algorithms (originally published in 1968) devotes 170 pages to random number generators, and has recently been updated in its 3rd edition (1997). It is brilliant, a classic. If you don’t own it, you should stop reading right now, run to the nearest bookstore, and buy it.

A good random number generator will satisfy both theoretical and statistical properties. Theoretical properties are often hard to obtain (they require real math!), but one prefers a random number generator with a long period, low serial correlation, and a tendency not to “fall mainly on the planes.” Statistical tests are performed with numerical simulations. Generally, a random number generator is used to estimate some quantity for which the theory of probability provides an exact answer. Comparison to this exact answer provides a measure of “randomness”.

16.2 The Random Number Generator Interface

It is important to remember that a random number generator is not a “real” function like sine or cosine. Unlike real functions, successive calls to a random number generator yield different return values. Of course that is just what you want for a random number generator, but to achieve this effect, the generator must keep track of some kind of “state” variable. Sometimes this state is just an integer (sometimes just the value of the previously generated random number), but often it is more complicated than that and may involve a whole array of numbers, possibly with some indices thrown in. To use the random number generators, you do not need to know the details of what comprises the state, and besides that varies from algorithm to algorithm.

The random number generator library uses two special structs, gsl_rng_type which holds static information about each type of generator and gsl_rng which describes an instance of a generator created from a given gsl_rng_type.

The functions described in this section are declared in the header file ‘gsl_rng.h’.

16.3 Random number generator initialization

Random gsl_rng * gsl rng alloc (const gsl_rng_type * T)

This function returns a pointer to a newly-created instance of a random number generator of type T. For example, the following code creates an instance of the Tausworthe generator,

     gsl_rng * r = gsl_rng_alloc (gsl_rng_taus);

If there is insufficient memory to create the generator then the function returns a null pointer and the error handler is invoked with an error code of GSL_ENOMEM. The generator is automatically initialized with the default seed, gsl_rng_default_ seed. This is zero by default but can be changed either directly or by using the environment variable GSL_RNG_SEED (see Section 16.6 [Random number environment variables], page 163). The details of the available generator types are described later in this chapter.

Random void gsl rng set (const gsl_rng * r, unsigned long int s)

This function initializes (or ‘seeds’) the random number generator. If the generator is seeded with the same value of s on two different runs, the same stream of random numbers will be generated by successive calls to the routines below. If different values of s are supplied, then the generated streams of random numbers should be completely different. If the seed s is zero then the standard seed from the original implementation is used instead. For example, the original Fortran source code for the ranlux generator used a seed of 314159265, and so choosing s equal to zero reproduces

this when using gsl_rng_ranlux.

Random void gsl rng free (gsl_rng * r)

This function frees all the memory associated with the generator r.

16.4 Sampling from a random number generator

The following functions return uniformly distributed random numbers, either as integers or double precision floating point numbers. To obtain non-uniform distributions see Chapter 18 [Random Number Distributions], page 177.

Random unsigned long int gsl rng get (const gsl_rng * r)

This function returns a random integer from the generator r. The minimum and maximum values depend on the algorithm used, but all integers in the range [min,max] are equally likely. The values of min and max can determined using the auxiliary

functions gsl_rng_max (r) and gsl_rng_min (r).

Random double gsl rng uniform (const gsl_rng * r)

This function returns a double precision floating point number uniformly distributed in the range [0,1). The range includes 0.0 but excludes 1.0. The value is typically obtained by dividing the result of gsl_rng_get(r) by gsl_rng_max(r) + 1.0 in double precision. Some generators compute this ratio internally so that they can provide floating point numbers with more than 32 bits of randomness (the maximum number of bits that can be portably represented in a single unsigned long int).

Random double gsl rng uniform pos (const gsl_rng * r)

This function returns a positive double precision floating point number uniformly distributed in the range (0,1), excluding both 0.0 and 1.0. The number is obtained by sampling the generator with the algorithm of gsl_rng_uniform until a non-zero value is obtained. You can use this function if you need to avoid a singularity at 0.0.

Random unsigned long int gsl rng uniform int (const gsl_rng * r,
    unsigned long int n)

This function returns a random integer from 0 to n-1 inclusive. All integers in the range [0,n-1] are equally likely, regardless of the generator used. An offset correction is applied so that zero is always returned with the correct probability, for any minimum value of the underlying generator. If n is larger than the range of the generator then the function calls the error handler with an error code of GSL_EINVAL and returns zero.

16.5 Auxiliary random number generator functions

The following functions provide information about an existing generator. You should use them in preference to hard-coding the generator parameters into your own code.

Random const char * gsl rng name (const gsl_rng * r)

This function returns a pointer to the name of the generator. For example,

     printf("r is a '%s' generator\n",
            gsl_rng_name (r));

would print something like r is a 'taus' generator.

Random unsigned long int gsl rng max (const gsl_rng * r)

gsl_rng_max returns the largest value that gsl_rng_get can return.

Random unsigned long int gsl rng min (const gsl_rng * r)

gsl_rng_min returns the smallest value that gsl_rng_get can return. Usually this value is zero. There are some generators with algorithms that cannot return zero, and for these generators the minimum value is 1.

Random void * gsl rng state (const gsl_rng * r)

Random size_t gsl rng size (const gsl_rng * r)

These function return a pointer to the state of generator r and its size. You can use this information to access the state directly. For example, the following code will write the state of a generator to a stream,

     void * state = gsl_rng_state (r);
     size_t n = gsl_rng_size (r);
     fwrite (state, n, 1, stream);

Random const gsl_rng_type ** gsl rng types setup (void)

This function returns a pointer to an array of all the available generator types, terminated by a null pointer. The function should be called once at the start of the program, if needed. The following code fragment shows how to iterate over the array of generator types to print the names of the available algorithms,

     const gsl_rng_type **t, **t0;
     t0 = gsl_rng_types_setup ();
     printf("Available generators:\n");
     for (t = t0; *t != 0; t++)
       {
         printf("%s\n", (*t)->name);
       }

16.6 Random number environment variables

The library allows you to choose a default generator and seed from the environment

variables GSL_RNG_TYPE and GSL_RNG_SEED and the function gsl_rng_env_setup. This

makes it easy try out different generators and seeds without having to recompile your program.

Function const gsl_rng_type * gsl rng env setup (void)

This function reads the environment variables GSL_RNG_TYPE and GSL_RNG_SEED and uses their values to set the corresponding library variables gsl_rng_default and gsl_rng_default_seed. These global variables are defined as follows,

     extern const gsl_rng_type *gsl_rng_default
     extern unsigned long int gsl_rng_default_seed

The environment variable GSL_RNG_TYPE should be the name of a generator, such as taus or mt19937. The environment variable GSL_RNG_SEED should contain the desired seed value. It is converted to an unsigned long int using the C library function strtoul. If you don’t specify a generator for GSL_RNG_TYPE then gsl_rng_mt19937 is used as the default. The initial value of gsl_rng_default_seed is zero.

Here is a short program which shows how to create a global generator using the environment

variables GSL_RNG_TYPE and GSL_RNG_SEED,
#include <stdio.h>
#include <gsl/gsl_rng.h>
gsl_rng * r;  /* global generator */
int
main (void)
{
  const gsl_rng_type * T;
  gsl_rng_env_setup();
  T = gsl_rng_default;
  r = gsl_rng_alloc (T);
  printf("generator type: %s\n", gsl_rng_name (r));
  printf("seed = %u\n", gsl_rng_default_seed);
  printf("first value = %u\n", gsl_rng_get (r));
  return 0;
}

Running the program without any environment variables uses the initial defaults, an mt19937 generator with a seed of 0,

bash$ ./a.out
generator type: mt19937
seed = 0
first value = 2867219139

By setting the two variables on the command line we can change the default generator and the seed,

bash$ GSL_RNG_TYPE="taus" GSL_RNG_SEED=123 ./a.out
GSL_RNG_TYPE=taus
GSL_RNG_SEED=123
generator type: taus
seed = 123
first value = 2720986350

16.7 Saving and restoring random number generator state

The above methods ignore the random number ‘state’ which changes from call to call. It is often useful to be able to save and restore the state. To permit these practices, a few somewhat more advanced functions are supplied. These include:

Random int gsl rng memcpy (gsl_rng * dest, const gsl_rng * src)

This function copies the random number generator src into the pre-existing generator dest, making dest into an exact copy of src. The two generators must be of the same type.

Random gsl_rng * gsl rng clone (const gsl_rng * r)

This function returns a pointer to a newly created generator which is an exact copy of the generator r.

Random void gsl rng print state (const gsl_rng * r)

This function prints a hex-dump of the state of the generator r to stdout. At the moment its only use is for debugging.

16.8 Random number generator algorithms

The functions described above make no reference to the actual algorithm used. This is deliberate so that you can switch algorithms without having to change any of your application source code. The library provides a large number of generators of different types, including simulation quality generators, generators provided for compatibility with other libraries and historical generators from the past.

The following generators are recommended for use in simulation. They have extremely long periods, low correlation and pass most statistical tests.

Generator gsl rng mt19937

The MT19937 generator of Makoto Matsumoto and Takuji Nishimura is a variant of the twisted generalized feedback shift-register algorithm, and is known as the "Mersenne Twister" generator. It has a Mersenne prime period of 2^{19937} − 1 (about 10^{6000}) and is equi-distributed in 623 dimensions. It has passed the diehard statistical tests. It uses 624 words of state per generator and is comparable in speed to the other generators. The original generator used a default seed of 4357 and choosing s equal to zero in gsl_rng_set reproduces this. For more information see,

Makoto Matsumoto and Takuji Nishimura, "Mersenne Twister: A 623dimensionally equidistributed uniform pseudorandom number generator". ACM Transactions on Modeling and Computer Simulation, Vol. 8, No. 1 (Jan. 1998), Pages 3-30

The generator gsl_rng_19937 uses the corrected version of the seeding procedure published later by the two authors above. The original seeding procedure suffered from low-order periodicity, but can be used by selecting the alternate generator gsl_

rng_mt19937_1998.

Generator gsl rng ranlxs0 Generator gsl rng ranlxs1 Generator gsl rng ranlxs2

The generator ranlxs0 is a second-generation version of the ranlux algorithm of Lüscher, which produces "luxury random numbers". This generator provides single precision output (24 bits) at three luxury levels ranlxs0, ranlxs1 and ranlxs2. It uses double-precision floating point arithmetic internally and can be significantly faster than the integer version of ranlux, particularly on 64-bit architectures. The period of the generator is about 10^{171}. The algorithm has mathematically proven properties and can provide truly decorrelated numbers at a known level of randomness. The higher luxury levels provide additional decorrelation between samples as an additional safety margin.

Generator gsl rng ranlxd1 Generator gsl rng ranlxd2

These generators produce double precision output (48 bits) from the ranlxs generator. The library provides two luxury levels ranlxd1 and ranlxd2.

Generator gsl rng ranlux Generator gsl rng ranlux389

The ranlux generator is an implementation of the original algorithm developed by Lüscher. It uses a lagged-fibonacci-with-skipping algorithm to produce "luxury random numbers". It is a 24-bit generator, originally designed for single-precision IEEE floating point numbers. This implementation is based on integer arithmetic, while the second-generation versions ranlxs and ranlxd described above provide floatingpoint implementations which will be faster on many platforms. The period of the generator is about 10^{171}. The algorithm has mathematically proven properties and it can provide truly decorrelated numbers at a known level of randomness. The default level of decorrelation recommended by Lüscher is provided by gsl_rng_ranlux, while gsl_rng_ranlux389 gives the highest level of randomness, with all 24 bits decorrelated. Both types of generator use 24 words of state per generator. For more information see,

M. Lüscher, "A portable high-quality random number generator for lattice field theory calculations", Computer Physics Communications, 79 (1994) 100-110. F. James, "RANLUX: A Fortran implementation of the high-quality pseudorandom number generator of Lüscher", Computer Physics Communications, 79 (1994) 111-114

Generator gsl rng cmrg

This is a combined multiple recursive generator by L’Ecuyer. Its sequence is,

z_{n} = (x_{n} − y_{n}) mod m_{1}

where the two underlying generators x_{n} and y_{n} are,

\begin{aligned}x_{n} & = (a_{1}x_{n−1} + a_{2}x_{n−2} + a_{3}x_{n−3}) mod m_{1} \\ y_{n} & = (b_{1}y_{n−1} + b_{2}y_{n−2} + b_{3}y_{n−3}) mod m_{2} \end{aligned}

with coefficients a_{1} = 0, a_{2} = 63308, a_{3} = −183326, b_{1} = 86098, b_{2} = 0, b_{3} = −539608, and moduli m_{1} = 2^{31} − 1 = 2147483647 and m_{2} = 2145483479. The period of this generator is 2^{205} (about 10^{61}). It uses 6 words of state per generator. For more information see,

P. L’Ecuyer, "Combined Multiple Recursive Random Number Generators," Operations Research, 44, 5 (1996), 816–822.

Generator gsl rng mrg

This is a fifth-order multiple recursive generator by L’Ecuyer, Blouin and Coutre. Its sequence is,

x_{n} = (a_{1}x_{n−1} + a_{5}x_{n−5}) mod m

with a_{1} = 107374182, a_{2} = a_{3} = a_{4} = 0, a_{5} = 104480 and m = 2^{31} − 1. The period of this generator is about 10^{46}. It uses 5 words of state per generator. More information can be found in the following paper, P. L’Ecuyer, F. Blouin, and R. Coutre, "A search for good multiple recursive random number generators", ACM Transactions on Modeling and Computer Simulation 3, 87-98 (1993).

Generator gsl rng taus

This is a maximally equidistributed combined Tausworthe generator by L’Ecuyer. The sequence is,

x_{n} = (s_{n}^{1} ⊕ s_{n}^{2} ⊕ s_{n}^{3})

where,

\begin{aligned}s_{n+1}^{1} & = (((s_{n}^{1}&4294967294) ≪ 12) ⊕ (((s_{n}^{1} ≪ 13) ⊕ s_{n}^{1}) ≫ 19)) \\ s_{n+1}^{2} & = (((s_{n}^{2}&4294967288) ≪ 4) ⊕ (((s_{n}^{2} ≪ 2) ⊕ s_{n}^{2}) ≫ 25)) \\ s_{n+1}^{3} & = (((s_{n}^{3}&4294967280) ≪ 17) ⊕ (((s_{n}^{3} ≪ 3) ⊕ s_{n}^{3}) ≫ 11)) \end{aligned}

computed modulo 2^{32}. In the formulas above ⊕ denotes “exclusive-or”. Note that the algorithm relies on the properties of 32-bit unsigned integers and has been implemented using a bitmask of 0xFFFFFFFF to make it work on 64 bit machines. The period of this generator is 2^{88} (about 10^{26}). It uses 3 words of state per generator. For more information see,

P. L’Ecuyer, "Maximally Equidistributed Combined Tausworthe Generators", Mathematics of Computation, 65, 213 (1996), 203–213.

Generator gsl rng gfsr4

The gfsr4 generator is like a lagged-fibonacci generator, and produces each number as an xor’d sum of four previous values.

r_{n} = r_{n−A} ⊕ r_{n−B} ⊕ r_{n−C} ⊕ r_{n−D}

Ziff (ref below) notes that "it is now widely known" that two-tap registers (such as R250, which is described below) have serious flaws, the most obvious one being the three-point correlation that comes from the definition of the generator. Nice mathematical properties can be derived for GFSR’s, and numerics bears out the claim that 4-tap GFSR’s with appropriately chosen offsets are as random as can be measured, using the author’s test. This implementation uses the values suggested the the example on p392 of Ziff’s article: A = 471, B = 1586, C = 6988, D = 9689. If the offsets are appropriately chosen (such the one ones in this implementation), then the sequence is said to be maximal. I’m not sure what that means, but I would guess that means all states are part of the same cycle, which would mean that the period for this generator is astronomical; it is (2^{K})^{D} ≈ 10^{93334} where K = 32 is the number of bits in the word, and D is the longest lag. This would also mean that any one random number could easily be zero; ie 0 ≤ r < 2^{32}. Ziff doesn’t say so, but it seems to me that the bits are completely independent here, so one could use this as an efficient bit generator; each number supplying 32 random bits. The quality of the generated bits depends on the underlying seeding procedure, which may need to be improved in some circumstances. For more information see,

Robert M. Ziff, "Four-tap shift-register-sequence random-number generators", Computers in Physics, 12(4), Jul/Aug 1998, pp 385-392.

16.9 Unix random number generators

The standard Unix random number generators rand, random and rand48 are provided as part of GSL. Although these generators are widely available individually often they aren’t all available on the same platform. This makes it difficult to write portable code using them and so we have included the complete set of Unix generators in GSL for convenience. Note that these generators don’t produce high-quality randomness and aren’t suitable for work requiring accurate statistics. However, if you won’t be measuring statistical quantities and just want to introduce some variation into your program then these generators are quite acceptable.

Generator gsl rng rand

This is the BSD rand() generator. Its sequence is

x_{n+1} = (ax_{n} + c) mod m

with a = 1103515245, c = 12345 and m = 2^{31}. The seed specifies the initial value, x_{1}. The period of this generator is 2^{31}, and it uses 1 word of storage per generator.

Generator gsl rng random bsd Generator gsl rng random libc5 Generator gsl rng random glibc2

These generators implement the random() family of functions, a set of linear feedback shift register generators originally used in BSD Unix. There are several versions of random() in use today: the original BSD version (e.g. on SunOS4), a libc5 version (found on older GNU/Linux systems) and a glibc2 version. Each version uses a different seeding procedure, and thus produces different sequences. The original BSD routines accepted a variable length buffer for the generator state, with longer buffers providing higher-quality randomness. The random() function implemented algorithms for buffer lengths of 8, 32, 64, 128 and 256 bytes, and the algorithm with the largest length that would fit into the user-supplied buffer was used. To support these algorithms additional generators are available with the following names,

     gsl_rng_random8_bsd
     gsl_rng_random32_bsd
     gsl_rng_random64_bsd
     gsl_rng_random128_bsd
     gsl_rng_random256_bsd

where the numeric suffix indicates the buffer length. The original BSD random function used a 128-byte default buffer and so gsl_rng_random_bsd has been made equivalent to gsl_rng_random128_bsd. Corresponding versions of the libc5 and glibc2 generators are also available, with the names gsl_rng_random8_libc5, gsl_rng_

random8_glibc2, etc.

Generator gsl rng rand48

This is the Unix rand48 generator. Its sequence is

x_{n+1} = (ax_{n} + c) mod m

defined on 48-bit unsigned integers with a = 25214903917, c = 11 and m = 2^{48}. The seed specifies the upper 32 bits of the initial value, x_{1}, with the lower 16 bits set to 0x330E. The function gsl_rng_get returns the upper 32 bits from each term of the sequence. This does not have a direct parallel in the original rand48 functions, but forcing the result to type long int reproduces the output of mrand48. The function gsl_rng_uniform uses the full 48 bits of internal state to return the double precision number x_{n}/m, which is equivalent to the function drand48. Note that some versions of the GNU C Library contained a bug in mrand48 function which caused it to produce different results (only the lower 16-bits of the return value were set).

16.10 Numerical Recipes generators

The following generators are provided for compatibility with Numerical Recipes. Note that the original Numerical Recipes functions used single precision while we use double precision. This will lead to minor discrepancies, but only at the level of single-precision rounding error. If necessary you can force the returned values to single precision by storing them in a volatile float, which prevents the value being held in a register with double or extended precision. Apart from this difference the underlying algorithms for the integer part of the generators are the same.

Generator gsl rng ran0

Numerical recipes ran0 implements Park and Miller’s minstd algorithm with a modified seeding procedure.

Generator gsl rng ran1

Numerical recipes ran1 implements Park and Miller’s minstd algorithm with a 32element Bayes-Durham shuffle box.

Generator gsl rng ran2

Numerical recipes ran2 implements a L’Ecuyer combined recursive generator with a 32-element Bayes-Durham shuffle-box.

Generator gsl rng ran3

Numerical recipes ran3 implements Knuth’s portable subtractive generator.

16.11 Other random number generators

The generators in this section are provided for compatibility with existing libraries. If you are converting an existing program to use GSL then you can select these generators to check your new implementation against the original one, using the same random number generator. After verifying that your new program reproduces the original results you can then switch to a higher-quality generator.

Note that most of the generators in this section are based on single linear congruence relations, which are the least sophisticated type of generator. In particular, linear congruences have poor properties when used with a non-prime modulus, as several of these routines do (e.g. with a power of two modulus, 2^{31} or 2^{32}). This leads to periodicity in the least significant bits of each number, with only the higher bits having any randomness. Thus if you want to produce a random bitstream it is best to avoid using the least significant bits.

Generator gsl rng ranf

This is the CRAY random number generator RANF. Its sequence is

x_{n+1} = (ax_{n}) mod m

defined on 48-bit unsigned integers with a = 44485709377909 and m = 2^{48}. The seed specifies the lower 32 bits of the initial value, x_{1}, with the lowest bit set to prevent the seed taking an even value. The upper 16 bits of x_{1} are set to 0. A consequence of this procedure is that the pairs of seeds 2 and 3, 4 and 5, etc produce the same sequences. The generator compatibile with the CRAY MATHLIB routine RANF. It produces double precision floating point numbers which should be identical to those from the original RANF. There is a subtlety in the implementation of the seeding. The initial state is reversed through one step, by multiplying by the modular inverse of a mod m. This is done for compatibility with the original CRAY implementation. Note that you can only seed the generator with integers up to 2^{32}, while the original CRAY implementation uses non-portable wide integers which can cover all 2^{48} states of the generator. The function gsl_rng_get returns the upper 32 bits from each term of the sequence. The function gsl_rng_uniform uses the full 48 bits to return the double precision number x_{n}/m. The period of this generator is 2^{46}.

Generator gsl rng ranmar

This is the RANMAR lagged-fibonacci generator of Marsaglia, Zaman and Tsang. It is a 24-bit generator, originally designed for single-precision IEEE floating point numbers. It was included in the CERNLIB high-energy physics library.

Generator gsl rng r250

This is the shift-register generator of Kirkpatrick and Stoll. The sequence is

x_{n} = x_{n−103} ⊕ x_{n−250}

where ⊕ denote “exclusive-or”, defined on 32-bit words. The period of this generator is about 2^{250} and it uses 250 words of state per generator. For more information see, S. Kirkpatrick and E. Stoll, "A very fast shift-register sequence random number generator", Journal of Computational Physics, 40, 517-526 (1981)

Generator gsl rng tt800

This is an earlier version of the twisted generalized feedback shift-register generator, and has been superseded by the development of MT19937. However, it is still an acceptable generator in its own right. It has a period of 2^{800} and uses 33 words of storage per generator. For more information see,

Makoto Matsumoto and Yoshiharu Kurita, "Twisted GFSR Generators II", ACM Transactions on Modelling and Computer Simulation, Vol. 4, No. 3, 1994, pages 254-266.

Generator gsl rng vax

This is the VAX generator MTH$RANDOM. Its sequence is,

x_{n+1} = (ax_{n} + c) mod m

with a = 69069, c = 1 and m = 2^{32}. The seed specifies the initial value, x_{1}. The period of this generator is 2^{32} and it uses 1 word of storage per generator.

Generator gsl rng transputer

This is the random number generator from the INMOS Transputer Development system. Its sequence is,

x_{n+1} = (ax_{n}) mod m

with a = 1664525 and m = 2^{32}. The seed specifies the initial value, x_{1}.

Generator gsl rng randu

This is the IBM RANDU generator. Its sequence is

x_{n+1} = (ax_{n}) mod m

with a = 65539 and m = 2^{31}. The seed specifies the initial value, x_{1}. The period of this generator was only 2^{29}. It has become a textbook example of a poor generator.

Generator gsl rng minstd

This is Park and Miller’s "minimal standard" minstd generator, a simple linear congruence which takes care to avoid the major pitfalls of such algorithms. Its sequence is,

x_{n+1} = (ax_{n}) mod m

with a = 16807 and m = 2^{31} − 1 = 2147483647. The seed specifies the initial value, x_{1}. The period of this generator is about 2^{31}. This generator is used in the IMSL Library (subroutine RNUN) and in MATLAB (the RAND function). It is also sometimes known by the acronym "GGL" (I’m not sure what that stands for). For more information see,

Park and Miller, "Random Number Generators: Good ones are hard to find", Communications of the ACM, October 1988, Volume 31, No 10, pages 1192-1201.

Generator gsl rng uni Generator gsl rng uni32

This is a reimplementation of the 16-bit SLATEC random number generator RUNIF. A generalization of the generator to 32 bits is provided by gsl_rng_uni32. The original source code is available from NETLIB.

Generator gsl rng slatec

This is the SLATEC random number generator RAND. It is ancient. The original source code is available from NETLIB.

Generator gsl rng zuf

This is the ZUFALL lagged Fibonacci series generator of Peterson. Its sequence is,

\begin{aligned}t & = u_{n−273} + u_{n−607} \\ u_{n} & = t − floor(t) \end{aligned}

The original source code is available from NETLIB. For more information see, W. Petersen, "Lagged Fibonacci Random Number Generators for the NEC SX-3", International Journal of High Speed Computing (1994).

16.12 Random Number Generator Performance

The following table shows the relative performance of a selection the available random number generators. The simulation quality generators which offer the best performance are

taus, gfsr4 and mt19937.
1754 k ints/sec,    870 k doubles/sec, taus
1613 k ints/sec,    855 k doubles/sec, gfsr4
1370 k ints/sec,    769 k doubles/sec, mt19937
565 k ints/sec,    571 k doubles/sec, ranlxs0
400 k ints/sec,    405 k doubles/sec, ranlxs1
490 k ints/sec,    389 k doubles/sec, mrg
407 k ints/sec,    297 k doubles/sec, ranlux
243 k ints/sec,    254 k doubles/sec, ranlxd1
251 k ints/sec,    253 k doubles/sec, ranlxs2
238 k ints/sec,    215 k doubles/sec, cmrg
247 k ints/sec,    198 k doubles/sec, ranlux389
141 k ints/sec,    140 k doubles/sec, ranlxd2
1852 k ints/sec,    935 k doubles/sec, ran3
813 k ints/sec,    575 k doubles/sec, ran0
787 k ints/sec,    476 k doubles/sec, ran1
379 k ints/sec,    292 k doubles/sec, ran2

16.13 References and Further Reading

The subject of random number generation and testing is reviewed extensively in Knuth’s Seminumerical Algorithms.

Donald E. Knuth, The Art of Computer Programming: Seminumerical Algorithms (Vol 2, 3rd Ed, 1997), Addison-Wesley, ISBN 0201896842.

Further information is available in the review paper written by Pierre L’Ecuyer,

P. L’Ecuyer, “Random Number Generation”, Chapter 4 of the Handbook on Simulation, Jerry Banks Ed., Wiley, 1998, 93–137.

    http://www.iro.umontreal.ca/~lecuyer/papers.html in the file ‘handsim.ps’.

On the World Wide Web, see the pLab home page (http://random.mat.sbg.ac.at/) for a lot of information on the state-of-the-art in random number generation, and for numerous links to various "random" WWW sites. The source code for the diehard random number generator tests is also available online.

DIEHARD source code G. Marsaglia,

    http://stat.fsu.edu/pub/diehard/

16.14 Acknowledgements

Thanks to Makoto Matsumoto, Takuji Nishimura and Yoshiharu Kurita for making the source code to their generators (MT19937, MM&TN; TT800, MM&YK) available under the GNU General Public License. Thanks to Martin Lüscher for providing notes and source code for the ranlxs and ranlxd generators.

17 Quasi-Random Sequences

This chapter describes functions for generating quasi-random sequences in arbitrary dimensions. A quasi-random sequence progressively covers a d-dimensional space with a set of points that are uniformly distributed. Quasi-random sequences are also known as low-discrepancy sequences. The quasi-random sequence generators use an interface that is similar to the interface for random number generators.

The functions described in this section are declared in the header file ‘gsl_qrng.h’.

17.1 Quasi-random number generator initialization

Function gsl_qrng * gsl qrng alloc (const gsl_qrng_type * T, unsigned int

d) This function returns a pointer to a newly-created instance of a quasi-random sequence generator of type T and dimension d. If there is insufficient memory to create the generator then the function returns a null pointer and the error handler is invoked with an error code of GSL_ENOMEM.

Function void gsl qrng free (gsl_qrng * q)

This function frees all the memory associated with the generator q.

Function void gsl qrng init (const gsl_qrng * q)

This function reinitializes the generator q to its starting point.

17.2 Sampling from a quasi-random number generator

Function int gsl qrng get (const gsl_qrng * q, double x[])

This function returns the next point x from the sequence generator q. The space available for x must match the dimension of the generator. The point x will lie in the range 0 < x_{i} < 1 for each x_{i}.

17.3 Auxiliary quasi-random number generator functions

Function const char * gsl qrng name (const gsl_qrng * q)

This function returns a pointer to the name of the generator.

Function size_t gsl qrng size (const gsl_qrng * q) Function void * gsl qrng state (const gsl_qrng * q)

These function return a pointer to the state of generator r and its size. You can use this information to access the state directly. For example, the following code will write the state of a generator to a stream,

     void * state = gsl_qrng_state (q);
     size_t n = gsl_qrng_size (q);
     fwrite (state, n, 1, stream);

17.4 Saving and resorting quasi-random number generator state

Function int gsl qrng memcpy (gsl_qrng * dest, const gsl_qrng * src)

This function copies the quasi-random sequence generator src into the pre-existing generator dest, making dest into an exact copy of src. The two generators must be of the same type.

Function gsl_qrng * gsl qrng clone (const gsl_qrng * q)

This function returns a pointer to a newly created generator which is an exact copy of the generator r.

17.5 Quasi-random number generator algorithms

The following quasi-random sequence algorithms are available,

Generator gsl qrng niederreiter 2

This generator uses the algorithm described in Bratley, Fox, Niederreiter, ACM Trans. Model. Comp. Sim. 2, 195 (1992). It is valid up to 12 dimensions.

Generator gsl qrng sobol

This generator uses the Sobol sequence described in Antonov, Saleev, USSR Comput. Maths. Math. Phys. 19, 252 (1980). It is valid up to 40 dimensions.

17.6 Examples

The following program prints the first 1024 points of the 2-dimensional Sobol sequence.

#include <stdio.h>
#include <gsl/gsl_qrng.h>
int
main (void)
{
  int i;
  gsl_qrng * q = gsl_qrng_alloc (gsl_qrng_sobol, 2);
  for (i = 0; i < 1024; i++)
    {
      double v[2];
      gsl_qrng_get(q, v);
      printf("%.5f %.5f\n", v[0], v[1]);
    }
  gsl_qrng_free(q);
  return 0;
}

Here is the output from the program,

$ ./a.out
0.50000 0.50000
0.75000 0.25000
0.25000 0.75000
0.37500 0.37500
0.87500 0.87500
0.62500 0.12500
0.12500 0.62500
....

It can be seen that successive points progressively fill-in the spaces between previous points. The following plot shows the distribution in the x-y plane of the first 1024 points from the Sobol sequence,

Distribution of the first 1024 points from the quasi-random Sobol sequence

17.7 References

The implementations of the quasi-random sequence routines are based on the algorithms described in the following paper,

P. Bratley and B.L. Fox and H. Niederreiter, “Algorithm 738: Programs to Generate Niederreiter’s Low-discrepancy Sequences”, Transactions on Mathematical Software, Vol. 20, No. 4, December, 1994, p. 494-495.

18 Random Number Distributions

This chapter describes functions for generating random variates and computing their probability distributions. Samples from the distributions described in this chapter can be obtained using any of the random number generators in the library as an underlying source of randomness. In the simplest cases a non-uniform distribution can be obtained analytically from the uniform distribution of a random number generator by applying an appropriate transformation. This method uses one call to the random number generator.

More complicated distributions are created by the acceptance-rejection method, which compares the desired distribution against a distribution which is similar and known analytically. This usually requires several samples from the generator.

The functions described in this section are declared in ‘gsl_randist.h’.

18.1 The Gaussian Distribution

Random double gsl ran gaussian (const gsl_rng * r, double sigma)

This function returns a Gaussian random variate, with mean zero and standard deviation sigma. The probability distribution for Gaussian random variates is,

p(x)dx = \frac{1}{\sqrt{2πσ^{2}}} exp(−x^{2}/2σ^{2})dx

for x in the range −∞ to +∞. Use the transformation z = μ + x on the numbers returned by gsl_ran_gaussian to obtain a Gaussian distribution with mean μ. This function uses the Box-Mueller algorithm which requires two calls the random number generator r.

Function double gsl ran gaussian pdf (double x, double sigma)

This function computes the probability density p(x) at x for a Gaussian distribution with standard deviation sigma, using the formula given above.

σ = 2 σ = 1 Gaussian Distribution x p(x)  5 4 3 2 1 0 -1 -2 -3 -4 -5 0.5 0.4 0.3 0.2 0.1 0

Function double gsl ran gaussian ratio method (const gsl_rng * r, const

    double sigma)

This function computes a gaussian random variate using the Kinderman-Monahan ratio method.

Random double gsl ran ugaussian (const gsl_rng * r) Random double gsl ran ugaussian pdf (double x) Random double gsl ran ugaussian ratio method (const gsl_rng * r,

    const double sigma)

These functions compute results for the unit Gaussian distribution. They are equivalent to the functions above with a standard deviation of one, sigma = 1.

18.2 The Gaussian Tail Distribution

Random double gsl ran gaussian tail (const gsl_rng * r, double a, double

sigma) This function provides random variates from the upper tail of a Gaussian distribution with standard deviation sigma. The values returned are larger than the lower limit a, which must be positive. The method is based on Marsaglia’s famous rectanglewedge-tail algorithm (Ann Math Stat 32, 894-899 (1961)), with this aspect explained in Knuth, v2, 3rd ed, p139,586 (exercise 11). The probability distribution for Gaussian tail random variates is,

p(x)dx = \frac{1}{N(a; σ)} exp(−x^{2}/2σ^{2})dx

for x > a where N(a; σ) is the normalization constant,

N(a; σ) = \frac{1}{2}erfc (\frac{a}{\sqrt{2σ^{2}}} ) .

Function double gsl ran gaussian tail pdf (double x, double a, double

sigma) This function computes the probability density p(x) at x for a Gaussian tail distribution with standard deviation sigma and lower limit a, using the formula given above.

σ = 1, a = 1.5 Gaussian Tail Distribution x p(x)  5 4 3 2 1 0 0.2 0.1 0

Random double gsl ran ugaussian tail (const gsl_rng * r, double a) Random double gsl ran ugaussian tail pdf (double x, double a)

These functions compute results for the tail of a unit Gaussian distribution. They are equivalent to the functions above with a standard deviation of one, sigma = 1.

18.3 The Bivariate Gaussian Distribution

Random void gsl ran bivariate gaussian (const gsl_rng * r, double

    sigma x, double sigma y, double rho, double * x, double * y)

This function generates a pair of correlated gaussian variates, with mean zero, correlation coefficient rho and standard deviations sigma x and sigma y in the x and y directions. The probability distribution for bivariate gaussian random variates is,

p(x, y)dxdy = \frac{1}{2πσ_{x}σ_{y}\sqrt{1 − ρ^{2}}} exp (\frac{−(x^{2} + y^{2} − 2ρxy)}{2σ_{x}^{2}σ_{y}^{2}(1 − ρ^{2})}) dxdy

for x, y in the range −∞ to +∞. The correlation coefficient rho should lie between 1 and −1.

Function double gsl ran bivariate gaussian pdf (double x, double y,

    double sigma x, double sigma y, double rho)

This function computes the probability density p(x, y) at (x,y) for a bivariate gaussian distribution with standard deviations sigma x, sigma y and correlation coefficient rho, using the formula given above.

y 2 1 0 -1 -2 x  2 1 0 -1 -2  σ_{x} = 1, σ_{y} = 1, ρ = 0.9 Bivariate Gaussian Distribution

18.4 The Exponential Distribution

Random double gsl ran exponential (const gsl_rng * r, double mu)

This function returns a random variate from the exponential distribution with mean mu. The distribution is,

p(x)dx = \frac{1}{μ} exp(−x/μ)dx

for x ≥ 0.

Function double gsl ran exponential pdf (double x, double mu)

This function computes the probability density p(x) at x for an exponential distribution with mean mu, using the formula given above.

μ = 2 μ = 1 Exponential Distribution x p(x)  3 2 1 0 1 0.5 0

18.5 The Laplace Distribution

Random double gsl ran laplace (const gsl_rng * r, double a)

This function returns a random variate from the the Laplace distribution with width a. The distribution is,

p(x)dx = \frac{1}{2a} exp(−|x/a|)dx

for −∞ < x < ∞.

Function double gsl ran laplace pdf (double x, double a)

This function computes the probability density p(x) at x for a Laplace distribution with mean a, using the formula given above.

a = 2 a = 1 Laplace Distribution (Two-sided Exponential) x p(x)  5 4 3 2 1 0 -1 -2 -3 -4 -5 0.5 0.4 0.3 0.2 0.1 0

18.6 The Exponential Power Distribution

Random double gsl ran exppow (const gsl_rng * r, double a, double b)

This function returns a random variate from the exponential power distribution with scale parameter a and exponent b. The distribution is,

p(x)dx = \frac{1}{2aΓ(1 + 1/b)} exp(−|x/a|^{b})dx

for x ≥ 0. For b = 1 this reduces to the Laplace distribution. For b = 2 it has the same form as a gaussian distribution, but with a = \sqrt{2}σ.

Function double gsl ran exppow pdf (double x, double a, double b)

This function computes the probability density p(x) at x for an exponential power distribution with scale parameter a and exponent b, using the formula given above.

a = 2, b = 0.5 a = 1, b = 2.5 Exponential Power Distribution x p(x)  5 4 3 2 1 0 -1 -2 -3 -4 -5 0.8 0.7 0.6 0.5 0.4 0.3 0.2 0.1 0

18.7 The Cauchy Distribution

Random double gsl ran cauchy (const gsl_rng * r, double a)

This function returns a random variate from the Cauchy distribution with scale parameter a. The probability distribution for Cauchy random variates is,

p(x)dx = \frac{1}{aπ(1 + (x/a)^{2})}dx

for x in the range −∞ to +∞. The Cauchy distribution is also known as the Lorentz distribution.

Function double gsl ran cauchy pdf (double x, double a)

This function computes the probability density p(x) at x for a Cauchy distribution with scale parameter a, using the formula given above.

a = 2 a = 1 Cauchy Distribution x p(x)  5 4 3 2 1 0 -1 -2 -3 -4 -5 0.4 0.3 0.2 0.1 0

18.8 The Rayleigh Distribution

Random double gsl ran rayleigh (const gsl_rng * r, double sigma)

This function returns a random variate from the Rayleigh distribution with scale parameter sigma. The distribution is,

p(x)dx = \frac{x}{σ^{2}} exp(−x^{2}/(2σ^{2}))dx

for x > 0.

Function double gsl ran rayleigh pdf (double x, double sigma)

This function computes the probability density p(x) at x for a Rayleigh distribution with scale parameter sigma, using the formula given above.

σ = 2 σ = 1 Rayleigh Distribution x p(x)  5 4 3 2 1 0 0.7 0.6 0.5 0.4 0.3 0.2 0.1 0

18.9 The Rayleigh Tail Distribution

Random double gsl ran rayleigh tail (const gsl_rng * r, double a double

sigma) This function returns a random variate from the tail of the Rayleigh distribution with scale parameter sigma and a lower limit of a. The distribution is,

p(x)dx = \frac{x}{σ^{2}} exp((a^{2} − x^{2})/(2σ^{2}))dx

for x > a.

Function double gsl ran rayleigh tail pdf (double x, double a, double

sigma) This function computes the probability density p(x) at x for a Rayleigh tail distribution with scale parameter sigma and lower limit a, using the formula given above.

a = 0.5, σ = 2 a = 1, σ = 1 Rayleigh Tail Distribution x p(x)  5 4 3 2 1 0 1 0.9 0.8 0.7 0.6 0.5 0.4 0.3 0.2 0.1 0

18.10 The Landau Distribution

Random double gsl ran landau (const gsl_rng * r)

This function returns a random variate from the Landau distribution. The probability distribution for Landau random variates is defined analytically by the complex integral,

p(x) = \frac{1}{2πi} ∫_{c−i∞}^{c+i∞} ds exp(s log(s) + xs)

For numerical purposes it is more convenient to use the following equivalent form of the integral,

p(x) = (1/π) ∫_{0}^{∞} dt exp(−t log(t) − xt) sin(πt).

Function double gsl ran landau pdf (double x)

This function computes the probability density p(x) at x for the Landau distribution using an approximation to the formula given above.

Landau Distribution x p(x)  10 9 8 7 6 5 4 3 2 1 0 -1 -2 -3 -4 -5 0.2 0.1 0

18.11 The Levy alpha-Stable Distributions

Random double gsl ran levy (const gsl_rng * r, double c, double alpha)

This function returns a random variate from the Levy symmetric stable distribution with scale c and exponent alpha. The symmetric stable probability distribution is defined by a fourier transform,

p(x) = \frac{1}{2π} ∫_{−∞}^{+∞} dt exp(−itx − |ct|^{α})

There is no explicit solution for the form of p(x) and the library does not define a corresponding pdf function. For α = 1 the distribution reduces to the Cauchy distribution. For α = 2 it is a Gaussian distribution with σ = \sqrt{2}c. For α < 1 the tails of the distribution become extremely wide. The algorithm only works for 0 < α ≤ 2.

c = 1, α = 2.0 c = 1, α = 1.0 Levy Distribution x p(x)  5 4 3 2 1 0 -1 -2 -3 -4 -5 0.4 0.3 0.2 0.1 0

18.12 The Levy skew alpha-Stable Distribution

Random double gsl ran levy skew (const gsl_rng * r, double c, double

alpha, double beta) This function returns a random variate from the Levy skew stable distribution with scale c, exponent alpha and skewness parameter beta. The skewness parameter must lie in the range [−1, 1]. The Levy skew stable probability distribution is defined by a fourier transform,

p(x) = \frac{1}{2π} ∫_{−∞}^{+∞} dt exp(−itx − |ct|^{α}(1 − iβsign(t) tan(πα/2)))

When α = 1 the term tan(πα/2) is replaced by −(2/π) log |t|. There is no explicit solution for the form of p(x) and the library does not define a corresponding pdf function. For α = 2 the distribution reduces to a Gaussian distribution with σ = \sqrt{2}c and the skewness parameter has no effect. For α < 1 the tails of the distribution become extremely wide. The symmetric distribution corresponds to β = 0. The algorithm only works for 0 < α ≤ 2.

The Levy alpha-stable distributions have the property that if N alpha-stable variates are drawn from the distribution p(c, α, β) then the sum Y = X_{1} + X_{2} + … + X_{N} will also be distributed as an alpha-stable variate, p(N^{1/α}c, α, β).

c = 1, α = 1.0, β = 1.0 Levy Skew Distribution x p(x)  5 4 3 2 1 0 -1 -2 -3 -4 -5 0

18.13 The Gamma Distribution

Random double gsl ran gamma (const gsl_rng * r, double a, double b)

This function returns a random variate from the gamma distribution. The distribution function is,

p(x)dx = \frac{1}{Γ(a)b^{a}} x^{a−1}e^{−x/b}dx

for x > 0.

Function double gsl ran gamma pdf (double x, double a, double b)

This function computes the probability density p(x) at x for a gamma distribution with parameters a and b, using the formula given above.

a = 3 a = 2 a = 1 Gamma Distribution x p(x)  5 4 3 2 1 0 1 0.5 0

18.14 The Flat (Uniform) Distribution

Random double gsl ran flat (const gsl_rng * r, double a, double b)

This function returns a random variate from the flat (uniform) distribution from a to b. The distribution is,

p(x)dx = \frac{1}{(b − a)}dx

if a ≤ x < b and 0 otherwise.

Function double gsl ran flat pdf (double x, double a, double b)

This function computes the probability density p(x) at x for a uniform distribution from a to b, using the formula given above.

a = 1.2, b = 4.8 a = 0.5, b = 2.5 Flat Distribution x p(x)  5 4 3 2 1 0 1 0.5 0

18.15 The Lognormal Distribution

Random double gsl ran lognormal (const gsl_rng * r, double zeta, double

sigma) This function returns a random variate from the lognormal distribution. The distribution function is,

p(x)dx = \frac{1}{x\sqrt{2πσ^{2}}} exp(−(ln(x) − ζ)^{2}/2σ^{2})dx

for x > 0.

Function double gsl ran lognormal pdf (double x, double zeta, double

sigma) This function computes the probability density p(x) at x for a lognormal distribution with parameters zeta and sigma, using the formula given above.

ζ = 1, σ = 1 ζ = 0, σ = 1 Lognormal Distribution x p(x)  3 2 1 0 0.7 0.6 0.5 0.4 0.3 0.2 0.1 0

18.16 The Chi-squared Distribution

The chi-squared distribution arises in statistics If Y_{i} are n independent gaussian random variates with unit variance then the sum-of-squares,

X_{i} = ∑_{i} Y_{i}^{2}

has a chi-squared distribution with n degrees of freedom.

Random double gsl ran chisq (const gsl_rng * r, double nu)

This function returns a random variate from the chi-squared distribution with nu degrees of freedom. The distribution function is,

p(x)dx = \frac{1}{Γ(ν/2)}(x/2)^{ν/2−1} exp(−x/2)dx

for x ≥ 0.

Function double gsl ran chisq pdf (double x, double nu)

This function computes the probability density p(x) at x for a chi-squared distribution with nu degrees of freedom, using the formula given above.

ν = 3 ν = 2 ν = 1 Chi-squared Distribution x p(x)  3 2 1 0 1 0.5 0

18.17 The F-distribution

The F-distribution arises in statistics. If Y_{1} and Y_{2} are chi-squared deviates with ν_{1} and ν_{2} degrees of freedom then the ratio,

X = \frac{(Y_{1}/ν_{1})}{(Y_{2}/ν_{2})}

has an F-distribution F (x; ν_{1}, ν_{2}).

Random double gsl ran fdist (const gsl_rng * r, double nu1, double nu2)

This function returns a random variate from the F-distribution with degrees of freedom nu1 and nu2. The distribution function is,

p(x)dx = \frac{Γ((ν_{1} + ν_{2})/2)}{Γ(ν_{1}/2)Γ(ν_{2}/2)}ν_{1}^{ν_{1}/2}ν_{2}^{ν_{2}/2}x^{ν_{1}/2−1}(ν_{2} + ν_{1}x)^{−ν_{1}/2−ν_{2}/2}

for x ≥ 0.

Function double gsl ran fdist pdf (double x, double nu1, double nu2)

This function computes the probability density p(x) at x for an F-distribution with nu1 and nu2 degrees of freedom, using the formula given above.

ν_{1} = 1, ν_{2} = 2 ν_{1} = 1, ν_{2} = 1 F-Distribution x p(x)  2 1 0 1 0.5 0

18.18 The t-distribution

The t-distribution arises in statistics. If Y_{1} has a normal distribution and Y_{2} has a chi-squared distribution with ν degrees of freedom then the ratio,

X = \frac{Y_{1}}{\sqrt{Y_{2}/ν}}

has a t-distribution t(x; ν) with ν degrees of freedom.

Random double gsl ran tdist (const gsl_rng * r, double nu)

This function returns a random variate from the t-distribution. The distribution function is,

p(x)dx = \frac{Γ((ν + 1)/2)}{\sqrt{πν}Γ(ν/2)}(1 + x^{2}/ν)^{−(ν+1)/2}dx

for −∞ < x < +∞.

Function double gsl ran tdist pdf (double x, double nu)

This function computes the probability density p(x) at x for a t-distribution with nu degrees of freedom, using the formula given above.

ν_{1} = 5 ν_{1} = 1 Student’s t distribution x p(x)  4 3 2 1 0 -1 -2 -3 -4 0.5 0.4 0.3 0.2 0.1 0

18.19 The Beta Distribution

Random double gsl ran beta (const gsl_rng * r, double a, double b)

This function returns a random variate from the beta distribution. The distribution function is,

p(x)dx = \frac{Γ(a + b)}{Γ(a)Γ(b)}x^{a−1}(1 − x)^{b−1}dx

for 0 ≤ x ≤ 1.

Function double gsl ran beta pdf (double x, double a, double b)

This function computes the probability density p(x) at x for a beta distribution with parameters a and b, using the formula given above.

a = 1, b = 4 a = 4, b = 1 a = 2, b = 2 Beta Distribution x p(x)  1 0.75 0.5 0.25 0 4 3 2 1 0

18.20 The Logistic Distribution

Random double gsl ran logistic (const gsl_rng * r, double a)

This function returns a random variate from the logistic distribution. The distribution function is,

p(x)dx = \frac{exp(−x/a)}{a(1 + exp(−x/a))^{2}} dx

for −∞ < x < +∞.

Function double gsl ran logistic pdf (double x, double a)

This function computes the probability density p(x) at x for a logistic distribution with scale parameter a, using the formula given above.

a = 2 a = 1 Logistic Distribution x p(x)  5 4 3 2 1 0 -1 -2 -3 -4 -5 0.3 0.2 0.1 0

18.21 The Pareto Distribution

Random double gsl ran pareto (const gsl_rng * r, double a, double b)

This function returns a random variate from the Pareto distribution of order a. The distribution function is,

p(x)dx = (a/b)/(x/b)^{a+1}dx

for x ≥ b.

Function double gsl ran pareto pdf (double x, double a, double b)

This function computes the probability density p(x) at x for a Pareto distribution with exponent a and scale b, using the formula given above.

a = 3, b = 2 a = 1, b = 1 Pareto Distribution x p(x)  5 4 3 2 1 0 2 1.5 1 0.5 0

18.22 The Spherical Distribution (2D & 3D)

The spherical distributions generate random vectors, located on a spherical surface. They can be used as random directions, for example in the steps of a random walk.

Random void gsl ran dir 2d (const gsl_rng * r, double *x, double *y)

Random void gsl ran dir 2d trig method (const gsl_rng * r, double *x,

    double *y)

This function returns a random direction vector v = (x,y) in two dimensions. The vector is normalized such that |v|^{2} = x^{2} + y^{2} = 1. The obvious way to do this is to take a uniform random number between 0 and 2π and let x and y be the sine and cosine respectively. Two trig functions would have been expensive in the old days, but with modern hardware implementations, this is sometimes the fastest way to go. This is the case for my home Pentium (but not the case for my Sun Sparcstation 20 at work). Once can avoid the trig evaluations by choosing x and y in the interior of a unit circle (choose them at random from the interior of the enclosing square, and then reject those that are outside the unit circle), and then dividing by \sqrt{x^{2} + y^{2}}. A much cleverer approach, attributed to von Neumann (See Knuth, v2, 3rd ed, p140, exercise 23), requires neither trig nor a square root. In this approach, u and v are chosen at random from the interior of a unit circle, and then x = (u^{2} − v^{2})/(u^{2} + v^{2}) and y = uv/(u^{2} + v^{2}).

Random void gsl ran dir 3d (const gsl_rng * r, double *x, double *y,
    double * z)

This function returns a random direction vector v = (x,y,z) in three dimensions. The vector is normalized such that |v|^{2} = x^{2} +y^{2} +z^{2} = 1. The method employed is due to Robert E. Knop (CACM 13, 326 (1970)), and explained in Knuth, v2, 3rd ed, p136. It uses the surprising fact that the distribution projected along any axis is actually uniform (this is only true for 3d).

Random void gsl ran dir nd (const gsl_rng * r, int n, double *x)

This function returns a random direction vector v = (x_{1}, x_{2}, …, x_{n}) in n dimensions. The vector is normalized such that |v|^{2} = x_{1}^{2} + x_{2}^{2} + ⋯ + x_{n}^{2} = 1. The method uses the fact that a multivariate gaussian distribution is spherically symmetric. Each component is generated to have a gaussian distribution, and then the components are normalized. The method is described by Knuth, v2, 3rd ed, p135-136, and attributed to G. W. Brown, Modern Mathematics for the Engineer (1956).

18.23 The Weibull Distribution

Random double gsl ran weibull (const gsl_rng * r, double a, double b)

This function returns a random variate from the Weibull distribution. The distribution function is,

p(x)dx = \frac{b}{a^{b}} x^{b−1} exp(−(x/a)^{b})dx

for x ≥ 0.

Function double gsl ran weibull pdf (double x, double a, double b)

This function computes the probability density p(x) at x for a Weibull distribution with scale a and exponent b, using the formula given above.

a = 2, b = 3 a = 1, b = 2 a = 1, b = 1 Weibull Distribution x p(x)  2 1.5 1 0.5 0 1.5 1 0.5 0

18.24 The Type-1 Gumbel Distribution

Random double gsl ran gumbel1 (const gsl_rng * r, double a, double b)

This function returns a random variate from the Type-1 Gumbel distribution. The Type-1 Gumbel distribution function is,

p(x)dx = ab exp(−(b exp(−ax) + ax))dx

for −∞ < x < ∞.

Function double gsl ran gumbel1 pdf (double x, double a, double b)

This function computes the probability density p(x) at x for a Type-1 Gumbel distribution with parameters a and b, using the formula given above.

Type 1, a = 1, b = 1 Type 1 Gumbel Distribution x p(x)  2 1.5 1 0.5 0 -0.5 -1 -1.5 -2 0.7 0.6 0.5 0.4 0.3 0.2 0.1 0

18.25 The Type-2 Gumbel Distribution

Random double gsl ran gumbel2 (const gsl_rng * r, double a, double b)

This function returns a random variate from the Type-2 Gumbel distribution. The Type-2 Gumbel distribution function is,

p(x)dx = abx^{−a−1} exp(−bx^{−a})dx

for 0 < x < ∞.

Function double gsl ran gumbel2 pdf (double x, double a, double b)

This function computes the probability density p(x) at x for a Type-2 Gumbel distribution with parameters a and b, using the formula given above.

Type 2, a = 1, b = 1 Type 2 Gumbel Distribution x p(x)  2 1.5 1 0.5 0 0.7 0.6 0.5 0.4 0.3 0.2 0.1 0

18.26 General Discrete Distributions

Given K discrete events with different probabilities P [k], produce a random value k consistent with its probability.

The obvious way to do this is to preprocess the probability list by generating a cumulative probability array with K + 1 elements:

\begin{aligned}C[0] & = 0 \\ C[k + 1] & = C[k] + P [k]. \end{aligned}

Note that this construction produces C[K] = 1. Now choose a uniform deviate u between 0 and 1, and find the value of k such that C[k] ≤ u < C[k + 1]. Although this in principle requires of order log K steps per random number generation, they are fast steps, and if you use something like ⌊uK⌋ as a starting point, you can often do pretty well.

But faster methods have been devised. Again, the idea is to preprocess the probability list, and save the result in some form of lookup table; then the individual calls for a random discrete event can go rapidly. An approach invented by G. Marsaglia (Generating discrete random numbers in a computer, Comm ACM 6, 37-38 (1963)) is very clever, and readers interested in examples of good algorithm design are directed to this short and well-written paper. Unfortunately, for large K, Marsaglia’s lookup table can be quite large.

A much better approach is due to Alastair J. Walker (An efficient method for generating discrete random variables with general distributions, ACM Trans on Mathematical Software 3, 253-256 (1977); see also Knuth, v2, 3rd ed, p120-121,139). This requires two lookup tables, one floating point and one integer, but both only of size K. After preprocessing, the random numbers are generated in O(1) time, even for large K. The preprocessing suggested by Walker requires O(K^{2}) effort, but that is not actually necessary, and the implementation provided here only takes O(K) effort. In general, more preprocessing leads to faster generation of the individual random numbers, but a diminishing return is reached pretty early. Knuth points out that the optimal preprocessing is combinatorially difficult for large K.

This method can be used to speed up some of the discrete random number generators below, such as the binomial distribution. To use if for something like the Poisson Distribution, a modification would have to be made, since it only takes a finite set of K outcomes.

Random gsl_ran_discrete_t * gsl ran discrete preproc (size_t K,

         const double * P)

This function returns a pointer to a structure that contains the lookup table for the discrete random number generator. The array P[] contains the probabilities of the discrete events; these array elements must all be positive, but they needn’t add up to one (so you can think of them more generally as "weights") – the preprocessor will normalize appropriately. This return value is used as an argument for the gsl_ran_ discrete function below.

Random size_t gsl ran discrete (const gsl_rng * r, const
         gsl_ran_discrete_t * g)

After the preprocessor, above, has been called, you use this function to get the discrete random numbers.

Random double gsl ran discrete pdf (size_t k, const gsl_ran_discrete_t

* g) Returns the probability P [k] of observing the variable k. Since P [k] is not stored as part of the lookup table, it must be recomputed; this computation takes O(K), so if K is large and you care about the original array P [k] used to create the lookup table, then you should just keep this original array P [k] around.

Random void gsl ran discrete free (gsl_ran_discrete_t * g)

De-allocates the lookup table pointed to by g.

18.27 The Poisson Distribution

Random unsigned int gsl ran poisson (const gsl_rng * r, double mu)

This function returns a random integer from the Poisson distribution with mean mu. The probability distribution for Poisson variates is,

p(k) = \frac{μ^{k}}{k!} exp(−μ)

for k ≥ 0.

Function double gsl ran poisson pdf (unsigned int k, double mu)

This function computes the probability p(k) of obtaining k from a Poisson distribution with mean mu, using the formula given above.

μ = 2.5 Poisson Distribution k p(k)  10 9 8 7 6 5 4 3 2 1 0 0.3 0.2 0.1 0

18.28 The Bernoulli Distribution

Random unsigned int gsl ran bernoulli (const gsl_rng * r, double p)

This function returns either 0 or 1, the result of a Bernoulli trial with probability p. The probability distribution for a Bernoulli trial is,

\begin{aligned}p(0) & = 1 − p \\ p(1) & = p \end{aligned}

Function double gsl ran bernoulli pdf (unsigned int k, double p)

This function computes the probability p(k) of obtaining k from a Bernoulli distribution with probability parameter p, using the formula given above.

p = 0.7 Bernoulli Trial k p(k)  1 0 1 0.9 0.8 0.7 0.6 0.5 0.4 0.3 0.2 0.1 0

18.29 The Binomial Distribution

Random unsigned int gsl ran binomial (const gsl_rng * r, double p,
    unsigned int n)

This function returns a random integer from the binomial distribution, the number of successes in n independent trials with probability p. The probability distribution for binomial variates is,

p(k) = \frac{n!}{k!(n − k)!}p^{k}(1 − p)^{n−k}

for 0 ≤ k ≤ n.

Function double gsl ran binomial pdf (unsigned int k, double p, unsigned
    int n)

This function computes the probability p(k) of obtaining k from a binomial distribution with parameters p and n, using the formula given above.

p = 0.5, n = 9 Binomial Distribution k p(k)  10 9 8 7 6 5 4 3 2 1 0 0.3 0.2 0.1 0

18.30 The Negative Binomial Distribution

Random unsigned int gsl ran negative binomial (const gsl_rng * r,

    double p, double n)

This function returns a random integer from the negative binomial distribution, the number of failures occurring before n successes in independent trials with probability p of success. The probability distribution for negative binomial variates is,

p(k) = \frac{Γ(n + k)}{Γ(k + 1)Γ(n)}p^{n}(1 − p)^{k}

Note that n is not required to be an integer.

Function double gsl ran negative binomial pdf (unsigned int k, double

    p, double n)

This function computes the probability p(k) of obtaining k from a negative binomial distribution with parameters p and n, using the formula given above.

p = 0.5, n = 3.5 Negative Binomial Distribution k p(k)  10 9 8 7 6 5 4 3 2 1 0 0.3 0.2 0.1 0

18.31 The Pascal Distribution

Random unsigned int gsl ran pascal (const gsl_rng * r, double p,
    unsigned int k)

This function returns a random integer from the Pascal distribution. The Pascal distribution is simply a negative binomial distribution with an integer value of n.

p(k) = \frac{(n + k − 1)!}{k!(n − 1)!}p^{n}(1 − p)^{k}

for k ≥ 0

Function double gsl ran pascal pdf (unsigned int k, double p, unsigned
    int n)

This function computes the probability p(k) of obtaining k from a Pascal distribution with parameters p and n, using the formula given above.

p = 0.5, n = 3 Pascal Distribution k p(k)  10 9 8 7 6 5 4 3 2 1 0 0.3 0.2 0.1 0

18.32 The Geometric Distribution

Random unsigned int gsl ran geometric (const gsl_rng * r, double p)

This function returns a random integer from the geometric distribution, the number of independent trials with probability p until the first success. The probability distribution for geometric variates is,

p(k) = p(1 − p)^{k−1}

for k ≥ 1.

Function double gsl ran geometric pdf (unsigned int k, double p)

This function computes the probability p(k) of obtaining k from a geometric distribution with probability parameter p, using the formula given above.

p = 0.5 Geometric Distribution k p(k)  5 4 3 2 1 0 0.7 0.6 0.5 0.4 0.3 0.2 0.1 0

18.33 The Hypergeometric Distribution

Random unsigned int gsl ran hypergeometric (const gsl_rng * r,

    unsigned int n1, unsigned int n2, unsigned int t)

This function returns a random integer from the hypergeometric distribution. The probability distribution for hypergeometric random variates is,

p(k) = C(n_{1}, k)C(n_{2}, t − k)/C(n_{1} + n_{2}, k)

where C(a, b) = a!/(b!(a − b)!). The domain of k is max(0 t − n_{2}), …, max(t, n_{1}).

Function double gsl ran hypergeometric pdf (unsigned int k, unsigned

    int n1, unsigned int n2, unsigned int t)

This function computes the probability p(k) of obtaining k from a hypergeometric distribution with parameters n1, n2, n3, using the formula given above.

n1 = 5, n2 = 20, t = 3 Hypergeometric Distribution k p(k)  10 9 8 7 6 5 4 3 2 1 0 0.7 0.6 0.5 0.4 0.3 0.2 0.1 0

18.34 The Logarithmic Distribution

Random unsigned int gsl ran logarithmic (const gsl_rng * r, double p)

This function returns a random integer from the logarithmic distribution. The probability distribution for logarithmic random variates is,

p(k) = \frac{−1}{log(1 − p)}(\frac{p^{k}}{k} )

for k ≥ 1.

Function double gsl ran logarithmic pdf (unsigned int k, double p)

This function computes the probability p(k) of obtaining k from a logarithmic distribution with probability parameter p, using the formula given above.

p = 0.7 Logarithmic Distribution k p(k)  10 9 8 7 6 5 4 3 2 1 0 0.7 0.6 0.5 0.4 0.3 0.2 0.1 0

18.35 Shuffling and Sampling

The following functions allow the shuffling and sampling of a set of objects. The algorithms rely on a random number generator as source of randomness and a poor quality generator can lead to correlations in the output. In particular it is important to avoid generators with a short period. For more information see Knuth, v2, 3rd ed, Section 3.4.2, “Random Sampling and Shuffling”.

Random void gsl ran shuffle (const gsl_rng * r, void * base, size_t n,
    size_t size)

This function randomly shuffles the order of n objects, each of size size, stored in the array base[0..n-1]. The output of the random number generator r is used to produce the permutation. The algorithm generates all possible n! permutations with equal probability, assuming a perfect source of random numbers. The following code shows how to shuffle the numbers from 0 to 51,

     int a[52];
     for (i = 0; i < 52; i++)
       {
         a[i] = i;
       }
     gsl_ran_shuffle (r, a, 52, sizeof (int));
Random int gsl ran choose (const gsl_rng * r, void * dest, size_t k, void
    * src, size_t n, size_t size)

This function fills the array dest[k] with k objects taken randomly from the n elements of the array src[0..n-1]. The objects are each of size size. The output of the random number generator r is used to make the selection. The algorithm ensures all possible samples are equally likely, assuming a perfect source of randomness. The objects are sampled without replacement, thus each object can only appear once in dest[k]. It is required that k be less than or equal to n. The objects in dest will be in the same relative order as those in src. You will need to call gsl_ran_shuffle(r, dest, n, size) if you want to randomize the order. The following code shows how to select a random sample of three unique numbers from the set 0 to 99,

     double a[3], b[100];
     for (i = 0; i < 100; i++)
       {
         b[i] = (double) i;
       }
     gsl_ran_choose (r, a, 3, b, 100, sizeof (double));

Random void gsl ran sample (const gsl_rng * r, void * dest, size_t k,
  void * src, size_t n, size_t size)

This function is like gsl_ran_choose but samples k items from the original array of n items src with replacement, so the same object can appear more than once in the output sequence dest. There is no requirement that k be less than n in this case.

18.36 Examples

The following program demonstrates the use of a random number generator to produce variates from a distribution. It prints 10 samples from the Poisson distribution with a mean of 3.

#include <stdio.h>
#include <gsl/gsl_rng.h>
#include <gsl/gsl_randist.h>
int
main (void)
{
const gsl_rng_type * T;
gsl_rng * r;
int i, n = 10;
double mu = 3.0;
/* create a generator chosen by the
   environment variable GSL_RNG_TYPE */
gsl_rng_env_setup();
T = gsl_rng_default;
r = gsl_rng_alloc (T);
/* print n random variates chosen from
   the poisson distribution with mean
   parameter mu */
for (i = 0; i < n; i++)
  {
    unsigned int k = gsl_ran_poisson (r, mu);
    printf(" %u", k);
  }
printf("\n");
return 0;
}

If the library and header files are installed under ‘/usr/local’ (the default location) then the program can be compiled with these options,

gcc demo.c -lgsl -lgslcblas -lm

Here is the output of the program,

$ ./a.out
 4 2 3 3 1 3 4 1 3 5

The variates depend on the seed used by the generator. The seed for the default generator type gsl_rng_default can be changed with the GSL_RNG_SEED environment variable to produce a different stream of variates,

$ GSL_RNG_SEED=123 ./a.out
GSL_RNG_SEED=123
 1 1 2 1 2 6 2 1 8 7

The following program generates a random walk in two dimensions.

#include <stdio.h>
#include <gsl/gsl_rng.h>
#include <gsl/gsl_randist.h>
int
main (void)
{
  const gsl_rng_type * T;
  gsl_rng * r;
  gsl_rng_env_setup();
  T = gsl_rng_default;
  r = gsl_rng_alloc (T);
  int i;
  double x = 0, y = 0, dx, dy;
  printf("%g %g\n", x, y);
  for (i = 0; i < 10; i++)
    {
      gsl_ran_dir_2d (r, &dx, &dy);
      x += dx; y += dy;
      printf("%g %g\n", x, y);
    }
  return 0;
}

Example output from the program, three 10-step random walks from the origin.

Random walk k p(k)  7 6 5 4 3 2 1 0 -1 -2 -3 -4 -5 -6 -7 7 6 5 4 3 2 1 0 -1 -2 -3 -4 -5 -6 -7

18.37 References and Further Reading

For an encyclopaedic coverage of the subject readers are advised to consult the book Non-Uniform Random Variate Generation by Luc Devroye. It covers every imaginable distribution and provides hundreds of algorithms.

Luc Devroye, Non-Uniform Random Variate Generation, Springer-Verlag, ISBN 0-387-96305-7.

The subject of random variate generation is also reviewed by Knuth, who describes algorithms for all the major distributions.

Donald E. Knuth, The Art of Computer Programming: Seminumerical Algorithms (Vol 2, 3rd Ed, 1997), Addison-Wesley, ISBN 0201896842.

The Particle Data Group provides a short review of techniques for generating distributions of random numbers in the “Monte Carlo” section of its Annual Review of Particle Physics.

Review of Particle Properties R.M. Barnett et al., Physical Review D54, 1 (1996)

    http://pdg.lbl.gov/.

The Review of Particle Physics is available online in postscript and pdf format.

19 Statistics

This chapter describes the statistical functions in the library. The basic statistical functions include routines to compute the mean, variance and standard deviation. More advanced functions allow you to calculate absolute deviations, skewness, and kurtosis as well as the median and arbitrary percentiles. The algorithms use recurrence relations to compute average quantities in a stable way, without large intermediate values that might overflow.

The functions are available in versions for datasets in the standard floating-point and integer types. The versions for double precision floating-point data have the prefix gsl_ stats and are declared in the header file ‘gsl_stats_double.h’. The versions for integer data have the prefix gsl_stats_int and are declared in the header files ‘gsl_stats_int.h’.

19.1 Mean, Standard Deviation and Variance

Statistics double gsl stats mean (const double data[], size_t stride, size_t

n) This function returns the arithmetic mean of data, a dataset of length n with stride stride. The arithmetic mean, or sample mean, is denoted by μ̂ and defined as,

μ̂ = \frac{1}{N} ∑ x_{i}

where x_{i} are the elements of the dataset data. For samples drawn from a gaussian distribution the variance of μ̂ is σ^{2}/N.

Statistics double gsl stats variance (const double data[], size_t stride,

    size_t n)

This function returns the estimated, or sample, variance of data, a dataset of length n with stride stride. The estimated variance is denoted by σ̂^{2} and is defined by,

σ̂^{2} = \frac{1}{(N − 1)} ∑(x_{i} − μ̂)^{2}

where x_{i} are the elements of the dataset data. Note that the normalization factor of 1/(N −1) results from the derivation of σ̂^{2} as an unbiased estimator of the population variance σ^{2}. For samples drawn from a gaussian distribution the variance of σ̂^{2} itself is 2σ^{4}/N. This function computes the mean via a call to gsl_stats_mean. If you have already computed the mean then you can pass it directly to gsl_stats_variance_m.

Statistics double gsl stats variance m (const double data[], size_t stride,

    size_t n, double mean)

This function returns the sample variance of data relative to the given value of mean. The function is computed with μ̂ replaced by the value of mean that you supply,

σ̂^{2} = \frac{1}{(N − 1)} ∑(x_{i} − mean)^{2}

Statistics double gsl stats sd (const double data[], size_t stride, size_t n) Statistics double gsl stats sd m (const double data[], size_t stride, size_t

    n, double mean)

The standard deviation is defined as the square root of the variance. These functions return the square root of the corresponding variance functions above.

Statistics double gsl stats variance with fixed mean (const double

    data[], size_t stride, size_t n, double mean)

This function computes an unbiased estimate of the variance of data when the population mean mean of the underlying distribution is known a priori. In this case the estimator for the variance uses the factor 1/N and the sample mean μ̂ is replaced by the known population mean μ,

σ̂^{2} = \frac{1}{N} ∑(x_{i} − μ)^{2}

Statistics double gsl stats sd with fixed mean (const double data[],

    size_t stride, size_t n, double mean)

This function calculates the standard deviation of data for a a fixed population mean mean. The result is the square root of the corresponding variance function.

19.2 Absolute deviation

Statistics double gsl stats absdev (const double data[], size_t stride,

    size_t n)

This function computes the absolute deviation from the mean of data, a dataset of length n with stride stride. The absolute deviation from the mean is defined as,

absdev = \frac{1}{N} ∑ |x_{i} − μ̂|

where x_{i} are the elements of the dataset data. The absolute deviation from the mean provides a more robust measure of the width of a distribution than the variance. This function computes the mean of data via a call to gsl_stats_mean.

Statistics double gsl stats absdev m (const double data[], size_t stride,

    size_t n, double mean)

This function computes the absolute deviation of the dataset data relative to the given value of mean,

absdev = \frac{1}{N} ∑ |x_{i} − mean|

This function is useful if you have already computed the mean of data (and want to avoid recomputing it), or wish to calculate the absolute deviation relative to another value (such as zero, or the median).

19.3 Higher moments (skewness and kurtosis)

Statistics double gsl stats skew (const double data[], size_t stride, size_t

n) This function computes the skewness of data, a dataset of length n with stride stride. The skewness is defined as,

skew = \frac{1}{N} ∑ (\frac{x_{i} − μ̂}{σ̂})^{3}

where x_{i} are the elements of the dataset data. The skewness measures the asymmetry of the tails of a distribution. The function computes the mean and estimated standard deviation of data via calls

to gsl_stats_mean and gsl_stats_sd.

Statistics double gsl stats skew m sd (const double data[], size_t stride,

    size_t n, double mean, double sd)

This function computes the skewness of the dataset data using the given values of the mean mean and standard deviation sd,

skew = \frac{1}{N} ∑ (\frac{x_{i} − mean}{sd})^{3}

These functions are useful if you have already computed the mean and standard deviation of data and want to avoid recomputing them.

Statistics double gsl stats kurtosis (const double data[], size_t stride,

    size_t n)

This function computes the kurtosis of data, a dataset of length n with stride stride. The kurtosis is defined as,

kurtosis = (\frac{1}{N} ∑ (\frac{x_{i} − μ̂}{σ̂})^{4}) − 3

The kurtosis measures how sharply peaked a distribution is, relative to its width. The kurtosis is normalized to zero for a gaussian distribution.

Statistics double gsl stats kurtosis m sd (const double data[], size_t

    stride, size_t n, double mean, double sd)

This function computes the kurtosis of the dataset data using the given values of the mean mean and standard deviation sd,

kurtosis = \frac{1}{N} (∑ (\frac{x_{i} − mean}{sd})^{4}) − 3

This function is useful if you have already computed the mean and standard deviation of data and want to avoid recomputing them.

19.4 Autocorrelation

Function double gsl stats lag1 autocorrelation (const double data[],

    const size_t stride, const size_t n)

This function computes the lag-1 autocorrelation of the dataset data.

a_{1} = \frac{∑_{i=1}^{n}(x_{i} − μ̂)(x_{i−1} − μ̂)}{∑_{i=1}^{n}(x_{i} − μ̂)(x_{i} − μ̂)}

Function double gsl stats lag1 autocorrelation m (const double data[],

    const size_t stride, const size_t n, const double mean)

This function computes the lag-1 autocorrelation of the dataset data using the given value of the mean mean.

19.5 Covariance

Function double gsl stats covariance (const double data1[], const size_t

    stride1, const double data2[], const size_t stride2, const size_t n)

This function computes the covariance of the datasets data1 and data2 which must both be of the same length n.

covar = \frac{1}{(n − 1)} ∑_{i=1}^{n}(x_{i} − x̂)(y_{i} − ŷ)

Function double gsl stats covariance m (const double data1[], const

    size_t stride1, const double data2[], const size_t n, const double
    mean1, const double mean2)

This function computes the covariance of the datasets data1 and data2 using the given values of the means, mean1 and mean2.

19.6 Weighted Samples

The functions described in this section allow the computation of statistics for weighted samples. The functions accept an array of samples, x_{i}, with associated weights, w_{i}. Each sample x_{i} is considered as having been drawn from a Gaussian distribution with variance σ_{i}^{2}. The sample weight w_{i} is defined as the reciprocal of this variance, w_{i} = 1/σ_{i}^{2}. Setting a weight to zero corresponds to removing a sample from a dataset.

Statistics double gsl stats wmean (const double w[], size_t wstride, const

    double data[], size_t stride, size_t n)

This function returns the weighted mean of the dataset data with stride stride and length n, using the set of weights w with stride wstride and length n. The weighted mean is defined as,

μ̂ = \frac{∑ w_{i}x_{i}}{∑ w_{i}}

Statistics double gsl stats wvariance (const double w[], size_t wstride,

    const double data[], size_t stride, size_t n)

This function returns the estimated variance of the dataset data with stride stride and length n, using the set of weights w with stride wstride and length n. The estimated variance of a weighted dataset is defined as,

σ̂^{2} = \frac{∑ w_{i}}{(∑ w_{i})^{2} − ∑(w_{i}^{2})} ∑ w_{i}(x_{i} − μ̂)^{2}

Note that this expression reduces to an unweighted variance with the familiar 1/(N − 1) factor when there are N equal non-zero weights.

Statistics double gsl stats wvariance m (const double w[], size_t wstride,

    const double data[], size_t stride, size_t n, double wmean)

This function returns the estimated variance of the weighted dataset data using the given weighted mean wmean.

Statistics double gsl stats wsd (const double w[], size_t wstride, const

    double data[], size_t stride, size_t n)

The standard deviation is defined as the square root of the variance. This function returns the square root of the corresponding variance function gsl_stats_wvariance above.

Statistics double gsl stats wsd m (const double w[], size_t wstride, const

    double data[], size_t stride, size_t n, double wmean)

This function returns the square root of the corresponding variance function gsl_

stats_wvariance_m above.

Statistics double gsl stats wvariance with fixed mean (const double

    w[], size_t wstride, const double data[], size_t stride, size_t n)

This function computes an unbiased estimate of the variance of weighted dataset data when the population mean mean of the underlying distribution is known a priori. In this case the estimator for the variance replaces the sample mean μ̂ by the known population mean μ,

σ̂^{2} = \frac{∑ w_{i}(x_{i} − μ)^{2}}{∑ w_{i}}

Statistics double gsl stats wsd with fixed mean (const double w[],

    size_t wstride, const double data[], size_t stride, size_t n)

The standard deviation is defined as the square root of the variance. This function returns the square root of the corresponding variance function above.

Statistics double gsl stats wabsdev (const double w[], size_t wstride,

    const double data[], size_t stride, size_t n)

This function computes the weighted absolute deviation from the weighted mean of data. The absolute deviation from the mean is defined as,

absdev = \frac{∑ w_{i}|x_{i} − μ̂|}{∑ w_{i}}

Statistics double gsl stats wabsdev m (const double w[], size_t wstride,

    const double data[], size_t stride, size_t n, double wmean)

This function computes the absolute deviation of the weighted dataset data about the given weighted mean wmean.

Statistics double gsl stats wskew (const double w[], size_t wstride, const

    double data[], size_t stride, size_t n)

This function computes the weighted skewness of the dataset data.

skew = \frac{∑ w_{i}((x_{i} − xbar)/σ)^{3}}{∑ w_{i}}

Statistics double gsl stats wskew m sd (const double w[], size_t wstride,

    const double data[], size_t stride, size_t n, double wmean, double wsd)

This function computes the weighted skewness of the dataset data using the given values of the weighted mean and weighted standard deviation, wmean and wsd.

Statistics double gsl stats wkurtosis (const double w[], size_t wstride,

    const double data[], size_t stride, size_t n)

This function computes the weighted kurtosis of the dataset data.

kurtosis = \frac{∑ w_{i}((x_{i} − xbar)/sigma)^{4}}{∑ w_{i}} − 3

Statistics double gsl stats wkurtosis m sd (const double w[], size_t

    wstride, const double data[], size_t stride, size_t n, double wmean,
    double wsd)

This function computes the weighted kurtosis of the dataset data using the given values of the weighted mean and weighted standard deviation, wmean and wsd.

19.7 Maximum and Minimum values

Statistics double gsl stats max (const double data[], size_t stride, size_t

n) This function returns the maximum value in data, a dataset of length n with stride stride. The maximum value is defined as the value of the element x_{i} which satisfies x_{i} ≥ x_{j} for all j. If you want instead to find the element with the largest absolute magnitude you will need to apply fabs or abs to your data before calling this function.

Statistics double gsl stats min (const double data[], size_t stride, size_t n)

This function returns the minimum value in data, a dataset of length n with stride stride. The minimum value is defined as the value of the element x_{i} which satisfies x_{i} ≤ x_{j} for all j. If you want instead to find the element with the smallest absolute magnitude you will need to apply fabs or abs to your data before calling this function.

Statistics void gsl stats minmax (double * min, double * max, const

    double data[], size_t stride, size_t n)

This function finds both the minimum and maximum values min, max in data in a single pass.

Statistics size_t gsl stats max index (const double data[], size_t stride,

    size_t n)

This function returns the index of the maximum value in data, a dataset of length n with stride stride. The maximum value is defined as the value of the element x_{i} which satisfies x_{i} ≥ x_{j} for all j. When there are several equal maximum elements then the first one is chosen.

Statistics size_t gsl stats min index (const double data[], size_t stride,

    size_t n)

This function returns the index of the minimum value in data, a dataset of length n with stride stride. The minimum value is defined as the value of the element x_{i} which satisfies x_{i} ≥ x_{j} for all j. When there are several equal minimum elements then the first one is chosen.

Statistics void gsl stats minmax index (size_t * min index, size_t *

    max index, const double data[], size_t stride, size_t n)

This function returns the indexes min index, max index of the minimum and maximum values in data in a single pass.

19.8 Median and Percentiles

The median and percentile functions described in this section operate on sorted data. For convenience we use quantiles, measured on a scale of 0 to 1, instead of percentiles (which use a scale of 0 to 100).

Statistics double gsl stats median from sorted data (const double

sorted data[], size_t stride, size_t n) This function returns the median value of sorted data, a dataset of length n with stride stride. The elements of the array must be in ascending numerical order. There are no checks to see whether the data are sorted, so the function gsl_sort should always be used first. When the dataset has an odd number of elements the median is the value of element (n − 1)/2. When the dataset has an even number of elements the median is the mean of the two nearest middle values, elements (n − 1)/2 and n/2. Since the algorithm for computing the median involves interpolation this function always returns a floatingpoint number, even for integer data types.

Statistics double gsl stats quantile from sorted data (const double

    sorted data[], size_t stride, size_t n, double f )

This function returns a quantile value of sorted data, a double-precision array of length n with stride stride. The elements of the array must be in ascending numerical order. The quantile is determined by the f, a fraction between 0 and 1. For example, to compute the value of the 75th percentile f should have the value 0.75.

There are no checks to see whether the data are sorted, so the function gsl_sort should always be used first. The quantile is found by interpolation, using the formula

quantile = (1 − δ)x_{i} + δx_{i+1}

where i is floor((n − 1)f) and δ is (n − 1)f − i. Thus the minimum value of the array (data[0*stride]) is given by f equal to zero, the maximum value (data[(n-1)*stride]) is given by f equal to one and the median value is given by f equal to 0.5. Since the algorithm for computing quantiles involves interpolation this function always returns a floating-point number, even for integer data types.

19.9 Example statistical programs

Here is a basic example of how to use the statistical functions:

#include <stdio.h>
#include <gsl/gsl_statistics.h>
int
main(void)
{
  double data[5] = {17.2, 18.1, 16.5, 18.3, 12.6};
  double mean, variance, largest, smallest;
  mean     = gsl_stats_mean(data, 1, 5);
  variance = gsl_stats_variance(data, 1, 5);
  largest  = gsl_stats_max(data, 1, 5);
  smallest = gsl_stats_min(data, 1, 5);
  printf("The dataset is %g, %g, %g, %g, %g\n",
        data[0], data[1], data[2], data[3], data[4]);
  printf("The sample mean is %g\n", mean);
  printf("The estimated variance is %g\n", variance);
  printf("The largest value is %g\n", largest);
  printf("The smallest value is %g\n", smallest);
  return 0;
}

The program should produce the following output,

The dataset is 17.2, 18.1, 16.5, 18.3, 12.6
The sample mean is 16.54
The estimated variance is 4.2984
The largest value is 18.3
The smallest value is 12.6

Here is an example using sorted data,

#include <stdio.h>
#include <gsl/gsl_sort.h>
#include <gsl/gsl_statistics.h>
int
main(void)
{
  double data[5] = {17.2, 18.1, 16.5, 18.3, 12.6};
  double median, upperq, lowerq;
  printf("Original dataset:  %g, %g, %g, %g, %g\n",
        data[0], data[1], data[2], data[3], data[4]);
  gsl_sort (data, 1, 5);
  printf("Sorted dataset: %g, %g, %g, %g, %g\n",
        data[0], data[1], data[2], data[3], data[4]);
  median
    = gsl_stats_median_from_sorted_data (data,
                                         1, 5);
  upperq
    = gsl_stats_quantile_from_sorted_data (data,
                                           1, 5,
                                           0.75);
  lowerq
    = gsl_stats_quantile_from_sorted_data (data,
                                           1, 5,
                                           0.25);
  printf("The median is %g\n", median);
  printf("The upper quartile is %g\n", upperq);
  printf("The lower quartile is %g\n", lowerq);
  return 0;
}

This program should produce the following output,

Original dataset: 17.2, 18.1, 16.5, 18.3, 12.6
Sorted dataset: 12.6, 16.5, 17.2, 18.1, 18.3
The median is 17.2
The upper quartile is 18.1
The lower quartile is 16.5

19.10 References and Further Reading

The standard reference for almost any topic in statistics is the multi-volume Advanced Theory of Statistics by Kendall and Stuart.

Maurice Kendall, Alan Stuart, and J. Keith Ord. The Advanced Theory of Statistics (multiple volumes) reprinted as Kendall’s Advanced Theory of Statistics. Wiley, ISBN 047023380X.

Many statistical concepts can be more easily understood by a Bayesian approach. The following book by Gelman, Carlin, Stern and Rubin gives a comprehensive coverage of the subject.

Andrew Gelman, John B. Carlin, Hal S. Stern, Donald B. Rubin. Bayesian Data Analysis. Chapman & Hall, ISBN 0412039915.

For physicists the Particle Data Group provides useful reviews of Probability and Statistics in the "Mathematical Tools" section of its Annual Review of Particle Physics.

Review of Particle Properties R.M. Barnett et al., Physical Review D54, 1 (1996)

The Review of Particle Physics is available online at http://pdg.lbl.gov/.

20 Histograms

This chapter describes functions for creating histograms. Histograms provide a convenient way of summarizing the distribution of a set of data. A histogram consists of a set of bins which count the number of events falling into a given range of a continuous variable x. In GSL the bins of a histogram contain floating-point numbers, so they can be used to record both integer and non-integer distributions. The bins can use arbitrary sets of ranges (uniformly spaced bins are the default). Both one and two-dimensional histograms are supported.

Once a histogram has been created it can also be converted into a probability distribution function. The library provides efficient routines for selecting random samples from probability distributions. This can be useful for generating simulations based real data.

The functions are declared in the header files ‘gsl_histogram.h’ and

‘gsl_histogram2d.h’.

20.1 The histogram struct

A histogram is defined by the following struct,

Data Type gsl histogram

size_t n This is the number of histogram bins

     double * range

The ranges of the bins are stored in an array of n+1 elements pointed to by range.

     double * bin

The counts for each bin are stored in an array of n elements pointed to by bin. The bins are floating-point numbers, so you can increment them by non-integer values if necessary.

The range for bin[i] is given by range[i] to range[i+1]. For n bins there are n + 1 entries in the array range. Each bin is inclusive at the lower end and exclusive at the upper end. Mathematically this means that the bins are defined by the following inequality,

bin[i] corresponds to range[i] ≤ x < range[i+1]

Here is a diagram of the correspondence between ranges and bins on the number-line for x,

          [ bin[0] )[ bin[1] )[ bin[2] )[ bin[3] )[ bin[5] )
       ---|---------|---------|---------|---------|---------|---  x
        r[0]      r[1]      r[2]      r[3]      r[4]      r[5]

In this picture the values of the range array are denoted by r. On the left-hand side of each bin the square bracket "[" denotes an inclusive lower bound (r ≤ x), and the round parentheses ")" on the right-hand side denote an exclusive upper bound (x < r). Thus any samples which fall on the upper end of the histogram are excluded. If you want to include this value for the last bin you will need to add an extra bin to your histogram.

The gsl_histogram struct and its associated functions are defined in the header file

‘gsl_histogram.h’.

20.2 Histogram allocation

The functions for allocating memory to a histogram follow the style of malloc and free. In addition they also perform their own error checking. If there is insufficient memory available to allocate a histogram then the functions call the error handler (with an error number of GSL_ENOMEM) in addition to returning a null pointer. Thus if you use the library error handler to abort your program then it isn’t necessary to check every histogram alloc.

Function gsl_histogram * gsl histogram alloc (size_t n)

This function allocates memory for a histogram with n bins, and returns a pointer to a newly created gsl_histogram struct. If insufficient memory is available a null pointer is returned and the error handler is invoked with an error code of GSL_ENOMEM. The bins and ranges are not initialized, and should be prepared using one of the rangesetting functions below in order to make the histogram ready for use.

Function int gsl histogram set ranges (gsl_histogram * h, const double

range[], size_t size) This function sets the ranges of the existing histogram h using the array range of size size. The values of the histogram bins are reset to zero. The range array should contain the desired bin limits. The ranges can be arbitrary, subject to the restriction that they are monotonically increasing. The following example shows how to create a histogram with logarithmic bins with ranges [1,10), [10,100) and [100,1000).

     gsl_histogram * h = gsl_histogram_alloc (3);
     /* bin[0] covers the range 1 <= x < 10 */
     /* bin[1] covers the range 10 <= x < 100 */
     /* bin[2] covers the range 100 <= x < 1000 */
     double range[4] = { 1.0, 10.0, 100.0, 1000.0 };
     gsl_histogram_set_ranges (h, range, 4);

Note that the size of the range array should be defined to be one element bigger than the number of bins. The additional element is required for the upper value of the final bin.

Function int gsl histogram set ranges uniform (gsl_histogram * h,

    double xmin, double xmax)

This function sets the ranges of the existing histogram h to cover the range xmin to xmax uniformly. The values of the histogram bins are reset to zero. The bin ranges are shown in the table below,

bin[0] corresponds to  xmin ≤ x < xmin + d
bin[1] corresponds to xmin + d ≤ x < xmin + 2d
… … …
bin[n-1] corresponds to xmin + (n − 1)d ≤ x < xmax

where d is the bin spacing, d = (xmax − xmin)/n.

Function void gsl histogram free (gsl_histogram * h)

This function frees the histogram h and all of the memory associated with it.

20.3 Copying Histograms

Function int gsl histogram memcpy (gsl_histogram * dest, const

    gsl_histogram * src)

This function copies the histogram src into the pre-existing histogram dest, making dest into an exact copy of src. The two histograms must be of the same size.

Function gsl_histogram * gsl histogram clone (const gsl_histogram *

src) This function returns a pointer to a newly created histogram which is an exact copy of the histogram src.

20.4 Updating and accessing histogram elements

There are two ways to access histogram bins, either by specifying an x coordinate or by using the bin-index directly. The functions for accessing the histogram through x coordinates use a binary search to identify the bin which covers the appropriate range.

Function int gsl histogram increment (gsl_histogram * h, double x)

This function updates the histogram h by adding one (1.0) to the bin whose range contains the coordinate x. If x lies in the valid range of the histogram then the function returns zero to indicate success. If x is less than the lower limit of the histogram then the function returns GSL_EDOM, and none of bins are modified. Similarly, if the value of x is greater than or equal to the upper limit of the histogram then the function returns GSL_EDOM, and none of the bins are modified. The error handler is not called, however, since it is often necessary to compute histogram for a small range of a larger dataset, ignoring the values outside the range of interest.

Function int gsl histogram accumulate (gsl_histogram * h, double x,

double weight) This function is similar to gsl_histogram_increment but increases the value of the appropriate bin in the histogram h by the floating-point number weight.

Function double gsl histogram get (const gsl_histogram * h, size_t i)

This function returns the contents of the ith bin of the histogram h. If i lies outside the valid range of indices for the histogram then the error handler is called with an error code of GSL_EDOM and the function returns 0.

Function int gsl histogram get range (const gsl_histogram * h, size_t i,

    double * lower, double * upper)

This function finds the upper and lower range limits of the ith bin of the histogram h. If the index i is valid then the corresponding range limits are stored in lower and upper. The lower limit is inclusive (i.e. events with this coordinate are included in the bin) and the upper limit is exclusive (i.e. events with the coordinate of the upper limit are excluded and fall in the neighboring higher bin, if it exists). The function returns 0 to indicate success. If i lies outside the valid range of indices for the histogram then the error handler is called and the function returns an error code of GSL_EDOM.

Function double gsl histogram max (const gsl_histogram * h) Function double gsl histogram min (const gsl_histogram * h)

Function size_t gsl histogram bins (const gsl_histogram * h)

These functions return the maximum upper and minimum lower range limits and the number of bins of the histogram h. They provide a way of determining these values without accessing the gsl_histogram struct directly.

Function void gsl histogram reset (gsl_histogram * h)

This function resets all the bins in the histogram h to zero.

20.5 Searching histogram ranges

The following functions are used by the access and update routines to locate the bin which corresponds to a given x coordinate.

Function int gsl histogram find (const gsl_histogram * h, double x,
    size_t * i)

This function finds and sets the index i to the bin number which covers the coordinate x in the histogram h. The bin is located using a binary search. The search includes an optimization for histograms with uniform range, and will return the correct bin immediately in this case. If x is found in the range of the histogram then the function sets the index i and returns GSL_SUCCESS. If x lies outside the valid range of the histogram then the function returns GSL_EDOM and the error handler is invoked.

20.6 Histogram Statistics

Function double gsl histogram max val (const gsl_histogram * h)

This function returns the maximum value contained in the histogram bins.

Function size_t gsl histogram max bin (const gsl_histogram * h)

This function returns the index of the bin containing the maximum value. In the case where several bins contain the same maximum value the smallest index is returned.

Function double gsl histogram min val (const gsl_histogram * h)

This function returns the minimum value contained in the histogram bins.

Function size_t gsl histogram min bin (const gsl_histogram * h)

This function returns the index of the bin containing the minimum value. In the case where several bins contain the same maximum value the smallest index is returned.

Function double gsl histogram mean (const gsl_histogram * h)

This function returns the mean of the histogrammed variable, where the histogram is regarded as a probability distribution. Negative bin values are ignored for the purposes of this calculation. The accuracy of the result is limited by the bin width.

Function double gsl histogram sigma (const gsl_histogram * h)

This function returns the standard deviation of the histogrammed variable, where the histogram is regarded as a probability distribution. Negative bin values are ignored for the purposes of this calculation. The accuracy of the result is limited by the bin width.

20.7 Histogram Operations

Function int gsl histogram equal bins p (const gsl_histogram *h1, const

    gsl_histogram *h2)

This function returns 1 if the all of the individual bin ranges of the two histograms are identical, and 0 otherwise.

Function int gsl histogram add (gsl_histogram *h1, const gsl_histogram

*h2) This function adds the contents of the bins in histogram h2 to the corresponding bins of histogram h1, i.e. h′_{1}(i) = h_{1}(i) + h_{2}(i). The two histograms must have identical bin ranges.

Function int gsl histogram sub (gsl_histogram *h1, const gsl_histogram

*h2) This function subtracts the contents of the bins in histogram h2 from the corresponding bins of histogram h1, i.e. h′_{1}(i) = h_{1}(i) − h_{2}(i). The two histograms must have identical bin ranges.

Function int gsl histogram mul (gsl_histogram *h1, const gsl_histogram

*h2) This function multiplies the contents of the bins of histogram h1 by the contents of the corresponding bins in histogram h2, i.e. h′_{1}(i) = h_{1}(i) ∗ h_{2}(i). The two histograms must have identical bin ranges.

Function int gsl histogram div (gsl_histogram *h1, const gsl_histogram

*h2) This function divides the contents of the bins of histogram h1 by the contents of the corresponding bins in histogram h2, i.e. h′_{1}(i) = h_{1}(i)/h_{2}(i). The two histograms must have identical bin ranges.

Function int gsl histogram scale (gsl_histogram *h, double scale)

This function multiplies the contents of the bins of histogram h by the constant scale, i.e. h′_{1}(i) = h_{1}(i) ∗ scale.

Function int gsl histogram shift (gsl_histogram *h, double offset)

This function shifts the contents of the bins of histogram h by the constant offset, i.e. h′_{1}(i) = h_{1}(i) + offset.

20.8 Reading and writing histograms

The library provides functions for reading and writing histograms to a file as binary data or formatted text.

Function int gsl histogram fwrite (FILE * stream, const gsl_histogram *

h) This function writes the ranges and bins of the histogram h to the stream stream in binary format. The return value is 0 for success and GSL_EFAILED if there was a problem writing to the file. Since the data is written in the native binary format it may not be portable between different architectures.

Function int gsl histogram fread (FILE * stream, gsl_histogram * h)

This function reads into the histogram h from the open stream stream in binary format. The histogram h must be preallocated with the correct size since the function uses the number of bins in h to determine how many bytes to read. The return value is 0 for success and GSL_EFAILED if there was a problem reading from the file. The data is assumed to have been written in the native binary format on the same architecture.

Function int gsl histogram fprintf (FILE * stream, const gsl_histogram *

    h, const char * range format, const char * bin format)

This function writes the ranges and bins of the histogram h line-by-line to the stream stream using the format specifiers range format and bin format. These should be one of the %g, %e or %f formats for floating point numbers. The function returns 0 for success and GSL_EFAILED if there was a problem writing to the file. The histogram output is formatted in three columns, and the columns are separated by spaces, like this,

     range[0] range[1] bin[0]
     range[1] range[2] bin[1]
     range[2] range[3] bin[2]
     ....
     range[n-1] range[n] bin[n-1]

The values of the ranges are formatted using range format and the value of the bins are formatted using bin format. Each line contains the lower and upper limit of the range of the bins and the value of the bin itself. Since the upper limit of one bin is the lower limit of the next there is duplication of these values between lines but this allows the histogram to be manipulated with line-oriented tools.

Function int gsl histogram fscanf (FILE * stream, gsl_histogram * h)

This function reads formatted data from the stream stream into the histogram h. The data is assumed to be in the three-column format used by gsl_histogram_fprintf. The histogram h must be preallocated with the correct length since the function uses the size of h to determine how many numbers to read. The function returns 0 for success and GSL_EFAILED if there was a problem reading from the file.

20.9 Resampling from histograms

A histogram made by counting events can be regarded as a measurement of a probability distribution. Allowing for statistical error, the height of each bin represents the probability of an event where the value of x falls in the range of that bin. The probability distribution function has the one-dimensional form p(x)dx where,

p(x) = n_{i}/(Nw_{i})

In this equation n_{i} is the number of events in the bin which contains x, w_{i} is the width of the bin and N is the total number of events. The distribution of events within each bin is assumed to be uniform.

20.10 The histogram probability distribution struct

The probability distribution function for a histogram consists of a set of bins which measure the probability of an event falling into a given range of a continuous variable x. A probability distribution function is defined by the following struct, which actually stores the cumulative probability distribution function. This is the natural quantity for generating samples via the inverse transform method, because there is a one-to-one mapping between the cumulative probability distribution and the range [0,1]. It can be shown that by taking a uniform random number in this range and finding its corresponding coordinate in the cumulative probability distribution we obtain samples with the desired probability distribution.

Data Type gsl histogram pdf

size_t n This is the number of bins used to approximate the probability distribution function.

     double * range

The ranges of the bins are stored in an array of n+1 elements pointed to by range.

     double * sum

The cumulative probability for the bins is stored in an array of n elements pointed to by sum.

The following functions allow you to create a gsl_histogram_pdf struct which represents this probability distribution and generate random samples from it.

Function gsl_histogram_pdf * gsl histogram pdf alloc (size_t n)

This function allocates memory for a probability distribution with n bins and returns a pointer to a newly initialized gsl_histogram_pdf struct. If insufficient memory is available a null pointer is returned and the error handler is invoked with an error

     code of GSL_ENOMEM.

Function int gsl histogram pdf init (gsl_histogram_pdf * p, const

         gsl_histogram * h)

This function initializes the probability distribution p with with the contents of the histogram h. If any of the bins of h are negative then the error handler is invoked with an error code of GSL_EDOM because a probability distribution cannot contain negative values.

Function void gsl histogram pdf free (gsl_histogram_pdf * p)

This function frees the probability distribution function p and all of the memory associated with it.

Function double gsl histogram pdf sample (const gsl_histogram_pdf * p,

    double r)

This function uses r, a uniform random number between zero and one, to compute a single random sample from the probability distribution p. The algorithm used to compute the sample s is given by the following formula,

s = range[i] + δ ∗ (range[i + 1] − range[i])

where i is the index which satisfies sum[i] ≤ r < sum[i + 1] and delta is (r − sum[i])/(sum[i + 1] − sum[i]).

20.11 Example programs for histograms

The following program shows how to make a simple histogram of a column of numerical data supplied on stdin. The program takes three arguments, specifying the upper and lower bounds of the histogram and the number of bins. It then reads numbers from stdin, one line at a time, and adds them to the histogram. When there is no more data to read it prints out the accumulated histogram using gsl_histogram_fprintf.

#include <stdio.h>
#include <stdlib.h>
#include <gsl/gsl_histogram.h>
int
main (int argc, char **argv)
{
  double a, b;
  size_t n;
  if (argc != 4)
    {
      printf ("Usage: gsl-histogram xmin xmax n\n"
              "Computes a histogram of the data "
              "on stdin using n bins from xmin "
              "to xmax\n");
      exit (0);
    }
  a = atof (argv[1]);
  b = atof (argv[2]);
  n = atoi (argv[3]);
  {
    int status;
    double x;
    gsl_histogram * h = gsl_histogram_alloc (n);

         gsl_histogram_set_uniform (h, a, b);
         while (fscanf(stdin, "%lg", &x) == 1)
           {
             gsl_histogram_increment(h, x);
           }
         gsl_histogram_fprintf (stdout, h, "%g", "%g");
         gsl_histogram_free (h);
       }
       exit (0);
     }

Here is an example of the program in use. We generate 10000 random samples from a Cauchy distribution with a width of 30 and histogram them over the range -100 to 100, using 200 bins.

     $ gsl-randist 0 10000 cauchy 30
        | gsl-histogram -100 100 200 > histogram.dat

A plot of the resulting histogram shows the familiar shape of the Cauchy distribution and the fluctuations caused by the finite sample size.

     $ awk '{print $1, $3 ; print $2, $3}' histogram.dat
        | graph -T X

20.12 Two dimensional histograms

A two dimensional histogram consists of a set of bins which count the number of events falling in a given area of the (x, y) plane. The simplest way to use a two dimensional histogram is to record two-dimensional position information, n(x, y). Another possibility is to form a joint distribution by recording related variables. For example a detector might record both the position of an event (x) and the amount of energy it deposited E. These could be histogrammed as the joint distribution n(x, E).

20.13 The 2D histogram struct

Two dimensional histograms are defined by the following struct,

Data Type gsl histogram2d

     size_t nx, ny

This is the number of histogram bins in the x and y directions.

     double * xrange

The ranges of the bins in the x-direction are stored in an array of nx + 1 elements pointed to by xrange.

     double * yrange

The ranges of the bins in the y-direction are stored in an array of ny + 1 pointed to by yrange.

     double * bin

The counts for each bin are stored in an array pointed to by bin. The bins are floating-point numbers, so you can increment them by non-integer values if necessary. The array bin stores the two dimensional array of bins in a single block of memory according to the mapping bin(i,j) =

               bin[i * ny + j].

The range for bin(i,j) is given by xrange[i] to xrange[i+1] in the x-direction and yrange[j] to yrange[j+1] in the y-direction. Each bin is inclusive at the lower end and exclusive at the upper end. Mathematically this means that the bins are defined by the following inequality,

bin(i,j) corresponds to xrange[i] ≤ x < xrange[i + 1]
and  yrange[j] ≤ y < yrange[j + 1]

Note that any samples which fall on the upper sides of the histogram are excluded. If you want to include these values for the side bins you will need to add an extra row or column to your histogram.

The gsl_histogram2d struct and its associated functions are defined in the header file

‘gsl_histogram2d.h’.

20.14 2D Histogram allocation

The functions for allocating memory to a 2D histogram follow the style of malloc and free. In addition they also perform their own error checking. If there is insufficient memory available to allocate a histogram then the functions call the error handler (with an error number of GSL_ENOMEM) in addition to returning a null pointer. Thus if you use the library error handler to abort your program then it isn’t necessary to check every 2D histogram

alloc.

Function gsl_histogram2d * gsl histogram2d alloc (size_t nx, size_t

ny) This function allocates memory for a two-dimensional histogram with nx bins in the x direction and ny bins in the y direction. The function returns a pointer to a newly created gsl_histogram2d struct. If insufficient memory is available a null pointer is returned and the error handler is invoked with an error code of GSL_ENOMEM. The bins and ranges must be initialized with one of the functions below before the histogram is ready for use.

Function int gsl histogram2d set ranges (gsl_histogram2d * h, const

    double xrange[], size_t xsize, const double yrange[], size_t ysize)

This function sets the ranges of the existing histogram h using the arrays xrange and yrange of size xsize and ysize respectively. The values of the histogram bins are reset to zero.

Function int gsl histogram2d set ranges uniform (gsl_histogram2d * h,

    double xmin, double xmax, double ymin, double ymax)

This function sets the ranges of the existing histogram h to cover the ranges xmin to xmax and ymin to ymax uniformly. The values of the histogram bins are reset to zero.

Function void gsl histogram2d free (gsl_histogram2d * h)

This function frees the 2D histogram h and all of the memory associated with it.

20.15 Copying 2D Histograms

Function int gsl histogram2d memcpy (gsl_histogram2d * dest, const

    gsl_histogram2d * src)

This function copies the histogram src into the pre-existing histogram dest, making dest into an exact copy of src. The two histograms must be of the same size.

Function gsl_histogram2d * gsl histogram2d clone (const

    gsl_histogram2d * src)

This function returns a pointer to a newly created histogram which is an exact copy of the histogram src.

20.16 Updating and accessing 2D histogram elements

You can access the bins of a two-dimensional histogram either by specifying a pair of (x, y) coordinates or by using the bin indices (i, j) directly. The functions for accessing the histogram through (x, y) coordinates use binary searches in the x and y directions to identify the bin which covers the appropriate range.

Function int gsl histogram2d increment (gsl_histogram2d * h, double x,

    double y)

This function updates the histogram h by adding one (1.0) to the bin whose x and y ranges contain the coordinates (x,y). If the point (x, y) lies inside the valid ranges of the histogram then the function returns zero to indicate success. If (x, y) lies outside the limits of the histogram then the function returns GSL_EDOM, and none of bins are modified. The error handler is not called, since it is often necessary to compute histogram for a small range of a larger dataset, ignoring any coordinates outside the range of interest.

Function int gsl histogram2d accumulate (gsl_histogram2d * h, double

    x, double y, double weight)

This function is similar to gsl_histogram2d_increment but increases the value of the appropriate bin in the histogram h by the floating-point number weight.

Function double gsl histogram2d get (const gsl_histogram2d * h, size_t
    i, size_t j)

This function returns the contents of the (i,j)th bin of the histogram h. If (i,j) lies outside the valid range of indices for the histogram then the error handler is called with an error code of GSL_EDOM and the function returns 0.

Function int gsl histogram2d get xrange (const gsl_histogram2d * h,

    size_t i, double * xlower, double * xupper)

Function int gsl histogram2d get yrange (const gsl_histogram2d * h,

    size_t j, double * ylower, double * yupper)

These functions find the upper and lower range limits of the ith and jth bins in the x and y directions of the histogram h. The range limits are stored in xlower and xupper or ylower and yupper. The lower limits are inclusive (i.e. events with these coordinates are included in the bin) and the upper limits are exclusive (i.e. events with the value of the upper limit are not included and fall in the neighboring higher bin, if it exists). The functions return 0 to indicate success. If i or j lies outside the valid range of indices for the histogram then the error handler is called with an error

code of GSL_EDOM.

Function double gsl histogram2d xmax (const gsl_histogram2d * h) Function double gsl histogram2d xmin (const gsl_histogram2d * h)

Function size_t gsl histogram2d nx (const gsl_histogram2d * h)

Function double gsl histogram2d ymax (const gsl_histogram2d * h) Function double gsl histogram2d ymin (const gsl_histogram2d * h)

Function size_t gsl histogram2d ny (const gsl_histogram2d * h)

These functions return the maximum upper and minimum lower range limits and the number of bins for the x and y directions of the histogram h. They provide a way of determining these values without accessing the gsl_histogram2d struct directly.

Function void gsl histogram2d reset (gsl_histogram2d * h)

This function resets all the bins of the histogram h to zero.

20.17 Searching 2D histogram ranges

The following functions are used by the access and update routines to locate the bin which corresponds to a given (x y) coordinate.

Function int gsl histogram2d find (const gsl_histogram2d * h, double x,
    double y, size_t * i, size_t * j)

This function finds and sets the indices i and j to the to the bin which covers the coordinates (x,y). The bin is located using a binary search. The search includes an optimization for histogram with uniform ranges, and will return the correct bin immediately in this case. If (x, y) is found then the function sets the indices (i,j) and returns GSL_SUCCESS. If (x, y) lies outside the valid range of the histogram then the function returns GSL_EDOM and the error handler is invoked.

20.18 2D Histogram Statistics

Function double gsl histogram2d max val (const gsl_histogram2d * h)

This function returns the maximum value contained in the histogram bins.

Function void gsl histogram2d max bin (const gsl_histogram2d * h,

    size_t * i, size_t * j)

This function returns the indices (i,j) of the bin containing the maximum value in the histogram h. In the case where several bins contain the same maximum value the first bin found is returned.

Function double gsl histogram2d min val (const gsl_histogram2d * h)

This function returns the minimum value contained in the histogram bins.

Function void gsl histogram2d min bin (const gsl_histogram2d * h,

    size_t * i, size_t * j)

This function returns the indices (i,j) of the bin containing the minimum value in the histogram h. In the case where several bins contain the same maximum value the first bin found is returned.

20.19 2D Histogram Operations

Function int gsl histogram2d equal bins p (const gsl_histogram2d *h1,

    const gsl_histogram2d *h2)

This function returns 1 if the all of the individual bin ranges of the two histograms are identical, and 0 otherwise.

Function int gsl histogram2d add (gsl_histogram2d *h1, const

    gsl_histogram2d *h2)

This function adds the contents of the bins in histogram h2 to the corresponding bins of histogram h1, i.e. h′_{1}(i, j) = h_{1}(i, j) + h_{2}(i, j). The two histograms must have identical bin ranges.

Function int gsl histogram2d sub (gsl_histogram2d *h1, const

    gsl_histogram2d *h2)

This function subtracts the contents of the bins in histogram h2 from the corresponding bins of histogram h1, i.e. h′_{1}(i, j) = h_{1}(i, j) − h_{2}(i, j). The two histograms must have identical bin ranges.

Function int gsl histogram2d mul (gsl_histogram2d *h1, const

    gsl_histogram2d *h2)

This function multiplies the contents of the bins of histogram h1 by the contents of the corresponding bins in histogram h2, i.e. h′_{1}(i, j) = h_{1}(i, j) ∗ h_{2}(i, j). The two histograms must have identical bin ranges.

Function int gsl histogram2d div (gsl_histogram2d *h1, const

    gsl_histogram2d *h2)

This function divides the contents of the bins of histogram h1 by the contents of the corresponding bins in histogram h2, i.e. h′_{1}(i, j) = h_{1}(i, j)/h_{2}(i, j). The two histograms must have identical bin ranges.

Function int gsl histogram2d scale (gsl_histogram2d *h, double scale)

This function multiplies the contents of the bins of histogram h by the constant scale, i.e. h′_{1}(i, j) = h_{1}(i, j) ∗ scale.

Function int gsl histogram2d shift (gsl_histogram2d *h, double offset)

This function shifts the contents of the bins of histogram h by the constant offset, i.e. h′_{1}(i, j) = h_{1}(i, j) + offset.

20.20 Reading and writing 2D histograms

The library provides functions for reading and writing two dimensional histograms to a file as binary data or formatted text.

Function int gsl histogram2d fwrite (FILE * stream, const

    gsl_histogram2d * h)

This function writes the ranges and bins of the histogram h to the stream stream in binary format. The return value is 0 for success and GSL_EFAILED if there was a problem writing to the file. Since the data is written in the native binary format it may not be portable between different architectures.

Function int gsl histogram2d fread (FILE * stream, gsl_histogram2d * h)

This function reads into the histogram h from the stream stream in binary format. The histogram h must be preallocated with the correct size since the function uses the number of x and y bins in h to determine how many bytes to read. The return value is 0 for success and GSL_EFAILED if there was a problem reading from the file. The data is assumed to have been written in the native binary format on the same architecture.

Function int gsl histogram2d fprintf (FILE * stream, const

    gsl_histogram2d * h, const char * range format, const char * bin format)

This function writes the ranges and bins of the histogram h line-by-line to the stream stream using the format specifiers range format and bin format. These should be one of the %g, %e or %f formats for floating point numbers. The function returns 0 for success and GSL_EFAILED if there was a problem writing to the file. The histogram output is formatted in five columns, and the columns are separated by spaces, like this,

     xrange[0] xrange[1] yrange[0] yrange[1] bin(0,0)
     xrange[0] xrange[1] yrange[1] yrange[2] bin(0,1)
     xrange[0] xrange[1] yrange[2] yrange[3] bin(0,2)
     ....
     xrange[0] xrange[1] yrange[ny-1] yrange[ny] bin(0,ny-1)
     xrange[1] xrange[2] yrange[0] yrange[1] bin(1,0)
     xrange[1] xrange[2] yrange[1] yrange[2] bin(1,1)
     xrange[1] xrange[2] yrange[1] yrange[2] bin(1,2)
     ....
     xrange[1] xrange[2] yrange[ny-1] yrange[ny] bin(1,ny-1)
     ....
     xrange[nx-1] xrange[nx] yrange[0] yrange[1] bin(nx-1,0)
     xrange[nx-1] xrange[nx] yrange[1] yrange[2] bin(nx-1,1)
     xrange[nx-1] xrange[nx] yrange[1] yrange[2] bin(nx-1,2)
     ....
     xrange[nx-1] xrange[nx] yrange[ny-1] yrange[ny] bin(nx-1,ny-1)

Each line contains the lower and upper limits of the bin and the contents of the bin. Since the upper limits of the each bin are the lower limits of the neighboring bins there is duplication of these values but this allows the histogram to be manipulated with line-oriented tools.

Function int gsl histogram2d fscanf (FILE * stream, gsl_histogram2d * h)

This function reads formatted data from the stream stream into the histogram h. The data is assumed to be in the five-column format used by gsl_histogram_fprintf. The histogram h must be preallocated with the correct lengths since the function uses the sizes of h to determine how many numbers to read. The function returns 0 for success and GSL_EFAILED if there was a problem reading from the file.

20.21 Resampling from 2D histograms

As in the one-dimensional case, a two-dimensional histogram made by counting events can be regarded as a measurement of a probability distribution. Allowing for statistical error, the height of each bin represents the probability of an event where (x,y) falls in the range of that bin. For a two-dimensional histogram the probability distribution takes the form p(x, y)dxdy where,

p(x, y) = n_{ij}/(NA_{ij})

In this equation n_{ij} is the number of events in the bin which contains (x, y), A_{ij} is the area of the bin and N is the total number of events. The distribution of events within each bin is assumed to be uniform.

Data Type gsl histogram2d pdf

size_t nx, ny

This is the number of histogram bins used to approximate the probability distribution function in the x and y directions.

double * xrange

The ranges of the bins in the x-direction are stored in an array of nx + 1 elements pointed to by xrange.

double * yrange

The ranges of the bins in the y-direction are stored in an array of ny + 1 pointed to by yrange.

double * sum

The cumulative probability for the bins is stored in an array of nx*ny elements pointed to by sum.

The following functions allow you to create a gsl_histogram2d_pdf struct which represents a two dimensional probability distribution and generate random samples from it.

Function gsl_histogram2d_pdf * gsl histogram2d pdf alloc (size_t nx,

    size_t ny)

This function allocates memory for a two-dimensional probability distribution of size nx-by-ny and returns a pointer to a newly initialized gsl_histogram2d_pdf struct. If insufficient memory is available a null pointer is returned and the error handler is invoked with an error code of GSL_ENOMEM.

Function int gsl histogram2d pdf init (gsl_histogram2d_pdf * p, const

    gsl_histogram2d * h)

This function initializes the two-dimensional probability distribution calculated p from the histogram h. If any of the bins of h are negative then the error handler is invoked with an error code of GSL_EDOM because a probability distribution cannot contain negative values.

Function void gsl histogram2d pdf free (gsl_histogram2d_pdf * p)

This function frees the two-dimensional probability distribution function p and all of the memory associated with it.

Function int gsl histogram2d pdf sample (const gsl_histogram2d_pdf *

    p, double r1, double r2, double * x, double * y)

This function uses two uniform random numbers between zero and one, r1 and r2, to compute a single random sample from the two-dimensional probability distribution p.

20.22 Example programs for 2D histograms

This program demonstrates two features of two-dimensional histograms. First a 10 by 10 2d-histogram is created with x and y running from 0 to 1. Then a few sample points are added to the histogram, at (0.3,0.3) with a height of 1, at (0.8,0.1) with a height of 5 and at (0.7,0.9) with a height of 0.5. This histogram with three events is used to generate a random sample of 1000 simulated events, which are printed out.

#include <stdio.h>
#include <gsl/gsl_rng.h>
#include <gsl/gsl_histogram2d.h>
int
main (void)
{
  const gsl_rng_type * T;
  gsl_rng * r;
  gsl_histogram2d * h = gsl_histogram2d_alloc (10, 10)
  gsl_histogram2d_set_ranges_uniform (h,
                                      0.0, 1.0,
                                      0.0, 1.0);
  gsl_histogram2d_accumulate (h, 0.3, 0.3, 1);
  gsl_histogram2d_accumulate (h, 0.8, 0.1, 5);
  gsl_histogram2d_accumulate (h, 0.7, 0.9, 0.5);
  gsl_rng_env_setup();

{
int i;
gsl_histogram2d_pdf * p
= gsl_histogram2d_pdf_alloc (h->n);
gsl_histogram2d_pdf_init (p, h);
for (i = 0; i < 1000; i++) {
double x, y;
double u = gsl_rng_uniform (r);
double v = gsl_rng_uniform (r);
int status
 = gsl_histogram2d_pdf_sample (p, u, v, &x, &y);
printf("%g %g\n", x, y);
}
}
return 0;
}

The following plot shows the distribution of the simulated events. Using a higher resolution grid we can see the original underlying histogram and also the statistical fluctuations caused by the events being uniformly distributed over the the area of the original bins.

21 N-tuples

This chapter describes functions for creating and manipulating ntuples, sets of values associated with events. The ntuples are stored in files. Their values can be extracted in any combination and booked in an histogram using a selection function.

The values to be stored are held in a user-defined data structure, and an ntuple is created associating this data structure with a file. The values are then written to the file (normally inside a loop) using the ntuple functions described below.

A histogram can be created from ntuple data by providing a selection function and a value function. The selection function specifies whether an event should be included in the subset to be analyzed or not. The value function computes the entry to be added to the histogram entry for each event.

All the ntuple functions are defined in the header file ‘gsl_ntuple.h’

21.1 The ntuple struct

Ntuples are manipulated using the gsl_ntuple struct. This struct contains information on the file where the ntuple data is stored, a pointer to the current ntuple data row and the size of the user-defined ntuple data struct.

typedef struct {
    FILE * file;
    void * ntuple_data;
    size_t size;
} gsl_ntuple;

21.2 Creating ntuples

Function gsl_ntuple * gsl ntuple create (char * filename, void *

ntuple data, size_t size) This function creates a new write-only ntuple file filename for ntuples of size size and returns a pointer to the newly created ntuple struct. Any existing file with the same name is truncated to zero length and overwritten. A pointer to memory for the current ntuple row ntuple data must be supplied – this is used to copy ntuples in and out of the file.

21.3 Opening an existing ntuple file

Function gsl_ntuple * gsl ntuple open (char * filename, void *

ntuple data, size_t size) This function opens an existing ntuple file filename for reading and returns a pointer to a corresponding ntuple struct. The ntuples in the file must have size size. A pointer to memory for the current ntuple row ntuple data must be supplied – this is used to copy ntuples in and out of the file.

21.4 Writing ntuples

Function int gsl ntuple write (gsl_ntuple * ntuple)

This function writes the current ntuple ntuple->ntuple data of size ntuple->size to the corresponding file.

Function int gsl ntuple bookdata (gsl_ntuple * ntuple)

This function is a synonym for gsl_ntuple_write

21.5 Reading ntuples

Function int gsl ntuple read (gsl_ntuple * ntuple)

This function reads the current row of the ntuple file for ntuple and stores the values in ntuple->data

21.6 Closing an ntuple file

Function int gsl ntuple close (gsl_ntuple * ntuple)

This function closes the ntuple file ntuple and frees its associated allocated memory.

21.7 Histogramming ntuple values

Once an ntuple has been created its contents can be histogrammed in various ways using the function gsl_ntuple_project. Two user-defined functions must be provided, a function to select events and a function to compute scalar values. The selection function and the value function both accept the ntuple row as a first argument and other parameters as a second argument.

The selection function determines which ntuple rows are selected for histogramming. It is defined by the following struct,

     typedef struct {
       int (* function) (void * ntuple_data, void * params);
       void * params;
     } gsl_ntuple_select_fn;

The struct component function should return a non-zero value for each ntuple row that is to be included in the histogram.

The value function computes scalar values for those ntuple rows selected by the selection function,

     typedef struct {
       double (* function) (void * ntuple_data, void * params);
       void * params;
     } gsl_ntuple_value_fn;

In this case the struct component function should return the value to be added to the histogram for the ntuple row.

Function int gsl ntuple project (gsl_histogram * h, gsl_ntuple * ntuple,

    gsl_ntuple_value_fn *value func, gsl_ntuple_select_fn *select func)

This function updates the histogram h from the ntuple ntuple using the functions value func and select func. For each ntuple row where the selection function select func is non-zero the corresponding value of that row is computed using the function value func and added to the histogram. Those ntuple rows where select func returns zero are ignored. New entries are added to the histogram, so subsequent calls can be used to accumulate further data in the same histogram.

21.8 Example programs

The following example programs demonstrate the use of ntuples in managing a large dataset. The first program creates a set of 100,000 simulated "events", each with 3 associated values (x, y, z). These are generated from a gaussian distribution with unit variance, for demonstration purposes, and written to the ntuple file ‘test.dat’.

#include <config.h>
#include <gsl/gsl_ntuple.h>
#include <gsl/gsl_rng.h>
#include <gsl/gsl_randist.h>
struct data
{
  double x;
  double y;
  double z;
};
int
main (void)
{
  const gsl_rng_type * T;
  gsl_rng * r;
  struct data ntuple_row;
  int i;
  gsl_ntuple *ntuple
    = gsl_ntuple_create ("test.dat", &ntuple_row,
                         sizeof (ntuple_row));
  gsl_rng_env_setup();
  T = gsl_rng_default;
  r = gsl_rng_alloc (T);
  for (i = 0; i < 10000; i++)
    {
      ntuple_row.x = gsl_ran_ugaussian (r);
      ntuple_row.y = gsl_ran_ugaussian (r);
      ntuple_row.z = gsl_ran_ugaussian (r);

    gsl_ntuple_write (ntuple);
  }
gsl_ntuple_close(ntuple);
return 0;
}

The next program analyses the ntuple data in the file ‘test.dat’. The analysis procedure is to compute the squared-magnitude of each event, E^{2} = x^{2} +y^{2} +z^{2}, and select only those which exceed a lower limit of 1.5. The selected events are then histogrammed using their E^{2} values.

#include <config.h>
#include <math.h>
#include <gsl/gsl_ntuple.h>
#include <gsl/gsl_histogram.h>
struct data
{
double x;
double y;
double z;
};
int sel_func (void *ntuple_data, void *params);
double val_func (void *ntuple_data, void *params);
int
main (void)
{
struct data ntuple_row;
int i;
gsl_ntuple *ntuple
  = gsl_ntuple_open ("test.dat", &ntuple_row,
                     sizeof (ntuple_row));
double lower = 1.5;
gsl_ntuple_select_fn S;
gsl_ntuple_value_fn V;
gsl_histogram *h = gsl_histogram_alloc (100);
gsl_histogram_set_ranges_uniform(h, 0.0, 10.0);
S.function = &sel_func;
S.params = &lower;
V.function = &val_func;
V.params = 0;
gsl_ntuple_project (h, ntuple, &V, &S);

gsl_histogram_fprintf (stdout, h, "%f", "%f");
gsl_histogram_free (h);
gsl_ntuple_close (ntuple);
return 0;
}
int
sel_func (void *ntuple_data, void *params)
{
double x, y, z, E, scale;
scale = *(double *) params;
x = ((struct data *) ntuple_data)->x;
y = ((struct data *) ntuple_data)->y;
z = ((struct data *) ntuple_data)->z;
E2 = x * x + y * y + z * z;
return E2 > scale;
}
double
val_func (void *ntuple_data, void *params)
{
double x, y, z;
x = ((struct data *) ntuple_data)->x;
y = ((struct data *) ntuple_data)->y;
z = ((struct data *) ntuple_data)->z;
return x * x + y * y + z * z;
}

The following plot shows the distribution of the selected events. Note the cut-off at the lower bound.

Chapter 21: N-tuples 250

21.9 References and Further Reading

Further information on the use of ntuples can be found in the documentation for the cern packages paw and hbook (available online).

22 Monte Carlo Integration

This chapter describes routines for multidimensional Monte Carlo integration. These include the traditional Monte Carlo method and adaptive algorithms such as vegas and miser which use importance sampling and stratified sampling techniques. Each algorithm computes an estimate of a multidimensional definite integral of the form,

I = ∫_{x_{l}}^{x_{u}} dx ∫_{y_{l}}^{y_{u}} dy ...f(x, y, ...)

over a hypercubic region ((x_{l}, x_{u}), (y_{l}, y_{u}), ...) using a fixed number of function calls. The routines also provide a statistical estimate of the error on the result. This error estimate should be taken as a guide rather than as a strict error bound — random sampling of the region may not uncover all the important features of the function, resulting in an underestimate of the error.

The functions are defined in separate header files for each routine, gsl_monte_plain.h,

‘gsl_monte_miser.h’ and ‘gsl_monte_vegas.h’.

22.1 Interface

All of the Monte Carlo integration routines use the same interface. There is an allocator to allocate memory for control variables and workspace, a routine to initialize those control variables, the integrator itself, and a function to free the space when done.

Each integration function requires a random number generator to be supplied, and returns an estimate of the integral and its standard deviation. The accuracy of the result is determined by the number of function calls specified by the user. If a known level of accuracy is required this can be achieved by calling the integrator several times and averaging the individual results until the desired accuracy is obtained.

Random sample points used within the Monte Carlo routines are always chosen strictly within the integration region, so that endpoint singularities are automatically avoided.

The function to be integrated has its own datatype, defined in the header file

‘gsl_monte.h’.

Data Type gsl monte function

This data type defines a general function with parameters for Monte Carlo integration.

     double (* function) (double * x, size_t dim, void * params)

this function should return the value f(x, params) for argument x and parameters params, where x is an array of size dim giving the coordinates of the point where the function is to be evaluated.

     size_t dim

the number of dimensions for x

void * params a pointer to the parameters of the function Here is an example for a quadratic function in two dimensions,

f(x, y) = ax^{2} + bxy + cy^{2}

with a = 3, b = 2, c = 1. The following code defines a gsl_monte_function F which you could pass to an integrator:

     struct my_f_params { double a; double b; double c; };
     double
     my_f (double x, size_t dim, void * p) {
        struct my_f_params * fp = (struct my_f_params *)p;
        if (dim != 2)
           {
             fprintf(stderr, "error: dim != 2");
             abort();
           }
        return  fp->a * x[0] * x[0]
                  + fp->b * x[0] * x[1]
                    + fp->c * x[1] * x[1];
     }
     gsl_monte_function F;
     struct my_f_params params = { 3.0, 2.0, 1.0 };
     F.function = &my_f;
     F.dim = 2;
     F.params = &params;

The function f(x) can be evaluated using the following macro,

     #define GSL_MONTE_FN_EVAL(F,x)
         (*((F)->function))(x,(F)->dim,(F)->params)

22.2 PLAIN Monte Carlo

The plain Monte Carlo algorithm samples points randomly from the integration region to estimate the integral and its error. Using this algorithm the estimate of the integral E(f; N) for N randomly distributed points x_{i} is given by,

E(f; N) = V ⟨f⟩ = \frac{V}{N} ∑_{i}^{N} f(x_{i}).

where V is the volume of the integration region. The error on this estimate σ(E; N) is calculated from the estimated variance of the mean,

σ^{2}(E; N) = \frac{V}{N} ∑_{i}^{N}(f(x_{i}) − ⟨f⟩)^{2}

For large N this variance decreases asymptotically as var(f)/N, where var(f) is the true variance of the function over the integration region. The error estimate itself should decrease as σ(f)/\sqrt{N}. The familiar law of errors decreasing as 1/\sqrt{N} applies — to reduce the error by a factor of 10 requires a 100-fold increase in the number of sample points.

The functions described in this section are declared in the header file

‘gsl_monte_plain.h’.

Function gsl_monte_plain_state * gsl monte plain alloc (size_t dim)

This function allocates and initializes a workspace for Monte Carlo integration in dim dimensions.

Function int gsl monte plain init (gsl_monte_plain_state* s)

This function initializes a previously allocated integration state. This allows an existing workspace to be reused for different integrations.

Function int gsl monte plain integrate (gsl_monte_function * f, double *

         xl, double * xu, size_t dim, size_t calls, gsl_rng * r,
         gsl_monte_plain_state * s, double * result, double * abserr)

This routines uses the plain Monte Carlo algorithm to integrate the function f over the dim-dimensional hypercubic region defined by the lower and upper limits in the arrays xl and xu, each of size dim. The integration uses a fixed number of function calls calls, and obtains random sampling points using the random number generator r. A previously allocated workspace s must be supplied. The result of the integration is returned in result, with an estimated absolute error abserr.

Function void gsl monte plain free (gsl_monte_plain_state* s), This function frees the memory associated with the integrator state s.

22.3 MISER

The miser algorithm of Press and Farrar is based on recursive stratified sampling. This technique aims to reduce the overall integration error by concentrating integration points in the regions of highest variance.

The idea of stratified sampling begins with the observation that for two disjoint regions a and b with Monte Carlo estimates of the integral E_{a}(f) and E_{b}(f) and variances σ_{a}^{2}(f) and σ_{b}^{2}(f), the variance V ar(f) of the combined estimate E(f) = \frac{1}{2}(E_{a}(f) + E_{b}(f)) is given by,

V ar(f) = \frac{σ_{a}^{2}(f)}{4N_{a}} + \frac{σ_{b}^{2}(f)}{4N_{b}}

It can be shown that this variance is minimized by distributing the points such that,

\frac{N_{a}}{N_{a} + N_{b}} = \frac{σ_{a}}{σ_{a} + σ_{b}}

Hence the smallest error estimate is obtained by allocating sample points in proportion to the standard deviation of the function in each sub-region.

The miser algorithm proceeds by bisecting the integration region along one coordinate axis to give two sub-regions at each step. The direction is chosen by examining all d possible bisections and selecting the one which will minimize the combined variance of the two subregions. The variance in the sub-regions is estimated by sampling with a fraction of the total number of points available to the current step. The same procedure is then repeated recursively for each of the two half-spaces from the best bisection. The remaining sample points are allocated to the sub-regions using the formula for N_{a} and N_{b}. This recursive allocation of integration points continues down to a user-specified depth where each subregion is integrated using a plain Monte Carlo estimate. These individual values and their error estimates are then combined upwards to give an overall result and an estimate of its error.

The functions described in this section are declared in the header file

‘gsl_monte_miser.h’.

Function gsl_monte_miser_state * gsl monte miser alloc (size_t dim)

This function allocates and initializes a workspace for Monte Carlo integration in dim dimensions. The workspace is used to maintain the state of the integration.

Function int gsl monte miser init (gsl_monte_miser_state* s)

This function initializes a previously allocated integration state. This allows an existing workspace to be reused for different integrations.

Function int gsl monte miser integrate (gsl_monte_function * f, double *

    xl, double * xu, size_t dim, size_t calls, gsl_rng * r,
    gsl_monte_miser_state * s, double * result, double * abserr)

This routines uses the miser Monte Carlo algorithm to integrate the function f over the dim-dimensional hypercubic region defined by the lower and upper limits in the arrays xl and xu, each of size dim. The integration uses a fixed number of function calls calls, and obtains random sampling points using the random number generator r. A previously allocated workspace s must be supplied. The result of the integration is returned in result, with an estimated absolute error abserr.

Function void gsl monte miser free (gsl_monte_miser_state* s), This function frees the memory associated with the integrator state s.

The miser algorithm has several configurable parameters. The following variables can be accessed through the gsl_monte_miser_state struct,

Variable double estimate frac

This parameter specifies the fraction of the currently available number of function calls which are allocated to estimating the variance at each recursive step. The default value is 0.1.

Variable size_t min calls

This parameter specifies the minimum number of function calls required for each estimate of the variance. If the number of function calls allocated to the estimate using estimate frac falls below min calls then min calls are used instead. This ensures that each estimate maintains a reasonable level of accuracy. The default value of min calls

is 16 * dim.

Variable size_t min calls per bisection

This parameter specifies the minimum number of function calls required to proceed with a bisection step. When a recursive step has fewer calls available than min calls per bisection it performs a plain Monte Carlo estimate of the current subregion and terminates its branch of the recursion. The default value of this parameter

is 32 * min_calls.

Variable double alpha

This parameter controls how the estimated variances for the two sub-regions of a bisection are combined when allocating points. With recursive sampling the overall variance should scale better than 1/N, since the values from the sub-regions will be obtained using a procedure which explicitly minimizes their variance. To accommodate this behavior the miser algorithm allows the total variance to depend on a scaling parameter α,

V ar(f) = \frac{σ_{a}}{N_{a}^{α}} + \frac{σ_{b}}{N_{b}^{α}}

The authors of the original paper describing miser recommend the value α = 2 as a good choice, obtained from numerical experiments, and this is used as the default value in this implementation.

Variable double dither

This parameter introduces a random fractional variation of size dither into each bisection, which can be used to break the symmetry of integrands which are concentrated near the exact center of the hypercubic integration region. The default value of dither is zero, so no variation is introduced. If needed, a typical value of dither is around 0.1.

22.4 VEGAS

The vegas algorithm of Lepage is based on importance sampling. It samples points from the probability distribution described by the function |f|, so that the points are concentrated in the regions that make the largest contribution to the integral.

In general, if the Monte Carlo integral of f is sampled with points distributed according to a probability distribution described by the function g, we obtain an estimate E_{g}(f; N),

E_{g}(f; N) = E(f/g; N)

with a corresponding variance,

V ar_{g}(f; N) = V ar(f/g; N)

If the probability distribution is chosen as g = |f|/I(|f|) then it can be shown that the variance V_{g}(f; N) vanishes, and the error in the estimate will be zero. In practice it is not possible to sample from the exact distribution g for an arbitrary function, so importance sampling algorithms aim to produce efficient approximations to the desired distribution.

The vegas algorithm approximates the exact distribution by making a number of passes over the integration region while histogramming the function f. Each histogram is used to define a sampling distribution for the next pass. Asymptotically this procedure converges to the desired distribution. In order to avoid the number of histogram bins growing like K^{d} the probability distribution is approximated by a separable function: g(x_{1}, x_{2}, …) = g_{1}(x_{1})g_{2}(x_{2}) … so that the number of bins required is only Kd. This is equivalent to locating the peaks of the function from the projections of the integrand onto the coordinate axes. The efficiency of vegas depends on the validity of this assumption. It is most efficient when the peaks of the integrand are well-localized. If an integrand can be rewritten in a form which is approximately separable this will increase the efficiency of integration with vegas.

vegas incorporates a number of additional features, and combines both stratified sampling and importance sampling. The integration region is divided into a number of “boxes”, with each box getting in fixed number of points (the goal is 2). Each box can then have a fractional number of bins, but if bins/box is less than two, Vegas switches to a kind variance reduction (rather than importance sampling).

Function gsl_monte_vegas_state * gsl monte vegas alloc (size_t dim)

This function allocates and initializes a workspace for Monte Carlo integration in dim dimensions. The workspace is used to maintain the state of the integration.

Function int gsl monte vegas init (gsl_monte_vegas_state* s)

This function initializes a previously allocated integration state. This allows an existing workspace to be reused for different integrations.

Function int gsl monte vegas integrate (gsl_monte_function * f, double *

         xl, double * xu, size_t dim, size_t calls, gsl_rng * r,
         gsl_monte_vegas_state * s, double * result, double * abserr)

This routines uses the vegas Monte Carlo algorithm to integrate the function f over the dim-dimensional hypercubic region defined by the lower and upper limits in the arrays xl and xu, each of size dim. The integration uses a fixed number of function calls calls, and obtains random sampling points using the random number generator r. A previously allocated workspace s must be supplied. The result of the integration is returned in result, with an estimated absolute error abserr. The result and its error estimate are based on a weighted average of independent samples. The chi-squared per degree of freedom for the weighted average is returned via the state struct component, s->chisq, and must be consistent with 1 for the weighted average to be reliable.

Function void gsl monte vegas free (gsl_monte_vegas_state* s), This function frees the memory associated with the integrator state s.

The vegas algorithm computes a number of independent estimates of the integral internally, according to the iterations parameter described below, and returns their weighted average. Random sampling of the integrand can occasionally produce an estimate where the error is zero, particularly if the function is constant in some regions. An estimate with zero error causes the weighted average to break down and must be handled separately. In the original Fortran implementations of vegas the error estimate is made non-zero by substituting a small value (typically 1e-30). The implementation in GSL differs from this and avoids the use of an arbitrary constant – it either assigns the value a weight which is the average weight of the preceding estimates or discards it according to the following procedure,

current estimate has zero error, weighted average has finite error The current estimate is assigned a weight which is the average weight of the preceding estimates.

current estimate has finite error, previous estimates had zero error The previous estimates are discarded and the weighted averaging procedure begins with the current estimate.

current estimate has zero error, previous estimates had zero error The estimates are averaged using the arithmetic mean, but no error is computed.

The vegas algorithm is highly configurable. The following variables can be accessed

through the gsl_monte_vegas_state struct,

Variable double result Variable double sigma

These parameters contain the raw value of the integral result and its error sigma from the last iteration of the algorithm.

Variable double chisq

This parameter gives the chi-squared per degree of freedom for the weighted estimate of the integral. The value of chisq should be close to 1. A value of chisq which differs significantly from 1 indicates that the values from different iterations are inconsistent. In this case the weighted error will be under-estimated, and further iterations of the algorithm are needed to obtain reliable results.

Variable double alpha

The parameter alpha controls the stiffness of the rebinning algorithm. It is typically set between one and two. A value of zero prevents rebinning of the grid. The default value is 1.5.

Variable size_t iterations

The number of iterations to perform for each call to the routine. The default value is 5 iterations.

Variable int stage

Setting this determines the stage of the calculation. Normally, stage = 0 which begins with a new uniform grid and empty weighted average. Calling vegas with stage = 1 retains the grid from the previous run but discards the weighted average, so that one can “tune” the grid using a relatively small number of points and then do a large run with stage = 1 on the optimized grid. Setting stage = 2 keeps the grid and the weighted average from the previous run, but may increase (or decrease) the number of histogram bins in the grid depending on the number of calls available. Choosing stage = 3 enters at the main loop, so that nothing is changed, and is equivalent to performing additional iterations in a previous call.

Variable int mode

The possible choices are GSL_VEGAS_MODE_IMPORTANCE, GSL_VEGAS_MODE_
STRATIFIED, GSL_VEGAS_MODE_IMPORTANCE_ONLY. This determines whether vegas

will use importance sampling or stratified sampling, or whether it can pick on its own. In low dimensions vegas uses strict stratified sampling (more precisely, stratified sampling is chosen if there are fewer than 2 bins per box).

Variable int verbose Variable FILE * ostream

These parameters set the level of information printed by vegas. All information is written to the stream ostream. The default setting of verbose is -1, which turns off all output. A verbose value of 0 prints summary information about the weighted average and final result, while a value of 1 also displays the grid coordinates. A value of 2 prints information from the rebinning procedure for each iteration.

22.5 Examples

The example program below uses the Monte Carlo routines to estimate the value of the following 3-dimensional integral from the theory of random walks,

I = ∫_{−π}^{+π} \frac{dk_{x}}{2π} ∫_{−π}^{+π} \frac{dk_{y}}{2π} ∫_{−π}^{+π} \frac{dk_{z}}{2π} \frac{1}{(1 − cos(k_{x}) cos(k_{y}) cos(k_{z}))}

The analytic value of this integral can be shown to be I = Γ(1/4)^{4}/(4π^{3}) = 1.393203929685676859.... The integral gives the mean time spent at the origin by a random walk on a body-centered cubic lattice in three dimensions.

For simplicity we will compute the integral over the region (0, 0, 0) to (π, π, π) and multiply by 8 to obtain the full result. The integral is slowly varying in the middle of the region but has integrable singularities at the corners (0, 0, 0), (0, π, π), (π, 0, π) and (π, π, 0). The Monte Carlo routines only select points which are strictly within the integration region and so no special measures are needed to avoid these singularities.

#include <stdlib.h>
#include <gsl/gsl_math.h>
#include <gsl/gsl_monte.h>
#include <gsl/gsl_monte_plain.h>
#include <gsl/gsl_monte_miser.h>
#include <gsl/gsl_monte_vegas.h>
/* Computation of the integral,
      I = int (dx dy dz)/(2pi)^3  1/(1-cos(x)cos(y)cos(z))
   over (-pi,-pi,-pi) to (+pi, +pi, +pi).  The exact answer
   is Gamma(1/4)^4/(4 pi^3).  This example is taken from
   C.Itzykson, J.M.Drouffe, "Statistical Field Theory -
   Volume 1", Section 1.1, p21, which cites the original
   paper M.L.Glasser, I.J.Zucker, Proc.Natl.Acad.Sci.USA 74
   1800 (1977) */
/* For simplicity we compute the integral over the region
   (0,0,0) -> (pi,pi,pi) and multiply by 8 */
double exact = 1.3932039296856768591842462603255;
double
g (double *k, size_t dim, void *params)
{
  double A = 1.0 / (M_PI * M_PI * M_PI);
  return A / (1.0 - cos (k[0]) * cos (k[1]) * cos (k[2]));
}
void
display_results (char *title, double result, double error)
{
  printf ("%s ==================\n", title);
  printf ("result = % .6f\n", result);
  printf ("sigma  = % .6f\n", error);

printf ("exact  = % .6f\n", exact);
printf ("error  = % .6f = %.1g sigma\n", result - exact,
        fabs (result - exact) / error);
}
int
main (void)
{
double res, err;
double xl[3] = { 0, 0, 0 };
double xu[3] = { M_PI, M_PI, M_PI };
const gsl_rng_type *T;
gsl_rng *r;
gsl_monte_function G = { &g, 3, 0 };
size_t calls = 500000;
gsl_rng_env_setup ();
T = gsl_rng_default;
r = gsl_rng_alloc (T);
{
  gsl_monte_plain_state *s = gsl_monte_plain_alloc (3);
  gsl_monte_plain_integrate (&G, xl, xu, 3, calls, r, s,
                             &res, &err);
  gsl_monte_plain_free (s);
  display_results ("plain", res, err);
}
{
  gsl_monte_miser_state *s = gsl_monte_miser_alloc (3);
  gsl_monte_miser_integrate (&G, xl, xu, 3, calls, r, s,
                             &res, &err);
  gsl_monte_miser_free (s);
  display_results ("miser", res, err);
}
{
  gsl_monte_vegas_state *s = gsl_monte_vegas_alloc (3);
  gsl_monte_vegas_integrate (&G, xl, xu, 3, 10000, r, s,
                             &res, &err);
  display_results ("vegas warm-up", res, err);
  printf ("converging...\n");

    do
      {
        gsl_monte_vegas_integrate (&G, xl, xu, 3, calls/5, r, s,
                                   &res, &err);
        printf ("result = % .6f sigma = % .6f "
                "chisq/dof = %.1f\n", res, err, s->chisq);
      }
    while (fabs (s->chisq - 1.0) > 0.5);
    display_results ("vegas final", res, err);
    gsl_monte_vegas_free (s);
  }
  return 0;
}

With 500,000 function calls the plain Monte Carlo algorithm achieves a fractional error of 0.6%. The estimated error sigma is consistent with the actual error, and the computed result differs from the true result by about one standard deviation,

plain ==================
result =  1.385867
sigma  =  0.007938
exact  =  1.393204
error  = -0.007337 = 0.9 sigma

The miser algorithm reduces the error by a factor of two, and also correctly estimates the error,

miser ==================
result =  1.390656
sigma  =  0.003743
exact  =  1.393204
error  = -0.002548 = 0.7 sigma

In the case of the vegas algorithm the program uses an initial warm-up run of 10,000 function calls to prepare, or "warm up", the grid. This is followed by a main run with five iterations of 100,000 function calls. The chi-squared per degree of freedom for the five iterations are checked for consistency with 1, and the run is repeated if the results have not converged. In this case the estimates are consistent on the first pass.

vegas warm-up ==================
result =  1.386925
sigma  =  0.002651
exact  =  1.393204
error  = -0.006278 = 2 sigma
converging...
result =  1.392957 sigma =  0.000452 chisq/dof = 1.1
vegas final ==================
result =  1.392957
sigma  =  0.000452
exact  =  1.393204
error  = -0.000247 = 0.5 sigma

If the value of chisq had differed significantly from 1 it would indicate inconsistent results, with a correspondingly underestimated error. The final estimate from vegas (using a similar number of function calls) is significantly more accurate than the other two algorithms.

22.6 References and Further Reading

The miser algorithm is described in the following article, W.H. Press, G.R. Farrar, Recursive Stratified Sampling for Multidimensional Monte Carlo Integration, Computers in Physics, v4 (1990), pp190-195.

The vegas algorithm is described in the following papers, G.P. Lepage, A New Algorithm for Adaptive Multidimensional Integration, Journal of Computational Physics 27, 192-203, (1978) G.P. Lepage, VEGAS: An Adaptive Multi-dimensional Integration Program, Cornell preprint CLNS 80-447, March 1980

23 Simulated Annealing

Stochastic search techniques are used when the structure of a space is not well understood or is not smooth, so that techniques like Newton’s method (which requires calculating Jacobian derivative matrices) cannot be used.

In particular, these techniques are frequently used to solve combinatorial optimization problems, such as the traveling salesman problem.

The basic problem layout is that we are looking for a point in the space at which a real valued energy function (or cost function) is minimized.

Simulated annealing is a technique which has given good results in avoiding local minima; it is based on the idea of taking a random walk through the space at successively lower temperatures, where the probability of taking a step is given by a Boltzmann distribution.

The functions described in this chapter are declared in the header file ‘gsl_siman.h’.

23.1 Simulated Annealing algorithm

We take random walks through the problem space, looking for points with low energies; in these random walks, the probability of taking a step is determined by the Boltzmann distribution

p = e^{−(E_{i+1}−E_{i})/(kT)}

if E_{i+1} > E_{i}, and p = 1 when E_{i+1} ≤ E_{i}.

In other words, a step will occur if the new energy is lower. If the new energy is higher, the transition can still occur, and its likelihood is proportional to the temperature T and inversely proportional to the energy difference E_{i+1} − E_{i}.

The temperature T is initially set to a high value, and a random walk is carried out at that temperature. Then the temperature is lowered very slightly (according to a cooling schedule) and another random walk is taken.

This slight probability of taking a step that gives higher energy is what allows simulated annealing to frequently get out of local minima.

An initial guess is supplied. At each step, a point is chosen at a random distance from the current one, where the random distance r is distributed according to a Boltzmann distribution r = exp^{−E/kT}. After a few search steps using this distribution, the temperature T is lowered according to some scheme, for example T → T/μ_{T} where μ_{T} is slightly greater than 1.

23.2 Simulated Annealing functions

Simulated annealing gsl siman Efunc t

     double (*gsl_siman_Efunc_t) (void *xp);

Simulated annealing gsl siman step t

     void (*gsl_siman_step_t) (void *xp, double step_size);

Simulated annealing gsl siman metric t

     double (*gsl_siman_metric_t) (void *xp, void *yp);

Simulated annealing gsl siman print t

     void (*gsl_siman_print_t) (void *xp);

Simulated annealing gsl siman params t

These are the parameters that control a run of gsl_siman_solve.

     /* this structure contains all the information
        needed to structure the search, beyond the
        energy function, the step function and the
        initial guess. */
     struct s_siman_params {
       /* how many points to try for each step */
       int n_tries;
       /* how many iterations at each temperature? */
       int iters_fixed_T;
       /* max step size in the random walk */
       double step_size;
       /* the following parameters are for the
          Boltzmann distribution */
       double k, t_initial, mu_t, t_min;
     };
     typedef struct s_siman_params gsl_siman_params_t;

Simulated annealing void gsl siman solve (void *x0 p, gsl_siman_Efunc_t

    Ef, gsl_siman_metric_t distance, gsl_siman_print_t print position,
    size_t element size, gsl_siman_params_t params)

Does a simulated annealing search through a given space. The space is specified by providing the functions Ef, distance, print position, element size. The params structure (described above) controls the run by providing the temperature schedule and other tunable parameters to the algorithm (see Section 23.1 [Simulated Annealing algorithm], page 262). p The result (optimal point in the space) is placed in *x0 p. If print position is not null, a log will be printed to the screen with the following columns:

     number_of_iterations temperature x x-(*x0_p) Ef(x)

If print position is null, no information is printed to the screen.

23.3 Examples with Simulated Annealing

GSL’s Simulated Annealing package is clumsy, and it has to be because it is written in C, for C callers, and tries to be polymorphic at the same time. But here we provide some examples which can be pasted into your application with little change and should make things easier.

23.3.1 Trivial example

The first example, in one dimensional cartesian space, sets up an energy function which is a damped sine wave; this has many local minima, but only one global minimum, somewhere between 1.0 and 1.5. The initial guess given is 15.5, which is several local minima away from the global minimum.

/* set up parameters for this simulated annealing run */
/* how many points do we try before stepping */
#define N_TRIES 200
/* how many iterations for each T? */
#define ITERS_FIXED_T 10
/* max step size in random walk */
#define STEP_SIZE 10
/* Boltzmann constant */
#define K 1.0
/* initial temperature */
#define T_INITIAL 0.002
/* damping factor for temperature */
#define MU_T 1.005
#define T_MIN 2.0e-6
gsl_siman_params_t params
  = {N_TRIES, ITERS_FIXED_T, STEP_SIZE,
     K, T_INITIAL, MU_T, T_MIN};
/* now some functions to test in one dimension */
double E1(void *xp)
{
  double x = * ((double *) xp);
  return exp(-square(x-1))*sin(8*x);
}
double M1(void *xp, void *yp)

{
double x = *((double *) xp);
double y = *((double *) yp);
return fabs(x - y);
}
void S1(void *xp, double step_size)
{
double r;
double old_x = *((double *) xp);
double new_x;
r = gsl_ran_uniform();
new_x = r*2*step_size - step_size + old_x;
memcpy(xp, &new_x, sizeof(new_x));
}
void P1(void *xp)
{
printf("%12g", *((double *) xp));
}
int
main(int argc, char *argv[])
{
Element x0; /* initial guess for search */
double x_initial = 15.5;
gsl_siman_solve(&x_initial, E1, S1, M1, P1,
                sizeof(double), params);
return 0;
}

Here are a couple of plots that are generated by running siman_test in the following way:

./siman_test | grep -v "^#"
| xyplot -xyil -y -0.88 -0.83 -d "x...y"
| xyps -d > siman-test.eps
./siman_test | grep -v "^#"
| xyplot -xyil -xl "generation" -yl "energy" -d "x..y"
| xyps -d > siman-energy.eps

Example of a simulated annealing run: at higher temperatures (early in the plot) you see that the solution can fluctuate, but at lower temperatures it converges.

23.3.2 Traveling Salesman Problem

The TSP (Traveling Salesman Problem) is the classic combinatorial optimization problem. I have provided a very simple version of it, based on the coordinates of twelve cities in the southwestern United States. This should maybe be called the Flying Salesman Problem, since I am using the great-circle distance between cities, rather than the driving distance. Also: I assume the earth is a sphere, so I don’t use geoid distances.

The gsl_siman_solve() routine finds a route which is 3490.62 Kilometers long; this is confirmed by an exhaustive search of all possible routes with the same initial city.

The full code can be found in ‘siman/siman_tsp.c’, but I include here some plots generated with in the following way:

./siman_tsp > tsp.output
grep -v "^#" tsp.output
  | xyplot -xyil -d "x................y"
           -lx "generation" -ly "distance"
           -lt "TSP -- 12 southwest cities"
  | xyps -d > 12-cities.eps
grep initial_city_coord tsp.output
  | awk '{print $2, $3, $4, $5}'
  | xyplot -xyil -lb0 -cs 0.8
           -lx "longitude (- means west)"
           -ly "latitude"
           -lt "TSP -- initial-order"
  | xyps -d > initial-route.eps
grep final_city_coord tsp.output
  | awk '{print $2, $3, $4, $5}'
  | xyplot -xyil -lb0 -cs 0.8
           -lx "longitude (- means west)"
           -ly "latitude"
           -lt "TSP -- final-order"
  | xyps -d > final-route.eps

This is the output showing the initial order of the cities; longitude is negative, since it is west and I want the plot to look like a map.

# initial coordinates of cities (longitude and latitude)
###initial_city_coord: -105.95 35.68 Santa Fe
###initial_city_coord: -112.07 33.54 Phoenix
###initial_city_coord: -106.62 35.12 Albuquerque
###initial_city_coord: -103.2 34.41 Clovis
###initial_city_coord: -107.87 37.29 Durango
###initial_city_coord: -96.77 32.79 Dallas
###initial_city_coord: -105.92 35.77 Tesuque
###initial_city_coord: -107.84 35.15 Grants
###initial_city_coord: -106.28 35.89 Los Alamos
###initial_city_coord: -106.76 32.34 Las Cruces
###initial_city_coord: -108.58 37.35 Cortez
###initial_city_coord: -108.74 35.52 Gallup
###initial_city_coord: -105.95 35.68 Santa Fe

The optimal route turns out to be:

# final coordinates of cities (longitude and latitude)
###final_city_coord: -105.95 35.68 Santa Fe
###final_city_coord: -106.28 35.89 Los Alamos
###final_city_coord: -106.62 35.12 Albuquerque
###final_city_coord: -107.84 35.15 Grants
###final_city_coord: -107.87 37.29 Durango
###final_city_coord: -108.58 37.35 Cortez
###final_city_coord: -108.74 35.52 Gallup
###final_city_coord: -112.07 33.54 Phoenix
###final_city_coord: -106.76 32.34 Las Cruces
###final_city_coord: -96.77 32.79 Dallas
###final_city_coord: -103.2 34.41 Clovis
###final_city_coord: -105.92 35.77 Tesuque

     ###final_city_coord: -105.95 35.68 Santa Fe

Initial and final (optimal) route for the 12 southwestern cities Flying Salesman Problem.

Here’s a plot of the cost function (energy) versus generation (point in the calculation at which a new temperature is set) for this problem:

Chapter 23: Simulated Annealing 269 Example of a simulated annealing run for the 12 southwestern cities Flying Salesman Problem.

24 Ordinary Differential Equations

This chapter describes functions for solving ordinary differential equation (ODE) initial value problems. The library provides a variety of low-level methods, such as Runge-Kutta and Bulirsch-Stoer routines, and higher-level components for adaptive step-size control. The components can be combined by the user to achieve the desired solution, with full access to any intermediate steps.

These functions are declared in the header file ‘gsl_odeiv.h’.

24.1 Defining the ODE System

The routines solve the general n-dimensional first-order system,

\frac{dy_{i}(t)}{dt} = f_{i}(t, y_{1}(t), … y_{n}(t))

for i = 1, …, n. The stepping functions rely on the vector of derivatives f_{i} and the Jacobian matrix, J_{ij} = ∂f_{i}(t, y(t))/∂y_{j}. A system of equations is defined using the gsl_odeiv_system datatype.

Data Type gsl odeiv system

This data type defines a general ODE system with arbitrary parameters.

     int (* function) (double t, const double y[], double dydt[], void * params)

This function should store the elements of f(t, y, params) in the array dydt, for arguments (t,y) and parameters params

     int (* jacobian) (double t, const double y[], double * dfdy, double dfdt[],
     void * params);

This function should store the elements of f(t, y, params) in the array dfdt and the Jacobian matrix J_{ij} in the the array dfdy regarded as a rowordered matrix J(i,j) = dfdy[i * dim + j] where dim is the dimension of the system.

size_t dimension; This is the dimension of the system of equations

void * params This is a pointer to the arbitrary parameters of the system.

24.2 Stepping Functions

The lowest level components are the stepping functions which advance a solution from time t to t + h for a fixed step-size h and estimate the resulting local error.

Function gsl_odeiv_step * gsl odeiv step alloc (const

    gsl_odeiv_step_type * T, size_t dim)

This function returns a pointer to a newly allocated instance of a stepping function of type T for a system of dim dimensions.

Function int gsl odeiv step reset (gsl_odeiv_step * s)

This function resets the stepping function s. It should be used whenever the next use of s will not be a continuation of a previous step.

Function void gsl odeiv step free (gsl_odeiv_step * s)

This function frees all the memory associated with the stepping function s.

Function const char * gsl odeiv step name (const gsl_odeiv_step * s)

This function returns a pointer to the name of the stepping function. For example,

     printf("step method is '%s'\n",
             gsl_odeiv_step_name (s));

would print something like step method is 'rk4'.

Function unsigned int gsl odeiv step order (const gsl_odeiv_step * s)

This function returns the order of the stepping function on the previous step. This order can vary if the stepping function itself is adaptive.

Function int gsl odeiv step apply (gsl_odeiv_step * s, double t, double
    h, double y[], double yerr[], const double dydt_in[], double dydt_out[],
    const gsl_odeiv_system * dydt)

This function applies the stepping function s to the system of equations defined by dydt, using the step size h to advance the system from time t and state y to time t+h. The new state of the system is stored in y on output, with an estimate of the absolute error in each component stored in yerr. If the argument dydt in is not null it should point an array containing the derivatives for the system at time t on input. This is optional as the derivatives will be computed internally if they are not provided, but allows the reuse of existing derivative information. On output the new derivatives of the system at time t+h will be stored in dydt out if it is not null.

The following algorithms are available,

Step Type gsl odeiv step rk2

Embedded 2nd order Runge-Kutta with 3rd order error estimate.

Step Type gsl odeiv step rk4

4th order (classical) Runge-Kutta.

Step Type gsl odeiv step rkf45

Embedded 4th order Runge-Kutta-Fehlberg method with 5th order error estimate. This method is a good general-purpose integrator.

Step Type gsl odeiv step rkck

Embedded 4th order Runge-Kutta Cash-Karp method with 5th order error estimate.

Step Type gsl odeiv step rk8pd

Embedded 8th order Runge-Kutta Prince-Dormand method with 9th order error estimate.

Step Type gsl odeiv step rk2imp

Implicit 2nd order Runge-Kutta at Gaussian points

Step Type gsl odeiv step rk4imp

Implicit 4th order Runge-Kutta at Gaussian points

Step Type gsl odeiv step bsimp

Implicit Bulirsch-Stoer method of Bader and Deuflhard.

Step Type gsl odeiv step gear1

M=1 implicit Gear method

Step Type gsl odeiv step gear2

M=2 implicit Gear method

24.3 Adaptive Step-size Control

The control function examines the proposed change to the solution and its error estimate produced by a stepping function and attempts to determine the optimal step-size for a userspecified level of error.

Function gsl_odeiv_control * gsl odeiv control standard new (double

    eps abs, double eps rel, double a y, double a dydt)

The standard control object is a four parameter heuristic based on absolute and relative errors eps abs and eps rel, and scaling factors a y and a dydt for the system state y(t) and derivatives y′(t) respectively. The step-size adjustment procedure for this method begins by computing the desired error level D_{i} for each component,

D_{i} = ϵ_{abs} + ϵ_{rel} ∗ (a_{y}|y_{i}| + a_{dydt}h|y′_{i}|)

and comparing it with the observed error E_{i} = |yerr_{i}|. If the observed error E exceeds the desired error level D by more than 10% for any component then the method reduces the step-size by an appropriate factor,

h_{new} = h_{old} ∗ S ∗ (D/E)^{1/q}

where q is the consistency order of method (e.g. q = 4 for 4(5) embedded RK), and S is a safety factor of 0.9. The ratio D/E is taken to be the maximum of the ratios D_{i}/E_{i}. If the observed error E is less than 50% of the desired error level D for the maximum ratio D_{i}/E_{i} then the algorithm takes the opportunity to increase the step-size to bring the error in line with the desired level,

h_{new} = h_{old} ∗ S ∗ (E/D)^{1/(q+1)}

This encompasses all the standard error scaling methods.

Function gsl_odeiv_control * gsl odeiv control y new (double eps abs,

double eps rel) This function creates a new control object which will keep the local error on each step within an absolute error of eps abs and relative error of eps rel with respect to the solution y_{i}(t). This is equivalent to the standard control object with a y=1 and a dydt=0.

Function gsl_odeiv_control * gsl odeiv control yp new (double eps abs,

double eps rel) This function creates a new control object which will keep the local error on each step within an absolute error of eps abs and relative error of eps rel with respect to the derivatives of the solution y′_{i}(t) . This is equivalent to the standard control object with a y=0 and a dydt=1.

Function gsl_odeiv_control * gsl odeiv control alloc (const

    gsl_odeiv_control_type * T)

This function returns a pointer to a newly allocated instance of a control function of type T. This function is only needed for defining new types of control functions. For most purposes the standard control functions described above should be sufficient.

Function int gsl odeiv control init (gsl_odeiv_control * c, double

    eps abs, double eps rel, double a y, double a dydt)

This function initializes the control function c with the parameters eps abs (absolute error), eps rel (relative error), a y (scaling factor for y) and a dydt (scaling factor for derivatives).

Function void gsl odeiv control free (gsl_odeiv_control * c)

This function frees all the memory associated with the control function c.

Function int gsl odeiv control hadjust (gsl_odeiv_control * c,

    gsl_odeiv_step * s, const double y0[], const double yerr[], const double
    dydt[], double * h)

This function adjusts the step-size h using the control function c, and the current values of y, yerr and dydt. The stepping function step is also needed to determine the order of the method. If the error in the y-values yerr is found to be too large then the step-size h is reduced and the function returns GSL_ODEIV_HADJ_DEC. If the error is sufficiently small then h may be increased and GSL_ODEIV_HADJ_INC is returned. The function returns GSL_ODEIV_HADJ_NIL if the step-size is unchanged. The goal of the function is to estimate the largest step-size which satisfies the user-specified accuracy requirements for the current point.

Function const char * gsl odeiv control name (const gsl_odeiv_control

* c) This function returns a pointer to the name of the control function. For example,

     printf("control method is '%s'\n",
            gsl_odeiv_control_name (c));
would print something like control method is 'standard'

24.4 Evolution

The highest level of the system is the evolution function which combines the results of a stepping function and control function to reliably advance the solution forward over an interval (t_{0}, t_{1}). If the control function signals that the step-size should be decreased the evolution function backs out of the current step and tries the proposed smaller step-size. This is process is continued until an acceptable step-size is found.

Function gsl_odeiv_evolve * gsl odeiv evolve alloc (size_t dim)

This function returns a pointer to a newly allocated instance of an evolution function for a system of dim dimensions.

Function int gsl odeiv evolve apply (gsl_odeiv_evolve * e,

    gsl_odeiv_control * con, gsl_odeiv_step * step, const gsl_odeiv_system
    * dydt, double * t, double t1, double * h, double y[])

This function advances the system (e, dydt) from time t and position y using the stepping function step. The new time and position are stored in t and y on output. The initial step-size is taken as h, but this will be modified using the control function c to achieve the appropriate error bound if necessary. The routine may make several calls to step in order to determine the optimum step-size. If the step-size has been changed the value of h will be modified on output. The maximum time t1 is guaranteed not to be exceeded by the time-step. On the final time-step the value of t will be set to t1 exactly.

Function int gsl odeiv evolve reset (gsl_odeiv_evolve * e)

This function resets the evolution function e. It should be used whenever the next use of e will not be a continuation of a previous step.

Function void gsl odeiv evolve free (gsl_odeiv_evolve * e)

This function frees all the memory associated with the evolution function e.

24.5 Examples

The following program solves the second-order nonlinear Van der Pol oscillator equation,

x′′(t) + μx′(t)(x(t)^{2} − 1) + x(t) = 0

This can be converted into a first order system suitable for use with the library by introducing a separate variable for the velocity, y = x′(t),

\begin{aligned}x′ & = y \\ y′ & = −x + μy(1 − x^{2}) \end{aligned}

The program begins by defining functions for these derivatives and their Jacobian,

#include <stdio.h>
#include <gsl/gsl_errno.h>
#include <gsl/gsl_matrix.h>
#include <gsl/gsl_odeiv.h>
int
func (double t, const double y[], double f[],
    void *params)
{
double mu = *(double *)params;
f[0] = y[1];
f[1] = -y[0] - mu*y[1]*(y[0]*y[0] - 1);
return GSL_SUCCESS;
}
int
jac (double t, const double y[], double *dfdy,
   double dfdt[], void *params)
{
double mu = *(double *)params;
gsl_matrix_view dfdy_mat
  = gsl_matrix_view_array (dfdy, 2, 2);
gsl_matrix * m = &dfdy_mat.matrix;
gsl_matrix_set (m, 0, 0, 0.0);
gsl_matrix_set (m, 0, 1, 1.0);
gsl_matrix_set (m, 1, 0, -2.0*mu*y[0]*y[1] - 1.0);
gsl_matrix_set (m, 1, 1, -mu*(y[0]*y[0] - 1.0));
dfdt[0] = 0.0;
dfdt[1] = 0.0;
return GSL_SUCCESS;
}
int
main (void)
{
const gsl_odeiv_step_type * T
  = gsl_odeiv_step_rk8pd;
gsl_odeiv_step * s
  = gsl_odeiv_step_alloc (T, 2);

gsl_odeiv_control * c
  = gsl_odeiv_control_y_new (1e-6, 0.0);
gsl_odeiv_evolve * e
  = gsl_odeiv_evolve_alloc (2);
double mu = 10;
gsl_odeiv_system sys = {func, jac, 2, &mu};
double t = 0.0, t1 = 100.0;
double h = 1e-6;
double y[2] = { 1.0, 0.0 };
gsl_ieee_env_setup();
while (t < t1)
  {
    int status = gsl_odeiv_evolve_apply (e, c, s,
                                         &sys,
                                         &t, t1,
                                         &h, y);
    if (status != GSL_SUCCESS)
        break;
    printf("%.5e %.5e %.5e\n", t, y[0], y[1]);
  }
gsl_odeiv_evolve_free(e);
gsl_odeiv_control_free(c);
gsl_odeiv_step_free(s);
return 0;
}

The main loop of the program evolves the solution from (y, y′) = (1, 0) at t = 0 to t = 100. The step-size h is automatically adjusted by the controller to maintain an absolute accuracy of 10^{−6} in the function values y.

Numerical solution of the Van der Pol oscillator equation using Prince-Dormand 8th order Runge-Kutta.

To obtain the values at regular intervals, rather than the variable spacings chosen by the control function, the main loop can be modified to advance the solution from one point to the next. For example, the following main loop prints the solution at the fixed points t = 0, 1, 2, …, 100,

       for (i = 1; i <= 100; i++)
         {
           double ti = i * t1 / 100.0;
           while (t < ti)
             {
               gsl_odeiv_evolve_apply (e, c, s,
                                       &sys,
                                       &t, ti, &h,
                                       y);
             }
           printf("%.5e %.5e %.5e\n", t, y[0], y[1]);
         }

24.6 References and Further Reading

Many of the the basic Runge-Kutta formulas can be found in the Handbook of Mathematical Functions,

Abramowitz & Stegun (eds.), Handbook of Mathematical Functions, Section 25.5.

The implicit Bulirsch-Stoer algorithm bsimp is described in the following paper, G. Bader and P. Deuflhard, “A Semi-Implicit Mid-Point Rule for Stiff Systems of Ordinary Differential Equations.”, Numer. Math. 41, 373-398, 1983.

25 Interpolation

This chapter describes functions for performing interpolation. The library provides a variety of interpolation methods, including Cubic splines and Akima splines. The interpolation types are interchangeable, allowing different methods to be used without recompiling. Interpolations can be defined for both normal and periodic boundary conditions. Additional functions are available for computing derivatives and integrals of interpolating functions.

The functions described in this section are declared in the header files ‘gsl_interp.h’

and ‘gsl_spline.h’.

25.1 Introduction

Given a set of data points (x_{1}, y_{1}) … (x_{n}, y_{n}) the routines described in this section compute a continuous interpolating function y(x) such that y_{i} = y(x_{i}). The interpolation is piecewise smooth, and its behavior at the points is determined by the type of interpolation used.

25.2 Interpolation Functions

The interpolation function for a given dataset is stored in a gsl_interp object. These are created by the following functions.

Function gsl_interp * gsl interp alloc (const gsl_interp_type * T,
         size_t size)

This function returns a pointer to a newly allocated interpolation object of type T for size data-points.

Function int gsl interp init (gsl_interp * interp, const double xa[], const
         double ya[], size_t size)

This function initializes the interpolation object interp for the data (xa,ya) where xa and ya are arrays of size size. The interpolation object (gsl_interp) does not save the data arrays xa and ya and only stores the static state computed from the data. The xa data array is always assumed to be strictly ordered; the behavior for other arrangements is not defined.

Function void gsl interp free (gsl_interp * interp)

This function frees the interpolation object interp.

25.3 Interpolation Types

The interpolation library provides five interpolation types:

Interpolation Type gsl interp linear

Linear interpolation. This interpolation method does not require any additional memory.

Interpolation Type gsl interp cspline

Cubic spline with natural boundary conditions

Interpolation Type gsl interp cspline periodic

Cubic spline with periodic boundary conditions

Interpolation Type gsl interp akima

Akima spline with natural boundary conditions

Interpolation Type gsl interp akima periodic

Akima spline with periodic boundary conditions

The following related functions are available,

Function const char * gsl interp name (const gsl_interp * interp)

This function returns the name of the interpolation type used by interp. For example,

     printf("interp uses '%s' interpolation\n",
            gsl_interp_name (interp));

would print something like,

     interp uses 'cspline' interpolation.

Function unsigned int gsl interp min size (const gsl_interp * interp)

This function returns the minimum number of points required by the interpolation type of interp. For example, cubic interpolation requires a minimum of 3 points.

25.4 Index Look-up and Acceleration

The state of searches can be stored in a gsl_interp_accel object, which is a kind of iterator for interpolation lookups. It caches the previous value of an index lookup. When the subsequent interpolation point falls in the same interval its index value can be returned immediately.

Function size_t gsl interp bsearch (const double x_array[], double x,

size_t index lo, size_t index hi) This function returns the index i of the array x array such that x_array[i] <= x < x_array[i+1]. The index is searched for in the range [index lo,index hi].

Function gsl_interp_accel * gsl interp accel alloc (void)

This function returns a pointer to an accelerator object, which is a kind of iterator for interpolation lookups. It tracks the state of lookups, thus allowing for application of various acceleration strategies.

Function size_t gsl interp accel find (gsl_interp_accel * a, const double
    x_array[], size_t size, double x)

This function performs a lookup action on the data array x array of size size, using the given accelerator a. This is how lookups are performed during evaluation of an interpolation. The function returns an index i such that xarray[i] <= x <

xarray[i+1].

Function void gsl interp accel free (gsl_interp_accel* a)

This function frees the accelerator object a.

25.5 Evaluation of interpolating functions

Function double gsl interp eval (const gsl_interp * interp, const double
    xa[], const double ya[], double x, gsl_interp_accel * a)
Function int gsl interp eval e (const gsl_interp * interp, const double
    xa[], const double ya[], double x, gsl_interp_accel * a, double * y)

These functions return the interpolated value of y for a given point x, using the interpolation object interp, data arrays xa and ya and the accelerator a.

Function double gsl interp eval deriv (const gsl_interp * interp, const

    double xa[], const double ya[], double x, gsl_interp_accel * a)

Function int gsl interp eval deriv e (const gsl_interp * interp, const

    double xa[], const double ya[], double x, gsl_interp_accel * a, double *

d) These functions return the derivative d of an interpolated function for a given point x, using the interpolation object interp, data arrays xa and ya and the accelerator a.

Function double gsl interp eval deriv2 (const gsl_interp * interp, const

         double xa[], const double ya[], double x, gsl_interp_accel * a)

Function int gsl interp eval deriv2 e (const gsl_interp * interp, const

         double xa[], const double ya[], double x, gsl_interp_accel * a, double *

d2) These functions return the second derivative d2 of an interpolated function for a given point x, using the interpolation object interp, data arrays xa and ya and the accelerator a.

Function double gsl interp eval integ (const gsl_interp * interp, const

         double xa[], const double ya[], double a, double bdouble x,
         gsl_interp_accel * a)

Function int gsl interp eval integ e (const gsl_interp * interp, const

         double xa[], const double ya[], , double a, double b, gsl_interp_accel *
         a, double * result)

These functions return the numerical integral result of an interpolated function over the range [a, b], using the interpolation object interp, data arrays xa and ya and the accelerator a.

25.6 Higher-level interface

The functions described in the previous sections required the user to supply pointers to the x and y arrays on each call. The following functions are equivalent to the corresponding gsl_interp functions but maintain a copy of this data in the gsl_spline object. This removes the need to pass both xa and ya as arguments on each evaluation. These functions are defined in the header file ‘gsl_spline.h’.

Function gsl_spline * gsl spline alloc (const gsl_interp_type * T, size_t

n)

Function int gsl spline init (gsl_spline * spline, const double xa[], const
         double ya[], size_t size)

Function void gsl spline free (gsl_spline * spline)

Function double gsl spline eval (const gsl_spline * spline, double x,
         gsl_interp_accel * a)

Function int gsl spline eval e (const gsl_spline * spline, double x,

         gsl_interp_accel * a, double * y)

Function double gsl spline eval deriv (const gsl_spline * spline, double x,

         gsl_interp_accel * a)

Function int gsl spline eval deriv e (const gsl_spline * spline, double x,

         gsl_interp_accel * a, double * d)

Function double gsl spline eval deriv2 (const gsl_spline * spline, double

         x, gsl_interp_accel * a)

Function int gsl spline eval deriv2 e (const gsl_spline * spline, double x,

         gsl_interp_accel * a, double * d2)

Function double gsl spline eval integ (const gsl_spline * spline, double a,

    double b, gsl_interp_accel * acc)

Function int gsl spline eval integ e (const gsl_spline * spline, double a,

    double b, gsl_interp_accel * acc, double * result)

25.7 Examples

The following program demonstrates the use of the interpolation and spline functions. It computes a cubic spline interpolation of the 10-point dataset (x_{i}, y_{i}) where x_{i} = i + sin(i)/2 and y_{i} = i + cos(i^{2}) for i = 0 … 9.

#include <config.h>
#include <stdlib.h>
#include <stdio.h>
#include <math.h>
#include <gsl/gsl_errno.h>
#include <gsl/gsl_spline.h>
int
main (void)
{
  int i;
  double xi, yi, x[10], y[10];
  printf ("#m=0,S=2\n");
  for (i = 0; i < 10; i++)
    {
      x[i] = i + 0.5 * sin (i);
      y[i] = i + cos (i * i);
      printf ("%g %g\n", x[i], y[i]);
    }
  printf ("#m=1,S=0\n");
  {
    gsl_interp_accel *acc
      = gsl_interp_accel_alloc ();
    gsl_spline *spline
      = gsl_spline_alloc (gsl_interp_cspline, 10);
    gsl_spline_init (spline, x, y, 10);
    for (xi = x[0]; xi < x[9]; xi += 0.01)
      {
        double yi = gsl_spline_eval (spline, xi, acc);
        printf ("%g %g\n", xi, yi);
      }
    gsl_spline_free (spline);
    gsl_interp_accel_free(acc);
  }
  return 0;

     }

The output is designed to be used with the gnu plotutils graph program,

     $ ./a.out > interp.dat
     $ graph -T ps < interp.dat > interp.ps

The result shows a smooth interpolation of the original points. The interpolation method can changed simply by varying the first argument of gsl_spline_alloc.

25.8 References and Further Reading

Descriptions of the interpolation algorithms and further references can be found in the following book,

C.W. Ueberhuber, Numerical Computation (Volume 1), Chapter 9 "Interpolation", Springer (1997), ISBN 3-540-62058-3.

26 Numerical Differentiation

The functions described in this chapter compute numerical derivatives by finite differencing. An adaptive algorithm is used to find the best choice of finite difference and to estimate the error in the derivative. These functions are declared in the header file ‘gsl_diff.h’

26.1 Functions

Function int gsl diff central (const gsl_function *f, double x, double

*result, double *abserr) This function computes the numerical derivative of the function f at the point x using an adaptive central difference algorithm. The derivative is returned in result and an estimate of its absolute error is returned in abserr.

Function int gsl diff forward (const gsl_function *f, double x, double

*result, double *abserr) This function computes the numerical derivative of the function f at the point x using an adaptive forward difference algorithm. The function is evaluated only at points greater than x and at x itself. The derivative is returned in result and an estimate of its absolute error is returned in abserr. This function should be used if f(x) has a singularity or is undefined for values less than x.

Function int gsl diff backward (const gsl_function *f, double x, double

*result, double *abserr) This function computes the numerical derivative of the function f at the point x using an adaptive backward difference algorithm. The function is evaluated only at points less than x and at x itself. The derivative is returned in result and an estimate of its absolute error is returned in abserr. This function should be used if f(x) has a singularity or is undefined for values greater than x.

26.2 Example

The following code estimates the derivative of the function f(x) = x^{3/2} at x = 2 and at x = 0. The function f(x) is undefined for x < 0 so the derivative at x = 0 is computed

using gsl_diff_forward.
#include <stdio.h>
#include <gsl/gsl_math.h>
#include <gsl/gsl_diff.h>
double f (double x, void * params)
{
  return pow (x, 1.5);
}
int
main (void)
{
  gsl_function F;

double result, abserr;
F.function = &f;
F.params = 0;
printf("f(x) = x^(3/2)\n");
gsl_diff_central (&F, 2.0, &result, &abserr);
printf("x = 2.0\n");
printf("f'(x) = %.10f +/- %.5f\n", result, abserr);
printf("exact = %.10f\n\n", 1.5 * sqrt(2.0));
gsl_diff_forward (&F, 0.0, &result, &abserr);
printf("x = 0.0\n");
printf("f'(x) = %.10f +/- %.5f\n", result, abserr);
printf("exact = %.10f\n", 0.0);
return 0;
}

Here is the output of the program,

$ ./demo
f(x) = x^(3/2)
x = 2.0
f'(x) = 2.1213203435 +/- 0.01490
exact = 2.1213203436
x = 0.0
f'(x) = 0.0012172897 +/- 0.05028
exact = 0.0000000000

26.3 References and Further Reading

The algorithms used by these functions are described in the following book, S.D. Conte and Carl de Boor, Elementary Numerical Analysis: An Algorithmic Approach, McGraw-Hill, 1972.

27 Chebyshev Approximations

This chapter describes routines for computing Chebyshev approximations to univariate functions. A Chebyshev approximation is a truncation of the series f(x) = ∑ c_{n}T_{n}(x), where the Chebyshev polynomials T_{n}(x) = cos(n arccos x) provide an orthogonal basis of polynomials on the interval [−1, 1] with the weight function 1/\sqrt{1 − x^{2}}. The first few Chebyshev polynomials are, T_{0}(x) = 1, T_{1}(x) = x, T_{2}(x) = 2x^{2} − 1.

The functions described in this chapter are declared in the header file ‘gsl_chebyshev.h’.

27.1 The gsl cheb series struct

A Chebyshev series is stored using the following structure,

     typedef struct
     {
       double * c;   /* coefficients  c[0] .. c[order] */
       int order;    /* order of expansion             */
       double a;     /* lower interval point           */
       double b;     /* upper interval point           */
     } gsl_cheb_struct

The approximation is made over the range [a, b] using order+1 terms, including the coefficient c[0].

27.2 Creation and Calculation of Chebyshev Series

Function gsl_cheb_series * gsl cheb alloc (const size_t n)

This function allocates space for a Chebyshev series of order n and returns a pointer

     to a new gsl_cheb_series struct.

Function void gsl cheb free (gsl_cheb_series * cs)

This function frees a previously allocated Chebyshev series cs.

Function int gsl cheb init (gsl_cheb_series * cs, const gsl_function * f,
         const double a, const double b)

This function computes the Chebyshev approximation cs for the function f over the range (a, b) to the previously specified order. The computation of the Chebyshev approximation is an O(n^{2}) process, and requires n function evaluations.

27.3 Chebyshev Series Evaluation

Function double gsl cheb eval (const gsl_cheb_series * cs, double x)

This function evaluates the Chebyshev series cs at a given point x.

Function int gsl cheb eval err (const gsl_cheb_series * cs, const double
    x, double * result, double * abserr)

This function computes the Chebyshev series cs at a given point x, estimating both the series result and its absolute error abserr. The error estimate is made from the first neglected term in the series.

Function double gsl cheb eval n (const gsl_cheb_series * cs, size_t order,
    double x)

This function evaluates the Chebyshev series cs at a given point n, to (at most) the given order order.

Function int gsl cheb eval n err (const gsl_cheb_series * cs, const
    size_t order, const double x, double * result, double * abserr)

This function evaluates a Chebyshev series cs at a given point x, estimating both the series result and its absolute error abserr, to (at most) the given order order. The error estimate is made from the first neglected term in the series.

27.4 Derivatives and Integrals

The following functions allow a Chebyshev series to be differentiated or integrated, producing a new Chebyshev series. Note that the error estimate produced by evaluating the derivative series will be underestimated due to the contribution of higher order terms being neglected.

Function int gsl cheb calc deriv (gsl_cheb_series * deriv, const

    gsl_cheb_series * cs)

This function computes the derivative of the series cs, storing the derivative coefficients in the previously allocated deriv. The two series cs and deriv must have been allocated with the same order.

Function int gsl cheb calc integ (gsl_cheb_series * integ, const

    gsl_cheb_series * cs)

This function computes the integral of the series cs, storing the integral coefficients in the previously allocated integ. The two series cs and integ must have been allocated with the same order. The lower limit of the integration is taken to be the left hand end of the range a.

27.5 Examples

The following example program computes Chebyshev approximations to a step function. This is an extremely difficult approximation to make, due to the discontinuity, and was chosen as an example where approximation error is visible. For smooth functions the Chebyshev approximation converges extremely rapidly and errors would not be visible.

#include <stdio.h>
#include <gsl/gsl_math.h>
#include <gsl/gsl_chebyshev.h>
double
f (double x, void *p)
{
  if (x < 0.5)
    return 0.25;
  else
    return 0.75;
}
int
main (void)
{
  int i, n = 10000;
  gsl_cheb_series *cs = gsl_cheb_alloc (40);
  gsl_function F;
  F.function = f;
  F.params = 0;
  gsl_cheb_init (cs, &F, 0.0, 1.0);
  for (i = 0; i < n; i++)
    {
      double x = i / (double)n;
      double r10 = gsl_cheb_eval_n (cs, 10, x);
      double r40 = gsl_cheb_eval (cs, x);
      printf ("%g %g %g %g\n",
              x, GSL_FN_EVAL (&F, x), r10, r40);
    }
  gsl_cheb_free (cs);
  return 0;
}

The output from the program gives the original function, 10-th order approximation and 40-th order approximation, all sampled at intervals of 0.001 in x.

Chapter 27: Chebyshev Approximations 289

27.6 References and Further Reading

The following paper describes the use of Chebyshev series, R. Broucke, “Ten Subroutines for the Manipulation of Chebyshev Series [C1] (Algorithm 446)”. Communications of the ACM 16(4), 254-256 (1973)

28 Series Acceleration

The functions described in this chapter accelerate the convergence of a series using the Levin u-transform. This method takes a small number of terms from the start of a series and uses a systematic approximation to compute an extrapolated value and an estimate of its error. The u-transform works for both convergent and divergent series, including asymptotic series.

These functions are declared in the header file ‘gsl_sum.h’.

28.1 Acceleration functions

The following functions compute the full Levin u-transform of a series with its error estimate. The error estimate is computed by propagating rounding errors from each term through to the final extrapolation.

These functions are intended for summing analytic series where each term is known to high accuracy, and the rounding errors are assumed to originate from finite precision. They are taken to be relative errors of order GSL_DBL_EPSILON for each term.

The calculation of the error in the extrapolated value is an O(N^{2}) process, which is expensive in time and memory. A faster but less reliable method which estimates the error from the convergence of the extrapolated value is described in the next section For the method described here a full table of intermediate values and derivatives through to O(N) must be computed and stored, but this does give a reliable error estimate. .

Function gsl_sum_levin_u_workspace * gsl sum levin u alloc (size_t n)

This function allocates a workspace for a Levin u-transform of n terms. The size of the workspace is O(2n^{2} + 3n).

Function int gsl sum levin u free (gsl_sum_levin_u_workspace * w)

This function frees the memory associated with the workspace w.

Function int gsl sum levin u accel (const double * array, size_t

         array size, gsl_sum_levin_u_workspace * w, double * sum accel, double *

abserr) This function takes the terms of a series in array of size array size and computes the extrapolated limit of the series using a Levin u-transform. Additional working space must be provided in w. The extrapolated sum is stored in sum accel, with an estimate of the absolute error stored in abserr. The actual term-by-term sum is returned in w->sum_plain. The algorithm calculates the truncation error (the difference between two successive extrapolations) and round-off error (propagated from the individual terms) to choose an optimal number of terms for the extrapolation.

28.2 Acceleration functions without error estimation

The functions described in this section compute the Levin u-transform of series and attempt to estimate the error from the "truncation error" in the extrapolation, the difference between the final two approximations. Using this method avoids the need to compute an intermediate table of derivatives because the error is estimated from the behavior of the extrapolated value itself. Consequently this algorithm is an O(N) process and only requires O(N) terms of storage. If the series converges sufficiently fast then this procedure can be acceptable. It is appropriate to use this method when there is a need to compute many extrapolations of series with similar converge properties at high-speed. For example, when numerically integrating a function defined by a parameterized series where the parameter varies only slightly. A reliable error estimate should be computed first using the full algorithm described above in order to verify the consistency of the results.

Function gsl_sum_levin_utrunc_workspace *

gsl sum levin utrunc alloc (size_t n)

This function allocates a workspace for a Levin u-transform of n terms, without error estimation. The size of the workspace is O(3n).

Function int gsl sum levin utrunc free (gsl_sum_levin_utrunc_workspace

* w) This function frees the memory associated with the workspace w.

Function int gsl sum levin utrunc accel (const double * array, size_t

    array size, gsl_sum_levin_utrunc_workspace * w, double * sum accel,

double * abserr trunc) This function takes the terms of a series in array of size array size and computes the extrapolated limit of the series using a Levin u-transform. Additional working space must be provided in w. The extrapolated sum is stored in sum accel. The actual term-by-term sum is returned in w->sum_plain. The algorithm terminates when the difference between two successive extrapolations reaches a minimum or is sufficiently small. The difference between these two values is used as estimate of the error and is stored in abserr trunc. To improve the reliability of the algorithm the extrapolated values are replaced by moving averages when calculating the truncation error, smoothing out any fluctuations.

28.3 Example of accelerating a series

The following code calculates an estimate of ζ(2) = π^{2}/6 using the series,

ζ(2) = 1 + 1/2^{2} + 1/3^{2} + 1/4^{2} + …

After N terms the error in the sum is O(1/N), making direct summation of the series converge slowly.

#include <stdio.h>
#include <gsl/gsl_math.h>
#include <gsl/gsl_sum.h>
#define N 20

int
main (void)
{
double t[N];
double sum_accel, err;
double sum = 0;
int n;
gsl_sum_levin_u_workspace * w
  = gsl_sum_levin_u_alloc (N);
const double zeta_2 = M_PI * M_PI / 6.0;
/* terms for zeta(2) = \sum_{n=0}^{\infty} 1/n^2 */
for (n = 0; n < N; n++)
  {
    double np1 = n + 1.0;
    t[n] = 1.0 / (np1 * np1);
    sum += t[n];
  }
gsl_sum_levin_u_accel (t, N, w, &sum_accel, &err);
printf("term-by-term sum = % .16f using %d terms\n",
       sum, N);
printf("term-by-term sum = % .16f using %d terms\n",
       w->sum_plain, w->terms_used);
printf("exact value      = % .16f\n", zeta_2);
printf("accelerated sum  = % .16f using %d terms\n",
       sum_accel, w->terms_used);
printf("estimated error  = % .16f\n", err);
printf("actual error     = % .16f\n",
       sum_accel - zeta_2);
gsl_sum_levin_u_free (w);
return 0;
}

The output below shows that the Levin u-transform is able to obtain an estimate of the sum to 1 part in 10^{10} using the first eleven terms of the series. The error estimate returned by the function is also accurate, giving the correct number of significant digits.

bash$ ./a.out
term-by-term sum =  1.5961632439130233 using 20 terms
term-by-term sum =  1.5759958390005426 using 13 terms
exact value      =  1.6449340668482264
accelerated sum  =  1.6449340668166479 using 13 terms
estimated error  =  0.0000000000508580
actual error     = -0.0000000000315785

Note that a direct summation of this series would require 10^{10} terms to achieve the same precision as the accelerated sum does in 13 terms.

28.4 References and Further Reading

The algorithms used by these functions are described in the following papers, T. Fessler, W.F. Ford, D.A. Smith, hurry: An acceleration algorithm for scalar sequences and series ACM Transactions on Mathematical Software, 9(3):346–354, 1983. and Algorithm 602 9(3):355–357, 1983.

The theory of the u-transform was presented by Levin, D. Levin, Development of Non-Linear Transformations for Improving Convergence of Sequences, Intern. J. Computer Math. B3:371–388, 1973

A review paper on the Levin Transform is available online, Herbert H. H. Homeier, Scalar Levin-Type Sequence Transformations,

http://xxx.lanl.gov/abs/math/0005209

29 Discrete Hankel Transforms

This chapter describes functions for performing Discrete Hankel Transforms (DHTs). The functions are declared in the header file ‘gsl_dht.h’.

29.1 Definitions

The discrete Hankel transform acts on a vector of sampled data, where the samples are assumed to have been taken at points related to the zeroes of a Bessel function of fixed order; compare this to the case of the discrete Fourier transform, where samples are taken at points related to the zeroes of the sine or cosine function.

Specifically, let f(t) be a function on the unit interval. Then the finite ν-Hankel transform of f(t) is defined to be the set of numbers g_{m} given by

g_{m} = ∫_{0}^{1} tdt J_{ν}(j_{ν,m}t)f(t),

so that

f(t) = ∑_{m=1}^{∞} \frac{2J_{ν}(j_{ν,m}x)}{J_{ν+1}(j_{ν,m})^{2}} g_{m}.

Suppose that f is band-limited in the sense that g_{m} = 0 for m > M. Then we have the following fundamental sampling theorem.

g_{m} = \frac{2}{j_{ν,M}^{2}} ∑_{k=1}^{M−1} f(\frac{j_{ν,k}}{j_{ν,M}})\frac{J_{ν}(j_{ν,m}j_{ν,k}/j_{ν,M})}{J_{ν+1}(j_{ν,k})^{2}}.

It is this discrete expression which defines the discrete Hankel transform. The kernel in the summation above defines the matrix of the ν-Hankel transform of size M − 1. The coefficients of this matrix, being dependent on ν and M, must be precomputed and stored; the gsl_dht object encapsulates this data. The allocation function gsl_dht_alloc returns a gsl_dht object which must be properly initialized with gsl_dht_init before it can be used to perform transforms on data sample vectors, for fixed ν and M, using the gsl_ dht_apply function. The implementation allows a scaling of the fundamental interval, for convenience, so that one take assume the function is defined on the interval [0, X], rather than the unit interval.

Notice that by assumption f(t) vanishes at the endpoints of the interval, consistent with the inversion formula and the sampling formula given above. Therefore, this transform corresponds to an orthogonal expansion in eigenfunctions of the Dirichlet problem for the Bessel differential equation.

29.2 Functions

Function gsl_dht * gsl dht alloc (size_t size)

This function allocates a Discrete Hankel transform object of size size.

Function int gsl dht init (gsl_dht * t, double nu, double xmax)

This function initializes the transform t for the given values of nu and x.

Function gsl_dht * gsl dht new (size_t size, double nu, double xmax)

This function allocates a Discrete Hankel transform object of size size and initializes it for the given values of nu and x.

Function void gsl dht free (gsl_dht * t)

This function frees the transform t.

Function int gsl dht apply (const gsl_dht * t, double * f in, double * f out)

This function applies the transform t to the array f in whose size is equal to the size of the transform. The result is stored in the array f out which must be of the same length.

Function double gsl dht x sample (const gsl_dht * t, int n)

This function returns the value of the n’th sample point in the unit interval, \frac{j_{ν,n+1}}{j_{ν,M}} X. These are the points where the function f(t) is assumed to be sampled.

Function double gsl dht k sample (const gsl_dht * t, int n)

This function returns the value of the n’th sample point in "k-space", \frac{j_{ν,n+1}}{X}.

29.3 References and Further Reading

The algorithms used by these functions are described in the following papers, H. Fisk Johnson, Comp. Phys. Comm. 43, 181 (1987). D. Lemoine, J. Chem. Phys. 101, 3936 (1994).

Chapter 30: One dimensional Root-Finding 296

30 One dimensional Root-Finding

この章では任意の 1 次元関数の解を探索するルーチンについて述べる. このライブラリは様々な反復的解探査と収束テストの低レベルのコンポーネントを実装している. これらをユーザーが組みあわせて, 反復の途中段階すべてを見て望ましい解を得ることができる. メソッドの各クラスで同じ枠組が使われているので, プログラムを再コンパイルせずに実行時に解法を切り替えることができる. 解法の各インスタンスで状態を保持しているので, マルチスレッドプログラムでも使うことができる.

ヘッダファイル ‘gsl_roots.h’に解探査関数のプロトタイプと関連する定義がある.

30.1 Overview

1 次元解探査アルゴリズムは 2 つのクラスに分けることができる. 解の囲い込みと解の洗練である. 解を囲い込むアルゴリズムは収束を保証できる. 囲い込みアルゴリズムは解のあることを知っている区間が出発点となる. 区間の幅は反復的に縮小され, 望む許容誤差以下に区間幅が縮小するまで探査が続く. これは解の位置の誤差評価が厳密に行える利点がある.

解の洗練の技法は, 解の初期推測値を改良しようとする. このアルゴリズムは解に「十分近い」出発点を与えたときのみ収束する. またスピードのため厳密な誤差評価を犠牲にする. 解近傍で関数の挙動を近似することで, 初期値の高次の改良を試みる. 関数の挙動がアルゴリズムのものに近く, 初期値がよく推定できる時には洗練アルゴリズムは素早く収束する.

GSL ではどちらのタイプのアルゴリズムも同じ枠組みに収めた. ユーザーはアルゴリズムの高レベルドライバを実装し, ライブラリは各ステップに必要な個々の関数を実装する. 反復には 3 つのメインフェーズが存在する:

• アルゴリズム T の状態 s を初期化する

• 反復 T を用いて s を更新する

• s の収束を確かめ必要であれば繰り返す

囲い込み法の状態は gsl_root_fsolver構造体に保持される. 更新プロシージャは関数値のみ用いる (導関数値は用いない). 洗練法の状態は gsl_root_fdfsolver構造体に保持される. 更新には関数とその導関数が必要となる (このため fdfという名前になっている).

30.2 Caveats

解探査関数は 1 度に 1 つの解しか探査しない. 探査領域内にいくつかの解があるとき, 始めに見付かった解を返す. しかしどの解が見付かったのかを調べるのは難しい. 多くの場合, 複数の解がある領域で解を探している場合にも何も警告されない.

重解を持つ関数の取り扱いには注意を要する (たとえばf̧(x) = (x − x_{0})^{2} やf̧(x) = (x − x_{0})^{3} のような). 偶数次の重解に対しては囲い込み法は使えない. これらの方法では, 初期範囲は 0 をよぎらなければならない. つまり両端点で片方は正, もう片方は負の値をとらなければならない. 偶数次の重解は 0 をよぎらず, 接するだけである. 囲い込み法は奇数次の重解 (3 次, 5 次…) では機能する. 解の洗練法は一般的に高次の重解を探査することができるが, 収束が悪くなる. この場合収束の加速に Steffenson アルゴリズムを用いるとよい.

探査領域に f が解を持つことは絶対必要というわけではないが, 解の存在を調べるのに数値解探査関数を使うべきではない. もっとよい方法がある. 数値解探査が間違う可能性があるのだから, よく知らない関数を丸投げするのはよくない. 一般的に解を探す前にグラフを描いてチェックするのがよいだろう.

30.3 Initializing the Solver

Function gsl_root_fsolver * gsl root fsolver alloc (const

gsl_root_fsolver_type * T)

この関数はタイプ T のソルバのインスタンスを新しく割りあててポインタを返す. 例えば, 次のコードは bisection ソルバのインスタンスを作成する:

 const gsl_root_fsolver_type * T
   = gsl_root_fsolver_bisection;
 gsl_root_fsolver * s
   = gsl_root_fsolver_alloc (T);

ソルバを作るのに十分なメモリがない場合, 関数はヌルポインタを返し, エラーコード GSL_ ENOMEMによりエラーハンドラが呼びだされる.

Function gsl_root_fdfsolver * gsl root fdfsolver alloc (const

gsl_root_fdfsolver_type * T)

この関数はタイプ T の導関数ベースのソルバのインスタンスを新しく割りあててポインタを返す. 例えば, 次のコードは Newton-Raphson ソルバのインスタンスを作成する:

 const gsl_root_fdfsolver_type * T
   = gsl_root_fdfsolver_newton;
 gsl_root_fdfsolver * s
   = gsl_root_fdfsolver_alloc (T);

ソルバを作るのに十分なメモリがない場合, 関数はヌルポインタを返し, エラーコード GSL_ ENOMEMによりエラーハンドラが呼びだされる.

Function int gsl root fsolver set (gsl_root_fsolver * s, gsl_function * f,
double x lower, double x upper)

この関数は存在するソルバ s が関数 f の解を初期探査範囲 [x lower, x upper] で探査するよう (再) 初期化する.

Function int gsl root fdfsolver set (gsl_root_fdfsolver * s,

gsl_function_fdf * fdf, double root)

この関数は存在するソルバ s が関数とその微分 fdf および初期値 root で探査するように (再) 初期化する.

Function void gsl root fsolver free (gsl_root_fsolver * s) Function void gsl root fdfsolver free (gsl_root_fdfsolver * s)

この関数はソルバ s に割りあてられたメモリを解放する.

Function const char * gsl root fsolver name (const gsl_root_fsolver * s)

Function const char * gsl root fdfsolver name (const

gsl_root_fdfsolver * s)

これらの関数はソルバの名前のポインタを返す. 例えば

 printf("s is a '%s' solver\n",
        gsl_root_fsolver_name (s)) ;
は, s is a 'bisection' solverのような出力を返す.

30.4 Providing the function to solve

解探査をさせるため, 1 変数の連続関数, および場合によりその導関数を与える必要がある. 一般のパラメータを許すため, 関数は次のデータ型で定義されている必要がある:

Data Type gsl function

このデータ型はパラメータをもつ一般の関数を定義する.

double (* function) (double x, void * params)

この関数は, 引数 x, パラメータ params の値f̧(x, params) を返す.

void * params

関数のパラメータのポインタ.

一般の 2 次関数を例にあげる:

f(x) = ax^{2} + bx + c

ここで a = 3, b = 2, c = 1 だとしよう. 次のコードは解探査が可能な関数 gsl_function Fを定義する:

struct my_f_params { double a; double b; double c; } ;
double
my_f (double x, void * p) {
   struct my_f_params * params
     = (struct my_f_params *)p;
   double a = (params->a);
   double b = (params->b);
   double c = (params->c);
   return  (a * x + b) * x + c;
}
gsl_function F;
struct my_f_params params = { 3.0, 2.0, 1.0 };
F.function = &my_f;
F.params = &params;

関数 f(x) は次のマクロにより評価できる：

#define GSL_FN_EVAL(F,x)
    (*((F)->function))(x,(F)->params)

Data Type gsl function fdf

このデータ型でパラメータをもつ一般の関数およびその導関数を定義する.

double (* f ) (double x, void * params)

この関数は引数 x, パラメータ params の値f̧(x, params) を返す.

double (* df ) (double x, void * params)

この関数は引数 x, パラメータ params で, 関数 f の x に関する微分値 f̧′(x, params) を返す.

void (* fdf ) (double x, void * params, double * f, double * df)

この関数は, 引数 x, パラメータ params での関数 f の値をf̧(x, params) に, そして導関数 df の値をf̧′(x, params) で求めることを宣言する. この関数は別々の関数 f(x), f′(x) を最適化する–関数値とその導関数値を同時に求めることにより高速化を図っている.

void * params

関数のパラメータへのポインタ.

次の関数を例に挙げる: f̧(x) = exp(2x)

double
my_f (double x, void * params)
{
   return exp (2 * x);
}
double
my_df (double x, void * params)
{
   return 2 * exp (2 * x);
}
void
my_fdf (double x, void * params,
        double * f, double * df)
{
   double t = exp (2 * x);
   *f = t;
   *df = 2 * t;   /* computing using existing values */
}
gsl_function_fdf FDF;
FDF.f = &my_f;
FDF.df = &my_df;
FDF.fdf = &my_fdf;
FDF.params = 0;

関数 f(x) は次のマクロで評価できる:

#define GSL_FN_FDF_EVAL_F(FDF,x)
     (*((FDF)->f))(x,(FDF)->params)

導関数 f′(x) は次のマクロで評価できる:

#define GSL_FN_FDF_EVAL_DF(FDF,x)
     (*((FDF)->df))(x,(FDF)->params)

そして関数値 y = f(x) および導関数値 dy = f′(x) を同時に求めるには次のマクロを使う:

#define GSL_FN_FDF_EVAL_F_DF(FDF,x,y,dy)
     (*((FDF)->fdf))(x,(FDF)->params,(y),(dy))

このマクロは f(x) を y に, そして f′(x) を dy に格納する—これらはともに double型のポインタである.

30.5 Search Bounds and Guesses

ユーザーは初期範囲か初期値を与えることになる. この章では探査範囲や推定値がどのように機能し, 関数の引数がどのようにそれらを制御するのかを解説する.

推定値は単に x の値である. 解に対して許容範囲内に近づくまで反復される. double型をもつ. 探査範囲は区間の両端点である. 区間幅が要求精度内になるまで反復される. 区間は上限と下限を

示す 2 つの値で指定される. 端点を区間に含むかどうかは区間の使われる状況による.

30.6 Iteration

以下の関数は各アルゴリズムでの反復を制御する. 各関数は相当する型のソルバの状態を更新する反復を 1 回行う. 同じ関数であらゆるソルバに対応するため, コードを変更することなく実行時にメソッドを変更することができる.

Function int gsl root fsolver iterate (gsl_root_fsolver * s) Function int gsl root fdfsolver iterate (gsl_root_fdfsolver * s)

これらの関数はソルバ s の反復を 1 回行う. 予期しない問題が生じた場合はエラーコードが返される.

  GSL_EBADFUNC

関数値や導関数値が Infまたは NaNとなる特異点が発生した.

  GSL_EZERODIV

導関数値が反復点で消滅し, 0 で割る操作が発生しそうになったため強制終了した.

ソルバはその時点での解の最も良い推定値を保持している. 囲い込み法ではその時点で最も良い解区間を保持している. この情報は次の関数で参照することができる:

Function double gsl root fsolver root (const gsl_root_fsolver * s) Function double gsl root fdfsolver root (const gsl_root_fdfsolver * s)

これらの関数はソルバ s のもつ現時点での解の推定値を返す.

Function double gsl root fsolver x lower (const gsl_root_fsolver * s) Function double gsl root fsolver x upper (const gsl_root_fsolver * s)

これらの関数はソルバ s のもつ現時点での解区間を返す.

30.7 Search Stopping Parameters

解探査プロシージャは以下の条件が真になったときに停止する:

• ユーザーの指定した精度で解が求まったとき.

• ユーザーの指定した反復回数の上限値に達したとき.

• エラーが発生したとき.

これらの条件の処理はユーザーに任せられている. 以下の関数はユーザーが標準的な方法により現在の結果の精度を調べるために用意されている.

Function int gsl root test interval (double x lower, double x upper, double

epsrel, double epsabs) この関数は区間 [x lower, x upper] の収束を絶対誤差 epsabs および相対誤差 epsrel で調べ

るものである. 以下の条件が満たされた場合には GSL_SUCCESSを返す.

|a − b| < epsabs + epsrel min(|a|, |b|)

ただし, 区間 x = [a, b] が原点を含んでいないことが必要である. 原点を含んでいる場合には min(|a|, |b|) は (区間の最小値である)0 に置きかえられる. これは原点近くの解での相対誤差の正確な評価に必要な操作である. 区間に対するこの条件により, 区間内の推定値 r が真の解 r^{∗} に対して同様の条件を満たすことを要請する:

|r − r^{∗}| < epsabs + epsrel r^{∗}

ただし区間内に真の解が存在すると仮定している.

Function int gsl root test delta (double x1, double x0, double epsrel,

double epsabs)

この関数は数列…, x0, x1 が絶対誤差 epsabs および相対誤差 epsrel で収束しているかどうかを調べる. 以下の条件が満たされている場合に GSL_SUCCESSを返す:

|x_{1} − x_{0}| < epsabs + epsrel |x_{1}|

満たされていない場合には GSL_CONTINUEを返す.

Function int gsl root test residual (double f, double epsabs)

この関数は絶対誤差 epsabs に対する残差 f を調べる. 以下の条件が満たされたときは GSL_

SUCCESSを返す:

|f| < epsabs

満たされない場合は GSL_CONTINUEを返す. この条件は残差 |f(x)| が十分小さいため, 真の解 x の正確な場所が重要でなくなる場合には適している.

30.8 Root Bracketing Algorithms

この章で説明する解の囲い込み法は解が含まれていることを知っている初期範囲が必要となる–a と b が範囲の端点であれば, f(a) と f(b) は符号が異なる必要がある. つまり与えられた区間で少なくとも 1 回は 0 をよぎる必要がある. 妥当な初期範囲が与えられれば, 関数の振舞いがよければこれらのアルゴリズムは失敗することはない.

囲い込み法では偶数次の重解は見つけることができない. x 軸に接するからである.

Solver gsl root fsolver bisection

2 分法は囲い込み法のうち最も簡単な方法である. 線型収束を示し, ライブラリ中で最も遅く収束するアルゴリズムである. 各段で, 区間は 2 等分され, 中点での関数値が求められる. この値の符号によりどちらの区間に解が含まれるかを決定する. 区間の半分が捨てられ残る半分が新しい推定区間となる. このプロシージャが, 区間幅が十分小さくなるまで繰り返される. どの時点でも, 解の推定値は中点の値である.

2 分法を 4 回繰り返した. a_{n} は区間始点の n 番目の位置, b_{n} は区間終点の n 番目の位置. 各段での中点の位置も示してある.

Solver gsl root fsolver falsepos

false position アルゴリズムは線型補完による解探査法である. 収束は線型だが, 一般に 2 分法より早い. 各段で, 端点 (a, f(a)), (b, f(b)) を結ぶ直線が引かれ, x 軸との交点を「中点」とする. この点での関数値が計算され, その符号でどちらの領域を採用するかを決定し, 片方を捨て残りを次の領域とする. このプロシージャが区間が十分に小さくなるまで繰り返される. 現在の推定値は線型補完により求められる.

Solver gsl root fsolver brent

Brent-Dekker 法(Brent 法と略すことにする) は 2 分法アルゴリズムと内挿を組みあわせたものである. この方法は荒っぽいが早いアルゴリズムである. 各段で Brent 法は内挿曲線で関数を近似する. 初段では 2 端点の直線近似だが, 段が進むにつれ最新の 3 点について逆 2 乗フィットを行う. 内挿曲線と x 軸の交点を解の推定値とする. 推定値が現在の推定区間にある場合は内挿点を用い, 区間を縮小する. 内挿点が求められない時は通常の 2 分法に切りかえる. 最新の推定値は最新の内挿もしくは 2 分法の値である.

30.9 Root Finding Algorithms using Derivatives

この章で説明する解の洗練法は解の位置をあらかじめ推定しておく必要がある. 収束するという絶対的な保証はない–関数はこの手法に適した形をしていなくてはならず, 初期値は真の解に十分近くなければならない. 条件が満たされたときには収束は 2 次である.

これらのアルゴリズムは関数とその導関数が必要となる.

Derivative Solver gsl root fdfsolver newton

Newton 法は標準的な解洗練アルゴリズムである. このアルゴリズムは解の位置の初期推定値から始まる. 各段で, 関数 f の接線がひかれる. 接線と x 軸の交点が新しい推定値となる. 繰り返しは次の数列で定義される:

x_{i+1} = x_{i} − \frac{f(x_{i})}{f′(x_{i})}

Newton 法は単一の解に対しては 2 次で, 複数の解には線型で収束する.

Newton 法での数回の繰り返し. g_{n} は n 番目の推定値である.

Derivative Solver gsl root fdfsolver secant

割線法は Newton 法を単純化したもので, 導関数を毎回求める必要がない. 初段ではアルゴリズムは Newton 法から始め, 導関数を用いる.

x_{1} = x_{0} − \frac{f(x_{0})}{f′(x_{0})}

続く繰り返しでは, 微分値を求める代わりに数値的な推定値で代用する. 前の 2 点の傾きを使うことになる:

x_{i+1} = x_{i} − \frac{f(x_{i})}{f′_{est}} where f′_{est} = \frac{f(x_{i}) − f(x_{i−1})}{x_{i} − x_{i−1}}

導関数が解の近傍でさほど変化しないときには, 割線法はかなりの省力化を図ることができる. 導関数の評価が関数自身の評価よりコストが 0.44 倍以上かかるのであれば, 割線法はニュートン法より早い. 数値微分の計算と同様, 2 点の間隔が小さくなりすぎると桁落ちが生じる. 単一の解であれば, 収束は (1 + \sqrt{5})/2 (約 1.62) 次である. 複数の解では線型で収束する.

Derivative Solver gsl root fdfsolver steffenson

Steffenson 法はこれらルーチンの中で最も早い収束を示す. これは基本的な Newton 法に Aitken の「デルタ 2 乗」加速を組みあわせたものである. Newton 法の各段を x_{i} としたとき, 加速により新しい数列 R_{i} を生成する:

R_{i} = x_{i} − \frac{(x_{i+1} − x_{i})^{2}}{(x_{i+2} − 2x_{i+1} + x_{i})}

これは合理的な条件のもとで元の数列よりも早く収束する. 加速数列を生成するには最低 3 項必要である. 初段では元の Newton 法の結果を返す. 加速項の分母が 0 になる場合も Newton 法の結果を返す. 加速プロシージャを用いているため, 関数の振舞いが悪い場合にはこの方法は不安定になる.

30.10 Examples

どの解探査法であっても, 解くべき方程式を用意する必要がある. この例では, 前に示した一般の 2 次方程式を解くことにしよう. まず関数のパラメータを定義するためにヘッダファイル (‘demo_fn.h’) が必要である.

struct quadratic_params
  {
    double a, b, c;
  };
double quadratic (double x, void *params);
double quadratic_deriv (double x, void *params);
void quadratic_fdf (double x, void *params,
                    double *y, double *dy);

関数の定義は別のファイル (‘demo_fn.c’) で行うことにしよう.

double
quadratic (double x, void *params)
{
  struct quadratic_params *p
    = (struct quadratic_params *) params;
  double a = p->a;
  double b = p->b;
  double c = p->c;
  return (a * x + b) * x + c;
}
double
quadratic_deriv (double x, void *params)
{
  struct quadratic_params *p
    = (struct quadratic_params *) params;
  double a = p->a;
  double b = p->b;
  double c = p->c;

return 2.0 * a * x + b;
}
void
quadratic_fdf (double x, void *params,
             double *y, double *dy)
{
struct quadratic_params *p
  = (struct quadratic_params *) params;
double a = p->a;
double b = p->b;
double c = p->c;
*y = (a * x + b) * x + c;
*dy = 2.0 * a * x + b;
}

最初のプログラムは Brent 法のソルバ gsl_root_fsolver_brentを用い, 次の方程式を解くもので

ある.

x^{2} − 5 = 0

この解は x = \sqrt{5} = 2.236068... である.

#include <stdio.h>
#include <gsl/gsl_errno.h>
#include <gsl/gsl_math.h>
#include <gsl/gsl_roots.h>
#include "demo_fn.h"
#include "demo_fn.c"
int
main (void)
{
int status;
int iter = 0, max_iter = 100;
const gsl_root_fsolver_type *T;
gsl_root_fsolver *s;
double r = 0, r_expected = sqrt (5.0);
double x_lo = 0.0, x_hi = 5.0;
gsl_function F;
struct quadratic_params params = {1.0, 0.0, -5.0};
F.function = &quadratic;
F.params = &params;
T = gsl_root_fsolver_brent;
s = gsl_root_fsolver_alloc (T);
gsl_root_fsolver_set (s, &F, x_lo, x_hi);
printf ("using %s method\n",
        gsl_root_fsolver_name (s));

printf ("%5s [%9s, %9s] %9s %10s %9s\n",
      "iter", "lower", "upper", "root",
      "err", "err(est)");
do
{
  iter++;
  status = gsl_root_fsolver_iterate (s);
  r = gsl_root_fsolver_root (s);
  x_lo = gsl_root_fsolver_x_lower (s);
  x_hi = gsl_root_fsolver_x_upper (s);
  status = gsl_root_test_interval (x_lo, x_hi,
                                   0, 0.001);
  if (status == GSL_SUCCESS)
    printf ("Converged:\n");
  printf ("%5d [%.7f, %.7f] %.7f %+.7f %.7f\n",
          iter, x_lo, x_hi,
          r, r - r_expected,
          x_hi - x_lo);
}
while (status == GSL_CONTINUE && iter < max_iter);
return status;
}

以下に結果を挙げる:

bash$ ./a.out
using brent method
iter [    lower,     upper]      root        err  err(est)
1 [1.0000000, 5.0000000] 1.0000000 -1.2360680 4.0000000
2 [1.0000000, 3.0000000] 3.0000000 +0.7639320 2.0000000
3 [2.0000000, 3.0000000] 2.0000000 -0.2360680 1.0000000
4 [2.2000000, 3.0000000] 2.2000000 -0.0360680 0.8000000
5 [2.2000000, 2.2366300] 2.2366300 +0.0005621 0.0366300
Converged:
6 [2.2360634, 2.2366300] 2.2360634 -0.0000046 0.0005666
Brent 法ではなく 2 分法を使うようにする場合は, gsl_root_fsolver_brentを gsl_root_

fsolver_bisectionに変更すればよい. この場合, 収束が遅くなる:

bash$ ./a.out
using bisection method
iter [    lower,     upper]      root        err  err(est)
1 [0.0000000, 2.5000000] 1.2500000 -0.9860680 2.5000000
2 [1.2500000, 2.5000000] 1.8750000 -0.3610680 1.2500000
3 [1.8750000, 2.5000000] 2.1875000 -0.0485680 0.6250000
4 [2.1875000, 2.5000000] 2.3437500 +0.1076820 0.3125000
5 [2.1875000, 2.3437500] 2.2656250 +0.0295570 0.1562500
6 [2.1875000, 2.2656250] 2.2265625 -0.0095055 0.0781250
7 [2.2265625, 2.2656250] 2.2460938 +0.0100258 0.0390625
8 [2.2265625, 2.2460938] 2.2363281 +0.0002601 0.0195312

  9 [2.2265625, 2.2363281] 2.2314453 -0.0046227 0.0097656
 10 [2.2314453, 2.2363281] 2.2338867 -0.0021813 0.0048828
 11 [2.2338867, 2.2363281] 2.2351074 -0.0009606 0.0024414
Converged:
 12 [2.2351074, 2.2363281] 2.2357178 -0.0003502 0.0012207

次のプログラムは同じ関数を導関数も用いて解を探査するものである.

#include <stdio.h>
#include <gsl/gsl_errno.h>
#include <gsl/gsl_math.h>
#include <gsl/gsl_roots.h>
#include "demo_fn.h"
#include "demo_fn.c"
int
main (void)
{
int status;
int iter = 0, max_iter = 100;
const gsl_root_fdfsolver_type *T;
gsl_root_fdfsolver *s;
double x0, x = 5.0, r_expected = sqrt (5.0);
gsl_function_fdf FDF;
struct quadratic_params params = {1.0, 0.0, -5.0};
FDF.f = &quadratic;
FDF.df = &quadratic_deriv;
FDF.fdf = &quadratic_fdf;
FDF.params = &params;
T = gsl_root_fdfsolver_newton;
s = gsl_root_fdfsolver_alloc (T);
gsl_root_fdfsolver_set (s, &FDF, x);
printf ("using %s method\n",
        gsl_root_fdfsolver_name (s));
printf ("%-5s %10s %10s %10s\n",
        "iter", "root", "err", "err(est)");
do
  {
    iter++;
    status = gsl_root_fdfsolver_iterate (s);
    x0 = x;
    x = gsl_root_fdfsolver_root (s);
    status = gsl_root_test_delta (x, x0, 0, 1e-3);
    if (status == GSL_SUCCESS)
      printf ("Converged:\n");
    printf ("%5d %10.7f %+10.7f %10.7f\n",

          iter, x, x - r_expected, x - x0);
}
while (status == GSL_CONTINUE && iter < max_iter);
return status;
}

Newton 法の結果を以下に挙げる:

bash$ ./a.out
using newton method
iter        root        err   err(est)
1  3.0000000 +0.7639320 -2.0000000
2  2.3333333 +0.0972654 -0.6666667
3  2.2380952 +0.0020273 -0.0952381
Converged:
4  2.2360689 +0.0000009 -0.0020263

誤差は前の段との差を取るより次の段との差をとる方がより正確となる. gsl_root_fdfsolver_

newtonを gsl_root_fdfsolver_secantや gsl_root_fdfsolver_steffensonに変えることで

他の方法に切りかえることができる.

30.11 References and Further Reading

Brent-Dekker アルゴリズムについては以下の 2 つの論文を参照すること.

R. P. Brent, "An algorithm with guaranteed convergence for finding a zero of a function", Computer Journal, 14 (1971) 422-425 J. C. P. Bus and T. J. Dekker, "Two Efficient Algorithms with Guaranteed Convergence for Finding a Zero of a Function", ACM Transactions of Mathematical Software, Vol. 1 No. 4 (1975) 330-345

31 One dimensional Minimization

This chapter describes routines for finding minima of arbitrary one-dimensional functions. The library provides low level components for a variety of iterative minimizers and convergence tests. These can be combined by the user to achieve the desired solution, with full access to the intermediate steps of the algorithms. Each class of methods uses the same framework, so that you can switch between minimizers at runtime without needing to recompile your program. Each instance of a minimizer keeps track of its own state, allowing the minimizers to be used in multi-threaded programs.

The header file ‘gsl_min.h’ contains prototypes for the minimization functions and related declarations. To use the minimization algorithms to find the maximum of a function simply invert its sign.

31.1 Overview

The minimization algorithms begin with a bounded region known to contain a minimum. The region is described by an lower bound a and an upper bound b, with an estimate of the minimum x.

The value of the function at x must be less than the value of the function at the ends of the interval,

f(a) > f(x) < f(b)

This condition guarantees that a minimum is contained somewhere within the interval. On each iteration a new point x′ is selected using one of the available algorithms. If the new point is a better estimate of the minimum, f(x′) < f(x), then the current estimate of the minimum x is updated. The new point also allows the size of the bounded interval to be reduced, by choosing the most compact set of points which satisfies the constraint f(a) > f(x) < f(b). The interval is reduced until it encloses the true minimum to a desired tolerance. This provides a best estimate of the location of the minimum and a rigorous error estimate.

Several bracketing algorithms are available within a single framework. The user provides a high-level driver for the algorithm, and the library provides the individual functions necessary for each of the steps. There are three main phases of the iteration. The steps are,

• initialize minimizer state, s, for algorithm T

• update s using the iteration T

• test s for convergence, and repeat iteration if necessary

The state for the minimizers is held in a gsl_min_fminimizer struct. The updating procedure uses only function evaluations (not derivatives).

31.2 Caveats

Note that minimization functions can only search for one minimum at a time. When there are several minima in the search area, the first minimum to be found will be returned; however it is difficult to predict which of the minima this will be. In most cases, no error will be reported if you try to find a minimum in an area where there is more than one.

With all minimization algorithms it can be difficult to determine the location of the minimum to full numerical precision. The behavior of the function in the region of the minimum x^{∗} can be approximated by a Taylor expansion,

y = f(x^{∗}) + \frac{1}{2}f′′(x^{∗})(x − x^{∗})^{2}

and the second term of this expansion can be lost when added to the first term at finite precision. This magnifies the error in locating x^{∗}, making it proportional to \sqrt{ϵ} (where ϵ is the relative accuracy of the floating point numbers). For functions with higher order minima, such as x^{4}, the magnification of the error is correspondingly worse. The best that can be achieved is to converge to the limit of numerical accuracy in the function values, rather than the location of the minimum itself.

31.3 Initializing the Minimizer

Function gsl_min_fminimizer * gsl min fminimizer alloc (const

         gsl_min_fminimizer_type * T)

This function returns a pointer to a a newly allocated instance of a minimizer of type T. For example, the following code creates an instance of a golden section minimizer,

          const gsl_min_fminimizer_type * T
            = gsl_min_fminimizer_goldensection;
          gsl_min_fminimizer * s
            = gsl_min_fminimizer_alloc (T);

If there is insufficient memory to create the minimizer then the function returns a null pointer and the error handler is invoked with an error code of GSL_ENOMEM.

Function int gsl min fminimizer set (gsl_min_fminimizer * s,

    gsl_function * f, double minimum, double x lower, double x upper)

This function sets, or resets, an existing minimizer s to use the function f and the initial search interval [x lower, x upper], with a guess for the location of the minimum minimum. If the interval given does not contain a minimum, then the function returns an error

code of GSL_FAILURE.

Function int gsl min fminimizer set with values (gsl_min_fminimizer *

    s, gsl_function * f, double minimum, double f minimum, double x lower,
    double f lower, double x upper, double f upper)

This function is equivalent to gsl_min_fminimizer_set but uses the values f minimum, f lower and f upper instead of computing f(minimum), f(x_lower) and

f(x_upper).

Function void gsl min fminimizer free (gsl_min_fminimizer * s)

This function frees all the memory associated with the minimizer s.

Function const char * gsl min fminimizer name (const

    gsl_min_fminimizer * s)

This function returns a pointer to the name of the minimizer. For example,

     printf("s is a '%s' minimizer\n",
            gsl_min_fminimizer_name (s));

would print something like s is a 'brent' minimizer.

31.4 Providing the function to minimize

You must provide a continuous function of one variable for the minimizers to operate on. In order to allow for general parameters the functions are defined by a gsl_function data type (see Section 30.4 [Providing the function to solve], page 298).

31.5 Iteration

The following functions drive the iteration of each algorithm. Each function performs one iteration to update the state of any minimizer of the corresponding type. The same functions work for all minimizers so that different methods can be substituted at runtime without modifications to the code.

Function int gsl min fminimizer iterate (gsl_min_fminimizer * s)

This function performs a single iteration of the minimizer s. If the iteration encounters an unexpected problem then an error code will be returned,

GSL_EBADFUNC

the iteration encountered a singular point where the function evaluated

          to Inf or NaN.
GSL_FAILURE

the algorithm could not improve the current best approximation or bounding interval.

The minimizer maintains a current best estimate of the minimum at all times, and the current interval bounding the minimum. This information can be accessed with the following auxiliary functions,

Function double gsl min fminimizer minimum (const

    gsl_min_fminimizer * s)

This function returns the current estimate of the minimum for the minimizer s.

Function double gsl interval gsl_min_fminimizer_x_upper (const
    gsl_min_fminimizer * s)
Function double gsl interval gsl_min_fminimizer_x_lower (const
    gsl_min_fminimizer * s)

These functions return the current upper and lower bound of the interval for the minimizer s.

31.6 Stopping Parameters

A minimization procedure should stop when one of the following conditions is true:

• A minimum has been found to within the user-specified precision.

• A user-specified maximum number of iterations has been reached.

• An error has occurred.

The handling of these conditions is under user control. The function below allows the user to test the precision of the current result.

Function int gsl min test interval (double x lower, double x upper, double

epsrel, double epsabs) This function tests for the convergence of the interval [x lower, x upper] with absolute error epsabs and relative error epsrel. The test returns GSL_SUCCESS if the following condition is achieved,

|a − b| < epsabs + epsrel min(|a|, |b|)

when the interval x = [a, b] does not include the origin. If the interval includes the origin then min(|a|, |b|) is replaced by zero (which is the minimum value of |x| over the interval). This ensures that the relative error is accurately estimated for minima close to the origin. This condition on the interval also implies that any estimate of the minimum x_{m} in the interval satisfies the same condition with respect to the true minimum x_{m}^{∗},

|x_{m} − x_{m}^{∗}| < epsabs + epsrel x_{m}^{∗}

assuming that the true minimum x_{m}^{∗} is contained within the interval.

31.7 Minimization Algorithms

The minimization algorithms described in this section require an initial interval which is guaranteed to contain a minimum — if a and b are the endpoints of the interval and x is an estimate of the minimum then f(a) > f(x) < f(b). This ensures that the function has at least one minimum somewhere in the interval. If a valid initial interval is used then these algorithm cannot fail, provided the function is well-behaved.

Minimizer gsl min fminimizer goldensection

The golden section algorithm is the simplest method of bracketing the minimum of a function. It is the slowest algorithm provided by the library, with linear convergence. On each iteration, the algorithm first compares the subintervals from the endpoints to the current minimum. The larger subinterval is divided in a golden section (using the famous ratio (3 − \sqrt{5})/2 = 0.3189660…) and the value of the function at this new point is calculated. The new value is used with the constraint f(a′) > f(x′) < f(b′) to a select new interval containing the minimum, by discarding the least useful point. This procedure can be continued indefinitely until the interval is sufficiently small. Choosing the golden section as the bisection ratio can be shown to provide the fastest convergence for this type of algorithm.

Minimizer gsl min fminimizer brent

The Brent minimization algorithm combines a parabolic interpolation with the golden section algorithm. This produces a fast algorithm which is still robust. The outline of the algorithm can be summarized as follows: on each iteration Brent’s method approximates the function using an interpolating parabola through three existing points. The minimum of the parabola is taken as a guess for the minimum. If it lies within the bounds of the current interval then the interpolating point is accepted, and used to generate a smaller interval. If the interpolating point is not accepted then the algorithm falls back to an ordinary golden section step. The full details of Brent’s method include some additional checks to improve convergence.

31.8 Examples

The following program uses the Brent algorithm to find the minimum of the function f(x) = cos(x) + 1, which occurs at x = π. The starting interval is (0, 6), with an initial guess for the minimum of 2.

#include <stdio.h>
#include <gsl/gsl_errno.h>
#include <gsl/gsl_math.h>
#include <gsl/gsl_min.h>
double fn1 (double x, void * params)
{
  return cos(x) + 1.0;
}
int
main (void)
{

int status;
int iter = 0, max_iter = 100;
const gsl_min_fminimizer_type *T;
gsl_min_fminimizer *s;
double m = 2.0, m_expected = M_PI;
double a = 0.0, b = 6.0;
gsl_function F;
F.function = &fn1;
F.params = 0;
T = gsl_min_fminimizer_brent;
s = gsl_min_fminimizer_alloc (T);
gsl_min_fminimizer_set (s, &F, m, a, b);
printf ("using %s method\n",
        gsl_min_fminimizer_name (s));
printf ("%5s [%9s, %9s] %9s %10s %9s\n",
        "iter", "lower", "upper", "min",
        "err", "err(est)");
printf ("%5d [%.7f, %.7f] %.7f %+.7f %.7f\n",
        iter, a, b,
        m, m - m_expected, b - a);
do
  {
    iter++;
    status = gsl_min_fminimizer_iterate (s);
    m = gsl_min_fminimizer_minimum (s);
    a = gsl_min_fminimizer_x_lower (s);
    b = gsl_min_fminimizer_x_upper (s);
    status
      = gsl_min_test_interval (a, b, 0.001, 0.0);
    if (status == GSL_SUCCESS)
      printf ("Converged:\n");
    printf ("%5d [%.7f, %.7f] "
            "%.7f %.7f %+.7f %.7f\n",
            iter, a, b,
            m, m_expected, m - m_expected, b - a);
  }
while (status == GSL_CONTINUE && iter < max_iter);
return status;
}

Here are the results of the minimization procedure.

bash$ ./a.out
0 [0.0000000, 6.0000000] 2.0000000 -1.1415927 6.0000000
1 [2.0000000, 6.0000000] 3.2758640 +0.1342713 4.0000000
2 [2.0000000, 3.2831929] 3.2758640 +0.1342713 1.2831929
3 [2.8689068, 3.2831929] 3.2758640 +0.1342713 0.4142862
4 [2.8689068, 3.2831929] 3.2758640 +0.1342713 0.4142862
5 [2.8689068, 3.2758640] 3.1460585 +0.0044658 0.4069572
6 [3.1346075, 3.2758640] 3.1460585 +0.0044658 0.1412565
7 [3.1346075, 3.1874620] 3.1460585 +0.0044658 0.0528545
8 [3.1346075, 3.1460585] 3.1460585 +0.0044658 0.0114510
9 [3.1346075, 3.1460585] 3.1424060 +0.0008133 0.0114510
10 [3.1346075, 3.1424060] 3.1415885 -0.0000041 0.0077985
Converged:
11 [3.1415885, 3.1424060] 3.1415927 -0.0000000 0.0008175

31.9 References and Further Reading

Further information on Brent’s algorithm is available in the following book, Richard Brent, Algorithms for minimization without derivatives, Prentice-Hall (1973)

32 Multidimensional Root-Finding

This chapter describes functions for multidimensional root-finding (solving nonlinear systems with n equations in n unknowns). The library provides low level components for a variety of iterative solvers and convergence tests. These can be combined by the user to achieve the desired solution, with full access to the intermediate steps of the iteration. Each class of methods uses the same framework, so that you can switch between solvers at runtime without needing to recompile your program. Each instance of a solver keeps track of its own state, allowing the solvers to be used in multi-threaded programs. The solvers are based on the original Fortran library minpack.

The header file ‘gsl_multiroots.h’ contains prototypes for the multidimensional root finding functions and related declarations.

32.1 Overview

The problem of multidimensional root finding requires the simultaneous solution of n equations, f_{i}, in n variables, x_{i},

f_{i}(x_{1}, …, x_{n}) = 0 for i = 1 … n.

In general there are no bracketing methods available for n dimensional systems, and no way of knowing whether any solutions exist. All algorithms proceed from an initial guess using a variant of the Newton iteration,

x → x′ = x − J^{−1}f(x)

where x, f are vector quantities and J is the Jacobian matrix J_{ij} = ∂f_{i}/∂x_{j}. Additional strategies can be used to enlarge the region of convergence. These include requiring a decrease in the norm |f| on each step proposed by Newton’s method, or taking steepestdescent steps in the direction of the negative gradient of |f|.

Several root-finding algorithms are available within a single framework. The user provides a high-level driver for the algorithms, and the library provides the individual functions necessary for each of the steps. There are three main phases of the iteration. The steps are,

• initialize solver state, s, for algorithm T

• update s using the iteration T

• test s for convergence, and repeat iteration if necessary

The evaluation of the Jacobian matrix can be problematic, either because programming the derivatives is intractable or because computation of the n^{2} terms of the matrix becomes too expensive. For these reasons the algorithms provided by the library are divided into two classes according to whether the derivatives are available or not.

The state for solvers with an analytic Jacobian matrix is held in a gsl_multiroot_ fdfsolver struct. The updating procedure requires both the function and its derivatives to be supplied by the user.

The state for solvers which do not use an analytic Jacobian matrix is held in a gsl_ multiroot_fsolver struct. The updating procedure uses only function evaluations (not derivatives). The algorithms estimate the matrix J or J^{−1} by approximate methods.

32.2 Initializing the Solver

The following functions initialize a multidimensional solver, either with or without derivatives. The solver itself depends only on the dimension of the problem and the algorithm and can be reused for different problems.

Function gsl_multiroot_fsolver * gsl multiroot fsolver alloc (const

    gsl_multiroot_fsolver_type * T, size_t n)

This function returns a pointer to a a newly allocated instance of a solver of type T for a system of n dimensions. For example, the following code creates an instance of a hybrid solver, to solve a 3-dimensional system of equations.

     const gsl_multiroot_fsolver_type * T
         = gsl_multiroot_fsolver_hybrid;
     gsl_multiroot_fsolver * s
         = gsl_multiroot_fsolver_alloc (T, 3);

If there is insufficient memory to create the solver then the function returns a null pointer and the error handler is invoked with an error code of GSL_ENOMEM.

Function gsl_multiroot_fdfsolver * gsl multiroot fdfsolver alloc

    (const gsl_multiroot_fdfsolver_type * T, size_t n)

This function returns a pointer to a a newly allocated instance of a derivative solver of type T for a system of n dimensions. For example, the following code creates an instance of a Newton-Raphson solver, for a 2-dimensional system of equations.

     const gsl_multiroot_fdfsolver_type * T
         = gsl_multiroot_fdfsolver_newton;
     gsl_multiroot_fdfsolver * s =
         gsl_multiroot_fdfsolver_alloc (T, 2);

If there is insufficient memory to create the solver then the function returns a null pointer and the error handler is invoked with an error code of GSL_ENOMEM.

Function int gsl multiroot fsolver set (gsl_multiroot_fsolver * s,

    gsl_multiroot_function * f, gsl_vector * x)

This function sets, or resets, an existing solver s to use the function f and the initial guess x.

Function int gsl multiroot fdfsolver set (gsl_multiroot_fdfsolver * s,

    gsl_function_fdf * fdf, gsl_vector * x)

This function sets, or resets, an existing solver s to use the function and derivative fdf and the initial guess x.

Function void gsl multiroot fsolver free (gsl_multiroot_fsolver * s) Function void gsl multiroot fdfsolver free (gsl_multiroot_fdfsolver * s)

These functions free all the memory associated with the solver s.

Function const char * gsl multiroot fsolver name (const

    gsl_multiroot_fsolver * s)

Function const char * gsl multiroot fdfsolver name (const

    gsl_multiroot_fdfsolver * s)

These functions return a pointer to the name of the solver. For example,

     printf("s is a '%s' solver\n",
            gsl_multiroot_fdfsolver_name (s));

would print something like s is a 'newton' solver.

32.3 Providing the function to solve

You must provide n functions of n variables for the root finders to operate on. In order to allow for general parameters the functions are defined by the following data types:

Data Type gsl multiroot function

This data type defines a general system of functions with parameters.

int (* f) (const gsl_vector * x, void * params, gsl_vector * f )

this function should store the vector result f(x, params) in f for argument x and parameters params, returning an appropriate error code if the function cannot be computed.

size_t n the dimension of the system, i.e. the number of components of the vectors x and f.

void * params a pointer to the parameters of the function.

Here is an example using Powell’s test function,

f_{1}(x) = Ax_{0}x_{1} − 1, f_{2}(x) = exp(−x_{0}) + exp(−x_{1}) − (1 + 1/A)

with A = 10^{4}. The following code defines a gsl_multiroot_function system F which you could pass to a solver:

struct powell_params { double A; };
int
powell (gsl_vector * x, void * p, gsl_vector * f) {
   struct powell_params * params
     = *(struct powell_params *)p;
   double A = (params->A);
   double x0 = gsl_vector_get(x,0);
   double x1 = gsl_vector_get(x,1);
   gsl_vector_set (f, 0, A * x0 * x1 - 1)
   gsl_vector_set (f, 1, (exp(-x0) + exp(-x1)
                          - (1.0 + 1.0/A)))
   return GSL_SUCCESS
}
gsl_multiroot_function F;
struct powell_params params = { 10000.0 };

F.function = &powell;
F.n = 2;
F.params = &params;

Data Type gsl multiroot function fdf

This data type defines a general system of functions with parameters and the corresponding Jacobian matrix of derivatives,

int (* f) (const gsl_vector * x, void * params, gsl_vector * f )

this function should store the vector result f(x, params) in f for argument x and parameters params, returning an appropriate error code if the function cannot be computed.

int (* df) (const gsl_vector * x, void * params, gsl_matrix * J)

this function should store the n-by-n matrix result J_{ij} = ∂f_{i}(x, params)/∂x_{j} in J for argument x and parameters params, returning an appropriate error code if the function cannot be computed.

int (* fdf) (const gsl_vector * x, void * params, gsl_vector * f,
gsl_matrix * J)

This function should set the values of the f and J as above, for arguments x and parameters params. This function provides an optimization of the separate functions for f(x) and J(x) – it is always faster to compute the function and its derivative at the same time.

size_t n the dimension of the system, i.e. the number of components of the vectors x and f.

void * params a pointer to the parameters of the function.

The example of Powell’s test function defined above can be extended to include analytic derivatives using the following code,

int
powell_df (gsl_vector * x, void * p, gsl_matrix * J)
{
   struct powell_params * params
     = *(struct powell_params *)p;
   double A = (params->A);
   double x0 = gsl_vector_get(x,0);
   double x1 = gsl_vector_get(x,1);
   gsl_vector_set (J, 0, 0, A * x1)
   gsl_vector_set (J, 0, 1, A * x0)
   gsl_vector_set (J, 1, 0, -exp(-x0))
   gsl_vector_set (J, 1, 1, -exp(-x1))
   return GSL_SUCCESS
}
int
powell_fdf (gsl_vector * x, void * p,
            gsl_matrix * f, gsl_matrix * J) {
   struct powell_params * params

  = *(struct powell_params *)p;
double A = (params->A);
double x0 = gsl_vector_get(x,0);
double x1 = gsl_vector_get(x,1);
double u0 = exp(-x0);
double u1 = exp(-x1);
gsl_vector_set (f, 0, A * x0 * x1 - 1)
gsl_vector_set (f, 1, u0 + u1 - (1 + 1/A))
gsl_vector_set (J, 0, 0, A * x1)
gsl_vector_set (J, 0, 1, A * x0)
gsl_vector_set (J, 1, 0, -u0)
gsl_vector_set (J, 1, 1, -u1)
return GSL_SUCCESS
}
gsl_multiroot_function_fdf FDF;
FDF.f = &powell_f;
FDF.df = &powell_df;
FDF.fdf = &powell_fdf;
FDF.n = 2;
FDF.params = 0;

Note that the function powell_fdf is able to reuse existing terms from the function when calculating the Jacobian, thus saving time.

32.4 Iteration

The following functions drive the iteration of each algorithm. Each function performs one iteration to update the state of any solver of the corresponding type. The same functions work for all solvers so that different methods can be substituted at runtime without modifications to the code.

Function int gsl multiroot fsolver iterate (gsl_multiroot_fsolver * s) Function int gsl multiroot fdfsolver iterate (gsl_multiroot_fdfsolver *

s) These functions perform a single iteration of the solver s. If the iteration encounters an unexpected problem then an error code will be returned,

GSL_EBADFUNC

the iteration encountered a singular point where the function or its derivative evaluated to Inf or NaN.

GSL_ENOPROG

the iteration is not making any progress, preventing the algorithm from continuing.

The solver maintains a current best estimate of the root at all times. This information can be accessed with the following auxiliary functions,

Function gsl_vector * gsl multiroot fsolver root (const

    gsl_multiroot_fsolver * s)

Function gsl_vector * gsl multiroot fdfsolver root (const

    gsl_multiroot_fdfsolver * s)

These functions return the current estimate of the root for the solver s.

32.5 Search Stopping Parameters

A root finding procedure should stop when one of the following conditions is true:

• A multidimensional root has been found to within the user-specified precision.

• A user-specified maximum number of iterations has been reached.

• An error has occurred.

The handling of these conditions is under user control. The functions below allow the user to test the precision of the current result in several standard ways.

Function int gsl multiroot test delta (const gsl_vector * dx, const

    gsl_vector * x, double epsabs, double epsrel)

This function tests for the convergence of the sequence by comparing the last step dx with the absolute error epsabs and relative error epsrel to the current position x. The test returns GSL_SUCCESS if the following condition is achieved,

|dx_{i}| < epsabs + epsrel |x_{i}|

for each component of x and returns GSL_CONTINUE otherwise.

Function int gsl multiroot test residual (const gsl_vector * f, double

epsabs) This function tests the residual value f against the absolute error bound epsabs. The test returns GSL_SUCCESS if the following condition is achieved,

∑_{i} |f_{i}| < epsabs

and returns GSL_CONTINUE otherwise. This criterion is suitable for situations where the the precise location of the root, x, is unimportant provided a value can be found where the residual is small enough.

32.6 Algorithms using Derivatives

The root finding algorithms described in this section make use of both the function and its derivative. They require an initial guess for the location of the root, but there is no absolute guarantee of convergence – the function must be suitable for this technique and the initial guess must be sufficiently close to the root for it to work. When the conditions are satisfied then convergence is quadratic.

Derivative Solver gsl multiroot fdfsolver hybridsj

This is a modified version of Powell’s Hybrid method as implemented in the hybrj algorithm in minpack. Minpack was written by Jorge J. Moré, Burton S. Garbow and Kenneth E. Hillstrom. The Hybrid algorithm retains the fast convergence of Newton’s method but will also reduce the residual when Newton’s method is unreliable. The algorithm uses a generalized trust region to keep each step under control. In order to be accepted a proposed new position x′ must satisfy the condition |D(x′ − x)| < δ, where D is a diagonal scaling matrix and δ is the size of the trust region. The components of D are computed internally, using the column norms of the Jacobian to estimate the sensitivity of the residual to each component of x. This improves the behavior of the algorithm for badly scaled functions. On each iteration the algorithm first determines the standard Newton step by solving the system Jdx = −f. If this step falls inside the trust region it is used as a trial step in the next stage. If not, the algorithm uses the linear combination of the Newton and gradient directions which is predicted to minimize the norm of the function while staying inside the trust region.

dx = −αJ^{−1}f(x) − β∇|f(x)|^{2}

This combination of Newton and gradient directions is referred to as a dogleg step. The proposed step is now tested by evaluating the function at the resulting point, x′. If the step reduces the norm of the function sufficiently then it is accepted and size of the trust region is increased. If the proposed step fails to improve the solution then the size of the trust region is decreased and another trial step is computed. The speed of the algorithm is increased by computing the changes to the Jacobian approximately, using a rank-1 update. If two successive attempts fail to reduce the residual then the full Jacobian is recomputed. The algorithm also monitors the progress of the solution and returns an error if several steps fail to make any improvement,

GSL_ENOPROG

the iteration is not making any progress, preventing the algorithm from continuing.

GSL_ENOPROGJ

re-evaluations of the Jacobian indicate that the iteration is not making any progress, preventing the algorithm from continuing.

Derivative Solver gsl multiroot fdfsolver hybridj

This algorithm is an unscaled version of hybridsj. The steps are controlled by a spherical trust region |x′ − x| < δ, instead of a generalized region. This can be useful if the generalized region estimated by hybridsj is inappropriate.

Derivative Solver gsl multiroot fdfsolver newton

Newton’s Method is the standard root-polishing algorithm. The algorithm begins with an initial guess for the location of the solution. On each iteration a linear approximation to the function F is used to estimate the step which will zero all the components of the residual. The iteration is defined by the following sequence,

x → x′ = x − J^{−1}f(x)

where the Jacobian matrix J is computed from the derivative functions provided by f. The step dx is obtained by solving the linear system,

J dx = −f(x)

using LU decomposition.

Derivative Solver gsl multiroot fdfsolver gnewton

This is a modified version of Newton’s method which attempts to improve global convergence by requiring every step to reduce the Euclidean norm of the residual, |f(x)|. If the Newton step leads to an increase in the norm then a reduced step of relative size,

t = (\sqrt{1 + 6r} − 1)/(3r)

is proposed, with r being the ratio of norms |f(x′)|^{2}/|f(x)|^{2}. This procedure is repeated until a suitable step size is found.

32.7 Algorithms without Derivatives

The algorithms described in this section do not require any derivative information to be supplied by the user. Any derivatives needed are approximated from by finite difference.

Solver gsl multiroot fsolver hybrids

This is a version of the Hybrid algorithm which replaces calls to the Jacobian function by its finite difference approximation. The finite difference approximation is computed

using gsl_multiroots_fdjac with a relative step size of GSL_SQRT_DBL_EPSILON.

Solver gsl multiroot fsolver hybrid

This is a finite difference version of the Hybrid algorithm without internal scaling.

Solver gsl multiroot fsolver dnewton

The discrete Newton algorithm is the simplest method of solving a multidimensional system. It uses the Newton iteration

x → x − J^{−1}f(x)

where the Jacobian matrix J is approximated by taking finite differences of the function f. The approximation scheme used by this implementation is,

J_{ij} = (f_{i}(x + δ_{j}) − f_{i}(x))/δ_{j}

where δ_{j} is a step of size \sqrt{ϵ}|x_{j}| with ϵ being the machine precision (ϵ ≈ 2.22×10^{−16}). The order of convergence of Newton’s algorithm is quadratic, but the finite differences require n^{2} function evaluations on each iteration. The algorithm may become unstable if the finite differences are not a good approximation to the true derivatives.

Solver gsl multiroot fsolver broyden

The Broyden algorithm is a version of the discrete Newton algorithm which attempts to avoids the expensive update of the Jacobian matrix on each iteration. The changes to the Jacobian are also approximated, using a rank-1 update,

J^{−1} → J^{−1} − (J^{−1}df − dx)dx^{T} J^{−1}/dx^{T} J^{−1}df

where the vectors dx and df are the changes in x and f. On the first iteration the inverse Jacobian is estimated using finite differences, as in the discrete Newton algorithm. This approximation gives a fast update but is unreliable if the changes are not small, and the estimate of the inverse Jacobian becomes worse as time passes. The algorithm has a tendency to become unstable unless it starts close to the root. The Jacobian is refreshed if this instability is detected (consult the source for details). This algorithm is not recommended and is included only for demonstration purposes.

32.8 Examples

The multidimensional solvers are used in a similar way to the one-dimensional root finding algorithms. This first example demonstrates the hybrids scaled-hybrid algorithm, which does not require derivatives. The program solves the Rosenbrock system of equations,

f_{1}(x, y) = a(1 − x), f_{2}(x, y) = b(y − x^{2})

with a = 1, b = 10. The solution of this system lies at (x, y) = (1, 1) in a narrow valley.

The first stage of the program is to define the system of equations,

#include <stdlib.h>
#include <stdio.h>
#include <gsl/gsl_vector.h>
#include <gsl/gsl_multiroots.h>
struct rparams
  {
    double a;
    double b;
  };
int
rosenbrock_f (const gsl_vector * x, void *params,
              gsl_vector * f)
{
  double a = ((struct rparams *) params)->a;
  double b = ((struct rparams *) params)->b;
  double x0 = gsl_vector_get (x, 0);
  double x1 = gsl_vector_get (x, 1);
  double y0 = a * (1 - x0);
  double y1 = b * (x1 - x0 * x0);
  gsl_vector_set (f, 0, y0);

gsl_vector_set (f, 1, y1);
return GSL_SUCCESS;
}

The main program begins by creating the function object f, with the arguments (x,y) and parameters (a,b). The solver s is initialized to use this function, with the hybrids method.

int
main (void)
{
const gsl_multiroot_fsolver_type *T;
gsl_multiroot_fsolver *s;
int status;
size_t i, iter = 0;
const size_t n = 2;
struct rparams p = {1.0, 10.0};
gsl_multiroot_function f = {&rosenbrock_f, n, &p};
double x_init[2] = {-10.0, -5.0};
gsl_vector *x = gsl_vector_alloc (n);
gsl_vector_set (x, 0, x_init[0]);
gsl_vector_set (x, 1, x_init[1]);
T = gsl_multiroot_fsolver_hybrids;
s = gsl_multiroot_fsolver_alloc (T, 2);
gsl_multiroot_fsolver_set (s, &f, x);
print_state (iter, s);
do
  {
    iter++;
    status = gsl_multiroot_fsolver_iterate (s);
    print_state (iter, s);
    if (status)   /* check if solver is stuck */
      break;
    status =
      gsl_multiroot_test_residual (s->f, 1e-7);
  }
while (status == GSL_CONTINUE && iter < 1000);
printf ("status = %s\n", gsl_strerror (status));
gsl_multiroot_fsolver_free (s);
gsl_vector_free (x);

       return 0;
     }

Note that it is important to check the return status of each solver step, in case the algorithm becomes stuck. If an error condition is detected, indicating that the algorithm cannot proceed, then the error can be reported to the user, a new starting point chosen or a different algorithm used.

The intermediate state of the solution is displayed by the following function. The solver state contains the vector s->x which is the current position, and the vector s->f with corresponding function values.

     int
     print_state (size_t iter, gsl_multiroot_fsolver * s)
     {
       printf ("iter = %3u x = % .3f % .3f "
               "f(x) = % .3e % .3e\n",
               iter,
               gsl_vector_get (s->x, 0),
               gsl_vector_get (s->x, 1),
               gsl_vector_get (s->f, 0),
               gsl_vector_get (s->f, 1));
     }

Here are the results of running the program. The algorithm is started at (−10, −5) far from the solution. Since the solution is hidden in a narrow valley the earliest steps follow the gradient of the function downhill, in an attempt to reduce the large value of the residual. Once the root has been approximately located, on iteration 8, the Newton behavior takes over and convergence is very rapid.

     iter =  0 x = -10.000  -5.000  f(x) = 1.100e+01 -1.050e+03
     iter =  1 x = -10.000  -5.000  f(x) = 1.100e+01 -1.050e+03
     iter =  2 x =  -3.976  24.827  f(x) = 4.976e+00  9.020e+01
     iter =  3 x =  -3.976  24.827  f(x) = 4.976e+00  9.020e+01
     iter =  4 x =  -3.976  24.827  f(x) = 4.976e+00  9.020e+01
     iter =  5 x =  -1.274  -5.680  f(x) = 2.274e+00 -7.302e+01
     iter =  6 x =  -1.274  -5.680  f(x) = 2.274e+00 -7.302e+01
     iter =  7 x =   0.249   0.298  f(x) = 7.511e-01  2.359e+00
     iter =  8 x =   0.249   0.298  f(x) = 7.511e-01  2.359e+00
     iter =  9 x =   1.000   0.878  f(x) = 1.268e-10 -1.218e+00
     iter = 10 x =   1.000   0.989  f(x) = 1.124e-11 -1.080e-01
     iter = 11 x =   1.000   1.000  f(x) = 0.000e+00  0.000e+00
     status = success

Note that the algorithm does not update the location on every iteration. Some iterations are used to adjust the trust-region parameter, after trying a step which was found to be divergent, or to recompute the Jacobian, when poor convergence behavior is detected.

The next example program adds derivative information, in order to accelerate the solution. There are two derivative functions rosenbrock_df and rosenbrock_fdf. The latter computes both the function and its derivative simultaneously. This allows the optimization of any common terms. For simplicity we substitute calls to the separate f and df functions at this point in the code below.

int
rosenbrock_df (const gsl_vector * x, void *params,
             gsl_matrix * J)
{
double a = ((struct rparams *) params)->a;
double b = ((struct rparams *) params)->b;
double x0 = gsl_vector_get (x, 0);
double df00 = -a;
double df01 = 0;
double df10 = -2 * b  * x0;
double df11 = b;
gsl_matrix_set (J, 0, 0, df00);
gsl_matrix_set (J, 0, 1, df01);
gsl_matrix_set (J, 1, 0, df10);
gsl_matrix_set (J, 1, 1, df11);
return GSL_SUCCESS;
}
int
rosenbrock_fdf (const gsl_vector * x, void *params,
              gsl_vector * f, gsl_matrix * J)
{
rosenbrock_f (x, params, f);
rosenbrock_df (x, params, J);
return GSL_SUCCESS;
}

The main program now makes calls to the corresponding fdfsolver versions of the functions,

int
main (void)
{
const gsl_multiroot_fdfsolver_type *T;
gsl_multiroot_fdfsolver *s;
int status;
size_t i, iter = 0;
const size_t n = 2;
struct rparams p = {1.0, 10.0};
gsl_multiroot_function_fdf f = {&rosenbrock_f,
                                &rosenbrock_df,
                                &rosenbrock_fdf,
                                n, &p};
double x_init[2] = {-10.0, -5.0};
gsl_vector *x = gsl_vector_alloc (n);

gsl_vector_set (x, 0, x_init[0]);
gsl_vector_set (x, 1, x_init[1]);
T = gsl_multiroot_fdfsolver_gnewton;
s = gsl_multiroot_fdfsolver_alloc (T, n);
gsl_multiroot_fdfsolver_set (s, &f, x);
print_state (iter, s);
do
  {
    iter++;
    status = gsl_multiroot_fdfsolver_iterate (s);
    print_state (iter, s);
    if (status)
      break;
    status = gsl_multiroot_test_residual (s->f, 1e-7);
  }
while (status == GSL_CONTINUE && iter < 1000);
printf ("status = %s\n", gsl_strerror (status));
gsl_multiroot_fdfsolver_free (s);
gsl_vector_free (x);
return 0;
}

The addition of derivative information to the hybrids solver does not make any significant difference to its behavior, since it able to approximate the Jacobian numerically with sufficient accuracy. To illustrate the behavior of a different derivative solver we switch to gnewton. This is a traditional newton solver with the constraint that it scales back its step if the full step would lead "uphill". Here is the output for the gnewton algorithm,

iter = 0 x = -10.000  -5.000 f(x) =  1.100e+01 -1.050e+03
iter = 1 x =  -4.231 -65.317 f(x) =  5.231e+00 -8.321e+02
iter = 2 x =   1.000 -26.358 f(x) = -8.882e-16 -2.736e+02
iter = 3 x =   1.000   1.000 f(x) = -2.220e-16 -4.441e-15
status = success

The convergence is much more rapid, but takes a wide excursion out to the point (−4.23, −65.3). This could cause the algorithm to go astray in a realistic application. The hybrid algorithm follows the downhill path to the solution more reliably.

32.9 References and Further Reading

The original version of the Hybrid method is described in the following articles by Powell, M.J.D. Powell, "A Hybrid Method for Nonlinear Equations" (Chap 6, p 87-114) and "A Fortran Subroutine for Solving systems of Nonlinear Algebraic Equations" (Chap 7, p 115-161), in Numerical Methods for Nonlinear Algebraic Equations, P. Rabinowitz, editor. Gordon and Breach, 1970.

The following papers are also relevant to the algorithms described in this section, J.J. Moré, M.Y. Cosnard, "Numerical Solution of Nonlinear Equations", ACM Transactions on Mathematical Software, Vol 5, No 1, (1979), p 64-85 C.G. Broyden, "A Class of Methods for Solving Nonlinear Simultaneous Equations", Mathematics of Computation, Vol 19 (1965), p 577-593 J.J. Moré, B.S. Garbow, K.E. Hillstrom, "Testing Unconstrained Optimization Software", ACM Transactions on Mathematical Software, Vol 7, No 1 (1981), p 17-41

33 Multidimensional Minimization

This chapter describes routines for finding minima of arbitrary multidimensional functions. The library provides low level components for a variety of iterative minimizers and convergence tests. These can be combined by the user to achieve the desired solution, while providing full access to the intermediate steps of the algorithms. Each class of methods uses the same framework, so that you can switch between minimizers at runtime without needing to recompile your program. Each instance of a minimizer keeps track of its own state, allowing the minimizers to be used in multi-threaded programs. The minimization algorithms can be used to maximize a function by inverting its sign.

The header file ‘gsl_multimin.h’ contains prototypes for the minimization functions and related declarations.

33.1 Overview

The problem of multidimensional minimization requires finding a point x such that the scalar function,

f(x_{1}, …, x_{n})

takes a value which is lower than at any neighboring point. For smooth functions the gradient g = ∇f vanishes at the minimum. In general there are no bracketing methods available for the minimization of n-dimensional functions. All algorithms proceed from an initial guess using a search algorithm which attempts to move in a downhill direction. A one-dimensional line minimisation is performed along this direction until the lowest point is found to a suitable tolerance. The search direction is then updated with local information from the function and its derivatives, and the whole process repeated until the true ndimensional minimum is found.

Several minimization algorithms are available within a single framework. The user provides a high-level driver for the algorithms, and the library provides the individual functions necessary for each of the steps. There are three main phases of the iteration. The steps are,

• initialize minimizer state, s, for algorithm T

• update s using the iteration T

• test s for convergence, and repeat iteration if necessary

Each iteration step consists either of an improvement to the line-mimisation in the current direction or an update to the search direction itself. The state for the minimizers is held in

a gsl_multimin_fdfminimizer struct.

33.2 Caveats

Note that the minimization algorithms can only search for one local minimum at a time. When there are several local minima in the search area, the first minimum to be found will be returned; however it is difficult to predict which of the minima this will be. In most cases, no error will be reported if you try to find a local minimum in an area where there is more than one.

It is also important to note that the minimization algorithms find local minima; there is no way to determine whether a minimum is a global minimum of the function in question.

33.3 Initializing the Multidimensional Minimizer

The following function initializes a multidimensional minimizer. The minimizer itself depends only on the dimension of the problem and the algorithm and can be reused for different problems.

Function gsl_multimin_fdfminimizer *

gsl multimin fdfminimizer alloc (const

    gsl_multimin_fdfminimizer_type *T, size_t n)

This function returns a pointer to a a newly allocated instance of a minimizer of type T for an n-dimension function. If there is insufficient memory to create the minimizer then the function returns a null pointer and the error handler is invoked with an error

code of GSL_ENOMEM.

Function int gsl multimin fdfminimizer set (gsl_multimin_fdfminimizer

    * s, gsl_multimin_function_fdf *fdf, const gsl_vector * x, double

step size, double tol) This function initializes the minimizer s to minimize the function fdf starting from the initial point x. The size of the first trial step is given by step size. The accuracy of the line minimization is specified by tol. The precise meaning of this parameter depends on the method used. Typically the line minimization is considered successful if the gradient of the function g is orthogonal to the current search direction p to a relative accuracy of tol, where p · g < tol|p||g|.

Function void gsl multimin fdfminimizer free

    (gsl_multimin_fdfminimizer *s)

This function frees all the memory associated with the minimizer s.

Function const char * gsl multimin fdfminimizer name (const

    gsl_multimin_fdfminimizer * s)

This function returns a pointer to the name of the minimizer. For example,

     printf("s is a '%s' minimizer\n",
            gsl_multimin_fdfminimizer_name (s));
would print something like s is a 'conjugate_pr' minimizer.

33.4 Providing a function to minimize

You must provide a parametric function of n variables for the minimizers to operate on. You also need to provide a routine which calculates the gradient of the function and a third routine which calculates both the function value and the gradient together. In order to allow for general parameters the functions are defined by the following data type:

Data Type gsl multimin function fdf

This data type defines a general function of n variables with parameters and the corresponding gradient vector of derivatives,

double (* f) (const gsl_vector * x, void * params)

this function should return the result f(x, params) for argument x and parameters params.

int (* df) (const gsl_vector * x, void * params, gsl_vector * g)

this function should store the n-dimensional gradient g_{i} = ∂f(x, params)/∂x_{i} in the vector g for argument x and parameters params, returning an appropriate error code if the function cannot be computed.

int (* fdf) (const gsl_vector * x, void * params, double * f, gsl_vector *
g)

This function should set the values of the f and g as above, for arguments x and parameters params. This function provides an optimization of the separate functions for f(x) and g(x) – it is always faster to compute the function and its derivative at the same time.

size_t n the dimension of the system, i.e. the number of components of the vectors x.

void * params a pointer to the parameters of the function.

The following example function defines a simple paraboloid with two parameters,

/* Paraboloid centered on (dp[0],dp[1]) */
double
my_f (const gsl_vector *v, void *params)
{
  double x, y;
  double *dp = (double *)params;
  x = gsl_vector_get(v, 0);
  y = gsl_vector_get(v, 1);
  return 10.0 * (x - dp[0]) * (x - dp[0]) +
           20.0 * (y - dp[1]) * (y - dp[1]) + 30.0;
}
/* The gradient of f, df = (df/dx, df/dy). */
void
my_df (const gsl_vector *v, void *params,

       gsl_vector *df)
{
  double x, y;
  double *dp = (double *)params;
  x = gsl_vector_get(v, 0);
  y = gsl_vector_get(v, 1);
  gsl_vector_set(df, 0, 20.0 * (x - dp[0]));
  gsl_vector_set(df, 1, 40.0 * (y - dp[1]));
}
/* Compute both f and df together. */
void
my_fdf (const gsl_vector *x, void *params,
        double *f, gsl_vector *df)
{
  *f = my_f(x, params);
  my_df(x, params, df);
}

The function can be initialized using the following code,

gsl_multimin_function_fdf my_func;
double p[2] = { 1.0, 2.0 }; /* center at (1,2) */
my_func.f = &my_f;
my_func.df = &my_df;
my_func.fdf = &my_fdf;
my_func.n = 2;
my_func.params = (void *)p;

33.5 Iteration

The following function drives the iteration of each algorithm. The function performs one iteration to update the state of the minimizer. The same function works for all minimizers so that different methods can be substituted at runtime without modifications to the code.

Function int gsl multimin fdfminimizer iterate

    (gsl_multimin_fdfminimizer *s)

These functions perform a single iteration of the minimizer s. If the iteration encounters an unexpected problem then an error code will be returned.

The minimizer maintains a current best estimate of the minimum at all times. This information can be accessed with the following auxiliary functions,

Function gsl_vector * gsl multiroot fdfsolver x (const

    gsl_multiroot_fdfsolver * s)

Function double gsl multiroot fdfsolver minimum (const

    gsl_multiroot_fdfsolver * s)

Function gsl_vector * gsl multiroot fdfsolver gradient (const

    gsl_multiroot_fdfsolver * s)

These functions return the current best estimate of the location of the minimum, the value of the function at that point and its gradient, for the minimizer s.

Function int gsl multimin fdfminimizer restart

    (gsl_multimin_fdfminimizer *s)

This function resets the minimizer s to use the current point as a new starting point.

33.6 Stopping Criteria

A minimization procedure should stop when one of the following conditions is true:

• A minimum has been found to within the user-specified precision.

• A user-specified maximum number of iterations has been reached.

• An error has occurred.

The handling of these conditions is under user control. The functions below allow the user to test the precision of the current result.

Function int gsl multimin test gradient (const gsl_vector * g,double

epsabs) This function tests the norm of the gradient g against the absolute tolerance epsabs. The gradient of a multidimensional function goes to zero at a minimum. The test returns GSL_SUCCESS if the following condition is achieved,

|g| < epsabs

and returns GSL_CONTINUE otherwise. A suitable choice of epsabs can be made from the desired accuracy in the function for small variations in x. The relationship between these quantities is given by δf = g δx.

33.7 Algorithms

There are several minimization methods available. The best choice of algorithm depends on the problem. Each of the algorithms uses the value of the function and its gradient at each evaluation point.

Minimizer gsl multimin fdfminimizer conjugate fr

This is the Fletcher-Reeves conjugate gradient algorithm. The conjugate gradient algorithm proceeds as a succession of line minimizations. The sequence of search directions to build up an approximation to the curvature of the function in the neighborhood of the minimum. An initial search direction p is chosen using the gradient and line minimization is carried out in that direction. The accuracy of the line minimization is specified by the parameter tol. At the minimum along this line the function gradient g and the search direction p are orthogonal. The line minimization terminates when p · g < tol|p||g|. The search direction is updated using the Fletcher-Reeves formula p′ = g′ − βg where β = −|g′|^{2}/|g|^{2}, and the line minimization is then repeated for the new search direction.

Minimizer gsl multimin fdfminimizer conjugate pr

This is the Polak-Ribiere conjugate gradient algorithm. It is similar to the Fletcher-Reeves method, differing only in the choice of the coefficient β. Both methods work well when the evaluation point is close enough to the minimum of the objective function that it is well approximated by a quadratic hypersurface.

Minimizer gsl multimin fdfminimizer vector bfgs

This is the vector Broyden-Fletcher-Goldfarb-Shanno conjugate gradient algorithm. It is a quasi-Newton method which builds up an approximation to the second derivatives of the function f using the difference between successive gradient vectors. By combining the first and second derivatives the algorithm is able to take Newton-type steps towards the function minimum, assuming quadratic behavior in that region.

Minimizer gsl multimin fdfminimizer steepest descent

The steepest descent algorithm follows the downhill gradient of the function at each step. When a downhill step is successful the step-size is increased by factor of two. If the downhill step leads to a higher function value then the algorithm backtracks and the step size is decreased using the parameter tol. A suitable value of tol for most applications is 0.1. The steepest descent method is inefficient and is included only for demonstration purposes.

33.8 Examples

This example program finds the minimum of the paraboloid function defined earlier. The location of the minimum is offset from the origin in x and y, and the function value at the minimum is non-zero. The main program is given below, it requires the example function given earlier in this chapter.

int
main (void)
{
size_t iter = 0;
int status;
const gsl_multimin_fdfminimizer_type *T;
gsl_multimin_fdfminimizer *s;
/* Position of the minimum (1,2). */
double par[2] = { 1.0, 2.0 };
gsl_vector *x;
gsl_multimin_function_fdf my_func;
my_func.f = &my_f;
my_func.df = &my_df;
my_func.fdf = &my_fdf;
my_func.n = 2;
my_func.params = &par;
/* Starting point, x = (5,7) */
x = gsl_vector_alloc (2);
gsl_vector_set (x, 0, 5.0);
gsl_vector_set (x, 1, 7.0);
T = gsl_multimin_fdfminimizer_conjugate_fr;
s = gsl_multimin_fdfminimizer_alloc (T, 2);
gsl_multimin_fdfminimizer_set (s, &my_func, x, 0.01, 1e-4);
do
  {
    iter++;
    status = gsl_multimin_fdfminimizer_iterate (s);
    if (status)
      break;
    status = gsl_multimin_test_gradient (s->gradient, 1e-3);
    if (status == GSL_SUCCESS)
      printf ("Minimum found at:\n");

  printf ("%5d %.5f %.5f %10.5f\n", iter,
          gsl_vector_get (s->x, 0),
          gsl_vector_get (s->x, 1),
          s->f);
}
while (status == GSL_CONTINUE && iter < 100);
gsl_multimin_fdfminimizer_free (s);
gsl_vector_free (x);
return 0;
}

The initial step-size is chosen as 0.01, a conservative estimate in this case, and the line minimization parameter is set at 0.0001. The program terminates when the norm of the gradient has been reduced below 0.001. The output of the program is shown below,

     x       y         f
1 4.99629 6.99072  687.84780
2 4.98886 6.97215  683.55456
3 4.97400 6.93501  675.01278
4 4.94429 6.86073  658.10798
5 4.88487 6.71217  625.01340
6 4.76602 6.41506  561.68440
7 4.52833 5.82083  446.46694
8 4.05295 4.63238  261.79422
9 3.10219 2.25548   75.49762
10 2.85185 1.62963   67.03704
11 2.19088 1.76182   45.31640
12 0.86892 2.02622   30.18555
Minimum found at:
13 1.00000 2.00000   30.00000

Note that the algorithm gradually increases the step size as it successfully moves downhill, as can be seen by plotting the successive points.

The conjugate gradient algorithm finds the minimum on its second direction because the function is purely quadratic. Additional iterations would be needed for a more complicated function.

33.9 References and Further Reading

A brief description of multidimensional minimization algorithms and further references can be found in the following book,

C.W. Ueberhuber, Numerical Computation (Volume 2), Chapter 14, Section 4.4 "Minimization Methods", p. 325—335, Springer (1997), ISBN 3-540-62057-5.

34 Least-Squares Fitting

This chapter describes routines for performing least squares fits to experimental data using linear combinations of functions. The data may be weighted or unweighted. For weighted data the functions compute the best fit parameters and their associated covariance matrix. For unweighted data the covariance matrix is estimated from the scatter of the points, giving a variance-covariance matrix. The functions are divided into separate versions for simple one- or two-parameter regression and multiple-parameter fits. The functions are declared in the header file ‘gsl_fit.h’

34.1 Linear regression

The functions described in this section can be used to perform least-squares fits to a straight line model, Y = c_{0} + c_{1}X. For weighted data the best-fit is found by minimizing the weighted sum of squared residuals, χ^{2},

χ^{2} = ∑_{i} w_{i}(y_{i} − (c_{0} + c_{1}x_{i}))^{2}

for the parameters c_{0}, c_{1}. For unweighted data the sum is computed with w_{i} = 1.

Function int gsl fit linear (const double * x, const size_t xstride, const
    double * y, const size_t ystride, size_t n, double * c0, double * c1,
    double * cov00, double * cov01, double * cov11, double * sumsq)

This function computes the best-fit linear regression coefficients (c0,c1) of the model Y = c_{0} + c_{1}X for the datasets (x, y), two vectors of length n with strides xstride and ystride. The variance-covariance matrix for the parameters (c0, c1) is estimated from the scatter of the points around the best-fit line and returned via the parameters (cov00, cov01, cov11). The sum of squares of the residuals from the best-fit line is returned in sumsq.

Function int gsl fit wlinear (const double * x, const size_t xstride, const
    double * w, const size_t wstride, const double * y, const size_t ystride,
    size_t n, double * c0, double * c1, double * cov00, double * cov01,
    double * cov11, double * chisq)

This function computes the best-fit linear regression coefficients (c0,c1) of the model Y = c_{0} + c_{1}X for the weighted datasets (x, y), two vectors of length n with strides xstride and ystride. The vector w, of length n and stride wstride, specifies the weight of each datapoint. The weight is the reciprocal of the variance for each datapoint in y. The covariance matrix for the parameters (c0, c1) is estimated from weighted data and returned via the parameters (cov00, cov01, cov11). The weighted sum of squares of the residuals from the best-fit line, χ^{2}, is returned in chisq.

Function int gsl fit linear est (double x, double c0, double c1, double c00,
    double c01, double c11, double *y, double *y err)

This function uses the best-fit linear regression coefficients c0,c1 and their estimated covariance cov00,cov01,cov11 to compute the fitted function y and its standard deviation y err for the model Y = c_{0} + c_{1}X at the point x.

34.2 Linear fitting without a constant term

The functions described in this section can be used to perform least-squares fits to a straight line model without a constant term, Y = c_{1}X. For weighted data the best-fit is found by minimizing the weighted sum of squared residuals, χ^{2},

χ^{2} = ∑_{i} w_{i}(y_{i} − c_{1}x_{i})^{2}

for the parameter c_{1}. For unweighted data the sum is computed with w_{i} = 1.

Function int gsl fit mul (const double * x, const size_t xstride, const
    double * y, const size_t ystride, size_t n, double * c1, double * cov11,
    double * sumsq)

This function computes the best-fit linear regression coefficient c1 of the model Y = c_{1}X for the datasets (x, y), two vectors of length n with strides xstride and ystride. The variance of the parameter c1 is estimated from the scatter of the points around the best-fit line and returned via the parameter cov11. The sum of squares of the residuals from the best-fit line is returned in sumsq.

Function int gsl fit wmul (const double * x, const size_t xstride, const
    double * w, const size_t wstride, const double * y, const size_t ystride,
    size_t n, double * c1, double * cov11, double * sumsq)

This function computes the best-fit linear regression coefficient c1 of the model Y = c_{1}X for the weighted datasets (x, y), two vectors of length n with strides xstride and ystride. The vector w, of length n and stride wstride, specifies the weight of each datapoint. The weight is the reciprocal of the variance for each datapoint in y. The variance of the parameter c1 is estimated from the weighted data and returned via the parameters cov11. The weighted sum of squares of the residuals from the best-fit line, χ^{2}, is returned in chisq.

Function int gsl fit mul est (double x, double c1, double c11, double *y,
    double *y err)

This function uses the best-fit linear regression coefficient c1 and its estimated covariance cov11 to compute the fitted function y and its standard deviation y err for the model Y = c_{1}X at the point x.

34.3 Multi-parameter fitting

The functions described in this section perform least-squares fits to a general linear model, y = Xc where y is a vector of n observations, X is an n by p matrix of predictor variables, and c are the p unknown best-fit parameters, which are to be estimated.

The best-fit is found by minimizing the weighted sums of squared residuals, χ^{2},

χ^{2} = (y − Xc)^{T} W (y − Xc)

with respect to the parameters c. The weights are specified by the diagonal elements of the n by n matrix W. For unweighted data W is replaced by the identity matrix.

This formulation can be used for fits to any number of functions and/or variables by preparing the n-by-p matrix X appropriately. For example, to fit to a p-th order polynomial in x, use the following matrix,

X_{ij} = x_{i}^{j}

where the index i runs over the observations and the index j runs from 0 to p − 1. To fit to a set of p sinusoidal functions with fixed frequencies ω_{1}, ω_{2}, …, ω_{p}, use,

X_{ij} = sin(ω_{j}x_{i})

To fit to p independent variables x_{1}, x_{2}, …, x_{p}, use,

X_{ij} = x_{j}(i)

where x_{j}(i) is the i-th value of the predictor variable x_{j}.

The functions described in this section are declared in the header file ‘gsl_multifit.h’. The solution of the general linear least-squares system requires an additional working

space for intermediate results, such as the singular value decomposition of the matrix X.

Function gsl_multifit_linear_workspace * gsl multifit linear alloc
         (size_t n, size_t p)

This function allocates a workspace for fitting a model to n observations using p parameters.

Function void gsl multifit linear free (gsl_multifit_linear_workspace *

work) This function frees the memory associated with the workspace w.

Function int gsl multifit linear (const gsl_matrix * X, const gsl_vector *
         y, gsl_vector * c, gsl_matrix * cov, double * chisq,
         gsl_multifit_linear_workspace * work)

This function computes the best-fit parameters c of the model y = Xc for the observations y and the matrix of predictor variables X. The variance-covariance matrix of the model parameters cov is estimated from the scatter of the observations about the best-fit. The sum of squares of the residuals from the best-fit, χ^{2}, is returned in chisq. The best-fit is found by singular value decomposition of the matrix X using the preallocated workspace provided in work. The modified Golub-Reinsch SVD algorithm is used, with column scaling to improve the accuracy of the singular values. Any components which have zero singular value (to machine precision) are discarded from the fit.

Function int gsl multifit wlinear (const gsl_matrix * X, const gsl_vector
  * w, const gsl_vector * y, gsl_vector * c, gsl_matrix * cov, double *
  chisq, gsl_multifit_linear_workspace * work)

This function computes the best-fit parameters c of the model y = Xc for the observations y and the matrix of predictor variables X. The covariance matrix of the model parameters cov is estimated from the weighted data. The weighted sum of squares of the residuals from the best-fit, χ^{2}, is returned in chisq. The best-fit is found by singular value decomposition of the matrix X using the preallocated workspace provided in work. Any components which have zero singular value (to machine precision) are discarded from the fit.

34.4 Examples

The following program computes a least squares straight-line fit to a simple (fictitious) dataset, and outputs the best-fit line and its associated one standard-deviation error bars.

#include <stdio.h>
#include <gsl/gsl_fit.h>
int
main (void)
{
int i, n = 4;
double x[4] = { 1970, 1980, 1990, 2000 };
double y[4] = {   12,   11,   14,   13 };
double w[4] = {  0.1,  0.2,  0.3,  0.4 };
double c0, c1, cov00, cov01, cov11, chisq;
gsl_fit_wlinear (x, 1, w, 1, y, 1, n,
                 &c0, &c1, &cov00, &cov01, &cov11,
                 &chisq);
printf("# best fit: Y = %g + %g X\n", c0, c1);
printf("# covariance matrix:\n");
printf("# [ %g, %g\n#   %g, %g]\n",
       cov00, cov01, cov01, cov11);
printf("# chisq = %g\n", chisq);
for (i = 0; i < n; i++)
  printf("data: %g %g %g\n",
                x[i], y[i], 1/sqrt(w[i]));
printf("\n");
for (i = -30; i < 130; i++)
  {
    double xf = x[0] + (i/100.0) * (x[n-1] - x[0]);
    double yf, yf_err;
    gsl_fit_linear_est (xf,

                          c0, c1,
                          cov00, cov01, cov11,
                          &yf, &yf_err);
      printf("fit: %g %g\n", xf, yf);
      printf("hi : %g %g\n", xf, yf + yf_err);
      printf("lo : %g %g\n", xf, yf - yf_err);
    }
  return 0;
}

The following commands extract the data from the output of the program and display it using the gnu plotutils graph utility,

$ ./demo > tmp
$ more tmp
# best fit: Y = -106.6 + 0.06 X
# covariance matrix:
# [ 39602, -19.9
#   -19.9, 0.01]
# chisq = 0.8
$ for n in data fit hi lo ;
   do
     grep "^$n" tmp | cut -d: -f2 > $n ;
   done
$ graph -T X -X x -Y y -y 0 20 -m 0 -S 2 -Ie data
     -S 0 -I a -m 1 fit -m 2 hi -m 2 lo

The next program performs a quadratic fit y = c_{0} + c_{1}x + c_{2}x^{2} to a weighted dataset using the generalised linear fitting function gsl_multifit_wlinear. The model matrix X for a quadratic fit is given by,

X = ( \begin{matrix}1 & x_{0} & x_{0}^{2} \\ 1 & x_{1} & x_{1}^{2} \\ 1 & x_{2} & x_{2}^{2} \\ … & … & … \end{matrix} )

where the column of ones corresponds to the constant term c_{0}. The two remaining columns corresponds to the terms c_{1}x and and c_{2}x^{2}.

The program reads n lines of data in the format (x, y, err) where err is the error (standard deviation) in the value y.

#include <stdio.h>
#include <gsl/gsl_multifit.h>
int
main (int argc, char **argv)
{
int i, n;
double xi, yi, ei, chisq;
gsl_matrix *X, *cov;
gsl_vector *y, *w, *c;
if (argc != 2)
  {
    fprintf(stderr,"usage: fit n < data\n");
    exit (-1);
  }
n = atoi(argv[1]);
X = gsl_matrix_alloc (n, 3);
y = gsl_vector_alloc (n);
w = gsl_vector_alloc (n);
c = gsl_vector_alloc (3);
cov = gsl_matrix_alloc (3, 3);
for (i = 0; i < n; i++)
  {
    int count = fscanf(stdin, "%lg %lg %lg",
                       &xi, &yi, &ei);
    if (count != 3)
      {
        fprintf(stderr, "error reading file\n");
        exit(-1);
      }
    printf("%g %g +/- %g\n", xi, yi, ei);

    gsl_matrix_set (X, i, 0, 1.0);
    gsl_matrix_set (X, i, 1, xi);
    gsl_matrix_set (X, i, 2, xi*xi);
    gsl_vector_set (y, i, yi);
    gsl_vector_set (w, i, 1.0/(ei*ei));
  }
{
  gsl_multifit_linear_workspace * work
    = gsl_multifit_linear_alloc (n, 3);
  gsl_multifit_wlinear (X, w, y, c, cov,
                        &chisq, work);
  gsl_multifit_linear_free (work);
}
#define C(i) (gsl_vector_get(c,(i)))
#define COV(i,j) (gsl_matrix_get(cov,(i),(j)))
{
  printf("# best fit: Y = %g + %g X + %g X^2\n",
         C(0), C(1), C(2));
  printf("# covariance matrix:\n");
  printf("[ %+.5e, %+.5e, %+.5e  \n",
            COV(0,0), COV(0,1), COV(0,2));
  printf("  %+.5e, %+.5e, %+.5e  \n",
            COV(1,0), COV(1,1), COV(1,2));
  printf("  %+.5e, %+.5e, %+.5e ]\n",
            COV(2,0), COV(2,1), COV(2,2));
  printf("# chisq = %g\n", chisq);
}
return 0;
}

A suitable set of data for fitting can be generated using the following program. It outputs a set of points with gaussian errors from the curve y = e^{x} in the region 0 < x < 2.

#include <stdio.h>
#include <math.h>
#include <gsl/gsl_randist.h>
int
main (void)
{
double x;
const gsl_rng_type * T;
gsl_rng * r;
gsl_rng_env_setup();

  for (x = 0.1; x < 2; x+= 0.1)
    {
      double y0 = exp(x);
      double sigma = 0.1*y0;
      double dy = gsl_ran_gaussian(r, sigma)
      printf("%g %g %g\n", x, y0 + dy, sigma);
    }
  return 0;
}

The data can be prepared by running the resulting executable program,

$ ./generate > exp.dat
$ more exp.dat
0.1 0.97935 0.110517
0.2 1.3359 0.12214
0.3 1.52573 0.134986
0.4 1.60318 0.149182
0.5 1.81731 0.164872
0.6 1.92475 0.182212
....

To fit the data use the previous program, with the number of data points given as the first argument. In this case there are 19 data points.

$ ./fit 19 < exp.dat
0.1 0.97935 +/- 0.110517
0.2 1.3359 +/- 0.12214
...
# best fit: Y = 1.02318 + 0.956201 X + 0.876796 X^2
# covariance matrix:
[ +1.25612e-02, -3.64387e-02, +1.94389e-02
  -3.64387e-02, +1.42339e-01, -8.48761e-02
  +1.94389e-02, -8.48761e-02, +5.60243e-02 ]
# chisq = 23.0987

The parameters of the quadratic fit match the coefficients of the expansion of e^{x}, taking into account the errors on the parameters and the O(x^{3}) difference between the exponential and quadratic functions for the larger values of x. The errors on the parameters are given by the square-root of the corresponding diagonal elements of the covariance matrix. The chi-squared per degree of freedom is 1.4, indicating a reasonable fit to the data.

34.5 References and Further Reading

A summary of formulas and techniques for least squares fitting can be found in the "Statistics" chapter of the Annual Review of Particle Physics prepared by the Particle Data Group.

Review of Particle Properties R.M. Barnett et al., Physical Review D54, 1 (1996)

    http://pdg.lbl.gov/

The Review of Particle Physics is available online at the website given above. The tests used to prepare these routines are based on the NIST Statistical Reference Datasets. The datasets and their documentation are available from NIST at the following website,

              http://www.nist.gov/itl/div898/strd/index.html.

35 Nonlinear Least-Squares Fitting

This chapter describes functions for multidimensional nonlinear least-squares fitting. The library provides low level components for a variety of iterative solvers and convergence tests. These can be combined by the user to achieve the desired solution, with full access to the intermediate steps of the iteration. Each class of methods uses the same framework, so that you can switch between solvers at runtime without needing to recompile your program. Each instance of a solver keeps track of its own state, allowing the solvers to be used in multi-threaded programs.

The header file ‘gsl_multifit_nlin.h’ contains prototypes for the multidimensional nonlinear fitting functions and related declarations.

35.1 Overview

The problem of multidimensional nonlinear least-squares fitting requires the minimization of the squared residuals of n functions, f_{i}, in p parameters, x_{i},

Φ(x) = \frac{1}{2} ∑_{i=1}^{n} f_{i}(x_{1}, …, x_{p})^{2} = \frac{1}{2}||F (x)||^{2}

All algorithms proceed from an initial guess using the linearization,

ψ(p) = ||F (x + p)|| ≈ ||F (x) + Jp ||

where x is the initial point, p is the proposed step and J is the Jacobian matrix J_{ij} = ∂f_{i}/∂x_{j}. Additional strategies are used to enlarge the region of convergence. These include requiring a decrease in the norm ||F || on each step or using a trust region to avoid steps which fall outside the linear regime.

35.2 Initializing the Solver

Function gsl_multifit_fsolver * gsl multifit fsolver alloc (const

         gsl_multifit_fsolver_type * T, size_t n, size_t p)

This function returns a pointer to a a newly allocated instance of a solver of type T for n observations and p parameters. If there is insufficient memory to create the solver then the function returns a null pointer and the error handler is invoked with an error code of GSL_ENOMEM.

Function gsl_multifit_fdfsolver * gsl multifit fdfsolver alloc (const

         gsl_multifit_fdfsolver_type * T, size_t n, size_t p)

This function returns a pointer to a a newly allocated instance of a derivative solver of type T for n observations and p parameters. For example, the following code creates an instance of a Levenberg-Marquardt solver for 100 data points and 3 parameters,

          const gsl_multifit_fdfsolver_type * T
              = gsl_multifit_fdfsolver_lmder;
          gsl_multifit_fdfsolver * s
              = gsl_multifit_fdfsolver_alloc (T, 100, 3);

If there is insufficient memory to create the solver then the function returns a null pointer and the error handler is invoked with an error code of GSL_ENOMEM.

Function int gsl multifit fsolver set (gsl_multifit_fsolver * s,

    gsl_multifit_function * f, gsl_vector * x)

This function initializes, or reinitializes, an existing solver s to use the function f and the initial guess x.

Function int gsl multifit fdfsolver set (gsl_multifit_fdfsolver * s,

    gsl_function_fdf * fdf, gsl_vector * x)

This function initializes, or reinitializes, an existing solver s to use the function and derivative fdf and the initial guess x.

Function void gsl multifit fsolver free (gsl_multifit_fsolver * s) Function void gsl multifit fdfsolver free (gsl_multifit_fdfsolver * s)

These functions free all the memory associated with the solver s.

Function const char * gsl multifit fsolver name (const

    gsl_multifit_fdfsolver * s)

Function const char * gsl multifit fdfsolver name (const

    gsl_multifit_fdfsolver * s)

These functions return a pointer to the name of the solver. For example,

     printf("s is a '%s' solver\n",
            gsl_multifit_fdfsolver_name (s));

would print something like s is a 'lmder' solver.

35.3 Providing the Function to be Minimized

You must provide n functions of p variables for the minimization algorithms to operate on. In order to allow for general parameters the functions are defined by the following data types:

Data Type gsl multifit function

This data type defines a general system of functions with parameters.

int (* f) (const gsl_vector * x, void * params, gsl_vector * f )

this function should store the vector result f(x, params) in f for argument x and parameters params, returning an appropriate error code if the function cannot be computed.

size_t n the number of functions, i.e. the number of components of the vector f

size_t p the number of independent variables, i.e. the number of components of the vectors x

void * params a pointer to the parameters of the function

Data Type gsl multifit function fdf

This data type defines a general system of functions with parameters and the corresponding Jacobian matrix of derivatives,

int (* f) (const gsl_vector * x, void * params, gsl_vector * f )

this function should store the vector result f(x, params) in f for argument x and parameters params, returning an appropriate error code if the function cannot be computed.

int (* df) (const gsl_vector * x, void * params, gsl_matrix * J)

this function should store the n-by-p matrix result J_{ij} = ∂f_{i}(x, params)/∂x_{j} in J for argument x and parameters params, returning an appropriate error code if the function cannot be computed.

int (* fdf) (const gsl_vector * x, void * params, gsl_vector * f,
gsl_matrix * J)

This function should set the values of the f and J as above, for arguments x and parameters params. This function provides an optimization of the separate functions for f(x) and J(x) – it is always faster to compute the function and its derivative at the same time.

size_t n the number of functions, i.e. the number of components of the vector f

size_t p the number of independent variables, i.e. the number of components of the vectors x

void * params a pointer to the parameters of the function

35.4 Iteration

The following functions drive the iteration of each algorithm. Each function performs one iteration to update the state of any solver of the corresponding type. The same functions work for all solvers so that different methods can be substituted at runtime without modifications to the code.

Function int gsl multifit fsolver iterate (gsl_multifit_fsolver * s) Function int gsl multifit fdfsolver iterate (gsl_multifit_fdfsolver * s)

These functions perform a single iteration of the solver s. If the iteration encounters an unexpected problem then an error code will be returned. The solver maintains a current estimate of the best-fit parameters at all times. This information can be accessed with the following auxiliary functions,

Function gsl_vector * gsl multifit fsolver position (const

    gsl_multifit_fsolver * s)

Function gsl_vector * gsl multifit fdfsolver position (const

    gsl_multifit_fdfsolver * s)

These functions return the current position (i.e. best-fit parameters) of the solver s.

35.5 Search Stopping Parameters

A minimization procedure should stop when one of the following conditions is true:

• A minimum has been found to within the user-specified precision.

• A user-specified maximum number of iterations has been reached.

• An error has occurred.

The handling of these conditions is under user control. The functions below allow the user to test the current estimate of the best-fit parameters in several standard ways.

Function int gsl multifit test delta (const gsl_vector * dx, const

    gsl_vector * x, double epsabs, double epsrel)

This function tests for the convergence of the sequence by comparing the last step dx with the absolute error epsabs and relative error epsrel to the current position x. The test returns GSL_SUCCESS if the following condition is achieved,

|dx_{i}| < epsabs + epsrel |x_{i}|

for each component of x and returns GSL_CONTINUE otherwise.

Function int gsl multifit test gradient (const gsl_vector * g, double

epsabs) This function tests the residual gradient g against the absolute error bound epsabs. Mathematically, the gradient should be exactly zero at the minimum. The test returns GSL_SUCCESS if the following condition is achieved,

∑_{i} |g_{i}| < epsabs

and returns GSL_CONTINUE otherwise. This criterion is suitable for situations where the the precise location of the minimum, x, is unimportant provided a value can be found where the gradient is small enough.

Function int gsl multifit gradient (const gsl_matrix * J, const

    gsl_vector * f, gsl_vector * g)

This function computes the gradient g of Φ(x) = (1/2)||F (x)||^{2} from the Jacobian matrix J and the function values f, using the formula g = J^{T} f.

35.6 Minimization Algorithms using Derivatives

The minimization algorithms described in this section make use of both the function and its derivative. They require an initial guess for the location of the minimum. There is no absolute guarantee of convergence – the function must be suitable for this technique and the initial guess must be sufficiently close to the minimum for it to work.

Derivative Solver gsl multifit fdfsolver lmsder

This is a robust and efficient version of the Levenberg-Marquardt algorithm as implemented in the scaled lmder routine in minpack. Minpack was written by Jorge J. Moré, Burton S. Garbow and Kenneth E. Hillstrom. The algorithm uses a generalized trust region to keep each step under control. In order to be accepted a proposed new position x′ must satisfy the condition |D(x′ − x)| < δ, where D is a diagonal scaling matrix and δ is the size of the trust region. The components of D are computed internally, using the column norms of the Jacobian to estimate the sensitivity of the residual to each component of x. This improves the behavior of the algorithm for badly scaled functions. On each iteration the algorithm attempts to minimize the linear system |F + Jp| subject to the constraint |Dp| < Δ. The solution to this constrained linear system is found using the Levenberg-Marquardt method. The proposed step is now tested by evaluating the function at the resulting point, x′. If the step reduces the norm of the function sufficiently, and follows the predicted behavior of the function within the trust region. then it is accepted and size of the trust region is increased. If the proposed step fails to improve the solution, or differs significantly from the expected behavior within the trust region, then the size of the trust region is decreased and another trial step is computed. The algorithm also monitors the progress of the solution and returns an error if the changes in the solution are smaller than the machine precision. The possible error codes are,

GSL_ETOLF

the decrease in the function falls below machine precision

GSL_ETOLX

the change in the position vector falls below machine precision

GSL_ETOLG

the norm of the gradient, relative to the norm of the function, falls below machine precision

These error codes indicate that further iterations will be unlikely to change the solution from its current value.

Derivative Solver gsl multifit fdfsolver lmder

This is an unscaled version of the lmder algorithm. The elements of the diagonal scaling matrix D are set to 1. This algorithm may be useful in circumstances where the scaled version of lmder converges too slowly, or the function is already scaled appropriately.

35.7 Minimization Algorithms without Derivatives

There are no algorithms implemented in this section at the moment.

35.8 Computing the covariance matrix of best fit parameters

Function int gsl multifit covar (const gsl_matrix * J, double epsrel,

    gsl_matrix * covar)

This function uses the Jacobian matrix J to compute the covariance matrix of the best-fit parameters, covar. The parameter epsrel is used to remove linear-dependent columns when J is rank deficient. The covariance matrix is given by,

C = (J^{T} J)^{−1}

and is computed by QR decomposition of J with column-pivoting. Any columns of R which satisfy

|R_{kk}| ≤ epsrel|R_{11}|

are considered linearly-dependent and are excluded from the covariance matrix (the corresponding rows and columns of the covariance matrix are set to zero).

35.9 Examples

The following example program fits a weighted exponential model with background to experimental data, Y = A exp(−λt) + b. The first part of the program sets up the functions expb_f and expb_df to calculate the model and its Jacobian. The appropriate fitting function is given by,

f_{i} = ((A exp(−λt_{i}) + b) − y_{i})/σ_{i}

where we have chosen t_{i} = i. The Jacobian matrix J is the derivative of these functions with respect to the three parameters (A, λ, b). It is given by,

J_{ij} = \frac{∂f_{i}}{∂x_{j}}

where x_{0} = A, x_{1} = λ and x_{2} = b.

#include <stdlib.h>
#include <stdio.h>
#include <gsl/gsl_rng.h>
#include <gsl/gsl_randist.h>
#include <gsl/gsl_vector.h>
#include <gsl/gsl_blas.h>
#include <gsl/gsl_multifit_nlin.h>
struct data {
  size_t n;
  double * y;
  double * sigma;
};

int
expb_f (const gsl_vector * x, void *params,
      gsl_vector * f)
{
size_t n = ((struct data *)params)->n;
double *y = ((struct data *)params)->y;
double *sigma = ((struct data *) params)->sigma;
double A = gsl_vector_get (x, 0);
double lambda = gsl_vector_get (x, 1);
double b = gsl_vector_get (x, 2);
size_t i;
for (i = 0; i < n; i++)
  {
    /* Model Yi = A * exp(-lambda * i) + b */
    double t = i;
    double Yi = A * exp (-lambda * t) + b;
    gsl_vector_set (f, i, (Yi - y[i])/sigma[i]);
  }
return GSL_SUCCESS;
}
int
expb_df (const gsl_vector * x, void *params,
       gsl_matrix * J)
{
size_t n = ((struct data *)params)->n;
double *sigma = ((struct data *) params)->sigma;
double A = gsl_vector_get (x, 0);
double lambda = gsl_vector_get (x, 1);
size_t i;
for (i = 0; i < n; i++)
  {
    /* Jacobian matrix J(i,j) = dfi / dxj, */
    /* where fi = (Yi - yi)/sigma[i],      */
    /*       Yi = A * exp(-lambda * i) + b  */
    /* and the xj are the parameters (A,lambda,b) */
    double t = i;
    double s = sigma[i];
    double e = exp(-lambda * t);
    gsl_matrix_set (J, i, 0, e/s);
    gsl_matrix_set (J, i, 1, -t * A * e/s);
    gsl_matrix_set (J, i, 2, 1/s);

  }
return GSL_SUCCESS;
}
int
expb_fdf (const gsl_vector * x, void *params,
        gsl_vector * f, gsl_matrix * J)
{
expb_f (x, params, f);
expb_df (x, params, J);
return GSL_SUCCESS;
}

The main part of the program sets up a Levenberg-Marquardt solver and some simulated random data. The data uses the known parameters (1.0,5.0,0.1) combined with gaussian noise (standard deviation = 0.1) over a range of 40 timesteps. The initial guess for the parameters is chosen as (0.0, 1.0, 0.0).

int
main (void)
{
const gsl_multifit_fdfsolver_type *T;
gsl_multifit_fdfsolver *s;
int status;
size_t i, iter = 0;
const size_t n = 40;
const size_t p = 3;
gsl_matrix *covar = gsl_matrix_alloc (p, p);
double y[n], sigma[n];
struct data d = { n, y, sigma};
gsl_multifit_function_fdf f;
double x_init[3] = { 1.0, 0.0, 0.0 };
gsl_vector_view x = gsl_vector_view_array (x_init, p);
const gsl_rng_type * type;
gsl_rng * r;
gsl_rng_env_setup();
type = gsl_rng_default;
r = gsl_rng_alloc (type);

f.f = &expb_f;
f.df = &expb_df;
f.fdf = &expb_fdf;
f.n = n;
f.p = p;
f.params = &d;
/* This is the data to be fitted */
for (i = 0; i < n; i++)
  {
    double t = i;
    y[i] = 1.0 + 5 * exp (-0.1 * t)
               + gsl_ran_gaussian(r, 0.1);
    sigma[i] = 0.1;
    printf("data: %d %g %g\n", i, y[i], sigma[i]);
  };

T = gsl_multifit_fdfsolver_lmsder;
s = gsl_multifit_fdfsolver_alloc (T, n, p);
gsl_multifit_fdfsolver_set (s, &f, &x.vector);
print_state (iter, s);
do
  {
    iter++;
    status = gsl_multifit_fdfsolver_iterate (s);
    printf ("status = %s\n", gsl_strerror (status));
    print_state (iter, s);
    if (status)
      break;
    status = gsl_multifit_test_delta (s->dx, s->x,
                                      1e-4, 1e-4);
  }
while (status == GSL_CONTINUE && iter < 500);
gsl_multifit_covar (s->J, 0.0, covar);
gsl_matrix_fprintf (stdout, covar, "%g");
#define FIT(i) gsl_vector_get(s->x, i)
#define ERR(i) sqrt(gsl_matrix_get(covar,i,i))
printf("A      = %.5f +/- %.5f\n", FIT(0), ERR(0));
printf("lambda = %.5f +/- %.5f\n", FIT(1), ERR(1));

  printf("b      = %.5f +/- %.5f\n", FIT(2), ERR(2));
  printf ("status = %s\n", gsl_strerror (status));
  gsl_multifit_fdfsolver_free (s);
  return 0;
}
int
print_state (size_t iter, gsl_multifit_fdfsolver * s)
{
  printf ("iter: %3u x = % 15.8f % 15.8f % 15.8f "
          "|f(x)| = %g\n",
          iter,
          gsl_vector_get (s->x, 0),
          gsl_vector_get (s->x, 1),
          gsl_vector_get (s->x, 2),
          gsl_blas_dnrm2 (s->f));
}

The iteration terminates when the change in x is smaller than 0.0001, as both an absolute and relative change. Here are the results of running the program,

iter: 0 x = 1.00000000 0.00000000 0.00000000 |f(x)| = 118.574
iter: 1 x = 1.64919392 0.01780040 0.64919392 |f(x)| = 77.2068
iter: 2 x = 2.86269020 0.08032198 1.45913464 |f(x)| = 38.0579
iter: 3 x = 4.97908864 0.11510525 1.06649948 |f(x)| = 10.1548
iter: 4 x = 5.03295496 0.09912462 1.00939075 |f(x)| = 6.4982
iter: 5 x = 5.05811477 0.10055914 0.99819876 |f(x)| = 6.33121
iter: 6 x = 5.05827645 0.10051697 0.99756444 |f(x)| = 6.33119
iter: 7 x = 5.05828006 0.10051819 0.99757710 |f(x)| = 6.33119
A      = 5.05828 +/- 0.05983
lambda = 0.10052 +/- 0.00309
b      = 0.99758 +/- 0.03944
status = success

The approximate values of the parameters are found correctly. The errors on the parameters are given by the square roots of the diagonal elements of the covariance matrix.

35.10 References and Further Reading

The minpack algorithm is described in the following article, J.J. Moré, The Levenberg-Marquardt Algorithm: Implementation and Theory, Lecture Notes in Mathematics, v630 (1978), ed G. Watson.

The following paper is also relevant to the algorithms described in this section, J.J. Moré, B.S. Garbow, K.E. Hillstrom, "Testing Unconstrained Optimization Software", ACM Transactions on Mathematical Software, Vol 7, No 1 (1981), p 17-41

36 Physical Constants

This chapter describes macros for the values of physical constants, such as the speed of light, c, and gravitational constant, G. The values are available in different unit systems, including the standard MKS system (meters, kilograms, seconds) and the CGS system (centimeters, grams, seconds), which is commonly used in Astronomy.

The definitions of constants in the MKS system are available in the file ‘gsl_const_mks.h’. The constants in the CGS system are defined in ‘gsl_const_cgs.h’. Dimensionless constants, such as the fine structure constant, which are pure numbers are

defined in ‘gsl_const_num.h’.

The full list of constants is described briefly below. Consult the header files themselves for the values of the constants used in the library.

36.1 Fundamental Constants

GSL_CONST_MKS_SPEED_OF_LIGHT

The speed of light in vacuum, c.

GSL_CONST_NUM_AVOGADRO

Avogadro’s number, N_{a}.

GSL_CONST_MKS_FARADAY

The molar charge of 1 Faraday.

GSL_CONST_MKS_BOLTZMANN

The Boltzmann constant, k.

GSL_CONST_MKS_MOLAR_GAS

The molar gas constant, R_{0}.

GSL_CONST_MKS_STANDARD_GAS_VOLUME

The standard gas volume, V_{0}.

GSL_CONST_MKS_GAUSS

The magnetic field of 1 Gauss.

GSL_CONST_MKS_MICRON

The length of 1 micron.

GSL_CONST_MKS_HECTARE

The area of 1 hectare.

GSL_CONST_MKS_MILES_PER_HOUR

The speed of 1 mile per hour.

GSL_CONST_MKS_KILOMETERS_PER_HOUR

The speed of 1 kilometer per hour.

36.2 Astronomy and Astrophysics

GSL_CONST_MKS_ASTRONOMICAL_UNIT

The length of 1 astronomical unit (mean earth-sun distance), au.

GSL_CONST_MKS_GRAVITATIONAL_CONSTANT

The gravitational constant, G.

GSL_CONST_MKS_LIGHT_YEAR

The distance of 1 light-year, ly.

GSL_CONST_MKS_PARSEC

The distance of 1 parsec, pc.

GSL_CONST_MKS_GRAV_ACCEL

The standard gravitational acceleration on Earth, g.

GSL_CONST_MKS_SOLAR_MASS

The mass of the Sun.

36.3 Atomic and Nuclear Physics

GSL_CONST_MKS_ELECTRON_CHARGE

The charge of the electron, e.

GSL_CONST_MKS_ELECTRON_VOLT

The energy of 1 electron volt, eV.

GSL_CONST_MKS_UNIFIED_ATOMIC_MASS

The unified atomic mass, amu.

GSL_CONST_MKS_MASS_ELECTRON

The mass of the electron, m_{e}.

GSL_CONST_MKS_MASS_MUON

The mass of the muon, m_{μ}.

GSL_CONST_MKS_MASS_PROTON

The mass of the proton, m_{p}.

GSL_CONST_MKS_MASS_NEUTRON

The mass of the neutron, m_{n}.

GSL_CONST_NUM_FINE_STRUCTURE

The electromagnetic fine structure constant α.

GSL_CONST_MKS_RYDBERG

The Rydberg constant, Ry.

GSL_CONST_MKS_ANGSTROM

The length of 1 angstrom.

GSL_CONST_MKS_BARN

The area of 1 barn.

GSL_CONST_MKS_BOHR_MAGNETON

The Bohr Magneton, μ_{B}.

GSL_CONST_MKS_NUCLEAR_MAGNETON

The Nuclear Magneton, μ_{N}.

GSL_CONST_MKS_ELECTRON_MAGNETIC_MOMENT

The magnetic moment of the electron, μ_{e}.

GSL_CONST_MKS_PROTON_MAGNETIC_MOMENT

The magnetic moment of the proton, μ_{p}.

36.4 Measurement of Time

GSL_CONST_MKS_MINUTE

The number of seconds in 1 minute.

GSL_CONST_MKS_HOUR

The number of seconds in 1 hour.

GSL_CONST_MKS_DAY

The number of seconds in 1 day.

GSL_CONST_MKS_WEEK

The number of seconds in 1 week.

36.5 Imperial Units

GSL_CONST_MKS_INCH

The length of 1 inch.

GSL_CONST_MKS_FOOT

The length of 1 foot.

GSL_CONST_MKS_YARD

The length of 1 yard.

GSL_CONST_MKS_MILE

The length of 1 mile.

GSL_CONST_MKS_MIL

The length of 1 mil (1/1000th of an inch).

36.6 Nautical Units

GSL_CONST_MKS_NAUTICAL_MILE

The length of 1 nautical mile.

GSL_CONST_MKS_FATHOM

The length of 1 fathom.

GSL_CONST_MKS_KNOT

The speed of 1 knot.

36.7 Printers Units

GSL_CONST_MKS_POINT

The length of 1 printer’s point (1/72 inch).

GSL_CONST_MKS_TEXPOINT

The length of 1 TeX point (1/72.27 inch).

36.8 Volume

GSL_CONST_MKS_ACRE

The area of 1 acre.

GSL_CONST_MKS_LITER

The volume of 1 liter.

GSL_CONST_MKS_US_GALLON

The volume of 1 US gallon.

GSL_CONST_MKS_CANADIAN_GALLON

The volume of 1 Canadian gallon.

GSL_CONST_MKS_UK_GALLON

The volume of 1 UK gallon.

GSL_CONST_MKS_QUART

The volume of 1 quart.

GSL_CONST_MKS_PINT

The volume of 1 pint.

36.9 Mass and Weight

GSL_CONST_MKS_POUND_MASS

The mass of 1 pound.

GSL_CONST_MKS_OUNCE_MASS

The mass of 1 ounce.

GSL_CONST_MKS_TON

The mass of 1 ton.

GSL_CONST_MKS_METRIC_TON

The mass of 1 metric ton (1000 kg).

GSL_CONST_MKS_UK_TON

The mass of 1 UK ton.

GSL_CONST_MKS_TROY_OUNCE

The mass of 1 troy ounce.

GSL_CONST_MKS_CARAT

The mass of 1 carat.

GSL_CONST_MKS_GRAM_FORCE

The force of 1 gram weight.

GSL_CONST_MKS_POUND_FORCE

The force of 1 pound weight.

GSL_CONST_MKS_KILOPOUND_FORCE

The force of 1 kilopound weight.

GSL_CONST_MKS_POUNDAL

The force of 1 poundal.

36.10 Thermal Energy and Power

GSL_CONST_MKS_CALORIE

The energy of 1 calorie.

GSL_CONST_MKS_BTU

The energy of 1 British Thermal Unit, btu.

GSL_CONST_MKS_THERM

The energy of 1 Therm.

GSL_CONST_MKS_HORSEPOWER

The power of 1 horsepower.

36.11 Pressure

GSL_CONST_MKS_BAR

The pressure of 1 bar.

GSL_CONST_MKS_STD_ATMOSPHERE

The pressure of 1 standard atmosphere.

GSL_CONST_MKS_TORR

The pressure of 1 torr.

GSL_CONST_MKS_METER_OF_MERCURY

The pressure of 1 meter of mercury.

GSL_CONST_MKS_INCH_OF_MERCURY

The pressure of 1 inch of mercury.

GSL_CONST_MKS_INCH_OF_WATER

The pressure of 1 inch of water.

GSL_CONST_MKS_PSI

The pressure of 1 pound per square inch.

36.12 Viscosity

GSL_CONST_MKS_POISE

The dynamic viscosity of 1 poise.

GSL_CONST_MKS_STOKES

The kinematic viscosity of 1 stokes.

36.13 Light and Illumination

GSL_CONST_MKS_STILB

The luminance of 1 stilb.

GSL_CONST_MKS_LUMEN

The luminous flux of 1 lumen.

GSL_CONST_MKS_LUX

The illuminance of 1 lux.

GSL_CONST_MKS_PHOT

The illuminance of 1 phot.

GSL_CONST_MKS_FOOTCANDLE

The illuminance of 1 footcandle.

GSL_CONST_MKS_LAMBERT

The luminance of 1 lambert.

GSL_CONST_MKS_FOOTLAMBERT

The luminance of 1 footlambert.

36.14 Radioactivity

GSL_CONST_MKS_CURIE

The activity of 1 curie.

GSL_CONST_MKS_ROENTGEN

The exposure of 1 roentgen.

GSL_CONST_MKS_RAD

The absorbed dose of 1 rad.

36.15 Examples

The following program demonstrates the use of the physical constants in a calculation. In this case, the goal is to calculate the range of light-travel times from Earth to Mars.

The required data is the average distance of each planet from the Sun in astronomical units (the eccentricities of the orbits will be neglected for the purposes of this calculation). The average radius of the orbit of Mars is 1.52 astronomical units, and for the orbit of Earth it is 1 astronomical unit (by definition). These values are combined with the MKS values of the constants for the speed of light and the length of an astronomical unit to produce a result for the shortest and longest light-travel times in seconds. The figures are converted into minutes before being displayed.

     #include <stdio.h>
     #include <gsl/gsl_const_mks.h>
     int
     main (void)
     {
       double c  = GSL_CONST_MKS_SPEED_OF_LIGHT;
       double au = GSL_CONST_MKS_ASTRONOMICAL_UNIT;

double minutes = GSL_CONST_MKS_MINUTE;
/* distance stored in meters */
double r_earth = 1.00 * au;
double r_mars  = 1.52 * au;
double t_min, t_max;
t_min = (r_mars - r_earth) / c;
t_max = (r_mars + r_earth) / c;
printf("light travel time from Earth to Mars:\n");
printf("minimum = %.1f minutes\n", t_min / minutes);
printf("maximum = %.1f minutes\n", t_max / minutes);
return 0;
}

Here is the output from the program,

light travel time from Earth to Mars:
minimum = 4.3 minutes
maximum = 21.0 minutes

36.16 References and Further Reading

Further information on the values of physical constants is available from the NIST website,

http://www.physics.nist.gov/cuu/Constants/index.html

37 IEEE floating-point arithmetic

This chapter describes functions for examining the representation of floating point numbers and controlling the floating point environment of your program. The functions described in this chapter are declared in the header file ‘gsl_ieee_utils.h’.

37.1 Representation of floating point numbers

The IEEE Standard for Binary Floating-Point Arithmetic defines binary formats for single and double precision numbers. Each number is composed of three parts: a sign bit (s), an exponent (E) and a fraction (f). The numerical value of the combination (s, E, f) is given by the following formula,

(−1)^{s}(1 · fffff …)2^{E}

The sign bit is either zero or one. The exponent ranges from a minimum value E_{min} to a maximum value E_{max} depending on the precision. The exponent is converted to an unsigned number e, known as the biased exponent, for storage by adding a bias parameter, e = E + bias. The sequence fffff... represents the digits of the binary fraction f. The binary digits are stored in normalized form, by adjusting the exponent to give a leading digit of 1. Since the leading digit is always 1 for normalized numbers it is assumed implicitly and does not have to be stored. Numbers smaller than 2^{E_{min}} are be stored in denormalized form with a leading zero,

(−1)^{s}(0 · fffff …)2^{E_{min}}

This allows gradual underflow down to 2^{E_{min}−p} for p bits of precision. A zero is encoded with the special exponent of 2^{E_{min}−1} and infinities with the exponent of 2^{E_{max}+1}. The format for single precision numbers uses 32 bits divided in the following way,

     seeeeeeeefffffffffffffffffffffff
     s = sign bit, 1 bit
     e = exponent, 8 bits  (E_min=-126, E_max=127, bias=127)
     f = fraction, 23 bits

The format for double precision numbers uses 64 bits divided in the following way,

     seeeeeeeeeeeffffffffffffffffffffffffffffffffffffffffffffffffffff
     s = sign bit, 1 bit
     e = exponent, 11 bits  (E_min=-1022, E_max=1023, bias=1023)
     f = fraction, 52 bits

It is often useful to be able to investigate the behavior of a calculation at the bit-level and the library provides functions for printing the IEEE representations in a human-readable form.

Function void gsl ieee fprintf float (FILE * stream, const float * x) Function void gsl ieee fprintf double (FILE * stream, const double * x)

These functions output a formatted version of the IEEE floating-point number pointed to by x to the stream stream. A pointer is used to pass the number indirectly, to avoid any undesired promotion from float to double. The output takes one of the following forms, NaN the Not-a-Number symbol

Inf, -Inf positive or negative infinity

1.fffff...*2^E, -1.fffff...*2^E

a normalized floating point number

0.fffff...*2^E, -0.fffff...*2^E

a denormalized floating point number

0, -0 positive or negative zero

The output can be used directly in GNU Emacs Calc mode by preceding it with 2# to indicate binary.

Function void gsl ieee printf float (const float * x) Function void gsl ieee printf double (const double * x)

These functions output a formatted version of the IEEE floating-point number pointed to by x to the stream stdout.

The following program demonstrates the use of the functions by printing the single and double precision representations of the fraction 1/3. For comparison the representation of the value promoted from single to double precision is also printed.

#include <stdio.h>
#include <gsl/gsl_ieee_utils.h>
int
main (void)
{
  float f = 1.0/3.0;
  double d = 1.0/3.0;
  double fd = f; /* promote from float to double */
  printf(" f="); gsl_ieee_printf_float(&f);
  printf("\n");
  printf("fd="); gsl_ieee_printf_double(&fd);
  printf("\n");
  printf(" d="); gsl_ieee_printf_double(&d);
  printf("\n");
  return 0;
}

The binary representation of 1/3 is 0.01010101.... The output below shows that the IEEE format normalizes this fraction to give a leading digit of 1,

 f= 1.01010101010101010101011*2^-2
fd= 1.0101010101010101010101100000000000000000000000000000*2^-2
 d= 1.0101010101010101010101010101010101010101010101010101*2^-2

The output also shows that a single-precision number is promoted to double-precision by adding zeros in the binary representation.

37.2 Setting up your IEEE environment

The IEEE standard defines several modes for controlling the behavior of floating point operations. These modes specify the important properties of computer arithmetic: the direction used for rounding (e.g. whether numbers should be rounded up, down or to the nearest number), the rounding precision and how the program should handle arithmetic exceptions, such as division by zero.

Many of these features can now be controlled via standard functions such as fpsetround, which should be used whenever they are available. Unfortunately in the past there has been no universal API for controlling their behavior – each system has had its own way of accessing them. For example, the Linux kernel provides the function __setfpucw (set-fpu-control-word) to set IEEE modes, while HP-UX and Solaris use the functions fpsetround and fpsetmask. To help you write portable programs GSL allows you to specify modes in a platform-independent way using the environment variable GSL_IEEE_MODE. The library then takes care of all the necessary machine-specific initializations for you when you call

the function gsl_ieee_env_setup.

Function void gsl ieee env setup ()

This function reads the environment variable GSL_IEEE_MODE and attempts to set up the corresponding specified IEEE modes. The environment variable should be a list of keywords, separated by commas, like this,

GSL_IEEE_MODE = "keyword,keyword,..." where keyword is one of the following mode-names,

single-precision
double-precision
extended-precision
round-to-nearest
round-down
round-up
round-to-zero
mask-all
mask-invalid
mask-denormalized
mask-division-by-zero
mask-overflow
mask-underflow
trap-inexact
trap-common

If GSL_IEEE_MODE is empty or undefined then the function returns immediately and no attempt is made to change the system’s IEEE mode. When the modes from GSL_ IEEE_MODE are turned on the function prints a short message showing the new settings to remind you that the results of the program will be affected. If the requested modes are not supported by the platform being used then the function calls the error handler and returns an error code of GSL_EUNSUP. The following combination of modes is convenient for many purposes,

     GSL_IEEE_MODE="double-precision,"\
                     "mask-underflow,"\
                       "mask-denormalized"

This choice ignores any errors relating to small numbers (either denormalized, or underflowing to zero) but traps overflows, division by zero and invalid operations.

To demonstrate the effects of different rounding modes consider the following program which computes e, the base of natural logarithms, by summing a rapidly-decreasing series,

e = 1 + \frac{1}{2!} + \frac{1}{3!} + \frac{1}{4!} + … = 2.71828182846...

#include <math.h>
#include <stdio.h>
#include <gsl/gsl_ieee_utils.h>
int
main (void)
{
  double x = 1, oldsum = 0, sum = 0;
  int i = 0;
  gsl_ieee_env_setup (); /* read GSL_IEEE_MODE */
  do
    {
      i++;
      oldsum = sum;
      sum += x;
      x = x / i;
      printf("i=%2d sum=%.18f error=%g\n",
             i, sum, sum - M_E);
      if (i > 30)
         break;
    }
  while (sum != oldsum);
  return 0;
}

Here are the results of running the program in round-to-nearest mode. This is the IEEE default so it isn’t really necessary to specify it here,

GSL_IEEE_MODE="round-to-nearest" ./a.out
i= 1 sum=1.000000000000000000 error=-1.71828
i= 2 sum=2.000000000000000000 error=-0.718282
....
i=18 sum=2.718281828459045535 error=4.44089e-16
i=19 sum=2.718281828459045535 error=4.44089e-16

After nineteen terms the sum converges to within 4 × 10^{−16} of the correct value. If we now change the rounding mode to round-down the final result is less accurate,

     GSL_IEEE_MODE="round-down" ./a.out
     i= 1 sum=1.000000000000000000 error=-1.71828
     ....
     i=19 sum=2.718281828459041094 error=-3.9968e-15

The result is about 4 × 10^{−15} below the correct value, an order of magnitude worse than the result obtained in the round-to-nearest mode.

If we change to rounding mode to round-up then the series no longer converges (the reason is that when we add each term to the sum the final result is always rounded up. This is guaranteed to increase the sum by at least one tick on each iteration). To avoid this problem we would need to use a safer converge criterion, such as while (fabs(sum - oldsum) > epsilon), with a suitably chosen value of epsilon.

Finally we can see the effect of computing the sum using single-precision rounding, in the default round-to-nearest mode. In this case the program thinks it is still using double precision numbers but the CPU rounds the result of each floating point operation to singleprecision accuracy. This simulates the effect of writing the program using single-precision float variables instead of double variables. The iteration stops after about half the number of iterations and the final result is much less accurate,

     GSL_IEEE_MODE="single-precision" ./a.out
     ....
     i=12 sum=2.718281984329223633 error=1.5587e-07

with an error of O(10^{−7}), which corresponds to single precision accuracy (about 1 part in 10^{7}). Continuing the iterations further does not decrease the error because all the subsequent results are rounded to the same value.

37.3 References and Further Reading

The reference for the IEEE standard is, ANSI/IEEE Std 754-1985, IEEE Standard for Binary Floating-Point Arithmetic

A more pedagogical introduction to the standard can be found in the paper “What Every Computer Scientist Should Know About Floating-Point Arithmetic”.

David Goldberg: What Every Computer Scientist Should Know About Floating-Point Arithmetic. ACM Computing Surveys, Vol. 23, No. 1 (March 1991), pages 5-48 Corrigendum: ACM Computing Surveys, Vol. 23, No. 3 (September 1991), page 413. See also the sections by B. A. Wichmann and Charles B. Dunham in Surveyor’s Forum: “What Every Computer Scientist Should Know About Floating-Point Arithmetic”. ACM Computing Surveys, Vol. 24, No. 3 (September 1992), page 319

Appendix A Debugging Numerical Programs

This chapter describes some tips and tricks for debugging numerical programs which use GSL.

A.1 Using gdb

Any errors reported by the library are passed to the function gsl_error. By running your programs under gdb and setting a breakpoint in this function you can automatically catch any library errors. You can add a breakpoint for every session by putting

     break gsl_error

into your ‘.gdbinit’ file in the directory where your program is started.

If the breakpoint catches an error then you can use a backtrace (bt) to see the calltree, and the arguments which possibly caused the error. By moving up into the calling function you can investigate the values of variable at that point. Here is an example from the program fft/test_trap, which contains the following line,

     status = gsl_fft_complex_wavetable_alloc (0, &complex_wavetable);

The function gsl_fft_complex_wavetable_alloc takes the length of an FFT as its first argument. When this line is executed an error will be generated because the length of an FFT is not allowed to be zero.

To debug this problem we start gdb, using the file ‘.gdbinit’ to define a breakpoint in

gsl_error,
     bash$ gdb test_trap
     GDB is free software and you are welcome to distribute copies
     of it under certain conditions; type "show copying" to see
     the conditions.  There is absolutely no warranty for GDB;
     type "show warranty" for details.  GDB 4.16 (i586-debian-linux),
     Copyright 1996 Free Software Foundation, Inc.
     Breakpoint 1 at 0x8050b1e: file error.c, line 14.

When we run the program this breakpoint catches the error and shows the reason for it.

     (gdb) run
     Starting program: test_trap
     Breakpoint 1, gsl_error (reason=0x8052b0d
         "length n must be positive integer",
         file=0x8052b04 "c_init.c", line=108, gsl_errno=1)
         at error.c:14
     14        if (gsl_error_handler)

The first argument of gsl_error is always a string describing the error. Now we can look at the backtrace to see what caused the problem,

(gdb) bt
#0  gsl_error (reason=0x8052b0d
    "length n must be positive integer",
    file=0x8052b04 "c_init.c", line=108, gsl_errno=1)
    at error.c:14
#1  0x8049376 in gsl_fft_complex_wavetable_alloc (n=0,
    wavetable=0xbffff778) at c_init.c:108
#2  0x8048a00 in main (argc=1, argv=0xbffff9bc)
    at test_trap.c:94
#3  0x80488be in ___crt_dummy__()

We can see that the error was generated in the function gsl_fft_complex_wavetable_ alloc when it was called with an argument of n=0. The original call came from line 94 in

the file ‘test_trap.c’.

By moving up to the level of the original call we can find the line that caused the error,

(gdb) up
#1  0x8049376 in gsl_fft_complex_wavetable_alloc (n=0,
    wavetable=0xbffff778) at c_init.c:108
108   GSL_ERROR ("length n must be positive integer", GSL_EDOM);
(gdb) up
#2  0x8048a00 in main (argc=1, argv=0xbffff9bc)
    at test_trap.c:94
94    status = gsl_fft_complex_wavetable_alloc (0,
        &complex_wavetable);

Thus we have found the line that caused the problem. From this point we could also print out the values of other variables such as complex_wavetable.

A.2 Examining floating point registers

The contents of floating point registers can be examined using the command info float (not available on all platforms).

(gdb) info float
     st0: 0xc4018b895aa17a945000  Valid Normal -7.838871e+308
     st1: 0x3ff9ea3f50e4d7275000  Valid Normal 0.0285946
     st2: 0x3fe790c64ce27dad4800  Valid Normal 6.7415931e-08
     st3: 0x3ffaa3ef0df6607d7800  Spec  Normal 0.0400229
     st4: 0x3c028000000000000000  Valid Normal 4.4501477e-308
     st5: 0x3ffef5412c22219d9000  Zero  Normal 0.9580257
     st6: 0x3fff8000000000000000  Valid Normal 1
     st7: 0xc4028b65a1f6d243c800  Valid Normal -1.566206e+309
   fctrl: 0x0272 53 bit; NEAR; mask DENOR UNDER LOS;
   fstat: 0xb9ba flags 0001; top 7; excep DENOR OVERF UNDER LOS
    ftag: 0x3fff
     fip: 0x08048b5c
     fcs: 0x051a0023
  fopoff: 0x08086820
  fopsel: 0x002b

Individual registers can be examined using the variables $reg, where reg is the register name.

(gdb) p $st1
$1 = 0.02859464454261210347719

A.3 Handling floating point exceptions

It is possible to stop the program whenever a SIGFPE floating point exception occurs. This can be useful for finding the cause of an unexpected infinity or NaN. The current handler settings can be shown with the command info signal SIGFPE.

     (gdb) info signal SIGFPE
     Signal  Stop  Print  Pass to program Description
     SIGFPE  Yes   Yes    Yes             Arithmetic exception

Unless the program uses a signal handler the default setting should be changed so that SIGFPE is not passed to the program, as this would cause it to exit. The command handle

SIGFPE stop nopass prevents this.
     (gdb) handle SIGFPE stop nopass
     Signal  Stop  Print  Pass to program Description
     SIGFPE  Yes   Yes    No              Arithmetic exception

Depending on the platform it may be necessary to instruct the kernel to generate signals for floating point exceptions. For programs using GSL this can be achieved using the GSL_ IEEE_MODE environment variable in conjunction with the function gsl_ieee_env_setup() as described in see Chapter 37 [IEEE floating-point arithmetic], page 366.

     (gdb) set env GSL_IEEE_MODE=double-precision

A.4 GCC warning options for numerical programs

Writing reliable numerical programs in C requires great care. The following GCC warning options are recommended when compiling numerical programs:

     gcc -ansi -pedantic -Werror -Wall -W
       -Wmissing-prototypes -Wstrict-prototypes
       -Wtraditional -Wconversion -Wshadow
       -Wpointer-arith -Wcast-qual -Wcast-align
       -Wwrite-strings -Wnested-externs
       -fshort-enums -fno-common -Dinline= -g -O4

For details of each option consult the manual Using and Porting GCC. The following table gives a brief explanation of what types of errors these options catch.

-ansi -pedantic

Use ANSI C, and reject any non-ANSI extensions. These flags help in writing portable programs that will compile on other systems.

-Werror Consider warnings to be errors, so that compilation stops. This prevents warnings from scrolling off the top of the screen and being lost. You won’t be able to compile the program until it is completely warning-free.

-Wall This turns on a set of warnings for common programming problems. You need -Wall, but it is not enough on its own.

-O4 Turn on optimization. The warnings for uninitialized variables in -Wall rely on the optimizer to analyze the code. If there is no optimization then the warnings aren’t generated.

-W This turns on some extra warnings not included in -Wall, such as missing return values and comparisons between signed and unsigned integers.

-Wmissing-prototypes -Wstrict-prototypes

Warn if there are any missing or inconsistent prototypes. Without prototypes it is harder to detect problems with incorrect arguments.

-Wtraditional

This warns about certain constructs that behave differently in traditional and ANSI C. Whether the traditional or ANSI interpretation is used might be unpredictable on other compilers.

-Wconversion

The main use of this option is to warn about conversions from signed to unsigned integers. For example, unsigned int x = -1. If you need to perform such a conversion you can use an explicit cast.

-Wshadow This warns whenever a local variable shadows another local variable. If two variables have the same name then it is a potential source of confusion.

-Wpointer-arith -Wcast-qual -Wcast-align

These options warn if you try to do pointer arithmetic for types which don’t have a size, such as void, if you remove a const cast from a pointer, or if you cast a pointer to a type which has a different size, causing an invalid alignment.

-Wwrite-strings

This option gives string constants a const qualifier so that it will be a compiletime error to attempt to overwrite them.

-fshort-enums

This option makes the type of enum as short as possible. Normally this makes an enum different from an int. Consequently any attempts to assign a pointerto-int to a pointer-to-enum will generate a cast-alignment warning.

-fno-common

This option prevents global variables being simultaneously defined in different object files (you get an error at link time). Such a variable should be defined in one file and referred to in other files with an extern declaration.

-Wnested-externs

This warns if an extern declaration is encountered within an function.

-Dinline=

The inline keyword is not part of ANSI C. Thus if you want to use -ansi with a program which uses inline functions you can use this preprocessor definition to remove the inline keywords.

-g It always makes sense to put debugging symbols in the executable so that you can debug it using gdb. The only effect of debugging symbols is to increase the size of the file, and you can use the strip command to remove them later if necessary.

A.5 References and Further Reading

The following books are essential reading for anyone writing and debugging numerical programs with gcc and gdb.

R.M. Stallman, Using and Porting GNU CC, Free Software Foundation, ISBN 1882114388 R.M. Stallman, R.H. Pesch, Debugging with GDB: The GNU Source-Level Debugger, Free Software Foundation, ISBN 1882114779

Appendix B Contributors to GSL

(See the AUTHORS file in the distribution for up-to-date information.)

Mark Galassi Conceived GSL (with James Theiler) and wrote the design document. Wrote the simulated annealing package and the relevant chapter in the manual.

James Theiler Conceived GSL (with Mark Galassi). Wrote the random number generators and the relevant chapter in this manual.

Jim Davies Wrote the statistical routines and the relevant chapter in this manual.

Brian Gough FFTs, numerical integration, random number generators and distributions, root finding, minimization and fitting, polynomial solvers, complex numbers, physical constants, permutations, vector and matrix functions, histograms, statistics, ieee-utils, revised cblas Level 2 & 3, matrix decompositions and eigensystems.

Reid Priedhorsky Wrote and documented the initial version of the root finding routines while at Los Alamos National Laboratory, Mathematical Modeling and Analysis Group.

Gerard Jungman Series acceleration, ODEs, BLAS, Linear Algebra, Eigensystems, Hankel Transforms.

Mike Booth Wrote the Monte Carlo library.

Jorma Olavi Tähtinen Wrote the initial complex arithmetic functions.

Thomas Walter Wrote the initial heapsort routines and cholesky decomposition.

Fabrice Rossi Multidimensional minimization.

Appendix C Autoconf Macros

The following autoconf test will check for extern inline,

dnl Check for "extern inline", using a modified version
dnl of the test for AC_C_INLINE from acspecific.mt
dnl
AC_CACHE_CHECK([for extern inline], ac_cv_c_extern_inline,
[ac_cv_c_extern_inline=no
AC_TRY_COMPILE([extern $ac_cv_c_inline double foo(double x);
extern $ac_cv_c_inline double foo(double x) { return x+1.0; };
double foo (double x) { return x + 1.0; };],
[  foo(1.0)  ],
[ac_cv_c_extern_inline="yes"])
])
if test "$ac_cv_c_extern_inline" != no ; then
  AC_DEFINE(HAVE_INLINE,1)
  AC_SUBST(HAVE_INLINE)
fi

Appendix D: GSL CBLAS Library 378

Appendix D GSL CBLAS Library

The prototypes for the low-level cblas functions are declared in the file gsl_cblas.h. For the definition of the functions consult the documentation available from Netlib (see Section 11.3 [BLAS References and Further Reading], page 117).

D.1 Level 1

Function float cblas sdsdot (const int N, const float alpha, const float
         *x, const int incx, const float *y, const int incy)
Function double cblas dsdot (const int N, const float *x, const int incx,
         const float *y, const int incy)
Function float cblas sdot (const int N, const float *x, const int incx,
         const float *y, const int incy)
Function double cblas ddot (const int N, const double *x, const int incx,
         const double *y, const int incy)
Function void cblas cdotu sub (const int N, const void *x, const int
         incx, const void *y, const int incy, void *dotu)
Function void cblas cdotc sub (const int N, const void *x, const int
         incx, const void *y, const int incy, void *dotc)
Function void cblas zdotu sub (const int N, const void *x, const int
         incx, const void *y, const int incy, void *dotu)
Function void cblas zdotc sub (const int N, const void *x, const int
         incx, const void *y, const int incy, void *dotc)
Function float cblas snrm2 (const int N, const float *x, const int incx)
Function float cblas sasum (const int N, const float *x, const int incx)
Function double cblas dnrm2 (const int N, const double *x, const int

incx)

Function double cblas dasum (const int N, const double *x, const int

incx)

Function float cblas scnrm2 (const int N, const void *x, const int incx)
Function float cblas scasum (const int N, const void *x, const int incx)
Function double cblas dznrm2 (const int N, const void *x, const int incx)

Function double cblas dzasum (const int N, const void *x, const int incx)

Function CBLAS_INDEX cblas isamax (const int N, const float *x, const

int incx)

Function CBLAS_INDEX cblas idamax (const int N, const double *x, const

int incx)

Function CBLAS_INDEX cblas icamax (const int N, const void *x, const int

incx)

Function CBLAS_INDEX cblas izamax (const int N, const void *x, const int

incx)

Function void cblas sswap (const int N, float *x, const int incx, float
         *y, const int incy)
Function void cblas scopy (const int N, const float *x, const int incx,
         float *y, const int incy)
Function void cblas saxpy (const int N, const float alpha, const float *x,
         const int incx, float *y, const int incy)
Function void cblas dswap (const int N, double *x, const int incx, double
         *y, const int incy)
Function void cblas dcopy (const int N, const double *x, const int incx,
         double *y, const int incy)
Function void cblas daxpy (const int N, const double alpha, const double
         *x, const int incx, double *y, const int incy)
Function void cblas cswap (const int N, void *x, const int incx, void *y,
         const int incy)
Function void cblas ccopy (const int N, const void *x, const int incx,
         void *y, const int incy)
Function void cblas caxpy (const int N, const void *alpha, const void *x,
         const int incx, void *y, const int incy)
Function void cblas zswap (const int N, void *x, const int incx, void *y,
         const int incy)
Function void cblas zcopy (const int N, const void *x, const int incx,
         void *y, const int incy)
Function void cblas zaxpy (const int N, const void *alpha, const void *x,
         const int incx, void *y, const int incy)

Appendix D: GSL CBLAS Library 380

Function void cblas srotg (float *a, float *b, float *c, float *s)
Function void cblas srotmg (float *d1, float *d2, float *b1, const float
         b2, float *P)
Function void cblas srot (const int N, float *x, const int incx, float *y,
         const int incy, const float c, const float s)
Function void cblas srotm (const int N, float *x, const int incx, float
         *y, const int incy, const float *P)
Function void cblas drotg (double *a, double *b, double *c, double *s)
Function void cblas drotmg (double *d1, double *d2, double *b1, const
         double b2, double *P)
Function void cblas drot (const int N, double *x, const int incx, double
         *y, const int incy, const double c, const double s)
Function void cblas drotm (const int N, double *x, const int incx, double
         *y, const int incy, const double *P)
Function void cblas sscal (const int N, const float alpha, float *x, const

int incx)

Function void cblas dscal (const int N, const double alpha, double *x,
         const int incx)
Function void cblas cscal (const int N, const void *alpha, void *x, const

int incx)

Function void cblas zscal (const int N, const void *alpha, void *x, const

int incx)

Function void cblas csscal (const int N, const float alpha, void *x, const

int incx)

Function void cblas zdscal (const int N, const double alpha, void *x,
         const int incx)

D.2 Level 2

Function void cblas sgemv (const enum CBLAS_ORDER order, const enum
CBLAS_TRANSPOSE TransA, const int M, const int N, const float alpha,
const float *A, const int lda, const float *x, const int incx, const
float beta, float *y, const int incy)
Function void cblas sgbmv (const enum CBLAS_ORDER order, const enum
CBLAS_TRANSPOSE TransA, const int M, const int N, const int KL,
const int KU, const float alpha, const float *A, const int lda, const
float *x, const int incx, const float beta, float *y, const int incy)
Function void cblas strmv (const enum CBLAS_ORDER order, const enum
CBLAS_UPLO Uplo, const enum CBLAS_TRANSPOSE TransA, const enum
CBLAS_DIAG Diag, const int N, const float *A, const int lda, float *x,
const int incx)
Function void cblas stbmv (const enum CBLAS_ORDER order, const enum
CBLAS_UPLO Uplo, const enum CBLAS_TRANSPOSE TransA, const enum
CBLAS_DIAG Diag, const int N, const int K, const float *A, const int
lda, float *x, const int incx)
Function void cblas stpmv (const enum CBLAS_ORDER order, const enum
CBLAS_UPLO Uplo, const enum CBLAS_TRANSPOSE TransA, const enum
CBLAS_DIAG Diag, const int N, const float *Ap, float *x, const int

incx)

Function void cblas strsv (const enum CBLAS_ORDER order, const enum
CBLAS_UPLO Uplo, const enum CBLAS_TRANSPOSE TransA, const enum
CBLAS_DIAG Diag, const int N, const float *A, const int lda, float *x,
const int incx)
Function void cblas stbsv (const enum CBLAS_ORDER order, const enum
CBLAS_UPLO Uplo, const enum CBLAS_TRANSPOSE TransA, const enum
CBLAS_DIAG Diag, const int N, const int K, const float *A, const int
lda, float *x, const int incx)
Function void cblas stpsv (const enum CBLAS_ORDER order, const enum
CBLAS_UPLO Uplo, const enum CBLAS_TRANSPOSE TransA, const enum
CBLAS_DIAG Diag, const int N, const float *Ap, float *x, const int

incx)

Function void cblas dgemv (const enum CBLAS_ORDER order, const enum
CBLAS_TRANSPOSE TransA, const int M, const int N, const double alpha,
const double *A, const int lda, const double *x, const int incx, const
double beta, double *y, const int incy)

Function void cblas dgbmv (const enum CBLAS_ORDER order, const enum
CBLAS_TRANSPOSE TransA, const int M, const int N, const int KL,
const int KU, const double alpha, const double *A, const int lda,
const double *x, const int incx, const double beta, double *y, const int

incy)

Function void cblas dtrmv (const enum CBLAS_ORDER order, const enum
CBLAS_UPLO Uplo, const enum CBLAS_TRANSPOSE TransA, const enum
CBLAS_DIAG Diag, const int N, const double *A, const int lda, double
*x, const int incx)
Function void cblas dtbmv (const enum CBLAS_ORDER order, const enum
CBLAS_UPLO Uplo, const enum CBLAS_TRANSPOSE TransA, const enum
CBLAS_DIAG Diag, const int N, const int K, const double *A, const int
lda, double *x, const int incx)
Function void cblas dtpmv (const enum CBLAS_ORDER order, const enum
CBLAS_UPLO Uplo, const enum CBLAS_TRANSPOSE TransA, const enum
CBLAS_DIAG Diag, const int N, const double *Ap, double *x, const int

incx)

Function void cblas dtrsv (const enum CBLAS_ORDER order, const enum
CBLAS_UPLO Uplo, const enum CBLAS_TRANSPOSE TransA, const enum
CBLAS_DIAG Diag, const int N, const double *A, const int lda, double
*x, const int incx)
Function void cblas dtbsv (const enum CBLAS_ORDER order, const enum
CBLAS_UPLO Uplo, const enum CBLAS_TRANSPOSE TransA, const enum
CBLAS_DIAG Diag, const int N, const int K, const double *A, const int
lda, double *x, const int incx)
Function void cblas dtpsv (const enum CBLAS_ORDER order, const enum
CBLAS_UPLO Uplo, const enum CBLAS_TRANSPOSE TransA, const enum
CBLAS_DIAG Diag, const int N, const double *Ap, double *x, const int

incx)

Function void cblas cgemv (const enum CBLAS_ORDER order, const enum
CBLAS_TRANSPOSE TransA, const int M, const int N, const void *alpha,
const void *A, const int lda, const void *x, const int incx, const void
*beta, void *y, const int incy)
Function void cblas cgbmv (const enum CBLAS_ORDER order, const enum
CBLAS_TRANSPOSE TransA, const int M, const int N, const int KL,
const int KU, const void *alpha, const void *A, const int lda, const
void *x, const int incx, const void *beta, void *y, const int incy)
Function void cblas ctrmv (const enum CBLAS_ORDER order, const enum
CBLAS_UPLO Uplo, const enum CBLAS_TRANSPOSE TransA, const enum
CBLAS_DIAG Diag, const int N, const void *A, const int lda, void *x,
const int incx)

Function void cblas ctbmv (const enum CBLAS_ORDER order, const enum
CBLAS_UPLO Uplo, const enum CBLAS_TRANSPOSE TransA, const enum
CBLAS_DIAG Diag, const int N, const int K, const void *A, const int
lda, void *x, const int incx)
Function void cblas ctpmv (const enum CBLAS_ORDER order, const enum
CBLAS_UPLO Uplo, const enum CBLAS_TRANSPOSE TransA, const enum
CBLAS_DIAG Diag, const int N, const void *Ap, void *x, const int incx)
Function void cblas ctrsv (const enum CBLAS_ORDER order, const enum
CBLAS_UPLO Uplo, const enum CBLAS_TRANSPOSE TransA, const enum
CBLAS_DIAG Diag, const int N, const void *A, const int lda, void *x,
const int incx)
Function void cblas ctbsv (const enum CBLAS_ORDER order, const enum
CBLAS_UPLO Uplo, const enum CBLAS_TRANSPOSE TransA, const enum
CBLAS_DIAG Diag, const int N, const int K, const void *A, const int
lda, void *x, const int incx)
Function void cblas ctpsv (const enum CBLAS_ORDER order, const enum
CBLAS_UPLO Uplo, const enum CBLAS_TRANSPOSE TransA, const enum
CBLAS_DIAG Diag, const int N, const void *Ap, void *x, const int incx)
Function void cblas zgemv (const enum CBLAS_ORDER order, const enum
CBLAS_TRANSPOSE TransA, const int M, const int N, const void *alpha,
const void *A, const int lda, const void *x, const int incx, const void
*beta, void *y, const int incy)
Function void cblas zgbmv (const enum CBLAS_ORDER order, const enum
CBLAS_TRANSPOSE TransA, const int M, const int N, const int KL,
const int KU, const void *alpha, const void *A, const int lda, const
void *x, const int incx, const void *beta, void *y, const int incy)
Function void cblas ztrmv (const enum CBLAS_ORDER order, const enum
CBLAS_UPLO Uplo, const enum CBLAS_TRANSPOSE TransA, const enum
CBLAS_DIAG Diag, const int N, const void *A, const int lda, void *x,
const int incx)
Function void cblas ztbmv (const enum CBLAS_ORDER order, const enum
CBLAS_UPLO Uplo, const enum CBLAS_TRANSPOSE TransA, const enum
CBLAS_DIAG Diag, const int N, const int K, const void *A, const int
lda, void *x, const int incx)
Function void cblas ztpmv (const enum CBLAS_ORDER order, const enum
CBLAS_UPLO Uplo, const enum CBLAS_TRANSPOSE TransA, const enum
CBLAS_DIAG Diag, const int N, const void *Ap, void *x, const int incx)
Function void cblas ztrsv (const enum CBLAS_ORDER order, const enum
CBLAS_UPLO Uplo, const enum CBLAS_TRANSPOSE TransA, const enum
CBLAS_DIAG Diag, const int N, const void *A, const int lda, void *x,
const int incx)

Function void cblas ztbsv (const enum CBLAS_ORDER order, const enum
CBLAS_UPLO Uplo, const enum CBLAS_TRANSPOSE TransA, const enum
CBLAS_DIAG Diag, const int N, const int K, const void *A, const int
lda, void *x, const int incx)
Function void cblas ztpsv (const enum CBLAS_ORDER order, const enum
CBLAS_UPLO Uplo, const enum CBLAS_TRANSPOSE TransA, const enum
CBLAS_DIAG Diag, const int N, const void *Ap, void *x, const int incx)
Function void cblas ssymv (const enum CBLAS_ORDER order, const enum
CBLAS_UPLO Uplo, const int N, const float alpha, const float *A, const
int lda, const float *x, const int incx, const float beta, float *y,
const int incy)
Function void cblas ssbmv (const enum CBLAS_ORDER order, const enum
CBLAS_UPLO Uplo, const int N, const int K, const float alpha, const
float *A, const int lda, const float *x, const int incx, const float
beta, float *y, const int incy)
Function void cblas sspmv (const enum CBLAS_ORDER order, const enum
CBLAS_UPLO Uplo, const int N, const float alpha, const float *Ap,
const float *x, const int incx, const float beta, float *y, const int

incy)

Function void cblas sger (const enum CBLAS_ORDER order, const int M, const
int N, const float alpha, const float *x, const int incx, const float
*y, const int incy, float *A, const int lda)
Function void cblas ssyr (const enum CBLAS_ORDER order, const enum
CBLAS_UPLO Uplo, const int N, const float alpha, const float *x, const
int incx, float *A, const int lda)
Function void cblas sspr (const enum CBLAS_ORDER order, const enum
CBLAS_UPLO Uplo, const int N, const float alpha, const float *x, const
int incx, float *Ap)
Function void cblas ssyr2 (const enum CBLAS_ORDER order, const enum
CBLAS_UPLO Uplo, const int N, const float alpha, const float *x, const
int incx, const float *y, const int incy, float *A, const int lda)
Function void cblas sspr2 (const enum CBLAS_ORDER order, const enum
CBLAS_UPLO Uplo, const int N, const float alpha, const float *x, const
int incx, const float *y, const int incy, float *A)
Function void cblas dsymv (const enum CBLAS_ORDER order, const enum
CBLAS_UPLO Uplo, const int N, const double alpha, const double *A,
const int lda, const double *x, const int incx, const double beta,
double *y, const int incy)

Function void cblas dsbmv (const enum CBLAS_ORDER order, const enum
CBLAS_UPLO Uplo, const int N, const int K, const double alpha, const
double *A, const int lda, const double *x, const int incx, const double
beta, double *y, const int incy)
Function void cblas dspmv (const enum CBLAS_ORDER order, const enum
CBLAS_UPLO Uplo, const int N, const double alpha, const double *Ap,
const double *x, const int incx, const double beta, double *y, const int

incy)

Function void cblas dger (const enum CBLAS_ORDER order, const int M,
const int N, const double alpha, const double *x, const int incx, const
double *y, const int incy, double *A, const int lda)
Function void cblas dsyr (const enum CBLAS_ORDER order, const enum
CBLAS_UPLO Uplo, const int N, const double alpha, const double *x,
const int incx, double *A, const int lda)
Function void cblas dspr (const enum CBLAS_ORDER order, const enum
CBLAS_UPLO Uplo, const int N, const double alpha, const double *x,
const int incx, double *Ap)
Function void cblas dsyr2 (const enum CBLAS_ORDER order, const enum
CBLAS_UPLO Uplo, const int N, const double alpha, const double *x,
const int incx, const double *y, const int incy, double *A, const int

lda)

Function void cblas dspr2 (const enum CBLAS_ORDER order, const enum
CBLAS_UPLO Uplo, const int N, const double alpha, const double *x,
const int incx, const double *y, const int incy, double *A)
Function void cblas chemv (const enum CBLAS_ORDER order, const enum
CBLAS_UPLO Uplo, const int N, const void *alpha, const void *A, const
int lda, const void *x, const int incx, const void *beta, void *y, const

int incy)

Function void cblas chbmv (const enum CBLAS_ORDER order, const enum
CBLAS_UPLO Uplo, const int N, const int K, const void *alpha, const
void *A, const int lda, const void *x, const int incx, const void *beta,
void *y, const int incy)
Function void cblas chpmv (const enum CBLAS_ORDER order, const enum
CBLAS_UPLO Uplo, const int N, const void *alpha, const void *Ap, const
void *x, const int incx, const void *beta, void *y, const int incy)
Function void cblas cgeru (const enum CBLAS_ORDER order, const int M,
const int N, const void *alpha, const void *x, const int incx, const
void *y, const int incy, void *A, const int lda)

Function void cblas cgerc (const enum CBLAS_ORDER order, const int M,
const int N, const void *alpha, const void *x, const int incx, const
void *y, const int incy, void *A, const int lda)
Function void cblas cher (const enum CBLAS_ORDER order, const enum
CBLAS_UPLO Uplo, const int N, const float alpha, const void *x, const
int incx, void *A, const int lda)
Function void cblas chpr (const enum CBLAS_ORDER order, const enum
CBLAS_UPLO Uplo, const int N, const float alpha, const void *x, const
int incx, void *A)
Function void cblas cher2 (const enum CBLAS_ORDER order, const enum
CBLAS_UPLO Uplo, const int N, const void *alpha, const void *x, const
int incx, const void *y, const int incy, void *A, const int lda)
Function void cblas chpr2 (const enum CBLAS_ORDER order, const enum
CBLAS_UPLO Uplo, const int N, const void *alpha, const void *x, const
int incx, const void *y, const int incy, void *Ap)
Function void cblas zhemv (const enum CBLAS_ORDER order, const enum
CBLAS_UPLO Uplo, const int N, const void *alpha, const void *A, const
int lda, const void *x, const int incx, const void *beta, void *y, const

int incy)

Function void cblas zhbmv (const enum CBLAS_ORDER order, const enum
CBLAS_UPLO Uplo, const int N, const int K, const void *alpha, const
void *A, const int lda, const void *x, const int incx, const void *beta,
void *y, const int incy)
Function void cblas zhpmv (const enum CBLAS_ORDER order, const enum
CBLAS_UPLO Uplo, const int N, const void *alpha, const void *Ap, const
void *x, const int incx, const void *beta, void *y, const int incy)
Function void cblas zgeru (const enum CBLAS_ORDER order, const int M,
const int N, const void *alpha, const void *x, const int incx, const
void *y, const int incy, void *A, const int lda)
Function void cblas zgerc (const enum CBLAS_ORDER order, const int M,
const int N, const void *alpha, const void *x, const int incx, const
void *y, const int incy, void *A, const int lda)
Function void cblas zher (const enum CBLAS_ORDER order, const enum
CBLAS_UPLO Uplo, const int N, const double alpha, const void *x, const
int incx, void *A, const int lda)
Function void cblas zhpr (const enum CBLAS_ORDER order, const enum
CBLAS_UPLO Uplo, const int N, const double alpha, const void *x, const
int incx, void *A)

Function void cblas zher2 (const enum CBLAS_ORDER order, const enum
CBLAS_UPLO Uplo, const int N, const void *alpha, const void *x, const
int incx, const void *y, const int incy, void *A, const int lda)
Function void cblas zhpr2 (const enum CBLAS_ORDER order, const enum
CBLAS_UPLO Uplo, const int N, const void *alpha, const void *x, const
int incx, const void *y, const int incy, void *Ap)

D.3 Level 3

Function void cblas sgemm (const enum CBLAS_ORDER Order, const enum
CBLAS_TRANSPOSE TransA, const enum CBLAS_TRANSPOSE TransB, const int
M, const int N, const int K, const float alpha, const float *A, const
int lda, const float *B, const int ldb, const float beta, float *C,
const int ldc)
Function void cblas ssymm (const enum CBLAS_ORDER Order, const enum
CBLAS_SIDE Side, const enum CBLAS_UPLO Uplo, const int M, const int N,
const float alpha, const float *A, const int lda, const float *B, const
int ldb, const float beta, float *C, const int ldc)
Function void cblas ssyrk (const enum CBLAS_ORDER Order, const enum
CBLAS_UPLO Uplo, const enum CBLAS_TRANSPOSE Trans, const int N, const
int K, const float alpha, const float *A, const int lda, const float
beta, float *C, const int ldc)
Function void cblas ssyr2k (const enum CBLAS_ORDER Order, const enum
CBLAS_UPLO Uplo, const enum CBLAS_TRANSPOSE Trans, const int N, const
int K, const float alpha, const float *A, const int lda, const float
*B, const int ldb, const float beta, float *C, const int ldc)
Function void cblas strmm (const enum CBLAS_ORDER Order, const enum
CBLAS_SIDE Side, const enum CBLAS_UPLO Uplo, const enum
CBLAS_TRANSPOSE TransA, const enum CBLAS_DIAG Diag, const int M,
const int N, const float alpha, const float *A, const int lda, float
*B, const int ldb)
Function void cblas strsm (const enum CBLAS_ORDER Order, const enum
CBLAS_SIDE Side, const enum CBLAS_UPLO Uplo, const enum
CBLAS_TRANSPOSE TransA, const enum CBLAS_DIAG Diag, const int M,
const int N, const float alpha, const float *A, const int lda, float
*B, const int ldb)
Function void cblas dgemm (const enum CBLAS_ORDER Order, const enum
CBLAS_TRANSPOSE TransA, const enum CBLAS_TRANSPOSE TransB, const int
M, const int N, const int K, const double alpha, const double *A,
const int lda, const double *B, const int ldb, const double beta,
double *C, const int ldc)

Function void cblas dsymm (const enum CBLAS_ORDER Order, const enum
CBLAS_SIDE Side, const enum CBLAS_UPLO Uplo, const int M, const int N,
const double alpha, const double *A, const int lda, const double *B,
const int ldb, const double beta, double *C, const int ldc)
Function void cblas dsyrk (const enum CBLAS_ORDER Order, const enum
CBLAS_UPLO Uplo, const enum CBLAS_TRANSPOSE Trans, const int N, const
int K, const double alpha, const double *A, const int lda, const double
beta, double *C, const int ldc)
Function void cblas dsyr2k (const enum CBLAS_ORDER Order, const enum
CBLAS_UPLO Uplo, const enum CBLAS_TRANSPOSE Trans, const int N, const
int K, const double alpha, const double *A, const int lda, const double
*B, const int ldb, const double beta, double *C, const int ldc)
Function void cblas dtrmm (const enum CBLAS_ORDER Order, const enum
CBLAS_SIDE Side, const enum CBLAS_UPLO Uplo, const enum
CBLAS_TRANSPOSE TransA, const enum CBLAS_DIAG Diag, const int M,
const int N, const double alpha, const double *A, const int lda, double
*B, const int ldb)
Function void cblas dtrsm (const enum CBLAS_ORDER Order, const enum
CBLAS_SIDE Side, const enum CBLAS_UPLO Uplo, const enum
CBLAS_TRANSPOSE TransA, const enum CBLAS_DIAG Diag, const int M,
const int N, const double alpha, const double *A, const int lda, double
*B, const int ldb)
Function void cblas cgemm (const enum CBLAS_ORDER Order, const enum
CBLAS_TRANSPOSE TransA, const enum CBLAS_TRANSPOSE TransB, const int
M, const int N, const int K, const void *alpha, const void *A, const
int lda, const void *B, const int ldb, const void *beta, void *C, const

int ldc)

Function void cblas csymm (const enum CBLAS_ORDER Order, const enum
CBLAS_SIDE Side, const enum CBLAS_UPLO Uplo, const int M, const int N,
const void *alpha, const void *A, const int lda, const void *B, const
int ldb, const void *beta, void *C, const int ldc)
Function void cblas csyrk (const enum CBLAS_ORDER Order, const enum
CBLAS_UPLO Uplo, const enum CBLAS_TRANSPOSE Trans, const int N, const
int K, const void *alpha, const void *A, const int lda, const void
*beta, void *C, const int ldc)
Function void cblas csyr2k (const enum CBLAS_ORDER Order, const enum
CBLAS_UPLO Uplo, const enum CBLAS_TRANSPOSE Trans, const int N, const
int K, const void *alpha, const void *A, const int lda, const void *B,
const int ldb, const void *beta, void *C, const int ldc)

Function void cblas ctrmm (const enum CBLAS_ORDER Order, const enum
CBLAS_SIDE Side, const enum CBLAS_UPLO Uplo, const enum
CBLAS_TRANSPOSE TransA, const enum CBLAS_DIAG Diag, const int M,
const int N, const void *alpha, const void *A, const int lda, void *B,
const int ldb)
Function void cblas ctrsm (const enum CBLAS_ORDER Order, const enum
CBLAS_SIDE Side, const enum CBLAS_UPLO Uplo, const enum
CBLAS_TRANSPOSE TransA, const enum CBLAS_DIAG Diag, const int M,
const int N, const void *alpha, const void *A, const int lda, void *B,
const int ldb)
Function void cblas zgemm (const enum CBLAS_ORDER Order, const enum
CBLAS_TRANSPOSE TransA, const enum CBLAS_TRANSPOSE TransB, const int
M, const int N, const int K, const void *alpha, const void *A, const
int lda, const void *B, const int ldb, const void *beta, void *C, const

int ldc)

Function void cblas zsymm (const enum CBLAS_ORDER Order, const enum
CBLAS_SIDE Side, const enum CBLAS_UPLO Uplo, const int M, const int N,
const void *alpha, const void *A, const int lda, const void *B, const
int ldb, const void *beta, void *C, const int ldc)
Function void cblas zsyrk (const enum CBLAS_ORDER Order, const enum
CBLAS_UPLO Uplo, const enum CBLAS_TRANSPOSE Trans, const int N, const
int K, const void *alpha, const void *A, const int lda, const void
*beta, void *C, const int ldc)
Function void cblas zsyr2k (const enum CBLAS_ORDER Order, const enum
CBLAS_UPLO Uplo, const enum CBLAS_TRANSPOSE Trans, const int N, const
int K, const void *alpha, const void *A, const int lda, const void *B,
const int ldb, const void *beta, void *C, const int ldc)
Function void cblas ztrmm (const enum CBLAS_ORDER Order, const enum
CBLAS_SIDE Side, const enum CBLAS_UPLO Uplo, const enum
CBLAS_TRANSPOSE TransA, const enum CBLAS_DIAG Diag, const int M,
const int N, const void *alpha, const void *A, const int lda, void *B,
const int ldb)
Function void cblas ztrsm (const enum CBLAS_ORDER Order, const enum
CBLAS_SIDE Side, const enum CBLAS_UPLO Uplo, const enum
CBLAS_TRANSPOSE TransA, const enum CBLAS_DIAG Diag, const int M,
const int N, const void *alpha, const void *A, const int lda, void *B,
const int ldb)
Function void cblas chemm (const enum CBLAS_ORDER Order, const enum
CBLAS_SIDE Side, const enum CBLAS_UPLO Uplo, const int M, const int N,
const void *alpha, const void *A, const int lda, const void *B, const
int ldb, const void *beta, void *C, const int ldc)

Function void cblas cherk (const enum CBLAS_ORDER Order, const enum
CBLAS_UPLO Uplo, const enum CBLAS_TRANSPOSE Trans, const int N, const
int K, const float alpha, const void *A, const int lda, const float
beta, void *C, const int ldc)
Function void cblas cher2k (const enum CBLAS_ORDER Order, const enum
CBLAS_UPLO Uplo, const enum CBLAS_TRANSPOSE Trans, const int N, const
int K, const void *alpha, const void *A, const int lda, const void *B,
const int ldb, const float beta, void *C, const int ldc)
Function void cblas zhemm (const enum CBLAS_ORDER Order, const enum
CBLAS_SIDE Side, const enum CBLAS_UPLO Uplo, const int M, const int N,
const void *alpha, const void *A, const int lda, const void *B, const
int ldb, const void *beta, void *C, const int ldc)
Function void cblas zherk (const enum CBLAS_ORDER Order, const enum
CBLAS_UPLO Uplo, const enum CBLAS_TRANSPOSE Trans, const int N, const
int K, const double alpha, const void *A, const int lda, const double
beta, void *C, const int ldc)
Function void cblas zher2k (const enum CBLAS_ORDER Order, const enum
CBLAS_UPLO Uplo, const enum CBLAS_TRANSPOSE Trans, const int N, const
int K, const void *alpha, const void *A, const int lda, const void *B,
const int ldb, const double beta, void *C, const int ldc)
Function void cblas xerbla (int p, const char *rout, const char *form, ...)

D.4 Examples

The following program computes the product of two matrices using the Level-3 blas function sgemm,

( \begin{matrix}0.11 & 0.12 & 0.13 \\ 0.21 & 0.22 & 0.23 \end{matrix} ) ( \begin{matrix}1011 & 1012 \\ 1021 & 1022 \\ 1031 & 1031 \end{matrix} ) = ( \begin{matrix}367.76 & 368.12 \\ 674.06 & 674.72 \end{matrix} )

The matrices are stored in row major order but could be stored in column major order if the first argument of the call to cblas_sgemm was changed to CblasColMajor.

#include <stdio.h>
#include <gsl/gsl_cblas.h>
int
main (void)
{
int lda = 3;
float A[] = { 0.11, 0.12, 0.13,
            0.21, 0.22, 0.23 };
int ldb = 2;

float B[] = { 1011, 1012,
              1021, 1022,
              1031, 1032 };
int ldc = 2;
float C[] = { 0.00, 0.00,
              0.00, 0.00 };
/* Compute C = A B */
cblas_sgemm (CblasRowMajor,
             CblasNoTrans, CblasNoTrans, 2, 2, 3,
             1.0, A, lda, B, ldb, 0.0, C, ldc);
printf("[ %g, %g\n", C[0], C[1]);
printf("  %g, %g ]\n", C[2], C[3]);
return 0;
}

To compile the program use the following command line,

gcc demo.c -lgslcblas

There is no need to link with the main library -lgsl in this case as the cblas library is an independent unit. Here is the output from the program,

$ ./a.out
[ 367.76, 368.12
674.06, 674.72 ]

GNU General Public License

Version 2, June 1991

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1989, 1991 Free Software Foundation, Inc. 59 Temple Place - Suite 330, Boston, MA 02111-1307, USA

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

Preamble

The licenses for most software are designed to take away your freedom to share and change it. By contrast, the GNU General Public License is intended to guarantee your freedom to share and change free software—to make sure the software is free for all its users. This General Public License applies to most of the Free Software Foundation’s software and to any other program whose authors commit to using it. (Some other Free Software Foundation software is covered by the GNU Library General Public License instead.) You can apply it to your programs, too.

When we speak of free software, we are referring to freedom, not price. Our General Public Licenses are designed to make sure that you have the freedom to distribute copies of free software (and charge for this service if you wish), that you receive source code or can get it if you want it, that you can change the software or use pieces of it in new free programs; and that you know you can do these things.

To protect your rights, we need to make restrictions that forbid anyone to deny you these rights or to ask you to surrender the rights. These restrictions translate to certain responsibilities for you if you distribute copies of the software, or if you modify it.

For example, if you distribute copies of such a program, whether gratis or for a fee, you must give the recipients all the rights that you have. You must make sure that they, too, receive or can get the source code. And you must show them these terms so they know their rights.

We protect your rights with two steps: (1) copyright the software, and (2) offer you this license which gives you legal permission to copy, distribute and/or modify the software.

Also, for each author’s protection and ours, we want to make certain that everyone understands that there is no warranty for this free software. If the software is modified by someone else and passed on, we want its recipients to know that what they have is not the original, so that any problems introduced by others will not reflect on the original authors’ reputations.

Finally, any free program is threatened constantly by software patents. We wish to avoid the danger that redistributors of a free program will individually obtain patent licenses, in effect making the program proprietary. To prevent this, we have made it clear that any patent must be licensed for everyone’s free use or not licensed at all.

The precise terms and conditions for copying, distribution and modification follow.

TERMS AND CONDITIONS FOR COPYING, DISTRIBUTION AND MODIFICATION

0. This License applies to any program or other work which contains a notice placed by the copyright holder saying it may be distributed under the terms of this General Public License. The “Program”, below, refers to any such program or work, and a “work based on the Program” means either the Program or any derivative work under copyright law: that is to say, a work containing the Program or a portion of it, either verbatim or with modifications and/or translated into another language. (Hereinafter, translation is included without limitation in the term “modification”.) Each licensee is addressed as “you”. Activities other than copying, distribution and modification are not covered by this License; they are outside its scope. The act of running the Program is not restricted, and the output from the Program is covered only if its contents constitute a work based on the Program (independent of having been made by running the Program). Whether that is true depends on what the Program does.

1. You may copy and distribute verbatim copies of the Program’s source code as you receive it, in any medium, provided that you conspicuously and appropriately publish on each copy an appropriate copyright notice and disclaimer of warranty; keep intact all the notices that refer to this License and to the absence of any warranty; and give any other recipients of the Program a copy of this License along with the Program. You may charge a fee for the physical act of transferring a copy, and you may at your option offer warranty protection in exchange for a fee.

2. You may modify your copy or copies of the Program or any portion of it, thus forming a work based on the Program, and copy and distribute such modifications or work under the terms of Section 1 above, provided that you also meet all of these conditions: a. You must cause the modified files to carry prominent notices stating that you

changed the files and the date of any change. b. You must cause any work that you distribute or publish, that in whole or in part

contains or is derived from the Program or any part thereof, to be licensed as a whole at no charge to all third parties under the terms of this License.

c. If the modified program normally reads commands interactively when run, you must cause it, when started running for such interactive use in the most ordinary way, to print or display an announcement including an appropriate copyright notice and a notice that there is no warranty (or else, saying that you provide a warranty) and that users may redistribute the program under these conditions, and telling the user how to view a copy of this License. (Exception: if the Program itself is interactive but does not normally print such an announcement, your work based on the Program is not required to print an announcement.)

These requirements apply to the modified work as a whole. If identifiable sections of that work are not derived from the Program, and can be reasonably considered independent and separate works in themselves, then this License, and its terms, do not apply to those sections when you distribute them as separate works. But when you distribute the same sections as part of a whole which is a work based on the Program, the distribution of the whole must be on the terms of this License, whose permissions for other licensees extend to the entire whole, and thus to each and every part regardless of who wrote it.

Thus, it is not the intent of this section to claim rights or contest your rights to work written entirely by you; rather, the intent is to exercise the right to control the distribution of derivative or collective works based on the Program. In addition, mere aggregation of another work not based on the Program with the Program (or with a work based on the Program) on a volume of a storage or distribution medium does not bring the other work under the scope of this License.

3. You may copy and distribute the Program (or a work based on it, under Section 2) in object code or executable form under the terms of Sections 1 and 2 above provided that you also do one of the following: a. Accompany it with the complete corresponding machine-readable source code,

which must be distributed under the terms of Sections 1 and 2 above on a medium customarily used for software interchange; or,

b. Accompany it with a written offer, valid for at least three years, to give any third party, for a charge no more than your cost of physically performing source distribution, a complete machine-readable copy of the corresponding source code, to be distributed under the terms of Sections 1 and 2 above on a medium customarily used for software interchange; or,

c. Accompany it with the information you received as to the offer to distribute corresponding source code. (This alternative is allowed only for noncommercial distribution and only if you received the program in object code or executable form with such an offer, in accord with Subsection b above.)

The source code for a work means the preferred form of the work for making modifications to it. For an executable work, complete source code means all the source code for all modules it contains, plus any associated interface definition files, plus the scripts used to control compilation and installation of the executable. However, as a special exception, the source code distributed need not include anything that is normally distributed (in either source or binary form) with the major components (compiler, kernel, and so on) of the operating system on which the executable runs, unless that component itself accompanies the executable. If distribution of executable or object code is made by offering access to copy from a designated place, then offering equivalent access to copy the source code from the same place counts as distribution of the source code, even though third parties are not compelled to copy the source along with the object code.

4. You may not copy, modify, sublicense, or distribute the Program except as expressly provided under this License. Any attempt otherwise to copy, modify, sublicense or distribute the Program is void, and will automatically terminate your rights under this License. However, parties who have received copies, or rights, from you under this License will not have their licenses terminated so long as such parties remain in full compliance.

5. You are not required to accept this License, since you have not signed it. However, nothing else grants you permission to modify or distribute the Program or its derivative works. These actions are prohibited by law if you do not accept this License. Therefore, by modifying or distributing the Program (or any work based on the Program), you indicate your acceptance of this License to do so, and all its terms and conditions for copying, distributing or modifying the Program or works based on it.

6. Each time you redistribute the Program (or any work based on the Program), the recipient automatically receives a license from the original licensor to copy, distribute or modify the Program subject to these terms and conditions. You may not impose any further restrictions on the recipients’ exercise of the rights granted herein. You are not responsible for enforcing compliance by third parties to this License.

7. If, as a consequence of a court judgment or allegation of patent infringement or for any other reason (not limited to patent issues), conditions are imposed on you (whether by court order, agreement or otherwise) that contradict the conditions of this License, they do not excuse you from the conditions of this License. If you cannot distribute so as to satisfy simultaneously your obligations under this License and any other pertinent obligations, then as a consequence you may not distribute the Program at all. For example, if a patent license would not permit royalty-free redistribution of the Program by all those who receive copies directly or indirectly through you, then the only way you could satisfy both it and this License would be to refrain entirely from distribution of the Program. If any portion of this section is held invalid or unenforceable under any particular circumstance, the balance of the section is intended to apply and the section as a whole is intended to apply in other circumstances. It is not the purpose of this section to induce you to infringe any patents or other property right claims or to contest validity of any such claims; this section has the sole purpose of protecting the integrity of the free software distribution system, which is implemented by public license practices. Many people have made generous contributions to the wide range of software distributed through that system in reliance on consistent application of that system; it is up to the author/donor to decide if he or she is willing to distribute software through any other system and a licensee cannot impose that choice. This section is intended to make thoroughly clear what is believed to be a consequence of the rest of this License.

8. If the distribution and/or use of the Program is restricted in certain countries either by patents or by copyrighted interfaces, the original copyright holder who places the Program under this License may add an explicit geographical distribution limitation excluding those countries, so that distribution is permitted only in or among countries not thus excluded. In such case, this License incorporates the limitation as if written in the body of this License.

9. The Free Software Foundation may publish revised and/or new versions of the General Public License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns. Each version is given a distinguishing version number. If the Program specifies a version number of this License which applies to it and “any later version”, you have the option of following the terms and conditions either of that version or of any later version published by the Free Software Foundation. If the Program does not specify a version number of this License, you may choose any version ever published by the Free Software Foundation.

10. If you wish to incorporate parts of the Program into other free programs whose distribution conditions are different, write to the author to ask for permission. For software which is copyrighted by the Free Software Foundation, write to the Free Software Foundation; we sometimes make exceptions for this. Our decision will be guided by the two goals of preserving the free status of all derivatives of our free software and of promoting the sharing and reuse of software generally.

NO WARRANTY

11. BECAUSE THE PROGRAM IS LICENSED FREE OF CHARGE, THERE IS NO WARRANTY FOR THE PROGRAM, TO THE EXTENT PERMITTED BY APPLI-CABLE LAW. EXCEPT WHEN OTHERWISE STATED IN WRITING THE COPY-RIGHT HOLDERS AND/OR OTHER PARTIES PROVIDE THE PROGRAM “AS IS” WITHOUT WARRANTY OF ANY KIND, EITHER EXPRESSED OR IMPLIED, INCLUDING, BUT NOT LIMITED TO, THE IMPLIED WARRANTIES OF MER-CHANTABILITY AND FITNESS FOR A PARTICULAR PURPOSE. THE ENTIRE RISK AS TO THE QUALITY AND PERFORMANCE OF THE PROGRAM IS WITH YOU. SHOULD THE PROGRAM PROVE DEFECTIVE, YOU ASSUME THE COST OF ALL NECESSARY SERVICING, REPAIR OR CORRECTION.

12. IN NO EVENT UNLESS REQUIRED BY APPLICABLE LAW OR AGREED TO IN WRITING WILL ANY COPYRIGHT HOLDER, OR ANY OTHER PARTY WHO MAY MODIFY AND/OR REDISTRIBUTE THE PROGRAM AS PERMITTED ABOVE, BE LIABLE TO YOU FOR DAMAGES, INCLUDING ANY GENERAL, SPECIAL, INCIDENTAL OR CONSEQUENTIAL DAMAGES ARISING OUT OF THE USE OR INABILITY TO USE THE PROGRAM (INCLUDING BUT NOT LIMITED TO LOSS OF DATA OR DATA BEING RENDERED INACCURATE OR LOSSES SUSTAINED BY YOU OR THIRD PARTIES OR A FAILURE OF THE PROGRAM TO OPERATE WITH ANY OTHER PROGRAMS), EVEN IF SUCH HOLDER OR OTHER PARTY HAS BEEN ADVISED OF THE POSSIBILITY OF SUCH DAMAGES.

END OF TERMS AND CONDITIONS Appendix: How to Apply These Terms to Your New Programs

If you develop a new program, and you want it to be of the greatest possible use to the public, the best way to achieve this is to make it free software which everyone can redistribute and change under these terms.

To do so, attach the following notices to the program. It is safest to attach them to the start of each source file to most effectively convey the exclusion of warranty; and each file should have at least the “copyright” line and a pointer to where the full notice is found.

one line to give the program’s name and a brief idea of what it does. Copyright (C) yyyy name of author

This program is free software; you can redistribute it
and/or modify it under the terms of the GNU General Public
License as published by the Free Software Foundation; either
version 2 of the License, or (at your option) any later
version.
This program is distributed in the hope that it will be
useful, but WITHOUT ANY WARRANTY; without even the implied
warranty of MERCHANTABILITY or FITNESS FOR A PARTICULAR
PURPOSE.  See the GNU General Public License for more
details.
You should have received a copy of the GNU General Public
License along with this program; if not, write to the Free
Software Foundation, Inc., 59 Temple Place - Suite 330,
Boston, MA 02111-1307, USA.

Also add information on how to contact you by electronic and paper mail. If the program is interactive, make it output a short notice like this when it starts in an

interactive mode:

Gnomovision version 69, Copyright (C) 19yy name of author
Gnomovision comes with ABSOLUTELY NO WARRANTY; for details
type `show w'.  This is free software, and you are welcome
to redistribute it under certain conditions; type `show c'
for details.

The hypothetical commands ‘show w’ and ‘show c’ should show the appropriate parts of the General Public License. Of course, the commands you use may be called something other than ‘show w’ and ‘show c’; they could even be mouse-clicks or menu items—whatever suits your program.

You should also get your employer (if you work as a programmer) or your school, if any, to sign a “copyright disclaimer” for the program, if necessary. Here is a sample; alter the names:

Yoyodyne, Inc., hereby disclaims all copyright interest in
the program `Gnomovision' (which makes passes at compilers)
written by James Hacker.

signature of Ty Coon, 1 April 1989

     Ty Coon, President of Vice

This General Public License does not permit incorporating your program into proprietary programs. If your program is a subroutine library, you may consider it more useful to permit linking proprietary applications with the library. If this is what you want to do, use the GNU Library General Public License instead of this License.

GNU Free Documentation License

Version 1.1, March 2000

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2000 Free Software Foundation, Inc. 59 Temple Place, Suite 330, Boston, MA 02111-1307, USA

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

0. PREAMBLE The purpose of this License is to make a manual, textbook, or other written document free in the sense of freedom: to assure everyone the effective freedom to copy and redistribute it, with or without modifying it, either commercially or noncommercially. Secondarily, this License preserves for the author and publisher a way to get credit for their work, while not being considered responsible for modifications made by others. This License is a kind of “copyleft”, which means that derivative works of the document must themselves be free in the same sense. It complements the GNU General Public License, which is a copyleft license designed for free software. We have designed this License in order to use it for manuals for free software, because free software needs free documentation: a free program should come with manuals providing the same freedoms that the software does. But this License is not limited to software manuals; it can be used for any textual work, regardless of subject matter or whether it is published as a printed book. We recommend this License principally for works whose purpose is instruction or reference.

1. APPLICABILITY AND DEFINITIONS This License applies to any manual or other work that contains a notice placed by the copyright holder saying it can be distributed under the terms of this License. The “Document”, below, refers to any such manual or work. Any member of the public is a licensee, and is addressed as “you”. A “Modified Version” of the Document means any work containing the Document or a portion of it, either copied verbatim, or with modifications and/or translated into another language. A “Secondary Section” is a named appendix or a front-matter section of the Document that deals exclusively with the relationship of the publishers or authors of the Document to the Document’s overall subject (or to related matters) and contains nothing that could fall directly within that overall subject. (For example, if the Document is in part a textbook of mathematics, a Secondary Section may not explain any mathematics.) The relationship could be a matter of historical connection with the subject or with related matters, or of legal, commercial, philosophical, ethical or political position regarding them. The “Invariant Sections” are certain Secondary Sections whose titles are designated, as being those of Invariant Sections, in the notice that says that the Document is released under this License. The “Cover Texts” are certain short passages of text that are listed, as Front-Cover Texts or Back-Cover Texts, in the notice that says that the Document is released under this License.

A “Transparent” copy of the Document means a machine-readable copy, represented in a format whose specification is available to the general public, whose contents can be viewed and edited directly and straightforwardly with generic text editors or (for images composed of pixels) generic paint programs or (for drawings) some widely available drawing editor, and that is suitable for input to text formatters or for automatic translation to a variety of formats suitable for input to text formatters. A copy made in an otherwise Transparent file format whose markup has been designed to thwart or discourage subsequent modification by readers is not Transparent. A copy that is not “Transparent” is called “Opaque”. Examples of suitable formats for Transparent copies include plain ascii without markup, Texinfo input format, LaTEX input format, sgml or xml using a publicly available dtd, and standard-conforming simple html designed for human modification. Opaque formats include PostScript, pdf, proprietary formats that can be read and edited only by proprietary word processors, sgml or xml for which the dtd and/or processing tools are not generally available, and the machine-generated html produced by some word processors for output purposes only. The “Title Page” means, for a printed book, the title page itself, plus such following pages as are needed to hold, legibly, the material this License requires to appear in the title page. For works in formats which do not have any title page as such, “Title Page” means the text near the most prominent appearance of the work’s title, preceding the beginning of the body of the text.

2. VERBATIM COPYING You may copy and distribute the Document in any medium, either commercially or noncommercially, provided that this License, the copyright notices, and the license notice saying this License applies to the Document are reproduced in all copies, and that you add no other conditions whatsoever to those of this License. You may not use technical measures to obstruct or control the reading or further copying of the copies you make or distribute. However, you may accept compensation in exchange for copies. If you distribute a large enough number of copies you must also follow the conditions in section 3. You may also lend copies, under the same conditions stated above, and you may publicly display copies.

3. COPYING IN QUANTITY If you publish printed copies of the Document numbering more than 100, and the Document’s license notice requires Cover Texts, you must enclose the copies in covers that carry, clearly and legibly, all these Cover Texts: Front-Cover Texts on the front cover, and Back-Cover Texts on the back cover. Both covers must also clearly and legibly identify you as the publisher of these copies. The front cover must present the full title with all words of the title equally prominent and visible. You may add other material on the covers in addition. Copying with changes limited to the covers, as long as they preserve the title of the Document and satisfy these conditions, can be treated as verbatim copying in other respects. If the required texts for either cover are too voluminous to fit legibly, you should put the first ones listed (as many as fit reasonably) on the actual cover, and continue the rest onto adjacent pages. If you publish or distribute Opaque copies of the Document numbering more than 100, you must either include a machine-readable Transparent copy along with each Opaque copy, or state in or with each Opaque copy a publicly-accessible computer-network location containing a complete Transparent copy of the Document, free of added material, which the general network-using public has access to download anonymously at no charge using public-standard network protocols. If you use the latter option, you must take reasonably prudent steps, when you begin distribution of Opaque copies in quantity, to ensure that this Transparent copy will remain thus accessible at the stated location until at least one year after the last time you distribute an Opaque copy (directly or through your agents or retailers) of that edition to the public. It is requested, but not required, that you contact the authors of the Document well before redistributing any large number of copies, to give them a chance to provide you with an updated version of the Document.

4. MODIFICATIONS You may copy and distribute a Modified Version of the Document under the conditions of sections 2 and 3 above, provided that you release the Modified Version under precisely this License, with the Modified Version filling the role of the Document, thus licensing distribution and modification of the Modified Version to whoever possesses a copy of it. In addition, you must do these things in the Modified Version: A. Use in the Title Page (and on the covers, if any) a title distinct from that of the

Document, and from those of previous versions (which should, if there were any, be listed in the History section of the Document). You may use the same title as a previous version if the original publisher of that version gives permission.

B. List on the Title Page, as authors, one or more persons or entities responsible for authorship of the modifications in the Modified Version, together with at least five of the principal authors of the Document (all of its principal authors, if it has less than five).

C. State on the Title page the name of the publisher of the Modified Version, as the

publisher. D. Preserve all the copyright notices of the Document. E. Add an appropriate copyright notice for your modifications adjacent to the other

copyright notices. F. Include, immediately after the copyright notices, a license notice giving the public

permission to use the Modified Version under the terms of this License, in the form shown in the Addendum below.

G. Preserve in that license notice the full lists of Invariant Sections and required Cover

Texts given in the Document’s license notice. H. Include an unaltered copy of this License. I. Preserve the section entitled “History”, and its title, and add to it an item stating

at least the title, year, new authors, and publisher of the Modified Version as given on the Title Page. If there is no section entitled “History” in the Document, create one stating the title, year, authors, and publisher of the Document as given on its Title Page, then add an item describing the Modified Version as stated in the previous sentence.

J. Preserve the network location, if any, given in the Document for public access to a Transparent copy of the Document, and likewise the network locations given in the Document for previous versions it was based on. These may be placed in the “History” section. You may omit a network location for a work that was published at least four years before the Document itself, or if the original publisher of the version it refers to gives permission.

K. In any section entitled “Acknowledgments” or “Dedications”, preserve the section’s title, and preserve in the section all the substance and tone of each of the contributor acknowledgments and/or dedications given therein.

L. Preserve all the Invariant Sections of the Document, unaltered in their text and in their titles. Section numbers or the equivalent are not considered part of the section titles.

M. Delete any section entitled “Endorsements”. Such a section may not be included

in the Modified Version. N. Do not retitle any existing section as “Endorsements” or to conflict in title with

any Invariant Section.

If the Modified Version includes new front-matter sections or appendices that qualify as Secondary Sections and contain no material copied from the Document, you may at your option designate some or all of these sections as invariant. To do this, add their titles to the list of Invariant Sections in the Modified Version’s license notice. These titles must be distinct from any other section titles. You may add a section entitled “Endorsements”, provided it contains nothing but endorsements of your Modified Version by various parties—for example, statements of peer review or that the text has been approved by an organization as the authoritative definition of a standard. You may add a passage of up to five words as a Front-Cover Text, and a passage of up to 25 words as a Back-Cover Text, to the end of the list of Cover Texts in the Modified Version. Only one passage of Front-Cover Text and one of Back-Cover Text may be added by (or through arrangements made by) any one entity. If the Document already includes a cover text for the same cover, previously added by you or by arrangement made by the same entity you are acting on behalf of, you may not add another; but you may replace the old one, on explicit permission from the previous publisher that added the old one. The author(s) and publisher(s) of the Document do not by this License give permission to use their names for publicity for or to assert or imply endorsement of any Modified Version.

5. COMBINING DOCUMENTS You may combine the Document with other documents released under this License, under the terms defined in section 4 above for modified versions, provided that you include in the combination all of the Invariant Sections of all of the original documents, unmodified, and list them all as Invariant Sections of your combined work in its license notice. The combined work need only contain one copy of this License, and multiple identical Invariant Sections may be replaced with a single copy. If there are multiple Invariant Sections with the same name but different contents, make the title of each such section unique by adding at the end of it, in parentheses, the name of the original author or publisher of that section if known, or else a unique number. Make the same adjustment to the section titles in the list of Invariant Sections in the license notice of the combined work. In the combination, you must combine any sections entitled “History” in the various original documents, forming one section entitled “History”; likewise combine any sections entitled “Acknowledgments”, and any sections entitled “Dedications”. You must delete all sections entitled “Endorsements.”

6. COLLECTIONS OF DOCUMENTS You may make a collection consisting of the Document and other documents released under this License, and replace the individual copies of this License in the various documents with a single copy that is included in the collection, provided that you follow the rules of this License for verbatim copying of each of the documents in all other respects. You may extract a single document from such a collection, and distribute it individually under this License, provided you insert a copy of this License into the extracted document, and follow this License in all other respects regarding verbatim copying of that document.

7. AGGREGATION WITH INDEPENDENT WORKS A compilation of the Document or its derivatives with other separate and independent documents or works, in or on a volume of a storage or distribution medium, does not as a whole count as a Modified Version of the Document, provided no compilation copyright is claimed for the compilation. Such a compilation is called an “aggregate”, and this License does not apply to the other self-contained works thus compiled with the Document, on account of their being thus compiled, if they are not themselves derivative works of the Document. If the Cover Text requirement of section 3 is applicable to these copies of the Document, then if the Document is less than one quarter of the entire aggregate, the Document’s Cover Texts may be placed on covers that surround only the Document within the aggregate. Otherwise they must appear on covers around the whole aggregate.

8. TRANSLATION Translation is considered a kind of modification, so you may distribute translations of the Document under the terms of section 4. Replacing Invariant Sections with translations requires special permission from their copyright holders, but you may include translations of some or all Invariant Sections in addition to the original versions of these Invariant Sections. You may include a translation of this License provided that you also include the original English version of this License. In case of a disagreement between the translation and the original English version of this License, the original English version will prevail.

9. TERMINATION You may not copy, modify, sublicense, or distribute the Document except as expressly provided for under this License. Any other attempt to copy, modify, sublicense or distribute the Document is void, and will automatically terminate your rights under this License. However, parties who have received copies, or rights, from you under this License will not have their licenses terminated so long as such parties remain in full compliance.

10. FUTURE REVISIONS OF THIS LICENSE The Free Software Foundation may publish new, revised versions of the GNU Free Documentation License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns.

See http://www.gnu.org/copyleft/.

Each version of the License is given a distinguishing version number. If the Document specifies that a particular numbered version of this License “or any later version” applies to it, you have the option of following the terms and conditions either of that specified version or of any later version that has been published (not as a draft) by the Free Software Foundation. If the Document does not specify a version number of this License, you may choose any version ever published (not as a draft) by the Free Software Foundation.

ADDENDUM: How to use this License for your documents

To use this License in a document you have written, include a copy of the License in the document and put the following copyright and license notices just after the title page:

Copyright (C) year your name.  Permission is
granted to copy, distribute and/or modify this document
under the terms of the GNU Free Documentation License,
Version 1.1 or any later version published by the Free
Software Foundation; with the Invariant Sections being
list their titles, with the Front-Cover Texts being
list, and with the Back-Cover Texts being
list.  A copy of the license is included in the
section entitled ``GNU Free Documentation License''.

If you have no Invariant Sections, write “with no Invariant Sections” instead of saying which ones are invariant. If you have no Front-Cover Texts, write “no Front-Cover Texts” instead of “Front-Cover Texts being list”; likewise for Back-Cover Texts.

If your document contains nontrivial examples of program code, we recommend releasing these examples in parallel under your choice of free software license, such as the GNU General Public License, to permit their use in free software.

Function Index 406

Function Index

C

cblas_caxpy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 379 cblas_ccopy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 379 cblas_cdotc_sub . . . . . . . . . . . . . . . . . . . . . . . . . . . 378 cblas_cdotu_sub . . . . . . . . . . . . . . . . . . . . . . . . . . . 378 cblas_cgbmv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 382 cblas_cgemm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 388 cblas_cgemv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 382 cblas_cgerc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 386 cblas_cgeru . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 385 cblas_chbmv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 385 cblas_chemm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 389 cblas_chemv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 385 cblas_cher . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 386 cblas_cher2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 386 cblas_cher2k . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 390 cblas_cherk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 390 cblas_chpmv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 385 cblas_chpr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 386 cblas_chpr2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 386 cblas_cscal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 380 cblas_csscal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 380 cblas_cswap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 379 cblas_csymm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 388 cblas_csyr2k . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 388 cblas_csyrk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 388 cblas_ctbmv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 383 cblas_ctbsv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 383 cblas_ctpmv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 383 cblas_ctpsv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 383 cblas_ctrmm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 389 cblas_ctrmv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 382 cblas_ctrsm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 389 cblas_ctrsv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 383 cblas_dasum . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 378 cblas_daxpy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 379 cblas_dcopy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 379 cblas_ddot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 378 cblas_dgbmv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 382 cblas_dgemm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 387 cblas_dgemv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 381 cblas_dger . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 385 cblas_dnrm2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 378 cblas_drot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 380 cblas_drotg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 380 cblas_drotm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 380 cblas_drotmg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 380 cblas_dsbmv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 385 cblas_dscal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 380 cblas_dsdot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 378 cblas_dspmv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 385 cblas_dspr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 385 cblas_dspr2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 385 cblas_dswap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 379 cblas_dsymm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 388 cblas_dsymv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 384 cblas_dsyr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 385 cblas_dsyr2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 385 cblas_dsyr2k . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 388 cblas_dsyrk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 388 cblas_dtbmv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 382 cblas_dtbsv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 382 cblas_dtpmv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 382 cblas_dtpsv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 382 cblas_dtrmm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 388 cblas_dtrmv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 382 cblas_dtrsm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 388 cblas_dtrsv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 382 cblas_dzasum . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 379 cblas_dznrm2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 378 cblas_icamax . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 379 cblas_idamax . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 379 cblas_isamax . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 379 cblas_izamax . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 379 cblas_sasum . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 378 cblas_saxpy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 379 cblas_scasum . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 378 cblas_scnrm2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 378 cblas_scopy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 379 cblas_sdot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 378 cblas_sdsdot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 378 cblas_sgbmv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 381 cblas_sgemm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 387 cblas_sgemv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 381 cblas_sger . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 384 cblas_snrm2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 378 cblas_srot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 380 cblas_srotg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 380 cblas_srotm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 380 cblas_srotmg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 380 cblas_ssbmv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 384 cblas_sscal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 380 cblas_sspmv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 384 cblas_sspr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 384 cblas_sspr2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 384 cblas_sswap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 379 cblas_ssymm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 387 cblas_ssymv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 384 cblas_ssyr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 384 cblas_ssyr2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 384 cblas_ssyr2k . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 387 cblas_ssyrk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 387 cblas_stbmv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 381 cblas_stbsv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 381 cblas_stpmv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 381 cblas_stpsv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 381 cblas_strmm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 387 cblas_strmv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 381 cblas_strsm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 387 cblas_strsv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 381 cblas_xerbla . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 390 cblas_zaxpy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 379 cblas_zcopy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 379 cblas_zdotc_sub . . . . . . . . . . . . . . . . . . . . . . . . . . . 378 cblas_zdotu_sub . . . . . . . . . . . . . . . . . . . . . . . . . . . 378 cblas_zdscal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 380 cblas_zgbmv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 383 cblas_zgemm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 389 cblas_zgemv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 383 cblas_zgerc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 386 cblas_zgeru . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 386 cblas_zhbmv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 386 cblas_zhemm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 390 cblas_zhemv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 386 cblas_zher . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 386 cblas_zher2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 387 cblas_zher2k . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 390 cblas_zherk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 390 cblas_zhpmv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 386 cblas_zhpr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 386 cblas_zhpr2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 387 cblas_zscal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 380 cblas_zswap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 379 cblas_zsymm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 389 cblas_zsyr2k . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 389 cblas_zsyrk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 389 cblas_ztbmv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 383 cblas_ztbsv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 384 cblas_ztpmv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 383 cblas_ztpsv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 384 cblas_ztrmm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 389 cblas_ztrmv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 383 cblas_ztrsm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 389 cblas_ztrsv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 383

G

gsl_acosh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 gsl_asinh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 gsl_atanh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 gsl_blas_caxpy . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 gsl_blas_ccopy . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 gsl_blas_cdotc . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 gsl_blas_cdotu . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 gsl_blas_cgemm . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 gsl_blas_cgemv . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 gsl_blas_cgerc . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 gsl_blas_cgeru . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 gsl_blas_chemm . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 gsl_blas_chemv . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 gsl_blas_cher . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 gsl_blas_cher2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 gsl_blas_cher2k . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 gsl_blas_cherk . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 gsl_blas_cscal . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 gsl_blas_csscal . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 gsl_blas_cswap . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 gsl_blas_csymm . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 gsl_blas_csyr2k . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 gsl_blas_csyrk . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 gsl_blas_ctrmm . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 gsl_blas_ctrmv . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 gsl_blas_ctrsm . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 gsl_blas_ctrsv . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 gsl_blas_dasum . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 gsl_blas_daxpy . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 gsl_blas_dcopy . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 gsl_blas_ddot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 gsl_blas_dgemm . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 gsl_blas_dgemv . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 gsl_blas_dger . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 gsl_blas_dnrm2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 gsl_blas_drot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 gsl_blas_drotg . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 gsl_blas_drotm . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 gsl_blas_drotmg . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 gsl_blas_dscal . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 gsl_blas_dsdot . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 gsl_blas_dswap . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 gsl_blas_dsymm . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 gsl_blas_dsymv . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 gsl_blas_dsyr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 gsl_blas_dsyr2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 gsl_blas_dsyr2k . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 gsl_blas_dsyrk . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 gsl_blas_dtrmm . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 gsl_blas_dtrmv . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 gsl_blas_dtrsm . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 gsl_blas_dtrsv . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 gsl_blas_dzasum . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 gsl_blas_dznrm2 . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 gsl_blas_icamax . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 gsl_blas_idamax . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 gsl_blas_isamax . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 gsl_blas_izamax . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 gsl_blas_sasum . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 gsl_blas_saxpy . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 gsl_blas_scasum . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 gsl_blas_scnrm2 . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 gsl_blas_scopy . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 gsl_blas_sdot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 gsl_blas_sdsdot . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 gsl_blas_sgemm . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 gsl_blas_sgemv . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 gsl_blas_sger . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 gsl_blas_snrm2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 gsl_blas_srot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 gsl_blas_srotg . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 gsl_blas_srotm . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 gsl_blas_srotmg . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 gsl_blas_sscal . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 gsl_blas_sswap . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 gsl_blas_ssymm . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 gsl_blas_ssymv . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 gsl_blas_ssyr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 gsl_blas_ssyr2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 gsl_blas_ssyr2k . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 gsl_blas_ssyrk . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 gsl_blas_strmm . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 gsl_blas_strmv . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 gsl_blas_strsm . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 gsl_blas_strsv . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 gsl_blas_zaxpy . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 gsl_blas_zcopy . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 gsl_blas_zdotc . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 gsl_blas_zdotu . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 gsl_blas_zdscal . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 gsl_blas_zgemm . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 gsl_blas_zgemv . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 gsl_blas_zgerc . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 gsl_blas_zgeru . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 gsl_blas_zhemm . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 gsl_blas_zhemv . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 gsl_blas_zher . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 gsl_blas_zher2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 gsl_blas_zher2k . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 gsl_blas_zherk . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 gsl_blas_zscal . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 gsl_blas_zswap . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 gsl_blas_zsymm . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 gsl_blas_zsyr2k . . . . . . . . . . . . . . . . . . . . . . . . . . . 115 gsl_blas_zsyrk . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 gsl_blas_ztrmm . . . . . . . . . . . . . . . . . . . . . . . . . . . . 113 gsl_blas_ztrmv . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 gsl_blas_ztrsm . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 gsl_blas_ztrsv . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 gsl_block_alloc . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 gsl_block_calloc . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 gsl_block_fprintf . . . . . . . . . . . . . . . . . . . . . . . . . . 72 gsl_block_fread . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 gsl_block_free . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 gsl_block_fscanf . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 gsl_block_fwrite . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 gsl_cheb_alloc . . . . . . . . . . . . . . . . . . . . . . . . . . . . 286 gsl_cheb_calc_deriv. . . . . . . . . . . . . . . . . . . . . . . 287 gsl_cheb_calc_integ. . . . . . . . . . . . . . . . . . . . . . . 287 gsl_cheb_eval . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 287 gsl_cheb_eval_err . . . . . . . . . . . . . . . . . . . . . . . . . 287 gsl_cheb_eval_n . . . . . . . . . . . . . . . . . . . . . . . . . . . 287 gsl_cheb_eval_n_err. . . . . . . . . . . . . . . . . . . . . . . 287 gsl_cheb_free . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 286 gsl_cheb_init . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 286 gsl_complex_abs . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 gsl_complex_abs2 . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 gsl_complex_add . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 gsl_complex_add_imag. . . . . . . . . . . . . . . . . . . . . . . 20 gsl_complex_add_real. . . . . . . . . . . . . . . . . . . . . . . 19 gsl_complex_arccos . . . . . . . . . . . . . . . . . . . . . . . . . 22

                                                gsl_complex_arccos_real . . . . . . . . . . . . . . . . . . . 22

gsl_complex_arccosh. . . . . . . . . . . . . . . . . . . . . . . . 24

                                                gsl_complex_arccosh_real . . . . . . . . . . . . . . . . . . 24

gsl_complex_arccot . . . . . . . . . . . . . . . . . . . . . . . . . 23 gsl_complex_arccoth. . . . . . . . . . . . . . . . . . . . . . . . 24 gsl_complex_arccsc . . . . . . . . . . . . . . . . . . . . . . . . . 23

                                                gsl_complex_arccsc_real . . . . . . . . . . . . . . . . . . . 23

gsl_complex_arccsch. . . . . . . . . . . . . . . . . . . . . . . . 24 gsl_complex_arcsec . . . . . . . . . . . . . . . . . . . . . . . . . 23

                                                gsl_complex_arcsec_real . . . . . . . . . . . . . . . . . . . 23

gsl_complex_arcsech. . . . . . . . . . . . . . . . . . . . . . . . 24 gsl_complex_arcsin . . . . . . . . . . . . . . . . . . . . . . . . . 22

                                                gsl_complex_arcsin_real . . . . . . . . . . . . . . . . . . . 22

gsl_complex_arcsinh. . . . . . . . . . . . . . . . . . . . . . . . 24 gsl_complex_arctan . . . . . . . . . . . . . . . . . . . . . . . . . 23 gsl_complex_arctanh. . . . . . . . . . . . . . . . . . . . . . . . 24

                                                gsl_complex_arctanh_real . . . . . . . . . . . . . . . . . . 24

gsl_complex_arg . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 gsl_complex_conjugate . . . . . . . . . . . . . . . . . . . . . 20 gsl_complex_cos . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 gsl_complex_cosh . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 gsl_complex_cot . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 gsl_complex_coth . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 gsl_complex_csc . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 gsl_complex_csch . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 gsl_complex_div . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 gsl_complex_div_imag. . . . . . . . . . . . . . . . . . . . . . . 20 gsl_complex_div_real. . . . . . . . . . . . . . . . . . . . . . . 20 gsl_complex_exp . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 gsl_complex_inverse. . . . . . . . . . . . . . . . . . . . . . . . 20

                                                gsl_complex_linalg_LU_invert . . . . . . . . . . . . . 119

gsl_complex_log . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 gsl_complex_log_b . . . . . . . . . . . . . . . . . . . . . . . . . . 21 gsl_complex_log10 . . . . . . . . . . . . . . . . . . . . . . . . . . 21 gsl_complex_logabs . . . . . . . . . . . . . . . . . . . . . . . . . 19 gsl_complex_mul . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 gsl_complex_mul_imag. . . . . . . . . . . . . . . . . . . . . . . 20 gsl_complex_mul_real. . . . . . . . . . . . . . . . . . . . . . . 20 gsl_complex_negative. . . . . . . . . . . . . . . . . . . . . . . 20 gsl_complex_polar . . . . . . . . . . . . . . . . . . . . . . . . . . 18 gsl_complex_pow . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 gsl_complex_pow_real. . . . . . . . . . . . . . . . . . . . . . . 21 gsl_complex_rect . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 gsl_complex_sec . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 gsl_complex_sech . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 gsl_complex_sin . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 gsl_complex_sinh . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 gsl_complex_sqrt . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 gsl_complex_sqrt_real . . . . . . . . . . . . . . . . . . . . . 21 gsl_complex_sub . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 gsl_complex_sub_imag. . . . . . . . . . . . . . . . . . . . . . . 20 gsl_complex_sub_real. . . . . . . . . . . . . . . . . . . . . . . 19 gsl_complex_tan . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 gsl_complex_tanh . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 gsl_dht_alloc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 295 gsl_dht_apply . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 295 gsl_dht_free . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 295 gsl_dht_init . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 295 gsl_dht_k_sample . . . . . . . . . . . . . . . . . . . . . . . . . . 295 gsl_dht_new . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 295 gsl_dht_x_sample . . . . . . . . . . . . . . . . . . . . . . . . . . 295 gsl_diff_backward . . . . . . . . . . . . . . . . . . . . . . . . . 284 gsl_diff_central . . . . . . . . . . . . . . . . . . . . . . . . . . 284 gsl_diff_forward . . . . . . . . . . . . . . . . . . . . . . . . . . 284 GSL_EDOM. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 gsl_eigen_herm . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 gsl_eigen_herm_alloc. . . . . . . . . . . . . . . . . . . . . . 131 gsl_eigen_herm_free. . . . . . . . . . . . . . . . . . . . . . . 131 gsl_eigen_hermv . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 gsl_eigen_hermv_alloc . . . . . . . . . . . . . . . . . . . . 131 gsl_eigen_hermv_free. . . . . . . . . . . . . . . . . . . . . . 131 gsl_eigen_hermv_sort. . . . . . . . . . . . . . . . . . . . . . 132 gsl_eigen_symm . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 gsl_eigen_symm_alloc. . . . . . . . . . . . . . . . . . . . . . 130 gsl_eigen_symm_free. . . . . . . . . . . . . . . . . . . . . . . 130 gsl_eigen_symmv . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 gsl_eigen_symmv_alloc . . . . . . . . . . . . . . . . . . . . 130 gsl_eigen_symmv_free. . . . . . . . . . . . . . . . . . . . . . 130 gsl_eigen_symmv_sort. . . . . . . . . . . . . . . . . . . . . . 132 GSL_EINVAL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 GSL_ENOMEM . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 GSL_ERANGE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 GSL_ERROR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 GSL_ERROR_VAL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 gsl_expm1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

gsl_fft_complex_backward . . . . . . . . . . . . . . . . . 141
gsl_fft_complex_forward . . . . . . . . . . . . . . . . . . 141
gsl_fft_complex_inverse . . . . . . . . . . . . . . . . . . 141
gsl_fft_complex_radix2_backward. . . . . . . . . . 137
gsl_fft_complex_radix2_dif_backward . . . . . 137
gsl_fft_complex_radix2_dif_forward . . . . . . 137
gsl_fft_complex_radix2_dif_inverse . . . . . . 137
gsl_fft_complex_radix2_dif_transform . . . . 137
gsl_fft_complex_radix2_forward . . . . . . . . . . . 137
gsl_fft_complex_radix2_inverse . . . . . . . . . . . 137
gsl_fft_complex_radix2_transform. . . . . . . . . 137
gsl_fft_complex_transform . . . . . . . . . . . . . . . . 141
gsl_fft_complex_wavetable_alloc. . . . . . . . . . 140
gsl_fft_complex_wavetable_free . . . . . . . . . . . 140
gsl_fft_complex_workspace_alloc. . . . . . . . . . 141

                                                gsl_fft_complex_workspace_free . . . . . . . . . . . 141
                                                gsl_fft_halfcomplex_radix2_backward . . . . . 144
                                                gsl_fft_halfcomplex_radix2_inverse . . . . . . 144
                                                gsl_fft_halfcomplex_transform . . . . . . . . . . . . 147
                                                gsl_fft_halfcomplex_unpack . . . . . . . . . . . . . . . 147
                                                gsl_fft_halfcomplex_wavetable_alloc . . . . . 146
                                                gsl_fft_halfcomplex_wavetable_free . . . . . . 146
                                                gsl_fft_real_radix2_transform . . . . . . . . . . . . 144
                                                gsl_fft_real_transform . . . . . . . . . . . . . . . . . . . 147

gsl_fft_real_unpack. . . . . . . . . . . . . . . . . . . . . . . 147

                                                gsl_fft_real_wavetable_alloc . . . . . . . . . . . . . 146
                                                gsl_fft_real_wavetable_free . . . . . . . . . . . . . . 146
                                                gsl_fft_real_workspace_alloc . . . . . . . . . . . . . 146
                                                gsl_fft_real_workspace_free . . . . . . . . . . . . . . 146

gsl_finite . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 gsl_fit_linear . . . . . . . . . . . . . . . . . . . . . . . . . . . . 339 gsl_fit_linear_est. . . . . . . . . . . . . . . . . . . . . . . . 339 gsl_fit_mul . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 340 gsl_fit_mul_est . . . . . . . . . . . . . . . . . . . . . . . . . . . 340 gsl_fit_wlinear . . . . . . . . . . . . . . . . . . . . . . . . . . . 339 gsl_fit_wmul . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 340 gsl_heapsort . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 gsl_heapsort_index. . . . . . . . . . . . . . . . . . . . . . . . 100

                                                gsl_histogram_accumulate . . . . . . . . . . . . . . . . . 229

gsl_histogram_add . . . . . . . . . . . . . . . . . . . . . . . . . 231 gsl_histogram_alloc. . . . . . . . . . . . . . . . . . . . . . . 228 gsl_histogram_bins. . . . . . . . . . . . . . . . . . . . . . . . 230 gsl_histogram_clone. . . . . . . . . . . . . . . . . . . . . . . 229 gsl_histogram_div . . . . . . . . . . . . . . . . . . . . . . . . . 231

                                                gsl_histogram_equal_bins_p . . . . . . . . . . . . . . . 231

gsl_histogram_find. . . . . . . . . . . . . . . . . . . . . . . . 230 gsl_histogram_fprintf . . . . . . . . . . . . . . . . . . . . 232 gsl_histogram_fread. . . . . . . . . . . . . . . . . . . . . . . 232 gsl_histogram_free. . . . . . . . . . . . . . . . . . . . . . . . 228 gsl_histogram_fscanf. . . . . . . . . . . . . . . . . . . . . . 232 gsl_histogram_fwrite. . . . . . . . . . . . . . . . . . . . . . 232 gsl_histogram_get . . . . . . . . . . . . . . . . . . . . . . . . . 229

                                                gsl_histogram_get_range . . . . . . . . . . . . . . . . . . 229
                                                gsl_histogram_increment . . . . . . . . . . . . . . . . . . 229

gsl_histogram_max . . . . . . . . . . . . . . . . . . . . . . . . . 230 gsl_histogram_max_bin . . . . . . . . . . . . . . . . . . . . 230 gsl_histogram_max_val . . . . . . . . . . . . . . . . . . . . 230 gsl_histogram_mean. . . . . . . . . . . . . . . . . . . . . . . . 230 gsl_histogram_memcpy. . . . . . . . . . . . . . . . . . . . . . 229 gsl_histogram_min . . . . . . . . . . . . . . . . . . . . . . . . . 230 gsl_histogram_min_bin . . . . . . . . . . . . . . . . . . . . 230 gsl_histogram_min_val . . . . . . . . . . . . . . . . . . . . 230 gsl_histogram_mul . . . . . . . . . . . . . . . . . . . . . . . . . 231

                                                gsl_histogram_pdf_alloc . . . . . . . . . . . . . . . . . . 233
                                                gsl_histogram_pdf_free . . . . . . . . . . . . . . . . . . . 234
                                                gsl_histogram_pdf_init . . . . . . . . . . . . . . . . . . . 233
                                                gsl_histogram_pdf_sample . . . . . . . . . . . . . . . . . 234

gsl_histogram_reset. . . . . . . . . . . . . . . . . . . . . . . 230 gsl_histogram_scale. . . . . . . . . . . . . . . . . . . . . . . 231

gsl_histogram_set_ranges . . . . . . . . . . . . . . . . . 228
gsl_histogram_set_ranges_uniform. . . . . . . . . 228

gsl_histogram_shift. . . . . . . . . . . . . . . . . . . . . . . 231 gsl_histogram_sigma. . . . . . . . . . . . . . . . . . . . . . . 230 gsl_histogram_sub . . . . . . . . . . . . . . . . . . . . . . . . . 231

gsl_histogram2d_accumulate . . . . . . . . . . . . . . . 238

gsl_histogram2d_add. . . . . . . . . . . . . . . . . . . . . . . 240 gsl_histogram2d_alloc . . . . . . . . . . . . . . . . . . . . 237 gsl_histogram2d_clone . . . . . . . . . . . . . . . . . . . . 237 gsl_histogram2d_div. . . . . . . . . . . . . . . . . . . . . . . 240

gsl_histogram2d_equal_bins_p . . . . . . . . . . . . . 240

gsl_histogram2d_find. . . . . . . . . . . . . . . . . . . . . . 239

gsl_histogram2d_fprintf . . . . . . . . . . . . . . . . . . 241

gsl_histogram2d_fread . . . . . . . . . . . . . . . . . . . . 241 gsl_histogram2d_free. . . . . . . . . . . . . . . . . . . . . . 237

gsl_histogram2d_fscanf . . . . . . . . . . . . . . . . . . . 242
gsl_histogram2d_fwrite . . . . . . . . . . . . . . . . . . . 241

gsl_histogram2d_get. . . . . . . . . . . . . . . . . . . . . . . 238

gsl_histogram2d_get_xrange . . . . . . . . . . . . . . . 238
gsl_histogram2d_get_yrange . . . . . . . . . . . . . . . 238
gsl_histogram2d_increment . . . . . . . . . . . . . . . . 238
gsl_histogram2d_max_bin . . . . . . . . . . . . . . . . . . 239
gsl_histogram2d_max_val . . . . . . . . . . . . . . . . . . 239
gsl_histogram2d_memcpy . . . . . . . . . . . . . . . . . . . 237
gsl_histogram2d_min_bin . . . . . . . . . . . . . . . . . . 239
gsl_histogram2d_min_val . . . . . . . . . . . . . . . . . . 239

gsl_histogram2d_mul. . . . . . . . . . . . . . . . . . . . . . . 240 gsl_histogram2d_nx. . . . . . . . . . . . . . . . . . . . . . . . 238 gsl_histogram2d_ny. . . . . . . . . . . . . . . . . . . . . . . . 238

gsl_histogram2d_pdf_alloc . . . . . . . . . . . . . . . . 242
gsl_histogram2d_pdf_free . . . . . . . . . . . . . . . . . 243
gsl_histogram2d_pdf_init . . . . . . . . . . . . . . . . . 243
gsl_histogram2d_pdf_sample . . . . . . . . . . . . . . . 243

gsl_histogram2d_reset . . . . . . . . . . . . . . . . . . . . 239 gsl_histogram2d_scale . . . . . . . . . . . . . . . . . . . . 240

gsl_histogram2d_set_ranges . . . . . . . . . . . . . . . 237
gsl_histogram2d_set_ranges_uniform . . . . . . 237

gsl_histogram2d_shift . . . . . . . . . . . . . . . . . . . . 240 gsl_histogram2d_sub. . . . . . . . . . . . . . . . . . . . . . . 240 gsl_histogram2d_xmax. . . . . . . . . . . . . . . . . . . . . . 238 gsl_histogram2d_xmin. . . . . . . . . . . . . . . . . . . . . . 238 gsl_histogram2d_ymax. . . . . . . . . . . . . . . . . . . . . . 238 gsl_histogram2d_ymin. . . . . . . . . . . . . . . . . . . . . . 238 gsl_hypot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 gsl_ieee_env_setup. . . . . . . . . . . . . . . . . . . . . . . . 368

gsl_ieee_fprintf_double . . . . . . . . . . . . . . . . . . 366
gsl_ieee_fprintf_float . . . . . . . . . . . . . . . . . . . 366
gsl_ieee_printf_double . . . . . . . . . . . . . . . . . . . 367

gsl_ieee_printf_float . . . . . . . . . . . . . . . . . . . . 367 GSL_IMAG. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 gsl_integration_qag. . . . . . . . . . . . . . . . . . . . . . . 152 gsl_integration_qagi. . . . . . . . . . . . . . . . . . . . . . 154 gsl_integration_qagil . . . . . . . . . . . . . . . . . . . . 154 gsl_integration_qagiu . . . . . . . . . . . . . . . . . . . . 154 gsl_integration_qagp. . . . . . . . . . . . . . . . . . . . . . 153 gsl_integration_qags. . . . . . . . . . . . . . . . . . . . . . 153 gsl_integration_qawc. . . . . . . . . . . . . . . . . . . . . . 154 gsl_integration_qawf. . . . . . . . . . . . . . . . . . . . . . 157 gsl_integration_qawo. . . . . . . . . . . . . . . . . . . . . . 156

                                                gsl_integration_qawo_table_alloc. . . . . . . . . 156
                                                gsl_integration_qawo_table_free. . . . . . . . . . 156
                                                gsl_integration_qawo_table_set . . . . . . . . . . . 156
                                                gsl_integration_qawo_table_set_length . . . 156

gsl_integration_qaws. . . . . . . . . . . . . . . . . . . . . . 155

                                                gsl_integration_qaws_table_alloc. . . . . . . . . 155
                                                gsl_integration_qaws_table_free. . . . . . . . . . 155
                                                gsl_integration_qaws_table_set . . . . . . . . . . . 155

gsl_integration_qng. . . . . . . . . . . . . . . . . . . . . . . 152

                                                gsl_integration_workspace_alloc. . . . . . . . . . 152
                                                gsl_integration_workspace_free . . . . . . . . . . . 152
                                                gsl_interp_accel_alloc . . . . . . . . . . . . . . . . . . . 280

gsl_interp_accel_find . . . . . . . . . . . . . . . . . . . . 280 gsl_interp_accel_free . . . . . . . . . . . . . . . . . . . . 280 gsl_interp_akima . . . . . . . . . . . . . . . . . . . . . . . . . . 279

                                                gsl_interp_akima_periodic . . . . . . . . . . . . . . . . 279

gsl_interp_alloc . . . . . . . . . . . . . . . . . . . . . . . . . . 278 gsl_interp_bsearch. . . . . . . . . . . . . . . . . . . . . . . . 280 gsl_interp_cspline. . . . . . . . . . . . . . . . . . . . . . . . 279

                                                gsl_interp_cspline_periodic . . . . . . . . . . . . . . 279

gsl_interp_eval . . . . . . . . . . . . . . . . . . . . . . . . . . . 280 gsl_interp_eval_deriv . . . . . . . . . . . . . . . . . . . . 280

                                                gsl_interp_eval_deriv_e . . . . . . . . . . . . . . . . . . 280
                                                gsl_interp_eval_deriv2 . . . . . . . . . . . . . . . . . . . 281
                                                gsl_interp_eval_deriv2_e . . . . . . . . . . . . . . . . . 281

gsl_interp_eval_e . . . . . . . . . . . . . . . . . . . . . . . . . 280 gsl_interp_eval_integ . . . . . . . . . . . . . . . . . . . . 281

                                                gsl_interp_eval_integ_e . . . . . . . . . . . . . . . . . . 281

gsl_interp_free . . . . . . . . . . . . . . . . . . . . . . . . . . . 278 gsl_interp_init . . . . . . . . . . . . . . . . . . . . . . . . . . . 278 gsl_interp_linear . . . . . . . . . . . . . . . . . . . . . . . . . 279 gsl_interp_min_size. . . . . . . . . . . . . . . . . . . . . . . 279 gsl_interp_name . . . . . . . . . . . . . . . . . . . . . . . . . . . 279 gsl_interval . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 312 GSL_IS_EVEN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 GSL_IS_ODD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 gsl_isinf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 gsl_isnan . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

                                                gsl_linalg_bidiag_decomp . . . . . . . . . . . . . . . . . 126
                                                gsl_linalg_bidiag_unpack . . . . . . . . . . . . . . . . . 126
                                                gsl_linalg_bidiag_unpack_B . . . . . . . . . . . . . . . 126
                                                gsl_linalg_bidiag_unpack2 . . . . . . . . . . . . . . . . 126
                                                gsl_linalg_cholesky_decomp . . . . . . . . . . . . . . . 124
                                                gsl_linalg_cholesky_solve . . . . . . . . . . . . . . . . 124
                                                gsl_linalg_cholesky_svx . . . . . . . . . . . . . . . . . . 124
                                                gsl_linalg_complex_LU_decomp . . . . . . . . . . . . . 118
                                                gsl_linalg_complex_LU_det . . . . . . . . . . . . . . . . 119
                                                gsl_linalg_complex_LU_lndet . . . . . . . . . . . . . . 119
                                                gsl_linalg_complex_LU_refine . . . . . . . . . . . . . 119

gsl_linalg_complex_LU_sgndet . . . . . . . . . . . . . 119
gsl_linalg_complex_LU_solve . . . . . . . . . . . . . . 118
gsl_linalg_complex_LU_svx . . . . . . . . . . . . . . . . 119
gsl_linalg_hermtd_decomp . . . . . . . . . . . . . . . . . 125
gsl_linalg_hermtd_unpack . . . . . . . . . . . . . . . . . 126
gsl_linalg_hermtd_unpack_dsd . . . . . . . . . . . . . 126

gsl_linalg_HH_solve. . . . . . . . . . . . . . . . . . . . . . . 127 gsl_linalg_HH_svx . . . . . . . . . . . . . . . . . . . . . . . . . 127 gsl_linalg_LU_decomp. . . . . . . . . . . . . . . . . . . . . . 118 gsl_linalg_LU_det . . . . . . . . . . . . . . . . . . . . . . . . . 119 gsl_linalg_LU_invert. . . . . . . . . . . . . . . . . . . . . . 119 gsl_linalg_LU_lndet. . . . . . . . . . . . . . . . . . . . . . . 119 gsl_linalg_LU_refine. . . . . . . . . . . . . . . . . . . . . . 119 gsl_linalg_LU_sgndet. . . . . . . . . . . . . . . . . . . . . . 119 gsl_linalg_LU_solve. . . . . . . . . . . . . . . . . . . . . . . 118 gsl_linalg_LU_svx . . . . . . . . . . . . . . . . . . . . . . . . . 119 gsl_linalg_QR_decomp. . . . . . . . . . . . . . . . . . . . . . 120 gsl_linalg_QR_lssolve . . . . . . . . . . . . . . . . . . . . 120 gsl_linalg_QR_QRsolve . . . . . . . . . . . . . . . . . . . . 121 gsl_linalg_QR_QTvec. . . . . . . . . . . . . . . . . . . . . . . 120 gsl_linalg_QR_Qvec. . . . . . . . . . . . . . . . . . . . . . . . 121 gsl_linalg_QR_Rsolve. . . . . . . . . . . . . . . . . . . . . . 121 gsl_linalg_QR_Rsvx. . . . . . . . . . . . . . . . . . . . . . . . 121 gsl_linalg_QR_solve. . . . . . . . . . . . . . . . . . . . . . . 120 gsl_linalg_QR_svx . . . . . . . . . . . . . . . . . . . . . . . . . 120 gsl_linalg_QR_unpack. . . . . . . . . . . . . . . . . . . . . . 121 gsl_linalg_QR_update. . . . . . . . . . . . . . . . . . . . . . 121

gsl_linalg_QRPT_decomp . . . . . . . . . . . . . . . . . . . 122
gsl_linalg_QRPT_decomp2 . . . . . . . . . . . . . . . . . . 122
gsl_linalg_QRPT_QRsolve . . . . . . . . . . . . . . . . . . 123
gsl_linalg_QRPT_Rsolve . . . . . . . . . . . . . . . . . . . 123

gsl_linalg_QRPT_Rsvx. . . . . . . . . . . . . . . . . . . . . . 123 gsl_linalg_QRPT_solve . . . . . . . . . . . . . . . . . . . . 122 gsl_linalg_QRPT_svx. . . . . . . . . . . . . . . . . . . . . . . 122

gsl_linalg_QRPT_update . . . . . . . . . . . . . . . . . . . 123

gsl_linalg_R_solve. . . . . . . . . . . . . . . . . . . . . . . . 121 gsl_linalg_R_svx . . . . . . . . . . . . . . . . . . . . . . . . . . 121

gsl_linalg_solve_symm_cyc_tridiag . . . . . . . 127
gsl_linalg_solve_symm_tridiag . . . . . . . . . . . . 127

gsl_linalg_SV_decomp. . . . . . . . . . . . . . . . . . . . . . 123

gsl_linalg_SV_decomp_jacobi . . . . . . . . . . . . . . 124
gsl_linalg_SV_decomp_mod . . . . . . . . . . . . . . . . . 124

gsl_linalg_SV_solve. . . . . . . . . . . . . . . . . . . . . . . 124

gsl_linalg_symmtd_decomp . . . . . . . . . . . . . . . . . 125
gsl_linalg_symmtd_unpack . . . . . . . . . . . . . . . . . 125
gsl_linalg_symmtd_unpack_dsd . . . . . . . . . . . . . 125

gsl_log1p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 gsl_matrix_add . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89

gsl_matrix_add_constant . . . . . . . . . . . . . . . . . . . 90

gsl_matrix_alloc . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 gsl_matrix_calloc . . . . . . . . . . . . . . . . . . . . . . . . . . 83 gsl_matrix_column . . . . . . . . . . . . . . . . . . . . . . . . . . 87

gsl_matrix_const_column . . . . . . . . . . . . . . . . . . . 87
gsl_matrix_const_diagonal . . . . . . . . . . . . . . . . . 87

gsl_matrix_const_row. . . . . . . . . . . . . . . . . . . . . . . 87

                                                gsl_matrix_const_subdiagonal . . . . . . . . . . . . . . 88
                                                gsl_matrix_const_submatrix . . . . . . . . . . . . . . . . 85
                                                gsl_matrix_const_superdiagonal . . . . . . . . . . . . 88
                                                gsl_matrix_const_view_array . . . . . . . . . . . . . . . 85
                                                gsl_matrix_const_view_array_with_tda . . . . . 86
                                                gsl_matrix_const_view_vector . . . . . . . . . . . . . . 86
                                                gsl_matrix_const_view_vector_with_tda . . . . 86

gsl_matrix_diagonal. . . . . . . . . . . . . . . . . . . . . . . . 87

                                                gsl_matrix_div_elements . . . . . . . . . . . . . . . . . . . 90

gsl_matrix_fprintf . . . . . . . . . . . . . . . . . . . . . . . . . 84 gsl_matrix_fread . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 gsl_matrix_free . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 gsl_matrix_fscanf . . . . . . . . . . . . . . . . . . . . . . . . . . 84 gsl_matrix_fwrite . . . . . . . . . . . . . . . . . . . . . . . . . . 84 gsl_matrix_get . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 gsl_matrix_get_col . . . . . . . . . . . . . . . . . . . . . . . . . 88 gsl_matrix_get_row . . . . . . . . . . . . . . . . . . . . . . . . . 88 gsl_matrix_isnull . . . . . . . . . . . . . . . . . . . . . . . . . . 91 gsl_matrix_max . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 gsl_matrix_max_index. . . . . . . . . . . . . . . . . . . . . . . 90 gsl_matrix_memcpy . . . . . . . . . . . . . . . . . . . . . . . . . . 88 gsl_matrix_min . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 gsl_matrix_min_index. . . . . . . . . . . . . . . . . . . . . . . 90 gsl_matrix_minmax . . . . . . . . . . . . . . . . . . . . . . . . . . 90

                                                gsl_matrix_minmax_index . . . . . . . . . . . . . . . . . . . 90
                                                gsl_matrix_mul_elements . . . . . . . . . . . . . . . . . . . 90

gsl_matrix_ptr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 gsl_matrix_row . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87 gsl_matrix_scale . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 gsl_matrix_set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 gsl_matrix_set_all . . . . . . . . . . . . . . . . . . . . . . . . . 84 gsl_matrix_set_col . . . . . . . . . . . . . . . . . . . . . . . . . 89

                                                gsl_matrix_set_identity . . . . . . . . . . . . . . . . . . . 84

gsl_matrix_set_row . . . . . . . . . . . . . . . . . . . . . . . . . 89 gsl_matrix_set_zero. . . . . . . . . . . . . . . . . . . . . . . . 84 gsl_matrix_sub . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89

                                                gsl_matrix_subdiagonal . . . . . . . . . . . . . . . . . . . . 88

gsl_matrix_submatrix. . . . . . . . . . . . . . . . . . . . . . . 85

                                                gsl_matrix_superdiagonal . . . . . . . . . . . . . . . . . . 88

gsl_matrix_swap . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88

                                                gsl_matrix_swap_columns . . . . . . . . . . . . . . . . . . . 89
                                                gsl_matrix_swap_rowcol . . . . . . . . . . . . . . . . . . . . 89

gsl_matrix_swap_rows. . . . . . . . . . . . . . . . . . . . . . . 89 gsl_matrix_transpose. . . . . . . . . . . . . . . . . . . . . . . 89

                                                gsl_matrix_transpose_memcpy . . . . . . . . . . . . . . . 89

gsl_matrix_view_array . . . . . . . . . . . . . . . . . . . . . 85

                                                gsl_matrix_view_array_with_tda . . . . . . . . . . . . 86
                                                gsl_matrix_view_vector . . . . . . . . . . . . . . . . . . . . 86
                                                gsl_matrix_view_vector_with_tda . . . . . . . . . . . 86

GSL_MAX. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 GSL_MAX_DBL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 GSL_MAX_INT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 GSL_MAX_LDBL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 GSL_MIN. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 GSL_MIN_DBL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

gsl_min_fminimizer_alloc . . . . . . . . . . . . . . . . . 310
gsl_min_fminimizer_brent . . . . . . . . . . . . . . . . . 313
gsl_min_fminimizer_free . . . . . . . . . . . . . . . . . . 311
gsl_min_fminimizer_goldensection. . . . . . . . . 313
gsl_min_fminimizer_iterate . . . . . . . . . . . . . . . 311
gsl_min_fminimizer_minimum . . . . . . . . . . . . . . . 312
gsl_min_fminimizer_name . . . . . . . . . . . . . . . . . . 311
gsl_min_fminimizer_set . . . . . . . . . . . . . . . . . . . 311
gsl_min_fminimizer_set_with_values . . . . . . 311

GSL_MIN_INT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 GSL_MIN_LDBL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 gsl_min_test_interval . . . . . . . . . . . . . . . . . . . . 312 gsl_monte_miser_alloc . . . . . . . . . . . . . . . . . . . . 254 gsl_monte_miser_free. . . . . . . . . . . . . . . . . . . . . . 254 gsl_monte_miser_init. . . . . . . . . . . . . . . . . . . . . . 254

gsl_monte_miser_integrate . . . . . . . . . . . . . . . . 254

gsl_monte_plain_alloc . . . . . . . . . . . . . . . . . . . . 253 gsl_monte_plain_free. . . . . . . . . . . . . . . . . . . . . . 253 gsl_monte_plain_init. . . . . . . . . . . . . . . . . . . . . . 253

gsl_monte_plain_integrate . . . . . . . . . . . . . . . . 253

gsl_monte_vegas_alloc . . . . . . . . . . . . . . . . . . . . 256 gsl_monte_vegas_free. . . . . . . . . . . . . . . . . . . . . . 256 gsl_monte_vegas_init. . . . . . . . . . . . . . . . . . . . . . 256

gsl_monte_vegas_integrate . . . . . . . . . . . . . . . . 256

gsl_multifit_covar. . . . . . . . . . . . . . . . . . . . . . . . 353

gsl_multifit_fdfsolver_alloc . . . . . . . . . . . . . 348
gsl_multifit_fdfsolver_free . . . . . . . . . . . . . . 349
gsl_multifit_fdfsolver_iterate . . . . . . . . . . . 350
gsl_multifit_fdfsolver_lmder . . . . . . . . . . . . . 352
gsl_multifit_fdfsolver_lmsder . . . . . . . . . . . . 352
gsl_multifit_fdfsolver_name . . . . . . . . . . . . . . 349
gsl_multifit_fdfsolver_position. . . . . . . . . . 350
gsl_multifit_fdfsolver_set . . . . . . . . . . . . . . . 349
gsl_multifit_fsolver_alloc . . . . . . . . . . . . . . . 348
gsl_multifit_fsolver_free . . . . . . . . . . . . . . . . 349
gsl_multifit_fsolver_iterate . . . . . . . . . . . . . 350
gsl_multifit_fsolver_name . . . . . . . . . . . . . . . . 349
gsl_multifit_fsolver_position . . . . . . . . . . . . 350
gsl_multifit_fsolver_set . . . . . . . . . . . . . . . . . 349

gsl_multifit_gradient . . . . . . . . . . . . . . . . . . . . 351 gsl_multifit_linear. . . . . . . . . . . . . . . . . . . . . . . 341

gsl_multifit_linear_alloc . . . . . . . . . . . . . . . . 341
gsl_multifit_linear_free . . . . . . . . . . . . . . . . . 341
gsl_multifit_test_delta . . . . . . . . . . . . . . . . . . 351
gsl_multifit_test_gradient . . . . . . . . . . . . . . . 351

gsl_multifit_wlinear. . . . . . . . . . . . . . . . . . . . . . 342

gsl_multimin_fdfminimizer_alloc. . . . . . . . . . 331
gsl_multimin_fdfminimizer_conjugate_fr . . 335
gsl_multimin_fdfminimizer_conjugate_pr . . 335
gsl_multimin_fdfminimizer_free . . . . . . . . . . . 331
gsl_multimin_fdfminimizer_iterate . . . . . . . 333
gsl_multimin_fdfminimizer_name . . . . . . . . . . . 331

                                                gsl_multimin_fdfminimizer_restart . . . . . . . 334
                                                gsl_multimin_fdfminimizer_set . . . . . . . . . . . . 331
                                                gsl_multimin_fdfminimizer_steepest_descent

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 335

                                                gsl_multimin_fdfminimizer_vector_bfgs . . . 335
                                                gsl_multimin_test_gradient . . . . . . . . . . . . . . . 334
                                                gsl_multiroot_fdfsolver_alloc . . . . . . . . . . . . 317
                                                gsl_multiroot_fdfsolver_free . . . . . . . . . . . . . 317
                                                gsl_multiroot_fdfsolver_gnewton. . . . . . . . . . 323
                                                gsl_multiroot_fdfsolver_gradient. . . . . . . . . 334
                                                gsl_multiroot_fdfsolver_hybridj. . . . . . . . . . 322
                                                gsl_multiroot_fdfsolver_hybridsj. . . . . . . . . 322
                                                gsl_multiroot_fdfsolver_iterate. . . . . . . . . . 320
                                                gsl_multiroot_fdfsolver_minimum. . . . . . . . . . 334
                                                gsl_multiroot_fdfsolver_name . . . . . . . . . . . . . 318
                                                gsl_multiroot_fdfsolver_newton . . . . . . . . . . . 323
                                                gsl_multiroot_fdfsolver_root . . . . . . . . . . . . . 321
                                                gsl_multiroot_fdfsolver_set . . . . . . . . . . . . . . 317
                                                gsl_multiroot_fdfsolver_x . . . . . . . . . . . . . . . . 334
                                                gsl_multiroot_fsolver_alloc . . . . . . . . . . . . . . 317
                                                gsl_multiroot_fsolver_broyden . . . . . . . . . . . . 324
                                                gsl_multiroot_fsolver_dnewton . . . . . . . . . . . . 323
                                                gsl_multiroot_fsolver_free . . . . . . . . . . . . . . . 317
                                                gsl_multiroot_fsolver_hybrid . . . . . . . . . . . . . 323
                                                gsl_multiroot_fsolver_hybrids . . . . . . . . . . . . 323
                                                gsl_multiroot_fsolver_iterate . . . . . . . . . . . . 320
                                                gsl_multiroot_fsolver_name . . . . . . . . . . . . . . . 318
                                                gsl_multiroot_fsolver_root . . . . . . . . . . . . . . . 321
                                                gsl_multiroot_fsolver_set . . . . . . . . . . . . . . . . 317
                                                gsl_multiroot_test_delta . . . . . . . . . . . . . . . . . 321
                                                gsl_multiroot_test_residual . . . . . . . . . . . . . . 321

GSL_NAN. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 GSL_NEGINF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 gsl_ntuple_bookdata. . . . . . . . . . . . . . . . . . . . . . . 246 gsl_ntuple_close . . . . . . . . . . . . . . . . . . . . . . . . . . 246 gsl_ntuple_create . . . . . . . . . . . . . . . . . . . . . . . . . 245 gsl_ntuple_open . . . . . . . . . . . . . . . . . . . . . . . . . . . 245 gsl_ntuple_project. . . . . . . . . . . . . . . . . . . . . . . . 247 gsl_ntuple_read . . . . . . . . . . . . . . . . . . . . . . . . . . . 246 gsl_ntuple_write . . . . . . . . . . . . . . . . . . . . . . . . . . 246

                                                gsl_odeiv_control_alloc . . . . . . . . . . . . . . . . . . 273
                                                gsl_odeiv_control_free . . . . . . . . . . . . . . . . . . . 273
                                                gsl_odeiv_control_hadjust . . . . . . . . . . . . . . . . 273
                                                gsl_odeiv_control_init . . . . . . . . . . . . . . . . . . . 273
                                                gsl_odeiv_control_name . . . . . . . . . . . . . . . . . . . 274
                                                gsl_odeiv_control_standard_new . . . . . . . . . . . 272
                                                gsl_odeiv_control_y_new . . . . . . . . . . . . . . . . . . 273
                                                gsl_odeiv_control_yp_new . . . . . . . . . . . . . . . . . 273
                                                gsl_odeiv_evolve_alloc . . . . . . . . . . . . . . . . . . . 274
                                                gsl_odeiv_evolve_apply . . . . . . . . . . . . . . . . . . . 274

gsl_odeiv_evolve_free . . . . . . . . . . . . . . . . . . . . 274

                                                gsl_odeiv_evolve_reset . . . . . . . . . . . . . . . . . . . 274

gsl_odeiv_step_alloc. . . . . . . . . . . . . . . . . . . . . . 271 gsl_odeiv_step_apply. . . . . . . . . . . . . . . . . . . . . . 271 gsl_odeiv_step_bsimp. . . . . . . . . . . . . . . . . . . . . . 272 gsl_odeiv_step_free. . . . . . . . . . . . . . . . . . . . . . . 271 gsl_odeiv_step_gear1. . . . . . . . . . . . . . . . . . . . . . 272 gsl_odeiv_step_gear2. . . . . . . . . . . . . . . . . . . . . . 272 gsl_odeiv_step_name. . . . . . . . . . . . . . . . . . . . . . . 271 gsl_odeiv_step_order. . . . . . . . . . . . . . . . . . . . . . 271 gsl_odeiv_step_reset. . . . . . . . . . . . . . . . . . . . . . 271 gsl_odeiv_step_rk2. . . . . . . . . . . . . . . . . . . . . . . . 271 gsl_odeiv_step_rk2imp . . . . . . . . . . . . . . . . . . . . 272 gsl_odeiv_step_rk4. . . . . . . . . . . . . . . . . . . . . . . . 271 gsl_odeiv_step_rk4imp . . . . . . . . . . . . . . . . . . . . 272 gsl_odeiv_step_rk8pd. . . . . . . . . . . . . . . . . . . . . . 272 gsl_odeiv_step_rkck. . . . . . . . . . . . . . . . . . . . . . . 272 gsl_odeiv_step_rkf45. . . . . . . . . . . . . . . . . . . . . . 272 gsl_permutation_alloc . . . . . . . . . . . . . . . . . . . . . 94

gsl_permutation_calloc . . . . . . . . . . . . . . . . . . . . 94

gsl_permutation_data. . . . . . . . . . . . . . . . . . . . . . . 95

gsl_permutation_fprintf . . . . . . . . . . . . . . . . . . . 97

gsl_permutation_fread . . . . . . . . . . . . . . . . . . . . . 96 gsl_permutation_free. . . . . . . . . . . . . . . . . . . . . . . 94

gsl_permutation_fscanf . . . . . . . . . . . . . . . . . . . . 97
gsl_permutation_fwrite . . . . . . . . . . . . . . . . . . . . 96

gsl_permutation_get. . . . . . . . . . . . . . . . . . . . . . . . 95 gsl_permutation_init. . . . . . . . . . . . . . . . . . . . . . . 94

gsl_permutation_inverse . . . . . . . . . . . . . . . . . . . 95

gsl_permutation_next. . . . . . . . . . . . . . . . . . . . . . . 95 gsl_permutation_prev. . . . . . . . . . . . . . . . . . . . . . . 95

gsl_permutation_reverse . . . . . . . . . . . . . . . . . . . 95

gsl_permutation_size. . . . . . . . . . . . . . . . . . . . . . . 95 gsl_permutation_swap. . . . . . . . . . . . . . . . . . . . . . . 95 gsl_permutation_valid . . . . . . . . . . . . . . . . . . . . . 95 gsl_permute . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 gsl_permute_inverse. . . . . . . . . . . . . . . . . . . . . . . . 96 gsl_permute_vector . . . . . . . . . . . . . . . . . . . . . . . . . 96

gsl_permute_vector_inverse . . . . . . . . . . . . . . . . 96
gsl_poly_complex_solve . . . . . . . . . . . . . . . . . . . . 27
gsl_poly_complex_solve_cubic . . . . . . . . . . . . . . 27
gsl_poly_complex_solve_quadratic. . . . . . . . . . 26
gsl_poly_complex_workspace_alloc. . . . . . . . . . 27
gsl_poly_complex_workspace_free . . . . . . . . . . . 27

gsl_poly_eval . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 gsl_poly_solve_cubic. . . . . . . . . . . . . . . . . . . . . . . 27

gsl_poly_solve_quadratic . . . . . . . . . . . . . . . . . . 26

GSL_POSINF . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 gsl_pow_2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 gsl_pow_3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 gsl_pow_4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 gsl_pow_5 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 gsl_pow_6 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 gsl_pow_7 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 gsl_pow_8 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 gsl_pow_9 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 gsl_pow_int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 gsl_qrng_alloc . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 gsl_qrng_clone . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 gsl_qrng_free . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 gsl_qrng_get . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 gsl_qrng_init . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 gsl_qrng_memcpy . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 gsl_qrng_name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174

                                                gsl_qrng_niederreiter_2 . . . . . . . . . . . . . . . . . . 175

gsl_qrng_size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 gsl_qrng_sobol . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 gsl_qrng_state . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 gsl_ran_bernoulli . . . . . . . . . . . . . . . . . . . . . . . . . 206 gsl_ran_bernoulli_pdf . . . . . . . . . . . . . . . . . . . . 206 gsl_ran_beta . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 gsl_ran_beta_pdf . . . . . . . . . . . . . . . . . . . . . . . . . . 196 gsl_ran_binomial . . . . . . . . . . . . . . . . . . . . . . . . . . 207 gsl_ran_binomial_pdf. . . . . . . . . . . . . . . . . . . . . . 207

                                                gsl_ran_bivariate_gaussian . . . . . . . . . . . . . . . 180
                                                gsl_ran_bivariate_gaussian_pdf . . . . . . . . . . . 180

gsl_ran_cauchy . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 gsl_ran_cauchy_pdf. . . . . . . . . . . . . . . . . . . . . . . . 184 gsl_ran_chisq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193 gsl_ran_chisq_pdf . . . . . . . . . . . . . . . . . . . . . . . . . 193 gsl_ran_choose . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 gsl_ran_dir_2d . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199

                                                gsl_ran_dir_2d_trig_method . . . . . . . . . . . . . . . 199

gsl_ran_dir_3d . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 gsl_ran_dir_nd . . . . . . . . . . . . . . . . . . . . . . . . . . . . 199 gsl_ran_discrete . . . . . . . . . . . . . . . . . . . . . . . . . . 203 gsl_ran_discrete_free . . . . . . . . . . . . . . . . . . . . 204 gsl_ran_discrete_pdf. . . . . . . . . . . . . . . . . . . . . . 204

                                                gsl_ran_discrete_preproc . . . . . . . . . . . . . . . . . 203

gsl_ran_exponential. . . . . . . . . . . . . . . . . . . . . . . 181

                                                gsl_ran_exponential_pdf . . . . . . . . . . . . . . . . . . 181

gsl_ran_exppow . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183 gsl_ran_exppow_pdf. . . . . . . . . . . . . . . . . . . . . . . . 183 gsl_ran_fdist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194 gsl_ran_fdist_pdf . . . . . . . . . . . . . . . . . . . . . . . . . 194 gsl_ran_flat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 191 gsl_ran_flat_pdf . . . . . . . . . . . . . . . . . . . . . . . . . . 191 gsl_ran_gamma . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190 gsl_ran_gamma_pdf . . . . . . . . . . . . . . . . . . . . . . . . . 190 gsl_ran_gaussian . . . . . . . . . . . . . . . . . . . . . . . . . . 178 gsl_ran_gaussian_pdf. . . . . . . . . . . . . . . . . . . . . . 178

                                                gsl_ran_gaussian_ratio_method . . . . . . . . . . . . 178

gsl_ran_gaussian_tail . . . . . . . . . . . . . . . . . . . . 179

                                                gsl_ran_gaussian_tail_pdf . . . . . . . . . . . . . . . . 179

gsl_ran_geometric . . . . . . . . . . . . . . . . . . . . . . . . . 210 gsl_ran_geometric_pdf . . . . . . . . . . . . . . . . . . . . 210 gsl_ran_gumbel1 . . . . . . . . . . . . . . . . . . . . . . . . . . . 201 gsl_ran_gumbel1_pdf. . . . . . . . . . . . . . . . . . . . . . . 201 gsl_ran_gumbel2 . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 gsl_ran_gumbel2_pdf. . . . . . . . . . . . . . . . . . . . . . . 202

                                                gsl_ran_hypergeometric . . . . . . . . . . . . . . . . . . . 211
                                                gsl_ran_hypergeometric_pdf . . . . . . . . . . . . . . . 211

gsl_ran_landau . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 gsl_ran_landau_pdf. . . . . . . . . . . . . . . . . . . . . . . . 187 gsl_ran_laplace . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 gsl_ran_laplace_pdf. . . . . . . . . . . . . . . . . . . . . . . 182 gsl_ran_levy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 188 gsl_ran_levy_skew . . . . . . . . . . . . . . . . . . . . . . . . . 189 gsl_ran_logarithmic. . . . . . . . . . . . . . . . . . . . . . . 212

gsl_ran_logarithmic_pdf . . . . . . . . . . . . . . . . . . 212

gsl_ran_logistic . . . . . . . . . . . . . . . . . . . . . . . . . . 197 gsl_ran_logistic_pdf. . . . . . . . . . . . . . . . . . . . . . 197 gsl_ran_lognormal . . . . . . . . . . . . . . . . . . . . . . . . . 192 gsl_ran_lognormal_pdf . . . . . . . . . . . . . . . . . . . . 192

gsl_ran_negative_binomial . . . . . . . . . . . . . . . . 208
gsl_ran_negative_binomial_pdf . . . . . . . . . . . . 208

gsl_ran_pareto . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 gsl_ran_pareto_pdf. . . . . . . . . . . . . . . . . . . . . . . . 198 gsl_ran_pascal . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 gsl_ran_pascal_pdf. . . . . . . . . . . . . . . . . . . . . . . . 209 gsl_ran_poisson . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 gsl_ran_poisson_pdf. . . . . . . . . . . . . . . . . . . . . . . 205 gsl_ran_rayleigh . . . . . . . . . . . . . . . . . . . . . . . . . . 185 gsl_ran_rayleigh_pdf. . . . . . . . . . . . . . . . . . . . . . 185 gsl_ran_rayleigh_tail . . . . . . . . . . . . . . . . . . . . 186

gsl_ran_rayleigh_tail_pdf . . . . . . . . . . . . . . . . 186

gsl_ran_sample . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 gsl_ran_shuffle . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 gsl_ran_tdist . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 gsl_ran_tdist_pdf . . . . . . . . . . . . . . . . . . . . . . . . . 195 gsl_ran_ugaussian . . . . . . . . . . . . . . . . . . . . . . . . . 178 gsl_ran_ugaussian_pdf . . . . . . . . . . . . . . . . . . . . 178

gsl_ran_ugaussian_ratio_method . . . . . . . . . . . 178
gsl_ran_ugaussian_tail . . . . . . . . . . . . . . . . . . . 179
gsl_ran_ugaussian_tail_pdf . . . . . . . . . . . . . . . 179

gsl_ran_weibull . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 gsl_ran_weibull_pdf. . . . . . . . . . . . . . . . . . . . . . . 200 GSL_REAL. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 gsl_rng_alloc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 gsl_rng_clone . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 gsl_rng_cmrg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 gsl_rng_env_setup . . . . . . . . . . . . . . . . . . . . . . . . . 163 gsl_rng_free . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 gsl_rng_get . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 gsl_rng_gfsr4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 gsl_rng_max . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 gsl_rng_memcpy . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 gsl_rng_min . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 gsl_rng_minstd . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 gsl_rng_mrg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 gsl_rng_mt19937 . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 gsl_rng_name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 gsl_rng_print_state. . . . . . . . . . . . . . . . . . . . . . . 165 gsl_rng_r250 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 gsl_rng_ran0 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 gsl_rng_ran1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 gsl_rng_ran2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 gsl_rng_ran3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 gsl_rng_rand . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 168 gsl_rng_rand48 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 gsl_rng_random_bsd. . . . . . . . . . . . . . . . . . . . . . . . 168 gsl_rng_random_glibc2 . . . . . . . . . . . . . . . . . . . . 168 gsl_rng_random_libc5. . . . . . . . . . . . . . . . . . . . . . 168 gsl_rng_randu . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 gsl_rng_ranf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 gsl_rng_ranlux . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 gsl_rng_ranlux389 . . . . . . . . . . . . . . . . . . . . . . . . . 166 gsl_rng_ranlxd1 . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 gsl_rng_ranlxd2 . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 gsl_rng_ranlxs0 . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 gsl_rng_ranlxs1 . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 gsl_rng_ranlxs2 . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 gsl_rng_ranmar . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 gsl_rng_set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 gsl_rng_size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 gsl_rng_slatec . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 gsl_rng_state . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 163 gsl_rng_taus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 gsl_rng_transputer. . . . . . . . . . . . . . . . . . . . . . . . 171 gsl_rng_tt800 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 gsl_rng_types_setup. . . . . . . . . . . . . . . . . . . . . . . 163 gsl_rng_uni . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 gsl_rng_uni32 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 gsl_rng_uniform . . . . . . . . . . . . . . . . . . . . . . . . . . . 162 gsl_rng_uniform_int. . . . . . . . . . . . . . . . . . . . . . . 162 gsl_rng_uniform_pos. . . . . . . . . . . . . . . . . . . . . . . 162 gsl_rng_vax . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 gsl_rng_zuf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172

                                                gsl_root_fdfsolver_alloc . . . . . . . . . . . . . . . . . 297
                                                gsl_root_fdfsolver_free . . . . . . . . . . . . . . . . . . 297
                                                gsl_root_fdfsolver_iterate . . . . . . . . . . . . . . . 300
                                                gsl_root_fdfsolver_name . . . . . . . . . . . . . . . . . . 297
                                                gsl_root_fdfsolver_newton . . . . . . . . . . . . . . . . 303
                                                gsl_root_fdfsolver_root . . . . . . . . . . . . . . . . . . 300
                                                gsl_root_fdfsolver_secant . . . . . . . . . . . . . . . . 303
                                                gsl_root_fdfsolver_set . . . . . . . . . . . . . . . . . . . 297
                                                gsl_root_fdfsolver_steffenson . . . . . . . . . . . . 304
                                                gsl_root_fsolver_alloc . . . . . . . . . . . . . . . . . . . 297
                                                gsl_root_fsolver_bisection . . . . . . . . . . . . . . . 302
                                                gsl_root_fsolver_brent . . . . . . . . . . . . . . . . . . . 302
                                                gsl_root_fsolver_falsepos . . . . . . . . . . . . . . . . 302

gsl_root_fsolver_free . . . . . . . . . . . . . . . . . . . . 297

                                                gsl_root_fsolver_iterate . . . . . . . . . . . . . . . . . 300

gsl_root_fsolver_name . . . . . . . . . . . . . . . . . . . . 297 gsl_root_fsolver_root . . . . . . . . . . . . . . . . . . . . 300 gsl_root_fsolver_set. . . . . . . . . . . . . . . . . . . . . . 297

                                                gsl_root_fsolver_x_lower . . . . . . . . . . . . . . . . . 300
                                                gsl_root_fsolver_x_upper . . . . . . . . . . . . . . . . . 300

gsl_root_test_delta. . . . . . . . . . . . . . . . . . . . . . . 301

                                                gsl_root_test_interval . . . . . . . . . . . . . . . . . . . 301

gsl_root_test_residual . . . . . . . . . . . . . . . . . . . 301

GSL_SET_COMPLEX . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 gsl_set_error_handler . . . . . . . . . . . . . . . . . . . . . 12

gsl_set_error_handler_off . . . . . . . . . . . . . . . . . 12

GSL_SET_IMAG . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 GSL_SET_REAL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 gsl_sf_airy_Ai . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 gsl_sf_airy_Ai_deriv. . . . . . . . . . . . . . . . . . . . . . . 32

gsl_sf_airy_Ai_deriv_e . . . . . . . . . . . . . . . . . . . . 32
gsl_sf_airy_Ai_deriv_scaled . . . . . . . . . . . . . . . 32
gsl_sf_airy_Ai_deriv_scaled_e . . . . . . . . . . . . . 32

gsl_sf_airy_Ai_e . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 gsl_sf_airy_Ai_scaled . . . . . . . . . . . . . . . . . . . . . 32

gsl_sf_airy_Ai_scaled_e . . . . . . . . . . . . . . . . . . . 32

gsl_sf_airy_Bi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 gsl_sf_airy_Bi_deriv. . . . . . . . . . . . . . . . . . . . . . . 32

gsl_sf_airy_Bi_deriv_e . . . . . . . . . . . . . . . . . . . . 32
gsl_sf_airy_Bi_deriv_scaled . . . . . . . . . . . . . . . 32
gsl_sf_airy_Bi_deriv_scaled_e . . . . . . . . . . . . . 32

gsl_sf_airy_Bi_e . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 gsl_sf_airy_Bi_scaled . . . . . . . . . . . . . . . . . . . . . 32

gsl_sf_airy_Bi_scaled_e . . . . . . . . . . . . . . . . . . . 32

gsl_sf_airy_zero_Ai. . . . . . . . . . . . . . . . . . . . . . . . 33

gsl_sf_airy_zero_Ai_deriv . . . . . . . . . . . . . . . . . 33
gsl_sf_airy_zero_Ai_deriv_e . . . . . . . . . . . . . . . 33

gsl_sf_airy_zero_Ai_e . . . . . . . . . . . . . . . . . . . . . 33 gsl_sf_airy_zero_Bi. . . . . . . . . . . . . . . . . . . . . . . . 33

gsl_sf_airy_zero_Bi_deriv . . . . . . . . . . . . . . . . . 33
gsl_sf_airy_zero_Bi_deriv_e . . . . . . . . . . . . . . . 33

gsl_sf_airy_zero_Bi_e . . . . . . . . . . . . . . . . . . . . . 33

gsl_sf_angle_restrict_pos . . . . . . . . . . . . . . . . . 67
gsl_sf_angle_restrict_pos_e . . . . . . . . . . . . . . . 67
gsl_sf_angle_restrict_symm . . . . . . . . . . . . . . . . 67
gsl_sf_angle_restrict_symm_e . . . . . . . . . . . . . . 67

gsl_sf_atanint . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 gsl_sf_atanint_e . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 gsl_sf_bessel_I0 . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 gsl_sf_bessel_I0_e . . . . . . . . . . . . . . . . . . . . . . . . . 34

gsl_sf_bessel_I0_scaled . . . . . . . . . . . . . . . . . . . 35
gsl_sf_bessel_i0_scaled . . . . . . . . . . . . . . . . . . . 38
gsl_sf_bessel_I0_scaled_e . . . . . . . . . . . . . . . . . 35
gsl_sf_bessel_i0_scaled_e . . . . . . . . . . . . . . . . . 38

gsl_sf_bessel_I1 . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 gsl_sf_bessel_I1_e . . . . . . . . . . . . . . . . . . . . . . . . . 34

gsl_sf_bessel_I1_scaled . . . . . . . . . . . . . . . . . . . 35
gsl_sf_bessel_i1_scaled . . . . . . . . . . . . . . . . . . . 38
gsl_sf_bessel_I1_scaled_e . . . . . . . . . . . . . . . . . 35
gsl_sf_bessel_i1_scaled_e . . . . . . . . . . . . . . . . . 38
gsl_sf_bessel_i2_scaled . . . . . . . . . . . . . . . . . . . 38
gsl_sf_bessel_i2_scaled_e . . . . . . . . . . . . . . . . . 38
gsl_sf_bessel_il_scaled . . . . . . . . . . . . . . . . . . . 38
gsl_sf_bessel_il_scaled_array . . . . . . . . . . . . . 38
gsl_sf_bessel_il_scaled_e . . . . . . . . . . . . . . . . . 38

gsl_sf_bessel_In . . . . . . . . . . . . . . . . . . . . . . . . . . . 34

                                                gsl_sf_bessel_In_array . . . . . . . . . . . . . . . . . . . . 35

gsl_sf_bessel_In_e . . . . . . . . . . . . . . . . . . . . . . . . . 34

                                                gsl_sf_bessel_In_scaled . . . . . . . . . . . . . . . . . . . 35
                                                gsl_sf_bessel_In_scaled_array . . . . . . . . . . . . . 35
                                                gsl_sf_bessel_In_scaled_e . . . . . . . . . . . . . . . . . 35

gsl_sf_bessel_Inu . . . . . . . . . . . . . . . . . . . . . . . . . . 40 gsl_sf_bessel_Inu_e. . . . . . . . . . . . . . . . . . . . . . . . 40

                                                gsl_sf_bessel_Inu_scaled . . . . . . . . . . . . . . . . . . 40
                                                gsl_sf_bessel_Inu_scaled_e . . . . . . . . . . . . . . . . 40

gsl_sf_bessel_J0 . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 gsl_sf_bessel_j0 . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 gsl_sf_bessel_J0_e . . . . . . . . . . . . . . . . . . . . . . . . . 33 gsl_sf_bessel_j0_e . . . . . . . . . . . . . . . . . . . . . . . . . 36 gsl_sf_bessel_J1 . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 gsl_sf_bessel_j1 . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 gsl_sf_bessel_J1_e . . . . . . . . . . . . . . . . . . . . . . . . . 33 gsl_sf_bessel_j1_e . . . . . . . . . . . . . . . . . . . . . . . . . 36 gsl_sf_bessel_j2 . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 gsl_sf_bessel_j2_e . . . . . . . . . . . . . . . . . . . . . . . . . 36 gsl_sf_bessel_jl . . . . . . . . . . . . . . . . . . . . . . . . . . . 37

                                                gsl_sf_bessel_jl_array . . . . . . . . . . . . . . . . . . . . 37

gsl_sf_bessel_jl_e . . . . . . . . . . . . . . . . . . . . . . . . . 37

                                                gsl_sf_bessel_jl_steed_array . . . . . . . . . . . . . . 37

gsl_sf_bessel_Jn . . . . . . . . . . . . . . . . . . . . . . . . . . . 33

                                                gsl_sf_bessel_Jn_array . . . . . . . . . . . . . . . . . . . . 34

gsl_sf_bessel_Jn_e . . . . . . . . . . . . . . . . . . . . . . . . . 33 gsl_sf_bessel_Jnu . . . . . . . . . . . . . . . . . . . . . . . . . . 39 gsl_sf_bessel_Jnu_e. . . . . . . . . . . . . . . . . . . . . . . . 39 gsl_sf_bessel_K0 . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 gsl_sf_bessel_K0_e . . . . . . . . . . . . . . . . . . . . . . . . . 35

                                                gsl_sf_bessel_K0_scaled . . . . . . . . . . . . . . . . . . . 36
                                                gsl_sf_bessel_k0_scaled . . . . . . . . . . . . . . . . . . . 38
                                                gsl_sf_bessel_K0_scaled_e . . . . . . . . . . . . . . . . . 36
                                                gsl_sf_bessel_k0_scaled_e . . . . . . . . . . . . . . . . . 38

gsl_sf_bessel_K1 . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 gsl_sf_bessel_K1_e . . . . . . . . . . . . . . . . . . . . . . . . . 35

                                                gsl_sf_bessel_K1_scaled . . . . . . . . . . . . . . . . . . . 36
                                                gsl_sf_bessel_k1_scaled . . . . . . . . . . . . . . . . . . . 38
                                                gsl_sf_bessel_K1_scaled_e . . . . . . . . . . . . . . . . . 36
                                                gsl_sf_bessel_k1_scaled_e . . . . . . . . . . . . . . . . . 38
                                                gsl_sf_bessel_k2_scaled . . . . . . . . . . . . . . . . . . . 39
                                                gsl_sf_bessel_k2_scaled_e . . . . . . . . . . . . . . . . . 39
                                                gsl_sf_bessel_kl_scaled . . . . . . . . . . . . . . . . . . . 39
                                                gsl_sf_bessel_kl_scaled_array . . . . . . . . . . . . . 39
                                                gsl_sf_bessel_kl_scaled_e . . . . . . . . . . . . . . . . . 39

gsl_sf_bessel_Kn . . . . . . . . . . . . . . . . . . . . . . . . . . . 35

                                                gsl_sf_bessel_Kn_array . . . . . . . . . . . . . . . . . . . . 36

gsl_sf_bessel_Kn_e . . . . . . . . . . . . . . . . . . . . . . . . . 35

                                                gsl_sf_bessel_Kn_scaled . . . . . . . . . . . . . . . . . . . 36
                                                gsl_sf_bessel_Kn_scaled_array . . . . . . . . . . . . . 36
                                                gsl_sf_bessel_Kn_scaled_e . . . . . . . . . . . . . . . . . 36

gsl_sf_bessel_Knu . . . . . . . . . . . . . . . . . . . . . . . . . . 40 gsl_sf_bessel_Knu_e. . . . . . . . . . . . . . . . . . . . . . . . 40

                                                gsl_sf_bessel_Knu_scaled . . . . . . . . . . . . . . . . . . 40

gsl_sf_bessel_Knu_scaled_e . . . . . . . . . . . . . . . . 40

gsl_sf_bessel_lnKnu. . . . . . . . . . . . . . . . . . . . . . . . 40 gsl_sf_bessel_lnKnu_e . . . . . . . . . . . . . . . . . . . . . 40

gsl_sf_bessel_sequence_Jnu_e . . . . . . . . . . . . . . 39

gsl_sf_bessel_Y0 . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 gsl_sf_bessel_y0 . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 gsl_sf_bessel_Y0_e . . . . . . . . . . . . . . . . . . . . . . . . . 34 gsl_sf_bessel_y0_e . . . . . . . . . . . . . . . . . . . . . . . . . 37 gsl_sf_bessel_Y1 . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 gsl_sf_bessel_y1 . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 gsl_sf_bessel_Y1_e . . . . . . . . . . . . . . . . . . . . . . . . . 34 gsl_sf_bessel_y1_e . . . . . . . . . . . . . . . . . . . . . . . . . 37 gsl_sf_bessel_y2 . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 gsl_sf_bessel_y2_e . . . . . . . . . . . . . . . . . . . . . . . . . 37 gsl_sf_bessel_yl . . . . . . . . . . . . . . . . . . . . . . . . . . . 37

gsl_sf_bessel_yl_array . . . . . . . . . . . . . . . . . . . . 37

gsl_sf_bessel_yl_e . . . . . . . . . . . . . . . . . . . . . . . . . 37 gsl_sf_bessel_Yn . . . . . . . . . . . . . . . . . . . . . . . . . . . 34

gsl_sf_bessel_Yn_array . . . . . . . . . . . . . . . . . . . . 34

gsl_sf_bessel_Yn_e . . . . . . . . . . . . . . . . . . . . . . . . . 34 gsl_sf_bessel_Ynu . . . . . . . . . . . . . . . . . . . . . . . . . . 39 gsl_sf_bessel_Ynu_e. . . . . . . . . . . . . . . . . . . . . . . . 39 gsl_sf_bessel_zero_J0 . . . . . . . . . . . . . . . . . . . . . 40

gsl_sf_bessel_zero_J0_e . . . . . . . . . . . . . . . . . . . 40

gsl_sf_bessel_zero_J1 . . . . . . . . . . . . . . . . . . . . . 40

gsl_sf_bessel_zero_J1_e . . . . . . . . . . . . . . . . . . . 40
gsl_sf_bessel_zero_Jnu . . . . . . . . . . . . . . . . . . . . 41
gsl_sf_bessel_zero_Jnu_e . . . . . . . . . . . . . . . . . . 41

gsl_sf_beta . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 gsl_sf_beta_e . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 gsl_sf_beta_inc . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 gsl_sf_beta_inc_e . . . . . . . . . . . . . . . . . . . . . . . . . . 56 gsl_sf_Chi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 gsl_sf_Chi_e . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 gsl_sf_choose . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 gsl_sf_choose_e . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 gsl_sf_Ci . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 gsl_sf_Ci_e . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 gsl_sf_clausen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 gsl_sf_clausen_e . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 gsl_sf_complex_cos_e. . . . . . . . . . . . . . . . . . . . . . . 66

gsl_sf_complex_dilog_e . . . . . . . . . . . . . . . . . . . . 45

gsl_sf_complex_log_e. . . . . . . . . . . . . . . . . . . . . . . 63

gsl_sf_complex_logsin_e . . . . . . . . . . . . . . . . . . . 66

gsl_sf_complex_sin_e. . . . . . . . . . . . . . . . . . . . . . . 66 gsl_sf_conicalP_0 . . . . . . . . . . . . . . . . . . . . . . . . . . 62 gsl_sf_conicalP_0_e. . . . . . . . . . . . . . . . . . . . . . . . 62 gsl_sf_conicalP_1 . . . . . . . . . . . . . . . . . . . . . . . . . . 62 gsl_sf_conicalP_1_e. . . . . . . . . . . . . . . . . . . . . . . . 62

gsl_sf_conicalP_cyl_reg . . . . . . . . . . . . . . . . . . . 62
gsl_sf_conicalP_cyl_reg_e . . . . . . . . . . . . . . . . . 62

gsl_sf_conicalP_half. . . . . . . . . . . . . . . . . . . . . . . 62

gsl_sf_conicalP_half_e . . . . . . . . . . . . . . . . . . . . 62

gsl_sf_conicalP_mhalf . . . . . . . . . . . . . . . . . . . . . 62

                                                gsl_sf_conicalP_mhalf_e . . . . . . . . . . . . . . . . . . . 62
                                                gsl_sf_conicalP_sph_reg . . . . . . . . . . . . . . . . . . . 62
                                                gsl_sf_conicalP_sph_reg_e . . . . . . . . . . . . . . . . . 62

gsl_sf_cos . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 gsl_sf_cos_e . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 gsl_sf_cos_err . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 gsl_sf_cos_err_e . . . . . . . . . . . . . . . . . . . . . . . . . . . 67

                                                gsl_sf_coulomb_CL_array . . . . . . . . . . . . . . . . . . . 43

gsl_sf_coulomb_CL_e. . . . . . . . . . . . . . . . . . . . . . . . 43

                                                gsl_sf_coulomb_wave_F_array . . . . . . . . . . . . . . . 42
                                                gsl_sf_coulomb_wave_FG_array . . . . . . . . . . . . . . 42
                                                gsl_sf_coulomb_wave_FG_e . . . . . . . . . . . . . . . . . . 42
                                                gsl_sf_coulomb_wave_FGp_array . . . . . . . . . . . . . 42
                                                gsl_sf_coulomb_wave_sphF_array . . . . . . . . . . . . 42

gsl_sf_coupling_3j . . . . . . . . . . . . . . . . . . . . . . . . . 43 gsl_sf_coupling_3j_e. . . . . . . . . . . . . . . . . . . . . . . 43 gsl_sf_coupling_6j . . . . . . . . . . . . . . . . . . . . . . . . . 43 gsl_sf_coupling_6j_e. . . . . . . . . . . . . . . . . . . . . . . 43 gsl_sf_coupling_9j . . . . . . . . . . . . . . . . . . . . . . . . . 44 gsl_sf_coupling_9j_e. . . . . . . . . . . . . . . . . . . . . . . 44 gsl_sf_dawson . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 gsl_sf_dawson_e . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 gsl_sf_debye_1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 gsl_sf_debye_1_e . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 gsl_sf_debye_2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 gsl_sf_debye_2_e . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 gsl_sf_debye_3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 gsl_sf_debye_3_e . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 gsl_sf_debye_4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 gsl_sf_debye_4_e . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 gsl_sf_dilog . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 gsl_sf_dilog_e . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 gsl_sf_doublefact . . . . . . . . . . . . . . . . . . . . . . . . . . 55 gsl_sf_doublefact_e. . . . . . . . . . . . . . . . . . . . . . . . 55 gsl_sf_ellint_D . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 gsl_sf_ellint_D_e . . . . . . . . . . . . . . . . . . . . . . . . . . 47 gsl_sf_ellint_E . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 gsl_sf_ellint_E_e . . . . . . . . . . . . . . . . . . . . . . . . . . 47 gsl_sf_ellint_Ecomp. . . . . . . . . . . . . . . . . . . . . . . . 46 gsl_sf_ellint_Ecomp_e . . . . . . . . . . . . . . . . . . . . . 46 gsl_sf_ellint_F . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 gsl_sf_ellint_F_e . . . . . . . . . . . . . . . . . . . . . . . . . . 47 gsl_sf_ellint_Kcomp. . . . . . . . . . . . . . . . . . . . . . . . 46 gsl_sf_ellint_Kcomp_e . . . . . . . . . . . . . . . . . . . . . 46 gsl_sf_ellint_P . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 gsl_sf_ellint_P_e . . . . . . . . . . . . . . . . . . . . . . . . . . 47 gsl_sf_ellint_RC . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 gsl_sf_ellint_RC_e . . . . . . . . . . . . . . . . . . . . . . . . . 47 gsl_sf_ellint_RD . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 gsl_sf_ellint_RD_e . . . . . . . . . . . . . . . . . . . . . . . . . 47 gsl_sf_ellint_RF . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 gsl_sf_ellint_RF_e . . . . . . . . . . . . . . . . . . . . . . . . . 48 gsl_sf_ellint_RJ . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 gsl_sf_ellint_RJ_e . . . . . . . . . . . . . . . . . . . . . . . . . 48 gsl_sf_elljac_e . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 gsl_sf_erf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 gsl_sf_erf_e . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 gsl_sf_erf_Q . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 gsl_sf_erf_Q_e . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 gsl_sf_erf_Z . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 gsl_sf_erf_Z_e . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 gsl_sf_erfc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 gsl_sf_erfc_e . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 gsl_sf_eta . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 gsl_sf_eta_e . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 gsl_sf_eta_int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 gsl_sf_eta_int_e . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 gsl_sf_exp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 gsl_sf_exp_e . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 gsl_sf_exp_e10_e . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 gsl_sf_exp_err_e . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 gsl_sf_exp_err_e10_e. . . . . . . . . . . . . . . . . . . . . . . 50 gsl_sf_exp_mult . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 gsl_sf_exp_mult_e . . . . . . . . . . . . . . . . . . . . . . . . . . 49 gsl_sf_exp_mult_e10_e . . . . . . . . . . . . . . . . . . . . . 49 gsl_sf_exp_mult_err_e . . . . . . . . . . . . . . . . . . . . . 50

gsl_sf_exp_mult_err_e10_e . . . . . . . . . . . . . . . . . 51

gsl_sf_expint_3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 gsl_sf_expint_3_e . . . . . . . . . . . . . . . . . . . . . . . . . . 52 gsl_sf_expint_E1 . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 gsl_sf_expint_E1_e . . . . . . . . . . . . . . . . . . . . . . . . . 51 gsl_sf_expint_E2 . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 gsl_sf_expint_E2_e . . . . . . . . . . . . . . . . . . . . . . . . . 51 gsl_sf_expint_Ei . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 gsl_sf_expint_Ei_e . . . . . . . . . . . . . . . . . . . . . . . . . 51 gsl_sf_expm1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 gsl_sf_expm1_e . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 gsl_sf_exprel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 gsl_sf_exprel_2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 gsl_sf_exprel_2_e . . . . . . . . . . . . . . . . . . . . . . . . . . 50 gsl_sf_exprel_e . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 gsl_sf_exprel_n . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 gsl_sf_exprel_n_e . . . . . . . . . . . . . . . . . . . . . . . . . . 50 gsl_sf_fact . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 gsl_sf_fact_e . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 gsl_sf_fermi_dirac_0. . . . . . . . . . . . . . . . . . . . . . . 52

gsl_sf_fermi_dirac_0_e . . . . . . . . . . . . . . . . . . . . 52

gsl_sf_fermi_dirac_1. . . . . . . . . . . . . . . . . . . . . . . 52

gsl_sf_fermi_dirac_1_e . . . . . . . . . . . . . . . . . . . . 52

gsl_sf_fermi_dirac_2. . . . . . . . . . . . . . . . . . . . . . . 53

gsl_sf_fermi_dirac_2_e . . . . . . . . . . . . . . . . . . . . 53
gsl_sf_fermi_dirac_3half . . . . . . . . . . . . . . . . . . 53
gsl_sf_fermi_dirac_3half_e . . . . . . . . . . . . . . . . 53
gsl_sf_fermi_dirac_half . . . . . . . . . . . . . . . . . . . 53
gsl_sf_fermi_dirac_half_e . . . . . . . . . . . . . . . . . 53
gsl_sf_fermi_dirac_inc_0 . . . . . . . . . . . . . . . . . . 53
gsl_sf_fermi_dirac_inc_0_e . . . . . . . . . . . . . . . . 53
gsl_sf_fermi_dirac_int . . . . . . . . . . . . . . . . . . . . 53

                                                gsl_sf_fermi_dirac_int_e . . . . . . . . . . . . . . . . . . 53

gsl_sf_fermi_dirac_m1 . . . . . . . . . . . . . . . . . . . . . 52

                                                gsl_sf_fermi_dirac_m1_e . . . . . . . . . . . . . . . . . . . 52
                                                gsl_sf_fermi_dirac_mhalf . . . . . . . . . . . . . . . . . . 53
                                                gsl_sf_fermi_dirac_mhalf_e . . . . . . . . . . . . . . . . 53

gsl_sf_gamma . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 gsl_sf_gamma_e . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 gsl_sf_gamma_inc_P . . . . . . . . . . . . . . . . . . . . . . . . . 56 gsl_sf_gamma_inc_P_e. . . . . . . . . . . . . . . . . . . . . . . 56 gsl_sf_gamma_inc_Q . . . . . . . . . . . . . . . . . . . . . . . . . 56 gsl_sf_gamma_inc_Q_e. . . . . . . . . . . . . . . . . . . . . . . 56 gsl_sf_gammainv . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 gsl_sf_gammainv_e . . . . . . . . . . . . . . . . . . . . . . . . . . 54 gsl_sf_gammastar . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 gsl_sf_gammastar_e . . . . . . . . . . . . . . . . . . . . . . . . . 54 gsl_sf_gegenpoly_1 . . . . . . . . . . . . . . . . . . . . . . . . . 57 gsl_sf_gegenpoly_1_e. . . . . . . . . . . . . . . . . . . . . . . 57 gsl_sf_gegenpoly_2 . . . . . . . . . . . . . . . . . . . . . . . . . 57 gsl_sf_gegenpoly_2_e. . . . . . . . . . . . . . . . . . . . . . . 57 gsl_sf_gegenpoly_3 . . . . . . . . . . . . . . . . . . . . . . . . . 57 gsl_sf_gegenpoly_3_e. . . . . . . . . . . . . . . . . . . . . . . 57

                                                gsl_sf_gegenpoly_array . . . . . . . . . . . . . . . . . . . . 57

gsl_sf_gegenpoly_n . . . . . . . . . . . . . . . . . . . . . . . . . 57 gsl_sf_gegenpoly_n_e. . . . . . . . . . . . . . . . . . . . . . . 57 gsl_sf_hydrogenicR . . . . . . . . . . . . . . . . . . . . . . . . . 41 gsl_sf_hydrogenicR_1. . . . . . . . . . . . . . . . . . . . . . . 41

                                                gsl_sf_hydrogenicR_1_e . . . . . . . . . . . . . . . . . . . . 41

gsl_sf_hydrogenicR_e. . . . . . . . . . . . . . . . . . . . . . . 41 gsl_sf_hyperg_0F1 . . . . . . . . . . . . . . . . . . . . . . . . . . 57 gsl_sf_hyperg_0F1_e. . . . . . . . . . . . . . . . . . . . . . . . 57 gsl_sf_hyperg_1F1 . . . . . . . . . . . . . . . . . . . . . . . . . . 58 gsl_sf_hyperg_1F1_e. . . . . . . . . . . . . . . . . . . . . . . . 58 gsl_sf_hyperg_1F1_int . . . . . . . . . . . . . . . . . . . . . 57

                                                gsl_sf_hyperg_1F1_int_e . . . . . . . . . . . . . . . . . . . 57

gsl_sf_hyperg_2F0 . . . . . . . . . . . . . . . . . . . . . . . . . . 59 gsl_sf_hyperg_2F0_e. . . . . . . . . . . . . . . . . . . . . . . . 59 gsl_sf_hyperg_2F1 . . . . . . . . . . . . . . . . . . . . . . . . . . 58

                                                gsl_sf_hyperg_2F1_conj . . . . . . . . . . . . . . . . . . . . 58
                                                gsl_sf_hyperg_2F1_conj_e . . . . . . . . . . . . . . . . . . 58
                                                gsl_sf_hyperg_2F1_conj_renorm . . . . . . . . . . . . . 59
                                                gsl_sf_hyperg_2F1_conj_renorm_e . . . . . . . . . . . 59

gsl_sf_hyperg_2F1_e. . . . . . . . . . . . . . . . . . . . . . . . 58

                                                gsl_sf_hyperg_2F1_renorm . . . . . . . . . . . . . . . . . . 58
                                                gsl_sf_hyperg_2F1_renorm_e . . . . . . . . . . . . . . . . 58

gsl_sf_hyperg_U . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 gsl_sf_hyperg_U_e . . . . . . . . . . . . . . . . . . . . . . . . . . 58 gsl_sf_hyperg_U_e10_e . . . . . . . . . . . . . . . . . . . . . 58 gsl_sf_hyperg_U_int. . . . . . . . . . . . . . . . . . . . . . . . 58 gsl_sf_hyperg_U_int_e . . . . . . . . . . . . . . . . . . . . . 58

                                                gsl_sf_hyperg_U_int_e10_e . . . . . . . . . . . . . . . . . 58

gsl_sf_hypot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 gsl_sf_hypot_e . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 gsl_sf_hzeta . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 gsl_sf_hzeta_e . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 gsl_sf_laguerre_1 . . . . . . . . . . . . . . . . . . . . . . . . . . 59 gsl_sf_laguerre_1_e. . . . . . . . . . . . . . . . . . . . . . . . 59 gsl_sf_laguerre_2 . . . . . . . . . . . . . . . . . . . . . . . . . . 59 gsl_sf_laguerre_2_e. . . . . . . . . . . . . . . . . . . . . . . . 59 gsl_sf_laguerre_3 . . . . . . . . . . . . . . . . . . . . . . . . . . 59 gsl_sf_laguerre_3_e. . . . . . . . . . . . . . . . . . . . . . . . 59 gsl_sf_laguerre_n . . . . . . . . . . . . . . . . . . . . . . . . . . 59 gsl_sf_laguerre_n_e. . . . . . . . . . . . . . . . . . . . . . . . 59 gsl_sf_lambert_W0 . . . . . . . . . . . . . . . . . . . . . . . . . . 60 gsl_sf_lambert_W0_e. . . . . . . . . . . . . . . . . . . . . . . . 60 gsl_sf_lambert_Wm1 . . . . . . . . . . . . . . . . . . . . . . . . . 60 gsl_sf_lambert_Wm1_e. . . . . . . . . . . . . . . . . . . . . . . 60

gsl_sf_legendre_array_size . . . . . . . . . . . . . . . . 61

gsl_sf_legendre_H3d. . . . . . . . . . . . . . . . . . . . . . . . 63 gsl_sf_legendre_H3d_0 . . . . . . . . . . . . . . . . . . . . . 63

gsl_sf_legendre_H3d_0_e . . . . . . . . . . . . . . . . . . . 63

gsl_sf_legendre_H3d_1 . . . . . . . . . . . . . . . . . . . . . 63

gsl_sf_legendre_H3d_1_e . . . . . . . . . . . . . . . . . . . 63
gsl_sf_legendre_H3d_array . . . . . . . . . . . . . . . . . 63

gsl_sf_legendre_H3d_e . . . . . . . . . . . . . . . . . . . . . 63 gsl_sf_legendre_P1 . . . . . . . . . . . . . . . . . . . . . . . . . 60 gsl_sf_legendre_P1_e. . . . . . . . . . . . . . . . . . . . . . . 60 gsl_sf_legendre_P2 . . . . . . . . . . . . . . . . . . . . . . . . . 60 gsl_sf_legendre_P2_e. . . . . . . . . . . . . . . . . . . . . . . 60 gsl_sf_legendre_P3 . . . . . . . . . . . . . . . . . . . . . . . . . 60 gsl_sf_legendre_P3_e. . . . . . . . . . . . . . . . . . . . . . . 60 gsl_sf_legendre_Pl . . . . . . . . . . . . . . . . . . . . . . . . . 60

gsl_sf_legendre_Pl_array . . . . . . . . . . . . . . . . . . 60

gsl_sf_legendre_Pl_e. . . . . . . . . . . . . . . . . . . . . . . 60 gsl_sf_legendre_Plm. . . . . . . . . . . . . . . . . . . . . . . . 61

gsl_sf_legendre_Plm_array . . . . . . . . . . . . . . . . . 61

gsl_sf_legendre_Plm_e . . . . . . . . . . . . . . . . . . . . . 61 gsl_sf_legendre_Q0 . . . . . . . . . . . . . . . . . . . . . . . . . 60 gsl_sf_legendre_Q0_e. . . . . . . . . . . . . . . . . . . . . . . 60 gsl_sf_legendre_Q1 . . . . . . . . . . . . . . . . . . . . . . . . . 61 gsl_sf_legendre_Q1_e. . . . . . . . . . . . . . . . . . . . . . . 61 gsl_sf_legendre_Ql . . . . . . . . . . . . . . . . . . . . . . . . . 61 gsl_sf_legendre_Ql_e. . . . . . . . . . . . . . . . . . . . . . . 61

gsl_sf_legendre_sphPlm . . . . . . . . . . . . . . . . . . . . 61
gsl_sf_legendre_sphPlm_array . . . . . . . . . . . . . . 61
gsl_sf_legendre_sphPlm_e . . . . . . . . . . . . . . . . . . 61

gsl_sf_lnbeta . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 gsl_sf_lnbeta_e . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 gsl_sf_lnchoose . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 gsl_sf_lnchoose_e . . . . . . . . . . . . . . . . . . . . . . . . . . 55 gsl_sf_lncosh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 gsl_sf_lncosh_e . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 gsl_sf_lndoublefact. . . . . . . . . . . . . . . . . . . . . . . . 55 gsl_sf_lndoublefact_e . . . . . . . . . . . . . . . . . . . . . 55 gsl_sf_lnfact . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 gsl_sf_lnfact_e . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 gsl_sf_lngamma . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54

gsl_sf_lngamma_complex_e . . . . . . . . . . . . . . . . . . 55

gsl_sf_lngamma_e . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 gsl_sf_lngamma_sgn_e. . . . . . . . . . . . . . . . . . . . . . . 54 gsl_sf_lnpoch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 gsl_sf_lnpoch_e . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 gsl_sf_lnpoch_sgn_e. . . . . . . . . . . . . . . . . . . . . . . . 56 gsl_sf_lnsinh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 gsl_sf_lnsinh_e . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 gsl_sf_log . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 gsl_sf_log_1plusx . . . . . . . . . . . . . . . . . . . . . . . . . . 64 gsl_sf_log_1plusx_e. . . . . . . . . . . . . . . . . . . . . . . . 64 gsl_sf_log_1plusx_mx. . . . . . . . . . . . . . . . . . . . . . . 64

                                                gsl_sf_log_1plusx_mx_e . . . . . . . . . . . . . . . . . . . . 64

gsl_sf_log_abs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 gsl_sf_log_abs_e . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 gsl_sf_log_e . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 gsl_sf_log_erfc . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 gsl_sf_log_erfc_e . . . . . . . . . . . . . . . . . . . . . . . . . . 48 gsl_sf_multiply_e . . . . . . . . . . . . . . . . . . . . . . . . . . 45 gsl_sf_multiply_err_e . . . . . . . . . . . . . . . . . . . . . 45 gsl_sf_poch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 gsl_sf_poch_e . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 gsl_sf_pochrel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 gsl_sf_pochrel_e . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 gsl_sf_polar_to_rect. . . . . . . . . . . . . . . . . . . . . . . 67 gsl_sf_pow_int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 gsl_sf_pow_int_e . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 gsl_sf_psi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 gsl_sf_psi_1_int . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 gsl_sf_psi_1_int_e . . . . . . . . . . . . . . . . . . . . . . . . . 65 gsl_sf_psi_1piy . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 gsl_sf_psi_1piy_e . . . . . . . . . . . . . . . . . . . . . . . . . . 64 gsl_sf_psi_e . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 gsl_sf_psi_int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 gsl_sf_psi_int_e . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 gsl_sf_psi_n . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 gsl_sf_psi_n_e . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 gsl_sf_rect_to_polar. . . . . . . . . . . . . . . . . . . . . . . 67 gsl_sf_Shi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 gsl_sf_Shi_e . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 gsl_sf_Si . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 gsl_sf_Si_e . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 gsl_sf_sin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 gsl_sf_sin_e . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 gsl_sf_sin_err . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 gsl_sf_sin_err_e . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 gsl_sf_sinc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 gsl_sf_sinc_e . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 gsl_sf_synchrotron_1. . . . . . . . . . . . . . . . . . . . . . . 65

                                                gsl_sf_synchrotron_1_e . . . . . . . . . . . . . . . . . . . . 65

gsl_sf_synchrotron_2. . . . . . . . . . . . . . . . . . . . . . . 65

                                                gsl_sf_synchrotron_2_e . . . . . . . . . . . . . . . . . . . . 65

gsl_sf_taylorcoeff . . . . . . . . . . . . . . . . . . . . . . . . . 55 gsl_sf_taylorcoeff_e. . . . . . . . . . . . . . . . . . . . . . . 55 gsl_sf_transport_2 . . . . . . . . . . . . . . . . . . . . . . . . . 65 gsl_sf_transport_2_e. . . . . . . . . . . . . . . . . . . . . . . 65 gsl_sf_transport_3 . . . . . . . . . . . . . . . . . . . . . . . . . 65 gsl_sf_transport_3_e. . . . . . . . . . . . . . . . . . . . . . . 65 gsl_sf_transport_4 . . . . . . . . . . . . . . . . . . . . . . . . . 65 gsl_sf_transport_4_e. . . . . . . . . . . . . . . . . . . . . . . 65 gsl_sf_transport_5 . . . . . . . . . . . . . . . . . . . . . . . . . 65 gsl_sf_transport_5_e. . . . . . . . . . . . . . . . . . . . . . . 65 gsl_sf_zeta . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 gsl_sf_zeta_e . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 gsl_sf_zeta_int . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 gsl_sf_zeta_int_e . . . . . . . . . . . . . . . . . . . . . . . . . . 68 GSL_SIGN. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 gsl_siman_solve . . . . . . . . . . . . . . . . . . . . . . . . . . . 263 gsl_sort . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 gsl_sort_index . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 gsl_sort_largest . . . . . . . . . . . . . . . . . . . . . . . . . . 101

gsl_sort_largest_index . . . . . . . . . . . . . . . . . . . 101

gsl_sort_smallest . . . . . . . . . . . . . . . . . . . . . . . . . 101

gsl_sort_smallest_index . . . . . . . . . . . . . . . . . . 101

gsl_sort_vector . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 gsl_sort_vector_index . . . . . . . . . . . . . . . . . . . . 100

gsl_sort_vector_largest . . . . . . . . . . . . . . . . . . 101
gsl_sort_vector_largest_index . . . . . . . . . . . . 102
gsl_sort_vector_smallest . . . . . . . . . . . . . . . . . 101
gsl_sort_vector_smallest_index . . . . . . . . . . . 102

gsl_spline_alloc . . . . . . . . . . . . . . . . . . . . . . . . . . 281 gsl_spline_eval . . . . . . . . . . . . . . . . . . . . . . . . . . . 281 gsl_spline_eval_deriv . . . . . . . . . . . . . . . . . . . . 281

gsl_spline_eval_deriv_e . . . . . . . . . . . . . . . . . . 281
gsl_spline_eval_deriv2 . . . . . . . . . . . . . . . . . . . 281
gsl_spline_eval_deriv2_e . . . . . . . . . . . . . . . . . 281

gsl_spline_eval_e . . . . . . . . . . . . . . . . . . . . . . . . . 281 gsl_spline_eval_integ . . . . . . . . . . . . . . . . . . . . 282

gsl_spline_eval_integ_e . . . . . . . . . . . . . . . . . . 282

gsl_spline_free . . . . . . . . . . . . . . . . . . . . . . . . . . . 281 gsl_spline_init . . . . . . . . . . . . . . . . . . . . . . . . . . . 281 gsl_stats_absdev . . . . . . . . . . . . . . . . . . . . . . . . . . 218 gsl_stats_absdev_m. . . . . . . . . . . . . . . . . . . . . . . . 218 gsl_stats_covariance. . . . . . . . . . . . . . . . . . . . . . 220

gsl_stats_covariance_m . . . . . . . . . . . . . . . . . . . 220

gsl_stats_kurtosis. . . . . . . . . . . . . . . . . . . . . . . . 219

gsl_stats_kurtosis_m_sd . . . . . . . . . . . . . . . . . . 219
gsl_stats_lag1_autocorrelation . . . . . . . . . . . 220
gsl_stats_lag1_autocorrelation_m. . . . . . . . . 220

gsl_stats_max . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 222 gsl_stats_max_index. . . . . . . . . . . . . . . . . . . . . . . 223 gsl_stats_mean . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217

gsl_stats_median_from_sorted_data . . . . . . . 223

gsl_stats_min . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 222 gsl_stats_min_index. . . . . . . . . . . . . . . . . . . . . . . 223 gsl_stats_minmax . . . . . . . . . . . . . . . . . . . . . . . . . . 223

gsl_stats_minmax_index . . . . . . . . . . . . . . . . . . . 223
gsl_stats_quantile_from_sorted_data . . . . . 223

gsl_stats_sd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218 gsl_stats_sd_m . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218

                                                gsl_stats_sd_with_fixed_mean . . . . . . . . . . . . . 218

gsl_stats_skew . . . . . . . . . . . . . . . . . . . . . . . . . . . . 219 gsl_stats_skew_m_sd. . . . . . . . . . . . . . . . . . . . . . . 219 gsl_stats_variance. . . . . . . . . . . . . . . . . . . . . . . . 217 gsl_stats_variance_m. . . . . . . . . . . . . . . . . . . . . . 217

                                                gsl_stats_variance_with_fixed_mean . . . . . . 218

gsl_stats_wabsdev . . . . . . . . . . . . . . . . . . . . . . . . . 221 gsl_stats_wabsdev_m. . . . . . . . . . . . . . . . . . . . . . . 222 gsl_stats_wkurtosis. . . . . . . . . . . . . . . . . . . . . . . 222

                                                gsl_stats_wkurtosis_m_sd . . . . . . . . . . . . . . . . . 222

gsl_stats_wmean . . . . . . . . . . . . . . . . . . . . . . . . . . . 220 gsl_stats_wsd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 221 gsl_stats_wsd_m . . . . . . . . . . . . . . . . . . . . . . . . . . . 221

                                                gsl_stats_wsd_with_fixed_mean . . . . . . . . . . . . 221

gsl_stats_wskew . . . . . . . . . . . . . . . . . . . . . . . . . . . 222 gsl_stats_wskew_m_sd. . . . . . . . . . . . . . . . . . . . . . 222 gsl_stats_wvariance. . . . . . . . . . . . . . . . . . . . . . . 221 gsl_stats_wvariance_m . . . . . . . . . . . . . . . . . . . . 221

                                                gsl_stats_wvariance_with_fixed_mean . . . . . 221

gsl_strerror . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 gsl_sum_levin_u_accel . . . . . . . . . . . . . . . . . . . . 290 gsl_sum_levin_u_alloc . . . . . . . . . . . . . . . . . . . . 290 gsl_sum_levin_u_free. . . . . . . . . . . . . . . . . . . . . . 290

                                                gsl_sum_levin_utrunc_accel . . . . . . . . . . . . . . . 291
                                                gsl_sum_levin_utrunc_alloc . . . . . . . . . . . . . . . 291
                                                gsl_sum_levin_utrunc_free . . . . . . . . . . . . . . . . 291

gsl_vector_add . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79

                                                gsl_vector_add_constant . . . . . . . . . . . . . . . . . . . 79

gsl_vector_alloc . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 gsl_vector_calloc . . . . . . . . . . . . . . . . . . . . . . . . . . 74

                                                gsl_vector_complex_const_imag . . . . . . . . . . . . . 77
                                                gsl_vector_complex_const_real . . . . . . . . . . . . . 77
                                                gsl_vector_complex_imag . . . . . . . . . . . . . . . . . . . 77
                                                gsl_vector_complex_real . . . . . . . . . . . . . . . . . . . 77
                                                gsl_vector_const_subvector . . . . . . . . . . . . . . . . 76
                                                gsl_vector_const_subvector_with_stride . . . 76
                                                gsl_vector_const_view_array . . . . . . . . . . . . . . . 77
                                                gsl_vector_const_view_array_with_stride . . 78

gsl_vector_div . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 gsl_vector_fprintf . . . . . . . . . . . . . . . . . . . . . . . . . 75 gsl_vector_fread . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 gsl_vector_free . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 gsl_vector_fscanf . . . . . . . . . . . . . . . . . . . . . . . . . . 76 gsl_vector_fwrite . . . . . . . . . . . . . . . . . . . . . . . . . . 75 gsl_vector_get . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 gsl_vector_isnull . . . . . . . . . . . . . . . . . . . . . . . . . . 80 gsl_vector_max . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 gsl_vector_max_index. . . . . . . . . . . . . . . . . . . . . . . 80 gsl_vector_memcpy . . . . . . . . . . . . . . . . . . . . . . . . . . 78 gsl_vector_min . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 gsl_vector_min_index. . . . . . . . . . . . . . . . . . . . . . . 80 gsl_vector_minmax . . . . . . . . . . . . . . . . . . . . . . . . . . 80

                                                gsl_vector_minmax_index . . . . . . . . . . . . . . . . . . . 80

gsl_vector_mul . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 gsl_vector_ptr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 gsl_vector_reverse . . . . . . . . . . . . . . . . . . . . . . . . . 79 gsl_vector_scale . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 gsl_vector_set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 gsl_vector_set_all . . . . . . . . . . . . . . . . . . . . . . . . . 75 gsl_vector_set_basis. . . . . . . . . . . . . . . . . . . . . . . 75 gsl_vector_set_zero. . . . . . . . . . . . . . . . . . . . . . . . 75 gsl_vector_sub . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 gsl_vector_subvector. . . . . . . . . . . . . . . . . . . . . . . 76

                                                gsl_vector_subvector_with_stride. . . . . . . . . . 76

gsl_vector_swap . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78

                                                gsl_vector_swap_elements . . . . . . . . . . . . . . . . . . 79

gsl_vector_view_array . . . . . . . . . . . . . . . . . . . . . 77

                                                gsl_vector_view_array_with_stride. . . . . . . . . 78

Variable Index 421

Variable Index

A

alpha . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 255, 257

C

chisq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257

D

dither. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 255

E

estimate_frac . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 254

I

iterations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257

M

min_calls . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 254

min_calls_per_bisection . . . . . . . . . . . . . . . . . . 254

mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257

O

ostream. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257

R

result. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257

S

sigma . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257 stage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257

V

verbose. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 257

Type Index 422

Type Index

G

gsl_error_handler_t. . . . . . . . . . . . . . . . . . . . . . . . 12

gsl_fft_complex_wavetable . . . . . . . . . . . . . . . . 140

gsl_function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 298 gsl_function_fdf . . . . . . . . . . . . . . . . . . . . . . . . . . 298 gsl_histogram . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227 gsl_histogram_pdf . . . . . . . . . . . . . . . . . . . . . . . . . 233 gsl_histogram2d . . . . . . . . . . . . . . . . . . . . . . . . . . . 236 gsl_histogram2d_pdf. . . . . . . . . . . . . . . . . . . . . . . 242 gsl_monte_function. . . . . . . . . . . . . . . . . . . . . . . . 251 gsl_multifit_function . . . . . . . . . . . . . . . . . . . . 349

gsl_multifit_function_fdf . . . . . . . . . . . . . . . . 350
gsl_multimin_function_fdf . . . . . . . . . . . . . . . . 332
gsl_multiroot_function . . . . . . . . . . . . . . . . . . . 318
gsl_multiroot_function_fdf . . . . . . . . . . . . . . . 319

gsl_odeiv_system . . . . . . . . . . . . . . . . . . . . . . . . . . 270 gsl_siman_Efunc_t . . . . . . . . . . . . . . . . . . . . . . . . . 263 gsl_siman_metric_t. . . . . . . . . . . . . . . . . . . . . . . . 263 gsl_siman_params_t. . . . . . . . . . . . . . . . . . . . . . . . 263 gsl_siman_print_t . . . . . . . . . . . . . . . . . . . . . . . . . 263 gsl_siman_step_t . . . . . . . . . . . . . . . . . . . . . . . . . . 263

Concept Index 423

Concept Index

2

2D histograms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 236 2D random direction vector. . . . . . . . . . . . . . . . . . 199

3

3-j symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 3D random direction vector. . . . . . . . . . . . . . . . . . 199

6

6-j symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43

9

9-j symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43

A

acceleration of series . . . . . . . . . . . . . . . . . . . . . . . . 290 acosh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Adaptive step-size control, differential equations

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 272 Ai(x) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 Airy functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 Akima splines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 279 aliasing of arrays. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 alternative optimized functions . . . . . . . . . . . . . . . . 7 AMAX, Level-1 BLAS . . . . . . . . . . . . . . . . . . . . . . 107 angular reduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 ANSI C, use of . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 Apell symbol, see Pochammer symbol . . . . . . . . . 55 argument of complex number . . . . . . . . . . . . . . . . . 19 arithmetic exceptions . . . . . . . . . . . . . . . . . . . . . . . 368 asinh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 astronomical constants . . . . . . . . . . . . . . . . . . . . . . 360 ASUM, Level-1 BLAS. . . . . . . . . . . . . . . . . . . . . . . 107 atanh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 atomic physics, constants. . . . . . . . . . . . . . . . . . . . 360 autoconf, using with GSL . . . . . . . . . . . . . . . . . . . . . 6 AXPY, Level-1 BLAS . . . . . . . . . . . . . . . . . . . . . . . 108

B

Bader and Deuflhard, Bulirsch-Stoer method.

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 272 Bernoulli trial, random variates . . . . . . . . . . . . . . 206 Bessel functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 Bessel Functions, Fractional Order . . . . . . . . . . . . 39 best-fit parameters, covariance . . . . . . . . . . . . . . . 353 Beta distribution random variates . . . . . . . . . . . . 196 Beta function. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 Beta function, incomplete normalized . . . . . . . . . 56 Bi(x) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 bias, IEEE format . . . . . . . . . . . . . . . . . . . . . . . . . . 366 bidiagonalization of real matrices . . . . . . . . . . . . 126 binning data . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227 Binomial random variates . . . . . . . . . . . . . . . . . . . 207 bisection algorithm for finding roots . . . . . . . . . . 302 BLAS. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 blocks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 breakpoints . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 371 brent’s method for finding minima . . . . . . . . . . . 313 brent’s method for finding roots . . . . . . . . . . . . . 302 Broyden algorithm for multidimensional roots

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 324 BSD random number generator . . . . . . . . . . . . . . 168 BSD random number generator, rand . . . . . . . . 168 Bulirsch-Stoer method . . . . . . . . . . . . . . . . . . . . . . 272

C

C extensions, compatible use of . . . . . . . . . . . . . . . . 4 C++, compatibility . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 Carlson forms of Elliptic integrals . . . . . . . . . . . . . 46 Cash-Karp, Runge-Kutta method . . . . . . . . . . . . 272 Cauchy random variates. . . . . . . . . . . . . . . . . . . . . 184 CBLAS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 CBLAS, Low-level interface . . . . . . . . . . . . . . . . . 378 Chebyshev series. . . . . . . . . . . . . . . . . . . . . . . . . . . . 286 Checkergcc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 checking permutation for validity . . . . . . . . . . . . . 95 Chi-squared random variates . . . . . . . . . . . . . . . . 193 Cholesky decomposition . . . . . . . . . . . . . . . . . . . . . 124 Clausen functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 Clenshaw-Curtis quadrature . . . . . . . . . . . . . . . . . 151 CMRG, combined multiple recursive random

number generator . . . . . . . . . . . . . . . . . . . . . . 166 code reuse in applications . . . . . . . . . . . . . . . . . . . . . 9 combinatorial factor C(m,n) . . . . . . . . . . . . . . . . . . 55 combinatorial optimization . . . . . . . . . . . . . . . . . . 262 combinatorial searches . . . . . . . . . . . . . . . . . . . . . . 262 comparison functions, definition. . . . . . . . . . . . . . . 99 compatibility . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 compiling programs, include paths . . . . . . . . . . . . . 4 compiling programs, library paths . . . . . . . . . . . . . . 4 complementary incomplete Gamma function … 56 complex arithmetic . . . . . . . . . . . . . . . . . . . . . . . . . . 19 complex cosine function, special functions . . . . . 66 Complex Gamma function . . . . . . . . . . . . . . . . . . . . 55 complex hermitian matrix, eigensystem . . . . . . . 131 complex log sine function, special functions …. 66 complex numbers . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 complex sinc function, special functions . . . . . . . 66 complex sine function, special functions . . . . . . . 66 confluent hypergeometric function. . . . . . . . . . . . . 59 confluent hypergeometric functions . . . . . . . . . . . . 57 conical functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 conjugate of complex number . . . . . . . . . . . . . . . . . 20 constants, fundamental. . . . . . . . . . . . . . . . . . . . . . 359 constants, mathematical — defined as macros . . 14 constants, physical . . . . . . . . . . . . . . . . . . . . . . . . . . 359 contacting the GSL developers . . . . . . . . . . . . . . . . . 3 convergence, accelerating a series . . . . . . . . . . . . 290 conversion of units . . . . . . . . . . . . . . . . . . . . . . . . . . 359 cooling schedule . . . . . . . . . . . . . . . . . . . . . . . . . . . . 262 COPY, Level-1 BLAS . . . . . . . . . . . . . . . . . . . . . . . 108 cosine function, special functions . . . . . . . . . . . . . . 66 cosine of complex number . . . . . . . . . . . . . . . . . . . . 22 cost function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 262 Coulomb wave functions. . . . . . . . . . . . . . . . . . . . . . 41 coupling coefficients . . . . . . . . . . . . . . . . . . . . . . . . . . 43 covariance of best-fit parameters . . . . . . . . . . . . . 353 covariance, of two datasets . . . . . . . . . . . . . . . . . . 220 CRAY random number generator, RANF . . . . . 170 cubic equation, solving . . . . . . . . . . . . . . . . . . . . . . . 27 cubic splines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 279 Cylindrical Bessel Functions . . . . . . . . . . . . . . . . . . 33

D

Dawson function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 debugging numerical programs . . . . . . . . . . . . . . . 371 Debye functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 denormalized form, IEEE format. . . . . . . . . . . . . 366 derivatives, calculating numerically. . . . . . . . . . . 284 determinant of a matrix, by LU decomposition

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 Deuflhard and Bader, Bulirsch-Stoer method.

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 272 DFTs, see FFT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 differential equations, initial value problems… 270 differentiation of functions, numeric . . . . . . . . . . 284 digamma function . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 dilogarithm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 direction vector, random 2D . . . . . . . . . . . . . . . . . 199 direction vector, random 3D . . . . . . . . . . . . . . . . . 199 direction vector, random N-dimensional . . . . . . 199 Discrete Fourier Transforms, see FFT . . . . . . . . 135 discrete Hankel transforms . . . . . . . . . . . . . . . . . . 294 Discrete newton algorithm for multidimensional

roots . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 323 Discrete Newton algorithm for nonlinear systems

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 323 Discrete random numbers . . . . . . . . . . . . . . . 203, 204 Discrete random numbers, preprocessing. . . . . . 203 division by zero, IEEE exceptions . . . . . . . . . . . . 368 DOT, Level-1 BLAS . . . . . . . . . . . . . . . . . . . . . . . . 106 double factorial . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 double precision, IEEE format . . . . . . . . . . . . . . . 366 downloading GSL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

E

e, defined as a macro. . . . . . . . . . . . . . . . . . . . . . . . . 14 eigenvalues and eigenvectors . . . . . . . . . . . . . . . . . 130 elementary functions . . . . . . . . . . . . . . . . . . . . . . . . . 14 elementary operations . . . . . . . . . . . . . . . . . . . . . . . . 45 elliptic functions (Jacobi). . . . . . . . . . . . . . . . . . . . . 48 elliptic integrals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 energy function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 262 energy, units of . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 363 erf(x) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 erfc(x) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 error codes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 error function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 Error handlers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 error handling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 error handling macros . . . . . . . . . . . . . . . . . . . . . . . . 13 Errors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 estimated standard deviation . . . . . . . . . . . . . . . . 217 estimated variance . . . . . . . . . . . . . . . . . . . . . . . . . . 217 euclidean distance function, hypot . . . . . . . . . . . . 15 Euler’s constant, defined as a macro . . . . . . . . . . . 14 evaluation of polynomials. . . . . . . . . . . . . . . . . . . . . 26 exceptions, IEEE arithmetic . . . . . . . . . . . . . . . . . 368 exchanging permutation elements . . . . . . . . . . . . . 95 exp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 expm1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 exponent, IEEE format . . . . . . . . . . . . . . . . . . . . . 366 exponential function . . . . . . . . . . . . . . . . . . . . . . . . . 49 exponential integrals . . . . . . . . . . . . . . . . . . . . . . . . . 51 Exponential power distribution, random variates

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 183 Exponential random variates . . . . . . . . . . . . . . . . 181 exponential, difference from 1 computed accurately . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 exponentiation of complex number . . . . . . . . . . . . 21

F

F-distribution random variates. . . . . . . . . . . . . . . 194 factorial. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 factorization of matrices. . . . . . . . . . . . . . . . . . . . . 118 false position algorithm for finding roots . . . . . . 302 Fast Fourier Transforms, see FFT . . . . . . . . . . . . 135 FDL, GNU Free Documentation License . . . . . . 399 Fehlberg method, differential equations . . . . . . . 272 Fermi-Dirac function . . . . . . . . . . . . . . . . . . . . . . . . . 52 FFT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 FFT mathematical definition . . . . . . . . . . . . . . . . 135 FFT of complex data, mixed-radix algorithm. . 139 FFT of complex data, radix-2 algorithm . . . . . . 137 FFT of real data . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 FFT of real data, mixed-radix algorithm. . . . . . 145 FFT of real data, radix-2 algorithm . . . . . . . . . . 144 FFT, complex data . . . . . . . . . . . . . . . . . . . . . . . . . 136 finding minima . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 309 finding roots . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 finding zeros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 fitting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 339 fitting, using Chebyshev polynomials . . . . . . . . . 286 flat distribution random variates . . . . . . . . . . . . . 191 Fortran range checking, equivalent in gcc . . . . . . 74 Four-tap Generalized Feedback Shift Register

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 Fourier integrals, numerical . . . . . . . . . . . . . . . . . . 157 Fourier Transforms, see FFT . . . . . . . . . . . . . . . . 135 Fractional Order Bessel Functions . . . . . . . . . . . . . 39 free software, explanation of . . . . . . . . . . . . . . . . . . . 1 functions, numerical differentiation . . . . . . . . . . . 284 fundamental constants . . . . . . . . . . . . . . . . . . . . . . 359

G

Gamma distribution random variates . . . . . . . . . 190 gamma function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 Gauss-Kronrod quadrature . . . . . . . . . . . . . . . . . . 151 Gaussian random variates . . . . . . . . . . . . . . . 178, 180 Gaussian Tail random variates . . . . . . . . . . . . . . . 179 gcc extensions, range-checking . . . . . . . . . . . . . . . . 74 gcc warning options . . . . . . . . . . . . . . . . . . . . . . . . . 373 gdb . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 371 Gear method, differential equations . . . . . . . . . . 272 Gegenbauer functions . . . . . . . . . . . . . . . . . . . . . . . . 57 GEMM, Level-3 BLAS . . . . . . . . . . . . . . . . . . . . . . 112 GEMV, Level-2 BLAS . . . . . . . . . . . . . . . . . . . . . . 109 general polynomial equations, solving . . . . . . . . . 27 Geometric random variates. . . . . . . . . . . . . . 210, 211 GER, Level-2 BLAS . . . . . . . . . . . . . . . . . . . . . . . . 111 GERC, Level-2 BLAS . . . . . . . . . . . . . . . . . . . . . . . 111 GERU, Level-2 BLAS . . . . . . . . . . . . . . . . . . . . . . . 111 Givens Rotation, BLAS . . . . . . . . . . . . . . . . . . . . . 108 Givens Rotation, Modified, BLAS . . . . . . . . . . . . 108 GNEWTON algorithm . . . . . . . . . . . . . . . . . . . . . . 323 GNU General Public License . . . . . . . . . . . . . . . . . . 1 golden section algorithm for finding minima… 313 gsl-announce mailing list . . . . . . . . . . . . . . . . . . . . . . 2 gsl sf result. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 gsl sf result e10 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 Gumbel distribution (Type 1), random variates

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201 Gumbel distribution (Type 2), random variates

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 202

H

Hankel transforms, discrete . . . . . . . . . . . . . . . . . . 294 HAVE INLINE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 HBOOK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 250 header files, including . . . . . . . . . . . . . . . . . . . . . . . . . 4 heapsort . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 HEMM, Level-3 BLAS . . . . . . . . . . . . . . . . . . . . . . 113 HEMV, Level-2 BLAS . . . . . . . . . . . . . . . . . . . . . . 110 HER, Level-2 BLAS . . . . . . . . . . . . . . . . . . . . . . . . 111 HER2, Level-2 BLAS . . . . . . . . . . . . . . . . . . . . . . . 112 HER2K, Level-3 BLAS. . . . . . . . . . . . . . . . . . . . . . 115 HERK, Level-3 BLAS . . . . . . . . . . . . . . . . . . . . . . . 115 hermitian matrix, complex, eigensystem . . . . . . 131 histogram, from ntuple . . . . . . . . . . . . . . . . . . . . . . 246 histograms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227 householder linear solver . . . . . . . . . . . . . . . . . . . . 127 HYBRID algorithm, unscaled without derivatives

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 323 HYBRID algorithms for nonlinear systems. … 322 HYBRIDJ algorithm . . . . . . . . . . . . . . . . . . . . . . . . 322 HYBRIDS algorithm, scaled without derivatives

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 323 HYBRIDSJ algorithm. . . . . . . . . . . . . . . . . . . . . . . 322 hydrogen atom . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 hyperbolic cosine, inverse . . . . . . . . . . . . . . . . . . . . . 15 hyperbolic functions, complex numbers . . . . . . . . 23 hyperbolic sine, inverse . . . . . . . . . . . . . . . . . . . . . . . 16 hyperbolic space . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 hyperbolic tangent, inverse . . . . . . . . . . . . . . . . . . . 16 hypergeometric functions . . . . . . . . . . . . . . . . . . . . . 57 hypot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 hypot function, special functions . . . . . . . . . . . . . . 66

I

identity permutation . . . . . . . . . . . . . . . . . . . . . . . . . 94 IEEE exceptions . . . . . . . . . . . . . . . . . . . . . . . . . . . . 368 IEEE floating point . . . . . . . . . . . . . . . . . . . . . . . . . 366 IEEE format for floating point numbers . . . . . . 366 IEEE infinity, defined as a macro . . . . . . . . . . . . . 15 IEEE NaN, defined as a macro. . . . . . . . . . . . . . . . 15 illumination, units of. . . . . . . . . . . . . . . . . . . . . . . . 364 imperial units . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 361 importance sampling, VEGAS . . . . . . . . . . . . . . . 255 including GSL header files . . . . . . . . . . . . . . . . . . . . . 4 incomplete Beta function, normalized . . . . . . . . . 56 incomplete Gamma function . . . . . . . . . . . . . . . . . . 56 indirect sorting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 indirect sorting, of vector elements . . . . . . . . . . . 100 infinity, defined as a macro . . . . . . . . . . . . . . . . . . . 15 infinity, IEEE format . . . . . . . . . . . . . . . . . . . . . . . 366 initial value problems, differential equations… 270 initializing matrices . . . . . . . . . . . . . . . . . . . . . . . . . . 84 initializing vectors . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 inline functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 integer powers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 integration, numerical (quadrature) . . . . . . . . . . 151 interpolation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 278 interpolation, using Chebyshev polynomials … 286 inverse complex trigonometric functions . . . . . . . 22 inverse hyperbolic cosine . . . . . . . . . . . . . . . . . . . . . 15 inverse hyperbolic functions, complex numbers

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 inverse hyperbolic sine . . . . . . . . . . . . . . . . . . . . . . . 16 inverse hyperbolic tangent . . . . . . . . . . . . . . . . . . . . 16 inverse of a matrix, by LU decomposition . . . . . 119 inverting a permutation . . . . . . . . . . . . . . . . . . . . . . 95 Irregular Cylindrical Bessel Functions . . . . . . . . . 34 Irregular Modified Bessel Functions, Fractional

Order. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 Irregular Modified Cylindrical Bessel Functions

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 Irregular Modified Spherical Bessel Functions . . 38 Irregular Spherical Bessel Functions . . . . . . . . . . . 37 iterating through permutations . . . . . . . . . . . . . . . 95 iterative refinement of solutions in linear systems

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119

J

Jacobi elliptic functions . . . . . . . . . . . . . . . . . . . . . . 48

L

Laguerre functions . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 Lambert function . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 Landau random variates. . . . . . . . . . . . . . . . . . . . . 187 LAPACK, recommended for linear algebra … 118,

130 Laplace distribution random variates . . . . . . . . . 182 leading dimension, matrices. . . . . . . . . . . . . . . . . . . 82 least squares fit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 339 least squares fitting, nonlinear . . . . . . . . . . . . . . . 348 least-squares, covariance of best-fit parameters

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 353 Legendre forms of elliptic integrals . . . . . . . . . . . . 46 Legendre functions . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 length, computed accurately using hypot . . . . . . 15 Levenberg-Marquardt algorithms. . . . . . . . . . . . . 352 Levin u-transform . . . . . . . . . . . . . . . . . . . . . . . . . . 290 Levy distribution, random variates . . . . . . 188, 189 libraries, linking with . . . . . . . . . . . . . . . . . . . . . . . . . 4 license of GSL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 light, units of. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 364 linear algebra. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 linear algebra, BLAS. . . . . . . . . . . . . . . . . . . . . . . . 105 linear interpolation . . . . . . . . . . . . . . . . . . . . . . . . . 279 linear systems, refinement of solutions . . . . . . . . 119 linear systems, solution of . . . . . . . . . . . . . . . . . . . 118 linking with GSL libraries . . . . . . . . . . . . . . . . . . . . . 4 LMDER algorithm. . . . . . . . . . . . . . . . . . . . . . . . . . 352 log1p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 logarithm and related functions . . . . . . . . . . . . . . . 63 logarithm of Beta function. . . . . . . . . . . . . . . . . . . . 56 logarithm of combinatorial factor C(m,n) . . . . . . 55 logarithm of complex number . . . . . . . . . . . . . . . . . 21 logarithm of cosh function, special functions … 67 logarithm of double factorial . . . . . . . . . . . . . . . . . . 55 logarithm of factorial. . . . . . . . . . . . . . . . . . . . . . . . . 55 logarithm of Gamma function. . . . . . . . . . . . . . . . . 54 logarithm of Pochhammer symbol . . . . . . . . . . . . . 56 logarithm of sinh function, special functions … 67 logarithm of the determinant of a matrix . . . . . 119 logarithm, computed accurately near 1 . . . . . . . . 15 Logarithmic random variates . . . . . . . . . . . . . . . . 212 Logistic random variates . . . . . . . . . . . . . . . . . . . . 197 Lognormal random variates. . . . . . . . . . . . . . . . . . 192 long double. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 low discrepancy sequences . . . . . . . . . . . . . . . . . . . 174 Low-level CBLAS. . . . . . . . . . . . . . . . . . . . . . . . . . . 378 LU decomposition . . . . . . . . . . . . . . . . . . . . . . . . . . 118

M

macros for mathematical constants . . . . . . . . . . . . 14 magnitude of complex number . . . . . . . . . . . . . . . . 19 mailing list archives . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 mailing list for GSL announcements . . . . . . . . . . . . 2 mantissa, IEEE format . . . . . . . . . . . . . . . . . . . . . . 366 mass, units of . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 362 mathematical constants, defined as macros. . . . . 14 mathematical functions, elementary . . . . . . . . . . . 14 matrices . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71, 82 matrices, initializing . . . . . . . . . . . . . . . . . . . . . . . . . 84 matrices, range-checking. . . . . . . . . . . . . . . . . . . . . . 83 matrix determinant . . . . . . . . . . . . . . . . . . . . . . . . . 119 matrix factorization. . . . . . . . . . . . . . . . . . . . . . . . . 118 matrix inverse . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 matrix square root, Cholesky decomposition . . 124 matrix, operations . . . . . . . . . . . . . . . . . . . . . . . . . . 105 max . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217 maximization, see minimization . . . . . . . . . . . . . . 309 maximum of two numbers . . . . . . . . . . . . . . . . . . . . 17 mean . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217 min. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217 Minimization, caveats . . . . . . . . . . . . . . . . . . . . . . . 310 minimization, multidimensional . . . . . . . . . . . . . . 330 minimization, one-dimensional . . . . . . . . . . . . . . . 309 minimization, overview . . . . . . . . . . . . . . . . . . . . . . 309 minimization, providing a function to minimize

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 311 minimization, stopping parameters . . . . . . . . . . . 312 minimum finding, brent’s method . . . . . . . . . . . . 313 minimum finding, golden section algorithm … 313 minimum of two numbers . . . . . . . . . . . . . . . . . . . . 17 MINPACK, minimization algorithms. . . . . 322, 352 MISCFUN . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 MISER monte carlo integration . . . . . . . . . . . . . . 253 Mixed-radix FFT, complex data . . . . . . . . . . . . . 139 Mixed-radix FFT, real data . . . . . . . . . . . . . . . . . 145 Modified Bessel Functions, Fractional Order … 40 Modified Clenshaw-Curtis quadrature . . . . . . . . 151 Modified Cylindrical Bessel Functions . . . . . . . . . 34 Modified Givens Rotation, BLAS . . . . . . . . . . . . 108 Modified Newton’s Method for nonlinear systems

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 323 Modified Spherical Bessel Functions . . . . . . . . . . . 38 Monte Carlo integration . . . . . . . . . . . . . . . . . . . . . 251 MRG, multiple recursive random number generator

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 MT19937 random number generator. . . . . . . . . . 165 multidimensional root finding, Broyden algorithm

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 324 multidimensional root finding, overview . . . . . . 316 multidimensional root finding, providing a function

to solve . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 318 Multimin, caveats. . . . . . . . . . . . . . . . . . . . . . . . . . . 330 multiplication . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45

N

N-dimensional random direction vector . . . . . . . 199 NaN, defined as a macro . . . . . . . . . . . . . . . . . . . . . 15 nautical units . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 361 Negative Binomial distribution, random variates

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 Newton algorithm, discrete . . . . . . . . . . . . . . . . . . 323 Newton’s Method algorithm for finding roots. . 303 Newton’s Method for systems of nonlinear

equations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 323 Niederreiter sequence . . . . . . . . . . . . . . . . . . . . . . . 174 non-linear equation, solutions of . . . . . . . . . . . . . 296 non-linear functions, minimization . . . . . . . . . . . 309 nonlinear fitting, stopping parameters . . . . . . . . 351 nonlinear least squares fitting . . . . . . . . . . . . . . . . 348 nonlinear least-squares fitting, overview . . . . . . 348 nonlinear systems of equations, solution of …. 316 normalized form, IEEE format . . . . . . . . . . . . . . . 366 normalized incomplete Beta function . . . . . . . . . . 56 Not-a-number, defined as a macro. . . . . . . . . . . . . 15 NRM2, Level-1 BLAS . . . . . . . . . . . . . . . . . . . . . . . 107 ntuples . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 245 nuclear physics, constants . . . . . . . . . . . . . . . . . . . 360 numerical constants, defined as macros . . . . . . . . 14 numerical derivatives. . . . . . . . . . . . . . . . . . . . . . . . 284 numerical integration (quadrature) . . . . . . . . . . . 151 Numerical recipes, random number generators

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169

O

obtaining GSL . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 ODEs, initial value problems . . . . . . . . . . . . . . . . 270 optimization – combinatorial . . . . . . . . . . . . . . . . 262 optimization, see minimization . . . . . . . . . . . . . . . 309 optimized functions, alternatives . . . . . . . . . . . . . . . 7 ordinary differential equations, initial value

problem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 270 overflow, IEEE exceptions . . . . . . . . . . . . . . . . . . . 368

P

Pareto random variates . . . . . . . . . . . . . . . . . . . . . 198 PAW . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 250 permutations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 physical constants . . . . . . . . . . . . . . . . . . . . . . . . . . 359 physical dimension, matrices . . . . . . . . . . . . . . . . . . 82 pi, defined as a macro . . . . . . . . . . . . . . . . . . . . . . . . 14 plain monte carlo . . . . . . . . . . . . . . . . . . . . . . . . . . . 252 Pochhammer symbol . . . . . . . . . . . . . . . . . . . . . . . . . 55 Poisson random numbers . . . . . . . . . . . . . . . . . . . . 205 polar form of complex numbers . . . . . . . . . . . . . . . 18 polar to rectangular conversion . . . . . . . . . . . . . . . 67 polygamma functions . . . . . . . . . . . . . . . . . . . . . . . . 64 polynomial evaluation . . . . . . . . . . . . . . . . . . . . . . . . 26 polynomials, roots of . . . . . . . . . . . . . . . . . . . . . . . . . 26 power function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 power of complex number . . . . . . . . . . . . . . . . . . . . 21 power, units of . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 363 precision, IEEE arithmetic . . . . . . . . . . . . . . . . . . 368 pressure, units of . . . . . . . . . . . . . . . . . . . . . . . . . . . 363 Prince-Dormand, Runge-Kutta method . . . . . . . 272 printers units . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 362 probability distributions, from histograms …. 233 projection of ntuples . . . . . . . . . . . . . . . . . . . . . . . . 246 psi function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64

Q

QR decomposition . . . . . . . . . . . . . . . . . . . . . . . . . . 120 QR decomposition with column pivoting. . . . . . 122 QUADPACK . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 quadratic equation, solving . . . . . . . . . . . . . . . . . . . 26 quadrature . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 quasi-random sequences . . . . . . . . . . . . . . . . . . . . . 174

R

R250 shift-register random number generator. . 170 Racah coefficients . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 radioactivity, units of . . . . . . . . . . . . . . . . . . . . . . . 364 Radix-2 FFT for real data . . . . . . . . . . . . . . . . . . . 144 Radix-2 FFT, complex data . . . . . . . . . . . . . . . . . 137 rand48 random number generator . . . . . . . . . . . . 169 random number distributions . . . . . . . . . . . . . . . . 177 random number generators . . . . . . . . . . . . . . . . . . 160 Random number generators, Numerical recipes

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 RANDU random number generator . . . . . . . . . . 171 RANF random number generator . . . . . . . . . . . . 170 range . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217 range-checking for matrices . . . . . . . . . . . . . . . . . . . 83 range-checking for vectors . . . . . . . . . . . . . . . . . . . . 74 RANLUX random number generator . . . . . . . . . 166 RANLXD random number generator . . . . . . . . . 166 RANLXS random number generator . . . . . . . . . 165 RANMAR random number generator …. 170, 171 Rayleigh random variates. . . . . . . . . . . . . . . . . . . . 185 Rayleigh Tail random variates . . . . . . . . . . . . . . . 186 real symmetric matrix, eigensystem . . . . . . . . . . 130 Reciprocal Gamma function . . . . . . . . . . . . . . . . . . 54 rectangular to polar conversion . . . . . . . . . . . . . . . 67 recursive stratified sampling, MISER . . . . . . . . . 253 reduction of angular variables. . . . . . . . . . . . . . . . . 67 refinement of solutions in linear systems . . . . . . 119 regression, least squares . . . . . . . . . . . . . . . . . . . . . 339 Regular Bessel Functions, Fractional Order …. 39 Regular Bessel Functions, Zeros of . . . . . . . . . . . . 40 Regular Cylindrical Bessel Functions . . . . . . . . . . 33 Regular Modified Bessel Functions - Fractional

Order. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 Regular Modified Cylindrical Bessel Functions. . 34 Regular Modified Spherical Bessel Functions … 38 Regular Spherical Bessel Functions . . . . . . . . . . . . 36 Regulated Gamma function. . . . . . . . . . . . . . . . . . . 54 relative Pochhammer symbol . . . . . . . . . . . . . . . . . 56 representations of complex numbers . . . . . . . . . . . 18 resampling from histograms . . . . . . . . . . . . . . . . . 233 residual, in nonlinear systems of equations … 321,

351 reversing a permutation . . . . . . . . . . . . . . . . . . . . . . 95 RK2, Runge-Kutta Method. . . . . . . . . . . . . . . . . . 271 RK4, Runge-Kutta Method. . . . . . . . . . . . . . . . . . 271 RKF45, Runge-Kutta-Fehlberg method. . . . . . . 272 root finding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 root finding, bisection algorithm . . . . . . . . . . . . . 302 root finding, brent’s method . . . . . . . . . . . . . . . . . 302 root finding, caveats . . . . . . . . . . . . . . . . . . . . . . . . 296 root finding, false position algorithm . . . . . . . . . 302 root finding, initial guess . . . . . . . . . . . . . . . . . . . . 300 root finding, Newton’s Method algorithm . . . . . 303 root finding, overview . . . . . . . . . . . . . . . . . . . . . . . 296 root finding, providing a function to solve …. 298 root finding, search bounds . . . . . . . . . . . . . . . . . . 300 root finding, Secant Method algorithm . . . . . . . 303 root finding, Steffenson’s Method . . . . . . . . . . . . 304 root finding, stopping parameters. . . . . . . . 301, 321 roots . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 ROTG, Level-1 BLAS. . . . . . . . . . . . . . . . . . . . . . . 108 rounding mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 368 Runge-Kutta Cash-Karp method. . . . . . . . . . . . . 272 Runge-Kutta methods, ordinary differential

equations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 271 Runge-Kutta Prince-Dormand method . . . . . . . 272

S

sampling from histograms . . . . . . . . . . . . . . . . . . . 233 SCAL, Level-1 BLAS . . . . . . . . . . . . . . . . . . . . . . . 108 schedule - cooling . . . . . . . . . . . . . . . . . . . . . . . . . . . 262 Secant Method algorithm for finding roots …. 303 selection function, ntuples . . . . . . . . . . . . . . . . . . . 246 series, acceleration . . . . . . . . . . . . . . . . . . . . . . . . . . 290 shift-register random number generator. . . . . . . 170 sign bit, IEEE format . . . . . . . . . . . . . . . . . . . . . . . 366 sign of the determinant of a matrix . . . . . . . . . . 119 simulated annealing . . . . . . . . . . . . . . . . . . . . . . . . . 262 sin of complex number . . . . . . . . . . . . . . . . . . . . . . . 22 sine function, special functions . . . . . . . . . . . . . . . . 66 single precision, IEEE format . . . . . . . . . . . . . . . . 366 singular value decomposition . . . . . . . . . . . . . . . . 123 Sobol sequence . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 solution of linear system by householder

transformations . . . . . . . . . . . . . . . . . . . . . . . . 127 solution of linear systems, Ax=b . . . . . . . . . . . . . 118 solving a non-linear equation . . . . . . . . . . . . . . . . 296 solving nonlinear systems of equations . . . . . . . . 316 sorting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 sorting eigenvalues and eigenvectors . . . . . . . . . . 132 sorting vector elements . . . . . . . . . . . . . . . . . . . . . . 100 source code, reuse in applications . . . . . . . . . . . . . . 9 Special Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 Spherical Bessel Functions . . . . . . . . . . . . . . . . . . . . 36 spherical harmonics . . . . . . . . . . . . . . . . . . . . . . . . . . 60 spherical random variates, 2D . . . . . . . . . . . . . . . 199 spherical random variates, 3D . . . . . . . . . . . . . . . 199 spherical random variates, N-dimensional . . . . . 199 spline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 278 square root of a matrix, Cholesky decomposition

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 square root of complex number . . . . . . . . . . . . . . . 21 standard deviation . . . . . . . . . . . . . . . . . . . . . . . . . . 217 standards conformance, ANSI C . . . . . . . . . . . . . . . 4 statistics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217 Steffenson’s Method for finding roots . . . . . . . . . 304 stride, of vector index . . . . . . . . . . . . . . . . . . . . . . . . 73 summation, acceleration. . . . . . . . . . . . . . . . . . . . . 290 SVD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123 SWAP, Level-1 BLAS . . . . . . . . . . . . . . . . . . . . . . . 107 swapping permutation elements . . . . . . . . . . . . . . . 95 SYMM, Level-3 BLAS . . . . . . . . . . . . . . . . . . . . . . 113 symmetric matrix, real, eigensystem. . . . . . . . . . 130 SYMV, Level-2 BLAS. . . . . . . . . . . . . . . . . . . . . . . 110 Synchrotron functions . . . . . . . . . . . . . . . . . . . . . . . . 65 SYR, Level-2 BLAS. . . . . . . . . . . . . . . . . . . . . . . . . 111 SYR2, Level-2 BLAS. . . . . . . . . . . . . . . . . . . . . . . . 111 SYR2K, Level-3 BLAS . . . . . . . . . . . . . . . . . . . . . . 115 SYRK, Level-3 BLAS . . . . . . . . . . . . . . . . . . . . . . . 114 systems of equations, nonlinear . . . . . . . . . . . . . . 316

T

t-distribution random variates . . . . . . . . . . . . . . . 195 t-test . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217 tangent of complex number . . . . . . . . . . . . . . . . . . . 22 Tausworthe random number generator. . . . . . . . 167 Taylor coefficients, computation of . . . . . . . . . . . . 55 testing permutation for validity . . . . . . . . . . . . . . . 95 thermal energy, units of . . . . . . . . . . . . . . . . . . . . . 363 time units . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 361 trailing dimension, matrices . . . . . . . . . . . . . . . . . . 82 transforms, Hankel. . . . . . . . . . . . . . . . . . . . . . . . . . 294 Transport functions . . . . . . . . . . . . . . . . . . . . . . . . . . 65 Traveling Salesman Problem . . . . . . . . . . . . . . . . . 266 tridiagonal decomposition . . . . . . . . . . . . . . . . . . . 125 tridiagonal systems . . . . . . . . . . . . . . . . . . . . . . . . . 127 Trigonometric functions . . . . . . . . . . . . . . . . . . . . . . 66 trigonometric functions of complex numbers … 22 TRMM, Level-3 BLAS . . . . . . . . . . . . . . . . . . . . . . 113 TRMV, Level-2 BLAS . . . . . . . . . . . . . . . . . . . . . . 109 TRSM, Level-3 BLAS . . . . . . . . . . . . . . . . . . . . . . . 114 TRSV, Level-2 BLAS . . . . . . . . . . . . . . . . . . . . . . . 110 TSP . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 266 TT800 random number generator . . . . . . . . . . . . 171 two dimensional histograms . . . . . . . . . . . . . . . . . 236 two-sided exponential random variates . . . . . . . 182 Type 1 Gumbel distribution, random variates

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 201 Type 2 Gumbel distribution, random variate . . 202

U

u-transform for series . . . . . . . . . . . . . . . . . . . . . . . 290 underflow, IEEE exceptions . . . . . . . . . . . . . . . . . 368 units, conversion of . . . . . . . . . . . . . . . . . . . . . . . . . 359 units, imperial. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 361 Unix random number generators, rand . . . . . . . 168 Unix random number generators, rand48 . . . . . 168 usage, compiling application programs . . . . . . . . . . 4

V

value function, ntuples . . . . . . . . . . . . . . . . . . . . . . 246 Van der Pol oscillator, example . . . . . . . . . . . . . . 275 variance. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217 VAX random number generator . . . . . . . . . . . . . . 171 vector, operations. . . . . . . . . . . . . . . . . . . . . . . . . . . 105 vector, sorting elements of . . . . . . . . . . . . . . . . . . . 100 vectors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71, 73 vectors, initializing. . . . . . . . . . . . . . . . . . . . . . . . . . . 75 vectors, range-checking . . . . . . . . . . . . . . . . . . . . . . . 74 VEGAS monte carlo integration . . . . . . . . . . . . . 255 viscosity, units of . . . . . . . . . . . . . . . . . . . . . . . . . . . 363 volume units . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 362

W

W function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 warning options . . . . . . . . . . . . . . . . . . . . . . . . . . . . 373 warranty (none) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 website, developer information . . . . . . . . . . . . . . . . . 3 Weibull distribution random variates . . . . . . . . . 200 weight, units of . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 362 Wigner coefficients . . . . . . . . . . . . . . . . . . . . . . . . . . . 43

Z

zero finding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 zero, IEEE format . . . . . . . . . . . . . . . . . . . . . . . . . . 366 Zeros of Regular Bessel Functions . . . . . . . . . . . . . 40 Zeta functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68

Concept Index 430

Table of Contents

1 Introduction. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1.1 Routines available in GSL. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.2 GSL is Free Software . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.3 Obtaining GSL. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.4 An Example Program . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.5 No Warranty. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 1.6 Further Information . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

2 Using the library . . . . . . . . . . . . . . . . . . . . . . . . . . 4

2.1 ANSI C Compliance . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 2.2 Compiling and Linking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 2.3 Shared Libraries. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 2.4 Autoconf macros . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 2.5 Inline functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 2.6 Long double . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 2.7 Portability functions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 2.8 Alternative optimized functions . . . . . . . . . . . . . . . . . . . . . . . . . . 7 2.9 Support for different numeric types . . . . . . . . . . . . . . . . . . . . . . . 8 2.10 Compatibility with C++. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 2.11 Aliasing of arrays . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 2.12 Code Reuse . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

3 Error Handling . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

3.1 Error Reporting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 3.2 Error Handlers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 3.3 Using GSL error reporting in your own functions . . . . . . . . . 13

4 Mathematical Functions. . . . . . . . . . . . . . . . . . . 14

4.1 Mathematical Constants . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 4.2 Infinities and Not-a-number . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 4.3 Elementary Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 4.4 Small integer powers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 4.5 Testing the Sign of Numbers . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 4.6 Testing for Odd and Even Numbers. . . . . . . . . . . . . . . . . . . . . . 17 4.7 Maximum and Minimum functions . . . . . . . . . . . . . . . . . . . . . . 17

5 Complex Numbers . . . . . . . . . . . . . . . . . . . . . . . . 18

5.1 Complex numbers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 5.2 Properties of complex numbers . . . . . . . . . . . . . . . . . . . . . . . . . . 19 5.3 Complex arithmetic operators . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 5.4 Elementary Complex Functions. . . . . . . . . . . . . . . . . . . . . . . . . . 21 5.5 Complex Trigonometric Functions . . . . . . . . . . . . . . . . . . . . . . . 22 5.6 Inverse Complex Trigonometric Functions . . . . . . . . . . . . . . . . 22 5.7 Complex Hyperbolic Functions . . . . . . . . . . . . . . . . . . . . . . . . . . 23 5.8 Inverse Complex Hyperbolic Functions . . . . . . . . . . . . . . . . . . . 24 5.9 References and Further Reading . . . . . . . . . . . . . . . . . . . . . . . . . 25

6 Roots of Polynomials . . . . . . . . . . . . . . . . . . . . . 26

6.1 Polynomial evaluation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 6.2 Quadratic equations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 6.3 Cubic equations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 6.4 General polynomial equations . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 6.5 Examples . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 6.6 References and Further Reading . . . . . . . . . . . . . . . . . . . . . . . . . 29

7 Special Functions . . . . . . . . . . . . . . . . . . . . . . . . . 30

7.1 Usage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 7.2 The gsl sf result struct. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 7.3 Modes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 7.4 Airy Functions and Derivatives . . . . . . . . . . . . . . . . . . . . . . . . . . 31

7.4.1 Airy Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 7.4.2 Derivatives of Airy Functions . . . . . . . . . . . . . . . . . . . 32 7.4.3 Zeros of Airy Functions . . . . . . . . . . . . . . . . . . . . . . . . 33 7.4.4 Zeros of Derivatives of Airy Functions . . . . . . . . . . . 33

7.5 Bessel Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 7.5.1 Regular Cylindrical Bessel Functions . . . . . . . . . . . . 33 7.5.2 Irregular Cylindrical Bessel Functions . . . . . . . . . . . 34 7.5.3 Regular Modified Cylindrical Bessel Functions … 34 7.5.4 Irregular Modified Cylindrical Bessel Functions . . 35 7.5.5 Regular Spherical Bessel Functions . . . . . . . . . . . . . . 36 7.5.6 Irregular Spherical Bessel Functions . . . . . . . . . . . . . 37 7.5.7 Regular Modified Spherical Bessel Functions . . . . . 38 7.5.8 Irregular Modified Spherical Bessel Functions …. 38 7.5.9 Regular Bessel Function - Fractional Order . . . . . . 39 7.5.10 Irregular Bessel Functions - Fractional Order … 39 7.5.11 Regular Modified Bessel Functions - Fractional

Order . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 7.5.12 Irregular Modified Bessel Functions - Fractional

Order . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 7.5.13 Zeros of Regular Bessel Functions . . . . . . . . . . . . . . 40

7.6 Clausen Functions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 7.7 Coulomb Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41

7.7.1 Normalized Hydrogenic Bound States . . . . . . . . . . . 41 7.7.2 Coulomb Wave Functions . . . . . . . . . . . . . . . . . . . . . . 42 7.7.3 Coulomb Wave Function Normalization Constant

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43

7.8 Coupling Coefficients . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 7.8.1 3-j Symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 7.8.2 6-j Symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 7.8.3 9-j Symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44

7.9 Dawson Function. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 7.10 Debye Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 7.11 Dilogarithm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45

7.11.1 Real Argument . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 7.11.2 Complex Argument . . . . . . . . . . . . . . . . . . . . . . . . . . . 45

7.12 Elementary Operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 7.13 Elliptic Integrals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46

7.13.1 Definition of Legendre Forms . . . . . . . . . . . . . . . . . . 46 7.13.2 Definition of Carlson Forms . . . . . . . . . . . . . . . . . . . 46 7.13.3 Legendre Form of Complete Elliptic Integrals … 46 7.13.4 Legendre Form of Incomplete Elliptic Integrals. . 47 7.13.5 Carlson Forms. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47

7.14 Elliptic Functions (Jacobi) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 7.15 Error Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48

7.15.1 Error Function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 7.15.2 Complementary Error Function . . . . . . . . . . . . . . . . 48 7.15.3 Log Complementary Error Function. . . . . . . . . . . . 48 7.15.4 Probability functions. . . . . . . . . . . . . . . . . . . . . . . . . . 49

7.16 Exponential Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 7.16.1 Exponential Function . . . . . . . . . . . . . . . . . . . . . . . . . 49 7.16.2 Relative Exponential Functions . . . . . . . . . . . . . . . . 50 7.16.3 Exponentiation With Error Estimate . . . . . . . . . . . 50

7.17 Exponential Integrals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 7.17.1 Exponential Integral . . . . . . . . . . . . . . . . . . . . . . . . . . 51 7.17.2 Ei(x) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 7.17.3 Hyperbolic Integrals . . . . . . . . . . . . . . . . . . . . . . . . . . 51 7.17.4 Ei 3(x) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 7.17.5 Trigonometric Integrals . . . . . . . . . . . . . . . . . . . . . . . 52 7.17.6 Arctangent Integral . . . . . . . . . . . . . . . . . . . . . . . . . . . 52

7.18 Fermi-Dirac Function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 7.18.1 Complete Fermi-Dirac Integrals . . . . . . . . . . . . . . . . 52 7.18.2 Incomplete Fermi-Dirac Integrals . . . . . . . . . . . . . . 53

7.19 Gamma Function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 7.20 Gegenbauer Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 7.21 Hypergeometric Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 7.22 Laguerre Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 7.23 Lambert W Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 7.24 Legendre Functions and Spherical Harmonics . . . . . . . . . . . . 60

7.24.1 Legendre Polynomials . . . . . . . . . . . . . . . . . . . . . . . . . 60 7.24.2 Associated Legendre Polynomials and Spherical

Harmonics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 7.24.3 Conical Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 7.24.4 Radial Functions for Hyperbolic Space . . . . . . . . . 63

7.25 Logarithm and Related Functions . . . . . . . . . . . . . . . . . . . . . . 63 7.26 Power Function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 7.27 Psi (Digamma) Function. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64

7.27.1 Digamma Function . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 7.27.2 Trigamma Function . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 7.27.3 Polygamma Function . . . . . . . . . . . . . . . . . . . . . . . . . 65

7.28 Synchrotron Functions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 7.29 Transport Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 7.30 Trigonometric Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66

7.30.1 Circular Trigonometric Functions . . . . . . . . . . . . . . 66 7.30.2 Trigonometric Functions for Complex Arguments

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 7.30.3 Hyperbolic Trigonometric Functions. . . . . . . . . . . . 67 7.30.4 Conversion Functions . . . . . . . . . . . . . . . . . . . . . . . . . 67 7.30.5 Restriction Functions . . . . . . . . . . . . . . . . . . . . . . . . . 67 7.30.6 Trigonometric Functions With Error Estimates. . 67

7.31 Zeta Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 7.31.1 Riemann Zeta Function . . . . . . . . . . . . . . . . . . . . . . . 68 7.31.2 Hurwitz Zeta Function . . . . . . . . . . . . . . . . . . . . . . . . 68 7.31.3 Eta Function . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68

7.32 Examples. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 7.33 References and Further Reading . . . . . . . . . . . . . . . . . . . . . . . . 70

8 Vectors and Matrices . . . . . . . . . . . . . . . . . . . . . 71

8.1 Data types. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 8.2 Blocks. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71

8.2.1 Block allocation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 8.2.2 Reading and writing blocks . . . . . . . . . . . . . . . . . . . . . 72 8.2.3 Example programs for blocks . . . . . . . . . . . . . . . . . . . 73

8.3 Vectors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 8.3.1 Vector allocation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 8.3.2 Accessing vector elements . . . . . . . . . . . . . . . . . . . . . . 74 8.3.3 Initializing vector elements . . . . . . . . . . . . . . . . . . . . . 75 8.3.4 Reading and writing vectors . . . . . . . . . . . . . . . . . . . . 75 8.3.5 Vector views . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 8.3.6 Copying vectors . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 8.3.7 Exchanging elements . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 8.3.8 Vector operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 8.3.9 Finding maximum and minimum elements of vectors

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 8.3.10 Vector properties . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 8.3.11 Example programs for vectors . . . . . . . . . . . . . . . . . 80

8.4 Matrices . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 8.4.1 Matrix allocation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 8.4.2 Accessing matrix elements. . . . . . . . . . . . . . . . . . . . . . 83 8.4.3 Initializing matrix elements. . . . . . . . . . . . . . . . . . . . . 84 8.4.4 Reading and writing matrices. . . . . . . . . . . . . . . . . . . 84 8.4.5 Matrix views . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 8.4.6 Creating row and column views . . . . . . . . . . . . . . . . . 87 8.4.7 Copying matrices . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 88 8.4.8 Copying rows and columns . . . . . . . . . . . . . . . . . . . . . 88 8.4.9 Exchanging rows and columns . . . . . . . . . . . . . . . . . . 89 8.4.10 Matrix operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 8.4.11 Finding maximum and minimum elements of

matrices. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 8.4.12 Matrix properties . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 91 8.4.13 Example programs for matrices . . . . . . . . . . . . . . . . 91

8.5 References and Further Reading . . . . . . . . . . . . . . . . . . . . . . . . . 93

9 Permutations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94

9.1 The Permutation struct. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 9.2 Permutation allocation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 9.3 Accessing permutation elements . . . . . . . . . . . . . . . . . . . . . . . . . 95 9.4 Permutation properties . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 9.5 Permutation functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 9.6 Applying Permutations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 9.7 Reading and writing permutations . . . . . . . . . . . . . . . . . . . . . . . 96 9.8 Examples . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 9.9 References and Further Reading . . . . . . . . . . . . . . . . . . . . . . . . . 98

10 Sorting. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99

10.1 Sorting objects. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 10.2 Sorting vectors. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 10.3 Selecting the k-th smallest or largest elements . . . . . . . . . . 101 10.4 Computing the rank. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 10.5 Examples. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 10.6 References and Further Reading . . . . . . . . . . . . . . . . . . . . . . . 104

11 BLAS Support . . . . . . . . . . . . . . . . . . . . . . . . . 105

11.1 GSL BLAS Interface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 11.1.1 Level 1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 106 11.1.2 Level 2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 11.1.3 Level 3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112

11.2 Examples. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 116 11.3 References and Further Reading . . . . . . . . . . . . . . . . . . . . . . . 117

12 Linear Algebra . . . . . . . . . . . . . . . . . . . . . . . . . 118

12.1 LU Decomposition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 12.2 QR Decomposition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 12.3 QR Decomposition with Column Pivoting . . . . . . . . . . . . . . 122 12.4 Singular Value Decomposition. . . . . . . . . . . . . . . . . . . . . . . . . 123 12.5 Cholesky Decomposition . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 124 12.6 Tridiagonal Decomposition of Real Symmetric Matrices. . 125 12.7 Tridiagonal Decomposition of Hermitian Matrices. . . . . . . 125 12.8 Bidiagonalization . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 12.9 Householder solver for linear systems . . . . . . . . . . . . . . . . . . 127 12.10 Tridiagonal Systems . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 12.11 Examples. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128 12.12 References and Further Reading . . . . . . . . . . . . . . . . . . . . . . 129

13 Eigensystems . . . . . . . . . . . . . . . . . . . . . . . . . . . 130

13.1 Real Symmetric Matrices . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 13.2 Complex Hermitian Matrices. . . . . . . . . . . . . . . . . . . . . . . . . . 131 13.3 Sorting Eigenvalues and Eigenvectors . . . . . . . . . . . . . . . . . . 132 13.4 Examples. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 132 13.5 References and Further Reading . . . . . . . . . . . . . . . . . . . . . . . 134

14 Fast Fourier Transforms (FFTs) . . . . . . . . . 135

14.1 Mathematical Definitions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 14.2 Overview of complex data FFTs. . . . . . . . . . . . . . . . . . . . . . . 136 14.3 Radix-2 FFT routines for complex data . . . . . . . . . . . . . . . . 137 14.4 Mixed-radix FFT routines for complex data . . . . . . . . . . . . 139 14.5 Overview of real data FFTs . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 14.6 Radix-2 FFT routines for real data . . . . . . . . . . . . . . . . . . . . 144 14.7 Mixed-radix FFT routines for real data . . . . . . . . . . . . . . . . 145 14.8 References and Further Reading . . . . . . . . . . . . . . . . . . . . . . . 150

15 Numerical Integration . . . . . . . . . . . . . . . . . . 151

15.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 15.2 QNG non-adaptive Gauss-Kronrod integration . . . . . . . . . 152 15.3 QAG adaptive integration . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 15.4 QAGS adaptive integration with singularities. . . . . . . . . . . 153 15.5 QAGP adaptive integration with known singular points. . 153 15.6 QAGI adaptive integration on infinite intervals . . . . . . . . . 154 15.7 QAWC adaptive integration for Cauchy principal values

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 15.8 QAWS adaptive integration for singular functions. . . . . . . 155 15.9 QAWO adaptive integration for oscillatory functions …. 156 15.10 QAWF adaptive integration for Fourier integrals. . . . . . . 157 15.11 Error codes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 157 15.12 Examples. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 158 15.13 References and Further Reading . . . . . . . . . . . . . . . . . . . . . . 159

16 Random Number Generation . . . . . . . . . . . . 160

16.1 General comments on random numbers . . . . . . . . . . . . . . . . 160 16.2 The Random Number Generator Interface . . . . . . . . . . . . . 160 16.3 Random number generator initialization . . . . . . . . . . . . . . . 161 16.4 Sampling from a random number generator . . . . . . . . . . . . 162 16.5 Auxiliary random number generator functions . . . . . . . . . . 162 16.6 Random number environment variables . . . . . . . . . . . . . . . . 163 16.7 Saving and restoring random number generator state … 164 16.8 Random number generator algorithms . . . . . . . . . . . . . . . . . 165 16.9 Unix random number generators . . . . . . . . . . . . . . . . . . . . . . 168 16.10 Numerical Recipes generators . . . . . . . . . . . . . . . . . . . . . . . . 169 16.11 Other random number generators . . . . . . . . . . . . . . . . . . . . 170 16.12 Random Number Generator Performance . . . . . . . . . . . . . 172 16.13 References and Further Reading . . . . . . . . . . . . . . . . . . . . . . 173 16.14 Acknowledgements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173

17 Quasi-Random Sequences . . . . . . . . . . . . . . . 174

17.1 Quasi-random number generator initialization . . . . . . . . . . 174 17.2 Sampling from a quasi-random number generator . . . . . . . 174 17.3 Auxiliary quasi-random number generator functions …. 174 17.4 Saving and resorting quasi-random number generator state

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 17.5 Quasi-random number generator algorithms . . . . . . . . . . . . 175 17.6 Examples. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 17.7 References . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176

18 Random Number Distributions . . . . . . . . . . 177

18.1 The Gaussian Distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . 178 18.2 The Gaussian Tail Distribution. . . . . . . . . . . . . . . . . . . . . . . . 179 18.3 The Bivariate Gaussian Distribution . . . . . . . . . . . . . . . . . . . 180 18.4 The Exponential Distribution . . . . . . . . . . . . . . . . . . . . . . . . . 181 18.5 The Laplace Distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 182 18.6 The Exponential Power Distribution . . . . . . . . . . . . . . . . . . . 183 18.7 The Cauchy Distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 18.8 The Rayleigh Distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185 18.9 The Rayleigh Tail Distribution . . . . . . . . . . . . . . . . . . . . . . . . 186 18.10 The Landau Distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . 187 18.11 The Levy alpha-Stable Distributions. . . . . . . . . . . . . . . . . . 188 18.12 The Levy skew alpha-Stable Distribution . . . . . . . . . . . . . 189 18.13 The Gamma Distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . 190 18.14 The Flat (Uniform) Distribution . . . . . . . . . . . . . . . . . . . . . 191 18.15 The Lognormal Distribution . . . . . . . . . . . . . . . . . . . . . . . . . 192 18.16 The Chi-squared Distribution . . . . . . . . . . . . . . . . . . . . . . . . 193 18.17 The F-distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 194 18.18 The t-distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 195 18.19 The Beta Distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 196 18.20 The Logistic Distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . 197 18.21 The Pareto Distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 198 18.22 The Spherical Distribution (2D & 3D) . . . . . . . . . . . . . . . . 199 18.23 The Weibull Distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . 200 18.24 The Type-1 Gumbel Distribution. . . . . . . . . . . . . . . . . . . . . 201 18.25 The Type-2 Gumbel Distribution. . . . . . . . . . . . . . . . . . . . . 202 18.26 General Discrete Distributions . . . . . . . . . . . . . . . . . . . . . . . 203 18.27 The Poisson Distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . 205 18.28 The Bernoulli Distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . 206 18.29 The Binomial Distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . 207 18.30 The Negative Binomial Distribution . . . . . . . . . . . . . . . . . . 208 18.31 The Pascal Distribution . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 18.32 The Geometric Distribution. . . . . . . . . . . . . . . . . . . . . . . . . . 210 18.33 The Hypergeometric Distribution. . . . . . . . . . . . . . . . . . . . . 211 18.34 The Logarithmic Distribution . . . . . . . . . . . . . . . . . . . . . . . . 212 18.35 Shuffling and Sampling . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 18.36 Examples. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 18.37 References and Further Reading . . . . . . . . . . . . . . . . . . . . . . 216

19 Statistics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217

19.1 Mean, Standard Deviation and Variance . . . . . . . . . . . . . . . 217 19.2 Absolute deviation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 218 19.3 Higher moments (skewness and kurtosis) . . . . . . . . . . . . . . . 219 19.4 Autocorrelation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 220 19.5 Covariance . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 220 19.6 Weighted Samples. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 220 19.7 Maximum and Minimum values . . . . . . . . . . . . . . . . . . . . . . . 222 19.8 Median and Percentiles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 223 19.9 Example statistical programs. . . . . . . . . . . . . . . . . . . . . . . . . . 224 19.10 References and Further Reading . . . . . . . . . . . . . . . . . . . . . . 226

20 Histograms. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227

20.1 The histogram struct . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 227 20.2 Histogram allocation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 228 20.3 Copying Histograms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 229 20.4 Updating and accessing histogram elements . . . . . . . . . . . . 229 20.5 Searching histogram ranges . . . . . . . . . . . . . . . . . . . . . . . . . . . 230 20.6 Histogram Statistics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 230 20.7 Histogram Operations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 231 20.8 Reading and writing histograms . . . . . . . . . . . . . . . . . . . . . . . 232 20.9 Resampling from histograms . . . . . . . . . . . . . . . . . . . . . . . . . . 233 20.10 The histogram probability distribution struct. . . . . . . . . . 233 20.11 Example programs for histograms . . . . . . . . . . . . . . . . . . . . 234 20.12 Two dimensional histograms . . . . . . . . . . . . . . . . . . . . . . . . . 236 20.13 The 2D histogram struct . . . . . . . . . . . . . . . . . . . . . . . . . . . . 236 20.14 2D Histogram allocation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 237 20.15 Copying 2D Histograms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 237 20.16 Updating and accessing 2D histogram elements . . . . . . . . 238 20.17 Searching 2D histogram ranges . . . . . . . . . . . . . . . . . . . . . . . 239 20.18 2D Histogram Statistics . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 239 20.19 2D Histogram Operations. . . . . . . . . . . . . . . . . . . . . . . . . . . . 240 20.20 Reading and writing 2D histograms . . . . . . . . . . . . . . . . . . 241 20.21 Resampling from 2D histograms. . . . . . . . . . . . . . . . . . . . . . 242 20.22 Example programs for 2D histograms . . . . . . . . . . . . . . . . . 243

21 N-tuples . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 245

21.1 The ntuple struct . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 245 21.2 Creating ntuples . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 245 21.3 Opening an existing ntuple file . . . . . . . . . . . . . . . . . . . . . . . . 245 21.4 Writing ntuples . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 246 21.5 Reading ntuples. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 246 21.6 Closing an ntuple file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 246 21.7 Histogramming ntuple values. . . . . . . . . . . . . . . . . . . . . . . . . . 246 21.8 Example programs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 247 21.9 References and Further Reading . . . . . . . . . . . . . . . . . . . . . . . 250

22 Monte Carlo Integration . . . . . . . . . . . . . . . . 251

22.1 Interface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 251 22.2 PLAIN Monte Carlo. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 252 22.3 MISER. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 253 22.4 VEGAS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 255 22.5 Examples. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 258 22.6 References and Further Reading . . . . . . . . . . . . . . . . . . . . . . . 261

23 Simulated Annealing. . . . . . . . . . . . . . . . . . . . 262

23.1 Simulated Annealing algorithm. . . . . . . . . . . . . . . . . . . . . . . . 262 23.2 Simulated Annealing functions . . . . . . . . . . . . . . . . . . . . . . . . 263 23.3 Examples with Simulated Annealing . . . . . . . . . . . . . . . . . . . 264

23.3.1 Trivial example. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 264 23.3.2 Traveling Salesman Problem. . . . . . . . . . . . . . . . . . 266

24 Ordinary Differential Equations . . . . . . . . . 270

24.1 Defining the ODE System . . . . . . . . . . . . . . . . . . . . . . . . . . . . 270 24.2 Stepping Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 271 24.3 Adaptive Step-size Control. . . . . . . . . . . . . . . . . . . . . . . . . . . . 272 24.4 Evolution. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 274 24.5 Examples. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 275 24.6 References and Further Reading . . . . . . . . . . . . . . . . . . . . . . . 277

25 Interpolation . . . . . . . . . . . . . . . . . . . . . . . . . . . 278

25.1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 278 25.2 Interpolation Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 278 25.3 Interpolation Types . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 279 25.4 Index Look-up and Acceleration . . . . . . . . . . . . . . . . . . . . . . . 280 25.5 Evaluation of interpolating functions. . . . . . . . . . . . . . . . . . . 280 25.6 Higher-level interface . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 281 25.7 Examples. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 282 25.8 References and Further Reading . . . . . . . . . . . . . . . . . . . . . . . 283

26 Numerical Differentiation . . . . . . . . . . . . . . . 284

26.1 Functions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 284 26.2 Example . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 284 26.3 References and Further Reading . . . . . . . . . . . . . . . . . . . . . . . 285

27 Chebyshev Approximations . . . . . . . . . . . . . 286

27.1 The gsl cheb series struct . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 286 27.2 Creation and Calculation of Chebyshev Series . . . . . . . . . . 286 27.3 Chebyshev Series Evaluation . . . . . . . . . . . . . . . . . . . . . . . . . . 287 27.4 Derivatives and Integrals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 287 27.5 Examples. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 288 27.6 References and Further Reading . . . . . . . . . . . . . . . . . . . . . . . 289

28 Series Acceleration . . . . . . . . . . . . . . . . . . . . . 290

28.1 Acceleration functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 290 28.2 Acceleration functions without error estimation. . . . . . . . . 291 28.3 Example of accelerating a series . . . . . . . . . . . . . . . . . . . . . . . 291 28.4 References and Further Reading . . . . . . . . . . . . . . . . . . . . . . . 293

29 Discrete Hankel Transforms . . . . . . . . . . . . . 294

29.1 Definitions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 294 29.2 Functions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 295 29.3 References and Further Reading . . . . . . . . . . . . . . . . . . . . . . . 295

30 One dimensional Root-Finding . . . . . . . . . . 296

30.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 30.2 Caveats . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 296 30.3 Initializing the Solver. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 297 30.4 Providing the function to solve . . . . . . . . . . . . . . . . . . . . . . . . 298 30.5 Search Bounds and Guesses . . . . . . . . . . . . . . . . . . . . . . . . . . . 300 30.6 Iteration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 300 30.7 Search Stopping Parameters . . . . . . . . . . . . . . . . . . . . . . . . . . 301 30.8 Root Bracketing Algorithms . . . . . . . . . . . . . . . . . . . . . . . . . . 302 30.9 Root Finding Algorithms using Derivatives. . . . . . . . . . . . . 303 30.10 Examples. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 304 30.11 References and Further Reading . . . . . . . . . . . . . . . . . . . . . . 308

31 One dimensional Minimization . . . . . . . . . . 309

31.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 309 31.2 Caveats . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 310 31.3 Initializing the Minimizer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 310 31.4 Providing the function to minimize . . . . . . . . . . . . . . . . . . . . 311 31.5 Iteration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 311 31.6 Stopping Parameters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 312 31.7 Minimization Algorithms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 313 31.8 Examples. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 313 31.9 References and Further Reading . . . . . . . . . . . . . . . . . . . . . . . 315

32 Multidimensional Root-Finding. . . . . . . . . . 316

32.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 316 32.2 Initializing the Solver. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 317 32.3 Providing the function to solve . . . . . . . . . . . . . . . . . . . . . . . . 318 32.4 Iteration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 320 32.5 Search Stopping Parameters . . . . . . . . . . . . . . . . . . . . . . . . . . 321 32.6 Algorithms using Derivatives . . . . . . . . . . . . . . . . . . . . . . . . . . 322 32.7 Algorithms without Derivatives . . . . . . . . . . . . . . . . . . . . . . . 323 32.8 Examples. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 324 32.9 References and Further Reading . . . . . . . . . . . . . . . . . . . . . . . 329

33 Multidimensional Minimization. . . . . . . . . . 330

33.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 330 33.2 Caveats . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 330 33.3 Initializing the Multidimensional Minimizer . . . . . . . . . . . . 331 33.4 Providing a function to minimize . . . . . . . . . . . . . . . . . . . . . . 332 33.5 Iteration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 333 33.6 Stopping Criteria . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 334 33.7 Algorithms . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 335 33.8 Examples. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 336 33.9 References and Further Reading . . . . . . . . . . . . . . . . . . . . . . . 338

34 Least-Squares Fitting . . . . . . . . . . . . . . . . . . . 339

34.1 Linear regression . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 339 34.2 Linear fitting without a constant term . . . . . . . . . . . . . . . . . 340 34.3 Multi-parameter fitting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 341 34.4 Examples. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 342 34.5 References and Further Reading . . . . . . . . . . . . . . . . . . . . . . . 347

35 Nonlinear Least-Squares Fitting . . . . . . . . . 348

35.1 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 348 35.2 Initializing the Solver. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 348 35.3 Providing the Function to be Minimized . . . . . . . . . . . . . . . 349 35.4 Iteration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 350 35.5 Search Stopping Parameters . . . . . . . . . . . . . . . . . . . . . . . . . . 351 35.6 Minimization Algorithms using Derivatives . . . . . . . . . . . . . 352 35.7 Minimization Algorithms without Derivatives . . . . . . . . . . 353 35.8 Computing the covariance matrix of best fit parameters

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 353 35.9 Examples. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 353 35.10 References and Further Reading . . . . . . . . . . . . . . . . . . . . . . 358

36 Physical Constants . . . . . . . . . . . . . . . . . . . . . 359

36.1 Fundamental Constants. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 359 36.2 Astronomy and Astrophysics . . . . . . . . . . . . . . . . . . . . . . . . . . 360 36.3 Atomic and Nuclear Physics . . . . . . . . . . . . . . . . . . . . . . . . . . 360 36.4 Measurement of Time . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 361 36.5 Imperial Units . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 361 36.6 Nautical Units . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 361 36.7 Printers Units . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 362 36.8 Volume. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 362 36.9 Mass and Weight. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 362 36.10 Thermal Energy and Power . . . . . . . . . . . . . . . . . . . . . . . . . . 363 36.11 Pressure. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 363 36.12 Viscosity . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 363 36.13 Light and Illumination . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 364 36.14 Radioactivity . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 364 36.15 Examples. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 364 36.16 References and Further Reading . . . . . . . . . . . . . . . . . . . . . . 365

37 IEEE floating-point arithmetic . . . . . . . . . . 366

37.1 Representation of floating point numbers. . . . . . . . . . . . . . . 366 37.2 Setting up your IEEE environment . . . . . . . . . . . . . . . . . . . . 368 37.3 References and Further Reading . . . . . . . . . . . . . . . . . . . . . . . 370

Appendix A Debugging Numerical Programs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 371

A.1 Using gdb . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 371 A.2 Examining floating point registers. . . . . . . . . . . . . . . . . . . . . . 372 A.3 Handling floating point exceptions . . . . . . . . . . . . . . . . . . . . . 373 A.4 GCC warning options for numerical programs. . . . . . . . . . . 373 A.5 References and Further Reading . . . . . . . . . . . . . . . . . . . . . . . 375

Appendix B Contributors to GSL . . . . . . . . . . 376 Appendix C Autoconf Macros . . . . . . . . . . . . . . 377

Appendix D GSL CBLAS Library . . . . . . . . . . 378

D.1 Level 1 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 378 D.2 Level 2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 381 D.3 Level 3 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 387 D.4 Examples . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 390

GNU General Public License . . . . . . . . . . . . . . . . 392

Preamble . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 392 TERMS AND CONDITIONS FOR COPYING, DISTRIBUTION

AND MODIFICATION . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 393 Appendix: How to Apply These Terms to Your New Programs

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 397

GNU Free Documentation License . . . . . . . . . . . 399

ADDENDUM: How to use this License for your documents. … 405

Function Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 406

Variable Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 421

Type Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 422

Concept Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 423

xiv